A Bibliography of Publications in Byte Magazine: 1990–1994

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
WWW URL: https://www.math.utah.edu/~beebe/

12 January 2024 Version 2.14

Title word cross-reference

/Views [Api94b].

040/120 [Ano93-60]. 040/200 [Ano93-60].

1 [Ano90w, Ano90-51, Ano90a, Ano90-110, Ano90-113, Ano91-29, Ano93-61, Ano93-130, Ano94a, Del93a, Del93b, Die90b, Far90, Nan90b, Pep91, Ref94]. 1-2-3 [AR91, Gas93, Ano90-51, Ano90-110, Ano90-113, Ano91-29, Ano93-61, Ano94a, Del93a, Del93b, Die90b, Far90, Pep91]. 1-2-3/G [Ano90-51, Die90b]. 1-megabyte [Tho90d, Ano90a]. 1/2 [Ano92-45]. 1/2-inch [Ano92-45]. 10 [Ano94-165, Nan94b]. 10-Mbps [Ano94-165, Nan94b]. 10-point [Ano94-168].

100 [Ano94-129, Bry93b, HK94]. 100-Mbps [Bry93b]. 100-MHz [Ano94-129, HK94]. 1000 [Ano91u]. 10A [Ano92-73]. 110 [Nad90a]. 1120NX [Ano91e]. 1200C [Rei93b]. 13-pound [Nad90d]. 144 [Tho93b]. 14th [Ano92-185]. 15-to [Ano94-89]. 15th [Ano90m]. 16 [Tho90e]. 16-and [Gre94e]. 16-bit [Ano93-48, Ano94-88, Ano94-119, Gla90b, Sha94b, Tho94b, Wsz90]. 16-Million-Color [Tho90e]. 16.7 [Ano91-131]. 16.7-Million-Color [Ano91-131]. 165c [Tho93a]. 17-ppm [Ano93-92]. 180 [Ano93-93]. 180-MBps [Ano93-93]. 1954 [Hal94d]. 1954-1994 [Hal94d]. 1990s [Cra91k, Lip90b, Osm90, Ras91b, Ras91o, Tho90f, VC90s, Woo91b]. 1994 [Hal94d].

2 [Ano91-117, Ano91-76, Ano91-131, Ano91-116, Ano91-132, Ano91-192, Ano92-83, Ano92-91, Ano92-132, Ano92-112, Ano92-125, Ano93-80, Ano93-39, Ano93-40, Ano93-41, Ano93-60, Ano93-51, Ano93-61, Ano93-55, Ano93-86, Ano93-110, Ano93-117, Ano93-78, Ano93-79, Ano93-132, Ano94r, Ano94-33, Ano94-126, Ano94-75, Ano94-97, Ano94-127, Api94b, Cal91a, CV90c, Dan90, Egl91c, Gre90g, Gre94e, Ham90, Ham91, Ham92, Har90a, Hel90c, Hel90d, Hel90e, Hel90b, Hel91c, Hol90, Jef93b, Ken93b, Kog94, Kro90, Mas90b, Mas91c, Mas92, Mas93, Min90c, Min90i, Min90d, Min90e, Min90h, Min90f, Min91b, Min92b, Min92a, Min93a, Min93b, Mit91b, Nan90b, Nan92d, Nan92e, Nan93a, Nan93c, Nan93f, Nan94a, Orr94, Pou90s, Pou92n, Pou93n, Rya90d, Smi93a, Smi94c, Ude90g, Ude91b, Ude91f, Vos90b, Yag90k, Yag92g, Yag93f]. 2-based [Ude91b]. 2-D [Cal91a, Orr94, Smi94c]. 2-inch [Ano92-45]. 2-pound [Ano92-128]. 20 [Ano90d, Ano91-94, Bon90c, Mia90d, Mit91b, Nad90d]. 20-MB [Mia90d]. 20-megabyte [Mia90d]. 20-MHz [Ano90d, Bon90c]. 20-volume [Ano92-136]. 200 [Ano91-193, Ano93-60]. 20CSX [Nad90a]. 21-inch [Ano94-89, Api94a]. 212A [Tur90]. 24 [Tho90e]. 24-bit [Ano90-95, Ano91-131, BF90, DE92, Lov90a, Lov92a, Lov94]. 25 [Alf92d, Ano92-78, Ano92-222, AMW90, Api93b, Egl90b, San90a, Smi90f]. 25-MHz [Ano92-78, AMW90, Egl90b, San90a]. 25c [Egl92c, Ano92u]. 25E [Lin91c]. 28.8 [Egl94b]. 28.8-Kbps [Egl94b]. 286 [Nad90e]. 2s [Ano91-131].

3 [Ano90-110, Ano90-113, Ano91-29, Ano91-103, Ano92c, Ano92d, Ano92-142, Ano92-149, Ano92-156, Ano92-167, Ano92-212, Ano93-61, Ano93-116, Ano93-125, Ano93-135, Ano94-71, Ano94a, Bri90, Cal91a, Cla91a, CWVS94, Cra91i, Del93a, Del93b, Far90, Gla90h, Har90a, Lov91b, McA90, Orr94, Pep91, Ran90, Rei90a, Rei94a, Sac90, Ups90, Way94h, Yag93i]. 3-D [Ano91-103, Ano92c, Ano92d, Ano92-142, Ano92-149, Ano92-156, Ano92-167, Ano92-212, Ano93-116, Ano93-125, Ano93-135, Ano94-71, Bri90, Cal91a, Cla91a, CWVS94, Cra91i, Gla90h, Lov91b, McA90, Orr94, Ran90, Rei90a, Rei94a, Sac90, Ups90, Way94h, Yag93i]. 3-Draw [Sac90]. 3-pound [Wau93b]. 3.0 [Gla91d]. 3.3 [Ano94-111]. 3.3-V [Ano94-111]. 3/G [Ano90-51, Die90b]. 300-megabyte [Gre90i]. 3000T [Ano93-60, Pou92o]. 3000T-040 [Ano93-60]. 3000T-040/200 [Ano93-60]. 3000UX [Ano90f]. 3125A [Rei91b]. 32-bit [Ano92-198, Ano94-126, Ano94-111, Gla90b, Gre91b, Gre94e, Hel90d, Hel92a, Lov91d, Min92a, Nan93c, One91, Pou92d, Sha94b, Ude90g]. 320L [Api93b]. 325NC [Ano92-39]. 32CSX [Ano90k]. 33 [Alf90c, Ano91b, AD91, EW92, MG90, Nad90c, VC90d]. 33-MHz [Alf90c, Ano91b, AD91, EW92, MG90, Nad90c, VC90d]. 33N [Ano92-72]. 35mm [Ano94c, Ano94c]. 35mm-Size [Ano94c]. 368SX [Rei90c]. 386 [Gla91a, Ude90f]. 386DX [Bon90c]. 386Max [Ano91-162, Gla91a, Lan90c]. 386SL [Ano91-108, Ano91-197, Car92]. 386SL-based [Ano91-197]. 386SLC [Alf92d]. 386SLC/50 [Alf92d]. 386SX [Ano90d, Ano91-161, Ano91-166, Ano92u, Cla90b, GAM90, Mit91a, Mit91b, VN91c]. 386SX/25C [Ano92u]. 386SXes [Cla90b, GW92]. 386SXL [Car92]. 3D [Lov91b].

4 [Ano94-126, Wau92]. 4-MB [Ano94-126]. 4-pound [Nad90f]. 4.0 [Ano92-123]. 4/50d [Wau92]. 4/AEB [Ano92-64]. 40 [Ano91-170]. 40-MHz [Ano91-170]. 4000-040 [Ano93-60]. 4000-040/120 [Ano93-60]. 4000D [Ano92-131]. 425e [Ano91-58]. 466D2 [Ano92-65]. 480-pixel [Ano94c]. 486-class [Row93]. 486/25E [Lin91c]. 486/33 [Ano93y, Pou91l]. 486/33N [Ano92-72]. 486/50 [Ano90-99, Ano92-53]. 486DX2 [Alf92a, Ano92-78, HP94]. 486DX4 [Ano94-83, HK94]. 486s [FJPR93, Gre93d, HP94, San90a, VC90d]. 486SLC [Alf92d]. 486SLC-25 [Alf92d]. 486SX [Alf91d, Ano91c, Ano92-222]. 486SX/25 [Ano92-222]. 48SX [Ano90-36]. 4D [Smi90f]. 4D/25 [Smi90f]. 4DC2 [Ano92-65]. 4DC2-66V [Ano92-65]. 4L [Egl93c]. 4M [Cry92a]. 4Si [Ano93-36]. 4SL [Api93b]. 4SL/25 [Api93b].

5 [Ano93-126]. 5.7-pound [Rei90c]. 50 [Alf92d, Ano92-78, Api94d]. 50-and [FJPR93]. 50-MHz [Ano92-78, Api94d]. 50d [Wau92].

600-dpi [Van93b]. 6000 [Daw94c, RS90]. 64-bit [Ano93-74, Ano94-129, Lin90b, Lov94, Mas91b, Mas91a]. 640 [Ras91i]. 640-by [Ano94c]. 640-KB [Ras91i]. 640K [Gre90h, Lan90c]. 640K-byte [Gre90h, Lan90c]. 64K [Min90c]. 64K-byte [Min90c]. 66 [Ano92e, Ano92-111, Api92b, FJPR93, HP94]. 66-MHz [Ano92e, Ano92-111, Api92b, FJPR93, HP94]. 66V [Ano92-65]. 68040-based [Ano92m]. 680x0 [Ga94, Tho94h]. 69040-based [Yag93c].

700 [Ano91-134]. 7100LC [Smi94g]. 712/60 [Smi94g].

8-and [Ano93-48]. 80 [Ano90-95, Api94a]. 80-HZ [Api94a]. 80286s [Hay90b]. 80386SX [VC90e]. 80x86 [Ano94d, Ano94-71, Ano94-82, Hal94b, Pou94f, Rya94b, Sta94b]. 80x86-based [Ano94-71]. 80x86-Compatible [Ano94d]. 840AV [Tho94f]. 8514 [Cah90]. 8514/A [Cah90]. 8514/Ultra [Ano91d]. 86C911 [Ano92-171]. ’86s [Hay90b].

90 [Ano90-68]. 9080i [Ano91d]. ’91 [Ano91v]. 95LX [AR91]. 9600-bps [Gib90a, Hur93a]. 9624e [Ano90-92].

A/UX [Ano90h, Tho92c, Yag91g]. A20 [Ano92-91, Min92d]. A4 [Ano92g]. abilities [Mia93b]. ability [Ano94-117, Bri90, MW93b, Rog91, Smi94c]. able [Fow91, Gut93]. Abound [Ran93]. abounds [Fri92]. above [Rya94a]. abroad [Red93]. abundance [Ano92-142]. AC [Cla90b]. accelerated [Ano93-74]. Accelerating [Wal94a]. Acceleration [Ano94-106, Ano94-171, Ano93y, Ano94-129]. Accelerator [Ano90-93, Ano92-162, Ano92-169, Ano92-171, Ano93-58, Ano94-106, Ano94-90, Gre93e, Ric92, Tho92h, Ano93-30]. accelerators [Ano93-137, Gre93c, Way93e]. Acceptance [Ano93-121, Ano94-79, Ano93s, Cla94b, Sta94a]. Access [Ano92-182, Ano94-63, Ano94-87, AG91, Cas92, Cla94b, Fri94b, Gre92, Nad94b, Ude92a, Vos92, Ano91u, Ano92r, Ano92-85, Ano92-175, Ano93e, Ano94h, Ano94-76, Ano94-77, Ano94-143, Ano94-148, Ano94-166, Ano94-181, Buc92, Cla93b, Cro93a, Der92, Egl93e, Fie91a, Fri94c, Gla90d, HS94, Kay94b, LDH94, Mul94, Obe90, Ras90f, Sch91b, She90b, Smi94e, Way94a, Ano94-114, Car94, Hal93f, Pou93s, Ude91a]. accessible [Bry93a, Ste94b]. accessing [Ano94-63, SE94]. accessory [Chi92, Woo90b, Wsz93c]. accommodate [Ano94m, Kor93]. accomplished [Cra90j]. According [Man90]. account [Hed94]. accuracy [Ano92-49, Die94, Egl94c]. ACE [Ano91-87]. AcerAnyWare [Ano91e]. AcerPac [Ano92f]. achieve [Smi94g, Sta94a]. Achievements [Ano93-48, Ano94-90, Ano94-87, Ano94-88, Fox93, HM94, HHP93, HP94, Len94b, Len94a].

Acorn [Ano92g, Pou90e]. acquire [Tho91a]. Acquiring [Rei92c]. Acquisition [Bar91a, ESN94, Egl94g, Ano90-103, Rei92c, Smi93b]. Acrobat [Ano93a, Die93a]. Across [SE94, Ano92-177, Ano94-171, AWGY90, Api94b, Dej94, Nag93, Sto93, Van93a]. act [Ano92-50, Ano93a, GP90]. action [HC93]. ActionNote [Row93]. Active [Ano93-83, Ano94-35, Pou93e, Ano92-199, Ano93-126, Ano94-112, Pou94q, Ras92c, VC90c, Yag93b, Pou93e]. active-matrix [Ano93-126, Ano94-35, Ras92c, Yag93b]. activity [Pou94l]. Actor [Url90]. acts [Nan91l]. actually [Egl90b, Tho93c]. Ad [Ano92-199]. Adaptec [Lab92b]. adapted [Pou94c]. Adapter [Ano91-131, Lov92a, Ano92-139, Ano93-119, Ano93-131, Fis90b, Nan94b]. Adapters [Ano92-204, Ano93-48, Ano94-88, Die90e, Fis90b, Ano94-177, Gre91b, Lov94, Ras92a]. Adapting [Ano94e]. Add [Ano91-163, Ano94f, Ano94-166, Hel90a, Alf91d, Ano90i, Ano94-111, Ano94-162, Bar91e, FT90, Kay94a, Ken94, Lie93, Pic93, Ras91a, Smi94c, Tho93c, VC90a, YG92]. add-in [Bar91e, Ken94]. add-ins [VC90a]. add-on [Kay94a, Tho93c]. add-ons [Ras91a]. added [Ano92o, Gre94e, SS94]. Adding [Bon90d, Ano92-49, Ano94-61, Ano94-132, Hyo94, You93]. addition [Ano91-84]. additions [Egl92c]. address [Ano91-149]. addressable [Way91c]. addresses [Gre93b, Ras90i, Sch91b]. Addressing [Ano93-115, Ano94-112]. Adds [Ano94-40, Cau93, Ano90-30, Ano91-135, Ano91-174, Ano94e, Hal93b, Rei90a, Ude93f, Ude93h]. adherence [Bry94e]. administration [Ano94-147, Min90f, Smi94e]. administrative [Cot93a]. administrator [Ano92-195, ES94]. Adobe [Ano90-76, Ano91f, Ano91n, Ano92-82, Ano92-206, Ano93d, Ano93a, Ano93-43, Ano94-136, Ano94-189, Lin94, Tho93d, Vor92]. adopt [Ano94-57, Way94g]. adoption [Ano94e]. advance [Ano94q]. Advanced [Ano90-98, Ano91g, Ano91h, BR93, Hal90, Ude94a, Ude94k, Ano91-110, Ano91-109, TS90, Ude94e]. Advancement [Pou90v]. Advances [Ano92-102, Ano92-202, CP94, Wal94c]. Advancing [Bur92, Egl94b, Ano92d]. Advantage [Ano94-142, Bac94, Cru90, Ude93k, Ano91-65, Hol93a, Tho93d, VC90k]. advantages [DM90, Dys92]. Adventures [Ros90b]. Advice [Ano91-55, Nan91h]. Advise [Ano90e]. AEB [Ano92-64]. Afar [ES94]. Affair [Ude90g, Hal94i]. affect [Ano94-81, Fie91d, Fis90a, Ude92b, VC90j]. Affects [Ano94-90]. afford [Tho90h]. Affordable [Tho94h, UE93, Yag91f, Zei93, Ano91u, Ano91-125, Ano92u, Ano92-71, Ano94-129, Daw94b, Joc90c, Lov91b, Lov91d, Lov94, MW94, Smi90g, Tho93g, Tho94c, Yag92c]. aficionados [Ano91-115]. Afloat [Ano92j]. After [Ano90-113, Ano94k, Ano94-85, Sei91, Ano92-205, Sto93]. Again [Ano91-57, KPUG93, Pou91q, Pou92j, STA92, Ano92-122, Pou94s, Ras91h, VC90t, Ano92z]. against [Ano90-50, Egl94c, Gre93d, Hog90, Min90d, Smi90d]. Age [Ano91-99, Ano92-77, Nad91, Pou94d, Ude90e, Vau90, War92, CV90d, Cox90, Lio90a, Rya90c]. agencies [Ano94z]. Agenda [Ano91-149]. Agents [Ano94g, Way94a, Rei94f]. Agfa [Van94a]. aggregate [Fri94c]. Aggressive [Sma93b, Ano91-187, Hal94h, Van93c]. Ago [Gre93g, DG91]. agreeing [LDH94]. agreement [Ano94h, Ano94-81]. Agrep [MW92b]. Ahead [Fie90g, Var94b, Ano91-54, Ude90g]. AI [Ano91k, Ano91s, Ano92-60, Hel91a, Hel91f, KPC93, KDG^{+}93, Nan92c, Ras90p, Rya91a, Tho91a]. aid [ADE90, Nag92]. Aids [SU93, Kor93]. Aim [Ano94w, Ano94-154]. aims [Ano92-183, Ano92-207, Car94, Cau93]. ain’t [Ano92-148]. air [Ano94-81, Ano92-60]. air-traffic [Ano94-81]. Airis [Nad90e]. AIX [Ano94-159, Nan91d, Nan91e]. Al [Ano91-98]. Aladdin [Tho94i]. Alan [Ano91-42]. Aldus [Ano90-70, Ano91i, Ano91-123, Die93c, Egl94f]. Alert [Gre91c]. AlertView [HC93]. algebra [Way90c]. algorithmic [Sin94]. Algorithms [Way91b, Ano91-155, Api91d, Lie93, MW92b]. alienated [Mos93]. alike [Fin90a]. alive [Yag91d]. all-around [Hur93a]. All-in-One [Ano91j]. All-Terrain [Cla93a]. allegedly [Ano92-108]. Alliance [Hay93, Way94k]. Alliances [Ano94h]. allocate [Way90a]. allocation [Min90i]. allow [Ano92d, Ano92-207, Ano94h, Bry94a]. allows [Egl94g, Min90b, Nan94b]. Almost [Ano90-46, Egl94a, Ano91-174, Ano91-176, Ano91-193, Gla91d, MW93b, Nan94d, Ras90h]. Alone [MYT91, Lie93, Rya90b, Smi94b]. along [Egl93c, Mia92b]. Alpha [Ano92-106, Ano92-176, Bri90, Gre93h, Hal94h, Rya94a]. Alpha-based [Gre93h]. Alphas [Gre93h]. Alps [Ano93-131]. ALR [Ano90k, Api92b, Ga 94d, Mia90a, Nad90a, Nad90c, Per94a, Ung91, VC90h]. already [Ano93-68, Ano94a, Hal93b, Tho93d]. Also [Ano94w, Ano94-72, Egl94c, Gre94b, Len94b, Pou94q, You94]. Altair [Egl91b]. alter [Ano92-144, Ano94-85]. Alternative [Ano91-122, Ano94j, HK94, Min92a, Nan91d, Nan91e, Smi90a, Smi90b, Smi91a, Yag90a, Yag90b, Yag90c, Alf90b, Ano91-177, Ano94-143, Dav90, Gla91c, Ude90d]. alternatives [Bar91b, Cal93, DE90, UM90]. Although [Ref94, Rei94a, Ric94, Rya90c, Ano92c]. Alto [Cla91a]. Altos [UY91]. always [Cro93b, Gru90, VC90l, VC90q]. Am386SC [Ano94i]. Amateur [Ano91k]. amazing [Ano91-173]. Ambiguity [Sco93]. Ambitious [Ano94-32, Ano94-77, Hal94f, Pou93a, Van93a]. Ambra [Per94a]. AMD [Ano93b, Ano93-45, Gre90b, Hal93e, Hal94c, MG91]. America [Ano91-70, Hal94e]. American [Ano92y, Pou90v]. Ami [Ano92i, Woo90d]. Amid [Ano94z]. Amiga [Hol91, Ano90f, Ano93-60, Ano93-53, Hol91, Pou92o]. Amigas [Yag93c]. Amok [KGTD92]. Among [Way94g, Alf90c, Ano93-75, Ano94l, Bry94d, Gre90b, Nan93c, Pet91, Rya91e, Tho94i]. amounts [Rei94b, Rya91b]. Ample [Ano92j, Tho92h]. Amstrad [Bar90e, Pou93a, Ano92k, Pou93a]. Analog [Ano93m, Ano94j, ASS92]. Analysis [Ano91-157, Gib90b, Sta91a, Ano92j, Ano92-190, ADE90, Bis91, Gre90j, Ung90b, Wal94e]. analysts [Ano94-79]. analytical [Ano91-157, Ref94]. analyze [Cro93b]. Analyzers [Nan92c]. analyzes [Smi94f]. analyzing [Sta91a]. anatomy [Alf90a]. Andersen [Hed93b]. Anderson [Wil94]. Angle [Egl91c, Egl91c]. angles [UY91]. anguish [Ano91-106]. animated [Ros90a]. Animation [Lov91b, Yag92d, You93, Ano94-71, Cry94, Van93a]. Animator [Ros90a, Yag92d]. Anniversary [Ano90m, Ano90-33, Ano92x]. Announce [Ano94-45]. announcements [Ano94-84]. Announces [Ano93-38]. announcing [Ano94-45]. Annual [Ano91-93, Pou92l, Pou92k, Pou93t, Pou94s]. Anonymous [DeH93, Fie91a]. answer [Cla93a, Hed93b, Jof93, Ude91d, VC90j, Way90b]. answered [Ano92-49]. Answers [Ano92-95, Fie90a, Led90, Min91d]. Ante [Ano91-159, Nad90b, Sch92c, YS90]. Antialiasing [Cov90]. anticipated [Hal94b]. Anticipating [Ano91-95]. antivirus [Ano94-149]. Any [Gre90c, Egl93b, Smi94b, Wsz90]. Anymore [Ano94-132, ADE90, Ano92-208, Mia93a]. anyone [Ras90h, Ros90a]. anything [Saf93]. anytime [MW93b]. anywhere [Ano94-148, MW93b, Ras90h, Yag93b]. apart [Nan93d, Van94b]. API [IC94, Orr94, Ude93a, Ude93c]. Apiki [Api94b, Api94d]. APIs [Ano94-126]. APL [Ano91z]. Apollo [Ano91-60, Ano91-58, Smi90g]. Appeal [Nan91c, BEW92, Egl93b, Van94a]. appealing [Mit90b]. appear [Ano92-196, Ano93-68, Hal94f, Min92d]. appearing [Ano94-34, Ano94-71]. appears [Ano92-108]. appetites [WB90]. Apple [Ano90-37, Ano90h, Ano91l, Ano91m, Ano91-42, Ano91-66, Ano91-121, Ano92m, Ano92l, Ano92-129, Ano92-63, Ano92-104, Ano92-107, Ano93c, Ano93i, Ano93-37, Ano93-67, Ano93-85, Ano93-90, Ano93-93, Ano94n, Ano94k, Ano94l, Ano94m, Ano94-45, Ano94-162, Bar90a, Bet94, Car91a, Cra90j, Cra90l, Cra90o, Cra91h, Cra91k, DB93, Daw94a, Die90e, Gib90a, Gre94c, Joc90b, Lin91a, LAN92, Mas90a, NTS93d, Oth92, TH92, Tho92d, Tho92e, Tho92h, Tho92c, Tho93a, TS93a, TH93, Tho93e, TR94a, Tho94b, Tho94f, Tho94i, VC90j]. Apple/IBM [Ano92-107]. AppleShare [Cra90a]. AppleTalk [Ano91m, Ano92-86, Gib90a, SE91, VC90j]. appliance [Daw94b]. Application [Ano92-163, Ano94-39, Ano90-88, Ano91-176, Ano92-66, Ano93-72, Ano93v, Ano94-40, Ano94-69, Api94b, Die94, Gea90, Gib90b, Gil94, HK94, Law94, Len94a, Mia90i, Pou93k, Ras92b, She93a, Ste94a, The94, VC90s, Vos90a, Api94b]. Applications [Ano94f, Ano94x, Ano94-45, Ano94-71, Bro93, Chr94, DW91, GFK94, Gre94c, HK94, Hub92, Osh90, You93, Ano90j, Ano90-36, Ano91z, Ano91-56, Ano91-86, Ano91-127, Ano91-175, Ano91-187, Ano92d, Ano92-89, Ano92-99, Ano92-109, Ano92-80, Ano92-81, Ano92-135, Ano92-127, Ano92-178, Ano92-159, Ano92-192, Ano92-187, Ano92-186, Ano92-200, Ano93-61, Ano93-64, Ano93-110, Ano94h, Ano94e, Ano94g, Ano94l, Ano94-60, Ano94-81, Ano94-89, Ano94-90, Ano94-88, Ano94-148, Ano94-145, Ano94-174, Ano94-184, Api94b, Api94c, Bar93d, BR93, Bus92, Coo90a, Cot93b, DM90, EW92, Egl93a, Egl94g, FJPR93, Fox93, Ga 94a, Ga 94b, Gre92, Gre93a, Gut93, Hal93b, Har93, HP93, Hel90e, Hel91f, HH94a, HM94, HHP93, Hol94, HP94, Hur93a, KM94, Kay94c, Kea91a, Kea91c, Kor93, Lab92a, Law94, McA90, Min92d, Nan93c, NTS93c, Nic93, Nie94]. applications [Ols94, One91, Ras91o, Rei94a, Rya90b, RT94b, Sha94c, Sha94b, Smi94b, The94, Top91, Ude90b, UMR^{+}90, Ude93a, Ude93m, Url90, VC90c, VC90g, Vel90, Wal94e, Wau93b, Woo90a, Woo90c, Yag90b]. applications-based [Ano94-90]. Applied [Ano94-86]. applies [Sch90c]. apply [Ano92-38]. Applying [Ano92n, Egl93a]. Approach [Alf90c, Ano92-59, ET93, Osh90, Ric92, Ano91-111, Ano92-130, Ano93h, Ano93-98, Ano93-123, Bar92g, Kay94c, Nan93f, Way94f]. Approaches [Loe90, Ano92-216, Day90, Har93, Rob91a, Vog90]. Approaching [UY91, Rya90f]. appropriate [Sin94]. Approval [Pou91i]. Approximate [MW92a, MW92b]. Apricot [Ano92o]. April [Ano90b]. Aptly [Tho94k]. Arabia [Tab93]. Arago [Sch92c]. arcana [Hal94i]. Arche [Ano91-170, Pou91l]. Architects [Ano93-77]. Architecture [Ano94-90, The94, Ano92-176, Ano93-85, Ano94m, Ano94-180, DB93, Dej94, Hal94h, Jef93b, Min90c, Nan91d, Nan91i, Pou93g, Pou94f, RH93, Sch93b, Sta94b, Ude93m, Var94a, Way92b, Osh90, Way94d]. architectures [Ano90j, Ano92-202, Cro93a, Pou93h, Rei94e, Rob91a, Rya90a, VC90g]. archives [Ste91a]. archiving [Pou90k]. ARCnet [Gla91c]. Area [Ano91-190, Ano93v, Ano94-113, Ano94-149, Cla93a, Gre91a, HS94, Ste91a, Ste91d, Ude94d]. areas [MH91, VC90s]. Arena [Ano93a, Ano94-189, Kro90, Van94b]. aren’t [Ano91-169, Ano92-208, Fin90b, HPPR93, HP93, Mia93a]. Ares [Ano91-170]. argues [Ano90-50, Ano90-116, Ano91-194, Gru90]. ARM [Ano92g, Pou92d]. ARM610 [Pou92d]. Armed [Ken90]. Army [Van92a]. Array [Ano92-126, Alf92e, Ano93-132, Ano93-126, Mur93, Ric91]. Arrays [Alf92e, Ano92-126, Ano93f]. arrival [Ano90-41]. Arrive [Ano92-45, Mas92, Ano94-106, Bar93d, HPPR93]. arrived [Ano92-199, BL90, Rei93e, Ude93h]. Arrives [Ano94-106, LK90, Ano92-36, Pou93t]. arsenide [Rob90b]. Art [Ano90-98, Ano90-116, BGE^{+}92, Lov93a, Nan91f, Pou94c, Ran90, Ros90a, Bar91b, DWN91, KDG^{+}93, Ols94, Van93a]. Artichoke [Bus92]. Artificial [Ano91-41, Dys92, KYCD92, Ste91b, Ken93b]. Artisoft [Ano90-44, Ano92-87, Nan93d]. Artist [Ano91n]. Asante [Die90e]. ASCII [She91]. Ashes [Cra90o]. Ashton [Ano90r, Ano90-30]. Ashton-Tate [Ano90r, Ano90-30]. Asia [Ano94-119]. Ask [Ano92s, Ano92p, Ano92q, Ano92r, Ano92t, Ano92-95, Ano93d, Ano93e, Ano93f, Ano93-64]. asked [Fie90a, Min91d]. asking [Way93f]. asks [Buc92]. aspects [Dan90]. Assault [Ano93-42]. Assembly [Ano92-191, Ano92-192, Bet94, Chr94, Cot93b, Gre92, Gri94, Hol93a, Hol93b, Jef93a, Ken90, Kiy93b, Kiy94, MW92b, Nag92, Nan93f, NHW93, Pie93, ST93, Sin94, You93, Ano90-116, Gre90f, Gre90g, Ano90-116]. assistant [O’M94]. associated [Alf93]. Association [Pou90v]. assorted [Pou90s]. assortment [Ano92-134]. assumes [Pou94r]. AST [Ano92u, Mit93b, Nad90c, Sma92, VC90h, Wau92]. astronomy [TGN94]. Asymetrix [Ano93p, Woo90c]. Asynchronous [Pou93b]. AT&T [Ano92-183, Ano93-45, Ano94o, Bry94b, GR90, Mit91b, Rei94f, Rya93a]. AT-class [GK91, Nad90f]. at-work [Pou93e]. AT1000 [Ano91-181]. Atari [Ano93-27, Pou91j]. ATE [Api91c]. ATI [Ano91d, Ano94-171]. Atlantix [UY91]. atlas [Man90]. ATM [Ano94-158, Cla93a, Rei94b, Way94g]. ATS [Ano90g, VC90d]. attach [Tho94d]. Attached [GP90, Ras91j]. attack [DeH93]. attacking [Ano92-108]. Attacks [DeH93]. attempt [Cra91h]. attend [Pou93m]. attending [Pou94k]. attention [Ano92-147]. attention-grabbing [Ano92-147]. Attitude [Ano91-194]. attorney [McG90b]. Attractions [Pou91t]. attractive [Tra92]. Attributes [Min93a, Ano91-65]. auctions [Way90a]. Audience [Ano94-114, YG92, Die94]. Audio [YG92, Ano91-86, Ano92-199, Ano94-176, Lin94, Tho90b, Ano91-68]. Audio-Video [Ano91-68]. AudioMan [Ano93h]. audiotape [Lio90a]. Audit [NTS92a]. Auld [You90]. Authentication [Ano94-187]. authenticity [Way91d]. author [Cri93]. Authoring [Har93, Vel90]. authority [Sta91a]. Authorware [Mia90g, Har93]. AutoCAD [Cra90b, Gre94e, Lov91c, Wau92, Wau93a]. Autodesk [Ros90a, Wau93a]. automate [Egl91d, Egl94b, Gil94]. automated [Ano92-129]. automates [Ano94-179, Dlu91, HC93]. Automatic [Ano92-144, NTS93c, Joc90b, WS91, Wsz93a]. Automating [Way94b, Rei93d, Smi94a]. Automation [Ano91-145, Ano92-184, Ga 94b]. automotive [Pou93f]. automotive-control [Pou93f]. autonomous [Smi93b]. AutoShade [Lov91c]. Availability [Ham92, Ano92-112]. Available [Ude94k, Ano90-65, Ano92-157, Ano93-29, Ano94w, Ano94-86, Daw94d, Fie90g, Fri90, Fri94b, GN93, Lov93b, Mil93, Rob90a, Ude94e, Ude94h, Way90c, Woo90a, Wsz90]. Avoid [Tho94g]. avoiding [Ano91-61]. await [Mur93]. awaited [Ano91-137, Ano91-151, Ude90g, Ude93i, Ude94l]. Awards [Ano90l, Ano91t, Ano91a, Ano92a, Ano92b, Ano93-94, Ano94p, Ano94b, Nad93, Nad94a, Pou90m, Pou91u, Pou92l, Pou92k, Pou93r, Pou94m, Pou94s]. Aware [Dav94b]. awareness [Ano92-199, Yav94]. Away [Fri94a, Ras91e, Way94a, Ano91-56, Ano92-134, Egl94f, Ras91j]. Axis [Smi94e]. AXP [Gre93h].

Babbly [Ano91s]. Babelware [Mil93]. Back [Ano91-173, Ano92-205, Egl94f, Fie90b, Hay93, Kay94a, Way94j, Ano90y, Ano90-88, Ano93-113, Ano94-179, EB90a, Pou90o, Smi93e, Ano92u]. Back-It [Ano92u]. Backing [Ras90a]. Backpack [Ano90i]. backs [Ano92-221]. Backup [Ano90-57, Ano90-83, Ano92-189, Ano94-169, Egl93f, Ano91-134, Ano92u, Api90c, Fie91k, Gla90c, Mia90c, Pou90k, Ras90a]. Backups [Pou90k]. backward [Ano91-29]. Badge [Pou93e]. Badges [Pou93e]. balance [Van94a, Pou91k]. balanced [Ano93-116]. Balancing [Hog90, Ano93a]. ball [Cra90o]. balloon [DG91]. BallPoint [Ano91o]. Band [Ano94-28]. bandwagon [You90]. bandwidth [Ano94-37, Fri94c, SU93]. Bane [Rya91c]. Bang [Pou94m, VC90f, VN90b]. bank [Ano92-149]. Banyan [Ryd93]. Baran [Bar91b]. bare [Daw94d]. bare-bones [Daw94d]. barely [Ano94-66]. Bargain [Tho93b]. Barlow [Bar91f]. Barrier [Ras91i, Ano92-39, Gre90h, Lan90c, Sch91b, Tho93b]. Barry [Nan91h, Nan91j]. base [Mic91, Ano90-87]. Based [Ano94w, Len94b, Ano92m, Ano92-119, Ano92-149, Ano93-48, Ano94-72, Ano94-90, Ano94-135, Ano94-127, Ano94-177, Bar92b, BL94, Cry94, Die94, Eva90, Gre93d, Lie93, Mar90a, Nan91a, Pou94a, Ryd93, San90a, Sch92a, Sma93a, Ude91b, Van93a, Var94b, VN90b, Wsz91, Yag93c, Ano91-62, Ano91-127, Ano91-197, Ano92k, Ano92-53, Ano94f, Ano94-32, Ano94-71, Ano94-119, Ano94-129, BBWR90, CS92, Cra90c, Dav94a, Daw94c, Gre91c, Gre93h, Hal94b, Lin93, Nag92, Nan91d, NTS93b, Slo91, The94, Tho93f, Ude94e, Ung90a, VC90i, Yag91c]. bases [Lu93b, Sum90]. Bash [STA92]. Basic [Ano91-180, Pou93c, Yag93k, May94a, Ano91-177, Ano91-180, Ano92q, Ano92-30, Ano92-66, Ga 94c, Tho93c, Wig90, Wig92, Woo91b, Yag93a, Ano90-81, Ano92-145, Ano94f, Ano94-176, Hal93f, Hur93b, KYCD92, Pou93c, Pou93k, Ude92a, Ude94c, Ude94g]. BASIC-to-C [Ano92-66]. Basics [Ano94-68, Yag90f, Wig92]. basis [Ano92-42]. Bastion [Pou94e]. batch [Ano92q, Ano92-184]. batch-file [Ano92q, Ano92-184]. batteries [Ano91-128]. Battery [Egl92e, Len94a, Daw94c, Rei90b]. battery-life [Egl92e]. Battle [Ano91-124, Ano92v, Ano92-55, KM94, SY90, Ano92i, Ano94-102, Hal93e, Rei90b, SH90]. battlefield [Ano92-77]. battles [Fie91d, Hal94h, Pou93j]. bays [CW92]. BBS [Pou92m, WB90]. BBSes [Ano94z]. Be [Ano91-59, Ano93-135, Ano93-142, Ano94-30, Ano94-165, Cra90n, Ras91d, Ude90d, VC90p, Way93g, Alf91f, Ano91-42, Ano92-28, Ano92-127, Ano93-29, Ano93v, Ano93-93, Ano93-100, Ano94w, Ano94-34, Ano94-79, Ano94-82, Ano94-86, Ano94-114, Ano94c, Ano94-166, Chi92, Cla93a, Coo90c, Egl90b, Egl94c, Eva90, Fow91, Gre93a, Hal94i, Har91b, Hog90, Jof93, Kay94a, Lip90b, Mar92a, Mas91a, Mel93b, Mur93, PS94, Pou94c, Rei91a, Rei94b, Rei94c, Rie91, Rya90a, RP90, Sat90, SE93a, Taz91, Tho93f, Ude94j, VC90k, VC90r, VN91d, Wau93c, Way93f, Way94d, Way94g, Way94h, Yag90d, Yag90j, YG92, Yag93e]. Beam [Egl94d, Smi91b]. Beame [Nan93g]. Beating [HK94]. Beauty [Hel90b]. because [Mul94]. BeckerTools [Ano91p]. Become [All92b, Ano92-30, Ano94-180, Eva94, Tho94c, VC90s]. Becomes [Dav94b, Nan94c, Zei93, Ano92-77, Ano92-138, Ano94-149, Ano94-174, Sch92c]. Becoming [Yag90i, Ano92-80, Mar90b]. Bee [Ano90-27]. bee-hunting [Ano90-27]. beefed [Ano90-109]. beefed-up [Ano90-109]. Beefier [Smi93a]. beefs [Alf91b]. been [Ano91-42, Ano92-226, Egl93a, Han92a, Pou90m, Ref94, Yag90j]. beer [LS91a]. Before [Bar92b, Smi90d, Ano93-68, Ano94c, Dio94, Pou93c, VC90s]. Beget [Ano90-38]. begin [Ano94-72]. beginners [KYCD92]. beginning [Ano92-196, Ano94-29, Hal94f, Ols94, Way94k]. begins [Ano93-31, Yag90f, Yag92b]. Behavior [Gib90b]. behaviors [Ano92-178]. Behind [Ano93i, Bar90g, Egl90a, Joc92b, Ano91-67, Ano92-102, Ano94-82, Egl91b, Way93f]. Being [Ano92-74, Mit91c, Mit93a, YM93, Ano92-49, Ano94-172, Egl94a, Joc92a, Kay94c, McG90a]. Believe [Yag90d, Rei94g]. believer [Spi90a]. Believing [Web93]. Bell [Ano91-164]. bells [Ras90e]. Ben [Smi94c]. Bench [Bar91a]. Benchmark [Ano90-106, Tho93a, Ano90o, Ano92-188]. Benchmarking [Pou94l]. Benchmarks [Ano90n, Ano90o, Ano94-27, Ano94-36, Ano94-190, Die90d, Smi90d, Ano90-106, Gre93f, SB94, Tho93h, Gre93d]. benefits [Ano93-72, Ano93-110, Ano94-57, Kay94c, Mur93, Rei92b]. Berkeley [Sto93, Ano90-69]. berserk [Ano92-67]. Besides [She90a]. Best [Ano91q, Ano92w, Ano93k, Ano93l, Ano93j, Ano93-111, Ano94p, Ano94q, Ano94-74, Ano94-87, Car94, Egl91d, Hol94, SY90, Ano91-70, Ano92a, Ano92-31, Ano92c, Ano92-223, Ano93-47, Ano93-48, Ano94-66, Ano94-89, Ano94-90, Ano94-88, Ano94-112, Ano94b, Chi92, Cla93e, CW92, Cry94, Die91, EW92, Egl94c, Fin90a, FJPR93, FJK^{+}93, Fox93, Gru90, HM94, HK94, HHP93, HP94, Hur93a, Isl92, KM94, Len94b, Len94a, Lov93b, Mal93, Nad93, Nad94a, Pou91u, Rei94b, Rya93b, Var94a, Wau93a, Way91b, Yag90d, Yag93h]. best-known [Fin90a]. best-selling [Wau93a]. bet [Egl94c]. Beta [Ano94-33, Ano94-126]. Bets [Hel91d]. Better [Alf91d, All92a, Ano92-121, Ano92-184, Ano93-38, Ano93-110, Ano94h, Ano94y, Ano94s, Ano94r, Ano94-40, Ano94-89, Ano94-90, Ano94-135, Bar92f, Ga 93c, Hur93a, Mit91c, Ras92d, Van92b, VN90a, YM93, Ano92-71, Ano92-115, Ano92-97, Ano92-150, Ano92-154, Ano92-207, Ano94-76, Cov90, Cra91c, Egl91a, Fow91, Kli90b, Lan90b, Mat94a, Min90a, Nan93d, Smi90e, Smi93a, Way93f]. Between [Gre90d, Ano94-81, Ano94-77, Api94d, Dys92, HR90, Hay90b, MPW^{+}90, Per94a, Ras91o, Rya90a, Yag90f]. Beyond [Ano91-61, Ano92-27, Ano93w, Bar91b, Bar91c, Bar92g, Bar90i, HP93, Hel90c, Hel91c, Hel92b, Key92, Kiy93a, Lip90a, Mas93, McN94, Mel93a, Mia93a, Min93a, Nan93a, Nan94a, Orr94, Pou91a, Ras91a, Ras92b, Sch92a, Sch93a, Ude93a, Ude93b, Ude93c, Ude94b, Ude94c, Ude94d, Ude94e, Vor93, Alf92e, Nan93c, Ude93h, Vos90a]. Bézier [NTS92b]. Bible [Pou90q]. BiCMOS [Rya90g]. Bicycles [Ano91r]. bidding [Sat90]. Big [Cra90e, Ess94, Gas93, Mit90a, Nan91b, Pou92g, Var94a, Ano92-215, Ano92-226, Ano93-91, Ano94-79, Ga 94d, Gut93, Hal93e, Hal94h, Ras91h, VC90r, Way91a, Way94j, Yag90l, Api94a, Per92, RUHG93]. big-iron [Yag90l]. Big-Screen [Api94a]. big-selling [Ano92-226]. Bigger [Ano94t, Ano94s, Ano94-89, Mal93, Ras90h, Ano93-141, Ano94-135, Coo90c, Gre94d, Min90b, Smi93a]. Biggies [Ras90a]. bill [Zei93, Ano92y, Ano92-108]. binary [Ano94-61, She90b]. binary-file-transfer [Ano94-61]. biography [Ano91-120]. BIOS [Ano94y]. Birmingham [Ano93-96]. Birth [Ano92x, Rya90c, She90b, Cox90, TGN94]. birthday [Tho94e]. birthdays [Rya90c]. Bit [Ano92-194, Ber93, Bus92, Dys92, Pic93, Sey93, Vor93, Ano90-95, Ano91-131, Ano92-198, Ano93-48, Ano93-74, Ano94-126, Ano94-88, Ano94-111, Ano94-119, Ano94-129, BF90, DE92, Gla90b, Gre91b, Gre94e, Hel90c, Hel90d, Hel92a, Lin90b, Lov90a, Lov91d, Lov92a, Lov94, Mas91b, Mas91a, Min92a, NTS92b, Nan93c, One91, Pou92d, Sha94b, Tho94b, Ude90g, Wsz90, Bur92, Ebe93, STS92]. bit-map [Vor93]. Bit-Mapped [STS92]. bite [Mia90a]. Bits [Ano90c, Ano90-94]. BIX [Ano91-132]. Black [Nan91f, Smi93e]. blackouts [Fie91e]. Blasts [Ano94u, Ano94v]. Blaze [AD91, TS93a]. Blazes [Tho94f]. Blazing [Smi94a]. blazingly [Lov94]. Bleepers [Ano92y]. blending [Rei94b]. Blessing [Ano92-132]. blessings [Ano90-64]. Blind [Ano94e, Ano91-191]. Blink [Pou94n]. Bloated [PTUM93, Ano90-88, Nic93]. BLOB [She90b]. Block [LB90]. Blocks [Ano90j, Ano92-187, Ano93-72, Way94g]. Blooms [She94]. blown [Bar93b, Gre94b]. Blue [Nan91b, Mia90a, Per92, RUHG93]. Blueprints [Osm90]. Blues [Ano91-131, Ano94-169, VC90c]. board [Ano91d, Ano92-116, Ano92-199, Ano92-225, Ano93-122, Ano94z, Api90b, Lov90b, Lov91d, Pou93a, Tho92h]. Boards [Ano91-160, Ano92-162, VN91d, Ano91-103, Ano93-58, Ano94-90, Bar91e, Gre93e, Pou90s, Pou94j, Tho90e, YG92]. Bob [Pou94l]. bolder [Ano92-115]. Boldly [Ano90-106]. bones [Daw94d]. Book [KUD92, KGTD92, KYT^{+}92, KRCT93, KPC93, KCU^{+}93, KAY93, Ano90-41, Ano90-69, Ano91r, Ano91-102, Ano91-179, Ano91-150, Ano91-164, KPUG93]. Books [Ano92-48, CWVS94, ESN94, GT93, GFK94, GTGW94, GTN94, Ken93b, Ken93a, Ken93c, KGEM93, KDG^{+}93, Nov94a, RUHG93, RNG94, TGN94, UPSW94, UFC94, VFGZ94, Ano90-116, Ano91-41, Ano91-92, Ano91-105, Ken92, Ude94b]. booming [Ano94-172]. Boon [Rya91c]. Boost [Ano92-204, Ano94-60, Cry92b, And90, Ano92-49, Ano92-181, Ga 94d, Hal94h]. boosted [Tho90f]. boosts [Ano91-50]. boot [Way94j]. Booting [Pou94l]. Borland [Ano90-72, Ano90-108, Ano90-109, Ano91-152, Ano91-174, Ano92-137, Ano93-118, Ano94w, Ano94-66, Ano94-117, Gas93, Gre93a, Gre93b, Han92a, Kea91a, Kli90b, MS93, Rei90a, Spi90b, Van92b]. Born [AD91, Gib90b, Ras90b, Smi93c, Ano92z]. Born-Again [Ano92z]. borrows [Daw94c]. botches [Ano92q]. Both [Ano93-111, Api94a, Car94, Nan93b, Ano92-223, Ano94-57, Ano94-112, Ano94-116, Car91a, ES94, Ga94, Gre94e, LM90, Smi94c, Smi94f, Tho94c, Van93c, Van94b, Zei93]. Bottleneck [Ano92-198, Lov94, Par90]. Bottlenecks [Fle93, AG91, Gre91b]. bought [Ano94-76]. Boundary [Nak94]. Bounds [Ano93a]. Bourne [STA92]. Bow [Yag92c]. box [Ano92-86, Ano93-109, Ano94-166, Gre93h, Ras90a, Ano94-181]. Boxes [Kiy93b]. boys [Way94j]. BPR [Ano94x]. bps [Gib90a, Hur93a]. brain [Cau93, SC92]. Braincel [Ano93-60, Cau93]. Brains [SC92]. Branch [Ano94-76, Rya93b]. Branches [Ano92d]. Brand [Ga 92a]. Brave [Way92a]. Brazil [Ban92, Pou93m]. Breach [Pou93r]. break [Bar93c, Gre90h, Gre91b, Tho93b]. breaking [Ano94-116]. Breaks [Fle93, Cry92b, Lan90c]. Breakthrough [Yag93a, Psa92]. breakthroughs [Pou94d, Smi94g]. breathes [Kli90c]. Bred [Gib90b, Yag92f]. Breed [Per93c, Ano93-111, Dav94a, Gla91b]. Brett [Gla90d]. brick [Pou91p, Ano90w, Mia91]. Bricklin [Vos90a]. bridge [Rya90e]. Bridges [Ano91-148, Ano91-34]. Bridging [Ude90a]. brief [Fie91i]. bright [Cra90q, Ano92-60]. brighter [Ano92-115]. brightest [Ano91-70]. Brilliant [Ano92-60, Rog91]. Bring [TR94a, Ano91w, Ano92-158, Ano92-183, Cox90, Mas90a, Mia93b, Pou91s, Rei92b, Spi90b]. Bringing [Ano93-64, Api92a, Ano91-86, Ano94-54, Lov93b]. Brings [Ano94-86, Gib90a, Smi90e, Ano90-85, Ano91u, Ano91-53, Ano92u, Ano92-99, Ano92-197, Ano92-199, Ano94-165, Ano94-180, Ano94-181, Api92a, Bar92d, Cal91b, Die94, Egl93c, Har91a, HL93, Joc90c, Kea91c, LDH94, Lov91b, Lov91c, Nan93c, Sch93b, Smi93e, Sta94b, Tho93g, Ull93, Ung91, VC90b, Yag92d, Yag92f]. brinks [Ude93h]. BriteLite [Ano93-127]. broadband [Fis90c]. broadcasts [Ano92-64]. Broaden [Nan91c]. broadens [Lav91]. broke [Sto93]. Broker [Ano94h]. brought [Ano92-161]. Browser [NTS93b]. Browsing [Ste91a]. Brute [Hal94h]. bubble [LB91, Joc90a]. Bubble-Jet [Joc90a]. bubbles [Joc90a]. Bubbly [Ano91s]. Buck [VC90f, VN90b]. Bucket [Tho91a]. Buckley [Ano92-92]. Budget [LS91a, Tho94c, VC90f]. Buffer [Bri90]. Bug [Cra90d, Sha94a]. Build [EWT90, GFK94, Hol93b, Key92, Par92, Yag91a, Ano92-147, Ano92-187, Ano94f, Api94b, Api94c, Cau91, Egl94g, Ga 94b, Gre93d, Lu93b, RH93, Smi90b, Ude93i]. builders [Hay94b]. Building [Ano90j, Ano91-38, Ano92-187, Ano93-72, Ano94y, Gre93a, Mil94b, Rei94b, Way93d, Ano92-44, Ano92-161, EB90b, Smi90a]. Builds [Van92b, Ano92-56, Bry94b, Gil94, Pou91t, Tho93c, Yag93f]. Built [Ano92-88, Ano92-202, Ano90-92, Ano91-42, Ano91-83, Ano91-172, Daw94a, MS93, RT91, Tho94b, Tho94f, Wsz93a]. built-in [Ano91-172, Daw94a, RT91, Tho94b, Tho94f]. bulge [Ano94-102]. Bullet [Cox90]. bulletin [Ano94z]. Bulletproof [Nan93c]. Bumbling [UT91]. bunch [Ano92-33]. bundle [Ano92-211, Lab92a, Van93a]. Bundled [Nan94d]. Buoyed [HS94]. burden [Ano92-195]. burgeoning [Wal94c]. Burn [Ano94-89, GN93]. burned [Ano91-55]. Bus [Ano92-94, Ano92-198, Ano93-85, Ano94t, Ano94-90, FJPR93, Gla90g, Gla90d, Kay94a, Ken94, WK90, Ano92-65, Ano92e, Ano92-169, Ano94-72, Ano94-111, Bry94d, Cla94b, Ess94, Fox93, RH93, Ric91, VC90a, VC90d, Ano93-137, Ano94-90, Ga 93e]. bus/accelerator [Ano92-169]. buses [Ano92-198, Bry94c, Cla94b, Kay94a, Mar92a]. Bush [KUD92]. Business [Ano90k, Ano91x, Ano93-105, Ano94-172, BGE^{+}92, Bar91g, Cra91a, Mia93b, Ras91l, Ste94b, Tur90, Yag90b, Ano90b, Ano90-50, Ano91-55, Ano91-123, Ano93s, Ano93-113, Ano94z, Ano94x, Ano94-64, Ano94-68, Ano94-89, Die94, Egl93b, Fie90e, GT93, HR91, HM94, HK94, HHP93, Joc91, KM94, Ras90c, Ras90o, Ras91c, Ras92b, Rei93d, Ste91d, TGN94, Way94e, Egl93b]. Business-Card [Egl93b]. Businesses [Ano94z, Die93b, Pou93o, Ric94]. businesspeople [Mos93, Ste94b]. Busters [Ano94-97]. busy [Ano94-150, Pou94q]. Buttons [STS92]. ButtonWare [Ano90-73]. Buy [Ano94-90, Ano91-55, HP94, Ras91l]. Buyer [Tho91b, Ano91-110, DW90]. Buyers [Yav94, Ras91n, Rei92b]. Buying [CV90a, Hal93a, Pac94, Red92a]. bypass [Ste91c]. Byte [Rei92a, VN90b, All92d, All93k, Ano90m, Ano90n, Ano90o, Ano90l, Ano90-112, Ano90c, Ano91t, Ano91-88, Ano91-89, Ano91-108, Ano91-115, Ano91-122, Ano91a, Ano91-170, Ano91-189, Ano91-186, Ano92p, Ano92q, Ano92r, Ano92a, Ano92w, Ano92-28, Ano92-27, Ano92-31, Ano92-47, Ano92-88, Ano92-126, Ano92b, Ano92-173, Ano92-188, Ano92-189, Ano93d, Ano93e, Ano93f, Ano93-50, Ano93-59, Ano93-57, Ano93-94, Ano94p, Ano94q, Ano94-27, Ano94-74, Ano94b, Ano94-147, Ano94-148, Ano94-149, Ano94-168, AMW90, AWGY90, AG91, Api93b, BGE^{+}92, BEW92, Cal93, CW92, CS92, Die91, DE92, EWT90, Egl91a, ET91, EW92, Egl92a, Egl93c, Egl94e, Fid92, FJPR93, GE92, GAW93, Ga 93d, Ga 94d, GR94, Gar94, Gre90b, GETY91, GD92b, GW92, Gre93a, Gre93b, Gre93d, Gre93c, Gre93e, Gre93f, GW93, Hur93a, Joc92b, Lab92b, Mit91a]. BYTE [Mit91b, Mit91c, NW92, Nad93, Nad94a, Pou90o, Smi90d, Smi93b, SE94, Tho91c, TG91, Tho92a, Tho93h, Tho94e, Wsz93a, Wsz93c, Yag92e, Yag92b, Ano92s, Ano92t, VN90f, Gre90h, Lan90c, Min90c]. Bytes [Ano90c, Ano94-150].

C [Ano90-108, Ano94-66, Han92a, Kea91a, Sha94b, Spi90b, Ano90-56, Ano91-78, Ano91-129, Ano91-163, Ano91-177, Ano92v, Ano92-66, Ano94f, Ano94-28, Ano94-66, Ano94-117, Ano94-180, Api94b, BR93, Cut91, Eug90, FH93, Ga 94c, GR90, Gre92, Gre93b, Gre94b, Gre94e, Han92a, Han92b, Hel90a, Hol93a, Hol93b, Hur93b, Kea91c, Kli90b, Mas90a, Min92c, Nag93, Nan90a, Pou90h, Sch90c, Sin94, Spi90a, Spi90b, Ude94c, Way94d, Wig90, Woo90c]. C-Cube [BR93, Way94d]. C-Phone [Ano94-28]. C/C [Ano94f, Gre93b, Gre94e]. CA [Ano94k]. Cable [Ano94-166, Rei94b]. Cables [Ano94-170, Smi94f]. Cabling [Ano94-56, Egl91b, Ras91j]. Cache [Ano92-215, Bac94, VN91a, RNG94, RNG94]. cache-memory [RNG94, RNG94]. Caching [AG91, Ano92-215]. CAD [Ano92-149, Ano93-137, Ano93-126, Ano94-29, Api90b, Bis91, Cal91a, Cra90f, Cra90b, HM94, HHP93, KM94, Ric92, Sac90, Ude90b, Ude90e, Wau93a, Woo91a]. CADVance [Ude90b]. Caere [Egl94c]. CalComp [Mia90b]. calculate [Ras90n]. Calculating [Mar90a]. calculator [Ano90-36]. calculus [UPSW94]. calendars [SE94]. Calera [Egl94c]. Call [Ano94-62, Len94b, Pou92d, Rei94c, Yag92c, Nan91l, RH93]. called [Fis90a, Way90a]. Calls [Nan91i, Ano94-37, Ano94-103, Kor93, Mal91a, Nie94, Vog90]. came [Ano94-168]. camera [Chi92]. cameras [Egl93d]. camps [SH90]. Can [Ano94-30, Car91a, Cra90f, Egl91a, Hog90, Min90i, Ras92b, Rya90e, Rya91a, VN91a, Ano90-94, Ano91-35, Ano91-66, Ano91-130, Ano92-34, Ano92-80, Ano92-126, Ano92-181, Ano92-186, Ano92-212, Ano92-213, Ano94f, Ano94-69, Ano94c, Ano94-148, Api94a, Api94b, Api94c, Bar92b, Bus92, CV90a, Chr94, Cla93a, Cra90h, Cro93b, Cry94, Dan90, Dav90, Der92, Die94, DH90, Egl90a, Egl93e, ES93, Egl94b, Eva90, Far90, Fie91a, GP90, Gib90a, Gil94, Gre91a, Gre92, Gre93a, Gre93d, Gre94b, Gre94e, Hal94a, Har91c, Hed93a, Hed93b, Kay94a, Ken90, Kor93, Laf91, Lie93, Lin90a, Mat94b, Mia90a, Nad90d, Nan91l, Nan94b, Nel91b, Nov94b, Pou90i, PS94, Pou94c, Ras90d, Ras90c, Ras90h, Ras90k, Ras90o, Ras91e, Ras91g, Red92b, Rei91c, RH93, Rei94a, Rob91b, ST93, Sha94c, SB94, Sin94, SE93a, Smi94a, Smi94b]. can [Smi94f, Ste91a, Tho90d, Tho90h, Tho91a, Tho93d, VC90s, VN90d, VN91d, Wal94d, Way91c, Way93i, Way94d, Way94e, Way94j, WB90, Wsz93a, Yag91d, Yag93b, Yag93e]. Canon [Ano93-76, Egl92b, Joc90a]. Can’t [Ano91-195, DE90, KPC93, Lin91b, Ano93g, Mic91, VC90l]. Cap [Ano94-113]. capabilities [Ano94-61, Bar90k, Bry94c, Fri90, Gre94e, Hay93, Hur93b, HC93, Lin94, McN94, Mul94, Ras90i, RT91, SL92, Tho91e]. capability [Ano91n, Ano92-64, Ano92-126, Rei93e]. Capable [Api94a, Ano94-143]. Capacities [Ano93k, Ano93l, Ano93j, Ano93-47, Ess94, Gla90c]. capacity [Ano91-63, Ano91-134, Ano92-102, AWGY90, CP94, Rei91a, Way94c]. capitalize [San90a]. Captains [GE92]. captivating [Yag91a]. capture [Ano93-122, Ano94-64, Smi91b]. capturing [Wal94d]. Card [Ano94-111, Alf92g, Ano93-74, Ano93-119, Ano94l, Ano94-57, Ano94-176, Api94a, Die93b, Lov90a, Nan94b, Ano93-30, Egl93b]. Cards [Ano93-48, Ano93-105, Ano94-90, Ano94-119, AG91, Ano94-106, Ano94-88, Api91c, Egl93b, Ken94, Pou94n, Rei91a]. care [Hal93d, Wal94b]. carefully [Ano94k]. Carls [Car94]. Carnegie [LHR90]. Carnot [Ano91-67]. carpet [Hel90e]. carried [Ras90g]. Carroll [HC93]. Carry [Alf91e, Ano93m, Yag93b]. cartridge [Egl90c]. Case [Ano94-136, Api93c, Min91a, Per93a, Ras92e]. cases [Ano94x]. Cash [Ano94-46]. Casio [Ano93i, Ano93-67]. Casting [Pou90l]. casual [Ano94g]. cat [MG90]. catastrophe [Ano91-169]. Catch [Bar90k, Har91a, Rei90d, VN91a, Cos94, Way93c]. Catch-Up [Rei90d]. catches [MB93]. catching [Pou94d]. Categorizing [Tho91a]. Category [Tho91a, Ano94-79, Eva94]. cathode [Ano94-35]. Cats [Pou92g, Hed93b]. caught [Dau91]. Cautionary [Ano94-43]. Cautiously [Ano93t]. cc [ES94, ES94]. CD [Alf93, All93f, Ano91u, Ano92-60, Ano92y, Ano92-136, Ano93s, Ano93x, Ano94-31, Ano94-87, Ano94-152, Bos93, CWVS94, Cra91l, Egl93e, ESN94, GT93, Ga 93d, GFK94, GTGW94, GTN94, Hal93a, Har91b, Hyo94, KYCD92, KUD92, KGTD92, KYT^{+}92, KRCT93, KPC93, KCU^{+}93, KAY93, Ken93b, Ken93a, Ken93c, KGEM93, KDG^{+}93, Min93b, Nov94a, Per93c, Pou94c, Pou90q, Pou90t, Pou90v, Pou91m, Pou91p, Pou93q, Ras90c, Ras90f, Ras91a, RUHG93, Rei94a, RNG94, TGN94, Tho94c, UE93, Ude93d, Ude93h, Ude93j, UPSW94, UFC94, VC90n, VFGZ94]. CD-R [Tho94c, UE93]. CD-recordable [Hyo94]. CD-ROM [Alf93, All93f, Ano91u, Ano92-60, Ano92y, Ano92-136, Ano93s, Ano93x, Ano94-31, Ano94-87, Ano94-152, Bos93, CWVS94, Cra91l, Egl93e, GT93, GFK94, GTGW94, GTN94, Hal93a, Har91b, KYCD92, KUD92, KGTD92, KYT^{+}92, KRCT93, KPC93, KCU^{+}93, KAY93, Ken93a, Ken93c, Min93b, Nov94a, Pou94c, Pou91p, Pou93q, Ras90c, Ras90f, Rei94a, RNG94, TGN94, Ude93d, Ude93h, Ude93j, UPSW94, UFC94, VFGZ94]. CD-ROMS [GFK94, CWVS94, ESN94, GT93, GTGW94, GTN94, Ken93b, KGEM93, KDG^{+}93, Nov94a, Pou90q, Pou90t, Pou90v, Pou91m, Pou91p, Ras90f, Ras91a, RUHG93, RNG94, TGN94, Tho94c, UPSW94, UFC94, VC90n, VFGZ94, Ken93a]. CDD521 [Ude93d]. CDPD [Ano93m]. CDR [Ano91u]. CDR-1000 [Ano91u]. CDs [Ano94-87, Ano94-152, Egl93e, Hyo94]. CeBIT [Ano91v]. Cedar [Bis91]. CEG [Ano91-106]. Ceiling [Ano92-44, Ano92-157]. Cellular [Bry94b, Ano93m, Ano93-134, Ano94-145, Ano94-174]. Center [Cla91a]. Central [Slo91]. centralized [MH91, Slo91]. centrally [Ano92-222]. centric [Ano94-113, Egl93a]. Centris [Tho93a]. centry [Sal94]. CEO [Ano91r]. CH [Kli90a]. chaining [EB90a]. Challenge [Ano93-71, Ber93, Egl93a, Hal94c, Ups90]. Challenged [Hay94b]. challenger [Del93b]. Challenges [Rya94b, Ano90-38, Ano92-133, Ano92-135, Ano94e, Osh90]. Champagne [LS91a]. champion [Ano94-100]. champs [Ano92l]. Chance [Cra90k]. change [Ano91-31, Ano92-41, Ano92-203, Ano94-125, Ga 93d, McA90, Ras91b, Ras91l, RUHG93, Rei93d, Rei94f, Rog91, Rya90e, Sey93, Sma93b, Way94a]. Changes [RH93, Cos94, Min90e, Rya91c, Slo91]. Changing [Fie91f, Wat94, Alf91a, Ano92d, Fis90c, Hog90, Mar90b, Rei94e, Swa93, Web92]. Channel [Bry94c, Way94c, Alf90c, Alf92b, Ano92-198, Mia90a]. channels [Fri94c, Pou90j]. Chaos [Pou91m, Ano91-179, Pou90m, Pou90r, Pou92j, Pou92n, Pou92k, Pou93j, Pou93l, Pou94l, Pou94q, Pou94s]. Chaotic [Pou90l, LDH94]. Character [EWT90, Way93a, Way93i, Pou93a, WS91, Gre90k]. characteristics [BL94]. characters [Gre90k, Smi90c]. Charge [Dan90, Tho90b, Pou93h]. chart [Bar92d]. chart-making [Bar92d]. Charting [Nie94, Ras90d]. charts [BGE^{+}92]. Chase [Ano93-45, Way94e]. chassis [Ano90k]. Cheap [AT92, Joc90c, Ras90e]. cheaper [Ano92-97]. Check [Ano91-33, Ano94-98]. Checker [Egl91a, Ano92-160]. Checking [Ano91-73]. checks [Hur93b]. Cheetah [MG90]. Cheshire [Hed93b]. chess [Ber93]. Chicago [Ano94-32, Ano94-33, Ano94-98, Bry94e, Hal94g, She94, Ude94i, Ude94j]. Child [YS92]. Children [Fou91]. China [Rei93a]. Chip [Ano90-46, Ano94-81, Ano94-108, Rob90b, Way93c, Ano91-106, Ano91-153, Ano91-165, Ano91-173, Ano92-105, Ano92-183, Ano93-75, BS91, Bar90j, GK91, Gut94, Hal93d, MG91, Pou94e, San90a, Sta94b, Ano91-153]. Chips [Ano91w, Ano92-105, Ano93-75, Bar90i, Hal93e, Lin90a, Mar90b, Rie91, Rya90f, Ano92-161, Ano92-202, Ano93-45, Ano94-129, Way91c, Way92b, Way94i, Ano92-202, Ano91-153]. Choice [Ano91a, Ano92b, Ano93-94, Ano94-30, Ano94b, Pou91u, Pou92l, Pou92k, Pou93r, Pou94m, Pou94s, Tho93f, Api90c, Mar90b, Mas92, Ude94k]. Choices [Alf92c, FJPR93, VC90l, Bra90]. choose [Ano92-31, Ano94-89, HM94, HHP93, Len94b, Mic91, Ras90g]. Choosing [Ano92p, Ano93-50, Nan91d, Ano91-195, Pet91]. chores [Fis90b, Gil94]. Chorus [Pou91i, Pou94a]. Chorus/MiX [Pou94a]. chosen [Ano90l]. Chris [Pou91k]. Christopher [Loc91]. chronicle [UFC94]. Cinema [Ebe93]. Circle [Coo90a]. circuit [Dio94]. circuits [Way94i]. Cirrus [Ano93m]. CISC [Ano94-65, Ano94-118, Ano94-153, Fie90l, Hal94c, Pou93g, Tho90f, Ude94k]. Cities [Pou90v, Ano92-60]. Citrix [Ude91b]. claiming [Ude93h]. clarifications [Ano94-95]. clarify [McN94]. clarifying [Ano92-91]. Clarion [Ste90]. Claris [Ano91-49, Ano92-54, Ano93-137, Gib92]. ClarisWorks [Hal93b]. Clarity [Ano92-204]. clash [Pou94j]. Class [Ano91-78, Cut91, Yag92d, Ano92-195, Ano92-196, Ano94-57, Eug90, Hol93a, Kea91c, Nag93, Pou93m, Row93, WE92, You93, GK91, Nad90f, Ras90b]. classes [BR93, Kiy93b, May94a]. Classic [Ano90g, Ano91x, Ano91y, Ano91z, Ano92-29, Ano92-30, App91, Tho93h, Ano90-69, Ano92-83, GT93]. classics [Pou90q]. Clean [Pou94h, Pou93l]. CleanSheet [Pou94h]. Clear [Par90, Ano94-166, VN91c]. clearer [Ano93-30]. Clearing [Fri94a]. clears [Wsz93b]. clever [Ano91o]. click [Yag90h, Cla93c]. Client [Cro93b, Smi94b, Ull93, Ano92-200, Bor93, Ga 93b, Gre92, Gre93a, HH94a, Kor93, Nan91i, Sch93b, Slo91]. Client/Server [Smi94b, Ull93, Cro93b, Ano92-200, Bor93, HH94a, Kor93, Nan91i, Sch93b]. Clients [Ano93-106, Ano91-88, ES93, Nad94b, Nan93g]. climb [Var94c]. clip [Van93a]. Clipboard [Ano92-199]. Cloak [Gre90a]. Clock [Alf92a, Yag90i, Ano92s]. Clocking [Api92b]. Clocks [Pou93b, Ano92-173]. ClockWise [Yag90i]. clonable [Gla91b]. clone [Ano92-32, Daw94d, Fie90j, MG91, Ude90c]. Clones [Ano92-96, Ano92-36, Ano94-159, Hay90b]. Close [Ano93u, Pie94, Ras90n, Ude92d, Ano94w]. Close-By [Ano93u]. Close-Up [Pie94]. Closer [Ano93-93, Gib90a, Har91a, HL93]. closes [Smi91a]. Closet [Ano94-76]. Clothing [Ano92-36, Yag91g]. Club [Ano91-170, Egl90b]. Cluster [Gut93]. clutter [Wsz93b]. cluttered [Ano93-105]. CMIP [Fis91a]. Co [Bar90j]. Co-ops [Bar90j]. coalitions [Ano94-171]. coaxial [Rei94b]. COBOL [Ano91x, Gre94a]. Code [Ano90-88, Ano92-31, Api93c, Nan90a, Tho94g, Ano91-43, Ano92-175, Ano93-141, Cro93a, Egl94g, Gre94b, Gre94e, Hay94b, Lan94b, Mat94b, Min90c]. codes [Gre90a]. CodeWarrior [Ga94]. codifying [Wal94d]. coexists [Ano94-175]. cofounder [Bar91f]. coherency [VN91a]. Coherent [Ano92-32, Fie90l]. Cohesion [Dej94]. Coin [Cra90q]. coincides [Ano90-41]. Cold [Red92b, Ano94-35]. Collaboraative [Cla93d]. collaborate [Smi94a]. Collaboration [GR94, Ano94-132]. Collaborative [HL93]. colleague [Pou93e]. Collecting [Mia94a]. collection [Pou94c, Smi93a]. collective [Bux90a]. collects [Ano92-220]. college [Pou92h, UFC94]. collide [Ano91-102]. Color [Alf91a, Ano91-27, Ano92-33, Ano92-90, Ano92-115, Ano92-179, Ano93n, Ano93-128, Ano94-34, Ano94-89, Api91a, BF90, Coo90b, Egl92b, Hal94g, HHP93, Joc91, Len94a, Lov90a, New93, Ras92c, Rei93b, Sma93b, Tho92a, Tho93b, Tho93g, Van94a, VC90b, Vau90, Yag93b, Zei93, Ano90-95, Ano91-62, Ano91-131, Ano92u, Ano92-27, Ano92-39, Ano92-71, Ano92-116, Ano92-131, Ano92-199, Ano93o, Ano93-90, Ano93-125, Ano94-182, Api94d, Bar90b, Bar91e, Cla91a, Die90a, DE92, HM94, KM94, Mac91, Nel91a, Rog91, Smi90f, Tho90h, Tho92b, Tho94b, Tho94a, Ano91-131, Ano94-89, Tho90e]. Color-Matching [Ano94-89]. Colorado [Ano90-57]. ColorCard [Tho90e]. ColorCard/24 [Tho90e]. ColorFrame [Ano93o]. colorful [Egl92b, Egl92c, Ras92c]. Colors [Nel91a, Tho92b, Vau90]. ColorScript [Tho93g]. ColorStudio [Ano91-124]. Colossal [Ano90-88]. colourful [Api90a]. Column [Pou92o, Pou92m, Pou92n, Pou93q, Pou93r, Pou93s, Pou93t, Pou93f]. columnist [Fie90a, Fie90f, Pou93i]. columnists [Ano91-189]. combination [Ano92-131, Ano94-31, Ano94-76, Cau91, Lin93].

Combine [Hel91b, Yag93b]. combined [SE94, Wal94a]. combines [Ano94-145, Rog91, Tho90a]. Combining [Hed93a, Ano92-183, Ano94-76]. combo [Ano92-169]. Comdex [Mal93, Ano92w, Ano94p, Pou90n]. Comdex/Spring [Ano92w]. Come [Ano91-99, Ano91-196, Ano90-84, Ano92-159, Ano93-138, Hel91f, Nan91j, Pou91d, Rya90g, RP90, Tho93f, Yag91d]. Comes [Api91a, Bar91f, Pou94d, Ude90e, Vau90, Ano92-58, Ano93-105, Cra91h, Egl94f, Fie90d, Gru90, Len94a, Lin90a, Lio90a, Per92, Sch93b, She93a, Ude93j, Van93a]. Cometh [Ras91h]. Coming [Ano92-68, Ano93-51, Ano94-71, Ano94-128, Nad91, Red92b, Rei93d, Rei94g, Sac90, Ano91-191, Ano92-48, Ano92-210, Bor93, Mas91a, Rei90b, SC92, Vas93b, Way94h]. command [Gla90d, Cra90d]. commands [Ano93-37, Die94]. Commentary [Cos94, Cri93, Eva94, Ga 94a, Hal94e, Hal94d, Ken93c, May94b, Mos93, Nov94b, Pac94, Rei94c, Sha94a, Sto93, Swa93, Ude94f]. Comments [Ano93-52, Ano93-54]. commercial [Ano94-35, Daw94d, Sal94, Sta94a]. Commercialize [Ano94-81]. committee [Ano94-111]. committees [She91]. Commodore [Hal94d, Ano90f, Ano90p, Ano93-53, Ano94-99, Yag93c]. Common [Ano91x, GD92a, Ano91s, Gla90d, Gre91a, Gre90g, May94b, Nan92e, Orr94, Die93a]. commonly [Ano92-30, Fie90a]. CommonView [Sch90c]. Communicate [Ano94-145, Ano91-33, Fie90i, MW93b, Ras90h, Ras91g]. communicates [Ano94-178]. communicating [ES94]. Communication [Ano92-185, Fri94c, Ano94-145, MPW^{+}90]. Communications [Die93b, Egl94b, Mel93a, MW93b, MW93a, Rya93a, Ano91-136, Ano92-151, Ano92-185, Ano93v, Ano94g, Ano94-77, Ano94-113, Fie90f, Fis90c, Ga 93b, Gre90d, Gre90g, HS94, Hay93, Hur93a, Jof93, Kre91, Mal91b, Mia93b, Mit91c, Pou93e, Rei93d, Rob91b, Sta91b, Sta91a, Tho94f, Tho94i, Ano90z, Ano94q, Ano94-77, Smi94e]. communications-centric [Ano94-113]. communications-oriented [Ano94g]. Communicator [Ano93-88, Ano94-179, Die93b]. communicators [Way94a]. communities [Ano94-96, Hal94e]. community [Ano93-98, Ano94-96, Ano93-98]. compact [Ano90-67, Cla90a]. companie [Ano94-84]. Companies [Ano90-58, Ano94-166, Ano90-54, Ano90-91, Ano92-194, Ano93-68, Ano94q, Ano94-57, Ano94-63, Ano94-113, Ano94-116, Ano94-165, Cra90h, Cra91l, Fri94b, Lov94, Ras91b, Rei94b, Sey93, She91, SE93a]. companion [Ano90v]. company [Ano92-108, Hal94d, Hel91a, Ras90i]. companywide [Ano91-38, Bon90a]. Compaq [Ano91-28, Ano92z, Api92b, Cla90a, Egl92c, Ga 94d, Lov92b, Mit91a, Nad90b, Nan93b, Per93b, Yag90d]. compare [Osm90]. Compared [Nan92e, CV90a, Egl94c]. compares [Ga 94d, Mit91b, NW92]. Comparing [Gre90c]. comparison [Ano94-94, Api92b, Van93b]. Comparisons [Ano93-87]. Compatibility [Ano91-29, Cra91j, Ano90-79, Ano91-151, Egl90c, Gla90b, Gre93g, The94]. Compatible [AMW90, Ano90-110, Gla91c, Ano94d, Die90e, Ano92-110, Hay90b, Pou90c, Sch92c]. Compel [Ano93p]. Compete [Yag91c, Ano93-67, Ano94-35, Rya94b]. Competing [Gla90c, Ano92-125]. Competition [Ano94-35, Ano91-187, Ano92z, Ano93-75, Ano94-31, Ano94-116, Bry93b, Cra90b, Hay90b]. competitive [Dys92, Ude90e, Van94b, You94]. Competitor [MG91, Egl94f, Gas93, Ros90b, Woo90a]. competitors [Hal94h, Hal94b, Pou93g, Van93c]. compiled [Tho94h]. Compiler [Ano94-36, Ano90-81, FH93, Gre94e, Han92b, Hay94b, Kea91a, Nan90a, Pou90h, Sch92c, Sha94c]. Compilers [Ano93q, Ano94-124, Hay94b, Mas90a, Sha94c, Gre93b, Lan94b]. complement [Ano94-76]. Complete [Die93b, Gre94e, Smi93d, Ste90, Die93b]. completely [Fie90b]. completing [Ano94-107]. Complex [Ano94-89, Ano94-147, Bur92, DH92, Laf91, Mul94, Ran90, Ric94, Smi94a]. complexities [Alf93, Ano94-165, Gre93b, Loc91]. complexity [KRCT93, Mos93]. complicate [Cla93b]. Complicated [Mia90b]. Component [Ano93-85, Ude94g]. components [Ano92-222, Ano94-76, Cox90, GK91]. Componentware [Ude94g]. composite [Ano91-167]. Compound [Ano94-78, Cro93a, Gre94b, Rei94e]. comprehensive [Ano90-56]. Compressing [Api93a, Wal94a]. Compression [Ano92-129, Ano93-121, Ano94-41, Ano94-94, Api91d, Bar90b, Hal94a, Way94d, Ano92r, Ano92-92, Fri94c, Pou93i, Ude93f, Ude93h, VN90b, VN90d, VN91e]. compressors [GW93]. Compromise [Ano94-32, Len94a, Ano92-201, Len94b, VN91c]. Compromises [Ras92c, Ano92-35]. Compton [KAY93]. CompuAdd [Ude90c]. CompUSA [Api93b]. computation [Gri94]. Compute [Ano90e, Sch92a, Ano94-107, Ras90n]. compute-intensive [Ano94-107]. Computer [Ano90q, Ano90y, Ano91-93, Ano92-222, Ano94m, Ano94-47, Ano94-119, Bar90j, Chi92, Ebe93, HL93, KGEM93, Pou92e, Pou92m, She91, TS90, Ude94h, AR91, Ano90e, Ano90a, Ano90d, Ano90-69, Ano90-84, Ano90-97, Ano90-100, Ano90c, Ano91-108, Ano91-186, Ano92k, Ano92-48, Ano92-52, Ano92-68, Ano92-69, Ano92-76, Ano92-109, Ano92-135, Ano92-127, Ano92-142, Ano92-138, Ano92-160, Ano93-37, Ano93-100, Ano93-116, Ano93-130, Ano94u, Ano94-65, Ano94-148, Ano94-149, CWVS94, Cra90k, Cry94, Die94, Fow91, GP90, Kay94a, Len94a, LAN92, MPW^{+}90, MR93, Mia92a, Mos93, Pet90, Pou90g, Pou91b, Pou92a, Pou91p, Pou91r, Pou91q, Pou91t, Pou92g, Pou92h, Pou93m, Pou94k, Pou94i, Red91a, Rei91a, Rei91b, RPRM92, Rie91, Rob90a, RP90, SC92, Sma93a, Sto93, TGN94, UFC94, Vas93b, VFGZ94, Way90c, Way90a, Way93f]. Computer [Pou94e]. computer-based [Cry94, Die94]. computer-graphics [Ano93-116]. computer-monitor [Ano92-68]. computer-output [Ano92-52]. Computerized [Wal94b, FT90]. Computers [All92a, Ano91-39, Ano92-34, Ban90, Ber93, GW92, KPC93, Laz93, Mit91b, Red93, RPRM92, Swa93, Yag91b, Alf92g, Ano90-27, Ano91r, Ano91-77, Ano91-146, Ano91-150, Ano91-170, Ano92l, Ano92d, Ano92-79, Ano92-77, Ano92-144, Ano92-194, Ano93-90, Ano94-135, Ano94-152, Bar91a, Cla94a, Cra91h, DeH93, DWN91, Egl90d, Egl93a, Fie90b, GP90, GK91, Har91c, Ken93a, Laz90, LHR90, Mil93, Pac94, Pou93f, Pou90w, Pou91s, Ras92d, Rya90a, Tho92d, TS93a, Tho94b, WS93, Yag90c, Mit90a]. ComputeServer [BBWR90]. Computing [ASS92, Ano91-31, Ano91-30, Ano92-63, Ano92-62, Ano92-61, Ano92-79, Ano92-135, Ano92-156, Ano94-95, Bar92b, Bar90h, Bar90g, Cou92, HL93, Mas91b, Mas91a, Mil94b, Nad91, Pou93b, Pou93f, Pou91p, Sel92, Smi90c, Way92b, Ano90w, Ano90-107, Ano91-38, Ano91-126, Ano92-35, Ano92-106, Ano92-194, Ano92-202, Ano92-209, Ano92-218, Ano93-56, Ano93-58, Ano94z, Ano94-101, Ano94-113, Ano94-149, Api94d, Bar90i, Cla93d, Cra91f, Cro93b, GW92, Hal94d, HK94, Lu92, MB93, Nel91b, Pou90u, Pou94p, Ras90b, Rei94e, Sey93, Smi94c, Tho92h, Tho94e, Ude90a, Ude91b, Ude93l, Ude94g, Way92a, Yag93d, Ano94q, Ano94-48, Ano94-189, Egl93a]. concentrating [Ano90-78]. concept [Cra91d, Smi94c]. Concepts [Web94, Ken90, UPSW94]. conceptual [Gib90b]. Concern [Ano91-32, Ano94-103]. concerned [Bos93]. concerns [Ano91-149, RPRM92]. concluding [Smi93e]. Concurrent [GR90, Sch93b, Wal94a]. conference [Pou94k]. conferences [Ano94-182]. Conferencing [Ano93u, Dav94a, Rei94a, Ano94-183, Egl94a, Ga 93b, YM93]. Confessions [Mas93]. configuration [Hal94i, IC94, Smi94e]. configures [Pou92g]. Configuring [Pou90h, Ken94]. Conflict [Cut91, HR90]. conflicting [Pet91]. conflicts [Ano93d, SU93]. confuse [Red92a]. Confused [Tur90]. confusing [Mur93]. Connect [Ano91-33, Api91b, Bry94a, CV90a, VC90q, Ano92-199]. Connected [Ano94-168]. Connecting [Ano94-37, Kay94a, SE93b, Ude93a, Ano92-199, Nad94b, SU93, Way94g]. Connection [Fis90d, Ano91-41, Api94c, Fri94c, UPSW94]. connection-oriented [Fri94c]. Connections [Ano90-53, Ano93-48, Gre91a, Nad94b, Ras91b, Ras92a, Ano92-177, Ano94-37, Bry94c, Bry94d, Tho94f]. Connectivity [Ano91-34, Cra90h, Hal93f, UY91, Ano92-87, Ano92-85, Nan91c, Nan93d, Ude93a]. connector [Gib90a]. connects [Ano92n, Ano94-180, Egl94d, Smi94e]. Conquer [Jef93a, WS91]. conquered [Bra90]. Conquers [Rei93e]. Conrad [Ano92-201]. cons [Ano91-90, Ano91-188]. conscious [TH93]. consider [Cla90b]. considerations [Pet91]. Consistency [Gru90]. consistent [Ras91c, VC90g]. Consortia [Bar90j, Ano92-125]. consortium [Ano91-87]. Conspiracy [Saf93]. construct [Kiy93b]. constructing [RUHG93]. construction [Gre93a, Pou94h]. Consultant [All93i]. consultants [Ano93-97]. Consulting [Hed93b, Wil94]. consumption [Ano93-132, Cla90b]. Contact [FW93]. contacts [Egl93b]. contains [Han92a]. contend [Cla94b]. contender [Ga 93b, Gib92, Way94g]. Contenders [Gla91a]. Content [Cro93a, Way91c]. content-addressable [Way91c]. Continental [Bar90h]. Continue [Ano93-45, Ano93s, Ano94-113, Ano94-159, Hal94b]. Continues [Pou91s, Tho91b, Ano92-109, EB90a, Hal94b, Pou94l, Zei93]. contracts [Ano94-98]. contribute [Ude93g]. contributed [Ude93g]. contribution [Ano91-42]. contributors [Ano91-189, Ano92a]. Control [Ano92-34, Ano94-62, Ano94-143, Cla94a, Coo91, Mit91c, NTS93d, Slo91, Yag93e, Ano92-66, Ano92-183, Ano94-81, Ano94-148, Ano94-181, Dan90, Dej94, IC94, Pou93f, Tho90b, Ano94-148]. controlled [Ano92-130]. Controller [San90b, AG91, Cra90c, Lab92b, Ano90-40]. controllers [Pou94j]. Controlling [Ano94-38, Cla94a, Kay94b]. Controls [Ude94c, Cry94, Ude94g, Ano94-117]. controversial [KPUG93]. controversy [MG91]. Contura [Egl92c]. convenience [Mia90e]. convenient [Ano94-149]. conventional [Ras91o]. Conventions [Ano94-122, Pou90v]. converge [Pou94f]. convergence [CP94]. converse [Ano94-182]. converses [Cra91e]. conversion [DH92, Tho94f]. convert [DM90]. Convertible [Ano90g, Ano92-35]. Converting [Ano94-39, Sha94c, Smi94b]. cook [Ano91-111]. cool [Ano93-44]. Cooperation [Daw94d]. Cooperative [Egl94d]. coordinate [Dan90]. copier [Ano92-43]. Copiers [Rei94g]. copies [Ano94-175]. Coping [Rya90a]. copper [Rei94b]. Coprocessor [Ano92-162].

Coprocessors [Gla90f, Mar90a, Ano92-196]. Copy [Mia92b, Bar91e]. copyright [Ano90e]. Copyrights [McG90b]. Cor [Mia90d]. Cord [Ano94-165, Nan94b]. core [Wal94e, Way93d, Ano90-57]. Corel [Ros90b, Van93a]. CorelDraw [Ano92-82, Ano94-40, Van92a, Van93a, Van94b]. CorelSCSI [Lab92b]. Corner [Ano92-184, Ano92-185, NTS92a, NTS92b, NTS93c, NTS93d, NTS93a, STA92]. Cornerstone [Fin90b]. Cornerstones [RLU^{+}91]. cornucopia [Ano92u]. Corp [Cra90d, Wal94a]. Corporate [KYCD92, Ras91b, Ras91c, Die94]. corporatewide [SE94]. corporation [Gre93a]. corporations [Ano94-149]. Correspondence [Wal93, Die94]. Cost [Ano91l, Ano91-79, Ano91-122, Ano94-57, Ano94-70, Ano94-106, Cal91a, Lov91d, Ano90z, Ano90a, Ano91-37, Ano91-194, Ano92-65, Ano92-72, Ano92-87, Ano92-199, Ano93-50, Ano93-70, Ano94-54, Ano94-112, Bry94a, Cla94a, Egl93c, Egl94c, Hed93a, Lab92a, Lin91a, Nad90e, Smi94g, Tho90h, Tho93f, Tho94a, UM90, Way94g, Yag90c, Api90c, Bar90c, Dio94, ESN94, Joc90b, Lov91b, Ude90c]. cost-effective [Cla94a]. cost-effectively [Way94g]. cost/performance [Hed93a]. Costs [Ano93-95, Gla90c, Hol94, Tho92g, Wal94b, Way94e]. Cote [Ga 94c]. Could [Ano94h, Ano94-165, Ano94-35, Ano94-65, Ano94-81, Ano94-135, Cla93a, Fie91d, Gre94a, Lie93, Par90, Ras90f, Rei94f, Sma93b, Ude92d, Way94a]. Counting [Pou90o]. countries [You94]. country [Pou94q, Pra93]. counts [Ano91-184]. Couple [Tho92c, Ano94-79]. Course [Ano90p, Pou94g, Ras90d]. courtesy [Gre91a]. cousins [Dav90]. cover [Ano93-44, She91]. cover-page [Ano93-44]. coverage [Ano94-101, Ano94-117]. covered [Ano94w]. covering [Ano94u, Ano94v]. covers [Hel92b]. covet [Rei93b]. Covox [Cra90d]. CP [Ano92t]. CP/M [Ano92t]. CPU [Alf92c, Ano91-90, Ano91-108, Ano92-97, Ano92-196, Ano93-50, Api94d, Gla91b, Gre94d, Hal93c, Lin90b, Mul94, Pou92d, Red92a, RH93, RT94a, Smi94g, Tho94f, VC90h, Way93f, Yag93b]. CPU-intensive [Mul94]. CPU-upgradable [VC90h]. CPUs [All92c, Ano92-78, Ano92-102, Ano94-116, Bac94, Fox93, Hay90b, Lan94b, Red92a, Sha94c, WK90, Len94a]. Crackdown [KGTD92]. cracking [UFC94]. Cracks [Ano91-137]. craftsmanship [Ste91b]. Cram [VC90c]. Crammed [Woo90a]. Crash [Hol94, Pou94m, NTS92a]. crashes [Cra91b]. crave [Red93]. Crawford [Pou91k]. Create [Ste91d, Ano91-159, Ano92-206, Ano92-207, Ano94g, Ano94x, Ano94-117, Ano94-174, BGE^{+}92, Hed93b, HWS91, Ref94, Ros90a, Vel90, Woo90c]. created [Gre90b, Pet90]. Creates [Ano94-92, Ano91-32]. Creating [Ano92-192, Ano93-116, Mar90b, STS92, Tra90, Ups90, Ano91-177, Ano93f, Ano94-135, Bur92, Eva90, Gib90b, Gre93a, Rob91a, Rya91e]. Creation [Bra90, Ano94-179, Smi94a, Ste91b, Ude93m]. creative [Ano90-112]. credible [Ude94f]. Crimson [Ano92-57]. Crisis [Rya90d, Rya91a]. criteria [Ano91-130]. critic [Cra91j, Ken93c]. critical [Ano92-178, Bac94, Egl94c, Gre93a, VN91a, Yav94]. criticisms [Min90d]. Croatia [Pra93]. crop [Ano92-153, Yag92a]. Cross [Ano92-133, Ga94, Ano90-95, Ano91-32, Ano93-31, Ano94-117, Ano94-174, Ano94-183, Bar92b, Cro93a, Gil94, Gre94e, Hel91d, SE94, UPSW94, Yag93e, Ano92-163, Ano94-38]. cross-application [Gil94]. Cross-Platform [Ga94, Ano92-133, Ano90-95, Ano91-32, Ano93-31, Ano94-117, Ano94-174, Ano94-183, Cro93a, Gre94e, Hel91d, SE94, UPSW94, Yag93e, Ano92-163, Ano94-38]. crossbreeds [Way91b]. Crossing [Cra90g, Nak94]. Crossroads [Min92b, Ano93-55]. crown [Gib92, Rya90d]. CRTs [Ano92-134]. crucial [Ano92-77, Pou94d]. crumble [Ras91o]. crumbled [Bar90g]. Crunch [GE92]. cruncher [Tho90c]. Crunching [Fri91]. Cruz [SY90]. cryptography [GTGW94]. Crystal [Par90, Cra90o]. CSprite [You93]. Cube [BR93, TS90, Way94d]. Cuckoo [Sto93, Ano90-69]. culminated [Bra90]. culmination [Lu92]. Cultural [Nak94, UFC94, SU93]. culture [Pet90]. cumbersome [Ras92c]. Cupertino [Ano94k]. Cure [Ano94-169]. Cures [Min92c]. curiosity [Rob90b]. Current [Woo90b, Ano93s, Ano94-168, Egl94g, Fie91d, Nan93c, VN91f, YG92]. currently [Ano94-135, Van93c]. Curtain [Yag92c, Bar90g]. curve [Ano94-82]. Curves [NTS92b]. Custom [Ano94-117, Mar90b, Ude94c, Ano94-166, Ude94g]. customize [Ano92-28, Coo90a]. Cut [Ano92-148, DE90, Yag92e, Ano94l, Way94e]. Cutting [Tho94c]. Cx486SLCs [Alf92d]. Cyber [Ken93a]. Cyberspace [Bar91f, KGTD92]. cycle [Hed94]. Cyclone [Bar92e]. Cyprus [Ano93-97]. Cyrix [Alf92d, Ano92-36, Fox93, Gre90b, Rya94b].

D [Ano91-64, Ano91-103, Ano92c, Ano92d, Ano92-142, Ano92-149, Ano92-156, Ano92-167, Ano92-212, Ano93-116, Ano93-125, Ano93-135, Ano94-71, Bri90, Cal91a, Cla91a, CWVS94, Cra91i, Gla90h, Lov91b, McA90, Orr94, Ran90, Rei90a, Rei94a, Sac90, Smi94c, Ups90, Way94h, Yag93i]. dabblers [Car94]. daemon [Fie91c]. damage [Gre91c]. dame [Way90c]. Damn [Voo93]. Dan [Vos90a]. Dangers [Ano91-35, Saf93]. Dariana [Ano90-114, UT91]. Dark [Loc91, Sch92c]. darker [Cra90q]. Darkest [Ano90-39]. Darkly [Ras90m]. darkrooms [Lov91a]. Dashboard [Ano92-37]. Data [Ano91-147, Ano92j, Ano92-126, Ano92-212, Ano93u, Ano93-108, Ano94-41, Ano94-42, Ano94-43, Ano94-58, Ano94-64, Ano94-67, Ano94-169, Api91d, AG91, Api93b, Bar91a, Bon90d, Bry94b, Cro93b, Cro93a, Cru90, Dav94a, DWKS91, Egl90d, Egl94b, ESN94, GD93, Gla90c, Gre90a, GD92b, Hal94a, Kor93, LM90, Mar92c, Nan94d, Rei92c, Rei94b, Rib93, Ric94, Rya91b, Smi93b, SE93b, Top91, Tre93, Ude93a, Ull93, VN90d, VN91f, Vog90, Vos90b, Way91d, Way94e, Ano90-103, Ano90-94, Ano91-130, Ano92-41, Ano92-95, Ano92-186, Ano92-220, Ano93m, Ano93-93, Ano93-130, Ano94h, Ano94-103, Ano94-98, Ano94-111, Ano94-174, Api94c, Cla92a, Cra90p, Egl93b, Egl94g, Fis90c, Fri94c, GW93, GTGW94, Hol94, Hur93a, Mul94, Nad94b, Ras90n, Ras90f, Ras90g, Ref94, Rei94c, Rei94d, Sch92b, SG92a, Smi93c, Smi94a, Sta91b, Sta91a]. data [Sta93, Ste91a, Ste94a, Sty91, Ude91a, Ude93h, VN90b, Way93g, Way94c, Way94g, Web93, WK90, Wsz93a, Yav94, Ano94q, Egl90c, Osh90, Tho90a, Ude91a]. data-analysis [Ano92j]. data-highway [Ano94-98]. data-only [Hur93a]. data-throughput [Hur93a]. Database [Ano92-38, Ano92-165, Ano92-187, Ano93-111, Ano93-105, Ano93-120, Ano94-120, Api94c, Ste90, SH90, Ude92a, Ude93a, Ano90-73, Ano91-49, Ano91-67, Ano92-59, Ano92-54, Ano92-137, Ano94w, Ano94-114, Car94, Cla93b, Cra90e, Die93c, Egl93b, GD92b, Gre94b, Hal93f, Kay94c, Mul94, Pou91m, Ras92f, Smi90b, Ude94k, Van93a, VN90c]. database-enabled [Ude93a]. database-independent [Api94c]. Databases [Ano92-59, Cra90e, Fin90a, MS93, Ste94a, VN90c, Ano92-84, Ano92-205, Ano94-101, CWVS94, Gre93a, Har93, Obe90, Pou91f, Ras90c, Ric94, She90b, TS93b]. Datafax [Ano93r]. Datasafe [Pou90c]. Dauphin [New93]. David [Ano92-114]. Day [Fie90g]. DaynaFile [Ano90q]. days [Ano90-88, Ano94z]. Daytona [Ude94i]. Daze [Pou90t]. dazzle [PP93]. Dazzles [Ano92m]. Dazzling [Tho92b, Tho92g]. dBASE [Ano90r, Ude91a, VN90a, Ano92-59, Ano94w, Sch92c]. dBase-compatible [Sch92c]. DBMS [Smi90b]. DBMSes [Ste94a]. DCE [Ano93-41, Mil94b]. DDE [Ano92-130]. DDK [Mas93]. Dead [Ano94a, YS92, Ano91x]. Deal [LM93]. Dealing [Cra91b, Led90, WK90]. Deals [Law94]. Death [Ano94-123, KCU^{+}93]. Debate [KPC93, Mul94, Ude94a, Ano91-132, Ano91-189, Ano91-186, Ano92-62, Ano94-65]. debug [Jef93a]. Debugger [Ano90-109, Pie93, Tho94k, Ude91c]. Debuggers [Kea91b, Ude91c]. Debugging [Kea91b, Ano92-175, Nag92, Pie93]. Debuts [BS91, Ude94l, VN91e]. DEC [Ano92-106, Ano92-176, Ano93-65, Ano94h, Ano94-44, Dej94, Gre93h, Osh90, Osm90, Pou93e, Smi94a, VC90m, VC90t, Yag90e]. decade [DG91, Hal94i, Pou94d]. Decades [Ano90b, Ano94u, Ano94v]. Deceptions [Ano92-40]. decide [Egl90a]. Decision [GB93, Ano91-157, Ras90j]. decision-support [Ras90j]. Decisions [Pou91l, Red92a]. Declared [Ano93b]. decline [Ano92y]. DECpc [Gre93h]. decrease [Way94i]. DECstation [Yag90e]. DECtalk [Ano94-178]. dedicated [Ano90-40, Ano94-143, Bri90, Nan94d, UM90]. Deep [Min90j, Hel90b]. default [Fie91f]. defeating [Ano92-108]. defended [Ano92-91, Ano93-60]. Defending [Ras91d]. defends [Min90d]. deficiencies [Ano93-69]. define [Bry93b]. defined [Sal94]. Defines [Cra91k, Ano92-176]. definition [Lip90b]. degree [IC94]. degrees [RM92]. delayed [Ano94w]. Delegation [Ano94-80]. Deliver [Ano91w, Ano93-90, Ano93-93, Ano94-106, Bar92b, Cra90f, Don92, Egl93d, Gla90b, Har93, Hel92a, Mar90a, RM92, Ref94, Ude93m, VC90a, Vor93]. delivered [Ano94-166]. Delivering [Ban91, Ano92-105, Ano93a]. Delivers [Tho91d, Ude90c, Ude93c, Ano90-62, Ano92-94, Api94c, Dej94, Die90a, Nan93a, Rei93d, Sha94b, Smi94a, Tho90h, Ude93i]. delivery [Ano94-77]. Dell [Ano92-39, Api92b, Ric92, Ung90a]. Delrina [Mia94b]. Delta [Ano91-167]. DeltaGraph [Bar92d]. Deltas [Ano92-203]. Deltasoft [Cra90b]. deluge [Vog90]. Demand [Ano93s, Rei93b]. demanding [EW92]. Demands [Ano92-49, Cla93d, Mel93b, RH93, Ude93l]. demise [Ano94-99]. demonstrate [Ano90-55, Dys92]. demonstrates [Pou94c]. depends [VC90j]. depersonalize [Ano91-35]. deploy [Smi93b]. depth [Ano91-165, GR94]. Depths [Smi93b, Tho91c]. derives [Ano91-187]. descendant [Alf90b]. describes [Bar91b, Ras91b]. Descriptions [Cra90m]. Deserve [Yag92c, Ano94-158, Kea90]. Design [Ano92-174, Bra90, Hed94, Ros90b, Vor93, Wal94a, Ano90-41, Ano92-34, Ano92-94, Ano92-207, Cal91a, Die90a, Gut94, Har91c, Hay90a, Hol91, KYT^{+}92, LB90, LAD91, Mar90b, McA90, Nak94, Pic93, Pou93d, RNG94, Sal94, Ste92, Tho93i, Van94b, VFGZ94, Way94f, Ano91-181, Bis91]. design-quality [Die90a]. DesignCAD [Cal91a]. designed [Ano94-63, Ano94-85, DE91b, Laf91]. Designer [Ros90b, Ano93-44, Ano93-100, Gut94]. designers [Ano91-165, Ano92-49, Ano92-62, Ano92-102, Bis91, Gre94d, Hed94, Sei91, SSF93]. Designing [Ano92-150, Cra92a, Jac91]. Designs [Ano91-90, Bry94d, Fox93, Hal94b, Hol93a, Mar92c]. Desk [Ano92-46, Yag92a, Ano93-105, Gre90j, Woo90b, Yag91e]. desk-accessory-style [Woo90b]. DeskJet [Rei93b]. DeskStation [Gre93d]. Desktop [Alf92d, And93, Ano91g, Ano91-107, Ano91-159, Ano92-68, Ano92-146, Ano92-147, Ano92-148, Ano93w, Ano93-107, Ano94-166, Bry92, Bry93a, Coo90b, Coo90c, Cru90, Dav94a, Die94, Don92, FJPR93, Ga 94b, Len94a, Lin93, Mil93, Mit91a, Nan92a, Pou91k, Rei93e, Rei94a, Tho91b, Tho92d, TR94a, Ude93d, Van93b, Way92a, Woo91a, Wsz93c, Yag91f, Yag92b, YM93, Alf90a, Alf91a, Ano90-102, Ano91w, Ano91-51, Ano91-68, Ano91-109, Ano92u, Ano92-38, Ano92-58, Ano92-123, Ano92-153, Ano93-75, Ano93-81, Ano93-138, Ano94x, Ano94-88, Ano94-147, Ano94-148, Ano94-171, Ano94-175, Bar93d, Cla90a, Cla93a, Cla94b, Cra90n, Egl91c, Egl93c, Fin90b, GR94, Hal94b, Joc90c, Mar90a, Nad90d, Nan93c, Nel91a, Pou92c, Pou94f, Rei94e, TS93a, Tho93f, Ude93j, Ull93, Way94d, WK90, Yag90h, Yag93d, Ano92u, Yag90j]. desktops [Alf92d, Ano93c, Mas91a, Pou92c]. DesktopTV [Ano91-36]. Despite [Ano91z, Ano92-30, Ano94-113, Hal93e, Ras91g]. Desqview [Ano92-213, Gre90d, Min90h]. Desqview/X [Ano92-213]. Destination [GK91]. destroys [Ano92-39]. detail [Ano92-152]. detailed [DB93, Oth92]. Details [Ano90-97, Ano92-174, Fie90f, Sch91a]. Detectives [UT91]. determine [Ano91-130, Cra91e, Gre93b, Ras90i]. determines [Ano92-69, Gas93, Van93c]. determining [Bac94]. detract [Smi94c]. Develop [Hal90, Sum90, Ano91-175, Ano94-113, New93, PS94, Sin94, Url90]. Developed [Pou93e, Ano92-135, Daw94d, Fie90d, McG90b]. Developer [Kro90, Mas93, Min93b, Ste90, Ano91-194, Ano94-69, Jef93b, Ude94l, Ste90]. Developers [Ano92-211, Ano93t, Ano94-45, Ano94-125, Han92a, Ano90-78, Ano92-62, Ano92-61, Ano92-89, Ano93-64, Ano94e, Ano94g, Ano94-132, Car94, Ham90, Hel91d, Kay94d, Kro90, Rei92a, Wig90]. Developing [Chr94, GTN94, Lan94a, Ols94, You94, Ano91-56, Ano92-42, Ano94-152, HH94a, SH91, Tho94k, VC90s]. Development [Ano90-34, Ano91-86, Ano92-163, Dej94, HWS91, Ano91-32, Ano91-187, Ano92-133, Ano92-197, Ano93-52, Ano93-56, Ano94w, Ano94-45, Ano94-98, Ano94-165, Ano94-174, Ano94-180, Car94, Cra90e, Cra90q, Cra91l, Ga94, Ga 92b, GFK94, Ga 94a, Gre94e, Kor93, Lan94a, Lin94, Mas91c, Mas92, Smi94a, Ste90, Tho93c, Ude93i, Ude93m, Wal94a]. developments [Ano90-29, McG90b, Rya91b]. Develops [Lov91a, Api94b]. Device [Ano93-80, Bar92g, HP93, Mas91c, Ano91-47, Ano94-147, Ano94-145, Ano94-174, Ano94-175, Hel90c, Mar90b, Mas90b, NTS93a, Par90, Pou93n]. device-driver [Mas90b, Pou93n]. device-independent [Hel90c]. Devices [Led90, NTS93a, Rei94g, Ano91o, Ano91-172, Ano92-139, Ano94-76, Ano94-77, Ano94-113, Egl90d, Egl92a, FT90, Gla90d, Gre90e, HR90, Kay94a, Lab92b, McA90, Pou90k, Rya93a, SC92]. Dex [Egl93b]. DGX [Ric92]. diagnose [Ano91-33]. Diagnostics [Api91c, Ano90-114]. Dial [ATU92, Daw94a]. Dial-Up [Daw94a]. Dialog [Kiy93b]. DIBs [Hel90c]. dicing [Ref94]. Diconix [Joc91]. Dictation [Ano94-86, Die94]. dictionary [GT93, Ano92-136]. Did [Pou91s, Rei92b]. didn’t [Pou94r]. Diet [Ano91-117, Ano92-129]. differ [Var94a]. Difference [Ano92-178]. differences [Gre90b, May94b]. Different [Ano94-117, Loe90, Mas90a, Pou94b, Ste90, Ano92-216, Ano94k, Ano94-84, Ano94-77, Api94d, Bry94a, Egl94b, Fis91a, Gre90j, Hal94f, Har93, HC93, Pie94, Pou94g, Rya90a, Ste91e, The94, Tho93i, UY91, VC90g]. difficult [Gil94, VN91d]. Difficulty [GD93]. diffusion [Nel91a]. Digging [Min90a]. Digibot [Ano93-133]. DigiDial [Ano94-181]. Digipro [Wsz90]. Digital [Ano92-40, Ano92-41, Ano92-211, Ano94-46, Fri94b, Lio90a, Lip90b, Mur93, Pou91o, Sch93c, Sch93d, Smi94c, Tho90a, Way94c, Way94d, ASS92, Ano91-47, Ano92-68, Ano92-206, Ano93m, Ano93-55, Ano93-134, Ano94-37, Ano94-171, O’M94, Ano90u, Ano91-40, Pou91l, Smi91a, Woo90a, Yag90e]. Digital-audiotape [Lio90a]. digital-image [Ano92-68]. Digital-Media [Smi94c]. Digital-read-channel [Way94c]. digital-video [Ano92-211]. Digitally [ZV92]. Digitizer [WS93, Mia90b]. dilemma [Ano93-64, Chi93, Ude91d]. Dimension [Yag91d]. dimensional [Ano91-64, Cra92a]. dimensions [Ano90-51, Ano91-53, Bri90, CV90b]. Dimmer [Ano94-89]. dinosaurs [Lie93]. DIP [Har91a, Loc91]. Direct [Ric92]. direction [McN94]. directions [Cra91e, Tho93i]. directly [Gre93d, Smi94e]. Director [Tho90b, Ano94-94, Tho90b]. directories [Hol93b]. Directory [Hol93b, Ude93h, Ano91-27, Ryd93]. DirectRoute [Fri94c]. DirectServe [Cra90a]. dired [STS92]. Dirty [Pou93l, Woo91b]. Disabled [Laz90, Laz93, Ano91-191, Sch91b]. disadvantages [Ano93-110]. Disappear [Ano91-131]. disappointing [Ano94-113, Api94e]. disappointments [Ude93i]. disarray [Bar90g]. Disasters [Pou90k]. discontinued [Ref94]. Discover [Ano92-178]. discovered [Ano94-162]. discoveries [Ano90t]. Discport [Bos93]. discs [Alf93]. Disctec [Mia90e]. discuss [Ano90-112, Ano92-61]. discussed [Ano92-108]. discusses [Ano91k, Gut94, Hay90a, Loc91, McG90b, Pou92f]. discussion [Ano94-166, Ano94-182, EB90a]. Disk [Alf91f, Alf92e, Ano91-63, Ano91-139, Ano92-129, Ano92-126, Mia90d, Mia90e, Nan91a, Pou90t, Rei91d, VN91b, Wsz90, Ano90-94, Ano91-156, Ano92-49, Ano92-141, Ano92-160, Ano93-132, AWGY90, AG91, BL90, Gre90i, LM90, Led90, Min90b, Min90i, NTS92a, Pou90n, Pou93i, Rai90, RP90, VN90b, VN90d, Way94c, WET90]. Disk-Array [Ano92-126, Ano93-132]. Disk-based [Nan91a]. diskless [Bar90c]. DiskPaper [Ano91-37]. Disks [Ano91-112, Ano90q, Ano92-28, Bar90f, Ras90a, VC90r]. disparate [Ude91a, VC90q]. Display [Ano92-134, Ano94c, Egl91e, Lov92a, Ano91-131, Ano92-33, Ano92c, Ano92-61, Ano92-199, Ano93o, Ano93-126, Api94d, Bar91b, Lov90a, New93, Tho91c, Ude90e]. display-list [Ude90e]. Displays [Ano92c, Ano92-42, Ano92-164, Ano94-135, Smi90c, Tho90a, Ano91-97, Ano92-115, Ano93-133, Ano94-35, Cov90, Mac91, Sma93b]. dissected [Min93b]. Dissenter [Min90d]. distance [Ano94-182, Ano92-177]. distill [Ste91a]. distinct [Ano93-110]. distinguishes [Gre90i]. Distributed [Ano91-38, Ano91-67, Ano92-186, Ano93-42, Ano94-47, Ano94-49, Kay94b, MH91, Mil94b, San90b, Sey93, Ano92-127, Ano92-190, Ano94h, Ano94-45, Dej94, Pou94a, Rei94d, Ric94, Ryd93, Smi94b, Wat94, Yav94, Ano94-48, Osh90]. Distributed-Computing [Ano94-48]. distributed-object [Ano94h]. distributed-objects [Dej94]. Distributing [Ric94]. distribution [Die93a]. divergent [DH92, LAN92]. diverse [Ano92-186, ES93, GD93, Rei94b, Ude90a]. diversions [Sel92]. Diversity [Rya90a]. Divide [Ano94-118, Jef93a, WS91]. dividends [Ano92-104]. divulged [Ano94n]. DLL [Nag92]. DLL-based [Nag92]. DLLs [Api94c, Key92]. DMI [Ano94-134]. Do [Ano90s, Ano90t, Ano91-39, Ano91-77, Ano92-89, Ano94-87, Cra90f, Fie90i, KPC93, Pet90, She93b, Ano91s, Ano91-63, Ano91-123, Ano91-169, Ano92-182, Ano92-205, Ano92-218, Ano93-72, Cra90h, Egl94c, Fie91h, Gla90f, HH94b, Law94, Lip90a, Min92d, Nan91l, Osm90, Ras91a, Ras91m, Rei91c, Rei94d, Rya90f, Sha94c, Smi90b, Ste91d, SH90, Wsz93b, Ano94-89, Ras90f]. Do-It-Yourself [Ano94-89, Ras90f]. doable [Min91c]. DOC.IT [Ano92-43]. docking [CW92]. Doctor [Min92c, WET90]. Doctors [Bar92c]. Document [Ano91-138, Ano93u, Ano94-49, Ano94-189, CS92, Day90, Har91a, Lu92, Rei94e, Ano92-180, Ano93d, Ano94-179, Ano94-183, Cro93a, HR91, Mic91, Pie94, War92, Yag92f, Zei93]. document-interchange [War92]. Document-oriented [Lu92]. document-retrieval [Ano92-180]. Documentation [Mur93]. Documents [Ano94-78, Hel91b, Mic91, Sch93b, War92, Ano92-150, Ano92-181, Ano93a, Ano93-133, Api92a, Cro93a, Day90, Die93a, Joc91, Loc91, Ras91c, Rei94e, Sch93d, Smi94a, Wsz93a]. Does [Ano90t, Ano91-153, Ano93-139, Ano94i, Die90a, Ga 92a, Gre93b, Tho93f, Ude93f, Way94k, Yag93g, Ano92-224, Ano93-88, Gri94, Ras91j, Ude91c]. Doesn’t [Ano91-189, Ano91-185, Ano92-201, Ano92-199, Eva94, Ken93c, Pou93g, Ras91f, Rya90a, Smi94c, VC90r]. Doing [Gil94, Ano90m, MH91, Saf93]. Dolch [Api91a, VC90i]. Dollars [Tho90h]. Domain [Sum90, Chr94, Fri90]. dominate [Hal94b, Pou93s]. domination [Pou93j]. Don [Cra90f, Cra90o, Cra90r, Cra91a, Cra91b, Cra91e, Cra91c, Cra91i, Cra92b, Lan90b]. done [Cra90m, McG90a]. Don’t [Ano94-50, Dau91, Fic90, Mos93, Pou94n, Alf91d, Ano92-149, Ano94-79, Bry92, Cau91, Cla93e, Egl94c, Hal93e, HH94b, Kor93, VC90k, VC90q, YG92]. Door [SL92]. Doors [Laz90]. DOS-based [Ude94e]. DOS-to-Unix [UY91]. dot [HM94, KM94]. dot-matrix [HM94, KM94]. Double [Nan92a, Pou90n, Rei92a, Yag90m, Ano91-156, Ano94-87, Hal94a, Nad90d, Way94c, Ano92-78, Ano94-87]. Double-Fast [Ano92-78]. double-speed [Ano94-87, Ano94-87]. doubled [Ano94-150]. Doublers [Alf92a]. doubles [Alf92a, Ano92-59, Hal93c]. DoubleSpace [SSF93]. Down [And93, Ano94-68, Fie90l, Fie91e, Ham92, Kli90a, Ras90j, Ude93i, Ano92t, Ano92-184, Dau91, Gla90c, PTUM93, Pou91s, Pou92f, Ran93]. Downloads [Ano91-79]. Downsizing [Ano92-45, Rya91c]. downward [Gla91c]. downward-compatible [Gla91c]. dozen [Ano92-27]. dpi [Van93b]. DPMI [Gla90e]. DR [Ano90u, Ano91-40, Woo90a]. Dr. [Ano93-107]. Draft [HM94, HHP93, KM94]. Dramatic [Nan91c, Pou94o]. dramatically [Mar92a, Rei94f, The94, Zei93]. drastically [HC93]. draw [Ano92-149, Ros90b, Sac90]. drawbacks [Kay94c]. Drawing [Cah90, Ano91i, Ano93-43, Ano94-68, Cra92a, Pou93a]. drawings [Woo91a]. DrawPerfect [Ano90v]. Draws [Wau93a, Way93c]. Dream [Smi90f]. Dreams [Ano91-41, Cla93a]. Dredging [Pou91m]. Dressing [Wsz93c]. Drive [Alf91f, Ano94-31, Egl93e, Hal93a, Mia90d, Mia90e, Alf92e, Ano90i, Ano91u, Ano92-37, Ano92-98, Ano92-126, Ano92-199, AG91, BL90, Pou90n, Pou91p, Pou94j, Rai90, Sat90, Way94c]. driven [Ano94-172]. Driver [Tho93d, Fis90b, Mas90b, Mas91c, Pou93n]. Drivers [Ano94-90, HP93, HHP93, Mas91c, Api94c, Bar92g, NTS93a, RF91, Tho93d]. Drives [Ano92-45, Ano93k, Ano93l, Ano93j, Ano93-47, Ano94-87, Ess94, Har91b, Rya93b, UE93, Way94c, Alf93, Ano90-57, Ano91-63, Ano92-49, Ano92-169, Ano93i, Ano94-129, AWGY90, CP94, Egl93f, Gre90i, LM90, Led90, Min90b, Tho94c, Tho94d]. Drivin [Gre90i]. driving [Ano92-202, Gla90c]. Drop [Ano94-31, Hal93e]. drops [Egl93b]. drove [Gut94]. Drowning [Vog90]. drudgery [Ano92-28]. Drum [Ano93-49, Van94a]. DSP [Ano92-183, Ano93-27, Bry92, Dav94a, Dav94b, Tho94f, Way92b]. DSP-Aware [Dav94b]. DSP-based [Dav94a]. DSPs [Ano94-51, Ano94-54]. DTP [Ano93-138, Cra90n, HM94, HHP93, KM94, PP90]. DTR [Ano93-130]. DTR-1 [Ano93-130]. Dual [Ano94-52, ET93, Fin90b, Len94a, Ano91l, Ano93-134, Ga 94d]. Dual-Page [Fin90b]. dual-Pentium [Ano93-134]. dual-processor [Ga 94d]. Dubious [Ano93-48, Ano94-90, Ano94-87, Ano94-88, Fox93, HM94, HHP93, HP94, Len94b, Len94a]. duds [She90a]. Due [Ano94-158, Egl94c]. Duel [Fin90b]. Dueling [Alf92a, Fis91a]. duke [Van93c]. dulled [Egl93a]. Duo [Tho93i, VN91e]. Duos [Ano93c]. during [Ano90c, Wsz91]. Dust [Ano94-107]. duty [Ano92l]. DVI [LK90]. DX2 [Ano92-65, Ano92-111, Api92b]. DX2s [Alf92a]. Dye [HHP93, Ano93-131, Nel91a, Tho92g]. dye-diffusion [Nel91a]. dye-sublimation [Ano93-131]. Dyes [Tho92g]. Dyna [San90a]. Dynabook [Ano91-42]. Dynamac [Loe90]. Dynamic [MH91, Sch93b, Vos90b, Ano93f, Ano93-76, Pic93, Ref94]. Dynamic-Viewing [Ref94]. dynamics [Nie94]. Dynamite [FJPR93, Ano92-54, Yag93b]. Dynapad [Sma93a]. DynaZIP [Ano94f].

E-Mail [Ano91-43, Ano93v, Ano94-30, Ano94-53, ES94, Fie91b, Rei93d, Ude93e, VN90e, Ano92-120, Ano93-58, Ano94-98, Ano94-145, ET91, Ga 93b, Sta94a]. Each [Cla94b, Ano91-184, Ano94-76, Api94b, Egl94c, Gas93, Gre91a, Gre93b, HC93, MW94, Rei94b, Slo91, Ude93l]. Early [Sha94b, Ano90-60, Ano91-137, Ano92-110, Ano94z, Bux90a]. earn [Ano91-170]. earns [Ano92-119]. earthbound [Dav90]. earthquake [Pou94m]. Ease [Ano94-63, Ano94-154, THN93, Api93c, Car94, Die94, Egl94e, Gre93a, Gre94c, Lan94a, Nan91c, Nan93d, Yag90h]. ease-of-use [Car94, Nan93d]. eases [Ano90-76, Ano91-106, Ano92-130, Ano92-195]. Easier [Ano93n, Cla91b, Cla91a, Eug90, Kea91b, Nan94b, Ano90r, Ano90-103, Ano90-116, Ano91-38, Ano92-95, Ano94-63, BGE^{+}92, Bry94d, Cas92, Cra91l, ES94, HWS91, Laf91, Mel93b, Min90f, Rya91b, Ude90a, VC90n]. easier-to-use [Mel93b]. easiest [Ano93-72]. easily [Ano90i, Ano92-206, Chr94, Gil94, PS94, Ras90k, Ude93c]. Easing [Lov94, VC90c]. Eastern [Bar90g]. Easy [Ano94e, Cra90b, LB91, NTS92a, Ras90e, Smi94b, Ude93f, Ano90-34, Ano90-42, Ano90-74, Ano91-175, Ano92-191, Ano92-214, Ano93h, Ano94-68, Ano94-114, Bar91e, Bos93, Bry93a, Coo90a, Cro93b, Die90e, Egl90c, Egl94e, Ga 94b, Hel91c, Lus90, Mul94, Nan94b, Pou91f, Tho93c, Yag92c, Yag93a]. easy-to-install [Nan94b]. easy-to-use [Ano94-114]. eavesdrop [Way93c]. Ebert [Cra91j]. Eccentric [Mia90b]. echo [Sto93]. ecological [RPRM92]. Economical [Len94a, Lov90a]. Ed [Ano92y, You90]. Eden [Pou91e]. Edge [Ano91-70, Ben91, Ber93, KRCT93, Ude93g, Ano91-102, Bar90j]. Edges [Egl91e, Ano93i]. EDI [Way94e]. Edit [Ano92-46]. edited [Ano92-148]. Editing [Ano91-32, Lov91a, Ano94-40, Egl94f, Lin93, Lin94, NTS93a, Sch93b, Yag93h, Ano94-40]. Edition [Ano90-111, Sha94b, Egl94c, Ude91a]. Editor [Ano91-62, Ano90-32, Ano91k, Ano92-46, Egl94b, NTS93c]. Editorial [All92d, All92a, All92b, All92c, All93f, All93g, All93h, All93i, All93j, All93k, All93e, All93a, All93b, All93c, All93d, All94a, All94b, All94c, All94d, All94e, All94f, All94g, All94h, All94i, All94j, All94k, Ano92-47, Ano93w, Ano93-59, Egl94e]. Editors [Ano91-176, Ano94q, Ano91-186, Ano92a, Ano92-90, Ano94p, Joc92b, Nad93, Smi93b, YG92]. Edsels [She90a]. Edsun [BS91].

Education [Fis91b, Cra90q, Pou94k]. Educational [Pou94k, Ano92-30, Pou90t]. effect [Ano94-65, Ano94-98]. effective [Ano94-149, Cla94a, Cov90, Sha94b, Smi94a, Ude93i, VN91e, WS91]. effectively [Way94g]. effectiveness [Ano94-149]. effects [Ano92-156]. efficiency [Hur93a, Rya91c]. efficient [Ano92-204, Mul94, Way93f]. Efforts [Ano94-154, Ano94s]. Egg [Ano90-69, Sto93]. Eglowstein [Egl94b, Egl94g]. Eight [Ano91b, Car92, Egl91d, SG92b, Alf92a, Ano92-189, AG91, Egl91a, Ess94, Fin90a, HHP93, Mit91c, TG91, Wsz93a]. EIS [Wat94]. EISA [Alf92b, Ano90-61, Ano92-198, Ano94-90, AD91, Fox93, Ga 93e, Gre91b, HP94, VC90a, WK90, Yag90d]. EISA-Bus [Ano94-90]. EISes [Cro93b]. either [Way94d]. elbows [Gib92]. ELC [Ano91-159]. Electric [Ano93-131]. electrical [Ano92-222]. electromagnetic [Saf93]. electromagnetic-field [Saf93]. electron [Smi91b]. electron-beam [Smi91b]. Electronic [Ano92-48, Ano92-51, Ken93c, Lin93, Lu93b, SE93a, Ano93a, Ano94z, Ano94-189, Buc92, Die93a, Egl93d, Ga 93d, Lov91a, Pou94c, Rob91b, Sch93d, UPSW94, Way93c, WB90, Bar91f]. electronic-document [Ano94-189]. Electronically [Lu93b, Mur93]. Electronics [Ano91-156]. Electrons [Smi91b]. Elegant [Kiy93b, Ga 93b]. element [Loc91, VN91a]. eliminate [Par90]. eliminates [Ano90u, Ano94-57]. Eliminating [Way94e]. elusive [Ano90-86, Yah92]. emacs [STS92]. embarking [Ano94m]. Embarrassment [Min91b]. Embedded [Ano92-49, Coo91, GTN94, Sch93b]. Embrace [Die90e]. embraces [VC90m]. Emerald [Mit90a]. emerge [Api90b, Hay90b]. emerges [Ano92-153, VN91c]. Emerging [Lie93, Ano91w]. eminent [Ken93c]. emission [Ano94-35]. Emissions [Ano94-89, McG90a]. emphasizes [Ano94-117]. Empire [Pou91s]. empowering [BR93]. empowers [Egl94g]. Emulating [Api94e]. Emulation [Hal94f, Joc90b]. enable [Api94c, Daw94a, Gre93a, MPW^{+}90, Sch93d, Ano94-184]. enabled [Ude93a]. enabling [Ano92-127]. Encarta [Ano93x, ESN94]. encode [Way94d]. encompasses [Ano92-100]. encounter [CV90a]. Encryption [Ano94-73, Way93c, Way93g]. encrypts [Ano93-130]. encyclopedia [Ano93x, KPC93, UFC94]. End [Alf92b, Ano91-44, Ano93-124, Cra90k, Cra90j, GAM90, Lu92, Per93b, Pou92f, Woo90c, Ano90s, Ano91l, Ano91-181, Ano92d, Ano92-59, Ano92-110, Ano94-32, Ano94-71, Ano94-86, Ano94-131, Ano94-127, Ano94-147, Car94, Del93a, Egl94c, Fox93, Fri90, GW92, Gut93, Hal94c, HH94a, HP94, KM94, Nan93b, Way94k, Yag90d, Yag90m, Ano94-107, Lov91d, LS91a]. end-user [Car94]. endorsement [Ano93-85]. Ends [Ano94-63, Gre93a, HH94a, Nan93b, Pou94o, Obe90, Rya90e, Zei93]. enduring [Ano91-42]. Energizes [Yag92d]. Energy [Ano94-89, Fox93, Ste94b]. enforcers [Way93g]. Engages [Ano91-50]. engine [Ano90k, Api94c, Ga 93b, Hal93f, Mit91a, Orr94, Oth92, Kli90b]. engineer [Cah90, Ga 94a, Gru90, YG92]. Engineering [EB90b, EB90a, Wal94a, Ano90-36, Ste91b]. Engineers [Smi93b, Ano92-161, Ols94]. Engines [Ano90-31, Ano91-50, Ano91-103, Ano94-54, Mar90a]. England [Pou94k]. English [Ano92-136, Swa93]. Enhanced [Bry94b, Hel92b, HHP93, Ano92-221, Ano94-129, Bry94a, Car94, Gre94e, Len94a]. enhancement [Ano92-143, LB91]. enhancements [Egl94a, Mit91a]. enhancer [Ano91-136]. enhancers [Pou92c]. Enhancing [Ano92-50]. enjoying [Ano94-96]. enjoys [Pou94f]. enliven [Cry94]. Enough [San90a, Ano94-82, Ras91f, She91, Woo90d]. Ensemble [Lin91b]. ensure [Ano92-178, Kre91]. ensuring [Way91d]. Enter [Mit91b]. entered [Ano93-105]. Entering [Rya90e]. Enterprise [Ano93-89, Ano94-84, GT93, Ras91k, SE94, Ano94-37, Ano94-92, Ano94-147, Lu93b, Ude93h, Way94g]. enterprise-wide [Ano94-147, Lu93b]. Enters [Ano90-95, Ano92-176, Ano94o, Ano94-33, Ano94-189, Gib92, Ano92-225, Lov92b]. Entertaining [CWVS94]. enthusiasm [Vor92]. entire [Hed94, Ras91b]. entree [Die90e]. Entrepreneurial [GT93]. entry [Ano92-130, Die94]. Envelope [Ano94-82]. Environment [Ano91-32, Jac91, Lin91b, Mil94b, Ano90-111, Ano92-123, Ano94-66, Ano94-95, Ano94-174, Api94b, Bor93, Car94, Cou92, Cra91f, Dan90, Dej94, Die94, Egl90a, Ga94, Gre94e, HH94a, Hog90, Nan94a, Rei94d, Smi93e, Ste90, Tho90c, Tho93c, Ude90a, Ude92b, Url90, Wat94, Way92a]. environmental [Ano94-103]. Environments [Ano91-38, Ano91-67, Ano92-203, Ano92-207, Egl94b, Ga 94c, Laf91]. Envoy [Ano94-113, Ano94-189]. EO [Ano93-88]. EPLDs [Mar90b]. Epson [Row93, Ano93y, Ano93-113]. equal [Gre90b, Hal93d, Smi94e]. equals [Ano92e]. equation [Way90b]. equipment [Ano90-84, Yag90e]. era [Ano92-133, Ano92-209, TS93a]. Ergo [Ano90w]. Ergonomic [Har91c]. Eriang [Ano94-55]. Error [Ano91-43]. error-code [Ano91-43]. Errors [Cra91b, Ano92-34, Egl94g]. escape [Ebe93]. ESDI [Alf90b]. especially [Ano92w, Pou94q, TR94a]. espionage [Ano90-69]. essays [Ano91-179]. essence [Ano94-127]. Essential [Ano93q, Len94a, NW92, Ude94b, Len93a, Loc91, Tho94k].

Establish [War92]. established [Hal93b, McN94]. Ether [Way93b, Way93h]. Ethernet [Ano90-44, Ano92-177, Ano92-199, Ano93-48, Ano94-56, Ano94-57, Ano94-88, Ano94-119, Ano94-158, Ano94-180, Ano94-177, Bry93b, Daw94a, Die90e, Egl91b, Fie92a, Gla91c, Gre91b, Nan94b, Nan94d, Ras92a, Sei91, Smi94e, Smi94f, Tho94b]. ethical [Cos94]. Ethics [Ano92-51, CWVS94]. Europe [Ano94-58, Ano94-60, Bar90g, Red91b, Vos92]. European [Ano93-98]. evaluate [Ano94-89, Fox93, Hol94, HP94]. evaluated [Ano91b, Ano94-119, Egl94e]. Evaluates [GW92, Ano94-147, Ano94-148, Ano94-149, AG91, Api94b, Car94, Cry94, DE92, EWT90, Egl94g, HH94a, Lin94, Ref94, SE94]. evaluating [Ano92-117, SB94]. Even [Ano92-52, Ano93z, Ano94-57, Ano92-154, Api94e, Cau91, Cra91c, Gre94b, Mal93, Pou94d, Van93a, Van92b, Way94h, Yag91e]. Event [Gre90c, Nag93]. Events [Nag93, Tho94i, Ano90c, HC93]. eventually [Ano93-37, Ano94m]. Ever [Mal93, Bry93a, Der92, Pou93c, Ude94j, Way93f]. Everex [Ano90-99]. every [Ano90-88, Ano91-47, Ken92, Ude93l]. everyday [UPSW94, VC90s]. Everyone [Ano93-120, Cra90m, Ude94h, Wsz93c]. Everything [Ano91-195, NTS93b, Rei90c, RH93]. Everywhere [Ude93l]. Evicts [Nan92d]. evidenced [SS94]. evident [Saf93]. Evil [Pou91s]. Evolution [Alf90b, Alf91b, Ano94-59, Web92, WK90, Ano90e, Cra91e, Gre93d]. evolutionary [Tho94j]. evolved [Ano94-92]. Evolves [Van90, Gla91b]. evolving [Mel93a, Sin94]. examination [Woo91b]. examine [Ano91-41, Min90a, Pou93f, SSF93]. examines [Ano90-100, Ano91-179, AMW90, Egl93c, Ga 93d, Gre93a, Mit91c, Per94a, Pou90q, Pou90p, Wsz93c]. exceed [Ess94]. Excel [Ano92-55, Ano93z, Del93b, Gas93, Gib92, Ref94, Rei91c]. excellent [Per94b, Pou93n]. excels [Ano92-132, Ano93-110]. Exception [Ano91-163, NHW93, Kiy93a]. exception-handling [Kiy93a]. Exchange [VN91f, Vos90b, Kor93]. excitement [Ken93c]. exciting [Ano92w, Ga94, Sch93b, Sta91b]. exclusive [Fri90, Len94a]. Exec [Ano92u, Hol91]. executables [Gre94e]. Executive [Ano91-157, Ano92e, Cro93b]. exemplified [Ude94g]. Exercise [Pou91n, Kay94a]. exhibit [Ano90-60]. exhibits [Cla93c]. exist [Pet91]. existed [Ano94z]. existing [DM90, GP90, Hal93b, Smi94b]. Exorcising [Min92d]. exotic [Man90]. expand [CV90a, Nel91a, RT94b]. expandability [Ano91-183]. Expanding [Hol90, Hal94i]. Expands [Ano93-93, Ano90-52, Lus90, Way93h]. expansion [Egl93c, RH93, Tho94b]. expect [Ano94-79, Daw94d, Gre93d, Hal93e]. Expectations [Ano91-55]. expected [Ano92-110]. expense [Ano94-57]. Expensive [AT92, Ano94-71, Ano94-107, Fri94b, Mul94, Pou93i]. Experience [Ga 93b, Wal94e, Ano94-175, Tho93a]. experiences [Nan92b]. experiencing [Ano91-81]. experiment [Sum90]. Experimenting [Ude94d]. experiments [Ano91-102, Dys92, Pou94p]. Expert [Cau91, Eva90, Hal90, Laf91, Sum90, Wal94d, Way93i, Ano91-141, EB90b, Lie93, Ras90p]. Expert-system [Wal94d, Hal90, Ras90p]. Experts [Ras91o, Wal94d, Ano93-138]. ExperVision [Bar92f]. Explained [Alf92e, Par92, Ano92p, Ano92-150]. explainer [KYT^{+}92]. explains [Ano91-98, Gre94b, Sta91b]. explanation [Api91d, Gre90a]. explicit [Api94c]. exploit [Ano92-80, Ano94-29, Cot93b, Kog94, Pie93]. exploration [Gre93h, Pou94l]. explore [Ano91-37, Ano91-105]. Exploring [Alf93, Ude93b, Ude94c, Ude94i, Ano92-161, Ano93x]. exposes [Ano92y, Cah90]. Exposing [Sha94a]. exposition [Pou93m]. exposure [Ref94]. Express [Ano94-178]. exquisite [Ano92-152]. Extend [Alf91c, Ano91-129]. Extended [Min93a, ST93, Ude91a, Gla91a, PS94, Ano91-48, Ric91]. extender [Nan90a]. Extenders [Ano92-44, Gla90b]. Extending [Ple94, Pou93c, Ano92-135, Ano92-160]. extends [Ano92-145, Ano92-177, Ano93-129]. Extensible [PS94]. extension [Kli90c, Lov91c]. Extensions [ST93, Yag93f, Pou93c, Yag93d, Jef93b, Yag92j]. external [Ano93-125]. Extra [Yag92c, Law94]. Extras [Rei91c]. Extremes [Smi93c]. Eye [Wsz93a, Yag91b, Ano90-82]. EZ [Smi94b]. EZ-RPC [Smi94b].

F [Ano92-92, Buc92, Ano90-32, Ano92-131]. F/X [Ano90-32, Ano92-131]. F340iW [Ano93-30]. fabrication [Ano94-135]. Face [Ano92-188, Ano93-43, Ano94-183, Cla94a, Gre94e, HHP93, Wau92, Yag93k, Ano90h, Ano91-49, Ano94-61, CV90b, Egl91e, Egl94f, Hed94, NTS92a, Ano94-61, Ano94-131, GR94, Gre90b]. Face-Life [Ano94-131]. Face-Lift [Wau92, Yag93k, Ano91-49, Ano94-61, Egl94f, Ano94-61]. Face-Off [Gre90b]. face-saving [NTS92a]. Face-to-Face [GR94]. Faces [Var94c, Way93i]. facilities [Gla90d]. Facility [Ano91-56, Vos90b]. Facing [Ano91-45]. Factor [Ano92-42]. factors [Ano92-105, Ano94-162, Len94a, Nak94]. factory [Cot93a]. Facts [Ano91-47, Cra90n, Man90, Mia94a]. Fail [Ano92-56]. Fail-Safe [Ano92-56]. failed [Ude94g]. failing [Cra90k]. fails [Ude93m]. failure [Ano92-126]. failures [Ano92r]. Fainthearted [VN91d]. fair [TS90]. fairly [Way90a]. Fairy [Ano90x]. Fairy-Tale [Ano90x]. faithful [Ude90c]. Falcon030 [Ano93-27]. fall [Ano94s, Ano94-31, Ano94-77, Ano94-106, Cla93a, Ano94p]. Falling [Ano94-60]. Falls [Ano91c]. familiar [Ano92-36]. Family [Ras90g, Ude90g, Ano92-36, RS90, TR94b, Var94c]. fancy [Min91c]. Fantasies [Pic93]. fantastic [RNG94]. Far [Ano92-190, Ano91x, Ano94-37, ES94, Nan91j]. Far-Flung [Ano92-190, Ano94-37, ES94]. Farallon [Ano90-55, Ano92-100, Ano94-189]. Faraway [Api91b, VC90k]. Farewell [Rya90f]. Fargo [Ano93-128, Tho93b]. fascinating [Ano91-102, Bar91b]. FAServer [Daw94b]. Fast [Ano92-53, Ano93-28, Ano94-79, Ano94-90, Ano94-87, Ano94-171, BL90, Cra90b, Cra90c, Ess94, FJPR93, GN93, Gre90e, LB91, Lin93, Mar92a, Nad90c, PGRA92, Rya91e, VC90a, Ano92-116, Ano92-111, Ano94-112, Bry94c, Cal91a, Daw94b, Fie91a, Gla91c, KM94, LM90, Lov94, Mar90b, MG90, Tho94b, TR94b, Woo90d, Ano92-78]. Faster [Ano93g, BL90, Cla90a, Ham91, Hol94, Jof93, Nad90c, San90a, Sta91b, Way93f, Ano91-117, Ano92-173, Ano94-107, Ano94-87, Cla93d, Del93a, Egl90b, Hay90b, Hay94b, Hur93a, Lan94b, Mel93b, Min90b, NTS93b, Pou93c, Sch93a, Tho94f]. Fastest [Alf91e, Api92b, VC90i, Alf90c, Ano91-60, Ano92-53, AD91, Api94e, Hal93c, LB91, Pou94b]. Father [Ano90y, Wsz90, Ano91-85]. Fatware [All93g, PTUM93, Ano93-53]. Fault [Ano91-46, Dau91, Nel91b, VN91b, Ano92-56, Ano93-132, Rie91, Ano91-139]. Fault-tolerant [Nel91b, Ano93-132, Ano91-139]. favorite [Ano93-94, Ano94-100, Pou90m, Smi90d]. favorites [Api90b]. Fax [Ano90-71, Ano91-47, Ano91-99, Ano93-44, Ano94-61, Egl94e, GAW93, Mar92b, Mia93b, Pou90k, RM92, Rei94g, Wsz93a, Ano91-48, Ano92-43, Ano92-64, Ano92-217, Ano93r, Ano94-145, Bar90f, Die93b, Hol94, Mia94a, Mia94b, Ras91a, Ano91-48, Mar92c]. Fax-O-Matic [Ano91-48]. Fax/Data [Mar92c]. Fax96 [Ano90z]. FaxConnection [Ano91-48]. faxed [Wsz93a]. faxes [Ano93-134, Ano94-175, New93]. faxing [Ano90z]. FaxModem [Ano93-130]. Faxxer [New93]. FD [Mia90f]. FDDI [Fis90c, Sta93]. Fearful [Kea90]. fearing [Ano94-57]. feats [Ano90-27]. Feature [Fie91b, Ano92i, Ano93-109, Hal94f, Rei94a, Vor93]. feature-packed [Ano93-109]. Feature-rich [Fie91b]. featured [Ano94-68, Ano94-92, Ano94-114, Nan93a]. Features [Ano93-43, Len94a, Ano90-30, Ano90-72, Ano90-87, Ano90-95, Ano91f, Ano91e, Ano91w, Ano91b, Ano91-151, Ano91-166, Ano92m, Ano92-154, Ano92-178, Ano93z, Ano93-141, Ano94w, Ano94a, Ano94-132, Ano94-148, Api94c, Api94d, Car94, Daw94d, Daw94c, Don92, Egl94f, Ga 93b, Gre94b, Hal93b, Han92a, Mat94b, MS93, Nad90e, Nan93d, Ras90b, Rya93b, Sch93b, Sma92, Tho90f, TS90, TS93a, Ude93h, Ude93i, Van93a, Van94b]. Federal [Fis90a]. Feds [Rei94c]. feed [WB90]. feel [Ano90-41, Ano91-172]. Fenasoft [Ban92, Pou93m]. Ferment [Pou92e]. Ferrari [Ano90k]. Fetch [Die93c, Die93c]. feud [Ano91-132]. Fever [Ano94-69, Pou91k]. few [Ano94-95, Cra90h, FH93, Ken90, Vor92]. Feynman [KYT^{+}92]. Fi2re [Ano92-60]. Fiber [Fis90c, Ano93-93, Bry94c, Min90j, Rei94b]. fiber-optic [Ano93-93, Fis90c, Min90j, Rei94b]. Fibre [Bry94c]. Field [Wal94d, Ano91-54, Ano91-154, Ano94-35, Gas93, Kay94c, Saf93]. fields [Ano90-58, McG90a]. Fifteen [Pou90o]. fifth [Ano92-29, Hal93d]. Fifties [Ano92-53]. Fighting [Ano93-53, PTUM93]. Fights [Egl94f]. figures [Hed93a]. File [Ano92-188, CV90b, Daw94b, DH92, Egl91e, Gre90f, Min90a, Min90b, Per93a, PGRA92, Ras91a, Ano90-73, Ano90-74, Ano91p, Ano91-74, Ano91-76, Ano92q, Ano92r, Ano92-141, Ano92-173, Ano92-175, Ano92-199, Ano94-61, Ano94-69, Bar90k, Cra90a, Mel93a, Min90i, NTS93c, Nan94d, Pou91m, STA92, STS92, Sty91, Tho90i, Ude93h, Ude94k, Way91a, Ano90-73, Ano92-184]. file-compression [Ano92r]. file-sharing [Bar90k]. file-tree [Ano91-74]. FileMaker [Ano91-49, Ano92-54]. files [Ano90-45, Ano92-28, Ano94f, GD92a, VC90r]. Filing [Ano91m]. Fill [UY91]. Filling [Ano90-57, Mia90c]. fills [Ano93-43, Yag92h]. Film [Cra91j]. films [Rei94a]. Filters [Ano91-61]. Finally [Coo90a, Mas92, Ano91-52, Ano91-175, Ano92h, Ano92-159, Ano92-194, Ano92-226, Cra91g, Die90b, Egl91e, Way90c]. finance [Ano92-154, Kof93]. finances [She93b]. Find [Pou90m, RARH92, Rib93, Ano91a, Ano92-95, Ano93-47, Ano94-88, Cra91f, Cro93b, Daw94d, Fox93, HP94, Len94a, Way91b, Zei93]. Finder [NTS93b]. Finding [Ano93-28, Dau91, Ano91z, Sta94a, Web94]. finds [Cra91i, Ga 94c, Gre94c, Lab92b, Pou93i, Pou94k, Pou94q, Ras92f, Smi94c, Tho94i, Ude93i, Van93a]. Fine [Gre93g, IC94, Lov93a, Pou94c]. Fine-Tune [IC94]. Fingers [Ano90t, Fie90i]. Fingertips [Ano92-75]. finite [Ano93-130]. Fire [Way93c, Cla94b]. Fires [Tho92b, Ano92z]. FireWire [Ano94-72]. Fireworks [Tho90e]. First [Alf92a, Ano91-131, Ano92-55, Ano92-56, Ano92-57, Ano93i, Ano93-29, Ano94-27, Ano94-72, Ano94-113, App91, Bar92a, Fie92a, Har91c, Min90b, Per92, Per94b, Pou92a, Ras91n, Sha94a, Sma93a, TH92, Tho92d, TH93, Ude93f, Wau92, Wau93a, Alf92d, Ano90-80, Ano91-108, Ano91-125, Ano91-197, Ano92-36, Ano92-104, Ano92-111, Ano93b, Ano93-68, Ano93-66, Ano93-88, Ano94-35, Ano94-45, Ano94-126, Ano94-86, Ano94-127, Daw94c, EB90b, GN93, Gib90b, Gre90d, GD92b, Gre93d, Hal93d, Pou91g, Pou93a, Ras91m, Sei91, TR94a, Ude93c, Ude93m, VC90i, Way94g, Woo90d, Yag90a, Yag92a, Ano94-62]. First-Party [Ano94-62]. first-release [Ude93c]. first-time [Ras91n]. firstborn [Ano91-58]. FirstClass [Ga 93b]. firsthand [Cra92a]. Fistful [Tho90h]. Fit [GW93, Mit91a, HR91, VC90q]. Fits [Nan90a, Ano92-31, Zei93]. Five [Ano91-162, Ano92j, Ano92-88, Ano92-188, Sco93, Sto93, Yag91c, Ano91-122, Ano94-184, Gre93b, Lov94, Mia90i]. Fix [Ude94j, Cra91b, Min93a]. fixes [Ano93-69]. flagship [Ga 93d, Ude90d]. flair [Egl92a]. flame [Ano92-57]. Flash [LM90, Tho93f]. flash-memory [LM90]. Flashdisk [Wsz90, Wsz90]. Flashdrives [Pou92j]. flashy [Ano92-101, Rei94a]. flat [Ano90-73]. Flatbed [Van94a]. Flatter [Ano92-121]. Flavors [Cla93b, Rei93c]. flaws [Cla93c, Sha94b]. fledged [Pou92a]. Flexes [Lov90a]. flexibility [Egl94g, Egl94e, Joc90a, Vor93]. flexible [Dav90, Fri94c]. FlexOS [Smi91a]. FlexScan [Ano91-97, Ano91d, Ano93-30]. flicker [Api94a]. flicker-free [Api94a]. Flight [Ano90-27]. flimsy [Ano94s]. flippy [Ano91-53]. flips [New93]. Flirting [Ken90]. floating [TR94b, Vor93]. floating-point [TR94b]. Flood [Rib93]. floor [Cot93a]. flop [Ham91]. Floppy [Ano93-108, Mia90d, Ano90q, Ano90-94, BL90, LM90, Led90, Mia90e, Ras90a]. Floptical [Ano92-199]. flout [Hol94]. Flow [HH94b, May94a, Ano94x, Smi94a, Ude93m, Wal94e]. Floyd [Ano93-107]. Fluency [Ano92-211]. Fluke [Smi94f]. Flung [Ano92-190, Ano94-37, ES94]. fly [Ano90-27, GW93, Hal94a, Ras90n, Way94d, Ano92-129]. flying [DG91]. Focus [Ano92-187, Ano92-188, Ano92-189, Ano92-190, Api90b, AWGY90, AG91, Bar93d, CS92, ES93, GW93, Lov93b, SG92b, SE93b, Wal94d, YG92, Ano91z, HC93]. Focuses [Ano92-47, Ano94-117]. fold [Ano93-64]. Folio [Sch93b, Ano91-149]. Following [Pou94d]. follows [Hol91]. Font [Gla90a, Ano93a, BL94, Egl91e, Orr94]. font-matching [BL94]. Fonts [Ano91-74, Cra90c, Min91c, Gla90a, Van93a]. Foolish [Gru90]. Fools [Ano90b]. Foot [SL92]. Footprint [Mit90a]. Force [Hal94h, Ude90d, Ano94-31, Ano94-106, Nie94]. Forces [Ude90b, Ano91-31, Cla93b]. Forecasting [Ste94b]. forefront [Yav94]. Foreign [Gre90f, Bar90j]. foremost [Ano91-187]. Forges [Var94b, Ude90g]. Forging [Mia93b]. Forgoing [Ano93-91]. Form [SE93a, Ano92-105, Len94a, MPW^{+}90, Way93f]. formal [Obe90, Ste92]. Format [HHP93, Ano92-52, Ano94-169]. formats [DH92, GD93, Yag92b]. former [Ano94-168]. formidable [Smi90c]. Forms [HM94, KM94, Ano91-111, Egl91d, Sch92a, SE93a, Van93a, You93]. FORTRAN [App91]. Forum [Ano91-93]. forums [Ano94-166, Ano94-182]. Forward [Ano92-205, EB90a, Nan93c]. found [Ano94-71, Ano94-135]. foundation [LDH94, Sch92a, Yag93f, You93, Bar91f]. Four [Cla90b, Cra90e, Pou92c, Rob90a, Sat90, VC90f, Yag92e, Ano92-44, Ano92-121, Ano94-85, Ano94-107, Ano94-184, AWGY90, CS92, Egl93f, Egl94b, Gar94, May94a, SU93, Ste94b, Ude94b]. FoxPro [Ano91-50, Ano93-31, Ano94w, MS93, VN90a]. FPU [Gre90b, San90a]. FPUs [Gre90b, San90a]. Fractal [Pic93, Vor93]. Fractals [Api93a, Pic93, Ano90-39, Ano91-179]. Fraction [Ano94-57]. fragmenting [Ano92-133]. Frame [MW94, Rya91e, Ano92-113, Ano93-125, Ano90-28, MW94]. Frame-relay [Rya91e, MW94]. FrameMaker [Ano90-28, Ano91-51, Ano92-58, Yag92f]. framework [VC90g]. Frameworks [Api94b]. France [Pou94a]. Frank [Hay90a]. Fray [Mit91b, Sta94b]. Free [Ano91-74, Ano91-43, Ano91-176, Fie90c, Fie90d, GP90, Hin92, Min90c, Nan91a, Ano91-33, Ano91-73, Ano92-91, Api94a, Die94, Dys92, MPW^{+}90, Nan90a, Nan91h, STA92, Sum90, Ano94-88]. free-form [MPW^{+}90]. free-market [Dys92]. Free-Space [Hin92]. Freedom [Tho93f]. FreeHand [Ano91i, Van94b]. Freelance [Ano93-61, Ano93-32]. freely [Daw94d, Fie90g]. Frees [Ano92-181, Ull93]. Freestyle [Ano92-218]. Freestyle/SL [Ano92-218]. Fremont [Ano90z]. frequently [Min91d]. Fresh [Ano92-59]. friendliest [Yag93j]. Friendly [Egl94g, Nan94c, Ano90h, Ano92-79, Tho91d]. Friends [Gre90d]. frills [Ano93-91]. Front [Ano94-63, Gre93a, HH94a, Ano92-59, Hed94, Obe90]. Frontier [Bar91f, Ano91-52, Nan93g]. Frontiers [Cra91g]. Fruits [Cra90h]. frustrated [Ano94-125]. frustration [Kay94a, SE94]. fry [Dio94]. FTSA [Ano92-56]. Fuels [Ano92-94]. fulfilled [Har91b]. Full [Ano90-30, Ano91-53, Api91a, Coo90a, Egl91b, Ras90l, Ano90-77, Ano90-92, Ano92-116, Ano93-133, Ano94-68, Ano94-92, Ano94-114, Bar90b, Bar93b, Daw94d, Dej94, Gre94b, Nan93a, Nan94b, Pou92a, Rei94a, Ric91, Ude93i, VN90d]. full-blown [Bar93b, Gre94b]. full-color [Ano92-116, Api91a, Bar90b]. full-featured [Ano94-68, Ano94-92, Ano94-114, Nan93a]. full-fledged [Pou92a]. full-length [Rei94a]. full-motion [Bar90b]. full-page [Ano90-77]. full-size [Ano93-133]. full-text [Ude93i]. full-time [Ano90-92]. fully [Die94]. fun [Ano93-82, Fie91l, Smi94c]. function [Ano91-129, Ano94-92]. Functional [Pou94d, Ano92-185]. functionality [Ano92-105, Law94, Ref94, Ric91, SS94, Woo90b]. functions [Ano92-183, Ano92-196, Gre92, Gut94, Way92b]. fundamental [RH93]. fundamentally [Ano94-117, Rei94e]. funny [Ano90b]. Furniture [Pou93m]. fuse [Ano92-186]. Fuss [Joc90b]. Future [Ano90x, Ano91-54, Ano91-91, Ano91-113, Ano92-63, Ano92-62, Ano92-61, Ano93-108, Cra91k, Fie90e, Hel91b, Mel93a, Mil94b, Nan91g, Pou93d, Pou94i, RLU^{+}91, RP90, Ste94b, Tho92e, Tho94j, Ano90-112, Ano91-153, Ano91-151, Ano91-172, Ano91-194, Ano94-65, Ano94-126, Ano94-150, Cra90o, GK91, GTN94, Hay90a, McN94, Min91b, Min92a, PP90, Pou94b, Pou94h, Rei94f, Sch92a, Taz91, The94, Way93f, Way94i, YS92]. FutureBasic [Tho93c]. futures [Ano94-116]. Fuzzy [Bar93d, She93c, UPSW94]. Fuzzy-logic [Bar93d]. fx [Bar90a].

G [Ano90-51, Die90b]. gadgets [Pou90s]. Gain [Ano90-62, Ano91-77, Ano93s, Ude90f]. gained [Ref94]. gains [Ano92-194]. Gallery [Pou94c]. gallium [Rob90b]. galore [Cra90f]. gambles [Mia90a]. Game [Ano92-103, Bet94, Pic93]. games [Ano91-150, Bus92, Pou91k, Way94h]. GammaFax [Ano92-64]. GammaLink [Ano92-64]. gap [Ano94-82, Rya90e]. Gaps [UY91, Wsz93c]. Gasteiger [Gas93]. Gates [Ano92y, Ano92-85, KPUG93, Ano92y, Ano92-108]. Gateway [Ano92-65, Per94a]. Gateways [Gla90b]. gathering [Vog90]. Gauges [Ano94-89]. Gazelle [Ano92u]. GB [Ano92-217, Ano93j, Ano93-47]. Gelernter [Ano92-114]. Gems [Bar92a]. General [Ano94-89, HM94, HHP93, Per94a, Ano94-90, Chr94, Die94, Egl93a, HK94, KM94, Ken90, MH91, Ano94g, Ano94-85, HHP93, Osh90]. general-purpose [Ano94-90, Egl93a, HHP93]. Generalized [Ano92-181]. generate [Ano91-39, Ano94-68, Gre94e, Ste94b]. generating [Api93c]. Generation [Ano90-57, Ano92d, Ano92-176, Ano94o, Api93c, Cra91e, Har91b, Ano90-62, Ano90-97, Ano91i, Ano91o, Ano91-133, Ano92-89, Ano92-161, Ano92-202, Ano92-199, Ano93-118, Ano94-82, Ano94-168, ADE90, Api93b, Bar90e, Fie91m, Mia90c, Mia93b, Pou90j, Pou93f, Ran90, Rya93a, Ude92d, Van93a, Ano92-125, Sch93a]. Generators [Api93c]. Genesis [TGN94]. Genetic [Sin94, Way91b, Lie93]. Genius [Ano91-105, Ano91-120]. Gentler [Ano92-79]. Gently [Fie91e]. geographic [Man90]. GeoWorks [Ano90-68, Lin91b]. German [Ano94-162]. Germany [Lin93]. gestation [Ano94-85]. Gestures [McA90]. Get [Ano94-64, Ano94-131, Ano94-158, BBWR90, Cas92, Cra90p, Gre94e, Jof93, Lan90c, Mos93, Rya93a, Ste91c, Ano90-94, Ano91-38, Ano91-55, Ano92-91, Ano92-211, Ano93-97, Ano94-76, Bac94, Bon90b, Cov90, Dau91, DW90, Law94, Lip90a, MW94, Ras91f, Rei94d, She93b, Smi90b, Sty91, Yag93b, Eug90]. Gets [Ano93z, Ano93-80, Ano93-93, Ano94-29, Ano94-61, Ano94-68, Ano94-126, Ano94-143, Cla90a, Cra90c, HS94, LAN92, MG91, SL92, Sma93b, Wau92, Yag93c, Yag93k, Ano91p, Ano91-180, Ano92-143, Cra91c, Cra91g, Die90b]. Getting [Ano92q, Ano92-66, Ano94-65, Fie90f, GETY91, Hal94i, Kiy93a, Min91c, Nad90c, Nan91h, Ras90h, VC90p, VN90f, VN90b, Ano92t, Ano92-175, Ano94-158, Len93a]. getty [Fie91c]. Gettying [Fie91c]. Ghost [Ano94-97]. Giant [Ano94-66, Nan93c]. Gibson [Ken93a]. Giga [Pet91]. Giga-Storage [Pet91]. gigabyte [Ano92-98, Ess94]. Gigabytes [Ano91-82, Ano91-142]. Gilbert [Ano91-85]. Gilmor [Gil94]. Give [Ano91-103, Ano94-168, Gre93f, Ano91-115, Ano94-61, Ano94-76, BEW92, Bar90j, Die90e, Mat94b]. Gives [Tho93f, Ano91-49, Ano93-130, Egl90c, Egl94e, Ga 94d, Hal94h, IC94, Pou94h, Pou90u, Pou94m, Sal94, VN90b]. giving [Fie91m]. Glance [Die90d]. Glasnost [Pou90u]. Glass [Yag90h]. Glimpse [Pou93d]. glimpses [Ano91-83]. Glitzers [Ano92-67]. Global [Kay94d, She91, Ude93h, Ano93-59, Ano93-129, Ano94q, Ano94-95, Wal93, Egl94e]. Globalization [Abr94, All92d]. Globally [Wal93]. globe [She91]. Glockenspiel [Sch90c]. glory [Bac94]. Glue [Ude91d, Way94a]. Gluttony [Nic93, Ano93-51]. Go [Ano91-57, Ano91-77, Ano92h, Api94d, Cla90b, Egl93f, Fie90g, KM94, Mia90h, RM92, Sma93a, Ups90, Vos90b, Yag91b, Ano90-88, Ano92-134, Ano92-205, Ano93-138, Ano94-178, Cla93e, Lie93, Nan91j, Nan92a, Rya90f, SBS91, Tho94c, Daw94b, Egl93a]. goal [Hal94f, Har91a, Lan94b, Ude93l]. goals [Ano92-174]. Godzilla [Ken93b]. Goes [Ano91-28, Ano91-117, Ano91-178, Ano92-158, Ano92-211, Ano92-223, Ano94-129, Cra90p, DB93, Ham92, Kay94d, Mit93b, Mur93, Pou92h, She91, Spi90b, Ude93h, Wau93c, Way94d, Ano94-178, Ga 93b, Hel90b, Pou93f, Wau93a, Yag90g]. Going [KRCT93, Smi93c, Ano92-144, Coo90c]. Going-On [KRCT93]. Gold [Ano92-199, Hal90]. Golden [MG90]. gone [Ano92-67, Rya91d]. Good [Fie90a, Har91b, Mit90b, SE93a, Sta94a, Wal93, BGE^{+}92, Bar92d, Egl93a, Egl94a, FH93, HH94b, Mas93, Per94b, Ras91i, Rya90g, Sca92, Ude90b]. good-looking [BGE^{+}92]. Gordon [Ano91-164]. Gore [Ano91-98]. GOSIP [Fis90a]. Gospels [Rob90a]. got [Api94a, Hel90c, Lin94]. gourmet [Ano91s]. Government [Ano94-67, Ano94z, Fis91b, Ran93, Rei94c, VFGZ94, Way93c, Fis90a]. Grab [YG92]. grabbers [Ano93-125]. grabbing [Ano92-147]. Grace [Ano91-130]. gracefully [Fie91e]. graces [Joc91]. grade [Ano92-82]. gradient [Ano93-43]. graduate [Ano90y]. Grail [KDG^{+}93]. Grammar [Egl91a, Pou90r]. grande [Way90c]. Graphic [Ros90b]. Graphical [Ano91-140, Car91a, EW92, Egl94g, Lin91b, Nic90, Ano91-188, Ano92-216, Ano94e, Die90b, Die90c, Gre94e, Hur93b, Kea91c, Rya90b, Sch90c, SBS91, Way93e, Wsz93a]. graphically [Ano90y, Ano94-178]. Graphics [Ano90y, Ano90-31, Ano90-40, Ano90-98, Ano91-103, Ano92d, Ano92-94, Ano92-147, Ano93-32, Ano93-74, Ano93-136, Ano94-68, Ano94-106, Ano94-89, Ano94-90, Bar90b, Bri90, BR93, Cra91i, Die90b, Gre90k, Lov91d, Lov94, Ric91, Ric92, Smi90f, Ups90, Yag92d, Yag92c, van91, Ano90f, Ano90v, Ano90-93, Ano91-97, Ano91-181, Ano92q, Ano92m, Ano92-116, Ano92-204, Ano93y, Ano93-116, Ano93-123, Ano93-125, Ano93-126, Ano94-129, Ano94-171, Ano94-182, ADE90, Api90b, Api94a, Cah90, Cla91a, Cla93e, CWVS94, Cra90p, DH92, Ga 93d, Gar94, Hel91b, Lov90a, Mel93a, NTS93c, Rei90a, Ros90a, Sch93b, Smi93c, Smi94c, Tho90a, Van94a, Van94b, Yag93i, Ano93-61, Cla93c, Smi93c, TS93a]. graphics-acceleration [Ano93y]. graphics-board [Api90b]. Graphing [Bar92d, Ano91n]. Graphite [Ano93-30]. graphs [BGE^{+}92, Ste94b]. grasp [Ano92-142]. gratification [Hel90a]. Gray [Ano90-32, Ano93-125]. gray-scale [Ano90-32, Ano93-125]. Great [Ano90-62, Ano90c, Ano91-55, Ano93-33, Mul94, Tho93i, Ude94a, Ano90t, Ano91-84, Ano92-60, GK91, Kay94c, Pou94r, Ude90f, VC90d]. greater [Lan94a]. greatest [Ano90-91]. greatly [Ano90-37]. green [KDG^{+}93]. Greening [RPRM92]. Greg [Lov94]. Grehan [Gre94b]. grep [NTS93b]. grew [Har90a]. GRiD [Ras90b]. Grip [Car92, She93b]. Grolier [KPC93]. ground [Ano92-56, Ano94-116, Cry92b, Die93a]. ground-breaking [Ano94-116]. groundwork [Dav94a]. Group [GB93, SE94, Ano94-132, GETY91, Ano91-32, UT91]. group-collaboration [Ano94-132]. Group-scheduling [SE94]. Groups [GETY91, Ano92-92, Ano94-54, Smi94a]. Groupware [Ras90h, Ras90i, Nan93h, Ude93i]. Grow [VC90h]. Growing [Ano93s, CV90a, Ano93-73, Ano93-75, Smi94d, Tho94h, Var94c]. Grows [Ano92-32, Ano92-119, Ano94-111, RT94b]. Growth [Ras90i, Ano91-151, Ano94-60, Eva94, Ran93]. guarding [Ano94k]. guessers [Ras91d]. GUI [Ano91-56, Ano91-127, Ano91-191, Ano92-171, Ano92-197, Ano94e, HWS91, Kea91b, Pep91, Sch91b, Wau93c]. GUI-based [Ano91-127]. Guide [Ano91-138, Ano91-139, Ano91-140, Ano91-141, Ano91-142, Ano91-144, Ano91-143, Ano91-145, Ano91-146, Ano91-147, Ano91-148, Ano92-162, Ano92-163, Ano92-164, Ano92-165, Ano92-167, Ano92-166, Ano92-168, Ano93-101, Ano93-102, Ano93-104, Ano93-103, Ano94-48, Fid92, Fie91k, NW92, Sta91a, VFGZ94, Ano90s, Ano90-54, Ano90-91, Ano91-101, Ano91-110, Cra90r, DW90, Ken93a, Len93b, Mas91c, Nie94, Ste94b, Yag92j]. Guideposts [Sma90]. guides [KPC93, Nov94a]. GUIs [Ano91-172, Ano91-188, Ano92-67, Hub92, Mia90b, Mia90h]. gulf [Ano94-77]. Gupta [Gre93a]. guru [Tho91c]. Gurus [Fie91l]. GW [Ano90-103].

Hacker [KGTD92, Ano91-105, Sto93].

hackers [KDG^{+}93]. Hand

[Ano90-29, Egl90c, Pou92m, Alf92g, AR91, Ano90-90, Ano92-179, Ano94-113, Egl90d, Mar92c, Pou92a, Van93c]. hand-held [Alf92g, Ano90-90, Ano94-113, Pou92a]. hand-helds [Mar92c]. handful [Ano91-78]. Handle [Ano92-88, Kiy93a, Ano90-97, Ano92-41, Api94d, Eug90, Laf91, Lin94, Nag93, Rya91b]. handles [Ano92-217]. Handling [Fie91e, Nag93, NHW93, Ano91n, Ano91-163, Kiy93a]. Hands [Ano94-189, FJK^{+}93, Die94, Smi94c, Tho93a]. Hands-On [FJK^{+}93, Smi94c, Tho93a]. handwriting [Ano92-61, Ano94-81, Ano94-162]. handwriting-recognition [Ano94-162]. Hannover [Ano93-98]. HanZon [Cra90c]. happens [Cos94]. happiness [Cra91d]. Happy [Ano90-33, Ude92a]. Hard [Alf91f, Ano91-45, Ano91-112, Ano93-47, Gre90i, Mia90e, Mia92b, RP90, VC90l, VC90p, Way94c, Ano91-63, Ano91-156, Ano92-112, Ano92-169, AWGY90, AG91, Bar91e, Bar90j, Fri94b, Ga 94c, LM90, Min90b, Pou90n, Pou94j, Rai90, VC90r, VN90b, VN90d, WET90]. hard-to-justify [Fri94b]. Hardcard [Ano90-34]. harder [Ga 94c, Mal93]. HardFacts [Ano90-35]. Hardware [VN90b, Ano90l, Ano90-35, Ano91-68, Ano91a, Ano94m, Ano94-88, Ano94-143, Ano94-159, Ano94-169, Bry94d, Cra92b, Gla91c, Hal94f, Kli90c, Kog94, Lin93, Lin94, Orr94, Pou91k, Rya91b, Smi93e, Smi94f, Smi94g, Tho93e, Way94h, Yag92a, Yag93l]. Hardware-based [VN90b]. hardware/software [Ano94m, Ano94-143]. harness [Nan91i, Pou94g]. harnessed [Ste91b]. Harvard [Cla93c]. Harvest [TH93]. hassle [Ano93a, VC90p, Ano94-88]. Hassle-Free [Ano94-88]. hassles [Ano93-51, Hal94a]. hauls [Cra90o]. haven’t [Ano93-105]. Hawk [Egl90b]. Hayes [Ano93-131, Hay90a]. Hazards [Ano94-133, ASS92]. HBO [Ano93-68]. HDTV [Ano92-68, Ano93-68, Fis90c, Lip90b]. Head [FT90, HM94, Mia90h, Ano92-211, Ano93-111, Ano94c, KM94, Pou93b, Rya94a, SBS91]. Head-mounted [FT90, Ano94c]. head-to-head [KM94]. headaches [Hal94a]. headlines [Ano94z]. headset [Ano90-82]. Health [Har91c, Wal94b]. Hear [Hed93b]. Heart [AMW90]. Heat [HK94]. heated [Egl94f]. heating [Rei90b]. heats [Ano94-116, Hay90b]. heaven [Gla91d]. Heavy [Tho93a, Ano90-92, Ano92l, Smi93d]. heavy-duty [Ano92l]. Heavy-Hitting [Smi93d]. Heavyweight [CV90c]. Heavyweights [Ano92v]. Hedging [Hel91d]. Heed [Fie91d]. Heights [Ano91-124]. Heirs [GAM90]. held [Alf92g, Ano90-90, Ano94-113, Pou92a]. helds [Mar92c]. Hell [Cri93, Pou94r, Ano93-56]. Help [Ano92j, Ano94x, Ano94-69, Pou93o, Way93f, Wig90, Ano90a, Ano91-81, Ano91-130, Ano91-136, Ano92-68, Ano92-89, Ano92-142, Ano92-206, Ano93c, Ano94-64, AG91, Cro93b, DE91b, DH90, Egl90a, Egl91d, FT90, GD92a, Laz93, Nel91b, PTUM93, Sta90, Tur90, VN90d, Way93i, Yag91d]. helpful [Tho93c]. helping [Ano92-194]. helps [Ano92-46, Ano92-181, HR91, New93, Pra93, Smi90a, SE93b, Yag90i]. herald [Ano91-153, TS93a]. Hercules [Lov90a]. Here [Ano91-57, Ano91-81, Ano91-129, Ano92-80, Ano94-79, Bry94c, Cla90b, Egl93a, Fin90a, Jef93a, Kay94c, Kiy93a, Lu93a, MPW^{+}90, Min92d, Mit93a, Nag92, Nie94, Rya91e, Rya91d, Sal94, Smi94d, Sty91, Tur90, Ude94j, Bry92, Cra90l, Der92, Fri92, HPPR93, Hal94i, Hay94a, Lip90a, MW93b, Rya90f, Ste94b]. Hero [Ano91-59]. Heroes [Pou93p]. hesitate [Ano94-57]. Heterogeneous [Ras91g, Ano91-67, Dej94, Egl90a, Pou91b]. Hewlett [Alf91b, Ano90-40, Ano90-63, Ano91-58, Ano92-122, Ano93-36, Cry92a, Egl92b, Joc90c, Lus90, Pou94b, Rei93b, Tho93b, Wau93b]. hexadecimal [NTS93a]. HiColor [Ano92-131]. Hidden [Ano92-69, Ano94-142]. hides [San90b]. hierarchy [Ano92-102]. High [Alf92b, Ano90x, Ano93-39, Ano93-47, Ano93-124, Ano94-70, Fin90b, HHP93, Hol94, Lan90a, Pou92f, Rya94a, Wil94, Ano90-106, Ano91-118, Ano91-134, Ano91-164, Ano91d, Ano91-181, Ano92q, Ano92d, Ano92-110, Ano92-119, Ano92-188, Ano92-218, Ano92-215, Ano93-65, Ano93-75, Ano94-37, Ano94-71, Ano94-72, Ano94-129, Ano94-131, Ano94-147, AWGY90, Bar90g, Bry94d, Cla94b, Del93a, DG91, Egl94c, FP93, Fox93, Fri90, Fri94b, Gut93, Hal94c, Hay90b, HH94a, HK94, HP94, Hur93a, KM94, Ken94, Len94b, LAD91, Lip90b, Nan92d, Nan93b, Pou92d, Ran93, Ras90d, Smi91b, Smi94g, Tho90h, Tho91d, Yag90d, Yag91b, Yag92c, Zei93, Ano94-71, Ano94-72, Bar90j, Car91a, Fox93, Kre91, Len94b, Lov91a, Lov91d, LS91a, Min90a, Ran93, Rob90b, VC90d, Yag92a]. high-bandwidth [Ano94-37]. high-capacity [Ano91-134, AWGY90]. high-definition [Lip90b]. high-end [Ano91-181, Ano92d, Ano92-110, Ano94-71, Ano94-131, Ano94-147, Del93a, Egl94c, Fox93, Fri90, Gut93, HH94a, HP94, KM94, Nan93b, Lov91d, LS91a]. high-impact [Yag92c]. high-level [Ano90-106, Nan92d]. High-Octane [Rob90b]. High-Performance [Fox93, Lan90a, Ano92-188, Ano93-75, HK94, Len94b, Pou92d, Tho91d, Ano94-71, Car91a, Len94b, VC90d]. high-power [Ano92-218, FP93]. high-powered [Ano93-65, Hay90b, LAD91, Yag92a]. High-Quality [HHP93, KM94, Ras90d, Tho90h, Lov91a]. high-res [Ano91d]. high-res/low-rad [Ano91d]. High-Resolution [Fin90b, Ano92q]. High-Speed [Ano93-47, Hol94, Ano91-118, Ano91-134, Ano94-72, Ano94-129, Bry94d, Cla94b, Fri94b, Hur93a, Ano94-72, Fox93, Kre91]. High-tech [Ano90x, Wil94, Ano91-164, Yag91b, Bar90j, Ran93]. high-technology [Bar90g]. high-volume [Egl94c, Smi91b]. Higher [Per93b, Ano92-49, Ano94-135, Bar90i, McN94, Mel93b]. higher-performance [Bar90i]. higher-quality [Mel93b]. higher-resolution [Ano94-135]. higher-speed [McN94]. Highlights [Ano94u, Ano94v, Pou94i]. highly [Ano94-117]. highs [Ano91-94]. Highway [Ano94-42, Ano94-67, Ano94-73, Rei94b, Ano94z, Ano94-98, Rei94c]. Highways [Ano94-58]. Hill [Ano94q, Hal90]. Hints [Ano94-141, Ano94-94, Fie91l, Tho90d]. hip [Ken93a]. History [Fie90e, YS92, Alf90a, Ano90c, Ano91-150]. Hit [Alf92d, Ano93-29, Cra90i, Fie90c, Var94a, Ano92h, Ano94s]. Hits [Daw94c, Fie92b, Mia94b, VC90b]. Hitter [Tho93a]. Hitting [Cla93d, Smi93d]. HLLAPI [Fic90]. HM [Pou90a]. Hobbit [Ano94o, Rya93a]. hoe [Gre93a]. hogging [Nic93]. Hold [Cla90b, Ano94-54, Ano94-182, Gre94a, Pou94b, Van94b]. holds [Gar94, Pou94h]. Hole [Smi93b, Min90j, Smi93b]. Holocaust [GTN94]. holostore [Par90]. Home [Ano92-158, Ras91e, TH92]. homes [Ano93-68]. Hong [Rei93a]. Honorable [Ano93-34, Ano93-48, Ano94-89, Ano94-90, Ano94-87, Ano94-88, Fox93, HM94, HK94, HHP93, Hol94, HP94, Len94b, Len94a]. honored [Pou90m]. honors [SH90]. Hood [Alf92g, Alf93, Ano92-208, Ano92-209, BL94, Hyo94, Jef93b, Ken94, Oth92, Pie94, Pou92d, Pou93f, Pou93h, Pou93g, Sch93d, SSF93, Sta93, The94, WS93, Way93i, Gre93h]. Hoops [Orr94]. Hope [Ano93-33, Woo90a, Yag93j, Ano92-63, Ano94-169, Eva94, Fri94a]. Hopes [Ano93-39, Ano92-106, Ga 93d, Rei93a]. Hopper [Ano91-130]. horizon [Bry94e, Sch93a]. horizons [Ano92-135, Yag93d]. horse [Sch92c]. Horsepower [Ga 94d, Ano92-187]. host [Ano90l, Gre94e, Tho93b]. hosts [Nan93g]. Hot [Ano94-155, Fie91i, Nel91a, Pou93t, Pou94s, Vos90b, Ano92-126, Ano93-75, Ano94-172, Cra90c, Ano92-57]. hot-swapping [Ano92-126]. HotDocs [Ano94-179]. Howard [Egl94b]. however [Hal94i, Ude93l]. HP [AR91, Ano90-36, Ano90-48, Ano91-60, Ano91-75, Ano91-76, Ano92-72, Ano93-36, Ano93-65, Api93b, Egl90c, Egl93c, ET93, Rei93b, Smi94g, Wau93b, Way94k]. HP/Apollo [Ano91-60]. HPFS [Min90a, Min90b, Min90i]. Hub [Ano94-158, Ano94-57]. Hubs [Ano94-76, Ano94-158]. Hugh [Ano90-39]. Human [Ano91-61, Ano92-42, Ber93, Yag91d, Ano90-66, Ano91s, Ano91-35, Ben91, SC92]. Hungarian [SH91]. Hungry [CV90d]. hunt [Ano92-184]. hunter [Ano90-100]. hunting [Ano90-27]. Hurd [Hur93b]. hurdles [Bar92a, Way93i]. hurry [Ras90c]. Hurwicz [HC93]. Hyatt [Ano91-85]. hybrid [Mia90b]. Hybrids [Lie93]. Hype [Lip90a, Ras92b, Ano94-101, Ude92d]. HyperCard [Ano90-37, Cra91l, Eva90]. HyperSpace [Ano92-217]. Hypertext [KUD92]. Hyundai [Ano92-65]. HZ [Api94a].

I/O [Ano92-79, Ano92-220, Gla90g, Pou93h, RH93, Van90, WK90, Par90]. I325VM [Ano92-199]. i486 [Ung91, Lin91c, San90a, VC90i]. i486-based [VC90i, San90a]. i860 [BBWR90, Fri91]. i860-based [BBWR90]. IBM [Alf91e, Ano90-86, Ano91-68, Ano91-132, Ano91-158, Ano91-196, Ano92-63, Ano92-73, Ano92-107, Ano92-108, Ano92-112, Ano92-153, Ano93-29, Ano93-38, Ano93-42, Ano93-37, Ano93-39, Ano93-40, Ano93-41, Ano93-61, Ano94m, Ano94-33, Ano94-126, Ano94-77, Ano94-127, Ano94-145, Bar90d, Bra90, Daw94c, Die94, Fie90h, Fox93, Ham92, Hel90b, Jef93b, Mas93, Min93b, Mit91b, Nan91b, Nan93a, Nan93c, Osh90, Osm90, Per92, Pou90c, Pou93n, RS90, RUHG93, Rei94f, Row93, RT94a, Sma91a, TR94a, Ude94f, Var94a, Woo90b, Yag92g, Yag93f].

IBM-compatible [Pou90c]. Icon [Gri94]. IDE [Alf91f, Ano94-129, Ano94-144, Bry94a, Ess94]. idea [VC90n]. ideal [Cla90b, Eva90, Pou91r, Pou93k]. ideas [Ano94-178, Sac90]. identifier [Ano91-43]. identify [Ano94-90, Ano94-87, HK94, KM94]. Identity [Rya91a]. idiosyncracies [Ano90u]. IDL [Ano92-218]. IEEE [LDH94]. If [Ano93s, Cau93, Der92, Fow91, Ras90f, Smi94a, Ano90-94, Ano91-81, Ano91-150, Ano91-169, Ano94-76, Api94a, Bis91, Cla93a, Egl94c, Lin94, Pou90m, Ude90a, Bis91]. II [Tho92g, Yag90i, Ano90y, Ano92o, Ano92-97, Ano92-222, Ano93-133, Egl90b, Egl90c, Fie90b, Fou91, Gar94, Gla90d, Sta93]. IIfx [Bar90a, Tho90c]. III [Ano90-48, Ano90-45, Ano91-118, Ano94-94, May94b, Tho90e, Tho92b]. III-MIS [Ano94-94]. IIISi [Ano91-75]. IIP [Joc90c]. IIsi [Tho91d]. IIT [Gre90b]. IIXL [Ano90-34]. Iliad [Ano91-32]. illustrate [Gre93d]. Illustration [Van94b, Van92a]. illustrations [Van93a]. Illustrator [Ano91f, Ano91n, Ano92-82, Ano93-43, Vor92]. Illustrators [Van92a]. Image [Ano91-62, Ano94-78, Ano94-89, Bar92c, Lov91a, Way93d, Ano90-76, Ano91-138, Ano91-184, Ano92-41, Ano92-68, Ano92-90, Ano92-143, Ano94-40, Ano94-131, Egl94f, Har91a, HR91, Ran90, Tho93b, Van94a, VN91e, Vor93, Wsz93a, Ano94-40]. image-compression [VN91e]. Image-editing [Lov91a, Ano94-40, Egl94f, Ano94-40]. image-processing [Ano94-131]. imaged [Loc91]. Images [Ano90-38, Min90j, VN91e, Ano92d, Ano92-142, Ano93-116, Ga 93d, Mel93b, Smi94a, Tho90h, Ups90, Way93f]. Imaginative [Smi94c]. Imaging [Ano90-32, Ano91-138, Ga 93d, Ano93-132, Ano94-176, Ano94-177, Mic91, Oth92, UPSW94, Ano92-167]. Imagining [Cot93a]. Imara [Day90]. immediate [Pou94b].

Impact [Ano90-30, GK91, Ken93b, Yag92c]. Impaired [Hol94, Ano94e, Sch91b]. impairment [Hur93a]. impediments [Ano91-39]. Impersonator [Ano92-99]. impinge [Nak94]. implement [Ano94x, Ste91d, Var94a]. implementation [Daw94d, Ude93m]. implementations [Fox93]. implementing [Kor93]. implied [Ano90-84]. Importance [Ano92-74]. Important [Ano90l, Ano93v, Ano94g, Ano94-89, HM94, HHP93, KM94, Lan94b, Rei94d, Sta94a, Ude94j]. imports [Ran93]. impossible [Gil94]. Impression [Mit90a]. Impressions [Ano92-55, Ano92-56, Per92, Pou92a, Sma93a, TH92, Tho92d, TH93, Ude93f, Wau92, Wau93a]. impressive [Ano90-107, Ano93-66, Ano93-140, Lin94, Van93a]. Improv [Ano90-113, Ref94, She93a]. Improve [Ano94-125, Ano94-152, Egl91a, Ano92-68, Ano92-112, Ano92-172, Ano93-46, Hed93a, Pou93h, Wal94b, Zei93]. Improved [Ano94-115, Ano90-62, Ano90-81, Ano91n, Ano94-40, Ano94-57, Ano94-106, ESN94, Nan93d, Rya90b, Ude94j, Yag93g]. improvements [Ano91-149, Ga 92b, Ken90, Nan91c, Nan93c, Ude93i]. improves [Ano90-37, Pou93e, Yag93k]. improving [Bry94e, Rei93d]. in-depth [Ano91-165, GR94]. incarnation [Ano90r, Ano91-51, Way94j]. inch [Ano92-45, Ano94-89, Api94a]. include [Ano91-172, Api94d, Fox93]. includes [Ano90-60, Ano91-126, Ano94a, Gre94b]. including [Ano91-162, Api93b, Fie90a, Pou90q, Pou91k, Sch93b, Ude93i, VN90a, Wau93b, Way93i, Wsz93c]. incompatibilities [Hal94f]. Incompatibility [Ano94-169, Rya90a, San90b]. inconsistencies [Ano92s]. incorporate [Alf92d]. incorporates [Ano92-152]. incorporating [Bry94c, Ste94b]. increase [Mar92a, Nel91b]. increases [Rya91c]. increasing [Gla90c]. increasingly [Alf90b, Ano94-149]. Incredible [Ano91-63]. indeed [Ras91k]. Independence [Ano93b, Api94b]. Independent [Ano94-169, Ano94-119, Api94c, Hel90c]. independently [McG90b]. index [Ano94a, Way91a]. indexes [Gre94b]. Indexing [Pou90b, Har93, Loc91, Ude93i, Way91a]. India [Ano91-120, Ano93-56, Ran93, Ude93g]. Indian [Ano91-120]. Indigo [Ano91-178, Ano92-155]. Indigo2 [Smi93c]. indirect [Ano92-156]. individual [Api90c]. individuals [Daw94d]. industrial [Cla94a]. industries [You94]. industry [Ano90-112, Ano90c, Ano91-186, Ano92y, Ano94m, Ano94-54, Ano94-65, Ano94-116, Bux90a, Cra90k, Pou94i, RPRM92, Saf93]. Indy [Smi94c]. inevitable [Dav94a]. Inexpensive [Ano94-79, Joc90c, Tho93b, VC90e, Way94h, Ano90-75, Ano91-181, Ano93h, Ano94-143, Api90c, Egl92c, Gar94, Nan94b, Smi93b]. InFax [Ano93-44]. Infini [Ano91-64]. Infini-D [Ano91-64]. influence [Tho94e]. influenced [Ano90t]. influential [Ano90-112]. info [Ano93-66, Wsz93b]. InfoAlliance [Ude91a]. Infoglut [Ano92-75, Ano92-92]. Infolio [Ano92-76]. infolock [Ano91-61]. Information [Ano91-67, Ano91-130, Ano92-51, Ano92-77, Ano92-181, Ano92-207, Ano93-97, Bon90b, Cla92a, Jac91, Rob91b, Yag91d, ASS92, Ano90-35, Ano90-53, Ano91u, Ano92j, Ano92d, Ano92-75, Ano92-186, Ano94g, Ano94n, Ano94z, Ano94-64, Bon90a, Cas92, Cox90, Cra90h, Cro93b, Gre91a, Gre93a, Har93, Mat94b, May94b, Nov94b, Pie94, Ras90c, She90b, SE93b, Ste91a, VFGZ94, Vog90, Web92, Woo90b]. Information-Age [Ano92-77]. information-highway [Ano94z]. information-sharing [Ano94z]. Informix [Die90c]. infrared [Egl94d].

Infrastructures [Jef93b]. Inherit [Ano91-65]. Inheritance [Ano94-80, Hol93a]. inherits [Hal93f]. initial [Ano94-113]. injuries [VFGZ94]. Ink [Egl92b, Rog91, Ano93-113, HM94, Joc91, KM94]. Ink-Jet [Egl92b, Ano93-113, HM94, Joc91, KM94, Rog91]. INMOS [Pou90h, Ano91-173]. inner [Cah90, SSF93]. Innovation [Ano90-50, Ano91-186, Ano92-181, Pou91h, Red93]. innovations [Vor93]. Innovative [SC92, Cau91, Cla91a, Cra90h, Egl93a, Hal94c, MG91, Ras90l, Rya94b, Smi94g, Tho94b]. Input [Ano93-37, Nag93, Bar92a, Die94, Egl90d, Gre90c, Pou93a, Tho90b, Ano92-140]. ins [VC90a, NTS93c]. Insider [KCU^{+}93, Lan90b]. Insignia [Ano94-157]. Inspiration [Cra90j]. inspirational [Cra90j]. install [Nan94b, VN91d]. installation [Cri93, Nan91c]. installed [Egl94e]. Installer [Cri93, Ano93-56]. installing [Fie90g]. installment [Ano91y, Ano92-29, Ano92-147, Yag90a]. installs [Nan94d, Pou94m]. instant [Egl93d, Hel90a, Hel90a]. Instant-C [Hel90a]. instantly [Ano91-156]. instead [Nan94b]. Instinct [Pou93k]. Institution [Smi93b]. instructions [Ano91-169]. Instruments [Ano90-103, Sch93d, Ano93-70, Nad90f]. Insurance [VN91b]. INT [Ano92-185]. integral [Hal93b, Hal94f]. integrate [Ano94h, Ano94-92, Len93b, SL92]. Integrated [Rei93e, Ano91-32, Ano91-67, Ano91-107, Ano92-196, Hal93b, Lin90b, San90a, TS93a, VC90f]. integrates [Ano92-217, Ano94-184]. Integrating [Ano91-67, Pou93q]. Integration [Ano94h, Ano92-197, Ano93z, Law94, Nan93d, Van93a]. integrity [Ano90-78, Kre91, Min90f, Nad94b, Top91]. Intel [Alf92a, Ano91-44, Ano91-68, Ano91-108, Ano91-165, Ano92-53, Ano92-78, Ano93-45, Ano93-46, Ano94-32, Ano94-81, Ano94-82, Ano94-83, Ano94-129, Ano94-171, Cla90b, Fox93, Fri91, Gre90b, Hal93c, Hal93e, Hal94c, Hal94b, Mat94b, Pou94f, RH93, Rya94b, Smi93d, Smi93e, Sta94b, Ung90a, Way94k]. Intel-ATI [Ano94-171]. Intel-based [Ano92-53, Ano94-32, Hal94b, Ung90a]. Intel/HP [Way94k]. Intel/IBM [Ano91-68]. Intel/VLSI [Ano93-45, Sta94b]. Intelligence [Ano92-49, Ano94-86]. Intelligent [Ano91-141, Rei93d, Rei94f, Ano90x, Ano90-52, Ano92-74, Ano94g, Ras91o, Rya91a, Ano90-40, Ano94-77]. intelligently [SB94]. intensifying [Bry93b]. intensive [Ano94-107, Mul94]. interact [Ano92-144, Ano93-37]. interaction [Ano91-35]. Interactive [Ano92c, Lu93a, MPW^{+}90, Ano94-174, Ano94-182, Cry94, Har93, Tho90b, Ano94-179, SY90]. interapplication [Gil94]. interchange [War92]. interchangeable [Cox90]. Interconnect [Ano93-85, SE91]. interconnected [Ano94-149]. Interconnecting [Lee92, Ude91d]. Interconnection [Ano91-34, Hin92]. interesting [Ano94-65]. Interface [Alf91f, Ano90-41, Ano92-79, Ano92-207, Bar92g, Cla91b, Cla91a, Gla90e, Nic90, Alf90b, Ano90-59, Ano91-43, Ano91-135, Ano92-31, Ano92-138, Ano92-216, Ano93-43, Ano93-48, Ano94-72, Egl90a, Ga 92b, Ga 93b, Hay90a, McN94, Per94b, Sch90c, Way93e, Wsz93c, Ano92-60]. Interface/Shplinterface [Ano90-41]. Interfaces [Ano91-140, Ano91-172, Kea91c, Yag90h, Ano90-29, Ano90-97, Ano91-188, Ano92-79, Ano94-54, ES94, Gru90, LHR90, SBS91]. Interfacing [Swa93]. internal [Ano93-125, Api94d]. internals [Ano91m]. international [Egl94c, Kay94d, She91, Ano90-57]. Internet [Ano92r, Ano92n, Ano93e, Ano94-50, Ano94-63, Ano94-150, Ano94-149, Ano94-172, Ano94-181, Der92, Fie91g, Hal94e, KPC93, Nov94a, Nov94b, Rei94b, SE91, Sto93, VFGZ94, Vos92]. Internet-In-A-Box [Ano94-181]. internetwork [ES93, Gib90a]. Interop [Ano94q]. Interoperability [Ano91-69, Ano94-154, Nan91j, Ano92-77, Cra90r]. interoperable [Fri94a]. interoperatibility [Ano92-119]. interpreter [Sch92c]. interprocess [Gre90d, Gre90g, Mal91b]. Interrupt [Alf92f, Ano92-70, Jef93a]. Interrupt-Processing [Alf92f]. Interrupts [Alf92f, Ano92-70, Pou92g]. intervention [Sha94c]. interview [Ano91-165]. interviewed [Ano94-168]. intriguing [Ano91-120, Ano93z, Rei93e]. introduce [Ano94-77]. introduced [Ano91-68, Hal94d]. Introduces [Ano91l, Ano90-68, Ano90-80, Ano91-125, Ano94-117, Daw94c, Die93c, Gre93f, Lin91a, Smi90g]. Introducing [Gre90g, Gre92, Tho92d]. Introduction [Ano91-69, Ano91-95, Ano91-104, Ude94a, Ano92x, Egl94f, Hal94b, Har90a]. introductions [VN91c]. introspection [Rya91a]. Introversion [Hal94e]. Intuit [Kof93]. invade [Ano92-77]. invades [Pou92n]. invented [Cra90l]. inventor [Ano91-85]. inventory [NTS92a, Ste94b]. investigates [Ano91r]. investment [Alf91c]. investments [Ano92-194]. Invisible [Nan93d, Nan93d]. invisibly [Kor93]. Invitation [Ano90-71, Joc92b]. involved [Ano90j, Ano93-68, Mil94a, Pet91, Rei94c]. involves [Slo91]. IP [Ano92-85, Ano92-200, Ano93-115, Daw94a, Daw94b, Fie92a, KYT^{+}92, Nan93g, Rei94b, Way94b]. IPC [Yag91e]. IPX [Ano91-159, Daw94a, Nan94a]. IQ [Ano90-42]. Iris [Smi90f, Ano92-155, Smi90f]. iRMX [Ano92-223]. iron [Yag90l, Bar90g]. ISA [Ano92-219, Ano94-90, Fox93, HP94, San90b]. ISA-Bus [Ano94-90]. ISDN [Ano94-60, Fis90c, Fri94a, Fri94b, Fri94c]. ish [Yag90h]. IsiCAD [Ude90b]. island [Ano94-132]. IslandDraw [Ano91-72]. IslandPaint [Ano91-72]. islands [Ano92-75]. IslandWrite [Ano91-72]. isle [Ano91-72]. Isn’t [Ano94a, Ano93-72, Daw94d, Fri94a, Gru90, Saf93, VC90n, VN91b, Wsz93c]. ISO [Hyo94, Sta90]. isolate [Smi94f]. Israel [Red92b]. issue [Ano93-50, Yav94]. issues [Ano91-189, Ano92s, Ano92-61, Ano92-91, Ano92-113, Ano93e, Ano93-55, Mil94a, Nad94b, Pou92l, SU93, Sta94a]. Italian [Ano92-128, Red93]. Italians [Red93]. Italy [Ano93-58]. Itself [Yag92g, Ano93-100, Nan94d]. ITV [Lu93a]. IV [Ano90r, VN90a]. I’ve [Hel90c].

Jack [Woo90b]. jaggies [Ano92-172]. jam [Ban92]. Jams [Ano94-150]. Japan [Ano94-42, Rei92a]. Japanese [She93c]. Jasmine [Cra90a, Gre90i]. Jerry [Ano90-60, Pou90k, Pou90o, Pou90q, Pou90p, Pou90s, Pou90t, Pou90u, Pou90w, Pou91o, Pou91j, Pou91l, Pou91k, Pou91m, Pou91n, Pou91p, Pou91r, Pou91q, Pou91t, Pou92e, Pou92f, Pou92g, Pou92h, Pou92i, Pou92l, Pou92o, Pou93m, Pou94l, Pou94m, Pou94n, Pou94k, Pou94i, Pou94o, Pou94p, Pou94q, Pou94r]. Jet [Ano92e, Egl92b, Rog91, Ano93-113, HM94, Joc91, KM94, Joc90a]. Jewels [Ras90g]. Jim [Car94, Hur93b]. Job [van91, Ano92-195, Nan91l, Sha94c, Way93e]. jobs [Ano91-71]. John [Ano90-100, Bar91f]. Join [Sta94b, Ude90b, WK90, Yag90m]. joins [Ano92-220, Ano94-189, Han92b]. Jordan [Tab93]. Journey [Ano91-116, Api91b, Ano90-86, Pou94k]. Joy [Nov94a]. JPEG [STA92]. Jr. [Ano92-92, Buc92]. juggle [Gre90c]. juggling [Mia93a]. Jukebox [Ano91-142, Pou91p]. jump [Smi90d, You90]. Jurassic [Cri93]. Just [Ano90-65, Ano91-73, Ano92-141, Ano94-85, Ano94-137, ADE90, Gre90d, Gre90e, Hel90a, Mos93, WET90, Ano91-85, Ano91-115, Ano92-33, Ano92-40, Ano92-208, Ano94-66, Ano94-148, Ano94-159, Bac94, Fin90b, Gre90j, HP93, Mia93a, Nov94b, Ols94, Rya90c, Tho90g, Van92a]. justify [Fri94b]. Justifying [Ude94k]. JVC [Tho94c].

K5 [Hal94c]. Kaleida [Ano92-107]. Kapor [Ano90-50]. Kay [Ano91-42]. KB [Ras91i]. Kbps [Hol94, Egl94b]. Keep [Ano92-150, Car92, Nad90c, Ano94-82, Fow91, Mia90i, RH93, Wal94c]. Keeping [Ano92-82, FW93, Day90]. Keeps [Ano93u]. Ken [Ano91k]. Kenner [Ano90-39, Ano90-100, Ano91-179]. Kensington [Kli90a]. Kernel [Ano91-68, Smi90a, Ano92-183, Hol91, Var94b]. Kernels [Var94a]. Key [Ano94-73, Api94c, Pac94, Ano90c, Ano92-127, Ano92-178, Ano94-54, Kli90b, Lov92c, SU93]. keyboard [Nag93]. Keyboarding [Bar93a]. Keyboardless [Mia92a]. Keyboards [Ano92-208, Pou93l, Gre90c]. Keys [Ano91-74, Ano92-208, Ano94-89]. keystrokes [Min92d]. Kicking [VC90m]. Kicks [Jof93]. kid [CV90a]. Kind [Pou94b, Ano92-199, Min92a, Rya91a]. Kinder [Ano92-79]. kinds [Sta91a]. King [Ano91-29, Ano94-165, RP90]. kiosks [Ano92-208, Har93]. Kit [Ano90-44, Min93b, FT90, Ude94l, Ano91-86]. Kits [Ano93s, Ano94-45, Pou92i]. Knife [Van92a]. Know [May94a, ESN94, Egl94c, Hay93, Man90, SB94, She93b, Ste91d]. Knowing [Ano92-42, Ken90]. Knowledge [EB90b, EB90a, Hed93a, Loc92, Eva90, Lie93, Lu93b, Sum90, Tho91a, Wal94d]. knowledge-based [Eva90, Lie93]. knowledge-engineering [EB90b]. KnowledgeSeeker [Ano91-157]. KnowledgeWare [Gre93a]. known [Ano91-47, Ano94-72, Fin90a]. Knows [Rei91b]. Kodak [Egl93d, Ga 93d, Joc91]. Kong [Rei93a]. Kopin [Ano94c]. Korn [STA92]. Korn-shell [STA92]. Kudos [Ano92-83, Ano92-84, Ano92-92]. Kurzweil [Ano90x, Ano94-86]. Kyocera [Red91a].

L [Ano92-86]. L40 [Mit91b]. Lab

[All93i, Ano93-47, Ano93-48, Ano94-89, Ano94-90, Ano94-87, Ano94-88, FJK^{+}93, Fox93, Ga 92a, HM94, HK94, HHP93, Hol94, HP94, Len94b, Len94a, Sac90, Ano94-119, Ano90o, Ano91-88, Ano91-89, Ano91-108, Ano91-122, Ano91-170, Ano92s, Ano92-28, Ano92-27, Ano92-31, Ano92-88, Ano92-126, Ano92-173, Ano92-189, Ano94-27, AMW90, AWGY90, AG91, BGE^{+}92, BEW92, CW92, CS92, Die91, DE92, EWT90, Egl91a, ET91, EW92, Egl92a, Egl93c, FJPR93, GE92, GAW93, Gre90b, GETY91, GD92b, GW92, Gre93b, Gre93c, Gre93e, Gre93f, GW93, Hur93a, Joc92b, Lab92b, Mit91a, Mit91b, Mit91c, NW92, Smi93b, SE94, TG91, Tho92a, Tho93h, Wsz93a, Wsz93c, Yag92e, Yag92b]. Laboratory [Sto93]. labs [Ano91-70]. Lack [Bar92a, Egl93a]. Lacks [Ude91c, Ude93i]. lag [Ano92-102]. Lags [Ano94-42]. LAN [Min90f, Alf91c, Ano90-62, Ano90-83, Ano91-88, Ano92-85, Ano92-139, Ano92-188, Ano92-189, Ano92-195, Ano92-222, Ano93-48, Ano93-131, Ano94q, Ano94-91, Ano94-143, Ano94-170, Ano94-177, CV90a, CV90c, Dav90, Egl91b, Egl94d, Egl94e, Fri94c, Gre91b, Gre91c, Min90f, Mit91c, Nan91k, Nan92c, Nan92d, NTS92a, Nan93d, NTS93d, Nan94c, Pou93j, Pou94m, Ras91d, Ras91e, Ras91f, Ras91j, Ras91n, Ras91m, SL92, Slo91, Sta90, Ude90d, UY91, Ude91b, Ude93h, Ude94k, VC90j, VC90k, VC90n, VC90p]. LAN-based [Gre91c, Slo91]. LANBit [Ano94-119]. Lancaster [Lan90b]. Land [Ano91-102, Rei92a, Tho90i, VC90t]. LANDesk [Ano93-127]. landmark [Ano90-108]. lands [Bar90g]. landscape [Ano90-77, Ano91-31]. Lane [Ano94-171, VC90a, Bry94c]. Lang [You90]. Language [Ano91x, Ano91-182, Ano94-55, Ano94-151, App91, Nan90b, Ano90-59, Ano90-115, Ano91z, Ano91-163, Ano91-187, Ano92-30, Ano92-117, Ano92-141, Ano94g, Ano94w, Bet94, Chr94, GR90, Gre94a, Har90a, Ken90, NTS92b, Nan93f, Obe90, Pet90, Pou90d, Pou91d, Ras91c, Swa93, Ano92-181, Gri94]. language-neutral [Nan93f]. Languages [Ano91x, Ano91y, Ano91z, Ano92-29, Ano92-30, Ano94-117, Woo91b, Ano91-182, Ano92-83, Ben91, Cro93a, DH90, Egl94b, GFK94, Mul94, Pet90, Pou94d, Ras90p, Rya90b]. LANlord [Nan92d]. LANMeter [Smi94f]. Lanmodem [ATU92]. LanRover [Ano92-86]. LanRover/L [Ano92-86]. LANs [Ano91-34, Ano91-122, Ano93-129, Ano94-92, Ano94-147, Ano94-149, Api91b, Bar90k, Bry94d, Cla93d, ET91, Fie91g, Gre91c, Kli90b, LDH94, MW94, Nan91f, Nan94d, Ras91a, Ras91g, Smi94e, Smi94f, SE91, Ste91e, UM90, Ude91d, VC90k, VC90s]. LanSafe [Ano92-222]. LANstation [Mit90a]. LANtastic [Ano90-44, Ano92-87, Ano92-173, Ano94-92, IC94, Nan93d, Pou93j]. LanWorks [VC90t]. LapCAD [Ano93-130]. LapLink [Ano90-45, Ano91-135]. Laptop [Ano90-46, Ano90-47, Ano93f, GK91, Mit91a, Ano94-135, Nad90d, TGN94, VC90b]. Laptops [Ano91-172, Pou91r, Pou92j, Cla90b, Ras90b]. Large [HHP93, Ano91-125, Ano92-180, Ano93-48, Cra90f, Har93, Rei94b, Rib93, She91, She90b, Way94g]. large-screen [Ano91-125]. largely [Ude93b]. Laser [Ano91l, Ano92-50, Ano92-88, Ano93-49, Ano94-34, DW90, DE90, Egl93c, Tho93g, Vos90a, Ano91-48, Ano92-224, Ano93-70, HM94, Joc90c, Joc90b, KM94, Lin91a, Mia92b, Per94b, Van93b, Tho93g]. Laser-Printer [Ano92-50]. laser-quality [Egl93c]. LaserJet [Ano93-36, Egl90c, Ano90-48, Ano91-75, Ano91-118, Cra90c, Cry92a, Egl93c, Joc90c]. LaserJets [Min91c]. LaserMaster [Ano92-224]. LaserPrinter [Ano92-73, Ano92-153]. Lasers [Van93b, Tho94a]. LaserWriters [Joc90b]. Lassoing [Bar92e]. Last [Ano92-108, Ano92-100, Min90c, Pou94e, Ano90c, Ano92k, Egl94f, Fie90e, LB90]. Late [Ano94-90, Ano92-104]. Later [Ano93g, Ano94-72, Fis90a]. Latest [Api93c, Fis90a, Han92a, Yag90e, Ano90x, Ano90-41, Ano90-91, Ano91-66, Ano91-135, Ano92-85, Ano92-132, Ano93-92, Ano93-131, Ano94a, Del93a, Egl94g, Ess94, Fie91m, Ga 92b, Hal93e, KM94, Kea91a, Lin94, Min90h, Min90f, Nan93h, Pou91d, Pou91n, Pou92i, RT94a, Smi93b, Tho92h, Tho92c, Tho94b, Ude91c, Van93b, Vor92, Way94j]. Lattice [Nan90a]. launcher [Ano92-66, Wsz93c]. Launches [Hal93c]. laurels [Bux90a]. Law [Bar91f, Ano90e, Cos94, Way93g]. Lawrence [Sto93]. lawsuit [Ano90-41]. lawyer [Ano90-84]. layering [Ano93-43]. Layers [Sco93, Vor93]. Laying [PP90]. Layoff [Eva94]. layout [Van93c]. lays [Dav94a]. LC [Ano92-97, Ano92-97]. LCD [Ano94-112, Ano94-135, Yag93b]. LCDs [Api91a, Bar91b]. Lead [Ano94-138, Pou93h]. Leader [Sha94b, Cla93c, KPUG93]. Leaders [She93c, Ano92v, Ano94-87, Joc92a]. leading [Ano94-116, Gas93, Hal93b, Len94b, SBS91, Way90b]. Leads [Ano93m, Mia90d]. Lean [Ano91-76, Ano92-119, Ano94-126]. leaner [Ano91-117, Hay94b]. Leap [Ano94-66, Nan93c, Ano92-155]. Learn [Hed93b, Ano92-147, Bus92, Par92, Pou90i]. Learning [Ano94-93, Fie92a, Pou91c, Sey93, Ano90-116, Ano91-191, Ken93a, Sch91b, Way93i, Woo90c]. learning-disabled [Ano91-191, Sch91b]. Learns [Ano91n, Ano94-175]. least [Ano94-86, Ano94-107, Api94d, Pou93i].

Leave [Nad90d]. Leaves [Ano94-107, Egl91b]. leaving [Ano91-70, Ano91-169, Tho93b]. LED [Ano90-67]. left [Ano92-105]. Legacy [HH94b, Ano90-49, Pou91l]. legal [Cos94, Hal93e, Sch93d, TGN94]. Legality [Cos94, Ano94-98]. Legend [Woo90d]. legitimate [Ras92b]. lend [Ano92-117]. length [Rei94a]. Lens [Rob91b]. Less [Ano90-62, AT92, Cal91a, Tho93f, VN91e, Ano92q, Ano94-71, Ano94-129, Ano94-143, Egl94c, Gre94d, Ham92, Pou93a, Sca92, Ude90f]. less-know [Egl94c]. Lesson [Pou90p]. Let [Ano91-77, Ano92-89, Eva94, Fie90i, Fie91e, NTS93c, Ano92d, Ano93-37, Ano94x, Bry94e, Egl90d, Kiy93b, MW92b, Mia90i, Mit91c, Pou93c, Ras90n, She93b, Tho94d, VC90g, Vel90, Way92a]. lets [Ano90i, Ano90q, Ano90-55, Ano91-156, Ano92-218, Ano94g, Ano94l, Ano94f, Ano94-174, Ano94-181, Cau91, Cov90, Dio94, Egl94g, Gre90h, Lu93b, Mia94b, Min90h, Ros90a, Sac90, Way93c, Way94a, Web93, WK90, Woo90c]. Letter [Min90d]. Letters [Ano92-90, Ano92-91, Ano92-92, Ano92-93, Ano93-60, Ano93-50, Ano93-52, Ano93-51, Ano93-53, Ano93-55, Ano93-56, Ano93-54, Ano93-59, Ano93-58, Ano93-57, Ano94-95, Ano94-94, Ano94-103, Ano94-96, Ano94-98, Ano94-97, Ano94-100, Ano94-99, Ano94-101, Ano94-102, Ano94-104]. letting [Way94c]. level [Ano90-106, Ken92, Nan92d, TG91, Yav94, Tho93d]. Levels [GTGW94, Ste94b]. Levy [KYCD92]. Lexmark [Per94b]. Liant [Api94b]. Lib [Ano92-199]. Liberty [Pou90q]. Librarian [Lab92a]. Libraries [Ano91-78, Ano92-31, Ano92-187, CWVS94, Orr94]. Library [Api94c, Ano92-183, Ga 92b, Kea91c]. License [Ude94f]. lie [Bar90g]. Lies [KYT^{+}92, Ano91-54, Ano92n]. Life [Ano90a, Ano94-105, Bet94, Dys92, KYCD92, Len94a, Min90e, Pou90q, Ste91b, Ano90g, Ano92-160, Daw94c, Egl92e, Hed94, Ken93b, Kli90c, VC90d, VC90s, Web94, Ano94-131, Bet94]. Lifetime [Hed94, Nel91b]. Lift [Wau92, Yag93k, Ano91-49, Ano94-61, Egl94f, Ano94-61]. Light [Ano93-33, Egl94d, Isl92, Ken93a, Pou94p, Gla90h]. lighting [Ano92-156]. lights [Ano92-57]. lightweight [Ano92l]. Like [Ano94-85, Ano90-69, Ano91-157, Ano92-50, Ano94w, Ano94-68, Fie90l, Gre94a, Kay94c, Min90e, Pou93a, Ras90e, Taz91, Tur90, Ude93h, VC90n]. liked [ES94]. likely [Ano93-68, Ano94-72, Hal94b, Pou94b]. Liken [Ano92-99]. Lilliput [VC90t]. limit [Way93g]. Limited [Ham92, Ano94-92, Ano94-125]. limited-function [Ano94-92]. Limits [Ano93-112, Hol90, Ano92-61, Ano92-102, Ano92-191, Rya90f]. Lindstrom [Lin94]. line [Ano92-217, Ano93b, Ano93c, Ano93-126, Ano93-134, Ano94-96, Ano94-129, Ano94-149, Ano94-166, Cry94, Daw94c, Del93a, Egl92c, Egl94b, Egl94g, Hur93a, Lav91, Lov92b, Ban91]. line-impairment [Hur93a]. Lines [Ano92-185, Hol94, Ano92-177, Rei94b]. Lineup [Ano93c, Ano91-126, Gar94]. Lingual [Wal93]. Link [Ano91-135, Ras91g, Ano93-134, ES93, Gre92, Ude94h, Yag90f]. Linking [Ano94-37, Dej94, Nan93g, SU93, Har93]. Links [VC90q, Vos90b, Ano94-149, Lin93]. LinkWorks [Smi94a]. Linux [Ano94-97, Daw94d]. Lisp [Ano91y]. List [Bar91g, Ano94-168, Ude90e]. Listen [Ano94-93]. listing [Ano91-138]. Listings [HM94, KM94, Ano92t, Ano92y, VFGZ94]. lists [Kay94c]. Lite [Ano92-119, Api93b, Ano90-75, Egl92c]. Lite/25c [Egl92c]. literally [Pou93f]. literature [Ano90-39]. Lithium [Len94a]. Litigation [Ano90-50]. little [Ano91-47, Ano93-109, Gre93c]. little-known [Ano91-47]. Littlest [Api93b, Yag91e]. Live [Ano91-158, Min90i, NHW93, Ude92d, VN90e]. Lives [Nad90f, Ano90t, Ano91-92, Ano92-93]. Living [FT90, Min90e, Ano92-198, Min91b, Smi90f]. living-color [Smi90f]. load [Ano94-159]. loaded [Ano91e, Ano91-197, Mia91]. Local [Ano92-94, Ano92-196, Bar93c, Ken94, Ano92-65, Ano92e, Ano92-169, Ano92-222, Ano93-85, Ano94-166, Api91b, Ess94, Fox93, Van94a]. local-bus [Ano92-65, Ano92-94, Ano93-85, Ess94, Fox93]. Localizing [Tab93, Ano94-162]. locally [Ano92-213]. LocalTalk [Egl94d]. Locating [Vos92]. location [MH91]. locations [ES94]. Locke [Loc91]. Lockheed [Wal94a]. locking [Api94c]. locks [Gre94b]. log [NTS93c, NTS93c]. Log-Ins [NTS93c]. loggers [Smi93b]. Logic [Bar93d, She93c, UPSW94, Way94i]. logical [Lu92]. Logistics [Pou91n]. Logitech [Ano90-89, Ano90-90, Kli90a]. Long [Ano91-196, Gea90, Ano91-37, Ano91-137, Ano91-151, Ano91-195, Ano93-138, Ano94w, Ano94c, BL90, HC93, McN94, Ude90g, Ude93i, Ude94l, Ano92-177]. long-awaited [Ano91-137, Ano91-151, Ude90g, Ude93i, Ude94l]. long-delayed [Ano94w]. Long-Distance [Ano92-177]. long-established [McN94]. long-promised [Ano91-37]. long-sought [BL90]. long-term [Ano91-195, HC93]. Longer [Len94a, Ano92-32, Ebe93, Fri94b, Min90c]. Look [Min90b, Pou94i, Alf91a, Ano90x, Ano90-41, Ano90c, Ano91-92, Ano91-160, Ano91-173, Ano92-44, Ano93-57, Ano94w, Ben91, BL94, CV90c, Cra90j, Cra92a, DB93, DE90, EB90a, Fie90l, Fie91d, Fie91i, Fin90a, GR94, Gla90f, Gla90g, Gre90e, GTGW94, GTN94, Gre94a, Gut93, Hal94d, HR90, Hyo94, Joc92b, Kay94c, KGTD92, KPUG93, Mal91a, McA90, Mia90i, Mil94a, Nan91b, Nov94a, Oth92, Pou91d, Pou94a, Pou90k, Pou90u, Pou90v, Ras90e, Ras90p, RT91, Sma93b, Smi90f, Smi94d, Ude94b, Vos90b, Way91d, Way93d, Yag92d, Yag92a, Yag92b]. look-and-feel [Ano90-41]. looked [Ano94-79]. Looking [Ano92-95, Fie90e, Nic90, Sca92, BGE^{+}92, Gre94d, Yag90h]. Looks [Kro90, Wal93, Ano90-84, Ano91-194, Ano92-173, Ano92-189, Ano94w, Cal93, CS92, Cra91a, Egl94b, Gar94, Gla90d, Gre90d, Gre93c, Kay94c, Min90a, Nan91j, Pou90o, Pou90s, Pou90t, Pou90w, Pou91o, Pou91j, Pou91k, Pou91m, Pou91n, Pou91p, Pou91q, Pou92e, Pou92i, Pou92o, Pou94o, Tho94e, Yag92e]. looms [Sch93a]. loopback [Ano92s]. Loose [Ano94g]. lose [Ano91-77]. losers [Ano91-164]. Losing [Ano91-70, Ber93]. Loss [Ano92-126, Ano94-43]. Lossless [Api91d]. Lost [Fie91f, Ano92-160]. lot [Ano91-38, Ano92-193, Ano92-199, Ano93i, Ano94-166, Min90e, Nan92e, Rie91, Ude90a, Wau93b]. Lotus [Ano90-51, Ano90-52, Ano90-110, Ano90-113, Ano91-149, Ano93-61, Ano94-84, Ano94a, Bon90a, Del93a, Del93b, Die90b, ES94, Far90, Gas93, Lip90a, Pep91, Ref94, She93a, Ude93i]. love [Ano91-185, Red93]. Loveria [Lov94]. Low [Ano91l, Ano91-122, Ano94-106, Cla90b, GAM90, Rei92b, Tho94a, UM90, Ano90z, Ano90a, Ano90-67, Ano91-37, Ano91-197, Ano92z, Ano92-65, Ano92-72, Ano92-87, Ano92-183, Ano93-50, Ano93-70, Ano94-32, Ano94-54, Ano94-112, Bry94a, Cla94b, Del93a, Egl93c, Egl94c, GW92, Hay93, Lab92a, Lin91a, Nad90e, Nan93b, Rya93b, Smi94c, Smi94g, Tho90h, Tho93f, TR94a, VC90e, Way90b, Yag90c, Yag90m, Zei93, Ano94-107, Ano94-108, Api90c, Bar90c, Dio94, ESN94, Joc90b, Lov91b, LS91a, Per94a, Ude90c]. Low-Cost [Ano91l, Ano91-122, Ano94-106, Tho94a, UM90, Ano90z, Ano91-37, Ano92-65, Ano92-72, Ano93-50, Ano93-70, Ano94-54, Bry94a, Egl94c, Lab92a, Lin91a, Tho90h, Yag90c, Api90c, Bar90c, Dio94, ESN94, Joc90b, Lov91b, Ude90c]. low-end [Ano94-32, Del93a, Nan93b, Yag90m, Ano94-107, LS91a]. low-overhead [Ano92-183]. low-power [Ano91-197, Ano94-108]. low-price [Rya93b, Smi94c]. Low-priced [Rei92b, Per94a]. low-rad [Ano91d]. low-speed [Cla94b]. lowcost [Joc92a]. Lower [Ano91-79, Hal93e, She92, BL90, Hol94, LB90]. Lower-Cost [Ano91-79]. Lowers [Lov91d]. lows [Ano91-94]. LS [Ano92o]. LT [Ano91-94, Ano93-132]. LT-20 [Ano91-94]. LTE [Egl92c]. Lucid [Ano92-150]. Lucky [Han92b]. luggable [Api90a]. lunacy [Ano91-102]. lunch [Fie90d]. Lunchbox [VC90i]. luxury [Fri94b]. LX [Ano93-109, Ano93-127].

M [Ano92t]. M1 [Rya94b]. Mac [Ano90-45, Cal91a, Cra90b, Ga 93a, Ano90-28, Ano90-45, Ano90-55, Ano90-77, Ano90-103, Ano90-90, Ano91i, Ano91f, Ano91-33, Ano91-29, Ano91-37, Ano91-52, Ano91-64, Ano91-73, Ano91-74, Ano91-88, Ano91-151, Ano91-167, Ano92m, Ano92-46, Ano92-54, Ano92-66, Ano92-73, Ano92-84, Ano92-95, Ano92-87, Ano92-85, Ano92-96, Ano92-99, Ano92-97, Ano92-109, Ano92-98, Ano92-100, Ano92-120, Ano92-141, Ano92-171, Ano92-175, Ano92-191, Ano92-192, Ano92-184, Ano92-185, Ano93-43, Ano93-58, Ano93-130, Ano93-138, Ano93-134, Ano94h, Ano94n, Ano94k, Ano94l, Ano94m, Ano94-45, Ano94-94, Ano94-107, Ano94-90, AWGY90, Api94b, Bar90a, Bar93b, Cal91a, Cal91b, Cot93b, Cra90a, Cra90e, Cra90f, Cra90b, Cra90i, Cra90m, Cra90l, Cra90p, Cra90r, Cra91a, Cra91e, Cra91c, Cra91d, Cra91f, Cra91g, Cra91i, Cra91j, Cra91k, Cra92a, Cra92b, Cry94, DB93, Die90e, Die91, DE92, Die93c, EWT90]. Mac [ET91, Egl93d, Egl94e, Eug90, GE92, GAW93, GFK94, Ga 94b, Ga 94c, Gre90c, Gre90i, Gre93e, Gre94c, HM94, Ken93b, Kof93, LM93, Loe90, Mas90a, Mia90b, Nan91a, NTS93c, Par92, Pou90k, Pou91k, Pou92h, Ras90p, RT91, ST93, Sty91, The94, Tho90a, Tho90b, Tho90d, Tho90c, Tho90g, Tho91c, Tho91d, Tho92f, Tho92c, Tho93a, TGN94, Tho94c, Tho94e, Tho94d, Tho94f, Tho94g, Tho94i, Tho94j, Van93c, VN91e, Way90b, Yag90h, Yag91g, Yag93g, Yag93h]. Mac-ish [Yag90h]. Mac/Windows [Cra90r]. MacFolk [Cra91e]. Macguffin [Ano92-194]. Machine [Ano91-76, Lin93, Ung90a, Vas93b, Ano90-111, Ano91-41, Ano91-42, Ano92-74, Ano93-51, Bar90f, Bra90, Cla90a, GK91, Hov93a, Mil93, MG90, Pou90c, Way93i, Ano93-101, Lin93]. machine-translation [Mil93, Ano93-101]. machinery [Pou94l]. Machines [Alf91d, Ano94-167, Pet90, VC90a, Ano90x, Ano90-60, Ano91-48, AD91, Api93b, BL90, Gre90e, Gre93d, Hay94b, Kea90, Nad90c, Smi94e, VC90g, Pou91g]. MacinStor [Ano92-98]. Macintosh [Ano90-93, Ano90-111, Ano92-87, Ano92-99, Ano92-192, Ano94k, Ano94m, Ano94-41, Ano94-90, Ano94-109, Cot93a, Cra90n, Cra91a, Cra91e, Cra91l, Ga94, GFK94, Ga 94c, Gib92, Gre90j, Gre93g, Hal93b, KGTD92, Lab92a, McN94, Ras90p, RT91, Smi94a, Tho90i, Tho91c, Tho92h, Tho93c, Tho93h, Tho94e, Ude90e, Ude93i, VC90t, Yag91g]. Macintoshes [Ano91-121, TH93]. MacRenderMan [Cal91b]. MacroMind [Tho90b, Yag91a]. Macs [Ano92-100, Ano94-183, Cra92a, Ano90g, Ano91-121, Ano91-184, Ano92-199, Ano93o, Ano94-107, Ano94-180, Die90e, Die93c, Egl94d, Gre90e, Gre93e, GW93, LB90, STA92, Smi94c, Smi94e, STS92, TG91, Tho92a, TH92, Tho92d, Tho93a, TH93, Tho94h, Woo91b, Yag93e]. MacSleuth [UT91]. Macsyma [Way90c]. Made [Ano91-85, Ano94-110, Rya90g, Smi94b, Ano90-74, Bry93a, Gla91d, Hel91c]. Madness [Ano92-103]. Magazines [Ano93-62, Ano94q]. Magellan [Ano90-52].

Magic [Ano94-85, Ano94-113, Cot93b, Ano91-157, Ano94g, Ano94-85]. Magnavox [Ano91-80]. magnetic [HR90, McG90a, Rya90e]. magneto [Ano92-45, CP94]. magneto-optical [Ano92-45, CP94]. Magnum [Yag90m]. Mail [VC90e, Ano93-60, Ano94-77, Die93b, ES93, Fie90k, Fie91b, Ga 93b, Ano91-43, Ano93v, Ano94-30, Ano94-53, ES94, Fie91b, Rei93d, Ude93e, VN90e, Ano92-120, Ano93-58, Ano94-98, Ano94-145, ET91, Ga 93b, Sta94a]. Mainframe [Ano90-80, Bor93, Way90c, Don92, Fis90d]. Mainframe-type [Bor93]. mainframe/minicomputer [Don92]. mainframes [And93, Ras90j]. mainland [Rei93a]. mainline [Hay94b]. Mainstay [Ga 94c]. Mainstream [Ano94-34, Ano94-51, Ano94-54, Ano94-129, Rei93b, Ano92h, Ano92d, Ano92-109, Die94, Rei94e]. maintain [Bry94d, New93]. maintains [MS93]. Maintenance [Pou90p, Ano92-28]. Major [Ano91-54, Ano93-61, Ano94-159, Egl94c, Gas93, Law94, MH91, Pou94i, Ras91g, Rei94b, Slo91]. Make [Ano94-54, Bry94d, Fie90g, Kea91b, KGEM93, Kor93, Ras91k, Red92a, Ano91-54, Ano91-107, Ano91-149, Ano92-50, Ano92-71, Ano92-82, Ano92-154, Ano92-157, Ano92-189, Ano92-204, Ano93-43, Ano94-63, Ano94-68, Ano94-169, Api90c, BS91, BGE^{+}92, Bar90i, Die94, DH90, Egl91a, Egl91e, Egl92c, Fie91h, Fis90c, Ga 93b, Hal94i, Har91c, HWS91, Ken90, Laf91, Lin91c, Mil94b, Mul94, Ras90f, Ras90g, Ras90k, Ras91i, Ras92d, Rei94e, Rya90b, Rya91a, Spi90a, Tho93f, Ude93l, Ude94k, Woo90a, Woo91a, Yag91d]. make-over [Ano93-43]. Maker [Smi90f]. Makers [Ano90-58, Ano93-75, Rob90b]. Makes [Ano93n, Ano93-40, San90b, Yag91f, Ano90-103, Ano91-111, Ano92-95, Ano93-76, BF90, Bos93, Cla90b, Cla91a, Cra91l, Dav94a, Egl91c, Egl93b, Ga 94b, LB91, Lov91d, Mia92a, Nan94d, Tho93e, Tho94f, Ude90b, Ude90e, Yag90e, Yag92c, Yag93a, Yag93e, Yag93i]. Making [Ano90-53, Ano91-81, Ano92-101, Ano92-125, Ano92-148, Cra91f, Fis90d, Gre93c, Lan94b, Loc92, Lov93b, LS91b, Ras92a, Sch91b, VN91e, Ano91-157, Ano94-76, Bar92d, Rya91b, She91]. Man [Ano90-69, Pou90u, Ras91h, Ano90-27, Ano90-86, Lip90a]. manage [Ano92-28, Ano92-182, Ano92-222, Hol93b, Nan92d, Rei94e, Yag90i]. Manageable [VN91e, Mil94b]. Management [Ano91-100, Ano91-101, Ano91-108, Ano91-128, Ano92-195, Ano93n, Ano94-49, Ano94-147, Day90, Hol93b, Ude91e, Ano92-186, Ano93-127, CS92, Fie91b, GD92a, Gre90h, Hel90c, Nan91b, Ras91f, Ras92f, SU93, Sta90, Ste91e, Top91, Ude93f, VC90l, Yah92]. Manager [Ano92-85, Ano92-182, Ano94-111, Api94b, CV90c, Gre92, Nan90b, SL92, Ude90d, Ano91p, Ano92-118, Ano93d, Cla93a, Nan92d, Ras91i, STS92, VN90c, Woo90b, Ano90-85, Ano93d, Cot93b, Min90d, Min90f]. Managers [Ano92-157, Ano92-165, Egl91e, FW93, VC90l, Ano91-162, Ano92-137, Ano94x, Bar93b, Bon90b, Rya91e]. Managing [Ano91-82, CS92, Cra92b, Hal94g, MH91, Min90f, Nan91b, Ras91f, Rei94e, Rei94d, Sta90, VC90r, Vel90, Wal94c, Ano91-61, Ano92-180, Ano94-147, HC93, Nan91f, VC90c]. manipulate [McA90, She90b]. manipulation [Ano90-76, Ano92-68]. manipulator [Ano92-173, Ano93-125]. manner [Ano92-130, Mul94]. Manor [Pou90m, Pou90r, Pou92j, Pou92n, Pou92k, Pou93j, Pou93l, Pou94l, Pou94q, Pou94s]. MANs [Ano91-34]. manufacturers [Ano90-31, Ano91-142, Ano93g, Ano94-159]. Many [Ano93-63, Ano94-58, Ano94-150, Cla93b, Mil93, Var94c, Ano90u, Ano90-87, Ano94-116, Ano94a, Bar92b, Chr94, Daw94c, Mil94a, Mos93, Pac94, Smi94e, Sta91a, TR94a, Way93i]. Map [Ano94n, SE93b, Vor93]. MAPI [She94, Ude93c]. Mapped [STS92]. Maps [Vor93, Hel90c, KUD92, NTS92b, SE93b]. March [Way94f]. marginal [Way93i]. mark [UT91, Min90d]. Market [Ano91-95, Ano94-185, Cla93b, KCU^{+}93, Nie94, Tho91b, Ano91-31, Ano92v, Ano92-225, Ano93-75, Ano94s, Ano94-71, Cla93c, Cry92a, DW91, DE91a, Dys92, Egl94f, Ga 93b, Gas93, GD92b, Hal94c, Law94, Lov92b, Mia90a, Min92b, Smi91b, TH92, TR94b, Wal93, Yag90m]. market-leading [Gas93]. marketing [Ano91-183, Ano93-64, Ham91, Ano90-110]. marketplaces [Ano94-116]. Markets [All92a, Dys92, Egl93a, Gre94d]. marks [Ano90-60, Ano92-119, Ga 92b]. Markup [Ano92-181]. Marriage [Rya90g, Bos93, Gla91d, Hal94c]. Marries [BR93, Min90h]. Marvin [GTGW94]. Mass [Ano91-142, Wal94c, Ano90-54, Ano92-102, HR90, Par90]. mass-storage [Ano92-102]. Masses [Ano90-54, Gre90j, Spi90b, Ung91]. Massive [Ano91-142, AWGY90, Rya91b]. master [Ols94, Cra90d]. masterful [Ano90x]. Mastering [Ham90, Ano94-111, Ric91, WK90]. MastersPort [Ano91-108, Ano91-197]. Mastery [Ham90]. Match [Ano91-88, VC90n, Kea90, Ran90, Fis90b]. Matching [MW92a, MW92b, Ste91e, Vau90, BL94, Ano94-89, BL94]. Material [Ano92-146, Mur93, Pou94c]. Materializes [Lov91b]. Math [Ano91-83, CWVS94, Gla90f, Way90c, Way90b, Ano91-92, Ano92-216, Bis91]. MathCAD [Ano92-216, Ga 93c]. Mathematica [Ano91-84, Ano92-216]. Mathematical [Ano91-92, CWVS94]. mathematicians [Ano90t]. Matic [Ano91-48]. Matrix [Ano94-35, Ano93-126, Ano94-112, Bry94d, HM94, KM94, Ras92c, Sma93b, Yag93b].

Matrox [Ano93-74, Yag92i]. Matter [Ano93-139, Ga 92a, Pou90r]. maturing [SS94]. Max [Kli90c, Lan90c]. maximize [Alf91c]. maximum [Ano92-58, Kre91]. may [And90, Ano91-42, Ano92-127, Ano94-35, Ano94-34, Ano94-54, Ano94-60, Ano94-82, Ano94c, Ano94-166, Bar90i, Del93a, Egl90b, Egl94c, Gea90, Hal93d, Jof93, Mas91a, NHW93, Nie94, PP93, Ras90e, Rei94f, Rya90b, Spi90a, Tho93d, Van94a, VN90a, Wau93c, Yag90d, Yag90j]. MB [Ano93k, Ano93l, Ano93-47, Ano94-126, Api94d, Mia90d]. MBps [Ano93-93, Ano94-165, Bry93b, Nan94b]. MC33 [Alf90c]. MCA [Ano91-131]. McAfee [Ano90-100]. MCC [Ano91-67]. McGraw [Ano94q]. McGraw-Hill [Ano94q]. Me [Ano90-97, Cra91j, Fie91e, Ano91-115]. Mean [Ano91-76, Ano92-119, Ano94h, Ano94-126, Hal93e, Ano91-191, Api94e, Lin90a, SL92]. meaner [Ano91-117]. meaning [Ano90-106, Web93]. Means [Ano94-172, Ano93-91, Ude94h]. Measure [SB94, Hur93a, Pou93g]. Mechanisms [Alf92f]. Media [Ano92-45, Har90c, HR90, Smi94c, Ano94-77, LDH94, Pou94c]. MediaMaker [Yag91a]. MediaTracks [Ano90-55]. medical [Ano94-101, Buc92]. medium [Ano93-108, Mel93a, RP90, Sta93]. Medley [Pou92i]. meet [Laz93, Ras90a]. Meets [Hel91f, Lov91c, Wsz93a, Yag92i, Min91a, Mit93b, Rei93b, Ryd93, Yag91b]. megabyte [Gre90i, Mia90d, Tho90d, Ano90a]. megabytes [Ano91-61, Wsz90]. Megafloppies [Sat90]. Melange [Cra90l]. Melding [Lin91d]. Mellon [LHR90]. member [TR94b]. members [Ano93z]. MemMaker [SSF93]. memo [Ano92-108]. Memories [Psa92, Rei91a, Way91c]. Memory [Ano90-57, Ano91-162, Ano91-168, Ano92-102, Ano92-157, Ano94-111, Gre90h, Min91a, Bri90, Bry94d, Gla91a, Hol90, LM90, Min90c, NTS93a, Pou93i, Ras91i, Rei91a, RH93, Rob91a, RNG94, Ste91c, Ude93f, VC90c, VN91a, Way91c, Woo90a, Wsz90, RNG94]. Memory-Card [Ano94-111]. Men [Ano90-66, Pet90]. mentality [Ano91-105]. Mention [Ano94-88, HK94]. Mentioned [Ano90-58]. Mentions [Ano93-34, Ano93-48, Ano94-89, Ano94-90, Ano94-87, Fox93, HM94, HHP93, Hol94, HP94, Len94b, Len94a]. Menus [Ano92-129, Ano92-192]. merely [Ano91-183]. merge [Ano94m, Fie90b]. merger [Ano91-58]. Merlin [HC93]. Message [Ste91c, Kor93, Nan94a]. Message-passing [Ste91c, Kor93]. Messages [Ano91-43, Gre91a, Rei93d, Way93c]. Messaging [ES93, Ude93c]. Metalis [Ano91-80]. Metalis/286 [Ano91-80]. metaphor [Ano93-111]. MetaWare [Gre93b, Hel92a]. method [Gla90d]. methodology [Wal94a]. Methods [Ano92-103, Ano94-56, Fie91k, Ano92-117, Ano92-180, Gre90g, Hel91a, Ste94b]. Metrowerk [Ga94]. Mexico [Bar93c]. MHz [Alf90c, Ano90d, Ano91b, Ano91-170, Ano92-78, Ano92e, Ano92-111, Ano94-129, AMW90, AD91, Api92b, Api94d, Bon90c, Egl90b, EW92, FJPR93, HK94, HP94, MG90, MM94, Nad90c, San90a, Tho90c, Tho92h, VC90d]. Mice [Kli90a, Gre90c]. Micro [Ano91-85, She90a, Ano92-149, Var94b, Alf90c, Alf92b, Ano92-198, Fis90d, Mia90a, Pou90b, San90a, Tho94c]. micro-based [Ano92-149]. micro-kernel-based [Var94b]. Micro-OCP [Pou90b]. Micro-to-Mainframe [Fis90d]. Microboards [Tho94c]. Microbytes [Ano92-104, Ano92-105, Ano92-111, Ano92-107, Ano92-108, Ano92-112, Ano92-109, Ano92-110, Ano93-64, Ano93-68, Ano93-65, Ano93-66, Ano93-67, Ano92-106]. Microchip [Ano92-161]. Microcom [Nan92d]. microcomputer [Ano90-112, Ano91-187, Ano92-161, Bra90, Lan90b, Ude93g]. microcomputers [Ano92-30, VC90t]. Microelectronics [Ano93-45]. Micrografx [Ros90b, Vor93]. Microkernel [Ano94-127, Pou94a, Var94a, Pou94a]. Microkernel-based [Ano94-127, Pou94a]. MicroNet [Gre90i]. Micronics [Tho93b]. MicroPaq [BS91]. MicroPhone [Ano92-171, Ano93-107]. Microprocessor [Ano92x, Pou94f, Tho90f, Ano91-85, Ano92-106, Pou91a]. Microprocessors [Ano92-105, Ano91-153, Ran90]. Microrim [Ano90-87]. Micros [Man90, Ras90j, Mas91a]. Microscope [Yag90c]. Microsoft [Ano90-56, Ano91-86, Ano91-132, Ano91-192, Ano92-85, Ano92-108, Ano92-112, Ano92-113, Ano92-145, Ano92-225, Ano93z, Ano93-46, Ano93-68, Ano93-69, Ano94h, Ano94n, Ano94w, Ano94-32, Ano94-33, Ano94-84, Ano94-114, Ano94-117, Car94, CV90c, Cry92b, Del93b, Egl93a, ES94, Ga 94d, Gas93, Gre90d, Gre93b, Hal93f, Hal90, Han92b, Har93, Hur93b, Jef93b, KPUG93, Mat94a, MS93, Pou93s, Ref94, Rei91c, Rei94g, Sch92a, Sha94b, She93b, Smi90e, Ude90g, Ude90d, Ude92a, Ude92d, Ude93c, Ude93f, Ude93l, Ude94j, Url90, Var94a, Wig90, Woo90d, Yag93d, Yag93k]. Microsoft/IBM [Ano91-132]. MicroSolutions [Ano90i]. MicroSparc [Api94d, Gar94]. MicroSpeed [Kli90a]. Microsystems [Ano92-56, Api91c, YS90]. Microtek [Egl93b, Van94a]. Microtest [ATU92, Bos93]. microWriter [Ano93-70]. middle [WE92]. Middleware [Kor93]. Mideast [Red92b]. Midrange [Tho93a, YS90]. Miffs [Ano93-46]. Might [Ano93-142, Ano92-40, Api94e, Chi92, Cro93b, Mur93, Ude93l, Way93f, Way94d]. migrate [MS93]. Migrates [Bis91]. migrating [Ras90j]. Migration [DH90, Moo90, Ude93h]. miles [Van94b]. Million [Ano91-131, Tho90e]. millions [Hal94d]. Mind [Ano91r, KAY93, Lin91d, Ras90p, Ano90-66, Ano91s, Ano91-41, Ano91-35]. mind-machine [Ano91-41]. mind-set [Ano91-35]. Minds [Ano90-66]. Mine [Cro93b, Loc92]. minichanger [Egl93e]. minicomputer [Don92]. minicomputers [And93]. MiniDisc [Ano93-108]. minimalist [Pou94p]. minimizing [Ano92-102]. Minimum [Tho90d]. minions [Han92b]. MinisPort [Mit90b]. Minix [Yag90c]. Minsky [GTGW94]. Minstrel [Pou90a]. minuses [Cra90i]. Minute [Gre93g]. MIPS [Yag90m, Ano91-87, Ano92-57, Ano93-71, Gre93d, Hal94h]. Mirror [Ano91-125, Ano92-114]. Mirrors [Bux90b, Bux90a, Way92b]. MIS [Ano94-94, Rya91c]. Misers [Len94a]. mishap [Pou90k]. mishmash [Ano94-85]. mismatch [Cut91]. miss [UT91]. Missile [Wal94a]. Missing [Min91a, Per93a, Ras91g, VC90q]. Mister [ET91]. Mitch [Ano90-50]. Mix [Ano91-88, Ano92-142, DWN91, Ga 93e, Ras90b, Rei94b, Pou94a, Fis90b]. Mix-and-Match [Fis90b]. Mixed [Ano92-132, ES93, Ano90-64, Ano90-111, Ano92-45]. mixed-machine [Ano90-111]. mixedmedia [Die93c]. mixes [Sta93]. Mixing [Ste91e]. MKS [Smi93a]. MLCP [Ano92-64]. MLCP-4 [Ano92-64]. MLCP-4/AEB [Ano92-64]. MMPM [Jef93b]. MMPM/2 [Jef93b]. MO [Ano92-45]. Mobile [Ano93-106, Len94a, MW93b, MW93a, Ano92-209, Ano94-88, Api94d, Rei94f]. Moby [Mia91]. Mode [Gla90e, Gla90b, Hel92b, Ano92-44, Gre90k, Orr94, Sch90b]. Model [Ano90w, Ano90-90, Ano92-153, Mit91b, Nad90a, Nan93f, Ano94-71, Gib90b, Rei94f, Sch92a, Ude93m, Gre93h, Smi94g]. modeling [Ano93-130, CWVS94, Wal94e].

Models

[CWVS94, Cla90b, KM94, Ref94, Woo91a]. Modem [Ano93g, Gib90a, Mar92c, Tur90, Ano90-92, Ano91-33, Ano91-135, Ano92r, Daw94a, Jof93]. Modems [Ano91-89, Ano93g, ATU92, Hol94, Hur93a, RM92, Ano93-128, Ano93-134, Fie91c, Mar92c, Mia94b]. Modern [Ano94-64, Coo90a, Lan94b, Tur90, Way93d]. moderns [Egl94b]. Modest [Pou93a]. modify [Kor93]. modifying [Sin94]. Modula [Fou91, Har90a, Pou90s]. Modula-2 [Har90a, Pou90s]. Modula-3 [Har90a]. Modular [Ano91-28, Mia90a, Mit93b, Alf92b, Ano94-117, Ano91-90]. Modular-CPU [Ano91-90]. Modularity [Ano91-183]. modules [Hal93b, PS94, Van93a]. Momenta [Ano91-91]. Moments [Ano90c]. Money [KCU^{+}93, Min90g, Way90a, Egl94g, Rei94a, Rya91c, She93b, She93b]. Monitor [Ano94-89, Ano90-82, Ano90-77, Ano91-53, Ano91-125, Ano91d, Ano92-68]. monitoring [Ano92-190, NTS93a]. Monitors [Ano92-27, Ano94-89, Fin90b, McG90a, Ano92-121, Api94a, Cra90f, Lin93, Lov90a]. Monochrome [Ano92-115]. Monolithic [BS91]. Monoliths [Ano94-123]. Monopoly [Ano91-44, Pou94f]. Monster [Ung91]. month [Ano91-126, Ano91-196, Pou93f, Pou94q, Pou94r]. months [Ano94w]. morning [Ano94p]. morphing [Nov94a, Pou94o]. MOs [CW92]. Mosaic [Ano90-110]. Moscow [Pou90u, Ano91-93, Fri92]. Most [Ano91-89, Nan91h, Ano90-112, Ano91-42, Ano91-85, Ano91a, Ano92-30, Ano94-148, Egl90b, Fie90a, Fie90c, Gre90g, Gre94d, Len93a, Orr94, Rei94d, Smi90g, Smi93d, Smi94a, Tho92h, Ude94e, Ude94j, Vos90a]. motherboards [AMW90, Ga 93e]. Motif [Ano92-109, Ano94-112, Api94b]. Motion [Gre94b, Ros90a, Bar90b, Cry94]. Motorola [Ano93-66, Ano94-113, Egl91b, RT94a, Tho90f]. mounds [Ano92-181]. mount [Ber93]. mountains [Ano92-95]. mounted [Ano90-82, Ano94c, FT90]. Mouse [Alf90a, Mia90b, Ano94-175, Bar93a, Nag93, Kli90a]. mouse-digitizer [Mia90b]. Mouseball [Ano90-86]. Mouth [Ano90-29]. Move [Ano92-31, Nan92c, Red92a, Ano90-45, Fis90c, Joc92a, Rob90b, Smi93e, Tho94f, Way94g]. Moves [Ano94-34, Kea91a, Ras92b, Ras92d, Way94e, Cra91k, Pou93s]. movie [Yag93h]. Movies [DB93, Ebe93, Ano92-206]. Moving [Ano92-133, Ano93-72, Ras90j, Yag92e, ASS92, And93, Gre94c, Mil94a, Yav94]. MP [Ano93-40, Nud91]. MPC [Lov93b, Yag92a]. MPEG [Way94d]. MPS [Mia90a]. MR [Ano92-153]. MS [Ano90u, Ano92t, Ano93-38, Ano93-53, Gre90e, Gre90h, Lan90c, Sch91a, SSF93, Ude93f, Woo90a]. MS-DOS [Ano92t, Ano93-38, Ano93-53, Gre90e, Gre90h, Lan90c, Sch91a, SSF93, Ude93f, Woo90a]. MT [Hov93a, Hov93b, Vas93a]. MT735 [Ano91-94]. much [Ano91-149, Ano92n, Ano93z, May94b, Sha94c, VC90p, YS92]. much-needed [Ano91-149]. Multi [Wal93]. Multi-Lingual [Wal93]. MultiAccess [Ung91]. multicomputers [Ste91c]. Multidimensional [Tre93, Ano93-133]. multifunction [Dav94a]. multilevel [Gre94b]. Multilingual [Ben91]. multilink [Pou91m]. multimasking [Nan93c]. Multimedia [All93j, Ano90p, Ano90-58, Ano91w, Ano91-86, Ano91-95, Ano91-143, Ano92-101, Ano92-183, Ano92-222, Ano93p, Ano93s, Ano93-39, Ano94-54, Cra90m, Cry94, Gut94, Jef93b, Len94a, LK90, Pou90s, Pou92i, Ras92b, Rob90a, She90b, Tho90b, Vel90, Yag91a, Yag91d, Yag92a, Yag92b, Yag92j, Yag93f, Ano91-36, Ano91-68, Ano91-66, Ano91-158, Ano92f, Ano92-107, Ano92-158, Ano93x, Ano93-126, Ano94-152, Ano94-180, Har93, Hed93b, KPC93, KAY93, Lin94, Lip90a, Lov90b, Lov93b, Pou91o, RNG94, Sch93b, Way94j]. multimedia-ready [Ano94-180]. Multiplatform [Ano93t, ES93, Ga 94a, Ano90-79, Ano92-141, Ano92-203, Api94b, Ga 93b, Mil94b, Smi94a]. Multiple [Ano94-87, AWGY90, Hay94a, Ols94, Vor93, Ano91-51, Ano92-89, Ano94-37, Api94b, Api94c, Fri94c, Hay94b, Hol93a, Mia90i, Min90h, Nan93h, Tho93f, Van93a]. multiple-platform [Ano91-51]. Multiprocessing [Ano91-96, Ano91-144, Ung91, Kea90, Kog94, Nud91]. Multiprocessor [Ano92-162, Ga 94d, Pou93h, Rya91d, GR90, VN91a]. Multiprotocol [Smi94e]. MultiScope [Kea91b, Ude91c]. multistage [Bry94d]. MultiSyncs [Ano92-121]. Multitasking [Ano91-35, Gre90j, Ano92-132, Ga 93b]. Multithreaded [Ano92-80, Ano92-81]. Multiuser [Fin90a, Ude91b, Ano92-187, Ano92-182, Die93c, Gre94b, Sch93b, Ude90b, Yag90b, Yag90g]. MultiView [Ano92-116]. Munich [Lin93]. Muscle [Ano92-88, Smi91a]. Muscles [Lov90a]. Muse [Ano90-59, Ano92-95]. Museum [Ano90-60]. Muss [Joc90b]. Must [Ano90-88, Ano92-215, Hay94b, Hed94, Lan94a, Mul94, Rei94c, VC90s]. MVP [Gut94]. MX [Ano93-36]. My [Ano91-77, Fie90g, Fie92a, Rei91a]. Mylex [Ano90-61]. Mysteries [Pou92k, Pou93o]. mysterious [Cri93]. mystery [Pou92k]. myth [Sha94a].

n [Ano91-88]. N40 [Daw94c]. Name [Nad90f, Nud91, Ano92-118, Ano92b, Ano93-94]. nameable [Ano94a]. Named [Tho94k]. naming [SH91]. Nanao [Ano91-97, Ano91d, Api94a]. Nance [Nan91j]. nanotechnology [Ken93b, Ken93a]. Narcissists [Ano91-105]. narrow [Ano91z]. NAS [Osm90]. Nasties [Ano90-100]. Nation [WB90]. National [Ano91-98, Ano91-181, Ano93-93, Fis91b, Pou94c]. Native [Nan90b, Tho94k]. Natural [Ano92-79, Ano92-117, Dys92, Obe90, VC90n, Ano90-59, Ano93-82, Pie94, SG92a]. Natural-language [Obe90, Ano90-59, Ano92-117]. Nature [Ste91b]. Navigating [Rib93]. navigational [Ano90-27]. NB [Ano91-166]. NBI [Ano90-49]. NC100 [Ano92k]. NCR [Alf90c, Ano90d, Ras91j, Rei91b]. nd [Hur93a]. NDIS [Fis90b]. Near [Cra90k, Sac90, Ano91-172, Ano92-48, Ano92-210, Cra90o, Mas91a, Mia92b, Pou94f, Way93f, Way94h, Yag91g]. near-laser-quality [Mia92b]. near-seamless [Yag91g]. nearing [Ano92-112]. nearly [Ano90-70, Ano91-157, Hal94i, Ude94h, VN90d]. NEC [Ano92-94, Ano92-121, Api92b, Nad90d, Ras92c, VC90b, Yag91b]. necessarily [Api94c]. necessary [ES94, Rei91a]. necessities [Mia90f]. necessity [VN91b]. Need [Ano94-87, Fis90c, Man90, Ano91-98, Ano91-183, Bar92a, Egl90b, HR91, Hog90, Kiy93a, Mat94b, Ras90e, Ras90f, Rei94g, Ste91d, Ude94f, Van94a, Way94g, Yag92h]. needed [Ano91w, Ano91-149, PS94, Pou93n, Rei94d, Tho93f, Ude93i]. Needless [Mos93]. Needs [Ano91-188, Ude93m, Ano92-31, Fie91m, FJK^{+}93, Jof93, Laz93, Ras90a, Ras90i, Ras91k, RT94b, VC90j]. neighborhood [Hal94e]. neighborhoods [Ano94-96]. neither [Pou94b]. Nestor [Ano94-81]. Net [Cau93, Fie91a, Ano94-81, Nan94d]. NetBIOS [Nan91h, Nan94a]. NetBlazer [Daw94a]. NetDDE [Ude93b]. Netherlands [Pou90h]. NetManage [Nan93g]. NetModem [ATU92, Gib90a]. NetModem/E [ATU92]. NetMounter [Ano92-199]. NetOctopus [Ano92-118]. NetRunner [Ano92-217]. NetWare [Ano90-62, Ano90-83, Ano91-88, Ano92-85, Ano92-119, Ano93-64, Ano93-73, Ano94-75, Ano94-149, AWGY90, Bos93, Gre94e, HC93, Ken93a, Nan91c, Nan91l, NTS93c, NTS93d, Ryd93, SL92, Ude90b, Ude90d, Ude90f, Ude93h, VC90q]. Network [Ano91-54, Ano91-99, Ano91-100, Ano91-101, Ano91-122, Ano91-148, Ano92-120, Ano93-48, Ano93-95, Ano94-77, Ano94-147, Ano94-165, ATU92, Bon90a, CV90c, Fie92a, Fis90b, GAW93, Lov92b, Mia94b, Pou90l, Ras91j, Rei94f, Smi94f, Ude91e, VC90l, VC90o, Ano90-44, Ano91-75, Ano91-77, Ano91-141, Ano92n, Ano92-88, Ano92-118, Ano92-189, Ano92-190, Ano92-201, Ano92-199, Ano92-217, Ano93-36, Ano93-73, Ano93-127, Ano94-57, Ano94-150, Api91b, Api94d, Cau93, Cla93a, Dau91, Egl94e, Fie90k, Fow91, GP90, Gib90a, Gre91a, Gut93, Hay93, IC94, Kor93, LS91b, Min90j, Nan91e, Nan91k, Nan91b, Pou93e, Pou94o, Ras91k, Ras91n, Rya91e, Smi94b, Sta90, Ste91d, Ste91e, Top91, Ude90d, Ude91a, UFC94, Yag90f, Fis91b, Hay93]. network-support [Nan91e]. networked [Ano94-182, DeH93]. Networking [Ano91-190, Ano92-209, Ano93-93, Cla93a, Daw94a, Egl91b, Egl94d, Fie91g, Fis91c, LS91b, Nan91f, Tho91e, Ude91d, Ude94d, Ano90-53, Ano92-108, Ano94q, Ano94-88, Ano94-165, Ken93b, Smi94f, Ude92c, Ude93h, VC90m, Yag91h]. Networks [Ano91-147, Ano92-88, Ano92-168, Ano92-194, Ano92-190, Ano93-89, Ano93-103, Ano93-79, Ano94-37, Cau91, Day90, Fie90k, Fow91, Gib90a, Gla91d, Sta91b, UM90, Ude91e, VC90p, Ano92-86, Ano92-182, Ano93-48, Ano93-93, Ano93-129, Ano94-85, CV90a, Cla93d, Daw94b, Dej94, Die90e, ES93, Fri94b, GAW93, HC93, Lie93, Mel93a, Ras91b, Rei94f, SU93, Ste91e, Way94g, WS91, Ano92-199, Mia94b]. Networkwide [Mal91b, VN91f]. Networld [Ano94q]. Neural [Cau93, WS91, Ano91-141, Ano94-81, Lie93, Pou94o, Ano94-81]. Neural-Net [Ano94-81]. neural-network [Ano91-141, Cau93]. neutral [Nan93f, Pie94]. Never [Ano91-102, Ano94-90, Way94h]. Never-Never [Ano91-102]. new-generation [Ano91i, ADE90]. Newcomer [Nad90e, Cla93c]. Newest [Ano91-159, Egl92c, Mat94a, Ano91-51, Ano93c, Ano93-36, Ano94-117, Hal93f, Han92b, Hel90b, Tho93a, TR94b, Ude94j, Yag90d]. News [Ano93-137, Ano92-129, Ano92-92, Ano94-166, Yag90m, Ano91-154]. newsletters [Pou92m]. Newtek [Yag91f]. Newton [Pou93a, Ano93i, Ano93-67, Ano94-162, Bet94, PP93]. NewtonScript [Bet94]. NewWave [Ano90-63, Lus90]. NexGen [Ano94-118]. Next [Ano91f, Ano92d, Ano93-142, Api93c, Cra91e, Egl91d, Ano90-62, Ano90-97, Ano91k, Ano91-55, Ano91-133, Ano92-106, Ano92-161, Ano92-202, Ano93-73, Ano94w, Ano94-32, Ano94-82, Ano94-168, Api93b, Fie90e, Mia93b, Pou90j, Rie91, Sat90, Tho93f, Ude92d, Ano90q, Ano90-113, BL90, Car91a, TS90, Ano92-125, Ano93d, Sch93a, Smi93e]. Next-Generation [Api93c, Ano90-62, Ano91-133, Ano94-82, Ano94-168, Api93b, Mia93b, Ude92d, Ano92-125, Sch93a]. NeXTstation [Car91a]. NextStep [Smi93e, Web94, Yag92h]. NFS [Daw94b]. NIC [Ano94-119]. nice [Daw94c]. niche [Gre94d]. Niches [RARH92]. Nichols [VN91f]. Nick [Bar91b]. NICs [Ano93-48, Ano94-88, Ano94-119]. nifty [Ano91e]. Night [Lin94, Ano94p]. nightmare [Ano92-108, VC90r]. Niklaus [Pou91d]. Nine [Ano92-157, Api91b, ET91, Egl93a]. Nineties [Ano90-95, Bar90d, Bar90i]. Ninety [KM94]. Ninety-two [KM94]. Nintendo [Ken93b]. NLMerlin [HC93]. NLMs [Ano94-149]. No [Ano92-126, Ano93-138, Ano94e, Fri94b, Ras91j, Tho90h, Ano92-32, Ano92c, Ano93a, Ano93-97, Ano94-132, Ebe93, Min90c, Ano94-189, Joc90b, Len94b, VN91c]. No-Compromise [Len94b, VN91c]. no-font-hassle [Ano93a]. No-Fuss [Joc90b]. No-Muss [Joc90b]. NobleNet [Smi94b]. Noisy [Ano91-105]. Nokia [Api94a]. non [Yag91c]. non-Sun [Yag91c]. noninterlaced [Ano92-27]. nonprogrammers [Egl94g, Vel90]. nonroman [Smi90c]. nonstandard [Ano93g, Rai90]. nontraditional [Ano94-64]. nonvolatile [Wsz90]. noon [Ano94p]. norm [Ano92-80]. Northgate [Ano91-106]. Northwestern [Hed93b]. Norton [Ano92u, Ano90-64, Ano91-107, Tho90g, Wsz93c, YT90]. Note [Ano90-65]. Notebook [Ano91-108, Ano91-149, Ano91-166, Ano92l, Ano92-218, Egl92e, Gre93f, Gre93g, Joc92b, Lab92b, Lab92a, Mit91b, Nad90e, NW92, Ras92c, Sch90a, Tho93h, Tho94b, Ano90-107, Ano91e, Ano91-161, Ano91-197, Ano92g, Ano92-39, Ano92-199, Ano93-76, Ano93-91, Ano93-90, Ano93-130, Ano93-126, Daw94c, DWN91, Egl92c, Len94a, LAN92, Pou91g, Ras92d, Rei91a, Rei90c, Sma92, Sma93a, Tho93e, VN91c, Yag93b, Ano92-173, Ano92-169, Ano92-172, Ano92-170, Ano92-171, Nad90b]. notebook-size [DWN91]. notebook/printer [Ano93-76]. Notebooks [Api93b, Car92, Egl92c, Len94a, Nad91, Nan92a, Per93b, VN91c, Ano94-180, FP93, Len94b, NW92]. NoteJet [Ano93-76]. Notepad [Ano92k]. Notepads [Rei91b]. Notes [Ano94-125, Bon90a, Ude93i, Yav94]. nothing [Nan91f]. notion [Egl93a]. notions [Ano90-38]. novel [Ano90-69, LB91, Way91b]. Novell [Ano90-62, Ano92-108, Ano93-64, Bos93, Gre94e, Mat94a, Nan94d, Rei94g]. NPS [Smi94e]. NREN [Ano91-98, Fis91b]. NSTL [Ano91-110, Ano91-109, HH94a, Ref94]. NSX [Mit91b]. NSX/20 [Mit91b]. NT [Ano92-108, Ano92-112, Ano92-125, Ano93-77, Ano93-75, Ano93-110, Ano93-139, Ano93-135, Ano94-32, Ano94-75, Ano94-90, Ano94-153, Ano94-177, Ano94-188, Ga 94d, Gre93d, Gre93h, Gre94e, Hal94h, KAY93, Kiy93a, Nan92e, NHW93, Nie94, One91, Sch92a, Sha94b, Ude92b, Ude92d, Ude94b, Ude94k, Way94b]. NuBus [Ano94-90]. nuclear [Ano92-103]. Number [Han92b, Ano94-63, Ano94-125, Ano94-150, Tho90c, Tho94h]. Numbers [And90, ADE90, Egl93d, Len93a, Nel91b, Ano91-83, Fri91]. numeric [Web93]. numerical [Gri94]. Numerous [Ano94-45, Ano94-40, Egl93b]. Nutek [Ano92-109]. Nx586 [Ano94-118].

O [Ano91-48, Ano92-79, Ano92-220, Gla90g, Par90, Pou93h, RH93, Van90, WK90]. oasis [Cra91i]. Oberon [Pou93d, Fou91, KPUG93, Pou91d]. Object [Ano92-117, Ano92-127, Ano94-121, Cla92c, Gib90b, Gre94a, Mul94, Nan93f, Ras92f, SG92a, Ste94a, Way92a, Ano90-65, Ano91-65, Ano91-111, Ano92-150, Ano92-186, Ano94h, Ano94-29, Ano94-101, Ano94-181, Ano94-177, DH90, DM90, GFK94, Gre94b, Hol91, Kay94c, Kiy94, Lu92, Mas90a, MS93, Nan90b, Par92, Pou93d, Pou94g, Sch92a, Smi93e, Spi90b, SH90, Ude94g, Van93a, Vor93, Way94f, You90, Ano91-111, Ano92-165, Ano94h, Ano94-120, Gre94a, HH94a, Hol91, Key92, Nan90b, Pou90i, Ras92e, Spi90b, Url90]. Object-1 [Nan90b]. object-based [Ano94-177, Sch92a]. Object-oriented [Ano92-117, Gre94a, Ste94a, Ano90-65, Ano91-65, Ano91-111, Ano92-150, Ano94-29, Ano94-181, DH90, DM90, GFK94, Gib90b, Gre94b, Hol91, Kay94c, Kiy94, Lu92, Mas90a, MS93, Mul94, Nan90b, Nan93f, Par92, Pou93d, Pou94g, Smi93e, Spi90b, SH90, Ude94g, Way94f, You90, Ano92-165, Gre94a, Hol91, Key92, Pou90i, Ras92e, Spi90b, Url90]. objections [Ano94-57]. Objective [Ano94-29, Sch92a]. Objects [Ano90-65, Ano92-127, Ano93-42, Ano94-38, Cla92b, DM90, Gib90b, Kay94c, Lu92, Pou94h, Ras92f, SG92a, Way94f, Ano91-184, Ano93-133, Ano94-182, Dej94, Key92, Mul94, She90b, Vor93, Ano93-118]. ObjectworksSmalltalk [Ano90-95]. Obscure [Ano92-34]. observers [Ano94-150]. Observing [Ano94-122]. Obsolete [Ano90-43, Ano94-169, Rya90b]. obstacle [Rya90a]. obstacles [Smi90c]. obtaining [Ano92-115]. Occam [Ano90-59, Ano92-95]. ocean [Min90j]. Oceanographic [Smi93b, Min90j]. OCP [Pou90b]. OCR [Ano93-44, Ano93-52, Bar92f, DE91b, Egl94c, Van93a, Wsz93a]. Octane [Rob90b]. October [Ano93-59]. Octocom [Ano93-128]. ODBC [Api94c, Ude93a]. Odd [Tho92c, Ano92-118, Way93i]. Odds [Pou94o, Ano93-46]. ODLI [Fis90b]. Off [Ano91-58, Ano92-188, Ano94-50, Ano94-83, Bry94a, Fin90b, Gas93, Gla91a, HHP93, Laz93, Rei94a, Rog91, Ano93-31, Ano94-79, Ano94-165, CV90b, Cla93c, Gil94, Wal94d, Wig92, Gre90b]. Off-the-shelf [Laz93, Gil94]. Offer [Ano91-122, Ano91-29, Ano91-72, Ano91-177, Ano93-110, Ano94w, Ano94-143, Bis91, Day90, Ebe93, Lan94a, Laz90, Mia92b, Nan93d, PP93, Ras90b, Sel92, Van94b, WS91]. offered [Ano94-171]. offering [Ano94-189, RT94b]. Offers [Ano94-112, Woo90a, Ano90-32, Ano90-79, Ano91n, Ano91-151, Ano92-54, Ano92-87, Ano93z, Ano93-43, Ano93-129, Ano94-40, Api91c, Cal91a, Cra90a, Joc90b, Kiy93a, LM90, Lin94, Lov90a, Mia90f, Min90f, Nad90e, Nan91h, Nan93f, Orr94, Pou93a, Ric91, Tho90f, Yag93h]. Office [Ano91-113, Ano91-119, Ano91-145, Ano93-83, Egl91d, Law94, Alf91b, Ano91r, Ano91-37, Cot93a, Har91a, Lin91b, Nan91k, Taz91, VC90k, Ano93z, Ano94-165, Nan93h]. offices [Ano94-34, ES94]. offs [Cal93, Egl93c]. Offshore [Ano94-142]. often [Ano94s, Ano94-76, Slo91]. Okidata [Ano90-67]. OkiLaser [Ano90-67]. Old [Ano91n, DM90, Alf91a, Ano90g, Ano92-100, Ano94-97, Gea90, Pou94e, Pou90p, Tho94h, VC90d]. oldest [Cra91c]. OLE [Ano93-72, Ano94h, Ano94-29, Ano94-49, Ano94-117, Ano94-123, Kiy94, Nan92b, Ple94, Sch93a, Ude94b, Ude94c]. OLE-supported [Ano94h]. Olivetti [Lav91, Ano92-128, Pou93e]. Omnibook [Wau93b, Api93b]. on-line [Ano92-217, Ano93-134, Ano94-96, Ano94-149, Ano94-166, Egl94b, Ban91]. on-screen [Min92d, Vau90]. on-the-fly [GW93, Ano92-129]. Once [Pou93r, RP90, Fri90, Gut93, Pou94s, Ras90j, VC90t]. One [Alf90b, Ano90-86, Ano91-105, Ano92-205, Ano93-110, BGE^{+}92, Egl94e, Nad92, Pou94r, Ude91d, Ude93i, Ano90-45, Ano90-60, Ano91-115, Ano92-31, Ano92-43, Ano92-149, Cla93e, Hal93d, Hog90, Hov93a, LB91, Lov94, Man90, May94a, MG90, Nie94, Sei91, SU93, Tho91a, Tho94i, Ude90h, VN91c, Ano91j, Egl94e, Nan90a]. one-fifth [Hal93d]. One-Size-Fits-All [Nan90a]. One-Ups [Nad92]. Ones [Ano91-114, HHP93, Tho92d]. Onion [Pou93t]. only [Ano94-82, Cau93, Daw94d, Egl93a, Hol90, Hur93a, Ras90j, Wal94e]. ons [Ras91a]. onto [GK91, Nan94b, Yag91e]. OOP [Har90a, Par92, Pou90i, Pou91c, Spi90a, Url90]. OOPS [Ano92-130]. Open [All92a, Ano91-38, Ano91-115, Osh90, Ano94-180, ESN94, Sch93b, Yag93d, Ano91-115, Ano94q, Ude93a, Yag90j]. OpenDoc [Ano93t, Ano94-49, Pie94]. OpenGL [Ano93-135]. Opening [Ano92-185, Laz90, Lin94, Bry94c]. openly [Red92b]. openness [Ano92-159]. Opens [Ano92-85, Ano94k, Alf92g]. OpenStep [Web94]. OpenVMS [Smi94a]. Operating [Ano91-54, Ano92-125, Ano92-166, Ano93-55, Ano94-159, Bar91c, Don92, Smi90a, Smi90b, Smi91a, Ude94a, Yag90a, Yag90b, Yag90c, Ano90-62, Ano90-68, Ano92-89, Ano92-80, Ano92-81, Ano92-178, Ano92-158, Ano92-223, Ano94n, Ano94-45, Ano94-126, Ano94-100, AWGY90, Api94b, Car91b, Cra91k, Fie90l, Hal94f, Hay94a, Hay94b, Hol91, Hyo94, Kea90, Nan93c, Nan93d, Nan93h, Nie94, Pie94, Pou91h, Pou93d, Pou94a, Pou94g, Rei94g, Sal94, Sch92a, Tho93f, Ude92d, Ude93f, Ude94f, Ude94j, VC90g, Var94b, Var94a, Way94f, Web94, Ano94-91]. Operating-System [Don92, Ano93-55, Hal94f, Hay94a, Hol91, Nan93d, Ude94f, Ano94-91]. operation [Ano94-111, Egl93c, SY90]. operational [Ano92-174]. operations [Gre93c, Way93e]. Opinions [Ano93-56, Var94a]. Opportunities [Ran93, Laz90]. opportunity [Bar90g]. opposing [Loe90]. opposite [Rya90e]. ops [Bar90j]. optic [Ano93-93, Fis90c, Min90j, Rei94b]. Optical [CP94, Har90b, Pou90t, Psa92, Rei91d, Rei93c, Rya90e, Web92, Ano92-45, Bry94c, Cas92, HR90, Isl92, Way93i]. optical-fiber [Bry94c]. Optimal [Way93a, Way93i]. Optimistic [Ano93t]. optimization [Pou93i]. Optimizations [Ano94-130, Tho94g]. optimize [Mat94b]. Optimizing [Ano94-124, Lan94b, Yav94]. Option [Alf92d, BF90, VN91b]. optional [Tho94b]. Options [Ano94-125, Nan91c, Lov93b, Nan93d, Nel91a, Rei90b]. Optiquest [Ano92-131]. opts [Ano93-141]. Opus [Ano90-80, Pou90c, Yag90l]. Oracle [Ano92-59]. orange [Ras91h]. Orchid [Pou93t]. Order [Ano91-31, Bur92, Ano92-149, LDH94, Wsz91]. ordered [WET90]. ordinary [Yag90c]. Organization [Sey93, Bac94, Bry94d, Loc92, Smi94a, SE94, Wal94d]. organizations [Ano90j]. Organize [Bon90a, GD92a]. organized [Bon90b]. organizers [Bar93b, Ras90n]. organizing [Ano91-155, Ano92-129, Ano92-102]. orientation [Ano92-127]. orientations [Ano90-77, HC93]. Oriented [Ano91x, Ano90y, Ano90-65, Ano91-65, Ano91-111, Ano92-117, Ano92-150, Ano94g, Ano94-29, Ano94-181, DH90, DM90, Fri94c, GFK94, Gib90b, Gre94a, Gre94b, Hol91, Kay94c, Kiy94, Lu92, Mas90a, MS93, Mul94, Nan90b, Nan93f, Par92, Pou93d, Pou94g, Smi93e, Spi90b, Ste94a, SH90, Ude94g, Way94f, You90, Ano92-165, Gre94a, Hol91, Key92, Pou90i, Ras92e, Spi90b, Url90]. Origins [KUD92, Cra90b]. Orlando [Ano93-99]. OS/2 [Ano91-117, Ano91-116, Ano91-132, Ano91-192, Ano92-83, Ano92-91, Ano92-132, Ano92-112, Ano92-125, Ano93-80, Ano93-39, Ano93-40, Ano93-41, Ano93-60, Ano93-51, Ano93-61, Ano93-55, Ano93-86, Ano93-110, Ano93-117, Ano93-78, Ano93-79, Ano93-132, Ano94r, Ano94-33, Ano94-126, Ano94-75, Ano94-97, Ano94-127, Api94b, CV90c, Dan90, Egl91c, Gre90g, Gre94e, Ham90, Ham91, Ham92, Hel90c, Hel90d, Hel90e, Hel90b, Hel91c, Hol90, Ken93b, Kog94, Kro90, Mas90b, Mas91c, Mas92, Mas93, Min90c, Min90i, Min90d, Min90e, Min90h, Min90f, Min91b, Min92b, Min92a, Min93a, Min93b, Nan90b, Nan92d, Nan92e, Nan93a, Nan93c, Nan93f, Nan94a, Pou92n, Pou93n, Rya90d, Smi93a, Ude90g, Ude91b, Ude91f, Vos90b, Yag90k, Yag93f]. OS/2-based [Ude91b]. OS/90 [Ano90-68]. Other [Ano90c, Fie90f, RARH92, Ano91-184, Ano92s, Ano92q, Ano92-60, Ano92y, Ano92-61, Ano92-74, Ano92-90, Ano92-91, Ano92-92, Ano92-113, Ano93e, Ano93-60, Ano93-50, Ano93-51, Ano93-53, Ano93-55, Ano93-59, Ano93-58, Ano94-76, Ano94-71, Ano94-166, Api94b, Bry94c, CV90a, Egl94c, Fri94b, Gre91a, Gre90c, Hed93a, Hol93b, KGTD92, KCU^{+}93, Ken93b, Ken93a, KGEM93, KDG^{+}93, MW94, Mit91a, Nie94, Ras91a, Ref94, Ude93h, VN90a, Vor93, Way91c, Way93i]. ought [Ano94-79]. Our [All93i, Ano90-69, Ano94-90, Fie90a, Fie90f, FJK^{+}93, Fox93, HK94, Ken93c, Pou90u, Pou93i, Ude93i, Ano90t, Ano90-38, Ano91y, Ano91-71, Ano92-29, Ano92-42, Bux90a, Cos94, GD92b, HM94, Joc92b, KM94, Len94a, She90a, Smi91a]. ours [Bux90a]. Out-Powered [Tho93f]. Outbound [Loe90]. outdone [VN90a]. Outguns [Lin90b]. outlining [Pou90w, Rei91c]. Outlook [Bar92b, Red91b]. Outpaces [Joc90a]. outperforms [RT94a]. output [Ano91-48, Ano92-52, Egl93c, Lan90b, LS91a, Mia92b, Vau90, Way93i, Zei93]. outs [Ano92s]. outstanding [Ano92m, Ano92-94, Kiy93a]. outweigh [Cra90i]. Over-the-wire [Nan94a]. overall [Sha94b]. overcome [Bar92a]. Overcoming [Min92b]. Overhaul [Gar94]. overhead [Ano92-183]. overloads [Ano94-150]. Overseas [You94]. overshadowed [Ano94-158]. Overturning [Tho91a]. Overview [ASS92, Ano92d, Ano92-79, Ano92-102, Ano92-134, Ano92-133, Ano92-135, Ano92-202, Ano94-89, Bor93, Don92, HL93, Loc92, Lu92, MW93b, Vas93b, Web92, Web93]. Own [Ano93g, Lie93, NTS93a, Ano91-39, Ano92-146, Ano94-176, Lin90b, Ras90f, Ras91k, RUHG93, Rei94b, Rei94g, Sel92, ST93, Sum90, Tho94c, Van94b]. Owns [McG90b]. Oxford [Ano92-136].

P1394 [Ano94-72, Cla94b]. P1394/Fire [Cla94b]. P5 [Ano92-111]. P54C [Ano94-27]. P75 [Alf91e]. PA [Pou94b, Smi94g]. PA-7100LC [Smi94g]. PA-RISC [Pou94b]. Pace [HP94, RH93, Wal94c, Yag91a]. paces [Die90c]. Pacific [Egl90c]. PacificPage [Ano91-118]. Pack [Api90b, Pou91r, Ano90-73, Ano91-63, Smi90d, Way94c]. package [Ano90-56, Ano90-114, Ano91-167, Ano92-36, Ano92-123, Ano93z, Ano93-76, Ano93-81, Cla93e, Hal93b, Lan90a, Mia90e, MS93, Nad90f, Pou92d, Ras90o, Rya93b, Sta94a, Ude93m, Van92a, Yag93k].

Packages [Mia93b, Ano91-63, Ano92-149, Ano94-92, BGE^{+}92, Egl91d, ET91, Egl93a, Egl94b, GD92a, HH94a, HC93, Man90, Mia92b, RM92, Tho94i, Van93c, VC90f, Wsz93a, Wsz93c, Yah92]. packaging [Smi94c, Smi94g]. Packard [Alf91b, Ano90-40, Ano90-63, Ano91-58, Ano92-122, Ano93-36, Cry92a, Egl92b, Joc90c, Lus90, Pou94b, Rei93b, Tho93b, Wau93b]. packed [Ano93-109, Ano93-119]. Packet [Sta91b, Bry94d, RF91]. packet-switching [Bry94d]. Packets [ATU92]. packs [Wau93b]. Pad [PP93, Pou93a]. Page [Ano91-53, Fin90b, Ano90-69, Ano90-77, Ano92-113, Ano93-44, Van93a, Van93c]. page-layout [Van93c]. page-turner [Ano90-69]. PageAhead [Ano92-38]. PageGarden [Vos90a]. PageMaker [Ano90-70, Van93c]. Pagemarq [Lov92b]. PagePlus [Ano93-81]. Pages [Way93f]. paging [Ano94-145]. Pain [Ano90-62, Ano91-56, Ano91-66, Ude90f]. painless [Hal94i]. painlessly [Ano93-32]. Pains [CV90a, Cra91h, Ano91-81]. Paint [TG91, Ano93-82]. Painter [Vor93, Ano93-82]. Pair [Ano92-137]. pairing [Ano94-171]. pairs [Rya90g]. Palatable [San90b, Ham92]. palette [Hel90c, Par92]. Palm [Egl90c, AR91]. Palmtop [Mar92c, AR91, Cal93]. Palmtops [Pou92m]. Palo [Cla91a]. Pandora [Ano93-83]. Panose [BL94]. Paper [Ano92-138, Ano93-62, Bar92e, DE91b, Ga 93c, Sty91, Ano90-85, Ano93-132, Hay93, Rog91, Way94e]. Paperless [Ano90-71, Ano91-119, Ano93a, Ano94-189, Egl91d, HR91, Ano91-37, Har91a, New93, Ano93-132]. Paperless-Document [Ano94-189]. paperwork [Egl91d]. Parade [Fie90c, Pou93t]. paradigms [Var94c]. Paradise [Ano93-30]. Paradox [Kli90b, Ano90-72, Ano91-152, Ano92-59, Ano94w, Kli90b, MS93]. Paradoxes [Ano92-137]. Paradyne [Bry94b]. Parallel [Ano92h, Ano94-39, Ano94-140, Ano94-167, Pou90h, Pou94g, Psa92, Rob91a, Sha94c, Ano92-169, Ano92-221, Ano93f, Ano94-88, Egl93f, GP90, Hay94b, Pou90g, Pou91b, Pou94d, Ran90, Rya91d]. parallel-port [Ano92-169, Ano93f, Ano94-88, Egl93f]. Parallel-processing [Ano92h]. Parallels [Dys92]. parameters [Fie91f]. Parametric [Bis91]. ParaSet [New93]. PARC [Ano91-42, Cla92a]. ParcPlace [Ano90-95]. Parents [YS92]. Park [Cri93]. Part [Ano90-105, Ano91x, Ano91y, Ano91z, Ano92-29, Ano92-30, Ano92-61, Ano92-80, Ano92-81, Ano92-146, Ano92-147, Ano92-148, Ano92-182, App91, EB90a, Fou91, Gla90g, Gla90d, Gre90g, Gre92, Nan91e, Pou90h, Smi90a, Smi90b, Smi91a, Yag90a, Yag90b, Yag90c, Yag92b, EB90b, Gre90d, Lov92c, VC90s, Wal94e]. participants [Ano91-132]. particular [MH91]. partitions [BBWR90, NTS93a]. partly [Ano94-172]. Partners [Ano93q, Ano93-84, Len93b]. Parts [Ano91-34]. party [Ano90-33, Tho94d, Ano94-62]. Pascal [Ano91-175, Ano91-174, Ano93-118, Har90a, Pou90i]. pass [Ano94-182]. Passage [Ano91-120]. passing [Kor93, Nan94a, Ste91c]. Passive [Sma93b, Ano94-112]. passive-matrix [Sma93b]. Past [Ano94u, Ano94v, Bry94e, Hay90a, Nad93, The94]. paste [Ano94l]. pasture [Bar93a]. patching [Ano91-176]. Patent [Ano94-69]. patents [Ano94-69]. Path [Ano94-135, Smi94a, Ano91-151, Ano94-111, Van93a]. path-based [Van93a]. Paths [Api94b, LAN92]. Pathways [All93h]. Patient [Wal94b]. Patrol [VC90r]. Pattern [MW92a, MW92b]. Patterns [DH90, RNG94]. Paulo [Ban92]. Pay [Loc92, Ano93-31, Wal94d]. paying [Ano92-104, Rie91]. PC [Alf92f, Alf92g, And90, AR91, Ano90-73, Ano90-74, Ano90-75, Ano91-31, Ano91-62, Ano91-125, Ano91-153, Ano91-196, Ano92-53, Ano92-56, Ano92-70, Ano92-87, Ano92-120, Ano92-210, Ano92-223, Ano92-221, Ano93-27, Ano93-38, Ano93-74, Ano93-84, Ano94l, Ano94v, Ano94-30, Ano94-51, Ano94-57, Ano94-90, Ano94-116, Ano94-166, Ano94-159, Ano94-181, AMW90, Api91c, Bar91e, Bra90, Bry94a, Bus92, Car91a, CS92, DWN91, DE92, EWT90, ET91, Egl93d, Fie91h, Fie92b, Fri90, Ga94, GAW93, Gre90c, Gre90j, Gre90k, Gre93d, Hal94c, Hal94i, Hay90b, HM94, HC93, Ken94, KDG^{+}93, LAN92, Lin93, Lov91a, Lov91b, Lov93b, McN94, Mia90b, Mia90h, Nad92, Pou90f, Pou91g, Ras90l, Red92a, RH93, Sch92b, Smi90e, Smi93d, Smi94a, Tho93f, Tho94c, Ung90b, VC90m, Way90c, Way94h, Yag90l, Yag91b, Yag91h, Yag92a, Yag92i]. PC-based [Ano91-62, CS92]. PC-compatible [Hay90b]. PC-DOS [Ano93-38]. PC-File [Ano90-73]. PC-Style [Smi90e]. PC-to-Mac [Ano92-87]. PC-to-PC [Ano90-74]. PC-VCR [Yag91b]. PC-Write [Ano90-75]. PC/Television [Ano92-221]. PCI [Ano93-85, Ano94-184, Ken94]. PCL [Alf91b, Pou90d]. PCMCIA [Alf92g, Ano92-139, Ano94-88, Ano94-111, Ano94-168, Ano94-176, Ano94-180, Bry94e, Pou92m, Sma92, Tho94b]. PCs [All92b, Ano90-60, Ano91-62, Ano91-77, Ano91b, Ano91-103, Ano92-41, Ano92-58, Ano92-100, Ano92-110, Ano92-211, Ano93-50, Ano93-75, Ano94-32, Ano94-54, Ano94-126, Ano94-71, Ano94-129, Ano94-166, Ano94-183, Api92b, Bar91a, Cal93, Day90, Fie91h, Fie91i, Fie92a, Fox93, GN93, Ga 94d, GAM90, GW93, Gut93, HP94, Lin91b, Lov91a, NW92, Per94a, Ras91l, Rei92b, Rya90c, San90b, Sch90a, Smi93d, Smi94c, Smi94e, Tho91b, TG91, Tho92a, Tra90, VC90p, Way90c, Woo91a, Woo91b, Yag93e]. PCs. [Smi93e]. PCSes [Ano94-128]. PDA [Ano93-45, Ano93-67, Ano94-162, Ano94-168, O’M94, Pou93a, Sta94b]. PDA600 [PP93]. PDAs [Ano93i, Ano94-102, Ano94-113, Ano94-168, HPPR93, Rei94g]. PDS [Wig90]. Peck [Ano91-121]. Pedestal [Yag90g]. Peek [Hal93d]. Peer [Ano91-122, Ano94-92, Nan94d, Ude92c, Nan93d, Tho91e, Ude93b]. peer-LAN [Nan93d]. Peer-to-peer [Ude92c, Tho91e, Ude93b]. Peers [Top91, UM90, GAW93]. Pen [Ano92-62, Ano92-61, Bar92b, Bar92a, CRR93, Egl93a, MB93, WS93, Ano92-35, Ano92-76, Ano93-130, Car91b, Pou93a, Sma93a, Ano92-140, Ano93-86, Egl93a, Pou93a, PP93]. pen-based [Bar92b, Sma93a]. pen-centric [Egl93a]. pen-input [Bar92a, Ano92-140]. Pen-software [Ano92-61]. penalty [Ano92-102]. Penlessly [Die90a]. Penn [SE91]. Penny [DE91a, Rei92b]. Penny-Pinching [Rei92b]. Penny-wise [DE91a]. PenPoint [Car91b, Egl93a, Rei91b, Cla92b]. Pentax [Ano90-42]. Pentium [Per94a, RH93, Ano93-54, Ano93-134, Ano94-82, Ano94-129, Ano94-130, Fox93, GN93, Ga 94d, Hal93c, Hal93d, Hal94h, HK94, HP94, Mat94b, Per94a, Pou93g, Pou94j, RH93, Rya94b, RT94a, Smi94g, Tho93f]. Pentium-based [Ano94-129]. Pentiums [HK94, HP94, MM94]. Pentop [New93]. PenWindows [Rei91b]. People [Ano91-92, Mel93b, Pou93e, Ano90-112, Ano91-155, Ban92, Egl90b, Pet90, Sch91b, SE94]. Perestroika [VC90o]. Perfect [Ano93-37, Die91, Fie90h, Ano90-70, Cra91f, Mit90b, Ran90, VC90l]. Perfectly [DWN91, VC90q]. perform [Hol93b, Hov93a, MW92b, Mul94, Way92b, Way94j]. Performance [Ano91-159, Ano91-170, Ano92-53, Ano93-87, Ano94-152, Api92b, Fox93, Hog90, Min90a, Min90b, Ric92, Tho91d, Alf91d, And90, Ano90-62, Ano91-50, Ano91-76, Ano91b, Ano92m, Ano92-49, Ano92-94, Ano92-102, Ano92-122, Ano92-178, Ano92-188, Ano92-215, Ano93-75, Ano94-57, Ano94-87, Ano94-148, Ano94-149, Api94d, Api94e, Bac94, Bar90i, Daw94c, Egl94e, Egl94f, Ess94, Gla90b, Gre90b, Gre91b, Hal93d, Hed93a, HH94a, HK94, Kli90b, Lan90a, Len94b, Mar90a, Nad94b, Nan93d, Par90, Per94b, Pou92d, Pou93h, Pou94b, Psa92, RM92, RS90, Smi94c, Smi94g, Tho90f, TS90, Tho93f, Tho94g, Tho94h, TR94b, Van94a, VC90d, Yag90l, Ano94-71, Car91a, Len94b, VC90d]. Performas [TH92]. Performer [MG90]. performers [Ano93-47, HP94]. Performs [Tho93f, Ano90-27, Mit91a]. Peripheral [Ano93-85, Bar90k]. peripheral-and [Bar90k]. Peripherals [Tho93e, Bry94a, Bry94e, Dav94a, Rei94g]. Perl [Ano92-141, Chr94, NTS93b]. Perl-based [NTS93b]. permit [Ano94g]. Perpetual [Fow91]. persistent [SG92a, Ste94a]. persistent-data [SG92a]. person [Ano94-132]. Personable [Bar93b]. Personal [All92b, All93h, Ano92-63, Ano93-88, Ano94-132, Bar90g, BBWR90, Bon90a, Bon90b, Cla94a, Cra91c, Die94, Fri91, Lab92a, LAN92, Nan94d, Ras90n, Rya93a, Smi90f, TS93b, Ano91-77, Ano92-77, Ano94z, Ano94-65, Ano94-145, Bar91a, Dav94a, DW90, Fie90b, Hal94d, Har91c, Kof93, Laz90, Lin91c, Mil93, Mos93, O’M94, Pou91s, She93b, Way94a, Woo90b, Yag90c, Nan94d, Smi90f]. personal-and [She93b]. personal-finance [Kof93]. personal-finances [She93b]. personalities [Hay94a]. Personality [Hay94a]. Perspective [Ano93-78, UFC94, Ano93-59, Ga 94a]. Perspectives [Ref94, Ano90-112]. Persuaders [Ano92-69]. Persuasion [Ano91-123]. Pesky [Ras91i]. PG [Tho93c]. PgC7600 [Pou91a]. PGP [Sta94a]. Phase [Rog91, Rya90e, VC90j, Yag90i]. Phase-change [Rog91, Rya90e]. Phaser [Tho92b, Tho92g]. Philip [Ude93d]. philosophies [Ano94-84]. Phone [Ano93-84, Ano92-177, Ano93m, Ano94-145, Fri94b, Ga 94b, Nan94b, Rei94b, Ude94h, Ano94-28]. PhonePro [Ga 94b]. Photo [Ano92-142, Cal91b, Lov91c, Ups90, Ano92-142, Cal91b, Ga 93d, Lov91c]. photo-realism [Lov91c, Ano92-142, Cal91b, Lov91c]. Photo-realistic [Ano92-142, Cal91b, Ups90]. PhotoEdge [Ga 93d]. Photographic [KYT^{+}92, Zei93]. photographic-quality [Zei93]. Photography [Egl93d, Ano93-55]. Photonic [Egl94d]. Photonics [Web92]. PhotoPaint [Ano94-40]. Photoshop [Ano94-136, Ano90-76, Ano91-124, Ano94-131, Egl94f, Yag93g]. PhotoStyler [Egl94f]. phototypesetter [Ano92-50]. physical [Egl91c]. physicist [KYT^{+}92]. physiology [Alf90a]. piano [Ano90-116]. Pick [Ken92, Nad93, Smi90b]. picked [Ano93-48]. Picking [Mal93, Ano94p]. picks [Ano91-122, Ano92w, Die91, Pou91r]. Picture [Ras90k, Vor93, Pou94h, Ano92-143, Ano94-131]. pictures [BGE^{+}92, Lip90b]. pie [Kay94d]. piece [Ano93-75, Kay94d]. Pieces [Lee92, Ano94-76, Cla93a, VC90q]. PIEdit [Ano91-32]. pigeonhole [Ga 94c]. Pilgrim [Ano91-116]. PIM [Wsz93b]. PIMS [Ano94-132, Bar93b]. pin [Ano92s]. pin-outs [Ano92s]. Pinching [Rei92b]. Pinnacle [Tho94c]. pioneer [Lan90b, Egl93e]. Pipeline [Ano94-133]. pipelines [Ano93-116]. pipelining [Rya93b]. Pitch [Die91, Cla93b]. pitfalls [Kro90, Tho94g]. Pivot [Ano91-53, Ano90-77]. Pixar [Cal91b]. pixel [Ano94c]. Pixels [Hin92, Lov90a, Ano92-33, Api94a]. PixelView [Ano91-125]. Place [Cra90n, Cla93a, Hal94e, Ras92f]. plain [Rog91]. Plan [Ano91-33, Ano92-146]. Planet [Api92b, Pou91k]. Plans [Ano94-77, Ano94n, Ano94-127, Tho92e]. Platform [Ano94m, Api94b, Cou92, Ga94, Ano90-95, Ano91-32, Ano91-51, Ano92-133, Ano93-31, Ano93-64, Ano94-45, Ano94-117, Ano94-174, Ano94-183, Cro93a, Egl94f, Gre94e, Hel91d, Ras91g, SE94, Ude93l, UPSW94, Yag93e, Ano92-163, Ano94-38]. Platforms [AWGY90, Ols94, Ano92-89, Ano92-133, Ano94h, Ano94k, Api94b, Ga94, Hal94f, McN94, Nag93, Pie94, Tho94c, Ude94k]. play [Ano90-116, Ano92-103, Ano94-32, Bry94e, Yag90e, Ano94-97, Ano94-134, Ano94-163, Ano94-173, Hal94i, Ung90a]. playback [Ano93-46, Ano94-106, Ano94-152]. Players [Gre91b, Per93c, RARH92, Ano91-54, Ano94-152, Rei94b]. playgrounds [Lov90b]. Playing [Rei90d, Ano91-66, Pou91c]. plays [Ano91-150]. Plea [Ano90-78]. pleasant [TR94a]. Please [ET91, MPW^{+}90]. Pleasure [Pou90n]. Pleasures [Cra91h]. plenty [Rya93b]. plots [Die90a]. plotter [Ano92-152]. ploy [Ano91-183]. Plug [Ano94-188, Nan94b, Ano92s, Bry94e, Egl94d, Wsz93c, Ano94-97, Ano94-134, Ano94-163, Ano94-173, Daw94b, Hal94i, Ung90a]. Plug-and-Go [Daw94b]. Plug-and-Play [Ung90a]. Plugging [Ano93-102, Der92]. plugs [GTGW94, Wsz90]. Plus [Ano90-34, Ano90-79, Ano92-151, Gla91c, Hay94a, Joc90b, Spi90a, Wsz93a, Ano92e, Ano91-129]. pluses [Cra90i, Cra90m]. PM [Ano90-51, Ano91-56, Ano91-65]. PM10 [Ano90-85]. PN1 [Daw94a]. Pocket [Ano94-53, Mia90d, Nad92, New93, Pou90f, Rei91a, Red91a, Ken93b]. Pocketable [Pou92a]. Poet [Gre94b, Gre94b]. Poets [Ano91-126]. point [Ano94-168, Car91b, Lip90a, TR94b, Yag90h, Ano91-167, Cla93c, Smi94d]. point-and-click [Yag90h, Cla93c]. Point-to-Point [Smi94d]. Pointers [Len94a]. pointing [Ano91o, Ano91-172, Ano92-139, Egl92a]. Points [Ano91-91]. poised [Ano92-108, Ano93-67, Ano94-85]. poker [Pou91c]. Polar [Sta94b]. Policy [Ano94-67, Gru90]. Polish [Ude91c, Ano92-38]. polished [Ano92-143]. Political [Ano93-89]. Politics [Ano93-114, Ano94-95]. Poltergeist [Min92d]. Polysilicon [Ano94-135]. Pondering [Pou92n]. pool [Smi90d]. Poor [FJPR93]. Popular [Rob91a, Alf90b, Ano92-154, Ano94w, DE90, Egl91a, Egl94b, Fie90c, Van93c]. popularity [Ano94-172, Smi94d]. Poqet [Nad92, Ras90l]. Port [Ano94-140, Gre94c, Ano92-169, Ano92-221, Ano93f, Ano94-88, Egl93f, Egl94d, Nie94]. Portability [Alf92g, Ano91-127, Ano92-159, Loe90, Mia90e, Ano91-195, Api94b, Api94d, DWN91, Hur93a, Mit93b]. Portable [Alf91e, Ano90-80, Ano91-146, Ano92-193, Ano94-88, Cra91f, Cra91h, DWN91, HHP93, Hol94, Joc90a, Lin90a, Mar92b, Nad91, Sch90a, VC90g, Ano91o, Ano91-197, Ano92u, Ano92-31, Ano92-98, Ano92-194, Ano92-200, Ano93o, Ano93-126, Ano94-189, Ano94-177, Egl92a, GK91, Len94b, Len94a, Mia92b, Pou90w, Pou91r, Pou91q, RM92, Ras92a, VC90i, Ano91o, Ano91-154, Cra90i, Gre93d]. portable-computer [Len94a]. Portables [Ano92-124, Len94b, Len94a, Rei90b, Ano91-128, Api91a, Api94d, Car92]. portend [Ano92-209]. Porthole [Hel90e]. Porting [Ano92-89, Ano94-136, Ano94-137, Sha94b]. portrait [Ano90-77]. Ports [Tho93f, Ano94-129, Ano94-184]. position [Cla93c]. Posix [Ano92-159, Ano92-159, Ano93-139]. possibilities [Rei93e, Ste91b]. possibility [Tho93f]. Possible [KYCD92, Fri94b, Mul94, Pou94g, Tho92d, Yag93e]. Post [Ano94-124]. Post-Compilers [Ano94-124]. Postman [ET91]. PostScript [Ano91-118, Ano92-153, Cra90c, DE91a, Egl90c, Joc90b, Lan90a, Lan90b, Tho92a, Tho93d, WE92, Ano92-90]. potential [Ano91-37, Ano92-198]. potentially [Yav94]. potpourri [Pou94l]. pound [Ano92-128, Nad90d, Nad90f, Rei90c, Wau93b]. Pounds [Row93]. Pountain [Pou94a]. Pournelle [Ano90-60, Ano93-52, Pou93l, Pou93m, Pou93n, Pou93o, Pou94l, Pou94m, Pou94n, Pou94k, Pou94i, Pou94o, Pou94j, Pou94p, Pou94q, Pou94r, Pou94s]. Power [Ano91-58, Ano91-103, Ano91-108, Ano91-128, Ano92-144, Ano92-145, Ano94-160, BR93, Car92, CV90b, Cry92b, Daw94d, Egl93a, Fie91b, GN93, Ga 94c, Gre91c, Gut93, Hol93a, KGEM93, Nan92a, Ras90l, Ras91h, Rei90b, She92, Smi94c, Smi94g, The94, Tho94h, VN91c, Yag90l, Yag91h, Yag92f, Ano91-166, Ano91-197, Ano92g, Ano92u, Ano92-58, Ano92-78, Ano92-87, Ano92-105, Ano92e, Ano92-142, Ano92-155, Ano92-183, Ano92-218, Ano92-222, Ano93-127, Ano94-82, Api90c, Api91c, BBWR90, Cla90b, Cra90f, DWN91, Fow91, FP93, Gre94d, HH94a, Kea91b, Len94a, Nad90e, Nad90f, Nan91i, Pou92d, Pou94g, Pou92j, Sal94, Tho92h, Ude93b, VC90e, Vor93, Way94i, Wsz91, Yag91e, Ano94-108, CV90d, Ga94, GFK94, Gre94c, The94, Tho94g, Tho94h, Wau92, Wsz93c]. Power-Hungry [CV90d]. power-on [Api91c]. power-saving [Ano91-108, Ano91-166, Pou92d]. Power2 [Ano94-138]. PowerBasic [Ano90-81]. PowerBook [Len94b, Tho93a, Tho93e, Tho93i]. PowerBooks [Ano92-86, Ano93-90, NTS92b, Tho93i, Tho94b]. PowerBuilder [HH94a]. powered [Ano93-65, Gar94, Hay90b, LAD91, Tho93f, Yag92a]. PowerExec [Ano93-91, Mit93b, Sma92]. PowerFlex [Nad90a]. Powerful [Len94a, Pou92a, Ras90j, Ano90-40, Ano92-187, Ano93-85, Ano94-68, Ano94-176, Bar90c, Bar91a, Coo90a, Fie91l, Fri94c, Ga94, GP90, Hur93b, Lin90a, Lin91c, Lov92b, MS93, Nad90d, NTS93b, Orr94, Smi93c, Tho90c, Tho92h, Tho93c, TH93, Ude94c, Vel90, Yag93a]. Powerfusion [Ude91d]. Powerhouse [Gut94, Ano92f, Yag92i]. Powerhouses [Ano91-159, SG92b]. PowerMate [Ano92-94]. PowerOpen [Ano92-125, Hal93d, Tho93f]. PowerPC [Ano92-63, Ano93-29, Ano93-52, Ano93-66, Ano94m, Ano94d, Ano94-45, Ano94-107, Ano94-127, Ano94-159, Daw94c, Gre94c, Hal93d, Rya93b, RT94a, Tho93f, TR94a, Tho94b, Tho94k, Tho94h, TR94b, UFC94]. PowerPC-based [Daw94c]. PowerPCs [Ano94-98]. PowerPlay [Ano92-219]. PowerPoint [Cry92b]. powers [Rya93a]. PowerSoft [Gre93a]. ppm [Ano93-92]. Practical [Ano92-146, Ano92-147, Ano92-148, Sch92b, Sta91a, Yag92b, Ano92-52, Ano92-101, Ras90l, Rei91a, Rya90f]. practicality [Rob90b]. Practically [HR91]. Practice [Hel91a, RNG94]. practices [Ano91-59]. practitioner [Mas91c]. Praise [Ano92-28, Nan91i, Ref94]. Pre [Ano94-124]. Pre- [Ano94-124]. precious [Way90a]. Precision [Ano92-149]. prediction [Rya93b]. predictions [Ano91-54]. predicts [Cra90o, Ste94b]. Preemptive [Pou93n]. prefer [Del93a]. preferred [Mel93a]. Preliminary [Ano93-66, Ano94-107, Gre93d]. Premiere [Yag93h, Ano92-206, Lin94, Yag93h]. Premium [Wau92, Ano92u]. PReP [Ano94m]. preparation [LM93]. prepare [Ano93-32]. prepared [Ras91d]. prerelease [Ano93i, Ano94w]. prescribes [Min92c]. Present [Bry94e, VC90m, Ano90x, Hay90a, Ude93m]. Presentation [Api94b, Nan90b, Yag92c, Ano91-123, Ano93p, ADE90, Cla93c, Cla93e, Cry94, Die91, Ras90d, Yag92d, Yag92e, Ano90-85, Min90d]. Presentations [Ano94-89, Cla93c, Cla93e, Cry94, Yag91a, Yag92e, Ano91-66, Ano93-32, Ras90k, Way94h, Yag91d, Yag92c, YG92, Ano93-137, Cla93c]. presenters [Cry94]. Presenting [Ano91a, EB90b]. presents [Ano92-133, Ano93-64, Cla92a, Cra90r, Lin93, Smi90c, Sta91a, Ste91b]. preserve [Ano94-96]. Press [HM94]. Presses [Ude93j]. Pressure [Ano94-92]. Pretty [Sta94a]. prevail [Rei94b]. prevent [Ano94-150, Gre91c]. Preventing [Ano91-169]. Prevention [Ano94-149]. preview [Ano93c]. previewing [Ano91-74]. previews [Pou93d]. Price [Alf91d, Ano91-170, Ano91-183, Car91a, Smi90g, Smi94g, Tho92a, Ano90-67, Ano91-76, Ano91-159, Ano92-53, Ano92-122, Ano94-87, Gre93d, Hal93d, Hal93e, Mit90a, Per94a, Ras91j, Rya93b, Smi94c, Tho93b, TH93, Tho93g, TR94a, Van94a]. price-conscious [TH93]. Price/Performance [Ano91-170, Ano91-76, Ano94-87, Smi94c, Van94a]. priced [Ano91-160, Ano92-110, Per94a, Rei92b, Van94b, Ano91-160, Per94a]. Prices [Ano94-31, Ano94-60, Hal93e, Yag91h, Ano94-106, Ano94-129, BL90, LB90, TH93, VC90e, Yag90l]. Pricey [Mia90e]. primary [RP90, Tho91a]. prime [HS94]. Primed [Hog90]. Primer [Ano93-89, Har90b, RNG94]. Primera [Ano93-128]. primers [KUD92, Tho93b]. primitive [Pou90g]. Princeton [SE91]. Principle [Pou93s]. principles [Pou90i]. Print [Ano92-150, Ano93-92, Nan91a, Smi94e, Way93f, Ano92-170, Len94b, Mel93a, Rog91]. printed [Vau90]. Printer [Alf91c, Ano91-104, Ano92-50, Ano92-224, Ano93-114, Ano93-128, Ano93-142, Joc90b, Lov92b, Per94b, Tho93g, Vos90a, Alf91b, Ano90-48, Ano90-67, Ano91-48, Ano91-75, Ano91-193, Ano92-43, Ano92-73, Ano92-152, Ano93-70, Ano93-76, Ano93-131, Ano94-104, Cry92a, DE90, DE91a, FJK^{+}93, Hay93, HM94, Joc90a, Joc91, Smi94e, Tho92b, Tho92g, Tho93d, Way93f]. printer/plotter [Ano92-152]. Printers [Ano92-88, DE90, Egl92b, FJK^{+}93, Hay93, HM94, HHP93, Joc90a, KM94, Mel93b, Rei94g, Tho93b, Ano91l, Ano91-27, Ano92-90, Ano92-153, Ano93-36, Ano93-54, Ano93-129, Ano94-34, Ano94-79, Bar91e, Cra90c, DW90, Egl90c, Fie91m, Joc90c, Led90, Len94b, Mia92b, Smi94e, Tho92a, Van93b, WE92]. printf [Ano91-129, Ano91-129]. Printing [Alf91a, Fie91m, Hay93, Joc91, Lov93a, LS91a, Smi91b, Ano91-94, Ano91-118, Die90a, Nel91a, Rog91, Tho93b, Tho93g, Vos90a, Zei93]. printout [Sty91]. Prints [Tho92g, Ano94-175]. prior [Ano94-69]. Priorities [Ano94-117]. Prioritizing [Ano91-130]. Privacy [Buc92, Sta94a, Sta94a]. Private [Ano90-82]. PRML [Ano94-139]. Pro [Ano91-49, Ano92-55, Ano93-107, Gas93, Mia94b, Ref94, Ras90e, Tho93c, Ano92i, Ano92-46, Ano92-54, Ano93-44, Mia94b, Pep91, Van92b]. probably [Ano93-66, VN90d]. probing [Min90j]. Problem [Smi94f, Ano92-91, DH92, Hog90, HH94b, MH91, Pou90p, Ste94a, Ste91e, UY91, Way93i]. Problems [Ano91-155, Fis91c, Gre91c, Nan92d, Ano92-160, Ano92-216, Ano93e, Ano93f, Ano94-94, Ano94-169, Api94d, CV90a, Chr94, Cra91f, Cra91j, Gre93g, Hol90, Mil94a, Min93a, Smi94f, Sta91a, Vor92, Way91c, Wsz91]. procedural [DH90]. Procedure [Nan91i, Kor93, Mal91a]. procedures [HC93]. Process [Cla94a, Wal94e, Ano92-117, Ano94x, Ano94-64, Ga 93d, Gre93a, Kay94a, Web92, Cot93b]. Process-analysis [Wal94e]. processes [Cla94a]. Processing [Alf92f, Ano92-41, Ano92-183, Ano93-48, Ano94-39, Cot93b, San90b, Smi90e, Woo90d, Ano91-138, Ano92h, Ano92-70, Ano94-82, Ano94-131, Ano94-181, Cas92, Cra90f, Gri94, Har91a, HR91, Lan90a, Len94a, Mia93a, Pou90h, Sha94c, Tho93b, Van94a, Vor93, Way93f, WK90]. Processor [Rya91f, Alf92d, Ano90-40, Ano90-49, Ano90-75, Ano91-87, Ano92g, Ano92-57, Ano92e, Ano92-202, Ano92-226, Ano94-81, Cah90, Ga 94d, Gar94, Gre93d, LAD91, Pou90f, Pou94b, Pou90r, Pou93k, Rya94b, Smi93c, Tho90f, Tho94b, Tho94h, Ude90e, Wal93, Wau93c]. Processors [Ano92-157, Ano93-63, EWT90, Gre94d, Smi93e, Ung91, Ano91w, Ano91-133, Ano92-41, Ano93b, BEW92, Ga 94d, Gar94, Hay94b, Mia93a, Pou93b, Pou94g, Sca92, Way93f, Woo90d]. Procomm [Ano92-151, Ano93e]. procurements [Fis90a]. Prodea [Gil94]. ProdeaSynergy [Gil94]. Prodes [Gil94]. produce [Ano91-59, Gre94e, Hay94b, Mel93b, Pou93c, Ras90d, Way93f]. producers [Ano91-101]. produces [Ano90d, Ano91-62, Kea91a]. producing [Ano91-66]. Product [Ano92-28, Ano92-27, Api90b, AWGY90, AG91, BGE^{+}92, BEW92, DE92, EW92, GE92, GD92b, GW92, Hed94, Ano90-108, Ano92-100, Ano93g, Ano94-57, Ano94-127, Ano94-162, Api94b, Cra90j, Egl94c, Gas93, Kay94c, Lav91, Nan93h, Pou90m, Ras91j]. production [Ano93-134, RNG94, Yag92b]. productive [Har91c, Yag90f]. Productivity [Ano92-194, Ano91-39, Ano91-72, Ano92-181, Ano94-145]. Products [Ano90l, Ano90a, Fid92, Ano90j, Ano90s, Ano90-35, Ano90-53, Ano90-65, Ano91-101, Ano92a, Ano92b, Ano92-157, Ano92-195, Ano93-94, Ano93-100, Ano94s, Ano94p, Ano94q, Ano94-61, Ano94-69, Ano94-72, Ano94-117, Ano94-158, Ano94-147, Ano94-149, DE91b, Egl92b, Egl94c, Fin90a, Ga 93d, HS94, Hay94b, Laz93, Len93b, Mal93, Nad93, Nad94a, Pou91o, Pou91u, Pou92i, Ude93g, VC90q, Var94c, Woo91b, Kli90a]. Professional [Api93c, Bar92d, Cra91i, Egl93d, FH93, Fri90, Ste90, Yag91f, Ano92-46, Ano92-146, Ano93-82, Ano94-68, Cra90n, Egl93b, Laz90, Lin94, Ras90k, TG91, Wig90, Har93, Min93b, Sch92c, Woo90d]. professional-level [TG91]. Professionals [Van94a, Ano90-56, Ga 93d]. Professor [Ano90-100, Ano91-179]. Profile [Ano92-175, Fis90a]. Profiles [CS92]. profit [Rei93a]. profound [Ano94-65, Tho94e, Day90]. Program [Ano91-177, Ano90-45, Ano90-59, Ano90-70, Ano91-51, Ano91-62, Ano91-136, Ano91-184, Ano92-66, Ano92-117, Ano92-151, Ano92-154, Ano92-191, Ano92-197, Ano93r, Ano93-43, Ano93-69, Ano94w, Cal91a, GFK94, Hel91e, KPUG93, Pou90w, Pou94o, SH91, Sum90, Tho94k, Wal94a, Way90a]. programmer [Ano92y, Ano92-145, Sha94c]. Programmers [Kor93, Ano91-32, Ano91-177, Bar93c, Gre94b, Mas92, Nan91h, Nic93, ST93, Smi94b]. Programming [Ano91-56, Ano91-182, Ano92-191, Ano92-197, Bet94, Cra91c, Dlu91, Fri92, Ga 94c, GTN94, Gri94, Hel90d, KGTD92, LAD91, Mas92, Pou90i, Pou91f, Pou94d, Pou93k, Sha94b, Sin94, Url90, Var94c, Ano90s, Ano90-116, Ano91-126, Ano91-163, Ano91-180, Ano92-84, Ano92-150, Ano93-68, Ano94g, Ano94-29, Ano94-54, Ano94-125, Ano94-177, Api94b, Chr94, Dej94, Egl90a, Egl94g, Fic90, GFK94, GR90, Gre90e, Hel91c, Hel92a, Hol91, Ken93b, Min92c, Mul94, Nan93f, Par92, PS94, Pou94h, Rei92c, Sch90c, Smi94b, Spi90b, Ude93g, Way91b, Woo91b, Yag93k]. Programs [Ano91g, Ano91-136, Pou90h, Pou93o, Tho93f, Yag92e, Ano91-33, Ano91-43, Ano91-109, Ano91-177, Ano92j, Ano92-28, Ano92-44, Ano92-157, Ano92-213, Ano94l, Ano94f, Ano94-68, Ano94-71, Ano94-86, Ano94-131, Ano94-143, Ano94-148, Cla93c, Cri93, Egl91a, Egl94b, Fie90d, Fie90i, Fie91b, GE92, Gre94c, HWS91, McG90b, Mia90i, Mit91c, NTS93d, Pou93c, Ref94, Sin94, Smi94b, Ste94b, TG91, Tho94k, Tho94h, Ude91c, Vos90a, Wig92, Yag93a, You93, van91]. progress [Ano91-68]. Progression [Ano93y]. project [Ano91-67, Ano93-134, Ano94m, Bar93b, Tho93c, Yah92]. projections [Ste94b]. Projects [Hed93b, Ano93-68, Ano94x, Gil94, van91]. Proliferate [Gre94d]. proliferating [Ano92-133]. proliferation [Coo91]. ProLine [Ano90-83]. Prometheus [Die93b]. Promise [Ano91-69, Bar92a, Bry94e, Ude91c, Ano91-68, Ano92j, Ano92-189, Ano92-204, Ano93i, Egl91a, Egl93a, Har91b, Hol94, Kay94a, NHW93, Ude90b, Ude93m, VN90e, Way94i, WS91, Wsz91, Yag91c]. promised [Ano91-37]. Promises [Ano90-84, Ano91-62, Ano93-40, Fis91c, Len94a, Bry94e, Hal94i, Hel90a, Spi90b, Way93c, Woo90a]. prompts [Pou90k]. proper [Fis91b]. prophet [Ano90x]. proponents [Rei94b]. proposed [Ano94-72]. proposing [Ras91d]. proposition [Ras92d]. proprietary [Ano91-50, Ano94m]. Pros [Ano91-90, TG91, Ano91-188, KYCD92]. ProSignia [Nan93b]. prospects [Pou90p]. ProSpeed [Nad90d]. prosperity [Lu93a]. Protable [Len94b]. protect [Way93c]. Protected [Gla90e, Gla90b, Ano92-44, Sch90b]. protected-mode [Ano92-44]. Protection [Ano93-103, Fie91k, Sta94a]. protects [Ano92-126]. protocol [LDH94, Rei94b, Ude93b, Ano91m, Hay93, Smi94d]. Protocols [Fis91a, Bry94c, Daw94a, Kor93, SU93, Smi94f, Ude94d]. Prototyping [Woo91a]. ProTracer [Ano92-152]. Prove [Ano92-121, SE93a]. proven [Ano94a, Bar90k]. Proves [Ano94a, Cra90f, Rya90g, Ude90d]. provide [Ano90-54, Ano92-79, Ano92-115, Ano92-126, Ano94-166, Ano94-181, Bar90k, Dav90, ES94, Ess94, Gre93a, Hay93, Kea91b, Nad94a, Ras90c, Rei91a, Rya91a, Smi94a, Tho92h, TR94a, Ude94k, Van94a, VC90e, Wal94e, Yag90h, Yag91d]. providers [Ano94-77, Joc92a]. Provides [Ano94l, Ude91d, Ano90-34, Ano91-157, Ano92-223, Ano93-132, Ano94-184, Ano94-183, Cra90b, Fri94c, Gib90b, Ken94, Mia90e, Mil94b, Per94b, Pie94, Ste90, The94, Tho90b, Tho91e, Tho93c, Way94a]. Providing [Slo91, Fis91b, Sch92c, Smi94e, VC90g]. Proving [Ano94-132]. prowess [Cra92a]. Prowling [Fie90k]. PS [Ano91-131, Ano92-153, Ano92-219, Mit91b, Yag92g]. PS/2 [Ano91-131, Mit91b, Yag92g]. PS/2s [Ano91-131]. Psion [Pou92a, Pou92b]. Public [Ano91-132, Ano91-147, GB93, Sum90, Chr94, Sta94a]. Publish [Lu93b, Mur93]. Publisher [Ano90-111, Ano92-38, Ano92-143, Ano94-131, Vor93]. publishers [Cra90n]. Publishing [Ano91g, Ano92-38, Ano93a, Pou94c, Pou91k, Ras90e, Ude93d, Yag92f, Ano91-51, Ano91-109, Ano92-58, Ano93-81, Ano93-138, Bry93a, Coo90c, Cra90n, Fin90b, GTN94, Lu93b, Ude93j, Ano90-81]. pulls [Ras91h]. pumped [Egl94f]. Pumping [Ano94-140, Bry93b, Lov90a]. pumps [Nad90a]. punch [Fie91m, Tho94h]. Purchasing [Ano94-141]. Purpose [HK94, Ano94-90, Egl93a, HHP93]. Pursuit [Pou90q, Pou93s]. Pursuits [Sel92]. Push [Ano93m, Api94a, Bon90c]. Pushes [Ano94-82]. Put [Ano94-44, UPSW94, Ano90-27, Ano94-165, Api91a, Api94d, Bar93a, Gre90j, Yag91c]. Puts [Mia90d, Sch90c, Ano91-88, Ano91-89, Ano92-78, Ano92-114, Ano92-221, Ano93-127, AWGY90, Die90c, Nad90f, Rya93b]. Putting [Ano91-133, Bar90b, Bar93d, Ras91o, SE93b, Sty91, Ude91f, Hed93b]. puzzle [Gre90a]. Puzzles [Ano92-60]. Pyramid [Ano92-102]. Pyramids [Top91]. Pyrotechnic [HK94].

Q [Api94c, Mia90d, Pou91n]. Q/Cor [Mia90d]. QEMM [Ano91-162, Gla91a]. QEMM-386 [Gla91a]. QFA [Ano91-134].

QFA-700 [Ano91-134]. QIC [VN90b]. QMS [Ano90-85, Ano92-153, Ano92-219, Ano93-92, Joc90b, Tho93g]. QMS-PS [Ano92-153, Ano92-219]. QMSWriter [Ano90-85]. QNX [Var94b, Yag90a]. Quad [Ano94-87]. Quaderno [Ano92-128]. Quadra [Tho94f, Ano92m, Ano94-107, Tho92h]. quadruple [Egl93e]. Quake [Pou94m]. Quality [Ano90-60, Ano91l, Ano93-49, Ano94-89, Ano94-142, FJPR93, HM94, HHP93, KM94, Die90a, Egl93c, Len94b, Lin91a, Mel93b, Mia92b, Pou93c, Ras90d, Tho90h, Zei93, Lov91a]. quantities [Ras90f]. Quark [Van93c]. quarter [Sal94]. Quartet [Ano94-184]. Quattro [Gas93, Ano92-55, Pep91, Ref94, Rei90a, Van92b]. query [Obe90]. QueryDOS [Ano92-173]. Quest [Ano90-86, Cra91j, KDG^{+}93]. Question [Buc92, Pou93i, HC93, Lip90a, Lu93a, You90]. Questions [Fie90a, LM93, Way94k, Ano92-83, Ano92-215, Hed93b, Led90, Min91d]. Queue [Ano92-150]. Queuing [Sta91a]. Quick [Dlu91, Ano90d, Ano91-123, Fie90c, Var94c, Woo91b, WS91, Yag92j, Yag93i]. quick-and-dirty [Woo91b]. QuickBasic [Pou93k]. QuickC [Ano92-197]. QuickDraw [Oth92]. Quicken [Ano92-154, Kof93, She93b]. Quicker [Kof93]. quickly [Ras90k, Smi94f, Way93f]. QuickRing [Ano93-93]. Quicksoft [Ano90-75]. QuickTime [Ano91-66, Ano92-206, DB93, Yag93h]. quiet [Ano90d]. quietly [Yag92h]. Quill [Tho90a]. Quite [Fie90j, Alf91d, HPPR93, Wsz93a]. Quo [Rei90a].

R [Ano90-87, Tho94c, UE93]. R.I.P. [Hal94d, Sma91a]. R4000 [Ano91-87, Ano92-57, Ano92-155]. R4000s [Ano92-57]. Race [MW94, Bry93b, Sch92c]. Racet [Gre90i]. Racing [Tho92h]. rad [Ano91d]. radiation [Saf93]. radical [She93a, Ude94j]. radically [Ano90-62, Ano92-144, Pou93d, Pou94g, Way94i]. Radio [Dav90]. Radiosity [Ano92-156]. Radius [Ano90-77, Fin90b, Tho92h, Ude90e]. rages [Ano92i]. RAID [Alf92e, And90, And93, Ano93-126]. RAID-5 [Ano93-126]. Raidion [Ano93-132]. Raise [Ano91-159, She92, Yav94]. Raises [Sch92c]. Raising [Ano92-44, Ano92-157]. RAM [Api94d, Cal91b, Mia90i, VC90c, Woo90a, Wsz90]. RAM-Cram [VC90c]. RAM-Crammed [Woo90a]. random [Min92d]. range [Ano91-80, Ano94-147, Tho93g]. RangeLAN [Ano92-219]. RangeLAN/ISA [Ano92-219]. ranging [Ano94-81, Bar93b, Egl90d, KM94]. ranking [Len94a]. rankings [Fox93, KM94]. rapidly [Api94d, Mar90b, Swa93, Tho92d]. Rapport [Ano94-181]. Rare [Len94a]. Rash [Ras90l, Ras91b]. raster [Ano90-32, Lov91d]. RasterOps [Ano90-93, Ano91-131]. RasterPro [Die90a]. rate [Api94a, Len94b]. Rated [Mar92b]. rather [WK90]. ratio [Ano91-76]. Rational [Hel90a]. Raw [Ano92-146]. Ray [Gla90h, Ran90]. Raymond [Ano90x]. rays [Gla90h]. RDI [Api94d]. Reach [Alf91c, Ano91-135, Egl90d, Ple94, Ano92-145, Nan93c, Nov94b, Ude93m]. Reaches [Ano91-124, Ano93w]. react [Ano93-57]. Reactions [Ano94-99]. Reacts [GB93]. read [Ano94f, Way94c]. Reader [Ano91a, Ano92b, Ano93-94, Ano94b, Ano93-60, Led90, Min90d]. Readers [Ano93-58, Ano93-57, Ano94-100, Ano94-101, Ano94-102, Ano90l, Ano91a, Ano92b, Ano93-54, Ano93-94, Ano94-31, Ano94b]. Readies [Ano94w]. Reading [Ude94b]. reads [Ano90-69, Ano93-133, Egl93b]. ready [Ano92-81, Ano94-159, Ano94-180, Fis90c, GN93, HS94, Hol94, Pou94e, Ras91h, VN90e]. Real [All93j, Ano92-127, Ano92-158, Ano92-223, HS94, Rya91d, Ste91b, VN90c, Way94d, Ano90-69, Ano90-84, Ano91-180, Ano91-182, Ano91-183, Ano92-135, Ano92-178, Ano92-159, Ano93-116, Ano94-103, Ano94-183, Ga 93b, Hel91a, KM94, Kay94c, Laf91, Orr94, Pou94d, Smi90a, Ups90, Ano92-135, Ano92-166, Ano92-159, Egl92e, Laf91, Way94d]. Real-time [Ano92-158, Way94d, Ano92-135, Ano92-127, Ano92-178, Ano92-159, Ano94-183, Laf91, Orr94, Pou94d, Smi90a, Ano92-135, Ano92-166, Ano92-159, Laf91, Way94d]. real-world [Hel91a, KM94, Kay94c, Egl92e]. Realism [Gla90h, Lov91c, Ano92-142, Cal91b, Lov91c]. Realistic [Ano92-167, Ano92d, Ano92-142, BF90, Cal91b, Cra90m, Ups90, Yag93i]. Realities [Cra91d]. Reality [Ano91-41, Ano94s, KCU^{+}93, Ano90-38, Ano92-40, Ano92c, Ano94-98, Ano94-174, Ebe93, Egl91c, Egl91e, Har91a, Hed93b, NHW93, Rei94a, Way94h, Yag91f]. realize [Ano92-194]. Realizer [Ano91-177]. really [Fis91a, Law94]. realm [Sta94b]. realms [ASS92, Pou94d]. Reason [Yag90d, Tho94i]. Reasonable [Tho92a]. reasons [Ano94-103, Kay94c]. rebound [Pra93]. rebuttal [Ano92-84]. receive [Ano92-45, Way91d]. receives [Ano93-85]. recently [Ano94n]. reception [Ano94-166]. Reckoned [Ude90d]. Reclaiming [Ano92-160]. Recognita [Egl94c]. Recognition [Bar92f, Cra90d, Egl94c, Way93a, Way93i, Ano92-61, Ano92-144, Ano94-81, Ano94-162, Pou93a, Tho94f, WS91]. recognized [Ano94q]. Recompile [Ano94-137]. reconfiguration [Hal94i]. reconfiguring [Api94d]. Reconstructed [Ano91-83]. Record [Wal94b, Api94c]. Recordable [Hyo94, Hyo94]. Recorders [Yag93l]. Recording [Tho94c, Fri90, YG92]. Records [Ano93-105, Buc92, Wal94b]. Recovery [Pra93]. recreational [Ano91-126]. recurring [Cra91b]. recursive [NTS93b]. red [Ano94-172, Hel90e, Ano92-57]. red-hot [Ano94-172, Ano92-57]. Redefine [Ano92-53, Ano92-40, Ano94m]. Redefines [Alf92g, Ano94m, Tho94b]. redefining [Nan92c]. redesigned [Ano90-90]. Redesigns [Ano91-112]. RediDockit [Ano94-181]. Redirected [Ano94-143]. reduce [Wal94b]. reduced [Ano94-129]. reduces [Ano93-132, Egl94g]. Reducing [Ano93-95]. Reeling [Gla90c]. reengineering [Ano93-100, Ano94x]. Refalo [Red91a]. Reference [Ano94m, Ano92-136, Mur93]. Refinement [Bar92a]. Reflect [Ano94-117, Ano94-84, Pet90]. Reflection [Ano90-82]. Reflections [Buc92, Gre93g]. refresh [Api94a]. regardless [Ano91-194]. register [Ano93-54]. registers [Rya93b]. Regulated [Way93g]. Reigns [Lov91c]. rein [Nic93]. Reinvents [Ano92l]. reinvigorating [WS93]. related [Ano90-58]. Relating [Ras92f]. Relational [VN90c, Ano92-59, Ano92-84, Ano92-205, Ano94-114, Cra90e, Mul94, SH90]. relative [THN93]. relay [MW94, Rya91e, MW94]. Release [Del93a, Han92a, Pep91, Ude93i, Ano90-110, Ano94w, Ano94-127, Api94c, Egl92b, Egl94f, Ham92, Ude93c, Vor92]. released [Ano94-126]. releases [Ano91-131, Ano93-38, Ano94-84, SY90, Van93a, Van93c]. releasing [Ano94-63]. Reliability [Alf90c, And90]. Reliable [Way93i, Api90c]. Relief [Dlu91, Yag90j, Ano92-172]. relieve [AG91]. relieves [SE94]. relying [Ric94]. remain [Ano92-75, Gre94a]. remains [Ude93b]. remarkable [Ano90-116, Ano91-153]. remarkably [Gre94d]. remembers [Pou90p]. reminders [Ano92-175]. Remote [Ano94-143, Fri94b, Gre91a, Mit91c, Nad94b, Nan91i, NTS93d, Ano92-85, Ano92-222, Ano94-148, Api91b, Daw94a, Fri94c, Kor93, Mal91a, Yag93e, Ano94-148].

Remote-Control [Mit91c, NTS93d, Ano94-148, Ano94-148]. remotely [Ano92-213, Ano92-220]. Removable [CW92, Lab92b]. removable-storage [CW92]. Remove [Ano91-136]. remover [Ano91-136]. Renaissance [WS93, Ga 93d]. render [Bri90]. Rendering [Ano92-167, Ano92-156, Cal91b]. RenderMan [Lov91c]. renewed [Ano94s]. repair [TGN94]. Repairing [Ano91-137]. repercussions [Sto93]. repetitive [VFGZ94]. Replacement [Len94a, Ano92-185, LM90, STA92]. Replication [Yav94]. Report [All93i, Ano92-28, Ano92-27, Ano93-47, Ano93-48, Ano93-96, Ano93-97, Ano93-98, Ano93-99, Ano93-100, Ano93-110, Ano94-89, Ano94-90, Ano94-87, Ano94-88, Ban92, BGE^{+}92, BEW92, Bar93c, DE92, EW92, FJPR93, FJK^{+}93, Fox93, Fri92, GE92, GD92b, GW92, HM94, HK94, HHP93, Hol94, HP94, Hur93a, Joc92a, Len94b, Len94a, Pra93, Ran93, Red92b, Rei92a, Rei93a, Tab93, LHR90, Min91b, Pou90n, Pou90v]. reports [Ano92-38, Ano92-108, Ste94b]. represents [Ano93-74, Egl94f, Kro90]. require [Rei94b]. Required [Ano92-191, Ano92-192, Bet94, Chr94, Cot93b, Gre92, Gri94, Hol93a, Hol93b, Jef93a, Kiy93b, Kiy94, MW92b, Nag92, Nan93f, NHW93, Pie93, ST93, Sin94, You93, Gre90f, Gre90g]. requirements [Hed94, TGN94, Wal94c, Way94i]. requires [Ano93-116, GD93]. res/low [Ano91d]. Rescue [Ras90c, Ano93-105]. Research [Cla91a, Fis91b, Sma92, Ano91-70, Ano92-207, Ano93-98, Pou94d, Ano90u, Ano91-40, Ano91-84, Pou91g, Pou91l, Smi91a, Woo90a]. Researchers [LHR90]. reservoir [Ste91b]. Reshaping [Ano92-161]. resident [Tho91c]. resist [Ano94-101, Ga 94c]. resistance [Min92b]. Resolution [Ano92-33, Ano92-50, Ano92-224, Ano94c, Fin90b, Van93b, Ano92q, Ano92-115, Ano94-135, Api94a, Cov90, Joc90a]. Resource [Ano91-138, Ano91-139, Ano91-140, Ano91-141, Ano91-142, Ano91-144, Ano91-143, Ano91-145, Ano91-146, Ano91-147, Ano91-148, Ano92-162, Ano92-163, Ano92-164, Ano92-165, Ano92-167, Ano92-166, Ano92-168, Ano92-182, Ano93-101, Ano93-102, Ano93-104, Ano93-103, Ano94-48, Gre92, VFGZ94, Nic93]. resource-hogging [Nic93]. resources [Ano90a, Ano92-182, Ano94-148, Fie91a, Hog90]. Respect [Ano91-62, Ano91-170, Ano93-97, Ano94-158, Ras91f]. respond [Ano93-58, Ano94-101, RT94b]. Responds [Ano92s, Ano92-113]. Response [Ano92-94, Ano94-112]. Responses [Ano93-59, Ano92-92]. responsibilities [Cos94]. Rest [Ste90, Tho93c, Ano94b, Bux90a, Rya94a, Smi90d]. restaurants [Ano91s]. resting [Hal93b]. restore [Ras90o]. result [Ano92-34, Ano94-129, Ano94-135]. Results [Hur93a, Ano94-107, Ano94-113, Egl93d, Gre93d, Tho93a]. rethinking [RH93]. Retooling [Tho93h]. Retrieval [Ano92-180, Ano94-78, Ano92-75]. retrieve [Ano94g]. retrieving [Loc91]. Retrofitting [Kog94]. retrospective [Ano92x]. Return [Ano90-88, Gla91c]. Reusable [Kiy93b, Cox90, Key92]. Revamps [Ano93c]. reveal [Ano93i, Hur93a]. Revealed [Ano92-191, Ano92-192, Gla91b, RT91, Tho91c, Pou92k, Pou93o]. reveals [Ano92m, Cra90n, Lan90b]. Reverse [Way94i]. Reversible [Way94i]. review [Ano91-110, Ano91-109, Car92, Egl93f, Fie90c, Per94a]. reviewed [Ras92a, SG92b, Ste94b]. Reviewer [Ano91-149, Car94, Egl92e, Gre93f, Gre93g, Joc92b, Lab92b, Lab92a, Tho93h, Ude93i, Ano92-173, Ano92-169, Ano92-172, Ano92-170, Ano92-171]. Reviews [KUD92, KGTD92, KYT^{+}92, KRCT93, KPC93, KCU^{+}93, KAY93, KGEM93, KDG^{+}93, Ano92-45, Bar91b, Wsz93a]. revise [She90b]. Revisited [Ano91-42, KPC93, Pou91j, Pou92o, Sto93, SH90]. Revitalizing [Nan91k]. Revolution [Lip90b, Pou91s, SH91, Web92, Ano92-161, Ano94u, Ano94v, Lan90b, Lu92, Mos93, Ude93d, UFC94]. revolutionizes [Cro93b]. Revs [Api91c]. revved [Yag90e]. revved-up [Yag90e]. rewards [Ham90, Sel92]. rework [Hay94b]. Rexx [Dan90]. Rhymer [Ano90-115]. rich [Fie91b, Kro90, Sei91]. Richard [Ano94-94, KYT^{+}92, May94b]. Riches [Min91b]. Ride [Ano92-174]. Rides [Lus90, Rya94a]. Right [Ano92-175, Red92a, Vas93a, van91, Ano91-99, Ano92-105, Ano92-192, Ano93-88, Api94a, Nan91e, Nan91l, Ras90b, Ude90a, Way90b]. Ring [Smi94f, Gla91c]. Rio [Yag92d]. RISC [Gre93d, Ano90-80, Ano91-87, Ano92-176, Ano93-75, Ano94-65, Ano94-82, Ano94-118, Ano94-137, Ano94-153, Api94e, Bar90c, BR93, Cra90c, Fie90l, Fie91m, Gla91b, Gre93d, Hal94c, Hal94f, Lan94a, Mar90a, Pou90e, Pou91a, Pou92d, Pou93g, Pou94b, RS90, Rya93b, RT94b, The94, Tho92e, Tho93f, Ude94k, Way94k, Yag90e, Yag90m, Yag91e]. RISC-based [Cra90c, The94, Tho93f]. RISC-CISC [Ano94-65]. RISC/CISC [Ano94-118]. Rising [Cra90o]. Risk [Ano93g]. risks [Rei92b]. rivaling [Ano94-33]. rivals [Ros90b]. rivers [Bar92b]. RLN [Ano92-177]. Road [Alf91e, Ano92-193, Ano94n, Ano94-87, Cra90i, Daw94c, Egl92a, Gea90, Pou91q, Pou92j, Tho92d, Tho92f, Way94g, Ano91-94, Ano93-127, Api91a, Api94d, Kiy94, Ras90l, Way90b]. Roadblocks [Fri94a]. Roads [MS93]. roaming [Pou94q]. Roared [Alf90a]. rob [Ano92-149]. robots [RUHG93]. robust [Ano92-123]. robustness [Lie93]. Rock [Gre93c]. Rocket [Hal93c, Tho92h]. rods [Cra90c]. Roger [Cra91j, Cra91j]. Role [Ano92-28, Fis91b, Nan92c, Rya91c, VC90c, Wat94]. Roll [Gre93c, Len94b, Lie93, GN93]. rolls [Hel90e]. ROM [Alf93, All93f, Ano91u, Ano92-60, Ano92y, Ano92-136, Ano93s, Ano93x, Ano94-31, Ano94-87, Ano94-152, Bos93, CWVS94, Cra91l, Egl93e, GT93, GFK94, GTGW94, GTN94, Hal93a, Har91b, KYCD92, KUD92, KGTD92, KYT^{+}92, KRCT93, KPC93, KCU^{+}93, KAY93, Ken93a, Ken93c, Min93b, Nov94a, Pou94c, Pou91p, Pou93q, Ras90c, Ras90f, Rei94a, RNG94, TGN94, Ude93d, Ude93h, Ude93j, UPSW94, UFC94, VFGZ94, Wau93b]. ROMS [GFK94, CWVS94, ESN94, GT93, GTGW94, GTN94, Ken93b, KGEM93, KDG^{+}93, Nov94a, Pou90q, Pou90t, Pou90v, Pou91m, Pou91p, Ras90f, Ras91a, RUHG93, RNG94, TGN94, Tho94c, UPSW94, UFC94, VC90n, VFGZ94, Ken93a]. roomful [Egl94d]. Roots [Ano92d, Ano90y, Ano92-30]. Rosie [Ano91-179]. Rough [Bar92a, Egl91e, Ano93i]. round [Hal93e, KM94]. Roundup [Ano94-147, Ano94-148, Ano94-149, HH94a, Lov92a, Ref94, Ano90-31, Bon90b]. routable [Ude94d]. Route [ATU92, Bry94c, Lie93, Pou93i, Ras90b]. Router [Ano93-95, Daw94a, Gib90a]. routers [Ano91-34, Daw94a]. routines [Jef93a, LB91]. routing [Ano94-97]. row [Gre93a]. royalty [Nan90a]. royalty-free [Nan90a]. RPC [Smi94b]. RPCs [Nan91i]. RS [Daw94c]. RS/6000 [Daw94c]. RTOS [Ano92-178]. rugged [LM90]. Rules [Ano94-87, Ano90t, Hol91]. Ruling [Hal93e]. Ruminations [KGEM93]. Run [Ano91-195, Ano94-113, Del93b, Ano92-111, Ano92-109, Ano92-188, Ano92-213, Ano94l, Ano94-126, Far90, Gut93, Hur93a, Lan94b, Min90h, Ols94, Tho94h, VC90g, The94].

Run-Time [The94]. Running [Fie90f, Ga 94d, MW94, Ano92-226, Ano93-110, Ano94-127, Ano94-148, Fie92a, Tho90d, VC90p]. Runs [Lin91b, Ano94-64, Gre93h, Rei91b, Yag90c]. Russia [Ano94-146]. rustles [Red91a]. Rx [DWKS91].

S3 [Ano92-171]. S4000 [Lin90b]. S486 [Alf90c]. S486/MC33 [Alf90c]. SAA [Osm90]. sacrificing [Cra90a]. Safari [Mit91b]. Safe [Hal94a, Ano92-189, Ano92-56]. Safer [DWKS91, CV90c, Har91c, Min90b]. Safety [And90, Ano94-73, Har91c, Kre91, Nel91b, Ste92]. saga [Pou90n]. sage [Nan91h]. Sake [BGE^{+}92]. sales [Ste94b]. Same [Ano94-151, Cra90q, Gil94, Gre94b, Lie93, Min90i, Rya90c, Way94h]. sample [Ano90-65, Ano92-212, GD92b, Sum90]. Sampler [Ano91-143, Tho94d]. sampling [Ano90j, Ano90-29, GR94]. sanity [Ras90o]. sans [STS92]. Santa [SY90]. Sao [Ban92]. satin [Nan94b]. Save [Ano91-39, Ano91-71, Kor93, Rei94a, Tho94h]. saves [Egl94g, Rya91c]. Saving [VN90d, Ano91-108, Ano91-166, NTS92a, Pou92d]. savings [Ano93-91]. Savvy [ST93]. Say [Ga 93e, Sma91b, Ano93-93, Ano93-138]. Saying [Ano94-173]. Says [Ano94-145, Ano91-115, Ano94-112, Cri93, YS92]. SBus [Ano92-174]. scalable [Rob91a, Rya93b]. scale [Ano90-32, Ano93-125]. scales [Ude91f]. Scaling [Ano92-102, Ste91c]. Scan [Ano90-42, Ano94-182, Len94a, Egl93b]. scan-in-Dex [Egl93b]. ScanMan [Ano90-89, Ano90-90, Ano92-179]. Scanner [Tho90h, Ano90-42, Ano90-89, Ano90-90, Ano92-43, Ano93-125, Egl93b, Van94a]. Scanners [Ano92-71, Ano93-105, Ano92-179, DE92, Van94a]. Scanning [DE92, Ano92-71]. scans [Ano94-175]. scattered [Loc92]. scene [Gla90h, Pou92m]. Scenes [Egl90a, Joc92b]. Schedule [GETY91]. Scheduler [Yag90i]. schedulers [GETY91, SE94]. Schedules [Ano91-155]. Scheduling [SE94, Ano93-134, Ras90o, Yav94]. schmopen [Ano91-115]. Scholar [Wal93]. School [Pou92h]. schools [Pac94]. Science [Ano91-102, Cra90p, Ano90-36, GFK94, KRCT93, Pou90v]. scientific [Ano92-218, Cra90p, GTGW94]. Scientists [Ano91-150, Ano91-70, Ano92-161, Min90j]. Scifert [Sei91]. SCO [Fie91i, Yag90j]. scramble [Kay94d]. scratch [Smi94b]. screamer [Ano93-92, Way94j]. Screaming [VC90m]. screen [Ano91-125, Ano92-218, Ano93a, Min92d, Sty91, Vau90, Api94a]. screens [Ano94-35, Ano94-135, BS91, Lov90b]. ScreenStar [Ano93-133]. Script [Woo91b, Ano94-181]. Scripting [Ano91-182, Ano91-52, Bet94, Cra91g, Egl94b, Rya90b]. Scripts [Rya90b, Ano91s, Fie91f]. Scrutiny [Del92]. SCSI [Ano91-76, Ano94-94, Ano94-144, Api94d, Gla90g, Gla90d, Gre90e, McN94, Pou94n, San90b, Tho94d, Van90]. SCSI-2 [Ano91-76]. SD [Tho92g]. SDK [Ano93-41]. sea [Ano92j]. Seamless [Ano94f, Yag91h, Ano92-186, Dej94, Yag91g]. seamlessly [Ano94h, Hal93b, Len93b]. Search [Ano92-180, Ano90-86, Ano92-75, Ano94-69, Cla93b, KYT^{+}92, Nan91e, Pou93k]. search-and-retrieval [Ano92-75]. searchable [Egl93b]. searches [MW92b, Way91c]. Searching [GD92a, Har93]. Season [Ano91-115]. Seat [Tho93d]. Second [Ano91-93, Ano94o, Yag93h, Ano92-97, Ano92-199, Cra91h, Ras91d, Ude93d]. second-generation [Ano92-199]. second-guessers [Ras91d]. Secrecy [GTGW94]. Secret [Hal94f, Gre90a]. Secrets [KGTD92, Lan90b, Mat94b, Pou93q, Ano91-47]. Secure [Kay94b, Ano90d, CV90c, Ras90g]. Securing [Kay94b]. Security [Ano94-161, Bor93, Min90f, Sch92b, Sta94a]. See [Ano92-216, Cas92, Hed93b, Len94a, Ano92-134, Ano94-125, Hol94, Ras90g, Ras91o, Smi90d, Vau90, Web93]. Seeing [Web93]. Seek [Ano93-84, Ano94g, Ano94-54, Gre94d]. Seeking [Bor93, Pou90r]. seem [Min92d, Nov94b]. seems [VC90n]. seen [AD91, She90a]. sees [Pou94i]. segments [Min90c]. select [Ano92a, Wsz93b]. Selected [Ano94-74]. Selection [Obe90, Ano94-159, Kay94c, UFC94]. selections [Ano92-74, KGTD92]. selects [CW92]. self [Api91c, Ken94, Sin94, Ano90-39]. self-configuring [Ken94]. self-modifying [Sin94]. Self-Similarity [Ano90-39]. self-test [Api91c]. sell [Red92b, VC90i]. sellers [Ano94-79]. selling [Ano92-226, Wau93a]. Semiconductor [Ano90-91, Rob90b, Rya90f, Ano93-93]. Senator [Ano91-98]. send [Gre91a, Rei94b, Way91d]. sends [New93, Pou92h]. Senior [Ano91k]. Sense [Ano92-101, Mia92a, Ano92-125, Ano94-96, Nan94d]. Senses [Ano92-79]. sentences [Ano92-117]. separate [Ano94b]. Separated [Rya90c]. separately [Ano94-76]. separations [Ano91-62]. Sequel [Ano92-97]. sequencers [YG92]. Serial [Cla94b, Ano94-72, Ano94-129, Bry94c, Kre91]. serial-bus [Ano94-72]. Series [Tho90e, Ano90-116, Ano91y, Ano92-29, App91, Egl93c, EB90b, Gre90d, Smi91a, Yag90a, Yag92b, Ano91-60, Pou92b, Smi94g]. Serious [Tho94i, Yag90b, Ano90b, Cla93c, Cra91g, Del93b, Egl94c, Egl94f, Smi94c, Ude93h, Yag90e]. Seriously [Cla94b]. serve [Ano91-142, Ano92-135, Gib90a, Sch93d]. Server [Ano94-75, Ano94-153, Daw94b, Ga 94d, Nan93b, Smi94b, Smi94e, Ude94k, Ull93, Wsz91, Ano91-76, Ano91-99, Ano92-119, Ano92-170, Ano92-200, Ano92-199, Bor93, Bos93, CV90d, Cro93b, Egl94e, Gre92, Gut93, HH94a, Kor93, Nan91i, Nan94d, Pou94m, Ryd93, Sch93b, UM90, VC90r, Nan94d]. Server-based [Wsz91, Ano92-119, Ryd93]. Servers [Ano91-99, Ano92-188, Ano93-120, CV90b, PGRA92, Ano93-75, Fin90a, GAW93, HC93, Nan93b, Ras91a, SG92a, Ste91a]. serves [BBWR90]. Service [Hov93b, Ude93h, Cra90a, Jef93a, Jof93, Van94a]. ServicePak [Min93b]. Services [Ano94-166, Ban91, Ano92-75, Ano92-85, Ano93m, Ano94-69, Ano94-77, Ano94-149, Die94, Fri94b, HS94, Kiy94, Mil94b, Rya91e, Slo91, Ude94k]. Serving [Ano93-106, CV90d, Nan93e]. SEs [Cra90m]. session [Ste94b]. sessions [Egl94b, Hyo94, Min90h]. Set [Ano91-170, HP94, Sch90a, Ano91-35, Ano91-68, Ano91-133, Ano92-63, Ano94e, Ano94-77, Ano94-117, Ano94-162, Ano94-166, Api94b, CP94, Cou92, Egl94f, Fie90f, Fie92a, Gla90d, Kli90c, Nan93d, Pou94h, Rei94f, Smi94a, Sta94b, Tho94b]. set-top [Ano94-166]. Sets [Ano90p, Ano92-105, Ano92-122, Ano93-68, RS90, Rib93, Wig90, Yag91a]. Setting [Fie92b, Hel91e, Ano92-113, Ano94s]. Seven [Ano91-151, Mia92b, Ras92a, Ano92-31, Ano94-149, Bon90b, Egl94b, Egl94g, Tho92a]. Several [Ano94-143, Pou94f, Ano94k, Ano94w, Ano94-33]. SGI [Ano92-57, Ano92-155, Smi94c, TS93a]. SGML [Ano92-90, Ano92-181, Ano93-56]. shake [Cry92a]. Shakespeare [Pou90q]. Shakespearean [May94b]. shame [Ano90-27]. shape [UMR^{+}90, Way92a]. shaping [Ano92-104]. Share [Yag93e, Ano94-102, Cra90h, Gre91a, Ken93c, Mia94b, She90a]. Shared [Ano92-182, Gre92, Bry94d, Egl94e, Gib90a, Hog90, Rob91a, Ste91c, VN91a]. shared-bus [Bry94d]. shared-memory [Bry94d, Rob91a, VN91a]. sharers [Alf91c]. Sharing [Ano91m, Ras91a, VC90n, Ano92-170, Ano94z, Bar90k, Gre93a, Mel93a, Rya90c, Tho90i, Ude93h]. sharp [Joc90a, Ano90-92, Ano93i, Ras90b, Tho90h]. sheer [Ano94-82]. Sheet [Pou94h]. Sheldon [Ano91k]. shelf [Gil94, Laz93]. shell [Ano90-52, Ano91-107, Hal90, STA92, Ude93e]. shells [Ras90p]. shield [Hal94f]. shift [Lip90b]. shine [BS91]. Shines [Ano92-224]. ship [Ano93-45, Ano93-66]. shipping [Ano93g]. Ships [Ano93-41, Ano94-183]. Shiva [ATU92, Gib90a]. shoebox [Ano92-114]. shoot [Ano92-146]. Shop [Hel91e, Mas90b]. shopping [Mia90g, UMR^{+}90]. Short [Ano91c, Ano93-107, Pou90f, Ano90-93]. shortage [Ras91g]. Shot [Yag90m]. Should [HP94, Rei94b, Ude94f, Way93g, Ano90-97, Ano93-29, Ano94-31, Ano94-71, Ano94-106, Ano94-125, Buc92, Cry92a, Van94a, Yav94, You90]. shoulders [Joc92b, Rya94a]. Shouldn’t [VC90p, Ano92-103]. Show [Ano94-83, Bar92a, Fie91j, Ano92-53, Ano94-35, Ban92, Mal93, Pou94k, Tho93i]. showcases [Ano92-85]. showing [Ano94-72, Cra90h, Mar92a]. showpiece [Ano90-99]. Shows [Ano91-58, Ude91c, Ano92g, Bon90b, Cra91c, FJK^{+}93, Gre90b, Sha94b, VC90d, Wal94e]. Shplinterface [Ano90-41]. Shrink [GW93, Yag90j]. shrink-wrapped [Yag90j]. Shrinking [Ano91-63]. Shrinks [Smi90g]. Shuffle [DH92, Egl93b]. Shuffles [Api90b]. shuttle [UPSW94, Ano92-217]. Sick [Wig90]. Sickness [Min92c]. Side [Ano90-94, Loc91, Ano90b, Cra90q, Die90a, Yag93l]. SideKick [Ano91-152]. Sides [Cra90q]. Sight [Egl91b, Yag92a, FT90]. sights [Wig90]. Sign [MPW^{+}90]. Signal [ASS92, Ano92-41, Ano92-183, Cou92, Way92b, Isl92, Sta90]. signal-computing [Cou92]. Signals [Bry92]. Signatures [Sch93c, Sch93d]. significant [Tho94j, Ude93i]. significantly [Ano92d]. Silence [Saf93]. Silicon [KDG^{+}93, Rya90g, SC92, Way94h, Way94i, Smi90f, Smi93c, TS93a]. Silver [Cox90, Nan94b, Nan94b]. silver-satin [Nan94b]. similar [Fis91a, HC93, Ref94]. Similarity [Ano90-39]. Similarly [Van94b, Ano92-110]. Simon [Ano94-145]. Simonizing [O’M94]. Simple [Egl93c, Ude93c, Ano90z, Bar93b, Daw94b, Egl94d, Egl94f, Hal94i, Nag92, Nan91f, Sta91a]. simpler [Sch93a]. simplicity [Ano92-87, Ano92-119]. simplified [Mul94]. Simplifies [Yag93i, Ano93-127, Mia90b, Smi94e]. simplify [Fis90b, Kay94a, Rob91b]. simply [Ano90-66]. Simpsons [Ano94-166]. Simulating [DH90, Gla90h, Pou90g, Cot93a]. Simulation [Dio94, Cra91a, Pou91k]. Simulations [Ano91-150]. simulator [Dio94]. simulators [Ano91-141]. Since [Egl94f, Ano94z]. Single [Yag90l, Ano92-223, Ano94-182, Api94b, GK91, Gre94e, Law94, Sta93, Ano91-153]. Single-Chip [Ano91-153]. Singular [Ano92-74]. SITcomm [Tho94i]. Site [KGTD92]. sites [Api91b, Gre91a]. situations [Api94d]. Six [Alf92b, Ano92-126, Ano92-204, Ung91, Ano91-121, Ano91-160, Ano92-190, Ano94-148, Egl93e, HM94, KM94, SE94, Yag90a]. six-part [Yag90a]. Sixty [Ano90-112]. Sixty-three [Ano90-112]. Size [Nan90a, Ano93-133, Ano94c, DWN91, Mit90a, Rei91a, VN90d, Ano94c]. Sizing [Fie90l, TS90]. Sizzles [Ano92m]. Sizzling [RS90]. SJ [Tho93b]. SJ-144 [Tho93b]. sketch [Sac90]. skill [Gre90a, Ken92, Ste91b]. skills [Smi94b]. skin [Hel90b]. SL-Enhanced [Len94a]. Slated [Ano94-127, Ano93-45]. Sleeper [Yag92h]. Sleepwalkers [Ano91-126].

Sleuth [Ano92-120]. slicing [Ref94]. slick [Ano94a, Yag92d]. slide [Ano94c, Way94h, Ano94c]. Slides [Ano91-106, Ras90d]. SlimLine [Ano91-106]. slimmer [Ano91-171]. slot [Bar90k, Wsz90]. slots [Sma92]. Slouching [Nov94b]. Slow [Ano94-79, Gre93e, Joc91, VC90a, Yag93e]. slower [Lin91b]. Slowly [Ano91-99, Cau93, VC90m]. slows [PTUM93]. SLT [Mit91a]. slug [Ano92v]. Small [Ano94s, Ano94-160, Api90c, Die93b, Gre94d, Mit90a, NTS93b, Tho93i, Var94a, Cra91l, Daw94a, Egl94d, Lin90a, Mia90e, Mia92b, Nan94d, Ras90l, Smi90a, Smi93b, Tho94h, Way94g, Ano94-165]. Small-Office [Ano94-165]. Smaller [Cla90a, Ham91, Rei90c, Ano91-63, Ano92-105, CV90c, Lan94b, Lin91b]. smallest [Ano91-193, Yag91e]. Smalltalk [Ano90-95, DG91]. Smalltalk-80 [Ano90-95]. Smart [Ano91-59, Ano93z, Fie91m, Gre91c, Hin92, KGEM93, Mac91, Pou90w, Way91c, Bar92g, Hed93b, Ano94c]. Smartcom [Ano93-131]. SmartConnex [San90b]. SmartConnex/ISA [San90b]. smarten [Cau93]. Smarter [Ano94x, Rei93d, Rei94g, Wal94e]. SmartIcons [Ano94a]. SmartLink [Ano93-130]. Smartly [Rob91b]. Smarts [Cau93, Rei94f]. Smile [Chi92]. Smith [Smi94c]. Smithsonian [Ano90-60]. Smoke [Bux90b, Bux90a]. Smooth [Cov90, Tho92e]. smoother [Law94]. smoothly [Fri94a]. SMP [Kog94, Ude94k]. SnapGrafx [Ano94-68]. Snapshots [KYCD92]. SNMP [Fis91a, SS94]. SNOBOL [Ano92-29]. Soars [Egl90b, TR94b]. sobering [Nan92b]. social [Bur92, Cos94]. Socket [Hal93c]. Sockets [Ano92-200]. SoftArc [Ga 93b]. Software [Ano90-78, Ano90-96, Ano91-141, Ano91-143, Ano91-145, Ano91-189, Ano92-28, Ano92-167, Ano92-184, Ano92-185, Ano92-186, Ano93-101, Ano93-104, Ano93-117, Ano94e, Ano94n, Ano94-146, Ano94-147, Ano94-148, Ano94-149, Ano94-186, Cri93, Eva94, Fie90c, Fie90d, HH94a, Kay94d, Kro90, Mar92b, Mia93b, Mit91c, Nan91g, NTS92a, NTS92b, NTS93c, NTS93d, NTS93a, Nic93, PS94, Ref94, STA92, Ude93g, Woo90c, You94, Ano90l, Ano90-55, Ano91-56, Ano91-68, Ano91-66, Ano91a, Ano91-158, Ano91-194, Ano92-34, Ano92-62, Ano92-61, Ano92-108, Ano92-203, Ano93p, Ano93-44, Ano93-51, Ano93-56, Ano93-64, Ano93-125, Ano94k, Ano94m, Ano94-29, Ano94-45, Ano94-54, Ano94-84, Ano94-63, Ano94-98, Ano94-102, Ano94-132, Ano94-131, Ano94-143, Ano94-169, Ano94-175, Ano94-176, Api94d, Ban92, Bar92f, Cot93a, Cox90, Cra90h, Cra90q, Cra91d, Cra91g, Cra91j, Cry94, DeH93, Dej94, Die91]. software [DH90, Egl93a, Egl94g, Fie90e, Fie90g, Ga 93d, Ga 94a, Gil94, Gre90c, Gre90i, Gru90, Hal93b, Hel90a, Kof93, LM93, LAD91, Lin93, Lov91a, Mil94a, Nak94, Nan91e, New93, Ols94, PTUM93, Pou90k, Pou90t, Pou91k, Pou92e, Pou93k, Ras90d, Rei92a, Rei94f, Rya91b, Smi94c, SE94, Ste94b, Tab93, Tho94c, Tho94f, Tho94h, Ude94g, VC90f, VC90g, Wau93a, Way93d, Way93f, Yag92a, Ano90-45, Ano94-189, Api94b, Nan93d, Ude93m, Yag90i]. software/hardware [Lin93]. SoftWindows [Ano94-157]. Solaris [Ano94-183, FH93, Smi93d]. Solbourne [Lin90b]. solely [Ano93-37]. Solid [Rei91a, FH93, MS93, Woo91a]. Solid-state [Rei91a]. solidifies [Cla93c]. Solution [Ano92-117, Ano94-28, Bos93, FH93, Lio90b, Lio90a, Pou93h, Ano90-83, Ano92-87, Ano94-149, Bar91e, Bry94a, Egl94e, Hog90, Lab92b, Pou90p, Smi93b, Ude91a, VC90l, Wal94e, Way91b]. Solutions [Ano91-95, Ano92-187, Ano92-188, Ano92-189, Ano92-190, Ano94-150, CS92, ES93, GW93, Lov93b, SG92b, SE93b, Wal94d, YG92, Ano90-54, Ano93e, Ano94-143, Ano94-189, CW92, Hol90, Ras91g, Ste91b, Ano94-157].

Solve [Ano92-216, Chr94, Ste94a, Way91c]. Solver [Smi94f]. solves [DH92]. solving [Way90b]. Some [Ano91-70, Ano91-195, Ano92-191, Ano92-192, Bar90k, Bet94, Chr94, Cot93b, Fie90a, Fie91l, Gre90f, Gre90g, Gre92, Gri94, Hol93a, Hol93b, Jef93a, Kiy93b, Kiy94, Lov93b, MW92b, Nag92, Nan92b, Nan93f, NHW93, Pie93, ST93, Sin94, Tho90d, You93, Ano91-56, Ano92-135, Ano93-69, Ano93-91, Ano94x, Ano94-65, Ano94-129, Bar91b, Bar92a, Cla93c, Cra92b, Fie91d, Hal94f, Hal94a, Han92a, Man90, Mia90f, Min90e, Min91d, Min92a, Nad94b, Pou90t, Pou91k, Pou94i, Ras90p, Sch93b, Smi90c, Smi94c, SE94, Sta91a, Sum90, Ude93i, Way93e]. someday [Ude93l]. Someone [Ano91-135]. Something [Ano91-185, Ano91-115]. sometimes [Gil94]. Somewhat [Ano94-158, Min90f, Pou93a]. Son [YS90, Ano91-58]. SONET [Cla92d]. Song [Cra90d]. Sony [Yag90m, Ano91-154, Ano93-108, Egl93d]. Soon [Alf91f, Ano92-134, Rya91d, Sac90, Ano92-48, Ano92-210, Ano93m, Ano93s, Ano93-29, Ano94-34, Ano94-166, Mar90a, Mar92a, Way94d]. Sooner [Fis90a]. Sophisticated [Bar92d, Kea90, Ras90d, Ano92-142, Ano93-116, Dej94, Die94, HH94a, Wal94a]. Sort [LB91, Ano93c]. Sorting [Ano91-155, LB91]. sorts [Gas93, Mel93a]. sought [BL90]. Soul [AMW90]. Sound [Ano92-225, Pou94j, RT91, Yag92a, Ano90-115, Ano92m, Ano92-41, Ano92-189, Ano93h, Cro93b, Cry94, FT90, Fri90, Pou94n, TS93a, Tho94b, YG92]. Sounded [Pou94j]. Sounds [Fri90]. source [Ano90-53, Ano91-138]. Sources [Ano90-102, Ano90-91, Ano91-101, Ano91-146, Ano91-148, Ano94-97, Api94b, Api94c, Gre90c, Tho90b, Ude91a]. Soviet [Ano90-104, Ano90-105, Pou90u]. Space [Hin92, VC90r, VN90d, Wal94a, Ano91-156, Ano92-160, Ano92-212, Hal94a, UPSW94]. span [Ano94-77, Cla93a, Ras91b]. SPARC [Ano91-159, Ano92-110, Ano93-127, Api94d, Gla91b, Way94j, Yag91c, Yag91e]. SPARC-based [Yag91c]. SPARC-compatible [Ano92-110]. SPARCard [Ano92-220]. Sparcbook [Ano94-183]. SPARCs [Ano92-193]. SPARCstation [Ano93-109, Bar90c, Lin90b, Ude90c, YS90, Yag91e, Gar94]. Sparcworks [FH93]. Spare [Ano90-97, Cal91b]. spark [Ano94-60]. sparking [Ano93-75]. Sparks [Lip90b, MG91]. Spawn [Way90a]. Speak [Ano92-52, Ano93-77]. Speaking [Ano94-151, Nan90b, Yag90f, ZV92]. Speaks [Ano91-132, Sta93, Yag92g, Yag92j, Ano91-130, Nan90b]. Spear [San90a]. spearhead [Ano91-87]. spec [Ras91d]. Special [Ano93-110, Bar90a, Cla94b, Gre90f, Kiy93b, Way91c]. specialist [Mas90b]. specialized [Ano92-196, Ano94-90, Way93e]. specifications [Alf92e, Fis90b]. Spectra [Ano90-81]. Spectrum [DE92, Nan93b, Rya90e, Way93h, Tho90e]. Spectrum/24 [Tho90e]. Speech [Ano92-52, Ano92-144, Ano94-178, Ano92-84, Tho94f]. Speed [Ano92-202, Ano92-204, Ano92-224, Ano93-47, AG91, Cla93d, Fie91c, HHP93, Ano91-68, Ano91-118, Ano91-134, Ano91-159, Ano92-49, Ano92-102, Ano92-201, Ano94-72, Ano94-87, Ano94-129, Ano94-171, Bry94d, Cla94b, Egl93c, Egl93e, Fri94b, Gre93d, Gre93c, Hal93c, Hol94, Hur93a, Joc90b, Len94b, Lov93a, Mar92a, McN94, MS93, Nan93d, Nan94b, Way91c, Way94e, Wsz93a, Ano94-72, Ano94-87, Fox93, Kre91]. speediest [Cla90a]. Speeding [Way93e, Jof93]. Speeds [Ano93-46, Bry94c, Ano90-44, Ano90-93, Ano92-173, Bri90, Ric92]. Speedy [Ano94-152, Ano90-75, Die90a, TR94b]. spell [Way94k]. spelling [Ano92-160]. spelling-checker [Ano92-160]. Spells [Ano92-177]. spend [WK90]. Spending [GP90]. spice [Zei93]. spiderwebs [Ano91-150]. Spinnaker [Ano90-79]. Spirit [Hay90b, Gre94a]. split [Ano94-97]. Spoken [LHR90, Ano93-37]. Spooler [Nan91a]. Sport [Mit90b]. sports [Ano94c]. Spot [Fie92b, Ano92-188]. Spots [Ano94-155]. spread [Ano94k]. Spreadsheet [Ano93-111, Gib92, Ano90-110, Ano94a, Del93a, Die90c, GE92, Pou94h, Rei90a, Rei91c, Van92b]. Spreadsheets [Ano91h, Ano92-55, Ano94-89, Cau93, Del92, Min90g, Ref94, Ano91-110, ADE90, Gas93, Pep91, Way94h]. Spring [Ano91q, TH93, Ano94-61, Ano92w]. Spyglass [Cra90p]. SQL [Ano92-59, Ano94-75, Ano94-153, Cla93b, Cru90, Fin90a, Ga 94d, Gre93a, HH94a, Min90g, Ude92a, Ude94k]. Square [Fin90b, Gas93, Gla91a, Cla93c, Wig92]. Squeeze [Ano91-89, Bar90b, Ano91-128, Min90c]. squeezes [Yag91e]. Squeezing [GK91]. SS1 [Ude90c]. ST506 [Alf90b]. Stac [Ano91-156]. stack [Ano91-124]. stackable [Ano94-177]. Stacker [Ano91-156]. Stacking [Nan93g]. stacks [Gas93, Nan93g, Smi90d]. staff [Ano91-189]. stage [Gib90b]. stake [Ano94-116]. Stakes [Nan93b]. Stalking [SG92b]. Stallings [Sta91b, Sta91a]. Stand [Lie93, Ano90-94, Ano93-67, Rya90b, Smi94b]. Stand-alone [Lie93, Smi94b]. Standard [Ano92-122, Ano93-74, Ano93-135, Ano94-59, Cla92d, Cry92a, Hyo94, Ric91, Sch90a, Van90, Alf91b, Alf91f, Alf92g, Ano91-97, Ano92-63, Ano92-196, Ano93g, Ano93-110, Ano94f, Ano94-111, Ano94-169, Api91c, Bry93b, Cla93b, Cou92, Gre90f, Lov90a, Mil94b, Nan90a, Nud91, Pie94, RS90, Sat90, SH91, Sta93, Tho94b, Way94c, Ano92-181, Ano94-120]. standardize [Ano94-54]. Standardizing [VN90e]. Standards [Ano91-170, Ano93g, Ano93-121, Ano94-154, Fie91d, Alf93, Ano91-68, Ano91-133, Ano91-159, Ano92-133, Ano92-209, Ano94m, Ano94-64, Ano94-171, Bry94e, HS94, Hol94, McN94, Sta90, Tur90, VC90m]. standout [Tho94i]. stands [Ano94-85]. staples [Nan93d]. Star [Fox93, Tho93b]. Stars [Ano94-89, Api94a, Fie91j]. Start [Ano93-88, Ude93j, VC90a, Ano92-211, Ano94-71, Gla90g, Mas93, Smi90a, Smi94c]. Start-up [Ano93-88, Smi94c]. started [Ano90-78, Ano92-104, Per94a]. Starter [Ano90-44]. Starting [GT93, Pou94h, Mel93b, Smi94b]. starts [Kay94b]. Stash [Ano92-215]. State [All93k, Ano90-98, Har90c, HR90, Ras90p, WB90, Ano91-186, Bar91b, DW91, DWN91, Rei91a, VN91f, SE91]. state-of-the-art [DWN91]. Station [Ano94-72, Yag90e, Ano92-220]. Statistical [Ano91-157]. Statistics [Pou93o, Ano94-104]. Status [Rei90a]. Stay [Ano92j, Lu93a, Hay94a]. staying [Sal94]. Stealing [Ano91-83]. Stealth [DeH93]. Stellar [Pou91t]. Step [Ano92-82, Ano92-205, Tho94j, Ano91k, Ano93-73, Bar92f, Ano90-99]. Stepping [Ga 92b]. Steps [Ano92-205, Ano94k, Ano94-33]. stereo [Tho94b]. stereoscopic [FT90]. Steve [VN91f]. Steven [KYCD92]. Still [Ano94-72, KPC93, Ano91z, Ano92-75, Ano93-105, Ano93-138, Ano94-96, Del93a, Fri94a, Pou93a, Sto93, VC90d]. sting [Smi94b]. stinting [Pac94]. stir [Rya91e]. stock [Ano94-71]. Stockholm [Pou93l]. Stomping [Ano90-100]. Stop [Ano92-194, Ber93, Bus92, Dys92, Ebe93, Pic93, Sey93, Bur92]. Stor [Nan94d]. Storage [Ano90-54, Ano91-142, Ano92-102, Ano92-168, Ano92-195, Ano93-112, Ano94-155, AWGY90, CW92, Har90b, Par90, Rei94d, Wal94c, And93, Ano90-34, Ano92-49, Ano92-217, Ano93-108, Ano93-132, Api94d, Hal94a, HR90, Lab92b, Mia90e, Mul94, Pet91, RP90, SG92a, Vog90, CV90b, Pet91]. Store [LM90, Ano90-94, Mul94, She90b, Way91d].

Stories [Fie90f]. storing [Ste94a]. storm [Law94, Rob90a, VN91e]. Story [Ano90-96, VN90c, Bar91f, GTN94, Pou92b, Sal94]. Storyboard [Ano91-158]. Straddles [Ano94-118]. straight [Ano90-94]. strain [VC90f, VFGZ94]. strange [Cri93, RNG94, Tho90i]. strangehold [Hal94c]. Stranger [Tho90i]. strategic [Ref94]. Strategies [All93g, Cra92b, Hel91d]. strategy [Ano91-195, Ano92-104, Ano93-31, Gre94d, Ras90g, Ude94f, War92]. stratosphere [CP94]. Streak [Nan94b]. streamline [HR91]. Street [Ano93-29, KUD92, TH93]. Streets [VC90b]. StreetTalk [Ryd93]. Strength [And90]. strengthen [Top91]. Strengthens [Hal93f]. Strengths [Mas90a, Ano94-84, Cla94b, Smi94c]. stress [Ano93-47, HHP93]. stress-tested [Ano93-47, HHP93]. stretch [Ano90y, Ano90a]. Stretching [Way93b, Way93h]. Strictly [Bon90b]. Strike [Pou93n]. Strikes [Ano91-173, Tho93g, Way94j]. string [Gri94, Nan94b]. Strings [Ras91j, Eug90]. stroke [Gre90k, Gre90k]. Stroke-Character [Gre90k]. strong [Ano94-31, Daw94c, Ga 93b, Gib92, Rei90b, Smi93e, Way94g, Yag93f]. stronger [Har93]. structure [Ano92-117, Gri94, Sin94]. Structured [NHW93]. Structures [DM90, Gre94b]. struggle [Wal94c]. struggled [Hal94i]. struggles [Cos94]. Struts [Ano90-61]. student [Ano90y, Ano90-115]. students [Cra90m]. studies [Ano91-126, SC92]. Studio [Coo90c, Lov91b, Yag92i]. StudioMaster [Ano92-46]. studios [Fri90]. Study [Ano94-136, Smi93b]. Stuff [Ano90-61, CV90a]. Style [Ano90-101, Ano91l, Ano92-38, BEW92, Egl91a, Joc91, Kea91c, Pou90r, Ras91c, Red93, Yag92i, Ano92-128, Lin91a, Pou91c, Woo90b, Yag90k, Smi90e]. stylish [Red93].

Stylus [Ano93-113]. Subclassing [Ano94-156, Kiy94]. Subjects [Ano93-107, Ano92y, Ken93b, KDG^{+}93]. Sublimation [HHP93, Ano93-131, Tho92g]. Subnote [Api93b]. Subnotebook [Ano94s, Cal93, Wau93b, Alf92g, Ano92-128, Api93b, Len94a]. Subnotebooks [Len94a, Len94b, Row93]. Subsidiary [Ano93m]. Substance [Ano90-101, SG92a, Yag92i]. Subsystems [Ano92-126, Ano92-102, RH93]. Subtyping [Ano94-156]. succeed [Ras92b]. succeeding [Ude94g]. Success [Ano91-151, Fri90, Ham91]. Succession [Rya90d]. successor [Ano90-113, Ano91-137, Nan93a]. such [Ano93z, Ano94-29, Ano94-71, Ano94a, Api94c, Api94d, Fie90d, Gre94a, Kor93, Lie93, Mul94]. Suddenly [Rei90c]. suffer [Cla93b]. suffering [Cra90l]. suggestions [Tho90d]. suitable [Ess94, Har93]. suite [Ano90o, Ano92-188, Ano94-85, Dej94]. suites [Law94]. Suits [Law94]. Sum [Ano91-34]. Summer [Ano93-90, Ano94s, Ano94-61, Ano94-125, Ude94l]. Summit [Ano90m, Ano90-112]. Sun [Ano91-159, Ano92-174, Ano93-65, Ano93-109, Api94d, Bar90c, Car91a, Fle93, FH93, Gar94, YS90, Yag91c, Yag91e]. SunConnect [Ano93-129]. Suns [Ano91-184]. SunSelect [Ano94-157]. Super [Ano90-102, Ano91-160, Ano92-55, Egl93e, Ano92-170, CV90d, Ano91-160, Kli90c]. Super-Priced [Ano91-160]. superalgorithm [Ano90-66]. SuperCalc [Ano92-172]. Supercharged [Ano92-172]. supercomputer [Mar90a, Par90, Pou94e]. Supercomputers [BBWR90, Rya90c]. Supercomputing [Fri91, Ano90-102]. superfast [Ano92-170]. superhighway [Ano94-103, Bry94c]. Superior [Wau93b, Ano92-42]. SuperMac [Ano92-206, Tho90e]. Superman [Hal94c]. SuperMatch [Ano93-122]. Supernote [Ano91-161]. superscalar [Hal94h, Pou94b, Rya94b, Smi94g]. SuperScope [Ano90-103]. SuperSparc [Gar94]. SuperSparcs [Ano94-52]. SupersPort [Ano91-197]. supplement [Ano91-110, Ano91-109]. supplies [Pou92j]. Support [Ano90a, Ano92-196, Ano93-80, Ano94f, Bry94b, GB93, KGTD92, Pie93, Ano90-95, Ano91-172, Ano94-45, Ano94-126, Ano94-103, Api94a, Api94c, Die94, Egl90a, Gre94b, Har93, Hyo94, Nan91e, Nan93c, Pou93n, Ras90j, Sch93b, Tho93f, Tho94i, Ude93h, Ude93i, Ude94c, Wal94e]. supported [Ano94h, Api94c, Chr94, Ude93c]. Supporting [Ano92-28, Ga 93d]. Supports [Ano93-139, Daw94a, Hyo94, Nan93h]. sure [Ano94-76, BGE^{+}92, Ras90g, Ras91k]. surely [Ano92k]. Surf [Rya91d]. surfaced [Ano92-108]. surprises [TR94a]. surprising [Ano92-53, Cra90n, Way91c]. surprisingly [Lov94, Mia92b, Tho94h]. survey [Cra90e, DE91a]. surveyed [Ano94-168]. Surveying [Ano92-190]. Surveys [Sma91b, Ano90-39, VN91f]. Survival [Hel91d, Cra90r, Lu93a]. Survivalist [Fie91k]. survive [Ude94f]. survives [Pou94m]. survivors [App91]. Svelte [Tho90h]. SVR4 [Nud91]. SVR4/MP [Nud91]. Swamp [Rya91b]. swapping [Ano92-126]. Sweet [Ano91-162]. Swift [Ano92-197, Die90e]. Swinging [KPUG93, Per92]. Swiss [Van92a]. Switch [Bry94d, Mia90i, Ano94-177, Fow91, Isl92]. Switched [Jof93]. switcher [Ano91-176]. switchers [Mia90i]. Switches [Isl92]. Switching [Ano94-57, Ano94-158, Fri94c, Bry94d]. swivels [Ano92-218]. SX [Bon90c, Mit91b, Nad90d, VN91c, VN91d]. SX/20 [Nad90d]. SXes [VC90e]. Symantec [Ano92u, Gre93b, Mas90a]. Symbiotics [Ano92-199]. Symbolic [Way90b]. symbols [Ano91z]. symmetric [Kog94]. Symmetry [Kea90, Nud91]. Symplex [Fri94c]. synchronization [Ano94-102]. synchronizing [NTS92b]. syndrome [Cra90l]. Syne [You90]. Synergy [Gil94, Gil94]. synthesis [Ano92-52, Ano92-84]. Synthesizer [Ano94-178]. System [Ano90-83, Ano92-39, Ano92-89, Ano92-198, Ano92-222, Ano92-225, Ano93-92, Ano94n, BL94, Cra91b, Don92, Hel90b, Mia91, Min90b, Nan93f, Sum90, Tho94j, Yag90a, Yag92g, Alf90c, Alf91e, Ano90-62, Ano90-68, Ano91-38, Ano91-55, Ano91-60, Ano91-66, Ano92c, Ano92-72, Ano92-78, Ano92-126, Ano92-223, Ano93-55, Ano93-132, Ano93-134, Ano94h, Ano94k, Ano94z, Ano94-45, Ano94-126, Ano94-148, Bac94, Car91b, Cra91c, Cra91k, Die94, Dys92, Egl94d, Fie90l, Fie91k, Ga 93b, Gre90h, Hal94f, Hal94i, Hal90, Hay94a, Hol91, Kea90, Kiy94, Lin93, Mic91, Nad90d, Nan93c, Nan93d, Nel91b, Nie94, Pou94a, Pou94g, Pou94j, Ras90p, Rei94g, Smi90a, Smi90b, Smi90d, Smi94a, Tho91a, Tho93b, Ude92d, Ude93f, Ude93m, Ude94f, Ude94j, Ung91, Var94b, VN91d, Wal94a, Wal94d, Web94, Yag90b, Yag90d, Ano93-134, Ano94-45, Ano94-91, Ano94-169]. System [Cot93b, Cra91b, Cra91k, Die94, Fie92b, Min90a, Moo90, NTS92a, NTS93a, RS90, Rei91b, Tho91e, Tho92e, Tho94j, YT90, Yag91g]. System/6000 [RS90]. Systempro [Nan93b, Yag90d]. Systempro/XL [Nan93b]. Systems [Alf91d, Alf92f, Ano90-95, Ano91k, Ano91-54, Ano91b, Ano91-138, Ano91-139, Ano91-144, Ano91-170, Ano92h, Ano92-53, Ano92-102, Ano92-125, Ano92-140, Ano92-166, Ano93-29, Ano94-27, Ano94-90, Ano94-147, Ano94-159, Ano94-187, Api93b, BL90, Bar91c, Bar92a, Coo91, EW92, Eva90, Fox93, GB93, Gre90f, GTN94, Hal90, HH94b, Per92, PS94, RS90, Sma93a, Smi90a, Smi90b, Smi91a, Ude94a, Yag90a, Yag90b, Yag90c, Alf92a, Alf92b, And90, Ano90y, Ano91-73, Ano91-74, Ano91-90, Ano91-141, Ano91-159, Ano92-34, Ano92-62, Ano92-61, Ano92-65, Ano92-70, Ano92-89, Ano92-108, Ano92-80, Ano92-81, Ano92-127, Ano92-178, Ano92-176, Ano92-158, Ano92-189, Ano92-182, Ano92-180, Ano92-186, Ano92-207, Ano93-134, Ano94-77, Ano94-92, Ano94-100, Ano94-162, AWGY90, Api90c, AD91, Api92b, Api94b, Api94d, Ben91, Bon90c, Bry94d, Cau91, CS92, Cot93a, Cra90e, Cra90h, Cro93b]. systems [Don92, Egl92a, Egl93a, ESN94, EB90b, Ess94, Fie90k, GN93, GR94, GR90, Gre93h, Hal93e, Hal94f, Hay94b, Hed93a, HWS91, HK94, HP94, Hyo94, Kay94b, Kor93, Laf91, Lie93, MH91, Mar92a, Mil93, Mul94, Nan91a, Nan93c, Nan93h, Per94a, Pou91h, Pou93a, Pou93d, Pou94d, Ras90j, Rei93e, Rya91a, Sal94, San90a, Sch92a, SB94, Sta94a, Sum90, TS93a, Tho93f, TR94a, VC90e, VC90q, VC90h, VC90g, Var94a, VN90a, VN91a, Vel90, Way93i, Way94f, Wsz93a, Yag90f, Yag91c, Yav94, Ano90-32, Ano90-57, Ano90-80, Ano91-48, Ano92-99, Die90e, Hel90a, Kli90a, Nan94b, Nan94d, Pou90a, SY90, Tho94i].

T1000 [Ano91-171]. T1000LE [Ano91-171]. T100X [Sma93a]. T1200XE [Ano90-107]. T1900C [Sma93b]. T4400SXC [Ano92-199]. T4500 [Sma93a]. T5 [Hal94h]. T560i [Ano91-97]. T9000 [Ano91-173]. table [Min90i]. tables [Ano92-108]. tablet [Rei91b, WS93]. tablets [MPW^{+}90]. tabs [Ano94a]. tackle [Way93i]. Tackles [You90, Lip90a, Pou94n, TH92]. tackling [Ano92-216]. tactic [Ano90-50]. Tadpole [Ano93-90, Api94d]. Tag [Car91a, Smi90g, Mia92b]. Tag-Along [Mia92b]. Tagging [Cro93a]. Taiwan [Ano90-104, Ano90-105, Ano93-100, Ano94-119, Joc92a]. Taiwan-based [Ano94-119]. Taiwanese [Ano94-159]. Take [Egl92b, Nad90d, VC90c, Ano91-66, Ano92-28, Ano92-106, Ano92-216, Ano94-79, CP94, Car91a, Hol93a, LAN92, Loe90, Mia90a, Ras91e, Tho93d, UMR^{+}90, VC90k]. taken [Ano94-117, Wau93c]. TAKES [Ano90-93, Ano94w, Ano94-138, Bar92f, Bry94a, Del93b, Die90c, ET93, Rei93b, Rei94a, Ric92, Rog91, Tho90b, Tho93d, Ano91-40, Ano91-56, Ano92-55, Ano93-98, Ano94-33, Die93b, GR94, HC93, Pou90f, Pou94a, Ras90a, Ras90l, Smi90e, Smi94b, Tho91c, Vos90a, Way90b]. Taking [Ano91-65, Ano91-163, Ano92-81, Ano94k, Ano94-65, Hal94e, Law94, Rob90a]. Tale [Ano94-43, Gre93h, Pou90v, Ano90x]. talent [Fri92, Ude93g]. Tales [Ano91-164, Ano92-67, KDG^{+}93, Mas90b, RUHG93]. Taligent [Ano92-125, Ude94l]. Talk [Ano91-165, Hay93, MR93, Sch91b, VC90j, Ano91-184, Way94a, NTS93d]. Talking [Ano93-114, Cra90n, Fie90i, Gre90d, Gre90g, Fie91c, MW94]. talks [Ano94-179, Mas90b, Sei91]. Tall [Ano91-48]. Tame [DE91b]. Tames [Mia90b]. Taming [Ano94-160]. Tandon [Alf92d, Ano91-166]. Tandy [Ano91u, Ano93-67, Mia90f]. Taos [Pou91h, Pou94g]. Tap [GAW93]. Tape [Ano90-57, Egl93f, Ano90-83, Ano91-134, Ano92-189, Gla90c, Mia90c]. Tapping [Ano92-200]. TARGA [Lov91d]. Target [Ano92-133, Ano92y]. Targets [Ano94-114, Han92a, DeH93, TR94b]. Task [Ano94e, Api93c, Gre94c]. Tasks [Gre90d, Gre90g, Ano94-107, Gil94, Hol93b, Mar90b, Smi94a]. Taste [Ano91-167]. Tate [Ano90r, Ano90-30]. Tax [LM93, Pou93k]. Tax-preparation [LM93]. taxes [Ham90]. Taxing [LM93]. TCNS [Ano92-201]. TCP [Ano92-85, Ano92-200, Ano93-115, Daw94b, Fie92a, KYT^{+}92, Nan93g, Rei94b, Way94b]. TCP/IP [Ano92-85, Ano92-200, Ano93-115, Daw94b, Fie92a, KYT^{+}92, Nan93g, Rei94b, Way94b]. Teacher [Pac94]. Teaching [Die93c]. team [Ras91i]. Teams [Dej94]. tech [Ano90x, Ano91-164, Wil94, Yag91b, Bar90j, Ran93]. TechNet [Ano94q]. technical [Ano94-103, Bar92a, Ham91, Mil94a]. technical-support [Ano94-103]. technically [Pou93n]. technique [Way91b]. techniques [Ano92-71, Ano92-135, Ano92-142, Ano94-135, DH90, DM90, Hal94h, Hel91a, Len93a, Par92, Way91d, Way93i]. Technological [UFC94, Ano90-99, Ano94-79, Cos94, RUHG93]. Technologies [Alf91a, Ano91-104, Ano92-134, Ano92d, Ano92-79, Ano92-115, Ano92-176, Ano92-161, Ano92-209, Ano93-37, Ano94p, Ano94-29, Ano94-85, Bar90b, Bar91b, BR93, CP94, Cau91, Gla90c, HR90, Hed93a, Lie93, Pet91, Rei94b, Sat90, WS93, Way92a, Web92, Woo91a, Yag91d, Ano91-125, Ano91-153, Die90e, Nan93g]. technologists [May94b]. Technology [Ano91-36, Ano93-105, Ano94-108, Bur92, Gre93d, Nan91g, Pra93, San90b, Tho94f, UT91, Way93h, And93, Ano91-50, Ano91-166, Ano92h, Ano92c, Ano92-61, Ano92-68, Ano92-135, Ano92-127, Ano92-150, Ano93-97, Ano94h, Ano94z, Ano94-31, Ano94-104, Ano94-135, Ano94c, Bar90g, Bar90j, Bry92, Cau93, DH90, Fis90c, Gre93a, Hal94c, HR91, Hed93b, HL93, Joc92a, Kay94c, Lio90a, Mic91, Pou92f, Ras92f, Rya90e, Rya90f, RP90, Rya91d, Sch91b, Smi94e, Sta91b, Wal94c, Wal94d, Wal94e, Way93e, Way94c, Zei93, Ano90-28, Ano90-82, Tho90e, VN91e]. tedium [SE94]. Teemar [Ano90-83]. Tektronix [Ano93-131, Ano93-129, Tho92b, Tho92g]. Telebit [Daw94a]. telecomm [Ano94-60]. Telecommute [Ano91-71]. telecommuting [Ano91-71, Daw94a]. Telephone [Ano94-165, Ano90t, Ano92y]. Telephony [Ano94-47, Ano94-109, Ga 94b, Ude94h, Ano94-54]. Telescript [Ano94-161, Ano94g, Way94a]. Television [Ano92-221]. tell [Ano93-64, FJPR93]. tells [Ano91-164, Ano92-31, Bar91f]. Telltale [McA90]. Telnet [STS92]. temper [Vor92]. Tempered [Sta90]. Ten [Bar93b, GD92a, Len93a, Sei91]. tenacious [Ano91-85]. tensions [Red92b]. tenth [Tho94e]. Terabytes [Ste91a]. term [Ano90-106, Ano91-195, HC93]. Terminals [Yag91h, Fie91i]. Terms [Yag90f, Ano93-51, Ano94-82, Nan90b]. Terrain [Cla93a]. Territory [Kea91a, Bon90c]. Test [Ano93i, Api90b, Egl92e, Gre93d, Ano91-89, Ano94-107, Ano94-87, Ano94-88, Ano94-148, Ano94-149, AWGY90, Api91c, Ess94, Gre90a, Yag91c]. Tested [Ano93-35, Ano94-75, Ano94-89, Ano94-90, Ano94-87, Ano94-88, Fox93, HM94, HK94, Hol94, HP94, KM94, Len94b, Len94a, Ano91-60, Ano93-47, Ano93-48, Egl94c, FJPR93, HH94a, HHP93]. Testing [All92c, Ano94-33, Ano94-91, Ano94-89, Egl94b, FJK^{+}93, Ga 93e, Nan93d, Ano94-119, Ano94-166, Egl92e, Joc92b, Lab92b, Smi94c]. Tests [Ano92-126, Ga 92a, Hol94, PGRA92, Ano91-108, Ano91-122, Ano92-27, Ano92-31, Ano92-88, Ano92-188, Ano94-90, Api93b, Api94d, Egl91a, ET91, Egl92a, GETY91, Gre93b, Gre93e, GW93, HM94, HK94, Hur93a, KM94, Lov94, Mit91a, Mit91b, Smi93e, Smi94f, TG91, Tho92a, Tho94b, Tho94c]. Teutonizing [Ano94-162]. Texas [Ano92-56, Ano93-70, Nad90f, Rei90c]. Text [Ano91-176, Ano93-28, Egl91e, Smi90c, Ano91n, BEW92, Die94, GD92a, Hel91b, Kea91c, MW92b, Mel93a, Mia93a, NTS93c, Tho90a, Tho94f, Ude93i, Wsz93a]. text-to-speech [Tho94f]. TextBridge [Egl94c]. textural [Har93]. TFT [Api94d]. tha [PS94]. Thanks [Ano91-168, Ano94c, Gib90a, Min90c, Way94d]. Theater [Cla92a]. Their [Ano91-124, Ano94-158, Ano90t, Ano90-112, Ano91-39, Ano91-54, Ano91-92, Ano91-128, Ano92b, Ano92-194, Ano92-198, Ano93-54, Ano93-94, Ano94-100, Ano94-102, Ano94-116, Ano94-117, Ano94-131, Bri90, Cla93b, Hay94b, Mul94, Rei94b, Rei94e, Rei94g, RT94b, Sel92, ST93, SE91, Tho94h, Van94b, VC90c, Var94a, WK90, Yav94]. them [Ano92-80, Egl94b, Kiy93a, Pet91, Pet90, SB94, WB90]. Theme [Ano94-112]. themselves [Ano92-75]. Theo [Yag90g]. theoretical [Ano92-102, Rya90f]. Theorist [Way90b]. theorists [Ano92-103]. Theory [Bus92, Ano91-126, Kay94c, RNG94]. Theos [Yag90b]. There [Ano92-33, Ano93-110, Cox90, Hov93a, Min92a, Mit91c, Mit93a, Nan91f, Smi94d, Yag91b, YM93, Ano91-185, Bon90b, Egl94a, Fie90d, Fri94a, Hal94i, Lip90a, May94b, VC90d, Way93c]. Thermal [Nel91a, Tho93b]. Thermal-transfer [Nel91a, Tho93b]. These [AD91, Api94d, DWN91, RM92, Alf91d, Ano91-33, Ano92-207, BGE^{+}92, Man90, TH93, Van93c, Wsz93a]. They’re [Ano92-48, Ano92-49, Ano94-159, Ano94-173, Rya90c]. thin [Nan94b]. Thing [Egl91d, Rei91d, Ano93-72, Fie90d, Gre94a]. Things [KGEM93, Rya90g]. Think [Ano91-169, Ga 94a, Ano92-69, Cla93e, Rob90a]. ThinkPad [Daw94c, Row93]. thinner [Tra92]. Third [Har91b, Ano91y, Tho94d, Ano94-62]. third-party [Tho94d, Ano94-62]. Thomas [Ano92-201]. Thomas-Conrad [Ano92-201]. Thompson [Tho91c, Tho93c, Tho94b, Tho94c, Tho94i]. thorniest [Ste91e]. thorough [Car92]. Thoroughly [Mia90f]. Those [Cal91b, Mel93b, Mur93, Tho90h]. Though [Ano94-68, Ano94-143, Daw94c]. Thought [Ano94-178]. Thousand [YS92]. Thousands [VFGZ94, VC90r]. Threads [Ano93-63, GD92a, Ham90]. threat [Ano94-149]. threaten [Pou94f]. Three [Ano91-73, Ano91-170, Ano92-53, Ano92-149, Ano93-116, Bon90c, Bry94d, Egl90d, Gas93, Law94, Mia93b, Pou90v, Ude90h, VC90a, VC90s, Wig92, Ano90-51, Ano90-112, Ano91-64, Ano91-165, Ano94m, Ano94-86, Ano94-107, Api93b, Bri90, Cra92a, Egl94c, Ga 94d, HH94a, Min92c, Per94a, Smi94e, Tho94c, VC90d, Wau93c]. three-dimensional [Ano91-64, Cra92a]. three-year [Ano94m]. Thriving [Ude90a]. throne [Ano90-113]. throughout [Gre93a, Loc92, Wal94d]. Throughput [Bry94b, Hur93a, Ken94, TR94b]. throw [Way93f]. thrown [Ano90-33]. Thumb [Ude93i]. Thumbelina [Ano91o]. Thumper [Egl92d]. Thy [Nud91]. TI [Ano93-90, Ano94-83, Gut94, Rei90c]. tie [Ste91e]. ties [Rei93a]. TIGA [Lin91c]. Tiger [DE91b]. Tight [Car92]. tighter [Ano93z, Van93a]. till [Ano92-148]. Timbuktu [Yag93e]. Time [All93f, Ano91-39, Ano91-71, Ano92-127, Ano92-158, Ano92-223, Cal91b, Cry94, GP90, Mia90i, Way90a, Ano90-78, Ano90-92, Ano91-33, Ano91-55, Ano91-155, Ano92-97, Ano92-135, Ano92-178, Ano92-159, Ano93-88, Ano93-116, Ano94-102, Ano94-183, Egl94g, HS94, Kor93, Laf91, Nic93, Orr94, Pou94d, Pou90r, Pou94s, Ras91h, Ras91n, Rei94a, Smi90a, Wal94a, Way94d, WK90, Yag90i, Ano92-135, Ano92-166, Ano92-159, Laf91, The94, Way94d]. Time-line-based [Cry94]. Times [Ano91-45, Ano92-149, Ano94q, Ano94-107, KYT^{+}92]. timesaving [Ano91-74, Nan91a]. Timing [Ano92-175]. Tiny [Smi90a, Pou92d]. Tips [Ano94-141, Ano94-163, Fie90g, Fie91c, Mil94a, Min91d, Ras91n, Ano91-195, Cra91d, KYT^{+}92, Len93a]. title [Pou94b]. titles [Ano92-60, Har93, KCU^{+}93, Ken93a, KGEM93]. TM [Ref94]. TM/1 [Ref94]. Toaster [Yag91f]. Today [Ano92-157, Ano92-215, Cla92d, DG91, Hay94b, Hel92a, Lan94b, Nan91j, Ano91-127, Ano94-85, Bry92, DH90, EW92, Har91a, Hed94, HM94, Nan91f, Nic93, One91, Ras90a, Rei94b, VC90r]. TOG [Ano92-60]. Together [Ano92-100, Ano94-44, Fie90h, Gib90a, Rya91b, Ano91-88, Min90i, Pou94g, Ste91e, Van93a]. Token [Gla91c, Smi94f]. Tokyo [Rei92a]. Tolerance [Ano91-46, Ano92-56, Rie91, VN91b]. tolerant [Ano93-132, Nel91b, Ano91-139]. Tom [Tho91c, Tho94b, Tho94c]. Tomorrow [DG91, HHP93, Taz91, Ano91-127, DH90, Ano92-202]. toner [Hay93]. Tongues [Pet90]. Too [Pac94, Ano93m, Ano94s, Bux90a, Gut93, VC90p, Way93f]. TookBook [Woo90c]. Tool [Ano91n, Mia93b, Tho93f, Woo90c, van91, Ano91i, Ano91-157, Ano92-172, Ano93-82, Car94, Cla94a, Fic90, FT90, Fri94c, Lin94, Par92, Ref94, Smi94f, Tho94k]. Toolkit [Ano92-184, EB90b, EB90a, Smi93a, Ano90-109, Ano91-195, Ano92-145, Smi93a]. toolkits [Ano92-44, Api94b, Hel92a, Min92c]. Tools [Ano92j, Ano92-145, Ano92-163, Ano92-189, Ano92-206, Ano94x, Ano94-40, Ano94-164, Bar92e, Ga 94c, Har93, Hed93a, HWS91, KGEM93, Mas92, May94a, Ano91-182, Ano92-190, Ano93-126, Ano94g, Ano94-64, Ano94-117, Ano94-125, Ano94-147, Ano94-176, Bar92d, Bis91, Car94, Cra90p, Dej94, Egl94g, ES94, FH93, Ga 92b, Gil94, Gre90j, Gre93a, Hol93b, Hur93b, Kiy93a, Lan94a, Mur93, Nan91b, Nan91l, PS94, Sha94b, Smi93a, Smi94a, Sta91a, Ude90a, Ude93i, Wal94e, Yag91d, Ano90-109]. Top [Ano93-117, Fid92, Ano91-165, Ano92b, Ano92-188, Ano92-190, Ano94-131, Ano94-166, Cra90e, GE92]. Topas [Yag93i]. topics [Ano92q, Ano92-90, Ano92-92, Ano93-50, Ano93-51, Ano93-53, Ano93-59, Ano93-58]. Torpedoes [Voo93]. torrent [VN91c]. tortuous [Ano90e]. Toshiba [Ano90-107, Ano91-171, Ano92-199, Ano94-152, Api91a, Sma93a, Sma93b]. Totable [Mia90f]. Touch [Egl90d, FW93, FT90, Ano91-172]. Touch-and-Feel [Ano91-172]. touchscreen [Mia92a]. Touchscreens [Ano92-208]. Tough [Yag93c, Ano93a, Gre93a]. toughest [Ano92-135]. tow [Ude90g]. town [Pou94b]. toy [Ano92-32]. Toys [Ano93-107]. Tracing [Gla90h, Ran90]. Track [Pou93e, Rya91e, Day90]. trackball [Pou91p]. trackballs [Kli90a]. Tracking [Ano92-203, Yah92, Pou93e]. TrackMan [Ano91o]. trackpad [Tho94b]. Trade [Ano92-206, Cal93, Egl93c]. Trade-offs [Cal93, Egl93c]. Trades [Woo90b]. Trading [HHP93]. traditional [Ano93-111, Cry94, HR90, LAN92, Pou94c, Ste94b]. traditionally [Ano92-102]. Traffic [Ano94-150, Ano92-217, Ano94-81]. tragedy [Ano92-34]. Trail [TS93a]. trailblazing [Ano90-48]. Trails [Tho94f]. Train [Ano94-72]. Training [Pac94, Ano91-59, Har93, Wil94]. Trakker [Ano92-221]. transaction [Ano93-48]. transaction-based [Ano93-48]. transactions [Api94c, Gre94b]. Transborder [Mil94a]. transceiver [Egl94d]. transfer [Ano93-93, Ano94-61, Bar90k, Nel91a, Ras90n, Tho93b]. transferred [Ano94-150]. transfers [Ano90-74, Ano92-185]. transformations [Mul94]. Transforming [Hal94i]. Transit [Mar92a]. Transition [Mel93b, DH90, Tho92e, Wsz93a]. translate [Ben91]. translated [Ano94s]. translating [Nie94]. Translation [Tho93f, Vas93b, Ano92-214, Ano93-51, Hov93a, Mil93, Ano93-101, Tho90a]. translator [Ano92-66]. Transmission [Ano93-95, Hol94]. transmits [Ano94-174]. transparency [VN91f]. Transparent [VC90g, VN91f, Ano91-38, Api94c, Mil94b]. transport [Rei94b]. transportable [Ano90w]. Transputer [Ano91-173, Pou90h]. transputers [Pou90j]. trapdoor [Way93c]. Traps [Mil94a, LAD91]. Travails [Pou94q]. Travel [Ras90b, Ano92-76]. Traveling [Pou94p, RM92, Ras92d, Ano90-45].

TravelMate [Nad90f]. Travels [Pou94q, Pou93m]. treats [Fie90j]. Tree [Ano91-48, Ano91-74, Var94c]. Trees [Ano91-74]. trek [Tho91c]. Trenches [Mas90b, Wal94e]. trend [Smi93b, Way94f]. Trends [Don92, Pou94f, Ude91e, Ano93s, Ano93-55, Coo91, Nad94a, Pou94i, Ste94b]. trial [DG91]. tribulations [Fie90j]. Tributaries [Ano92-203]. trick [Pou93q]. trickle [Pou92f]. trickle-down [Pou92f]. Tricks [Ano91n, Fie91l, Gre93e, LAD91, Min91d, Vos90a, Lan90b]. tricky [Min91c]. tries [Cou92, Cra91b]. Trim [Tra92]. trinity [Pou94n]. Trio [Alf91d, Alf91c, Ano91-133]. Trip [Pou94k]. Triumphs [Ano90-47]. Trivial [Sel92]. troff [Fie91e]. trouble [Mur93]. Troubled [Ude90a, Pra93]. Troubles [Ano90-47, Nan91l]. Troubleshooting [Ano94-141]. truck [Ras91h]. True [BF90, Lin93, Lov92c, Way91d, Ano90-79, Ano91-131, Cra91c, Egl91e, Gre94a, Gre94b, Tho93f, Mia90a]. true-color [Ano91-131]. TrueType [Gla90a, Lov92c]. Truevision [Lov91d]. Truly [Cra91f, Ano90w, Bry94e, Rob91a]. truth [KYT^{+}92]. Try [Man90, Dio94]. trying [CV90a, Cra91f]. TSR [Ano91-136]. TSRs [VC90c]. TT030 [Pou91j]. TTY [Hay90a]. tucked [Ano91-47]. Tune [IC94, Nan91k]. Tune-Up [Nan91k]. Tunes [Ano93-69]. Turbo [Ano92-131, Ano90-81, Ano90-108, Ano90-109, Ano91-174, Ano91-175, Ano93-118, Pou90i, Spi90b]. Turbocharged [Cra90c]. turf [Gib92]. Turn [Ano93-105, Ano94z, Ano92-46, Ano92-108, Ano92-181, Ano94-182, Lin93, Pou93b, Smi90e, Yag91g]. turnaround [Mar90b]. turner [Ano90-69]. turning [Ano94z, Law94]. Turns [Bon90c, Ano91-196, Ano93-111, Ano94-65, Cra90c, Kli90b, Lov90b, Lov91a, Yag92i]. Tut [Ano94-165, Nan94b]. TV [Ano92-221, Ano92-222, Ano93-142, Ano94-164, Ano94-166, Lu93a]. TWAIN [Ano92p]. Tweaking [Ano92-204]. twentieth [Ano92x]. twice [Ano92-111]. TWindows [Ano90-110]. Twinhead [Ano91-161]. Two [Ano91-92, Ano91-105, Ano91-176, Ano91-177, Ano92y, Ano92-115, Ano92-205, Ano92-206, Ano93-37, Ano93-61, Ano94-171, Cla93c, Cla94b, Cra90q, Fin90b, Fis90b, Ga 93e, Gre93h, Har93, HC93, Loe90, MS93, Pep91, Pou90v, Row93, Sma93a, SE93a, Tho93i, Tho94d, UT91, VC90e, VC90h, Yag90m, Yag93c, Ano92-27, Ano92-169, Ano93-133, Ano94m, Ano94u, Ano94v, Api91d, Day90, EB90b, Ga 93d, Ga 94d, Gla90g, Joc90b, KM94, Len94a, Nan93d, Sma92, Tho93a, Van93c]. two-part [EB90b, Gla90g]. two-to [Ano94m]. TyIN [Ano93-119]. Type [Ano93d, Lov92c, Ano92-215, Bac94, Bor93, Gla90a]. Typeface [BL94]. Typeface-Matching [BL94]. typefaces [Way93i]. typer [Ano92-141]. TypeReader [Bar92f]. Types [Ano94-167, Bry94d, GD93, Pou94g]. Typesetter [Ano90-43]. typesetting [Fie91e]. typical [Pou94b].

U.S.

[Ano91-70, Ano91-85, Ano91-98, Ano93-68, Ano94-95, Bar90j, Cra90k, Kay94d, Way93g]. U.S.S.R. [Ano91-30]. ubiquitous [Ano90t]. Ultimate [Ano92-207, SG92b, Ano92-79, Ano93-55, LAN92]. Ultimedia [Yag92g]. Ultra [BS91, Ano91d, BS91]. Ultraportable [Cal93]. Umax [Van94a]. unappreciated [Ude93b]. Unbounded [Rya90b, Sch90b]. uncertain [Ano94-126, Wat94]. Uncharted [Nie94]. Uncle [Pou94l]. unconventional [Bar90i]. under-the-hood [Gre93h]. undergone [Gas93]. underlying [Van93a]. underpinning [MS93]. underpowered [Ano94s]. Understanding [Egl90a, Ano92j, Mic91, Tho91a, Yag90f]. underwater [Smi93b]. Undocumented [Mat94b, Sch91a]. unfamiliarity [Ano94-57].

Unfortunately [Ude93m]. Unfulfilled [Ano91-69]. UniForum [Fie91j]. union [Ude92a, Ano90-104, Ano90-105]. Unique [Alf90c, Ano91z, Ano92-133, Ude93l, Yag90b]. unit [Ano91-134, Ano94-175, Egl94d]. Unite [Ano93-84, VC90s, CRR93]. unites [Ano92-86]. units [Bry94d, Gla90c]. unity [Hub92]. Univel [Tra92]. universe [Ano91-83, Ano92-114]. University [Hed93b]. UNIX [Ano93-64, FJPR93, Ano90f, Ano90h, Ano91-60, Ano91-73, Ano91-74, Ano91-88, Ano91-115, Ano91-154, Ano91-178, Ano91-185, Ano92p, Ano92-32, Ano92-66, Ano92-91, Ano92-93, Ano92-99, Ano92-125, Ano92-120, Ano92-122, Ano92-141, Ano92-175, Ano92-184, Ano92-185, Ano92-193, Ano92-220, Ano93-109, Ano93-120, Ano94-181, AWGY90, Daw94d, Far90, Fie90a, Fie90b, Fie90c, Fie90d, Fie90g, Fie90h, Fie90i, Fie90k, Fie90j, Fie90l, Fie91d, Fie91b, Fie91g, Fie91h, Fie91k, Fie91i, Fie91l, Fie92a, Fie92b, Fox93, GAW93, Gre90c, Gre90g, Gre93h, HK94, Hol90, HP94, Hub92, KUD92, KGEM93, McN94, Nan91d, Nan91a, NTS92a, NTS93a, Nan93g, Nie94, Nud91, Pou94a, Rya90d, Sal94, SY90, Smi90d, Smi90e, Smi93a, Smi93d, Smi94a, Smi94e, Smi94g, Tho90c, Tho92c, Tra92, UY91, Ude91d, Ung90a, Yag90c, Yag90h, Yag90j, Yag90m, Yag90k, Yag90f, Yag91h, YS92]. Unix [Yag93j]. Unix-based [Nan91d]. Unix-like [Fie90l]. Unix-Wary [Yag90j]. UnixWare [Yag93j]. Unleashes [Nan93a]. unleashing [Ano94-131]. unless [Sey93]. unlimited [Ste91b]. Unlocking [Cro93a]. Unplugged [Ano93v]. unprecedented [IC94]. unrealistic [Ude93l]. Unsolved [Pou92k]. Unsung [Pou93p]. Unthinkable [Ano91-169]. until [Ano94-72]. unveil [Ano93-65]. Unveils [Lov92b, Ano90o, Ano91-121, Ano92-199, Ano93b]. Update [Ano94-119, Len94b, Ude94l, Wau93c, Ano92-100, Api90b]. updated [Ano90-63, Ano91-192, Gar94, KPC93]. updates [Ano92-121, Tho93h]. updating [Tho92d]. Upgradability [Mit93b]. upgradable [Alf92b, Nad90c, Tho94b, VC90h]. Upgrade [Ano93s, Lov93b, VN91d, Ano91-152, Ano92-155, Ano93-140, Cra92b, Pou90r, Pou92i, Ude94j, Van93a]. Upgraded [Nan93d, Die93b]. Upgrades [Cra92b, Pou94r, Gas93, Led90]. Upgrading [Alf92b]. upon [Ano91-83, CV90d, LDH94]. upper [Van93c]. Ups [Nad90b, Nad92, YS90]. upscale [Ano92-72]. UPSes [Api90c, Gre91c, Wsz91]. Upside [Kli90a]. Upside-Down [Kli90a]. US$10 [Tho92g]. US$100 [Egl90d, Fie90j]. US$1000 [Tho93b, TH93]. US$139 [Gre93c]. US$15 [Egl90d]. US$2500 [Ano94-129]. US$3000 [Joc90b]. US$399 [Ano91u]. US$400 [Ano91-160]. US$4000 [Per94a, Tho94c]. US$4500 [Per94a, TH93]. US$500 [Ano94-106]. US$5000 [Ano92-39, Ano94-34]. US$59.95 [Ano93-81]. US$695 [Smi94e]. US$800 [Ano91-89]. US$99 [Nan94b]. US$999 [Nan94d]. USA [Ano94-110]. usability [Ano91b, Ano94-148, Ano94-149, HH94a]. usable [Ude93m, Wsz93a]. usage [Ste94b]. Use [Ano92-200, Ano93n, Kay94c, Nag93, Ano90r, Ano90q, Ano90-42, Ano90-92, Ano91-48, Ano91-108, Ano91-130, Ano92d, Ano92-69, Ano92-156, Ano92-212, Ano94-114, Car94, Cla91a, Daw94b, Die94, Egl94d, Egl94e, Fic90, Gre94b, Hed93b, Hel91a, HM94, Lan94a, Mac91, Mel93b, Mia90b, Mit91c, Nan93d, Pou90i, SB94, Tho93c, THN93, Ude93l, VC90c, Way93g, Yag90h]. used [Ano92-30, Ano93-93, Ano94-148, Kay94c, Ras90j]. Useful [Ano94-36, Ano91e, Ano91p, Ano91-107, Ano91a, Ano92-118, Hol93b, Nel91b, Ref94, Tho93e, Vor93]. Usenet [Ano92-129, Ano92-92]. User [Ano91-140, Ano92-207, Ano93-78, Kea91c, Nic90, Pou91u, Pou92l, Pou92o, Pou92m, Pou92n, Pou92k, Pou93q, Pou93r, Pou93s, Pou93t, Pou94m, Pou94s, Yag90l, Ano90s, Ano90-29, Ano91-149, Ano91-188, Ano92-31, Ano92-79, Ano94-77, Car94, Cla91a, Cra90r, Cra91g, Egl90a, ES94, Fie90d, Gru90, Hay90a, Per94b, Sch90c, SBS91, Wsz93c, Yag90e]. user-developed [Fie90d]. user-friendly [Ano92-79]. user-interface [Ano92-31, Hay90a]. Users [Ano94-168, Fie91b, Lov91b, Lu92, Tho93f, Woo90c, Ano91-59, Ano91-80, Ano92-54, Ano92-95, Ano94e, Ano94g, Ano94-76, Cla93b, Fie91d, Fie91f, Fri90, Gla91a, Hal94f, Hal94i, Ham90, Laz90, Lov91c, Mil94b, Rei93b, Sta94a, Ude91a, Yag90j]. Uses [Ung91, BS91, KM94, Lin90b, Nan92d, Sch90b, Smi94g, The94, Var94b, Way90a]. ushers [Ude93d]. Using [DM90, Gre90f, Gre90k, Hal94h, Hol93b, Min90j, Nag93, Rai90, Ude92b, Url90, Ano91-34, Ano92-216, Fie90j, Fie91i, Fri94b, Har91c, MW92b, Rya94b, VC90j, Way94d, Yag92j, van91]. USL [Var94a]. USRobotics [Mar92c]. USSR [Ban90]. usual [Mal93, Pou94n]. usually [Ano94-71]. utilities [Ano91-73, Ano91-74, Ano92r, Ano92-120, Ano92-141, Ano94-147, Nan91a, NTS92a, Pou91m, Tho94d, WET90, Ano90-64, Tho90g, YT90]. Utility [Gre90i, Tho94d, Ano91-176, Ano92-28, Ano92-129, Ano92-120, Ano92-141, Nan91h, STA92]. UTP [Nan94b]. UUCP [Fie90f]. UUNET [Ano92t]. UX [Ano90h, Tho92c, Yag91g].

V [Ano94-111, Api91c, SY90, YT90]. V-ATE [Api91c]. V.22 [Tur90]. V.32 [Ano91-89, Gib90a, Hur93a]. V.32/V.42bis [Ano91-89]. V.32bis [Daw94a]. V.32bits [Tur90]. V.32terbo [Hol94]. V.34 [Hol94]. V.42bis [Ano91-89]. V.fast [Ano93-128]. V/386 [SY90]. V32bis [Ano93-130]. Valuable [Sch91a, Ano92-52, Ras90c, SE93a]. Value [Ano94-166, Bon90d, Ano91-130, Ano92n, Ano92o]. value-added [Ano92o]. vanguard [Mia90d]. Vannevar [KUD92]. variables [SH91]. Variation [Ano94-112, Ano92-53]. variety [Ano94h, Api94b, Kay94c, Lan94a, Pet91, Ref94, RT94b]. Various [Rob91a, Ano91-130]. vary [May94a]. varying [RM92]. vast [Ras90f, Ste91a]. Vaughan [VN91f]. Vaughan-Nichols [VN91f]. VCR [Yag91b]. VCRs [Lin93]. VDI [Ano93-46]. Vectra [Ano92-72]. Vegas [Ano94p]. vehicles [Min90j]. VEISA [Ano90k]. Velox [Ano90-99]. Vendor [Hub92, Ano91-31, Ano92-72, Ano94s, Law94, Pie94, Ude94f]. vendor-neutral [Pie94]. Vendors [Ano94-159, Ano94-169, Joc92a, Ano94-29, Ano94-76, Ano94-131, Cla93b, Egl94c, LAN92, Ras91g, RT94b, VC90a, Yav94]. venerable [Ano92-151, Ude93f, Var94b]. Ventura [Ano90-111, Ano92-38]. Venture [Ano91-164, Ano92-107]. Verifying [Ano94-170]. versatility [Ano94-148, Ano94-149, HH94a]. Version [Alf91b, Hal93b, SS94, Yag93g, Ano90-28, Ano90-37, Ano90-64, Ano90-63, Ano91-137, Ano91-135, Ano92-66, Ano92-132, Ano92-111, Ano92-151, Ano92-173, Ano93-38, Ano93-43, Ano94w, Ano94-32, Ano94-33, Ano94-126, Ano94-162, Bar90c, CV90c, Daw94d, Dej94, Del93b, Gre94e, Hal93f, Kof93, Lin94, Min90e, Min90h, Min90f, Nan93h, Pou94c, Pou91n, Rei91c, Tho92c, Ude93l, Ude94e, Wau93c, Yag93k]. version-control [Ano92-66]. Versions [Nan91c, Ano91-29, Ano91-149, Ano91-192, Ano94-126, Ano94-131, Api93c, Cla93b, Nan93d, Tho90g]. versus [ATU92, Fie90l, Fis91a, Hol94, Osm90]. vertical [Egl93a]. Very [Ano90l, Mia90e, Way93e]. VESA [Ano94-171]. vessels [Smi93b]. VGA [Ano91-160, Ano92q, Ano92-27, Ano94c, Api91a, BS91, Cov90, Kli90c, Lov90a, Lov90b]. VH [Nad90e]. VH-286 [Nad90e]. via [Ano91-33, Ano91-135, Ano92-130, Mit91c, Nan93g, UFC94]. viability [HC93]. viable [Die94, Ude90d, Ude94k, War92]. Vibrant [Joc91]. Video [Ano92-46, Ano92-146, Ano92-147, Ano92-148, Ano92-211, Ano92-210, Ano93-46, Ano93-121, Ano94-106, Ano94-152, Ano94-171, BR93, Coo90c, Lin93, Pou90s, Rei93e, Tho90a, Tho91c, TS93a, Ung90b, Way94d, Yag91b, Yag91f, Yag92e, Yag92b, Yag93d, Yag93l, Ano91-86, Ano92-40, Ano93-56, Ano93-74, Ano93-122, Ano93-134, Ano94-54, Bar90b, Cry94, Lin91c, Lin94, Lov90b, Mel93a, Pou94n, Tho90b, Tho90e, Tho94f, Yag92i, YM93, Ano91-68, Lin93, Yag91f]. video-capture [Ano93-122]. video-presentation [Yag92e]. Videoconferencing [Ano94-70, Dav94a, GR94]. VideoLogic [Lov90b]. VideoQuill [Tho90a]. VideoRISC [Way94d]. videos [Ano92-147]. VideoSpigot [Ano92-206, Ano93-122]. videoteleconferencing [Ano94-60]. Vie [Ano93-121, Ano92-188]. View [Ano91-36, Ano92-179, Cla92a, Cra90p, Jef93b, Bis91, HH94a]. viewer [NTS93c, STA92, Har93]. viewers [FT90]. viewing [Ref94, Ref94]. viewpoint [Ano94-79]. Views [Ano93-137, Cov90, Sch93b, Ano94-102, Bon90a, Ano91-149, Api94b]. Village [Egl94e]. Vinci [Die90a]. Vines [Ude93k]. VIP [Ano90l]. Virtual [Cos94, FT90, Gre90h, HP93, KCU^{+}93, Ken93a, Pou90j, Tra90, Ano94-174, Egl91c, Hal94e, Hed93b, Rei94a, Smi94e, Way94h]. Virtually [Gre90h, Ano90t, Ste91b]. Virus [Ano93-103, Ano94-149, DeH93, Ano90-100, Pou92m]. Virus-Prevention [Ano94-149]. viruses [Ano92-184, Ano94-149]. Visio [Ano93-123, Ano94-68]. Vision [Ano91-98, Ano91-67, Ano92-42, Rei94b, Ano91-111]. visionary [Ano90y]. Visions [Taz91]. visit [Fie91j]. visits [VC90t]. Vista [Api91c]. Visual [Ano91-180, Cra90p, Hur93b, Pou93c, Rei92c, Ude94c, Var94c, Ano90-38, Ano94-125, Ano94-180, BEW92, BL94, KYT^{+}92, Ano91-177, Ano92-145, Ano94f, Ano94-178, Ano94-176, Hal93f, Hur93b, KYCD92, Pou93c, Sha94b, Ude92a, Ude94c, Ude94g, Yag93a, Yag93k]. visual-programming [Ano94-125]. Visualization [Ano90-38, Ano93-104, Bro93, GD93, Web93, Ano94-71, GTGW94, Smi93c, Way93d]. visualize [Ano92d, SE93b]. visually [Ano94e, Rib93, Sch91b]. vitality [Ano91-187]. VL [Ano93-137, Ano94-90, Ga 93e]. VL-Bus [Ano93-137, Ano94-90, Ga 93e]. VLIW [Way94k]. VLSI [Ano93-45, Sta94b, Way94i]. Voice [Ano93-37, Ano94-86, Cra90d, Ano92-217, Ano94-54, CRR93, Die93b, Die94, Hel91b, LHR90, Mel93a, Sta93, Tho94f, Ano94-86]. voice-dictation [Ano94-86]. voice-mail [Die93b]. Volante [Ano91-181]. Voltage [She92]. volume [Ano90-38, Ano90-100, Ano92-136, Egl94c, Smi91b, Ano92-167]. Volume-Imaging [Ano92-167]. Voxels [Ano92-212]. VPi386 [Pou91e]. VROOMM [Spi90b]. vs [Ano90-50, Ano91-132, Ano91-124, Ano94-62, Ano94-144, Ano94-157, Die93a, Hal94c, HHP93, Nan93d, Pep91, Van93c]. VTech [Ga 94d]. VUI [Hay90a]. Vulnerability [Ano91-191]. vulnerable [Ano94-149].

W [Api93c]. Wabi [Ano94-157]. Wait [Ano92-213, Ano94-159, One91, Ano92-226, Ano93g, Bar93c, Bry92, Tho90h, VC90l, Wau93c]. waiting [Han92a, Yag90j]. wakes [RPRM92]. Walk [Ano93-124]. Walking [Ano90t]. Wallet [Tho91d]. Wallet-Friendly [Tho91d]. Walls [Ano92-186, Ras91o, RUHG93]. WAN [Ano94-76, Ste91d]. wane [Red92b]. Wang [Ano94-44]. WANs [Ano91-34, Ano93v, Ano94-104]. Want [Bar90k, Ano93-75, ESN94, Joc92a, Mel93b, Mur93, Way93g, Zei93]. wanted [Ano94-114]. wants [Ude93g, Ude93l]. War [Ano90y, Ano92-74]. ward [EB90a]. Ware [VC90q, Nan94d]. wares [Red92b]. warmed [Tho90g]. warmed-over [Tho90g]. Warp [Ano91-50, Cla93d]. warranties [Ano90-84]. Warrior [Alf91e, Ga94]. Warriors [Ano92-77]. Wars [Ano94-121, Gib92, Hal94b, Pou93j, Rya91f, SBS91, SH90, Gla90a, Per94a]. Wary [Yag90j]. Was [Ano92-213, Mal93, Smi93c, Smi93e]. Watch [ASS92, Ano91-114, Ano92-222]. Watcom [Gre94e, Gre93b, Gre94e, Hel92a]. Water [Hel90a, Pou93l]. Waters [Ude90a]. Wave [Ano92-124, Ano92-123, CW92, Har91a, Lus90, Ano93-90, Ano94-61, Wsz93c]. Waveforms [Sty91]. WaveLAN [Ras91j]. Waves [Ano92j]. Wax [Egl92b]. Way [Ano91-196, Bar91d, Kor93, Sch92a, Ano91-37, Ano91-123, Ano91-175, Ano92-37, Ano92-184, Ano92-214, Ano93z, Ano93-138, Bar90k, Bon90b, Cla92a, Ga 93d, Gre94c, Hov93a, Isl92, MW94, PTUM93, Pie94, Pou91f, Ras91b, Ras91l, Rei94b, Rei94e, Rei94f, Rib93, She90b, Sma93b, Sta90, Sta94a, Way93f, Way94a, Web92, Wsz93b, Yag91g, YM93]. Wayne [Ras90l, Ras91b, Ras91h]. Ways [Ano91-177, Ano92-150, Ano92-214, Ano93-116, Ga 93e, Ano92d, Ano92-115, Ano94-64, Api91b, Bar90i, Cri93, Gre90c, Ref94, Tho91a]. weak [Daw94c]. Weapon [Hal94f]. weapons [Ano92-103]. wear [Ano94c]. WEB [Ano92-92, Ano94-172]. weds [Yag90l]. Weighs [RT94a]. weight [Rei91a, Way93f]. Weitek [Gre90b]. Welcome [Ano90-112, VC90t, Egl92c, Min90e]. well [Ano93-116, Ano94-32, Ano94-126, Fox93, Gre93b, Gre94d, Hog90, Kog94, Mit91a, Ric91, Tho94h, Sta90]. well-balanced [Ano93-116]. Well-Tempered [Sta90]. Went [HM94]. we’re [Ano90m, Gut93]. We’ve [Ano90-33, Ano91-60, AD91, Han92a, Nan91j, She90a]. WFW [Ano93-69]. Whacks [Egl92b]. Whale [Mia91]. what-if [Ano91-150, Bis91, Bis91]. Whatever [VN90d, WB90]. Wheel [Ano93i]. Wheels [Pou93f]. Where [Ano92-205, HR91, Lin91b, Rya90f, Ano94-178, Rei94d, She93b, Ude94g, VC90a]. whereabouts [Pou93e]. Whether [Smi94b, Zei93, Ga 94d, You90]. Which [Ano91-99, Fie90a, Ano91-83, Ano91-106, Ano92-104, Ano92-207, Ano93-64, Ano94-35, Ano94-127, Egl90a, FJPR93, FJK^{+}93, Gar94, Gre93h, Hol94, Pou93a, Van93c, Way94d]. which-chip [Ano91-106]. While [Cla93b, Cla93c, Gre94d, Ano94-96, Api94c, Gla90c, Har93, HC93, Pie94]. whistles [Ras90e]. white [Smi93e]. Whiteboarding [Ano94-186]. Whiteside [Nan93g]. Whither [Ano91-186, Ano91-187, Fis91b, Ude91f, Web94]. whizzes [Ano91-92]. Who [Ano90-102, Ano91-85, Ano91-188, Buc92, McG90b, Ras90o, Ano91-146, Ano93-64, Ano93-75, Egl93b, Fie90k, Mur93, Pet90, Rei93b, Tho90h]. whole [Ano94e, Ano94-162, Pou92b]. Whom [Nan91l, Ano92n]. Whose [Ano94-187, Ras91k, Rei94b]. Wide [Ano91-190, Ano94-172, Ste91a, Ano91-80, Ano93s, Ano93v, Ano94h, Ano94-147, Ano94-149, Cla93a, Gre91a, HS94, Lu93b, Mas92, Ste91d, Egl91c, Ude94d]. Wide-Area [Ano91-190, Ste91a, Ano93v, Ano94-149, Cla93a, Gre91a, HS94, Ste91d, Ude94d]. widely [Ano94-148, Chr94, May94a]. widening [Ano94-82]. Wider [Ano94-114, Ano94-159, RT94b]. widespread [Ano92-30, Ano94e, ES93, Pou93n]. Will [All92b, Ano91-31, Ano91-71, Ano93-135, Har91b, Rya90d, Smi91b, Tho93f, Ano91-172, Ano91-191, Ano92-41, Ano92-68, Ano92-76, Ano92-89, Ano92-106, Ano92-111, Ano92-109, Ano93m, Ano93s, Ano93-37, Ano93-66, Ano93v, Ano93-93, Ano94h, Ano94s, Ano94w, Ano94-32, Ano94-61, Ano94-126, Ano94-77, Ano94-86, Ano94-111, Ano94-166, Ano94-169, Bar92a, Bry94a, Cox90, Egl93b, Fis90a, GK91, Hal94c, Hal94b, Hay93, Lip90b, Mar90a, Mar92a, McA90, Ras91b, Ras91l, Ras91o, Rei93d, Rei94b, RP90, Rya91a, Rya94b, Sch93d, She90b, Taz91, Ude92b, Ude94f, Ude94j, VC90j, Wal94b, Way92a, Way93f, Way94c, Way94g, Way94h]. William [Ano92-92, Buc92, Ken93a, Sta91b, Sta91a]. willing [Sey93]. Win [Ano91-65]. Win32 [Ano94-97, One91, Ude94b, Ude94f]. Winding [Gea90]. Window [Ano93-134, Fie92b, HWS91, Lov90b, Moo90, Ras90m, SBS91, Yag91g, Yag91h, Nad94a]. Windowing [MYT91]. Windows [Lus90, Mia90g, Ude90h, UMR^{+}90, Ude93l, Sca92, Gea90, Abr94, Ano90-49, Ano90-89, Ano90-110, Ano91p, Ano91-29, Ano91-53, Ano91-65, Ano91-86, Ano91-84, Ano91-81, Ano91-107, Ano91-132, Ano91-133, Ano91-137, Ano91-174, Ano91-175, Ano91-177, Ano91-180, Ano91-184, Ano91-191, Ano92q, Ano92t, Ano92u, Ano92-37, Ano92-47, Ano92-59, Ano92-55, Ano92-54, Ano92-58, Ano92-82, Ano92-116, Ano92-108, Ano92-112, Ano92-125, Ano92-123, Ano92-143, Ano92-151, Ano92-150, Ano92-154, Ano92-169, Ano92-172, Ano92-197, Ano92-199, Ano92-204, Ano92-210, Ano92-213, Ano92-223, Ano92-218, Ano92-224, Ano92-225, Ano92-226, Ano92-214, Ano92-216, Ano93h, Ano93r, Ano93z, Ano93w, Ano93-30, Ano93-32, Ano93-41, Ano93-46, Ano93-50, Ano93-52, Ano93-72, Ano93-75, Ano93-110, Ano93-107, Ano93-117, Ano93-122, Ano93-136, Ano93-79, Ano93-131, Ano93-138, Ano93-139, Ano93-137, Ano93-134, Ano93-135, Ano93-140, Ano93-141, Ano94h]. Windows [Ano94i, Ano94l, Ano94f, Ano94w, Ano94-32, Ano94-33, Ano94-126, Ano94-79, Ano94-86, Ano94-90, Ano94-114, Ano94a, Ano94-148, Ano94-176, Ano94-184, Ano94-177, Ano94-188, Api92a, Api93c, Api94b, Api94e, Bar93b, Bar92g, Bar92d, BF90, Bis91, Bry93a, Cla93c, Cra90r, Cry94, Cut91, Dav94b, DW91, Del93b, Die90c, Die91, DH92, Dlu91, Egl91c, Egl91e, Egl92a, Egl93a, Egl93b, Egl94c, Egl94g, Egl94f, Fid92, FJPR93, Fox93, GE92, Ga 94d, Gas93, Gil94, Gla91a, Gla91d, Gre90d, Gre93b, Gre93c, Gre93f, Gre93h, Gre94e, Hal94h, Hal90, Han92a, Har93, HP93, Hel90c, Hel90e, Hel91c, Hel91b, Hel91e, Hel91f, Hel92b, Hel92a, HH94a, HK94, HP94, Jef93b, Kea91a, Ken92, KPUG93, KCU^{+}93, KAY93, Key92, Kiy93a, Kiy93b, LM93, Lab92a, Len94a, Len93a, Len93b, Lin91b, Lin93, Lov92c, Lov93a, Lov94]. Windows [LS91b, Lus90, Mar92b, Mia93b, MS93, Min91a, Min91b, Min91c, Min91d, Min92c, Min92d, Nag92, Nan92d, NTS92b, Nan92e, Nan93c, Nan93d, Nan93g, Nan94c, NHW93, Nie94, One91, Pie93, Pou92c, Pou92e, Pou92k, Pou93j, Pou93q, Pou93s, Pou94l, Ras91l, Ras92d, Rei92c, Ric92, Row93, Sch92a, Sch93b, Sha94b, She93b, She93a, STA92, STS92, Tho93f, Ude90g, Ude91c, Ude91f, Ude92a, Ude92b, Ude92d, Ude92c, Ude93a, Ude93b, Ude93l, Ude94b, Ude94d, Ude94e, Ude94j, Url90, Van93c, Van94b, Vor92, Vos90b, Wau93a, Wau93b, Way93e, Wig92, Woo90d, Woo90c, Wsz93b, Wsz93c, Yag92f, Yag92j, Yag93b, Yag93d, Yag93e, Yag93g, Yag93k, Yag93l, You93]. Windows- [Ano94h]. Windows-based [Ano94f, Lin93]. Windows/C [Cut91]. WindowsMaker [Api93c]. WinFax [Mia94b]. Wingine [Ano93y]. WingZ [Die90c, Gib92]. winner [Ano91-122, Cra91h, VN91c]. winners [Ano91-164, She90a]. Winning [Ano92-147, Ano91-76, Ano92-131, Hal93e]. Winnowing [Ano92t]. Winpro [Dlu91]. Winpro/3 [Dlu91]. wins [Pou94n]. WinSleuth [Ano90-114, UT91]. WinWriter [Per94b]. Wire [Cla94b, Egl91b, Ham92, Nan94a]. wired [Ano94-85]. Wireless [Ano92-209, Ano93-48, Ano93v, Ano93-102, HS94, MW93b, MW93a, Ano92-170, Ano93-131, Ano94-85, Ano94-175, Ano94-181, Ano94-177, Dav90, LDH94, Way93h]. wirelessly [Egl94d]. Wiring [Ano94-76, Ano94-165]. Wirth [Pou91d]. Wisdom [May94b]. Wise [Yag90i, DE91a, Smi93e]. Wish [Bar91g, Ano94-168, Kay94c]. within [Ano92-142, Ano92-191, Gre90f]. Without [Alf91c, Ano92-186, Ano92-208, Ano93-49, Ano93-62, Ano93-112, Cra90a, Egl91b, Nan94d, Pou93b, Rie91, Ano92-35, Egl94g, Gil94, Hal94a, Ken90, Mic91, Obe90, Sha94c, Smi94b, Ude91b, Woo90c]. WNDX [Api94b]. Wolf [Ano90-27]. Wolfram [Ano91-84]. wondered [Der92]. Wonderland [Lov90b]. wonders [Min90j, Rei92a]. Won’t [Hal93e, Ano92-134, Ano94w, Ano94-72, Hal94i, Sey93, VC90f]. Woods [Ano91-164, Fie91f, Min90j, Smi93b]. Word [EWT90, LHR90, Smi90e, Van93a, Woo90d, Ano90-49, Ano90-75, Ano91-133, Ano92-226, Ano94-181, BEW92, Lan90a, Mia93a, Pou90f, Pou90r, Pou93k, Sca92, Wal93, Wau93c, Ano91-192, Ano93z, Chi93, Smi90e, Woo90d]. WordPerfect [Ano90v, Ano90-115, Ano92-226, Ano93-137, Ano93-140, Ano93-141, Ano94-84, Ano94-189, Chi93, Cla93c, Nan93h, Wau93c]. Words [Ano91-133]. Wordtech [Sch92c]. Work [Alf92f, Ano91-81, Ano91-189, Ano92-70, Ano92-71, Ano93-37, Ano94x, Ano94-139, Ano94-169, GD92b, HH94b, May94a, Per94b, Pou93o, Ras91o, Rie91, Sch90c, Ude93m, UPSW94, Ano90x, Ano91-92, Ano92-69, Ano92-109, Ano92-207, Cra90m, Cri93, Egl94c, Fie91l, Gea90, Laf91, LS91b, Pie94, Pou93e, Ras91b, Rei92a, Rei94f, Rya91a, Sey93, Smi94a, Smi94c, Wal94e, Yag90e, Yag93i]. Work-Flow [HH94b, May94a, Ano94x, Smi94a, Ude93m, Wal94e]. workable [Ras92d]. Worked [Gre93g, Ano94p]. Workers [HWS91, Laz93]. workflow [Rei93d]. workgroup [Ano93-69, Ano94-182, Ude91b, Wal94e]. Workgroups [HM94, HHP93, KM94, Len93a, Len93b, Ano93-48, Ano93-129, Ano94-174, Ga 93a, HL93, Mia94b, Nan94d, Way94g, Len93a, Len93b, Pou93j, Pou93q, Pou93s, Ude92c, Ude93b, Ude94e]. Workhorse [GW92, McN94]. Working [Ano93-30, Cra90r, Pou90w, Wal94e, Ano90-58, Ano92-112, Ano94-33, Ano94-150, Ano94-162, Bar90j, Gre94d]. workings [Cah90, SSF93, Sta91b]. WorkMan [Ude93m]. Workout [Gre93f]. Workplace [Ano94-127, Ude93e, Ano94-188]. Works [Ano90-78, Ano92-138, Hov93a, Nan92b, Nan93e, Ano91-115, Ano91-126, Ano91-135, Ano91-157, Ano91-152, Ano92-42, Ano92-70, Ano94-111, Ano94-149, Cra92b, Gre94a, Hal90, Pou93a, PP93, Pou90q, Pou91l, Rei93d, Wsz93b, Ano93-137]. workshop [Pou91d]. workspaces [Ano92-28]. Workstation [Ano91-103, Car91a, Ras91l, Smi90g, SG92b, Smi94g, Yag91g, Yag91h, Ano90f, Ano90-80, Ano91-154, Ano92-122, Ano93-55, Ano93-126, Ano94-116, Api94d, Api94e, Fie91h, Gar94, Gla91b, HC93, Lin91c, Pou90a, Pou90e, Slo91, Smi94c, TR94b, Ung90a, YS90, Yag92f]. workstation-bred [Yag92f]. Workstations [Api94d, Fie90b, Fie91h, Yag91c, Ano92-99, Ano92-110, Ano93-65, Ano94-71, Bar91a, Bri90, Mil93, Mit91c, Nan92d, Yag90m]. World [Ano91-31, Ano92-80, Ano92-81, Bur92, Cot93a, EW92, Egl94e, FT90, Man90, Mic91, Smi90c, VC90i, Wat94, Yag92d, Alf92g, Ano90e, Ano90c, Ano91-64, Ano91-71, Ano91-193, Ano92-130, Ano93s, Ano93-72, Ano94-149, Ano94-190, Ber93, Bra90, Daw94d, Fie90i, Hel91a, KM94, Kay94c, LDH94, Nov94b, Pou93m, Rob90a, RNG94, Smi90g, Ude93g, Way94a, Ano90y, Ano94-172, Egl92e, Egl94e]. World-Class [Yag92d, Pou93m]. WorldPort [Mar92c]. Worlds [Car94, Ken93a, Ano90t, Ano92c, Ano92-223, Ano94-112, Ano92-114]. Worldwide [Cra91l, Fie91a]. worry [Fic90]. Worth [Ano92-213, Cal93, VN90f, VN90b, Ano92o, Ano92-226, Ano93x, Ano93-72, Mur93, VC90b, Wau93c]. worthwhile [Rya91a]. worthy [Pep91]. WorX [Dej94]. would [Daw94d, Gre94a]. wrap [Pou93r]. wrap-up [Pou93r]. wrapped [Yag90j]. wrapping [Pou92k]. wraps [MS93]. wrestle [She91]. Write [Ano94-50, Ano92-89, Ano92-200, Ano93-64, Ano94g, Ano94f, ST93, Smi94b, Yag93a, Ano90-75]. writer [Egl91a]. Writing [Ano93-72, BEW92, Bar92g, Egl91a, Hyo94, Nag92, Ols94, One91, Ano94-45, Egl94g, Smi94b]. Wrong [Ano91-185, Ano94-30]. WSP [Ano91-193]. WSP-200 [Ano91-193]. WYSIWYG [BEW92, EWT90, Egl91e, Vau90, Woo90d].

X [Ano90-32, Ano91-194, Ano92-131, Ano92-213, Ano93-134, Api94b, Fie91i, Fie92b, Moo90, Yag90h, Yag91g, Yag91h]. X.400 [VN90e]. X/Motif [Api94b]. X2 [Vor93]. Xbase [MS93, Sch92c, Ano91-187]. Xcelerated [Ano92-99]. XEN [Ano92o]. XEN-LS [Ano92o]. Xerox [Ano90-32, Ano91-42, Cla91a, Cla92a, Egl94c]. XGA [Ric91]. Xircom [Ano92-139]. XL [Ano91-118, Nan93b]. XT [Ras90b]. XT-class [Ras90b]. XVT [Api94b, Ga 92b].

Year [Ano90-96, Ano94-127, Cra90j, Ano92-108, Ano93g, Ano94m, Ano94-35, Ano94-72, Ano94-86, Nad93]. year-end [Ano94-86, Ano94-127]. Years [Ano90c, Pou90o, Sei91, Ano90-60, Ano92-106, Ano94k, Ano94-79, Ano94-85, Fie90e, RP90, She90a, Sto93, Wau93c]. Yesterday [DG91]. yield [Rya90d]. York [Ano92-84]. you’d [Ano94-76, Fow91, Rie91]. you’ll [Ano92-134, Ano92-147, Hay93, Zei93]. Yourdon [Ano92y, You90]. you’re [Ano91-81, RM92, Smi90a, VC90j]. Yourself [Ano90s, Cra91f, Nan91l, Ras91m, Ano94-89, Ras90f]. You’ve [Ano91-196, Api94a, Der92, Lin94]. Yugo [Ano90k].

Z

[Api93b, Bri90, Lin91c, Lov92c, Nan94d]. Z-486 [Lin91c]. Z-486/25E [Lin91c]. Z-Buffer [Bri90]. Z-Lite [Api93b]. Z-Stor [Nan94d]. Z80 [Ano92k]. Z80-based [Ano92k]. Zaftig [Hel90b]. Zedcor [Tho93c]. Zen [Way90b, Ano90-116]. Zenith [Ano91-108, Ano91-197, Api93b, CV90b, Mit90b, Nan94d]. Zeos [Alf92d, Nad92]. Zero [Bar90k]. Zeroslot [Bar90k]. Zinc [Api94b, Kea91c]. ZIP [Ano94f]. zippy [Ano93-126]. ZMODEM [Ano92-185]. Zoomer [Ano93i, Ano93i]. Zortech [Spi90a]. Z’s [Bar90k].

References

Abramson:1994:GW

[Abr94] Dean Abramson. Globalization of Windows. Byte Magazine, 19(11):177–??, November 1994. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Apiki:1991:EBB

[AD91] S. Apiki and S. Diehl. 486 EISA: Born to blaze: These 33-MHz 486 EISA systems are the fastest machines we’ve seen.

Byte Magazine, 16(5):216–220, 222, 224–226, 228, 230, 232–234, May 1991. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Apiki:1990:JNA

[ADE90] Steve Apiki, Stanford A. Diehl, and Howard Eglowstein. Not just for numbers anymore: The new-generation spreadsheets aid in analysis and graphics presentation. Byte Magazine, 15(2): 148–165, February 1990. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Apiki:1991:PFC

[AG91] S. Apiki and R. Grehan. Product focus: Caching cards speed data access: The BYTE Lab evaluates eight caching controller cards that help relieve hard disk drive bottlenecks. Byte Magazine, 16(1):168–172, 174, 177– 178, 180, 182, January 1991. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Alford:1990:MRH

[Alf90a] R. Alford. The mouse that roared: The history, anatomy, and physiology of the desktop mouse. Byte Magazine, 15 (12):395–398, 400–401, November 1990. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Alford:1990:EEO

[Alf90b] R. C. Alford. The evolution of ESDI: One increasingly popular alternative to the ST506 interface is a descendant of ST506, ESDI. Byte Magazine, 15(6): 297–298, 300, 302, 304, 306, June 1990. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Alford:1990:NSM

[Alf90c] R. C. Alford. NCR’s S486/MC33 has unique approach to reliability: NCR’s new 33-MHz 486 Micro Channel system is among the fastest. Byte Magazine, 15(13): 191–193, December 1990. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Alford:1991:CPT

[Alf91a] R. C. Alford. Color printing: Technologies old and new are changing the look of desktop color printing. Byte Magazine, 16 (10):149–150, 152, 155–156, 158, October 1991. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Alford:1991:EPV

[Alf91b] R. C. Alford. The evolution of PCL: Version 5 of Hewlett– Packard’s PCL beefs up the standard office printer. Byte Magazine, 16(6):325, 327–328, 330, 332, 336, 338, June 1991. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Alford:1991:EYP

[Alf91c] R. C. Alford. Extend your printer’s reach without a LAN: A trio of printer sharers that maximize your investment. Byte Magazine, 16(5):277–278, 280, 282, May 1991. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Alford:1991:TMB

[Alf91d] R. C. Alford. A trio of 486SX machines: Better than 386 systems: Price and performance don’t quite add up for these 486SX machines. Byte Magazine, 16(9):285–286, 288, 290, September 1991. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Alford:1991:FPI

[Alf91e] Roger C. Alford. The fastest portable: IBM’s P75 road warrior: Carry your 486 system with you. Byte Magazine, 16(4): 265–266, 268, April 1991. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Alford:1991:IHD

[Alf91f] Roger C. Alford. The IDE hard disk drive interface: Soon to be the AT standard? Byte Magazine, 16(3):317–322, 324, March 1991. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Alford:1992:DDF

[Alf92a] R. C. Alford. Dueling DX2s: The first 486 clock doublers: Intel’s new 486DX2 doubles the clock on eight new systems. Byte Magazine, 17(7):259–262, July 1992. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Alford:1992:UAH

[Alf92b] R. C. Alford. Upgrading at the high end: Six modular and upgradable EISA and Micro Channel systems. Byte Magazine, 17(7):246–248, 250, 252, 254, July 1992. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Alford:1992:CC

[Alf92c] Roger C. Alford. CPU choices. Byte Magazine, 17(13):??, ???? 1992. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Alford:1992:CCH

[Alf92d]  Roger C. Alford.  Cyrix

Cx486SLCs hit the desktop: The Tandon Option 386SLC/50 and Zeos 486SLC-25 are the first desktops to incorporate Cyrix’s new processor. Byte Magazine, 17(10):251–??, October 1992. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Alford:1992:DAE

[Alf92e] Roger C. Alford. Disk arrays explained: Inside RAID drive array specifications 0 through 5 and beyond. Byte Magazine, 17 (10):259–??, October 1992. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Alford:1992:HIW

[Alf92f] Roger C. Alford. How interrupts work: Interrupt-processing mechanisms in PC systems. Byte Magazine, 17(2):249–252, 254, 256, February 1992. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Alford:1992:UHP

[Alf92g] Roger C. Alford. Under the hood: The PCMCIA redefines portability: The PCMCIA 2.0 PC card standard opens a new world for subnotebook and handheld computers. Byte Magazine, 17(14):237–??, December 1992. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Alford:1993:UHC

[Alf93] Roger C. Alford. Under the hood: CD-ROM inside and out: Exploring the complexities of CD-ROM drives, discs, and associated standards. Byte Magazine, 18(3):197–??, March 1993. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Allen:1992:EOM

[All92a] Dennis Allen. Editorial: Open markets, better computers. Byte Magazine, 17(10):10–??, October 1992. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Allen:1992:EPW

[All92b] Dennis Allen. Editorial: PCs will become more personal. Byte Magazine, 17(12):12–??, November 1992. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Allen:1992:ETC

[All92c] Dennis Allen. Editorial: Testing and CPUs. Byte Magazine, 17 (14):12–??, December 1992. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Allen:1992:EGB

[All92d] Dennis Allen. Editorial: The globalization of BYTE. Byte Magazine, 17(9):16–??, September 1992. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Allen:1993:Eb

[All93a] Dennis Allen. Editorial. Byte Magazine, 18(9):10–??, August 1993. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Allen:1993:Ec

[All93b] Dennis Allen. Editorial. Byte Magazine, 18(10):10–??, September 1993. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Allen:1993:Ed

[All93c] Dennis Allen. Editorial. Byte Magazine, 18(12):14–??, November 1993. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Allen:1993:Ee

[All93d] Dennis Allen. Editorial. Byte Magazine, 18(13):14–??, December 1993. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Allen:1993:Ea

[All93e] Dennis Allen. Editorial 1993.

Byte Magazine, 18(1):12–??, January 1993. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Allen:1993:ECR

[All93f] Dennis Allen. Editorial CD-ROM: Now is the time. Byte Magazine, 18(2):12–??, February 1993. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Allen:1993:EFS

[All93g] Dennis Allen. Editorial: Fatware strategies. Byte Magazine, 18(4):12–??, April 1993. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Allen:1993:EMP

[All93h] Dennis Allen. Editorial: More personal pathways. Byte Magazine, 18(7):10–??, June 1993. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Allen:1993:EON

[All93i] Dennis Allen. Editorial: Our new lab report: Your new consultant. Byte Magazine, 18(6): 12–??, May 1993. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Allen:1993:ERM

[All93j] Dennis Allen. Editorial: The real multimedia. Byte Magazine, 18(8):10–??, July 1993. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Allen:1993:ESB

[All93k] Dennis Allen. Editorial: The state of BYTE. Byte Magazine, 18(3):12–??, March 1993. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Allen:1994:Ea

[All94a] Dennis Allen. Editorial. Byte Magazine, 19(1):10–??, January 1994. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Allen:1994:Eb

[All94b] Dennis Allen. Editorial. Byte Magazine, 19(2):10–??, February 1994. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Allen:1994:Ec

[All94c] Dennis Allen. Editorial. Byte Magazine, 19(3):10–??, March 1994. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Allen:1994:Ed

[All94d] Dennis Allen. Editorial. Byte Magazine, 19(4):10–??, April 1994. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Allen:1994:Ee

[All94e] Dennis Allen. Editorial. Byte Magazine, 19(5):10–??, May 1994. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Allen:1994:Ef

[All94f] Dennis Allen. Editorial. Byte Magazine, 19(6):9–??, June 1994. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Allen:1994:Eg

[All94g] Dennis Allen. Editorial. Byte Magazine, 19(7):14–??, July 1994. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Allen:1994:Eh

[All94h] Dennis Allen. Editorial. Byte Magazine, 19(8):10–??, August 1994. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Allen:1994:Ei

[All94i] Dennis Allen. Editorial. Byte Magazine, 19(9):14–??, September 1994. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Allen:1994:Ej

[All94j] Dennis Allen. Editorial. Byte Magazine, 19(10):10–??, October 1994. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Allen:1994:Ek

[All94k] Dennis Allen. Editorial. Byte Magazine, 19(11):10–??, November 1994. CODEN BYTEDJ.

ISSN 0360-5280 (print), 1082-7838 (electronic).

Apiki:1990:HSP

[AMW90] S. Apiki, R. Mitchell, and S. Wszola. The heart and soul of a PC compatible: The BYTE Lab examines 23 25-MHz 80386 motherboards. Byte Magazine, 15(4): 130–161, 138,140142, April 1990. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anderson:1990:SSN

[And90] M. H. Anderson. Strength (and safety) in numbers: RAID systems may boost PC performance and reliability. Byte Magazine, 15(13):337–339, December 1990. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Andrews:1993:RDS

[And93] Dave Andrews. RAID down to the desktop: This storage technology is moving from mainframes and minicomputers to the desktop. Byte Magazine, 18(8): 28–??, July 1993. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1990:MLS

[Ano90a] Anonymous. 1-Megabyte life support: Products that help you stretch the resources of a low cost computer. Byte Magazine, 15 (3):262–263, March 1990. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1990:DAF

[Ano90b] Anonymous. 1.5 decades of April fools: This is a serious business, but it has had its funny side. Byte Magazine, 15(4):350– ??, April 1990. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1990:YBB

[Ano90c] Anonymous. 15 years of bits, bytes, and other great moments: A look at key events in BYTE, the computer industry, and world history during the last 15 years. Byte Magazine, 15(9): 369–??, September 1990. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1990:MNP

[Ano90d] Anonymous. 20-MHz 386SX:

NCR produces a quick, quiet, and secure computer. Byte Magazine, 15(5):116–??, May 1990. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1990:ACT

[Ano90e] Anonymous. Advise and compute: The tortuous evolution of copyright law in the computer world. Byte Magazine, 15(4):352–??, April 1990. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1990:AUG

[Ano90f] Anonymous. Amiga 3000UX, a Unix graphics workstation from Commodore. Byte Magazine, 15(13):132–??, December 1990. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1990:ACC

[Ano90g] Anonymous. ATS convertible classic: new life for old Macs. Byte Magazine, 15(10):132–??, October 1990. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1990:UAU

[Ano90h] Anonymous. A/UX 2.0: Apple’s Unix with a friendly face. Byte Magazine, 15(8):106–??, August 1990. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1990:BML

[Ano90i] Anonymous. Backpack: MicroSolutions lets you add a drive easily. Byte Magazine, 15(9): 120–??, September 1990. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1990:BBS

[Ano90j] Anonymous. Building blocks: A sampling of products and organizations involved in applications architectures. Byte Magazine, 15(4):248–??, April 1990. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1990:BVAb

[Ano90k] Anonymous. Business VEISA 32CSX: ALR’s Yugo engine on a Ferrari chassis. Byte Magazine, 15(10):132–??, October 1990.

CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1990:BVAa

[Ano90l] Anonymous. The BYTE VIP awards: BYTE’s readers have chosen a host of hardware and software as very important products. Byte Magazine, 15(7):305– ??, July 1990. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1990:BAS

[Ano90m] Anonymous. BYTE’s 15th anniversary summit: What it is, why we’re doing it. Byte Magazine, 15(9):218–??, September 1990. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1990:BNBa

[Ano90n] Anonymous. BYTE’s new benchmarks. Byte Magazine, 15(8):10– ??, August 1990. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1990:BNBb

[Ano90o] Anonymous. BYTE’s new benchmarks: The BYTE Lab unveils a new DOS benchmark suite. Byte Magazine, 15(8):158– ??, August 1990. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1990:CSC

[Ano90p] Anonymous. Commodore sets course for multimedia. Byte Magazine, 15(5):122–??, May 1990.

CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1990:DLN

[Ano90q] Anonymous. DaynaFile: lets the NeXT computer use floppy disks. Byte Magazine, 15(7):110– ??, July 1990. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1990:DIA

[Ano90r] Anonymous. dBASE IV 1.1: Ashton-Tate’s new incarnation is easier to use. Byte Magazine, 15(5):116–??, May 1990. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1990:DIY

[Ano90s] Anonymous. Do it yourself: Your guide to end user programming products. Byte Magazine, 15 (8):254–??, August 1990. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1990:DWD

[Ano90t] Anonymous. Does the walking do the fingers: The worlds of the great mathematicians influenced their discoveries; the ubiquitous telephone virtually rules our lives. Byte Magazine, 15 (5):376–??, May 1990. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1990:DDR

[Ano90u] Anonymous. DR DOS 5.0: Digital Research eliminates many idiosyncracies of MS DOS. Byte Magazine, 15(8):106–??, August 1990. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1990:DWG

[Ano90v] Anonymous. DrawPerfect:

WordPerfect’s graphics companion. Byte Magazine, 15(3):114– ??, March 1990. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1990:EMB

[Ano90w] Anonymous. Ergo Model 1:

The Brick: truly transportable computing. Byte Magazine, 15 (6):116–??, June 1990. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1990:FTF

[Ano90x] Anonymous. A Fairy-Tale future: High-tech prophet Raymond Kurzweil’s latest work is a masterful look at the present and future of intelligent machines. Byte Magazine, 15(13): 418–??, December 1990. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1990:FCG

[Ano90y] Anonymous. The father of computer graphics: The roots of graphically oriented computer systems stretch back to World War II, MIT, and a visionary graduate student. Byte Magazine, 15(6):380–??, June 1990. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1990:FSL

[Ano90z] Anonymous. Fax96: simple and low-cost faxing from Fremont Communications. Byte Magazine, 15(1):81–??, January 1990. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1990:FBW

[Ano90-27] Anonymous. The flight of the bee wolf: A bee-hunting fly performs navigational feats that put man’s computers to shame. Byte Magazine, 15(6):384–??, June 1990. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1990:FFT

[Ano90-28] Anonymous. FrameMaker 2.0: Frame Technology’s version for the Mac. Byte Magazine, 15 (7):110–??, July 1990. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1990:HMS

[Ano90-29] Anonymous. From hand to mouth: A sampling of recent developments in user interfaces. Byte Magazine, 15(7):252– ??, July 1990. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1990:FIA

[Ano90-30] Anonymous. Full impact 2.0: Ashton-Tate adds features. Byte Magazine, 15(6):116–??, June 1990. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1990:GEM

[Ano90-31] Anonymous. Graphics engines:

A manufacturers roundup. Byte Magazine, 15(13):307–??, December 1990. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1990:GFX

[Ano90-32] Anonymous. Gray F/X: Xerox Imaging Systems offers a grayscale raster editor. Byte Magazine, 15(1):81–??, January 1990. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1990:HAW

[Ano90-33] Anonymous. Happy anniversary!: We’ve thrown you a party. Byte Magazine, 15(9): 10–??, September 1990. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1990:HIP

[Ano90-34] Anonymous. Hardcard IIXL, plus development provides easy storage. Byte Magazine, 15(13): 132–??, December 1990. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1990:HIH

[Ano90-35] Anonymous. HardFacts: information on 6000 hardware products. Byte Magazine, 15(9): 120–??, September 1990. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1990:HCE

[Ano90-36] Anonymous. HP 48SX: a calculator for engineering and science applications. Byte Magazine, 15(5):116–??, May 1990. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1990:HAG

[Ano90-37] Anonymous. HyperCard 2.0: Apple greatly improves version 1.2. Byte Magazine, 15(7):110– ??, July 1990. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1990:IBI

[Ano90-38] Anonymous. Images beget images: Visualization is a volume that challenges our notions of visual reality. Byte Magazine, 15 (9):518–??, September 1990. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1990:DSS

[Ano90-39] Anonymous. In darkest Self-Similarity: Hugh Kenner surveys the literature on fractals. Byte Magazine, 15(6):382–??, June 1990. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1990:IGC

[Ano90-40] Anonymous. Intelligent Graphics Controller 20: Hewlett– Packard’s powerful dedicated graphics processor. Byte Magazine, 15(1):81–??, January 1990. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1990:ISL

[Ano90-41] Anonymous. Interface/shplinterface:

The latest look-and-feel lawsuit coincides with the arrival of a new book on interface design. Byte Magazine, 15(10):366–??, October 1990. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1990:ISE

[Ano90-42] Anonymous. IQ scan: an easy to use scanner from Pentax. Byte Magazine, 15(8):106– ??, August 1990. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1990:TO

[Ano90-43] Anonymous. Is the typesetter obsolete? Byte Magazine, 15 (10):152–??, October 1990. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1990:LES

[Ano90-44] Anonymous. LANtastic Ethernet starter kit: Artisoft speeds up its network. Byte Magazine, 15(2):81–??, February 1990. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1990:LMI

[Ano90-45] Anonymous. LapLink Mac III:

move files from one Mac to another with this program from Traveling Software. Byte Magazine, 15(1):81–??, January 1990. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1990:LCA

[Ano90-46] Anonymous. A laptop on a chip: Almost. Byte Magazine, 15(13): 10–??, December 1990. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1990:LTT

[Ano90-47] Anonymous. Laptop troubles and triumphs. Byte Magazine, 15 (12):10–??, November 1990. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1990:LIH

[Ano90-48] Anonymous. LaserJet III: HP’s trailblazing printer. Byte Magazine, 15(4):102–??, April 1990. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1990:LWP

[Ano90-49] Anonymous. Legacy: a word processor for Windows 3.0 from NBI. Byte Magazine, 15(9): 120–??, September 1990. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1990:LVI

[Ano90-50] Anonymous. Litigation vs. innovation: Mitch Kapor argues against litigation as a business tactic. Byte Magazine, 15(9): 520–??, September 1990. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1990:LGT

[Ano90-51] Anonymous. Lotus 1-2-3/G: three dimensions for PM. Byte Magazine, 15(4):102–??, April 1990. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1990:MLE

[Ano90-52] Anonymous. Magellan 2.0: Lotus expands its intelligent DOS shell. Byte Magazine, 15(6):116– ??, June 1990. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1990:MCY

[Ano90-53] Anonymous. Making connections: Your source to networking products and information. Byte Magazine, 15(6):252– ??, June 1990. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1990:MSG

[Ano90-54] Anonymous. Masses of storage: A guide to companies that provide mass storage solutions. Byte Magazine, 15(12):338–??, November 1990. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1990:MFL

[Ano90-55] Anonymous. MediaTracks: Farallon lets you demonstrate software on the Mac. Byte Magazine, 15(10):132–??, October 1990. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1990:MCC

[Ano90-56] Anonymous. Microsoft C 6.0: a comprehensive package for professionals. Byte Magazine, 15 (3):114–??, March 1990. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1990:MFG

[Ano90-57] Anonymous. A “more filling” generation of tape backup: Tape drives from Colorado Memory Systems and Core International. Byte Magazine, 15(13):237–??, December 1990. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1990:MMM

[Ano90-58] Anonymous. Multimedia makers mentioned: Companies working in multimedia and related fields. Byte Magazine, 15(2): 236–??, February 1990. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1990:MON

[Ano90-59] Anonymous. Muse, Occam’s natural-language interface program. Byte Magazine, 15(13): 132–??, December 1990. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1990:MQN

[Ano90-60] Anonymous. Museum quality: A new Smithsonian exhibit marks 15 years of PCs and includes one of Jerry Pournelle’s early machines. Byte Magazine, 15 (8):286–??, August 1990. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1990:MSE

[Ano90-61] Anonymous. Mylex struts EISA’s stuff. Byte Magazine, 15(4):10–??, April 1990. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1990:NLP

[Ano90-62] Anonymous. NetWare 386: Less pain, great gain: Novell’s nextgeneration LAN operating system delivers radically improved performance. Byte Magazine, 15 (3):160F–??, March 1990. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1990:NUV

[Ano90-63] Anonymous. NewWave 3.0: an updated version from Hewlett– Packard. Byte Magazine, 15(12): 132–??, November 1990. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1990:NUN

[Ano90-64] Anonymous. Norton Utilities 5.0: a new version with mixed blessings. Byte Magazine, 15(9): 120–??, September 1990. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1990:ONJ

[Ano90-65] Anonymous. Objects of note:

Just a sample of the objectoriented products available. Byte Magazine, 15(10):265–??, October 1990. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1990:MMH

[Ano90-66] Anonymous. Of minds and men: Is the human mind simply a superalgorithm? Byte Magazine, 15(8):354–??, August 1990. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1990:OLP

[Ano90-67] Anonymous. OkiLaser 400: a low price compact LED printer from Okidata. Byte Magazine, 15(3):114–??, March 1990. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1990:OGI

[Ano90-68] Anonymous. OS/90: GeoWorks introduces an operating system. Byte Magazine, 15(8):106– ??, August 1990. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1990:OMB

[Ano90-69] Anonymous. Our man in Berkeley: A real page-turner, the Cuckoo’s Egg is a computer book that reads like a classic espionage novel. Byte Magazine, 15(3):360– ??, March 1990. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1990:PNP

[Ano90-70] Anonymous. PageMaker 4.0: a nearly perfect program from Aldus. Byte Magazine, 15(6):116– ??, June 1990. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1990:PFI

[Ano90-71] Anonymous. Paperless fax and an invitation for you. Byte Magazine, 15(5):10–??, May 1990. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1990:PNF

[Ano90-72] Anonymous. Paradox 3.5: new features from Borland. Byte Magazine, 15(10):132–??, October 1990. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1990:PFF

[Ano90-73] Anonymous. PC-File 5.0: a flat file database pack from Button-Ware. Byte Magazine, 15(3):114– ??, March 1990. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1990:PPF

[Ano90-74] Anonymous. PC-to-PC file transfers made easy. Byte Magazine, 15(6):64IS–18, 20, June 1990. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1990:PWL

[Ano90-75] Anonymous. PC-Write Lite: an inexpensive, speedy word processor from Quicksoft. Byte Magazine, 15(2):81–??, February 1990. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1990:PAE

[Ano90-76] Anonymous. Photoshop: Adobe eases image manipulation. Byte Magazine, 15(4):102–??, April 1990. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1990:PRF

[Ano90-77] Anonymous. Pivot: Radius’s full-page Mac monitor with portrait and landscape orientations. Byte Magazine, 15(5):116–??, May 1990. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1990:PSW

[Ano90-78] Anonymous. A plea for software that works: It’s time developers started concentrating on software integrity. Byte Magazine, 15(13): 420–??, December 1990. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1990:PSO

[Ano90-79] Anonymous. Plus: Spinnaker offers true multiplatform compatibility. Byte Magazine, 15(7):110– ??, July 1990. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1990:PMO

[Ano90-80] Anonymous. Portable mainframe: Opus Systems introduces the first portable RISC workstation. Byte Magazine, 15(1):81– ??, January 1990. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1990:PIT

[Ano90-81] Anonymous. PowerBasic 2.0: an improved Turbo Basic compiler from Spectra Publishing. Byte Magazine, 15(2):81–??, February 1990. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1990:PER

[Ano90-82] Anonymous. Private Eye: Reflection Technology’s headset mounted monitor. Byte Magazine, 15(8):106–??, August 1990. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1990:PBS

[Ano90-83] Anonymous. ProLine backup system. Teemar’s tape backup solution for a NetWare LAN. Byte Magazine, 15(13):132–??, December 1990. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1990:PPL

[Ano90-84] Anonymous. Promises, promises:

A lawyer looks at the warranties — real and implied — that come with computer equipment. Byte Magazine, 15(12):468–??, November 1990. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1990:QPQ

[Ano90-85] Anonymous. QMSWriter PM10:

QMS brings Presentation Manager to paper. Byte Magazine, 15(2):81–??, February 1990. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1990:QMO

[Ano90-86] Anonymous. Quest for a mouseball: One man’s journey in search of an elusive IBM mouseball. Byte Magazine, 15(7):360– ??, July 1990. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1990:RBM

[Ano90-87] Anonymous. R:base 3.0: many new features from Microrim. Byte Magazine, 15(4):102–??, April 1990. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1990:RCC

[Ano90-88] Anonymous. Return of the colossal code: Must we go back to the days of bloated code for every application? Byte Magazine, 15(5):380–??, May 1990. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1990:SLS

[Ano90-89] Anonymous. ScanMan 256: Logitech’s scanner for Windows 3.0. Byte Magazine, 15(12):132–??, November 1990. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1990:SML

[Ano90-90] Anonymous. ScanMan model 32: Logitech’s redesigned hand-held Mac scanner. Byte Magazine, 15(5):116–??, May 1990. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1990:SSYa

[Ano90-91] Anonymous. Semiconductor sources: Your guide to the companies with the latest and greatest. Byte Magazine, 15(1):282– ??, January 1990. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1990:SMB

[Ano90-92] Anonymous. Sharp 9624e: a modem built for heavy full-time use. Byte Magazine, 15(10):132– ??, October 1990. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1990:STR

[Ano90-93] Anonymous. SHORT TAKES RasterOps accelerator: speeds up Macintosh graphics. Byte Magazine, 15(9):120–??, September 1990. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1990:SSYc

[Ano90-94] Anonymous. Side by side: You can store more data on a floppy disk if you can get the bits to stand up straight. Byte Magazine, 15(12):304–??, November 1990. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1990:SEN

[Ano90-95] Anonymous. Smalltalk-80 enters the nineties: ObjectworksSmalltalk from ParcPlace systems features 24-bit color and cross-platform support. Byte Magazine, 15(10):138–??, October 1990. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1990:SSYb

[Ano90-96] Anonymous. The software story of the year. Byte Magazine, 15(7):10–??, July 1990. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1990:SMD

[Ano90-97] Anonymous. Spare me the details: The next generation of computer interfaces should handle details. Byte Magazine, 15 (10):368–??, October 1990. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1990:SAA

[Ano90-98] Anonymous. State of the art advanced graphics. Byte Magazine, 15(13):250–??, December 1990. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1990:STS

[Ano90-99] Anonymous. Step 486/50, a technological showpiece from Everex and Velox. Byte Magazine, 15(13):132–??, December 1990. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1990:SNP

[Ano90-100] Anonymous. Stomping the nasties: Professor Kenner examines a new volume by computer virus hunter John McAfee. Byte Magazine, 15(12):466–??, November 1990. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1990:SS

[Ano90-101] Anonymous. Style and substance. Byte Magazine, 15(3):10– ??, March 1990. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1990:SSW

[Ano90-102] Anonymous. Super sources:

Who’s who in desktop supercomputing. Byte Magazine, 15 (5):258–??, May 1990. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1990:SGI

[Ano90-103] Anonymous. SuperScope: GW instruments makes data acquisition with the Mac easier. Byte Magazine, 15(3):114–??, March 1990. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1990:TSUa

[Ano90-104] Anonymous. Taiwan, the Soviet Union, and you. Byte Magazine, 15(6):10–??, June 1990. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1990:TSUb

[Ano90-105] Anonymous. Taiwan, the Soviet Union, and you, part 2. Byte Magazine, 15(10):10–??, October 1990. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1990:BBN

[Ano90-106] Anonymous. To boldly benchmark: New meaning for the term “high-level benchmarks.”. Byte Magazine, 15(4):356–??, April 1990. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1990:TTI

[Ano90-107] Anonymous. Toshiba T1200XE: impressive notebook computing. Byte Magazine, 15(4):102–??, April 1990. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1990:TCA

[Ano90-108] Anonymous. Turbo C++: another landmark product from Borland. Byte Magazine, 15 (7):110–??, July 1990. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1990:TDT

[Ano90-109] Anonymous. Turbo Debugger and Tools 2.0: Borland’s beefedup toolkit. Byte Magazine, 15 (6):116–??, June 1990. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1990:TMM

[Ano90-110] Anonymous. TWindows: Mosaic Marketing’s spreadsheet for Windows is compatible with Lotus 1-2-3 release 2.01. Byte Magazine, 15(2):81–??, February 1990. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1990:VPM

[Ano90-111] Anonymous. Ventura Publisher, Macintosh edition 1.0: for a mixed-machine environment. Byte Magazine, 15(12):132–??, November 1990. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1990:WBS

[Ano90-112] Anonymous. Welcome to the BYTE summit: Sixty-three of the most creative and influential people in the industry discuss their perspectives on the microcomputer industry of the future. Byte Magazine, 15(9): 222–??, September 1990. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1990:WNA

[Ano90-113] Anonymous. What’s NeXT after 1-2-3: Improv is Lotus’s successor to the 1-2-3 throne. Byte Magazine, 15(10):147–??, October 1990. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1990:WDN

[Ano90-114] Anonymous. WinSleuth: Dariana’s new diagnostics package.

Byte Magazine, 15(12):132–??, November 1990. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1990:WRS

[Ano90-115] Anonymous. WordPerfect Rhymer: for the student of the sound of language. Byte Magazine, 15(12):132–??, November 1990. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1990:ZAA

[Ano90-116] Anonymous. Zen and the Art of Assembly: A remarkable series of books argues that assembly programming is easier than learning to play the piano. Byte Magazine, 15(7):358–??, July 1990. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1991:RCA

[Ano91a] Anonymous. 1991 readers’ choice awards: Presenting the hardware and software BYTE readers find most useful. Byte Magazine, 16(7):133–??, July 1991. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1991:MSE

[Ano91b] Anonymous. 33-MHz 386 systems: Eight PCs evaluated for performance, features, and usability. Byte Magazine, 16(8): 143–??, August 1991. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1991:FS

[Ano91c] Anonymous. The 486SX falls short. Byte Magazine, 16(6):36– ??, June 1991. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1991:UFA

[Ano91d] Anonymous. 8514/Ultra and FlexScan 9080i: an ATI 8514 board and a Nanao high-res/low-rad monitor. Byte Magazine, 16(4):52–??, April 1991. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1991:ANN

[Ano91e] Anonymous. AcerAnyWare 1120NX: a nifty notebook loaded with useful features. Byte Magazine, 16(9):48–??, September 1991. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1991:AIN

[Ano91f] Anonymous. Adobe Illustrator 3.0 for the next: all the Mac features, and more. Byte Magazine, 16(9):48–??, September 1991. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1991:ADP

[Ano91g] Anonymous. Advanced desktop publishing programs. Byte Magazine, 16(7):139–??, July 1991. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1991:AS

[Ano91h] Anonymous. Advanced spreadsheets. Byte Magazine, 16 (6):171–??, June 1991. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1991:AFN

[Ano91i] Anonymous. Aldus FreeHand 3.0: a new-generation Mac drawing tool. Byte Magazine, 16(5): 48–??, May 1991. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1991:AO

[Ano91j] Anonymous. The All-in-One DOS. Byte Magazine, 16(7):36– ??, July 1991. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1991:ASS

[Ano91k] Anonymous. Amateur systems: Senior editor Ken Sheldon discusses the next step in AI. Byte Magazine, 16(1):416–??, January 1991. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1991:AIL

[Ano91l] Anonymous. Apple introduces low-cost laser quality with style: A new dual low end for Apple printers. Byte Magazine, 16(5): 48–??, May 1991. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1991:ASI

[Ano91m] Anonymous. Apple sharing: The internals of the AppleTalk Filing Protocol. Byte Magazine, 16 (10):247–??, October 1991. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1991:AOT

[Ano91n] Anonymous. An artist’s old tool learns new tricks: Adobe’s Illustrator 3.0 offers improved text handling and a new graphing capability. Byte Magazine, 16(2): 178–??, February 1991. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1991:BTT

[Ano91o] Anonymous. BallPoint, Thumbelina, and TrackMan Portable: a clever new generation of portable pointing devices. Byte Magazine, 16(4):52–??, April 1991. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1991:BWG

[Ano91p] Anonymous. BeckerTools 2.0:

Windows 3.0 gets a useful file manager. Byte Magazine, 16 (4):52–??, April 1991. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1991:BS

[Ano91q] Anonymous. The best of spring.

Byte Magazine, 16(7):10–??, July 1991. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1991:BMN

[Ano91r] Anonymous. Bicycles for the mind: A new book investigates computers in the CEO’s office. Byte Magazine, 16(8):334– ??, August 1991. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1991:BBW

[Ano91s] Anonymous. The bubbly and the babbly: What do AI scripts, the human mind, and gourmet restaurants have in common? Byte Magazine, 16(13):371–??, December 1991. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1991:BA

[Ano91t] Anonymous. The BYTE awards. Byte Magazine, 16(1):147–??, January 1991. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1991:CTU

[Ano91u] Anonymous. CDR-1000: Tandy’s US$399 CD-ROM drive brings access to affordable information. Byte Magazine, 16(7):43– ??, July 1991. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1991:C

[Ano91v] Anonymous. CeBIT ’91. Byte Magazine, 16(5):10–??, May 1991. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1991:CDM

[Ano91w] Anonymous. Chips deliver multimedia: New processors are emerging that have the features needed to bring multimedia to the desktop. Byte Magazine, 16(13):163–??, December 1991. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1991:CLPa

[Ano91x] Anonymous. Classic languages, part 2: COBOL: The common business oriented language is far from dead. Byte Magazine, 16 (10):129–??, October 1991. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1991:CLPb

[Ano91y] Anonymous. Classic languages, part 3: Lisp: The third installment of our classic languages series. Byte Magazine, 16(12): 165–??, November 1991. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1991:CLPc

[Ano91z] Anonymous. Classic languages, part 4: APL: Despite its unique symbols and narrow focus, this language is still finding new applications. Byte Magazine, 16(13):141–??, December 1991. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1991:CPD

[Ano91-27] Anonymous. Color printers (directory). Byte Magazine, 16(10):

194–195, October 1991. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1991:CGM

[Ano91-28] Anonymous. Compaq goes modular. Byte Magazine, 16(10):36– ??, October 1991. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1991:CKW

[Ano91-29] Anonymous. Compatibility is king in 1-2-3 for Windows: Windows and Mac versions offer backward compatibility. Byte Magazine, 16(10):39–??, October 1991. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1991:CUR

[Ano91-30] Anonymous. Computing in the U.S.S.R. Byte Magazine, 16 (4):120–??, April 1991. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1991:CNW

[Ano91-31] Anonymous. Computing’s new world order: Will market forces change the PC vendor landscape? Byte Magazine, 16(9): 117–??, September 1991. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1991:CEE

[Ano91-32] Anonymous. Concern for the editing environment: The Iliad Group’s PIEdit creates an integrated development environment for cross-platform programmers. Byte Magazine, 16(2): 186–??, February 1991. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1991:CCP

[Ano91-33] Anonymous. Connect, check, and plan: Plan your time, communicate via modem, and diagnose your Mac with these free programs. Byte Magazine, 16(12):360–??, November 1991. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1991:CSP

[Ano91-34] Anonymous. Connectivity: The sum of its parts: Interconnection using LANs, MANs, WANs, bridges, and routers. Byte Magazine, 16(12):197–??, November 1991. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1991:DMM

[Ano91-35] Anonymous. The dangers of multitasking: A multitasking mind-set can depersonalize human interaction. Byte Magazine, 16(13):372–??, December 1991. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1991:DMV

[Ano91-36] Anonymous. DesktopTV: multimedia from A view technology. Byte Magazine, 16(3):52– ??, March 1991. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1991:DLC

[Ano91-37] Anonymous. DiskPaper 1.0: a low-cost way for you and your Mac to explore the potential of the long-promised “paperless office”. Byte Magazine, 16(7):43– ??, July 1991. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1991:DOE

[Ano91-38] Anonymous. Distributed open environments: Building a companywide, transparent, distributed computing system is about to get a lot easier. Byte Magazine, 16(12):229–??, November 1991. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1991:DCS

[Ano91-39] Anonymous. Do computers save time: Computers generate their own impediments to productivity. Byte Magazine, 16(7):101– ??, July 1991. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1991:DDR

[Ano91-40] Anonymous. DR DOS 6.0, Digital Research takes on DOS 5.0. Byte Magazine, 16(12): 68–??, November 1991. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1991:DAR

[Ano91-41] Anonymous. Dreams of artificial reality: Recent books examine the mind-machine connection. Byte Magazine, 16(9):

390–??, September 1991. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1991:DRA

[Ano91-42] Anonymous. Dynabook revisited with Alan Kay: From Xerox PARC to Apple, Alan Kay’s most enduring contribution may be a machine that has not yet been built. Byte Magazine, 16 (2):203–??, February 1991. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1991:MEM

[Ano91-43] Anonymous. E-Mail and error messages: Free E-mail programs, an E-mail interface, and an error-code identifier. Byte Magazine, 16(10):267–??, October 1991. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1991:EIM

[Ano91-44] Anonymous. The end of Intel’s monopoly? Byte Magazine, 16 (1):10–??, January 1991. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1991:FHT

[Ano91-45] Anonymous. Facing hard times?

Byte Magazine, 16(2):10–??, February 1991. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1991:FT

[Ano91-46] Anonymous. Fault tolerance.

Byte Magazine, 16(8):172–??, August 1991. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1991:FFL

[Ano91-47] Anonymous. Fax facts: The little-known digital secrets tucked inside every fax device. Byte Magazine, 16(2):301–??, February 1991. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1991:FMF

[Ano91-48] Anonymous. Fax-O-Matic and FaxConnection: fax machines that use a laser printer for output from Tall Tree Systems and Extended Systems. Byte Magazine, 16(2):127–??, February 1991. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1991:FPC

[Ano91-49] Anonymous. FileMaker pro: Claris gives a face-lift to its database. Byte Magazine, 16(2): 127–??, February 1991. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1991:FEW

[Ano91-50] Anonymous. FoxPro 2.0 engages the warp engines: A proprietary technology boosts FoxPro’s performance. Byte Magazine, 16(4):42–??, April 1991. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1991:FNI

[Ano91-51] Anonymous. FrameMaker 3.0:

the newest incarnation of a multiple-platform program for desktop publishing. Byte Magazine, 16(9):48–??, September 1991. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1991:FFS

[Ano91-52] Anonymous. Frontier 1.0: finally, scripting for the Mac. Byte Magazine, 16(9):48–??, September 1991. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1991:FPP

[Ano91-53] Anonymous. Full page pivot:

a “flippy” monitor brings new dimensions to Windows. Byte Magazine, 16(10):42–??, October 1991. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1991:FNO

[Ano91-54] Anonymous. The future of network operating systems: What lies ahead for network operating systems? major players in the field make their predictions. Byte Magazine, 16(2):268–??, February 1991. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1991:GEA

[Ano91-55] Anonymous. Great expectations:

Advice on how not to get burned the next time you buy a system for your business. Byte Magazine, 16(2):360–??, February 1991. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1991:GPF

[Ano91-56] Anonymous. GUI programming facility: software that takes away some of the pain of developing PM applications. Byte Magazine, 16(6):52–??, June 1991. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1991:HWG

[Ano91-57] Anonymous. Here we go again. Byte Magazine, 16(6):10–??, June 1991. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1991:HPS

[Ano91-58] Anonymous. Hewlett–Packard’s 425e shows off the power of the 68040: The firstborn son of the Hewlett–Packard and Apollo merger. Byte Magazine, 16(10): 239–??, October 1991. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1991:HHS

[Ano91-59] Anonymous. How to be a hero: Smart training practices produce smart users. Byte Magazine, 16 (9):392–??, September 1991. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1991:HAS

[Ano91-60] Anonymous. HP/Apollo 9000 Series 700: the fastest Unix system we’ve tested. Byte Magazine, 16(6):52–??, June 1991. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1991:HFB

[Ano91-61] Anonymous. Human filters: Beyond managing megabytes, to avoiding infolock. Byte Magazine, 16(5):388–??, May 1991. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1991:IEP

[Ano91-62] Anonymous. Image editor promises new respect for PCs: A new program produces PCbased color separations. Byte Magazine, 16(10):234–??, October 1991. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1991:ISD

[Ano91-63] Anonymous. The incredible shrinking disk: How do hard drives pack more and more capacity into smaller and smaller packages? Byte Magazine, 16 (10):255–??, October 1991. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1991:IDT

[Ano91-64] Anonymous. Infini-D: a threedimensional world for the Mac. Byte Magazine, 16(5):48–??, May 1991. CODEN BYTEDJ.

ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1991:IWT

[Ano91-65] Anonymous. Inherit the win: Taking advantage of the objectoriented attributes of PM and Windows. Byte Magazine, 16(9): 325–??, September 1991. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1991:IQA

[Ano91-66] Anonymous. Inside QuickTime: Apple’s latest system software can take the pain out of producing and playing multimedia presentations. Byte Magazine, 16(13):189–??, December 1991. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1991:IDI

[Ano91-67] Anonymous. Integrating distributed information: The vision behind MCC’s Carnot project is integrated information in distributed heterogeneous database environments. Byte Magazine, 16(12):247–??, November 1991. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1991:IIA

[Ano91-68] Anonymous. Intel/IBM’s Audio-Video kernel: A new set of hardware and software standards introduced by Intel and IBM promise to speed progress in desktop multimedia. Byte Magazine, 16(13):177–??, December 1991. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1991:IUP

[Ano91-69] Anonymous. Interoperability:

The unfulfilled promise: Introduction. Byte Magazine, 16(12): 184–??, November 1991. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1991:ALE

[Ano91-70] Anonymous. Is America losing its edge: Some of the best and brightest scientists are leaving U.S. research labs. Byte Magazine, 16(9):233–??, September 1991. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1991:ITT

[Ano91-71] Anonymous. Is it time to telecommute: Will telecommuting save the world and our jobs? Byte Magazine, 16(5):121– ??, May 1991. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1991:III

[Ano91-72] Anonymous. IslandWrite, IslandDraw, and IslandPaint offer an isle of productivity. Byte Magazine, 16(5):292–293, 295, May 1991. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1991:JCT

[Ano91-73] Anonymous. Just for checking up: Three free utilities for DOS, Mac, and Unix systems. Byte Magazine, 16(6):341– ??, June 1991. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1991:KTF

[Ano91-74] Anonymous. Keys, trees, and fonts: Free timesaving, filetree, and previewing utilities for DOS, Mac, and Unix systems. Byte Magazine, 16(8):277– ??, August 1991. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1991:LIH

[Ano91-75] Anonymous. LaserJet IIISi: HP’s new network printer. Byte Magazine, 16(4):52–??, April 1991. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1991:LMS

[Ano91-76] Anonymous. A lean, mean SCSI-2 machine: HP’s 486 file server has a winning price/performance ratio. Byte Magazine, 16(4):47– ??, April 1991. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1991:LMP

[Ano91-77] Anonymous. Let my PCs go: Do we lose more than we gain when we network personal computers? Byte Magazine, 16 (4):394–??, April 1991. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1991:LCH

[Ano91-78] Anonymous. Libraries with class:

A handful of C++ libraries. Byte Magazine, 16(2):164–??, February 1991. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1991:LCD

[Ano91-79] Anonymous. Lower-cost downloads. Byte Magazine, 16(4):10– ??, April 1991. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1991:MMWb

[Ano91-80] Anonymous. Magnavox Metalis/ 286: for a wide range of users. Byte Magazine, 16(8):42–??, August 1991. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1991:MWWb

[Ano91-81] Anonymous. Making Windows work: Here’s help if you’re experiencing pains with Windows 3.0. Byte Magazine, 16(2): 293–??, February 1991. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1991:MG

[Ano91-82] Anonymous. Managing gigabytes. Byte Magazine, 16(5):150– ??, May 1991. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1991:MRS

[Ano91-83] Anonymous. Math reconstructed: Stealing glimpses at the numbers upon which the universe is built. Byte Magazine, 16 (1):414–??, January 1991. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1991:MWWa

[Ano91-84] Anonymous. Mathematica for Windows 3.0: Wolfram Research’s great addition. Byte Magazine, 16(2):127–135, February 1991. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1991:MMWa

[Ano91-85] Anonymous. Micro, micro: Who made the micro: Is Gilbert Hyatt the father of the microprocessor, or just the most tenacious inventor in the U.S.? Byte Magazine, 16(1):304–306, 311–312, January 1991. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1991:MMD

[Ano91-86] Anonymous. Microsoft Multimedia Development Kit: bringing audio and video to Windows applications development. Byte Magazine, 16(8):42–??, August 1991. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1991:MRM

[Ano91-87] Anonymous. The Mips R4000: Mips’s new RISC processor is the spearhead of the ACE consortium. Byte Magazine, 16(13): 271–??, December 1991. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1991:MML

[Ano91-88] Anonymous. Mix ‘n’ match LAN:

The BYTE Lab puts together a LAN for Unix, Mac, and Net-Ware clients. Byte Magazine, 16(12):272–??, November 1991. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1991:MSM

[Ano91-89] Anonymous. Modems that squeeze the most out of V.32: The BYTE Lab puts 12 V.32/ V.42bis modems — all under US$800 — to the test. Byte Magazine, 16(10):214–??, October 1991. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1991:MCD

[Ano91-90] Anonymous. Modular-CPU designs: Pros and cons of modular-CPU systems. Byte Magazine, 16(12):351–??, November 1991. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1991:MPF

[Ano91-91] Anonymous. Momenta points to the future. Byte Magazine, 16 (12):48–??, November 1991. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1991:MMP

[Ano91-92] Anonymous. More mathematical people: Two new books look at the lives of math whizzes and their work. Byte Magazine, 16 (3):388–??, March 1991. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1991:MSAa

[Ano91-93] Anonymous. Moscow’s second annual computer forum. Byte Magazine, 16(8):10–??, August 1991. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1991:MLH

[Ano91-94] Anonymous. MT735 and LT-20: the highs and lows of printing on the road. Byte Magazine, 16(5): 48–??, May 1991. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1991:MSAb

[Ano91-95] Anonymous. Multimedia: Solutions anticipating a market: Introduction. Byte Magazine, 16(13):150–??, December 1991. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1991:M

[Ano91-96] Anonymous. Multiprocessing. Byte Magazine, 16(6):196–??, June 1991. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1991:NFT

[Ano91-97] Anonymous. Nanao FlexScan T560i: a new standard for graphics displays. Byte Magazine, 16 (13):58–??, December 1991. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1991:NVU

[Ano91-98] Anonymous. A national vision:

U.S. Senator Al Gore explains why we need NREN. Byte Magazine, 16(7):188–??, July 1991. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1991:NFS

[Ano91-99] Anonymous. Network fax servers come of age (slowly): Which fax server is right for you? Byte Magazine, 16(13):226–??, December 1991. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1991:NM

[Ano91-100] Anonymous. Network management. Byte Magazine, 16(3):154– ??, March 1991. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1991:NMS

[Ano91-101] Anonymous. Network management sources: Your guide to producers of network management products. Byte Magazine, 16 (3):217–??, March 1991. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1991:NNL

[Ano91-102] Anonymous. Never-never land:

Science and lunacy collide in a fascinating book about experiments over the edge. Byte Magazine, 16(2):358–??, February 1991. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1991:NDG

[Ano91-103] Anonymous. New 3-D graphics engines give PCs workstation power: 3-D graphics boards power PCs. Byte Magazine, 16(12):313–??, November 1991. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1991:NPT

[Ano91-104] Anonymous. New printer technologies: Introduction. Byte Magazine, 16(10):136–??, October 1991. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1991:NNO

[Ano91-105] Anonymous. Noisy narcissists, one genius: Two new books explore the hacker mentality. Byte Magazine, 16(7):351–??, July 1991. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1991:NSC

[Ano91-106] Anonymous. Northgate slides into CEG: Northgate’s Slim-Line eases the which-chip anguish. Byte Magazine, 16(7):43– ??, July 1991. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1991:NDW

[Ano91-107] Anonymous. Norton Desktop for Windows: make Windows more useful with an integrated shell. Byte Magazine, 16(9): 48–??, September 1991. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1991:NPM

[Ano91-108] Anonymous. Notebook power management at its zenith: The BYTE Lab tests the MastersPort, the first notebook computer to use Intel’s power-saving 386SL CPU. Byte Magazine, 16(13):249–??, December 1991. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1991:NRSb

[Ano91-109] Anonymous. NSTL review supplement: advanced desktop publishing programs. Byte Magazine, 16(7):139–142, 144, 146, 150, 152–154, 156–157, July 1991. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1991:NRSa

[Ano91-110] Anonymous. NSTL review supplement: advanced spreadsheets (buyer’s guide). Byte Magazine, 16(6):171–174, 176, 180, 182, 184, 188, 190–192, 194, June 1991. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1991:OVO

[Ano91-111] Anonymous. Object Vision: an object-oriented approach makes forms cook. Byte Magazine, 16(5):48–??, May 1991. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1991:HDR

[Ano91-112] Anonymous. Of hard disks and redesigns. Byte Magazine, 16 (3):10–??, March 1991. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1991:OF

[Ano91-113] Anonymous. The office of the future. Byte Magazine, 16(9): 204–??, September 1991. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1991:OW

[Ano91-114] Anonymous. The ones to watch. Byte Magazine, 16(11):154–??, ?? 1991. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1991:OSU

[Ano91-115] Anonymous. Open season on Unix: “Open, schmopen, just give me something that works,” says one of BYTE’s Unix aficionados. Byte Magazine, 16 (3):390–??, March 1991. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1991:PJ

[Ano91-116] Anonymous. OS/2 2.0: A pilgrim’s journey. Byte Magazine, 16(13):46–??, December 1991. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1991:GDL

[Ano91-117] Anonymous. OS/2 goes on a diet: OS/2 1.3: leaner, meaner, faster. Byte Magazine, 16(2): 143–??, February 1991. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1991:PXH

[Ano91-118] Anonymous. PacificPage XL:

high-speed PostScript printing for the LaserJet III. Byte Magazine, 16(8):42–??, August 1991. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1991:PO

[Ano91-119] Anonymous. The paperless office. Byte Magazine, 16(4):156– ??, April 1991. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1991:PII

[Ano91-120] Anonymous. A passage from India: An intriguing biography of an Indian genius. Byte Magazine, 16(6):414–??, June 1991. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1991:PNA

[Ano91-121] Anonymous. A peck of new Apple Macintoshes: Apple unveils six new Macs. Byte Magazine, 16 (12):50–??, November 1991. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1991:PLO

[Ano91-122] Anonymous. Peer LANs offer a low-cost network alternative: The BYTE Lab tests five peer LANs and picks a winner. Byte Magazine, 16(12):303–??, November 1991. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1991:PAQ

[Ano91-123] Anonymous. Persuasion 2.0: Aldus’s quick way to do a business presentation. Byte Magazine, 16(3):52–??, March 1991. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1991:PVC

[Ano91-124] Anonymous. Photoshop vs. ColorStudio: Their battle reaches new heights: How the new Photoshop and ColorStudio stack up. Byte Magazine, 16(12):327–??, November 1991. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1991:PPM

[Ano91-125] Anonymous. PixelView PC: Mirror Technologies introduces the first affordable large-screen monitor. Byte Magazine, 16(3):52– ??, March 1991. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1991:PSM

[Ano91-126] Anonymous. Poets and sleepwalkers: This month’s lineup includes works on recreational computing and studies in programming theory. Byte Magazine, 16(4):392–??, April 1991. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1991:PGP

[Ano91-127] Anonymous. Portability and the GUI: Portability for GUI-based applications — today and tomorrow. Byte Magazine, 16(12): 221–??, November 1991. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1991:PMH

[Ano91-128] Anonymous. Power management: How portables squeeze power from their batteries. Byte Magazine, 16(9):329–??, September 1991. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1991:PPH

[Ano91-129] Anonymous. Printf Plus: Here’s how to extend C’s printf function. Byte Magazine, 16(12): 335–??, November 1991. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1991:PIG

[Ano91-130] Anonymous. Prioritizing information: Grace Hopper speaks out on the value of data and various criteria you can use to help determine that value. Byte Magazine, 16(5):169–??, May 1991. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1991:PBD

[Ano91-131] Anonymous. PS/2 blues disappear with first 16.7-Million-Color MCA adapter: RasterOps releases the first true-color, 24-bit display adapter for PS/2s. Byte Magazine, 16(2):176–??, February 1991. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1991:PSV

[Ano91-132] Anonymous. The public speaks on OS/2 vs. Windows: BIX participants debate the Microsoft/ IBM feud. Byte Magazine, 16(12):141–??, November 1991. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1991:PWW

[Ano91-133] Anonymous. Putting words to Windows: A trio of nextgeneration word processors set new standards for Windows. Byte Magazine, 16(13):53–??, December 1991. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1991:QHC

[Ano91-134] Anonymous. QFA-700: a highcapacity, high-speed tape backup unit. Byte Magazine, 16(8):42– ??, August 1991. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1991:RLS

[Ano91-135] Anonymous. Reach out and link someone: LapLink’s latest version adds a new interface and works via a modem. Byte Magazine, 16(9):43–??, September 1991. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1991:RTP

[Ano91-136] Anonymous. Remove TSR DOS programs: A TSR remover, a help enhancer, and a communications program. Byte Magazine, 16(13):283–??, December 1991. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1991:RCW

[Ano91-137] Anonymous. Repairing the cracks in Windows: An early version of Windows 3.0’s longawaited successor. Byte Magazine, 16(9):38–??, September 1991. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1991:RGD

[Ano91-138] Anonymous. Resource guide:

Document imaging systems: A source listing for document image processing systems. Byte Magazine, 16(4):217–??, April 1991. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1991:RGF

[Ano91-139] Anonymous. Resource guide:

Fault-Tolerant disk systems.

Byte Magazine, 16(8):214–??, August 1991. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1991:RGG

[Ano91-140] Anonymous. Resource guide: Graphical user interfaces. Byte Magazine, 16(6):151–??, June 1991. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1991:RGI

[Ano91-141] Anonymous. Resource guide: Intelligent software: A guide to expert systems and neural-network simulators. Byte Magazine, 16 (1):300–??, January 1991. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1991:RGMa

[Ano91-142] Anonymous. Resource guide: Massive mass storage: Jukebox manufacturers serve up gigabytes of storage. Byte Magazine, 16 (5):213–??, May 1991. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1991:RGMc

[Ano91-143] Anonymous. Resource guide: Multimedia software sampler. Byte Magazine, 16(13):198–??, December 1991. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1991:RGMb

[Ano91-144] Anonymous. Resource guide: Multiprocessing systems. Byte Magazine, 16(6):255–??, June 1991. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1991:RGO

[Ano91-145] Anonymous. Resource guide: Office automation software. Byte Magazine, 16(9):245–??, September 1991. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1991:RGPa

[Ano91-146] Anonymous. Resource guide:

Portable sources: A who’s who in portable computers. Byte Magazine, 16(2):260–??, February 1991. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1991:RGPb

[Ano91-147] Anonymous. Resource guide:

Public data networks. Byte Magazine, 16(7):190–??, July 1991. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1991:RGS

[Ano91-148] Anonymous. Resource guide:

Sources for network bridges. Byte Magazine, 16(12):267–??, November 1991. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1991:RNN

[Ano91-149] Anonymous. Reviewer’s notebook: New versions of Lotus Agenda and Folio Views make much-needed improvements that address user concerns. Byte Magazine, 16(1):230–??, January 1991. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1991:SSN

[Ano91-150] Anonymous. Scientists and simulations: A new book plays whatif games with spiderwebs, computers, and history. Byte Magazine, 16(12):447–??, November 1991. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1991:SSL

[Ano91-151] Anonymous. Seven’s a success: A long-awaited Mac OS offers compatibility, new features, and a future growth path. Byte Magazine, 16(6):42–??, June 1991. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1991:SBU

[Ano91-152] Anonymous. SideKick 2.0: Borland’s upgrade works with Paradox. Byte Magazine, 16(3):52– ??, March 1991. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1991:SCP

[Ano91-153] Anonymous. The Single-Chip PC: Does Chips and Technologies’ remarkable new chip herald the future of microprocessors? Byte Magazine, 16(12): 148–??, November 1991. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1991:SPN

[Ano91-154] Anonymous. Sony’s Portable News: A Unix workstation for the field. Byte Magazine, 16 (3):46–??, March 1991. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1991:SSP

[Ano91-155] Anonymous. Sorting out schedules: Problems and algorithms of organizing people and time. Byte Magazine, 16(13):263–??, December 1991. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1991:SSE

[Ano91-156] Anonymous. Stacker: Stac Electronics lets you double your hard disk space instantly. Byte Magazine, 16(2):127–??, February 1991. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1991:SAE

[Ano91-157] Anonymous. Statistical analysis for the executive: Knowledge-Seeker provides an analytical statistical tool for executive decision making that nearly works like magic. Byte Magazine, 16 (2):183–??, February 1991. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1991:SLI

[Ano91-158] Anonymous. Storyboard live!:

IBM’s multimedia software. Byte Magazine, 16(2):127–??, February 1991. CODEN BYTEDJ.

ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1991:SND

[Ano91-159] Anonymous. Sun’s newest desktop powerhouses raise the SPARC performance ante: The ELC and IPX systems create new standards in speed and price for SPARC systems. Byte Magazine, 16(13):257–??, December 1991. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1991:SPS

[Ano91-160] Anonymous. Super-Priced Super VGA boards: A look at six Super VGA boards priced under US$400. Byte Magazine, 16 (10):225–??, October 1991. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1991:SNT

[Ano91-161] Anonymous. Supernote 386SX: a notebook from Twinhead. Byte Magazine, 16(4):52–??, April 1991. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1991:SMF

[Ano91-162] Anonymous. Sweet memory: Five new memory managers, including a new QEMM and 386Max. Byte Magazine, 16(12): 62–??, November 1991. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1991:TEC

[Ano91-163] Anonymous. Taking exception to C: Add exception handling to the C programming language. Byte Magazine, 16 (8):259–??, August 1991. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1991:TVW

[Ano91-164] Anonymous. Tales from the venture woods: Gordon Bell’s new book tells of high-tech winners and losers. Byte Magazine, 16 (10):330–??, October 1991. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1991:TID

[Ano91-165] Anonymous. A talk with Intel: An in-depth interview with three of Intel’s top chip designers. Byte Magazine, 16(4):131– ??, April 1991. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1991:TNN

[Ano91-166] Anonymous. Tandon NB 386SX notebook: features power-saving technology. Byte Magazine, 16 (3):52–??, March 1991. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1991:TDP

[Ano91-167] Anonymous. Taste, Delta Point’s composite package for the Mac. Byte Magazine, 16(1):126–??, January 1991. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1991:TM

[Ano91-168] Anonymous. Thanks for the memory. Byte Magazine, 16(9): 10–??, September 1991. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1991:TAU

[Ano91-169] Anonymous. Think about the unthinkable: Preventing a catastrophe by leaving instructions about what to do if you aren’t around. Byte Magazine, 16(12): 448–??, November 1991. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1991:TMS

[Ano91-170] Anonymous. Three 40-MHz 386 systems set new price/ performance standards: New computers from arche, ares, and club earn respect in the BYTE Lab. Byte Magazine, 16(12): 321–??, November 1991. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1991:TTS

[Ano91-171] Anonymous. Toshiba T1000LE: a slimmer T1000. Byte Magazine, 16(1):126–??, January 1991. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1991:TFI

[Ano91-172] Anonymous. Touch-and-Feel interfaces: Laptops of the near future will have to include builtin pointing devices to support GUIs. Byte Magazine, 16(2):

223–??, February 1991. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1991:TSB

[Ano91-173] Anonymous. The transputer strikes back: A look at Inmos’s amazing new T9000 transputer chip. Byte Magazine, 16(8):265– ??, August 1991. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1991:TPB

[Ano91-174] Anonymous. Turbo Pascal 6.0:

Borland almost adds Windows. Byte Magazine, 16(1):126–??, January 1991. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1991:TPW

[Ano91-175] Anonymous. Turbo Pascal for Windows: finally, an easy way to develop Windows applications. Byte Magazine, 16(6):52– ??, June 1991. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1991:TTE

[Ano91-176] Anonymous. Two text editors:

Free (or almost): two text editors, a patching utility, and an application switcher. Byte Magazine, 16(9):336–??, September 1991. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1991:TWP

[Ano91-177] Anonymous. Two ways to program BASIC for Windows: Visual Basic and realizer offer programmers an alternative to C for creating Windows programs. Byte Magazine, 16(10):221–??, October 1991. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1991:UGI

[Ano91-178] Anonymous. Unix goes Indigo. Byte Magazine, 16(9): 40–??, September 1991. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1991:RPK

[Ano91-179] Anonymous. Up from Rosie: Professor Kenner examines a new book of essays about fractals and chaos. Byte Magazine, 16 (5):386–??, May 1991. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1991:VBW

[Ano91-180] Anonymous. Visual basic: Windows programming gets real BA-SIC. Byte Magazine, 16(8):42– ??, August 1991. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1991:VAI

[Ano91-181] Anonymous. Volante AT1000: inexpensive high-end graphics from National Design. Byte Magazine, 16(1):126–??, January 1991. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1991:WPL

[Ano91-182] Anonymous. What is a programming language: Scripting tools are real programming languages. Byte Magazine, 16 (8):103–??, August 1991. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1991:WPM

[Ano91-183] Anonymous. What price modularity: Is expandability a real need or merely a marketing ploy? Byte Magazine, 16(10):103–??, October 1991. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1991:WNM

[Ano91-184] Anonymous. WHAT’s NEW:

Macs and Suns talk to each other, an image program for Windows 3.0 counts objects, and more. Byte Magazine, 16(7):52– ??, July 1991. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1991:WWUb

[Ano91-185] Anonymous. What’s wrong with Unix: Something there is that doesn’t love Unix. Byte Magazine, 16(13):113–??, December 1991. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1991:WIB

[Ano91-186] Anonymous. Whither innovation: BYTE editors debate the state of innovation in the computer industry. Byte Magazine, 16(4):111–??, April 1991. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1991:WXF

[Ano91-187] Anonymous. Whither xBase: The foremost applications development language for the microcomputer derives its vitality from aggressive competition. Byte Magazine, 16(13):131–??, December 1991. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1991:WNG

[Ano91-188] Anonymous. Who needs GUIs: The pros and cons of graphical user interfaces. Byte Magazine, 16(6):117–??, June 1991. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1991:WDS

[Ano91-189] Anonymous. Why doesn’t software work: BYTE columnists, staff, and contributors debate the issues. Byte Magazine, 16 (3):117–??, March 1991. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1991:WAN

[Ano91-190] Anonymous. Wide-area networking. Byte Magazine, 16(7):158– ??, July 1991. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1991:WVW

[Ano91-191] Anonymous. Windows of vulnerability: What will the coming of the GUI mean to the blind and the learning-disabled? Byte Magazine, 16(6):416–??, June 1991. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1991:WWUa

[Ano91-192] Anonymous. Word 5.5 and Word for OS/2: updated versions from Microsoft. Byte Magazine, 16 (1):126–??, January 1991. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1991:WAW

[Ano91-193] Anonymous. WSP-200: almost the world’s smallest printer. Byte Magazine, 16(3):52–??, March 1991. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1991:XAD

[Ano91-194] Anonymous. The X attitude:

A developer argues for software that looks to the future, regardless of the cost. Byte Magazine, 16(7):352–??, July 1991. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1991:YCR

[Ano91-195] Anonymous. You can’t run on everything: Some tips on choosing a portability toolkit or a long-term portability strategy. Byte Magazine, 16(12):255–??, November 1991. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1991:YCL

[Ano91-196] Anonymous. You’ve come a long way, PC: The IBM PC turns 10 this month. Byte Magazine, 16 (8):336–??, August 1991. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1991:ZMS

[Ano91-197] Anonymous. Zenith’s MastersPort 386SL and SupersPort 486: the first 386SL-based low-power notebook and a loaded 486 portable. Byte Magazine, 16(6):52– ??, June 1991. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1992:BAB

[Ano92a] Anonymous. The 1991 BYTE awards: BYTE editors and contributors select the best products of 1991. Byte Magazine, 17 (1):120–??, January 1992. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1992:RCA

[Ano92b] Anonymous. The 1992 readers’ choice awards: BYTE readers name their top products. Byte Magazine, 17(6):121–??, June 1992. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1992:DDI

[Ano92c] Anonymous. 3-D displays: Interactive 3-D display technology is a reality, although no system is the best of all worlds. Byte Magazine, 17(5):183–??, May 1992.

CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1992:DNG

[Ano92d] Anonymous. 3-D: The next generation of graphics overview: Roots and branches of 3-D: The use of 3-D graphics is significantly changing the ways computers let us visualize information. the high-end technologies that allow realistic images are advancing into mainstream applications. Byte Magazine, 17(5):152– ??, May 1992. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1992:MEJ

[Ano92e] Anonymous. A 66-MHz executive jet: A 66-MHz processor plus a local bus equals power. Byte Magazine, 17(6):53– ??, June 1992. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1992:AMP

[Ano92f] Anonymous. AcerPac 150: a multimedia powerhouse. Byte Magazine, 17(7):68–??, July 1992. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1992:AAN

[Ano92g] Anonymous. Acorn A4, this notebook shows the power of the ARM processor. Byte Magazine, 17(9):??, September 1992. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1992:ASG

[Ano92h] Anonymous. All systems go: Parallel-processing technology has finally hit the mainstream. Byte Magazine, 17(8):112–??, August 1992. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1992:APF

[Ano92i] Anonymous. Ami Pro 3.0, the feature battle rages. Byte Magazine, 17(9):??, September 1992. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1992:AWD

[Ano92j] Anonymous. Ample waves of data: Five tools to help you stay afloat: Five data-analysis programs promise understanding in a sea of information. Byte Magazine, 17(1):259–??, January 1992. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1992:ANN

[Ano92k] Anonymous. Amstrad Notepad NC100, surely the last Z80-based computer. Byte Magazine, 17 (10):??, October 1992. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1992:ARN

[Ano92l] Anonymous. Apple reinvents the notebook: Apple’s lightweight notebook computers are heavyduty champs. Byte Magazine, 17(3):253–??, March 1992. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1992:AQS

[Ano92m] Anonymous. Apple’s Quadra 900 sizzles and dazzles: Apple’s new 68040-based Mac reveals outstanding performance and new sound and graphics features. Byte Magazine, 17(2): 229–??, February 1992. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1992:AIV

[Ano92n] Anonymous. Applying the Internet: The value of a network lies as much in whom it connects as how. Byte Magazine, 17 (2):111–??, February 1992. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1992:AXL

[Ano92o] Anonymous. Apricot XEN-LS II, is value-added worth it? Byte Magazine, 17(14):52–??, December 1992. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1992:ABC

[Ano92p] Anonymous. Ask BYTE: Choosing a Unix; TWAIN explained: and more. Byte Magazine, 17 (9):309–??, September 1992. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1992:ABG

[Ano92q] Anonymous. Ask BYTE: Getting less from VGA, Windows batch-file botches, highresolution graphics for BASIC, and other topics. Byte Magazine, 17(4):299–??, April 1992. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1992:ABI

[Ano92r] Anonymous. Ask BYTE: Internet access; modem failures; filecompression utilities; and more. Byte Magazine, 17(10):282–??, October 1992. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1992:ABB

[Ano92s] Anonymous. Ask Byte: The BYTE Lab responds on clock inconsistencies, loopback plug pinouts, and other issues. Byte Magazine, 17(14):260–??, December 1992. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1992:ABW

[Ano92t] Anonymous. Ask Byte: Winnowing down Windows; BYTE listings on UUNET; getting from CP/M to MS-DOS; and more. Byte Magazine, 17(12):303–??, November 1992. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1992:BIW

[Ano92u] Anonymous. Back-It for Windows, Gazelle brings backup power to Windows; Premium Exec 386SX/25C, AST’s; affordable portable color 386SX Norton Desktop for DOS, Symantec’s desktop cornucopia. Byte Magazine, 17(8):46–??, August 1992. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1992:BHC

[Ano92v] Anonymous. Battle of the heavyweights: The C market leaders slug it out. Byte Magazine, 17(3):39–??, March 1992. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1992:BCS

[Ano92w] Anonymous. The best of Comdex/Spring: BYTE’s picks from an especially exciting Comdex. Byte Magazine, 17(6): 48–??, June 1992. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1992:BMT

[Ano92x] Anonymous. The birth of the microprocessor: On the twentieth anniversary of its introduction, a retrospective. Byte Magazine, 17(3):145–??, March 1992. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1992:BGT

[Ano92y] Anonymous. Bleepers of the gates: Two industry exposes target Bill Gates, Ed Yourdon on the decline of the American programmer, telephone listings on CD-ROM, and other subjects. Byte Magazine, 17(7):284– ??, July 1992. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1992:BAC

[Ano92z] Anonymous. Born-Again Compaq: Compaq fires low at the competition. Byte Magazine, 17(7):63–??, July 1992. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1992:BLPb

[Ano92-27] Anonymous. BYTE Lab product report: Monitors: Beyond VGA: The BYTE Lab tests two dozen noninterlaced color monitors. Byte Magazine, 17(5):208– ??, May 1992. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1992:BLPa

[Ano92-28] Anonymous. BYTE Lab product report: Software in a supporting role: They manage files and disks, customize workspaces, and take the drudgery out of maintenance: Praise be utility programs! Byte Magazine, 17 (2):178–??, February 1992. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1992:CLPa

[Ano92-29] Anonymous. Classic languages, part 5: SNOBOL: The fifth installment in our classic languages series. Byte Magazine, 17(1): 149–??, January 1992. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1992:CLPb

[Ano92-30] Anonymous. Classic languages, part 6: BASIC: Despite its educational roots, this language has become the most widespread and most commonly used on microcomputers. Byte Magazine, 17 (3):155–??, March 1992. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1992:CMB

[Ano92-31] Anonymous. Code on the move:

The BYTE Lab tests seven portable user-interface libraries and tells how to choose the one that best fits your needs. Byte Magazine, 17(7):206–??, July 1992. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1992:CGU

[Ano92-32] Anonymous. Coherent grows up:

Unix clone Coherent 4.0 is no longer a toy. Byte Magazine, 17(8):37–??, August 1992. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1992:CRT

[Ano92-33] Anonymous. Color and resolution: There is more to your display than just a bunch of pixels. Byte Magazine, 17(7):171– ??, July 1992. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1992:CCO

[Ano92-34] Anonymous. Computers out of control: Obscure errors in software and systems design can result in tragedy. Byte Magazine, 17(2):344–??, February 1992. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1992:CPC

[Ano92-35] Anonymous. Convertible, pen computing without compromises. Byte Magazine, 17(14):??, December 1992. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1992:CCF

[Ano92-36] Anonymous. Cyrix’s 486 in 386 clothing: The first family of 486 clones arrives in a familiar package. Byte Magazine, 17(5):37– ??, May 1992. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1992:DWN

[Ano92-37] Anonymous. Dashboard for Windows, a new way to drive Windows. Byte Magazine, 17(10):??, October 1992. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1992:DPS

[Ano92-38] Anonymous. Database publishing with style: Ventura Data-Base Publisher and PageAhead apply desktop publishing polish to database reports. Byte Magazine, 17(6):331–??, June 1992. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1992:DSC

[Ano92-39] Anonymous. Dell System 325NC: a color notebook that destroys the US$5000 barrier. Byte Magazine, 17(3):42–??, March 1992. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1992:DDD

[Ano92-40] Anonymous. Digital deceptions:

Digital video just might redefine reality. Byte Magazine, 17 (5):372–??, May 1992. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1992:DSP

[Ano92-41] Anonymous. Digital signal processing: The new digital signal processors will change how PCs handle sound and image data. Byte Magazine, 17(8):269– ??, August 1992. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1992:DHF

[Ano92-42] Anonymous. Displays: The human factor: Knowing how our vision works is the basis for developing superior displays. Byte Magazine, 17(7):195– ??, July 1992. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1992:DIP

[Ano92-43] Anonymous. DOC.IT, a printer, fax, copier, and scanner all in one. Byte Magazine, 17(12):??, November 1992. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1992:ERC

[Ano92-44] Anonymous. DOS extenders: Raising the ceiling: A look at four toolkits for building protected-mode DOS programs. Byte Magazine, 17(1):279–??, January 1992. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1992:DMI

[Ano92-45] Anonymous. Downsizing media: 3 1/2-inch MO drives arrive: New 3 1/2-inch magnetooptical drives receive mixed reviews. Byte Magazine, 17 (5):240–??, May 1992. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1992:EVA

[Ano92-46] Anonymous. Edit video at your desk: StudioMaster Pro helps turn a Mac into a professional video editor. Byte Magazine, 17(5):260–??, May 1992. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1992:EBF

[Ano92-47] Anonymous. Editorial: BYTE focuses on Windows. Byte Magazine, 17(11):6–??, November 1992. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1992:EBT

[Ano92-48] Anonymous. Electronic books: They’re coming soon to a computer near you. Byte Magazine, 17(6):263–??, June 1992. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1992:EID

[Ano92-49] Anonymous. Embedded intelligence: Demands for higher storage performance are being answered by disk designers: They’re adding intelligence to drives to boost speed and accuracy. Byte Magazine, 17(3):195– ??, March 1992. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1992:ELP

[Ano92-50] Anonymous. Enhancing laserprinter resolution: How to make a laser printer act like a phototypesetter. Byte Magazine, 17 (3):269–??, March 1992. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1992:EEI

[Ano92-51] Anonymous. Ethics of electronic information. Byte Magazine, 17 (8):10–??, August 1992. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1992:EWS

[Ano92-52] Anonymous. Even as we speak:

Speech synthesis is a practical and valuable computer-output format. Byte Magazine, 17 (4):165–??, April 1992. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1992:FFT

[Ano92-53] Anonymous. Fast fifties: Three 486/50 systems redefine PC performance: The fastest Intelbased systems show surprising variation in price and performance. Byte Magazine, 17 (5):255–??, May 1992. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1992:FPCd

[Ano92-54] Anonymous. Filemaker Pro 2.0, Claris offers Mac and Windows users a dynamite database. Byte Magazine, 17(10):??, October 1992. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1992:FIB

[Ano92-55] Anonymous. First impressions battle of the super spreadsheets: Quattro Pro for Windows takes on Excel 4.0. Byte Magazine, 17(7):40–??, July 1992. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1992:FIF

[Ano92-56] Anonymous. First impressions:the Fail-Safe PC: The Texas Microsystems FTSA PC builds in fault tolerance from the ground up. Byte Magazine, 17(5):34–??, May 1992. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1992:FRH

[Ano92-57] Anonymous. First of the Red-Hot R4000s: SGI Crimson lights the flame for the R4000 processor from Mips. Byte Magazine, 17(7):46–??, July 1992. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1992:FWM

[Ano92-58] Anonymous. FrameMaker for Windows: maximum desktop publishing power comes to PCs. Byte Magazine, 17(6):58–??, June 1992. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1992:FAD

[Ano92-59] Anonymous. A fresh approach to databases: Approach for Windows is a relational database that doubles as a front end to dBase, Paradox, and Oracle SQL. Byte Magazine, 17(8):255– ??, August 1992. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1992:APB

[Ano92-60] Anonymous. From AI to puzzles: Bright Air, brilliant Fi2re, TOG on Interface, great cities on CD-ROM, and other titles. Byte Magazine, 17(8):266– ??, August 1992. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1992:FPCc

[Ano92-61] Anonymous. The future of pen computing, part 2: Pen-software and systems developers discuss the limits of handwriting recognition, display technology, and other issues. Byte Magazine, 17(4):99–??, April 1992. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1992:FPCb

[Ano92-62] Anonymous. The future of pen computing: Pen software developers and systems designers debate the future of pen computing. Byte Magazine, 17(3):115– ??, March 1992. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1992:FPCa

[Ano92-63] Anonymous. The future of personal computing: With the PowerPC, IBM and Apple hope to set a new standard. Byte Magazine, 17(2):96–??, February 1992. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1992:GMA

[Ano92-64] Anonymous. GammaFax MLCP-4/AEB, GammaLink broadcasts fax capability. Byte Magazine, 17(9):??, September 1992. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1992:GHL

[Ano92-65] Anonymous. Gateway 2000 4DC2-66V and Hyundai 466D2, low-cost DX2 local-bus systems. Byte Magazine, 17(14):??, December 1992. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1992:GBC

[Ano92-66] Anonymous. Getting from BA-SIC to C: A BASIC-to-C translator, a Mac application launcher, and a Unix version-control program. Byte Magazine, 17(5):300– ??, May 1992. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1992:GAT

[Ano92-67] Anonymous. Glitzers anonymous: Tales of GUIs gone berserk. Byte Magazine, 17(2): 342–??, February 1992. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1992:HCD

[Ano92-68] Anonymous. HDTV is coming to desktop: HDTV will help improve computer-monitor technology and digital-image manipulation. Byte Magazine, 17(7):189– ??, July 1992. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1992:HPC

[Ano92-69] Anonymous. Hidden persuaders:

The computer you use determines how you work — and how you think. Byte Magazine, 17 (4):368–??, April 1992. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1992:HIW

[Ano92-70] Anonymous. How interrupts work: How interrupt processing works in PC systems. Byte Magazine, 17(2):247–??, February 1992. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1992:HSW

[Ano92-71] Anonymous. How scanners work: New techniques make color scanning better and more affordable. Byte Magazine, 17(6):347– ??, June 1992. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1992:HVL

[Ano92-72] Anonymous. HP Vectra 486/ 33N, a low-cost system from an upscale vendor. Byte Magazine, 17(14):??, December 1992. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1992:ILI

[Ano92-73] Anonymous. IBM LaserPrinter 10A: an IBM printer for your Mac. Byte Magazine, 17(7):68– ??, July 1992. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1992:IBS

[Ano92-74] Anonymous. The importance of being singular: War and the intelligent machine, and other selections. Byte Magazine, 17 (6):344–??, June 1992. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1992:IAY

[Ano92-75] Anonymous. Infoglut at your fingertips: All the information search-and-retrieval services still remain islands to themselves. Byte Magazine, 17(3): 364–??, March 1992. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1992:IPC

[Ano92-76] Anonymous. Infolio, have pen computer, will travel. Byte Magazine, 17(12):??, November 1992. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1992:IAW

[Ano92-77] Anonymous. Information-age warriors: As personal computers invade the battlefield, interoperability becomes crucial. Byte Magazine, 17(7):370–??, July 1992. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1992:IDF

[Ano92-78] Anonymous. Intel’s Double-Fast CPUs: Intel’s 486DX2 puts 50-MHz power in a 25-MHz system. Byte Magazine, 17(5):114– ??, May 1992. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1992:IAS

[Ano92-79] Anonymous. An interface for all senses overview: Kinder, gentler computing: Natural I/O technologies provide computers with the ultimate in user-friendly interfaces. Byte Magazine, 17 (4):134–??, April 1992. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1992:MWPa

[Ano92-80] Anonymous. It’s a multithreaded world, part 1: Multithreaded operating systems are becoming the norm. here’s how your applications can exploit them. Byte Magazine, 17(5):289– ??, May 1992. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1992:MWPb

[Ano92-81] Anonymous. It’s a multithreaded world, part 2: Multithreaded operating systems are taking over. are your applications ready? Byte Magazine, 17(6):351–??, June 1992. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1992:KSW

[Ano92-82] Anonymous. Keeping in step with Windows: CorelDraw and Adobe Illustrator make the grade under Windows. Byte Magazine, 17(8):42–??, August 1992. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1992:KCL

[Ano92-83] Anonymous. Kudos for classic languages, and questions about OS/2. Byte Magazine, 17(4):14– ??, April 1992. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1992:KMP

[Ano92-84] Anonymous. Kudos on Mac programming; rebuttal on relational databases; speech synthesis in York. Byte Magazine, 17(8):16– ??, August 1992. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1992:LMO

[Ano92-85] Anonymous. LAN manager 2.1 opens the gates: Microsoft’s latest showcases TCP/IP, remote access, Mac services, and Net-Ware connectivity. Byte Magazine, 17(4):259–??, April 1992. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1992:LBU

[Ano92-86] Anonymous. LanRover/L: a box that unites PowerBooks and AppleTalk networks. Byte Magazine, 17(7):68–??, July 1992. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1992:LMA

[Ano92-87] Anonymous. LANtastic for Macintosh: Artisoft’s PC-to-Mac connectivity solution offers power, simplicity, and low cost. Byte Magazine, 17(12): 58–??, November 1992. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1992:LMF

[Ano92-88] Anonymous. Laser muscle: Five printers built to handle networks: The BYTE Lab tests five network printers. Byte Magazine, 17 (2):198–??, February 1992. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1992:LSD

[Ano92-89] Anonymous. Let the system do the porting: A new generation of operating systems will help developers write applications for multiple platforms. Byte Magazine, 17(1):191–??, January 1992. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1992:LCP

[Ano92-90] Anonymous. Letters: Color Postscript printers; SGML image editors; and other topics. Byte Magazine, 17(9):20–??, September 1992. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1992:LHG

[Ano92-91] Anonymous. Letters: How to get Unix for free; OS/2 2.0 defended; clarifying the A20 problem; and other issues. Byte Magazine, 17 (12):22–??, November 1992. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1992:LIK

[Ano92-92] Anonymous. Letters: Infoglut kudos; responses to William F. Buckley, Jr.; Usenet news groups; WEB’s compression; and other topics. Byte Magazine, 17 (10):20–??, October 1992. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1992:LUL

[Ano92-93] Anonymous. Letters: Unix lives! Byte Magazine, 17(14): 22–??, December 1992. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1992:LBF

[Ano92-94] Anonymous. Local bus fuels PowerMate’s graphics response: NEC’s local-bus design delivers outstanding graphics performance. Byte Magazine, 17 (4):269–??, April 1992. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1992:LAA

[Ano92-95] Anonymous. Looking for answers? ask Muse: Occam’s Muse makes it easier for Mac users to find answers in mountains of data. Byte Magazine, 17(6):327– ??, June 1992. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1992:MC

[Ano92-96] Anonymous. Mac clones. Byte Magazine, 17(5):10–??, May 1992. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1992:MLI

[Ano92-97] Anonymous. Mac LC II: The sequel: With its new CPU, the Mac LC is better (and cheaper) the second time around. Byte Magazine, 17(8):261–??, August 1992. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1992:MPG

[Ano92-98] Anonymous. MacinStor, a portable gigabyte drive for the Mac. Byte Magazine, 17(9):??, September 1992. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1992:MIX

[Ano92-99] Anonymous. Macintosh impersonator: Xcelerated Systems’ Liken brings Mac applications to Unix workstations. Byte Magazine, 17(8):257–??, August 1992. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1992:MPT

[Ano92-100] Anonymous. Macs and PCs: Together at last: Farallon’s update of an old Mac product encompasses PCs. Byte Magazine, 17 (2):39–??, February 1992. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1992:MSM

[Ano92-101] Anonymous. Making sense of multimedia: Multimedia is flashy, but is it practical? Byte Magazine, 17(5):107–??, May 1992. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1992:MSA

[Ano92-102] Anonymous. Memory and storage advances overview: Scaling the memory pyramid: Memory and mass-storage subsystems traditionally lag behind the theoretical performance limits of CPUs. systems designers are minimizing the performance penalty by organizing storage in a hierarchy of speed and capacity. Byte Magazine, 17(3):160– ??, March 1992. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1992:MMG

[Ano92-103] Anonymous. The methods of madness: Game theorists shouldn’t play with nuclear weapons. Byte Magazine, 17 (5):370–??, May 1992. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1992:MAS

[Ano92-104] Anonymous. Microbytes: Apple’s strategy, which first started shaping up in late 1990, is paying dividends. Byte Magazine, 17(8):18–??, August 1992. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1992:MCL

[Ano92-105] Anonymous. Microbytes: Chips to the left, chips to the right: Microprocessors and chip sets are delivering more power and functionality in smaller form factors. Byte Magazine, 17(1):27–??, January 1992. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1992:MDH

[Ano92-106] Anonymous. MICROBYTES:

DEC hopes its new Alpha microprocessor will take it into the next 25 years of computing. Byte Magazine, 17(5):23– ??, May 1992. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1992:MKA

[Ano92-107] Anonymous. Microbytes: Kaleida, the Apple/IBM multimedia venture. Byte Magazine, 17(9): 24–??, September 1992. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1992:MLY

[Ano92-108] Anonymous. Microbytes: Last year, reports surfaced about a Bill Gates memo that allegedly discussed the “nightmare” of IBM “attacking” Microsoft in systems software and Novell “defeating” the company in networking. now, Windows NT appears poised to turn the tables. Byte Magazine, 17(7):25– ??, July 1992. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1992:MNC

[Ano92-109] Anonymous. Microbytes: Nutek continues to work on a computer that will run mainstream applications for the Mac under Motif. Byte Magazine, 17(6):25– ??, June 1992. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1992:MSC

[Ano92-110] Anonymous. Microbytes: SPARCcompatible workstations priced similarly to high-end PCs, are expected by early 1993. Byte Magazine, 17(14):30–??, December 1992. CODEN BYTEDJ.

ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1992:MFV

[Ano92-111] Anonymous. Microbytes: The first version of the P5 will run about twice as fast as a 66-MHz 486 DX2. Byte Magazine, 17 (10):26–??, October 1992. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1992:MMW

[Ano92-112] Anonymous. Microbytes: With Microsoft’s Windows NT nearing availability, IBM is working hard to improve OS/2 2.0. Byte Magazine, 17(12):28–??, November 1992. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1992:MRS

[Ano92-113] Anonymous. Microsoft responds; setting a page frame; and other issues. Byte Magazine, 17 (5):305–??, May 1992. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1992:MWD

[Ano92-114] Anonymous. Mirror Worlds:

David Gelernter’s Mirror Worlds puts the universe in a shoebox. Byte Magazine, 17(3):362– ??, March 1992. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1992:MCT

[Ano92-115] Anonymous. Monochrome to color: Two new technologies provide ways of obtaining brighter and bolder displays that have better resolution. Byte Magazine, 17(7):179–??, July 1992. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1992:MFF

[Ano92-116] Anonymous. MultiView 24: a fast full-color graphics board for Windows. Byte Magazine, 17(7): 68–??, July 1992. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1992:NSO

[Ano92-117] Anonymous. A natural solution: Object-oriented program methods lend structure to the process of evaluating natural-language sentences. Byte Magazine, 17 (2):237–??, February 1992. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1992:NUN

[Ano92-118] Anonymous. NetOctopus 1.1, a useful network manager with an odd name. Byte Magazine, 17 (9):??, September 1992. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1992:NGL

[Ano92-119] Anonymous. NetWare grows lean, not mean: NetWare lite 1.0 earns high marks for simplicity and interoperatibility with server-based NetWare. Byte Magazine, 17(3):246–??, March 1992. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1992:NSN

[Ano92-120] Anonymous. Network sleuth:

Network utilities for the Mac and PC; an E-mail utility for Unix. Byte Magazine, 17(3):279– ??, March 1992. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1992:NMP

[Ano92-121] Anonymous. New MultiSyncs prove that flatter is better: NEC updates the MultiSyncs with four new monitors. Byte Magazine, 17(4):262–??, April 1992. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1992:NUS

[Ano92-122] Anonymous. A new Unix standard: Hewlett–Packard again sets a standard for workstation price and performance. Byte Magazine, 17(3):36–??, March 1992. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1992:NWR

[Ano92-123] Anonymous. New Wave 4.0: a robust Windows desktop and environment package. Byte Magazine, 17(6):58–??, June 1992. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1992:NWP

[Ano92-124] Anonymous. A new wave of portables. Byte Magazine, 17(7): 16–??, July 1992. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1992:NGO

[Ano92-125] Anonymous. Next-Generation operating systems: Making sense of PowerOpen, Taligent, Windows NT, OS/2 2.0, and the competing Unix consortia. Byte Magazine, 17(2):91–??, February 1992. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1992:NMD

[Ano92-126] Anonymous. No more data loss: The BYTE Lab tests six diskarray subsystems: Disk arrays can provide a “hot-swapping” capability that protects your system from drive failure. Byte Magazine, 17(8):238–??, August 1992. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1992:ORTb

[Ano92-127] Anonymous. Objects in real time: Object orientation may be the key enabling computer technology for distributed realtime systems and applications. Byte Magazine, 17(8):187–??, August 1992. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1992:OQP

[Ano92-128] Anonymous. Olivetti Quaderno, a 2-pound subnotebook with Italian style. Byte Magazine, 17 (9):70–??, September 1992. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1992:FDC

[Ano92-129] Anonymous. On-the-Fly disk compression: Diet disk compression utility, automated Apple menus, and organizing Usenet news. Byte Magazine, 17(6):357– ??, June 1992. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1992:ODD

[Ano92-130] Anonymous. OOPS via DDE:

The DDE approach eases entry into the OOPS world in a controlled manner. Byte Magazine, 17(6):145–??, June 1992. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1992:OHT

[Ano92-131] Anonymous. Optiquest 4000D and HiColor Turbo F/X, a winning color combination. Byte Magazine, 17(12):??, November 1992. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1992:MBL

[Ano92-132] Anonymous. OS/2 2.0: A mixed blessing: The latest version excels at DOS multitasking. Byte Magazine, 17(8):247– ??, August 1992. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1992:OMT

[Ano92-133] Anonymous. Overview: A moving target: Cross-platform development presents unique challenges in this era of proliferating platforms and fragmenting standards. Byte Magazine, 17 (1):158–??, January 1992. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1992:ODT

[Ano92-134] Anonymous. Overview: Display technologies: Soon you’ll see an assortment of display technologies, but CRTs won’t go away. Byte Magazine, 17(7):158– ??, July 1992. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1992:ORTa

[Ano92-135] Anonymous. Overview: Real-Time computing: The techniques developed to serve realtime applications — some of the toughest challenges in computing — are extending the horizons of computer technology. Byte Magazine, 17(8):154– ??, August 1992. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1992:OED

[Ano92-136] Anonymous. Oxford English Dictionary, a 20-volume reference on CD-ROM. Byte Magazine, 17 (9):??, September 1992. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1992:PPN

[Ano92-137] Anonymous. A pair of paradoxes: A new pair of database managers from Borland. Byte Magazine, 17(8):32–??, August 1992. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1992:PWP

[Ano92-138] Anonymous. Paper works: paper becomes a computer interface. Byte Magazine, 17(6):58– ??, June 1992. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1992:PLA

[Ano92-139] Anonymous. PCMCIA LAN adapter from Xircom; new pointing devices; and more. Byte Magazine, 17(13):??, ?? 1992. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1992:PIS

[Ano92-140] Anonymous. Pen-Input systems. Byte Magazine, 17(4):10– ??, April 1992. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1992:PJU

[Ano92-141] Anonymous. Perl: Not just for Unix: A multiplatform utility language, DOS disk utilities, and a Mac file typer. Byte Magazine, 17(2):257–??, February 1992. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1992:PRP

[Ano92-142] Anonymous. Photo-Realism:

Photo-realistic 3-D images are now within your grasp with the help of a mix of sophisticated techniques and an abundance of computer power. Byte Magazine, 17(5):167–??, May 1992. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1992:PPWa

[Ano92-143] Anonymous. Picture Publisher 3.0: Windows image enhancement gets polished. Byte Magazine, 17(6):58–??, June 1992. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1992:PSA

[Ano92-144] Anonymous. The power of speech: Automatic speech recognition is going to radically alter how you interact with computers. Byte Magazine, 17(4):151– ??, April 1992. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1992:PTV

[Ano92-145] Anonymous. Power tools for Visual Basic: Microsoft’s new toolkit extends the Visual Basic programmer’s reach. Byte Magazine, 17(8):259–??, August 1992. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1992:PDVa

[Ano92-146] Anonymous. Practical desktop video, part 2: Raw material: Plan and shoot your own professional video. Byte Magazine, 17(5):129–??, May 1992. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1992:PDVb

[Ano92-147] Anonymous. Practical desktop video, part 3: Winning graphics: In this installment, you’ll learn how to build attentiongrabbing graphics into your desktop videos. Byte Magazine, 17 (6):131–??, June 1992. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1992:PDVc

[Ano92-148] Anonymous. Practical desktop video, part 4: Making the cut: It ain’t over till it’s edited. Byte Magazine, 17 (7):143–??, July 1992. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1992:PTT

[Ano92-149] Anonymous. Precision times three: With micro-based 3-D CAD packages, you don’t have to rob a bank in order to draw one. Byte Magazine, 17(5):192– ??, May 1992. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1992:PQW

[Ano92-150] Anonymous. Print queue ways to keep it lucid: Designing better documents, Windows programming, and object-oriented technology explained. Byte Magazine, 17(4):366–??, April 1992. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1992:PPWb

[Ano92-151] Anonymous. Procomm Plus for Windows: a Windows version of the venerable communications program. Byte Magazine, 17(6):

58–??, June 1992. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1992:PPP

[Ano92-152] Anonymous. ProTracer, this printer/plotter incorporates exquisite detail. Byte Magazine, 17(10): 58–??, October 1992. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1992:QPM

[Ano92-153] Anonymous. QMS-PS 815 MR and IBM LaserPrinter 10 model 30: a new crop of desktop Post-Script printers emerges. Byte Magazine, 17(2):42–??, February 1992. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1992:QWN

[Ano92-154] Anonymous. Quicken 2.0 for Windows, new features make this popular finance program even better. Byte Magazine, 17 (12):??, November 1992. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1992:RUL

[Ano92-155] Anonymous. R4000 upgrade, a leap in power for SGI’s Iris Indigo. Byte Magazine, 17(14):??, December 1992. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1992:RCE

[Ano92-156] Anonymous. Radiosity: Computing the effects of indirect lighting for use in 3-D rendering. Byte Magazine, 17(5):173– ??, May 1992. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1992:RCN

[Ano92-157] Anonymous. Raising the ceiling: Nine memory managers for today’s processors: Nine products that make more memory available to your DOS programs. Byte Magazine, 17 (3):240–??, March 1992. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1992:RTG

[Ano92-158] Anonymous. Real time goes home: Real-time operating systems bring multimedia into the home. Byte Magazine, 17(8):195– ??, August 1992. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1992:RTP

[Ano92-159] Anonymous. Real-Time posix:

Portability and openness finally come to real-time applications through Posix. Byte Magazine, 17(8):177–??, August 1992. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1992:RLD

[Ano92-160] Anonymous. Reclaiming lost disk space; extending computer life; and spelling-checker problems. Byte Magazine, 17(8):295– ??, August 1992. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1992:RMS

[Ano92-161] Anonymous. Reshaping the microchip: The scientists and engineers that brought you the microcomputer revolution are exploring new technologies for building the next generation of chips. Byte Magazine, 17(2): 137–??, February 1992. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1992:RGA

[Ano92-162] Anonymous. Resource guide: Accelerator, coprocessor, and multiprocessor boards. Byte Magazine, 17(2):158–??, February 1992. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1992:RGC

[Ano92-163] Anonymous. Resource guide: Cross-Platform application development tools. Byte Magazine, 17(1):201–??, January 1992. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1992:RGD

[Ano92-164] Anonymous. Resource guide: Displays. Byte Magazine, 17 (7):202–??, July 1992. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1992:RGO

[Ano92-165] Anonymous. Resource guide: Object-Oriented database managers. Byte Magazine, 17(14): 172–??, December 1992. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1992:RGRb

[Ano92-166] Anonymous. Resource guide:

Real-Time operating systems. Byte Magazine, 17(8):201–??, August 1992. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1992:RGRa

[Ano92-167] Anonymous. Resource guide:

Realistic 3-D rendering and Volume-Imaging software. Byte Magazine, 17(5):190–??, May 1992. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1992:RGS

[Ano92-168] Anonymous. Resource guide:

Storage for networks. Byte Magazine, 17(3):204–??, March 1992. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1992:RNL

[Ano92-169] Anonymous. REVIEWER’s NOTEBOOK: A local bus/ accelerator combo for Windows and two parallel-port hard drives. Byte Magazine, 17 (5):277–??, May 1992. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1992:RNSb

[Ano92-170] Anonymous. REVIEWER’s NOTEBOOK: A superfast super server and wireless print sharing. Byte Magazine, 17(4):272– ??, April 1992. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1992:RNSc

[Ano92-171] Anonymous. REVIEWER’s NOTEBOOK: S3’s 86C911 GUI accelerator, and a new Micro-Phone for the Mac. Byte Magazine, 17(6):342–??, June 1992. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1992:RNSa

[Ano92-172] Anonymous. REVIEWER’s NOTEBOOK: Supercharged SuperCalc, jaggies relief, and a tool to improve Windows. Byte Magazine, 17(1):288–??, January 1992. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1992:RNB

[Ano92-173] Anonymous. REVIEWER’s NOTEBOOK: The BYTE Lab clocks faster speeds for the new version of LANtastic and looks at QueryDOS, a new file manipulator. Byte Magazine, 17(8):264– ??, August 1992. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1992:RSD

[Ano92-174] Anonymous. A ride on the SBus: Design goals and operational details of Sun’s SBus. Byte Magazine, 17(5):283–??, May 1992. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1992:RPT

[Ano92-175] Anonymous. The right profile: Timing DOS file access; debugging Unix code; and getting reminders on the Mac. Byte Magazine, 17(8):288–??, August 1992. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1992:REN

[Ano92-176] Anonymous. RISC enters a new generation: DEC’s Alpha architecture defines a new generation for RISC technologies and systems. Byte Magazine, 17(8):141– ??, August 1992. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1992:RSL

[Ano92-177] Anonymous. RLN spells Long-Distance Ethernet: RLN extends Ethernet connections across phone lines. Byte Magazine, 17(6):339–??, June 1992. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1992:RDD

[Ano92-178] Anonymous. The RTOS difference: Discover how the key features and behaviors of realtime operating systems ensure performance in critical applications. Byte Magazine, 17(8):161– ??, August 1992. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1992:SCC

[Ano92-179] Anonymous. ScanMan color, color your view of hand scanners. Byte Magazine, 17(10):??, October 1992. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1992:SRN

[Ano92-180] Anonymous. Search and retrieval: New methods for managing large document-retrieval systems. Byte Magazine, 17(6):271– ??, June 1992. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1992:SFI

[Ano92-181] Anonymous. SGML frees information: Standard Generalized Markup Language helps you turn mounds of documents into information that can boost your productivity and innovation. Byte Magazine, 17(6):279– ??, June 1992. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1992:SRA

[Ano92-182] Anonymous. A shared resource access manager, part 1: How do you manage access to shared resources on networks or multiuser systems? Byte Magazine, 17 (8):279–??, August 1992. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1992:SPM

[Ano92-183] Anonymous. Signal processing for multimedia: By combining a DSP chip and a library of multimedia functions under the control of a low-overhead kernel, AT&T aims to bring the power of DSP to multimedia. Byte Magazine, 17(2):105–??, February 1992. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1992:SCB

[Ano92-184] Anonymous. Software corner Batch-File toolkit: Better automation for DOS and Unix, and a way to hunt down Mac viruses. Byte Magazine, 17 (4):295–??, April 1992. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1992:SCO

[Ano92-185] Anonymous. Software corner opening the lines of communication: A functional INT 14th replacement, communications for the Mac, and ZMODEM transfers from Unix. Byte Magazine, 17(7):301–??, July 1992. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1992:SWD

[Ano92-186] Anonymous. Software without walls: Distributed object management systems can fuse diverse distributed applications and data into seamless information systems. Byte Magazine, 17(3):122– ??, March 1992. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1992:SFD

[Ano92-187] Anonymous. Solutions focus database building blocks: Database libraries with the horsepower to build powerful multiuser applications. Byte Magazine, 17(1):204–??, January 1992. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1992:SFF

[Ano92-188] Anonymous. Solutions focus file servers face off: Five highperformance file servers vie for the top spot on tests run with BYTE’s new LAN benchmark suite. Byte Magazine, 17(2): 162–??, February 1992. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1992:SFN

[Ano92-189] Anonymous. Solutions focus new LAN backup tools: The BYTE Lab looks at eight tape backup systems that promise to make your network safe and sound. Byte Magazine, 17(4):192–196, 198, 200, 202, 204, 206, 208, April 1992. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1992:SFS

[Ano92-190] Anonymous. Solutions focus surveying far-flung networks: The top six tools for distributed network monitoring and analysis. Byte Magazine, 17(8):204– ??, August 1992. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1992:SARa

[Ano92-191] Anonymous. Some assembly required Mac programming revealed: The Mac is easy to program, within limits. Byte Magazine, 17(4):279–??, April 1992. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1992:SARb

[Ano92-192] Anonymous. Some assembly required Macintosh menus revealed: Creating the right menus for your Mac applications. Byte Magazine, 17(7):293–??, July 1992. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1992:SRP

[Ano92-193] Anonymous. SPARCs on the road: Portable Unix and a lot more. Byte Magazine, 17(5):40– ??, May 1992. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1992:SBP

[Ano92-194] Anonymous. Stop bit the productivity macguffin: Networks and portable computers are helping companies finally realize gains from their computing investments. Byte Magazine, 17 (8):360–??, August 1992. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1992:SMN

[Ano92-195] Anonymous. Storage management: A new class of products eases the burden of the LAN administrator’s job. Byte Magazine, 17(3):183–??, March 1992. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1992:SYL

[Ano92-196] Anonymous. Support your local CPU: With standard functions integrated into the CPU, a new class of specialized coprocessors is beginning to appear. Byte Magazine, 17(2):151–??, February 1992. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1992:SPW

[Ano92-197] Anonymous. Swift programming for Windows, in Windows: QuickC for Windows brings GUI integration to Windows program development. Byte Magazine, 17 (3):251–??, March 1992. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1992:SBS

[Ano92-198] Anonymous. System bus or system bottleneck: The 32-bit EISA and Micro Channel buses are not living up to their potential. Byte Magazine, 17(3):131– ??, March 1992. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1992:TTU

[Ano92-199] Anonymous. T4400SXC, Toshiba unveils its new 486 notebook with an active color display: NetMounter, connecting your Macs to a file server over Ethernet doesn’t have to cost a lot, Networks Connect, Symbiotics brings network awareness to the Windows clipboard I325VM Floptical, a new kind of drive has arrived Ad Lib Gold 1000, a second-generation audio board. Byte Magazine, 17(5):42– ??, May 1992. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1992:TSU

[Ano92-200] Anonymous. Tapping into sockets: Use TCP/IP sockets to write portable client/server applications. Byte Magazine, 17 (3):261–??, March 1992. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1992:TNT

[Ano92-201] Anonymous. TCNS: a network from Thomas-Conrad that doesn’t compromise speed. Byte Magazine, 17(6):58–??, June 1992. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1992:TCO

[Ano92-202] Anonymous. TOMORROW’s CHIPS overview: Built for speed: Advances in processor chips and architectures are driving the next generation of computing. Byte Magazine, 17(2):122–??, February 1992. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1992:TDT

[Ano92-203] Anonymous. Tributaries and deltas: Tracking software change in multiplatform environments. Byte Magazine, 17(1):177–??, January 1992. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1992:TWN

[Ano92-204] Anonymous. Tweaking Windows: New adapters boost speed and clarity: Six new graphics adapters promise to make Windows more efficient. Byte Magazine, 17(1):250–??, January 1992. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1992:TSF

[Ano92-205] Anonymous. Two steps forward, one step back: Where do we go after relational databases? Byte Magazine, 17(5):141–??, May 1992. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1992:TTQ

[Ano92-206] Anonymous. Two tools of the QuickTime trade: SuperMac’s VideoSpigot and Adobe’s Premiere help you create digital movies easily. Byte Magazine, 17(6):336–??, June 1992. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1992:UUI

[Ano92-207] Anonymous. The ultimate user interface: Information systems create environments in which we work. new research aims to allow better design of these information systems. Byte Magazine, 17(4):175–??, April 1992. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1992:UHK

[Ano92-208] Anonymous. Under the hood keyboards without keys: Touchscreens aren’t just for kiosks anymore. Byte Magazine, 17(7):287– ??, July 1992. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1992:UHW

[Ano92-209] Anonymous. Under the hood wireless networking: New technologies and standards portend the era of mobile computing. Byte Magazine, 17(4):291–??, April 1992. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1992:VWC

[Ano92-210] Anonymous. Video for Windows, coming soon to a PC near you. Byte Magazine, 17(14):??, December 1992. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1992:VGD

[Ano92-211] Anonymous. Video goes digital with fluency: Developers get a head start with this digital-video bundle for PCs. Byte Magazine, 17(4):267–??, April 1992. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1992:VDD

[Ano92-212] Anonymous. Voxels: Data in 3-D: Now you can use voxels to sample 3-D space. Byte Magazine, 17(5):177–??, May 1992.

CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1992:WDX

[Ano92-213] Anonymous. Was Desqview/X worth the wait: You can run DOS, X, and Windows programs locally or remotely. Byte Magazine, 17(8):249–??, August 1992. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1992:WWT

[Ano92-214] Anonymous. Ways for Windows, translation the easy way. Byte Magazine, 17(12):??, November 1992. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1992:WSC

[Ano92-215] Anonymous. What to stash in a cache: Today, caching is a must for high performance. now, the questions are: What type, and how big? Byte Magazine, 17 (3):175–??, March 1992. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1992:WYS

[Ano92-216] Anonymous. What you see is what you solve: Mathematica and MathCAD for Windows: Mathematica and Math-CAD take different approaches to tackling math problems using the graphical Windows interface. Byte Magazine, 17(5):263– ??, May 1992. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1992:WNH

[Ano92-217] Anonymous. Whats’ new: HyperSpace Shuttle handles up to 3 GB of on-line storage; NetRunner integrates voice, fax, and network traffic; and more. Byte Magazine, 17(9):88–??, September 1992. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1992:WNF

[Ano92-218] Anonymous. What’s new: The Freestyle/SL notebook’s screen swivels; IDL for Windows lets you do high-power scientific computing; and more. Byte Magazine, 17(12):76–??, November 1992. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1992:WNQ

[Ano92-219] Anonymous. WHAT’s NEW:

The QMS-PS 1700; RangeLAN/ ISA; PowerPlay; and more. Byte Magazine, 17(5):62–??, May 1992. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1992:WNS

[Ano92-220] Anonymous. WHAT’s NEW:

The SPARCard 2 joins DOS and Unix; the I/O Station 464 collects data remotely; and more. Byte Magazine, 17(7):76– ??, July 1992. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1992:WNT

[Ano92-221] Anonymous. What’s new: Trakker backs up from the enhanced parallel port; PC/ Television puts TV on your PC; and more. Byte Magazine, 17 (10):68–??, October 1992. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1992:WNW

[Ano92-222] Anonymous. What’s new: Watch TV on the 486SX/25 multimedia computer system; centrally manage electrical power to local or remote LAN components with LanSafe II; and more. Byte Magazine, 17(14):72–??, December 1992. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1992:WGR

[Ano92-223] Anonymous. Windows goes real time: The iRMX for Windows operating system provides the best of both worlds on a single PC. Byte Magazine, 17(4):119– ??, April 1992. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1992:WPS

[Ano92-224] Anonymous. Windows printer shines in speed, resolution: LaserMaster’s new laser printer does Windows. Byte Magazine, 17(5):269–??, May 1992. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1992:WSS

[Ano92-225] Anonymous. Windows sound system: Microsoft enters the sound board market. Byte Magazine, 17(12):??, November 1992. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1992:WWB

[Ano92-226] Anonymous. WordPerfect for Windows: The big-selling word processor is finally running under Windows. has it been worth the wait? Byte Magazine, 17(3):257– ??, March 1992. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1993:ABP

[Ano93a] Anonymous. Acrobat bounds into the paperless publishing arena: Adobe’s Acrobat has a tough balancing act in delivering electronic, no-font-hassle documents to the screen. Byte Magazine, 18(9):30–??, August 1993. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1993:ADI

[Ano93b] Anonymous. AMD declared independence: AMD unveils the first of a new line of processors. Byte Magazine, 18(10): 46–??, September 1993. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1993:ARL

[Ano93c] Anonymous. Apple revamps its lineup: We help you sort out Apple’s newest line of desktops and preview new Apple duos. Byte Magazine, 18(13):23–??, December 1993. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1993:ABA

[Ano93d] Anonymous. Ask BYTE: Adobe Type Manager conflicts, the Next as document manager, and more. Byte Magazine, 18(2): 239–??, February 1993. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1993:ABI

[Ano93e] Anonymous. Ask BYTE: Internet access solutions, Procomm problems, and other issues. Byte Magazine, 18(1):278–??, January 1993. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1993:ABL

[Ano93f] Anonymous. Ask BYTE: Laptop parallel-port problems and creating dynamic arrays. Byte Magazine, 18(4):215–??, April 1993. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1993:AYO

[Ano93g] Anonymous. At your own risk: Faster modems now, standards later: Modem manufacturers that can’t wait another year for a new standard are shipping nonstandard product. Byte Magazine, 18(12):38–??, November 1993. CODEN BYTEDJ.

ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1993:AEI

[Ano93h] Anonymous. AudioMan: An easy and inexpensive approach to Windows sound. Byte Magazine, 18(4):??, April 1993. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1993:BWF

[Ano93i] Anonymous. Behind the wheel of the first Newton and Zoomer PDAs: Test drives of prerelease Apple, Sharp, and Casio Zoomer PDAs reveal rough edges and a lot of promise. Byte Magazine, 18 (10):22–??, September 1993. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1993:BDCc

[Ano93j] Anonymous. The best drives in capacities from 1 to 2 GB. Byte Magazine, 18(10):188–??, September 1993. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1993:BDCa

[Ano93k] Anonymous. The best drives in capacities from 250 to 350 MB. Byte Magazine, 18(10): 178–??, September 1993. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1993:BDCb

[Ano93l] Anonymous. The best drives in capacities from 400 to 600 MB. Byte Magazine, 18(10):

184–??, September 1993. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1993:CSL

[Ano93m] Anonymous. Cirrus subsidiary leads CDPD push: Analog cellular phone services will soon carry digital data, too. Byte Magazine, 18(12):28–??, November 1993. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1993:CMM

[Ano93n] Anonymous. Color management makes color easier to use. Byte Magazine, 18(13):136–??, December 1993. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1993:CPC

[Ano93o] Anonymous. ColorFrame: A portable color display for Macs. Byte Magazine, 18(1):??, January 1993. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1993:CMP

[Ano93p] Anonymous. Compel: Multimedia presentation software from Asymetrix. Byte Magazine, 18 (6):??, May 1993. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1993:CEP

[Ano93q] Anonymous. Compilers: Essential partners. Byte Magazine, 18(12):135–??, November 1993.

CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1993:DWF

[Ano93r] Anonymous. Datafax: A Windows fax program. Byte Magazine, 18(1):??, January 1993. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1993:DMU

[Ano93s] Anonymous. Demand for multimedia upgrade kits growing: If current trends continue, CD-ROM will soon gain wide acceptance in the business world. Byte Magazine, 18(13):38–??, December 1993. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1993:DCO

[Ano93t] Anonymous. Developers cautiously optimistic about multiplatform OpenDoc. Byte Magazine, 18(12):130–??, November 1993. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1993:DCK

[Ano93u] Anonymous. Document conferencing keeps data close-by. Byte Magazine, 18(10):66–??, September 1993. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1993:MUW

[Ano93v] Anonymous. E-Mail unplugged by wireless WANs: E-mail will be an important application for wide-area, wireless communications. Byte Magazine, 18(12): 28–??, November 1993. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1993:EWR

[Ano93w] Anonymous. Editorial: Windows reaches beyond the desktop. Byte Magazine, 18(5):6– ??, Spring 1993. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1993:EMC

[Ano93x] Anonymous. Encarta: A multimedia CD-ROM encyclopedia worth exploring. Byte Magazine, 18(4):??, April 1993. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1993:EPG

[Ano93y] Anonymous. Epson Progression: A 486/33 with a graphicsacceleration Wingine. Byte Magazine, 18(1):??, January 1993. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1993:EGS

[Ano93z] Anonymous. Excel 5.0 gets smart: Excel 5.0 offers much in the way of new features. even more intriguing is the tighter integration with members of Microsoft’s Office package, such as Word 6.0 for Windows. Byte Magazine, 18(12):23–??, November 1993. CODEN BYTEDJ.

ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1993:FAP

[Ano93-27] Anonymous. Falcon030, Atari’s PC with a DSP. Byte Magazine, 18(2):??, February 1993. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1993:FTF

[Ano93-28] Anonymous. Finding text fast.

Byte Magazine, 18(10):128–??, September 1993. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1993:FPS

[Ano93-29] Anonymous. First PowerPC systems hit the street: The first IBM PowerPC systems should be available soon. Byte Magazine, 18(12):34–??, November 1993. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1993:FFG

[Ano93-30] Anonymous. Flexscan F340iW, Graphite Card, and Paradise Accelerator: Working toward clearer Windows. Byte Magazine, 18(1):??, January 1993. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1993:FCP

[Ano93-31] Anonymous. FoxPro 2.5, the cross-platform strategy begins to pay off. Byte Magazine, 18 (2):54–??, February 1993. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1993:FGW

[Ano93-32] Anonymous. Freelance Graphics for Windows 2.0, prepare presentations painlessly. Byte Magazine, 18(2):??, February 1993. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1993:GLH

[Ano93-33] Anonymous. The great light hope. Byte Magazine, 18(9):124– ??, August 1993. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1993:HM

[Ano93-34] Anonymous. Honorable mentions. Byte Magazine, 18(10): 190–??, September 1993. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1993:HWT

[Ano93-35] Anonymous. How we tested. Byte Magazine, 18(10):182–??, September 1993. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1993:HLM

[Ano93-36] Anonymous. HP LaserJet 4Si and 4Si MX: Hewlett–Packard’s newest network printers. Byte Magazine, 18(6):??, May 1993. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1993:IAW

[Ano93-37] Anonymous. IBM and Apple work to perfect voice input: Two technologies that will eventually let you interact with your computer solely through spoken commands. Byte Magazine, 18 (9):32–??, August 1993. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1993:IAB

[Ano93-38] Anonymous. IBM announces “better” DOS than MS-DOS: IBM releases a new version of PC-DOS. Byte Magazine, 18 (10):46–??, September 1993. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1993:IHH

[Ano93-39] Anonymous. IBM has high hopes for multimedia on OS/2. Byte Magazine, 18(12):110–??, November 1993. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1993:IMM

[Ano93-40] Anonymous. IBM makes MP promises for OS/2. Byte Magazine, 18(12):114–??, November 1993. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1993:ISD

[Ano93-41] Anonymous. IBM ships DCE SDK for OS/2 and Windows. Byte Magazine, 18(12):131–??, November 1993. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1993:IAD

[Ano93-42] Anonymous. IBM’s assault on distributed objects. Byte Magazine, 18(12):125–??, November 1993. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1993:INF

[Ano93-43] Anonymous. Illustrator 5.0: New face, new features: The new version of Adobe’s drawing program for the Mac offers layering, gradient fills, and an interface makeover. Byte Magazine, 18(9):22– ??, August 1993. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1993:IPF

[Ano93-44] Anonymous. InFax Pro 3.0: Fax software with OCR and a cool cover-page designer. Byte Magazine, 18(1):??, January 1993. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1993:IAA

[Ano93-45] Anonymous. Intel, AT&T, and AMD continue the chase: New PDA chips from AMD, AT&T microelectronics, and Intel/VLSI are slated to ship in 1994. Byte Magazine, 18(13):28–??, December 1993. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1993:IVS

[Ano93-46] Anonymous. Intel’s VDI speeds up video, miffs Microsoft: Microsoft and Intel are at odds on how to improve the video playback in video for Windows. Byte Magazine, 18(12):42–??, November 1993. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1993:LRH

[Ano93-47] Anonymous. Lab report: 32 high-speed hard drives: We stress-tested 32 hard drives to find the best performers in capacities from 250 MB to 2 GB. Byte Magazine, 18(10):176–??, September 1993. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1993:LRN

[Ano93-48] Anonymous. Lab report: Network connections: 100 Ethernet cards: We tested 100 Ethernet network interface cards and picked the best 8-and 16bit cards for workgroups, large networks, and transaction-based networks; NICs for workgroups; how we tested; NICs for large networks; wireless LAN adapters; NICs for transaction processing; honorable mentions; dubious achievements. Byte Magazine, 18 (9):172–??, August 1993. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1993:LQD

[Ano93-49] Anonymous. Laser quality without the drum. Byte Magazine, 18 (13):96–??, December 1993. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1993:LCC

[Ano93-50] Anonymous. Letters: Choosing a CPU; low-cost PCs; BYTE’s Windows issue; and other topics. Byte Magazine, 18(3):20– ??, March 1993. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1993:LCT

[Ano93-51] Anonymous. Letters: Coming to terms with software gluttony; OS/2 hassles; machine translation; and other topics. Byte Magazine, 18(6):20–??, May 1993. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1993:LCO

[Ano93-52] Anonymous. Letters: Comments on OCR, Windows development, Pournelle, and PowerPC. Byte Magazine, 18(12):18–??, November 1993. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1993:LFF

[Ano93-53] Anonymous. Letters: Fighting fatware, MS-DOS 6, the Commodore Amiga, and other topics. Byte Magazine, 18(8):19– ??, July 1993. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1993:LPP

[Ano93-54] Anonymous. Letters: From Pentium to printers, readers register their comments. Byte Magazine, 18(9):18–??, August 1993. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1993:LOS

[Ano93-55] Anonymous. Letters: Operatingsystem trends, OS/2 at a crossroads, digital photography, the ultimate workstation, and other issues. Byte Magazine, 18(2): 20–??, February 1993. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1993:LOV

[Ano93-56] Anonymous. Letters: Opinions on video computing, SGML, India’s software development, and installer hell. Byte Magazine, 18 (13):18–??, December 1993. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1993:LRRb

[Ano93-57] Anonymous. Letters: Readers react to BYTE’s new look. Byte Magazine, 18(10):18–??, September 1993. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1993:LRRa

[Ano93-58] Anonymous. Letters: Readers respond on E-mail, Mac accelerator boards, computing in Italy, and other topics. Byte Magazine, 18(7):18–??, June 1993. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1993:LRB

[Ano93-59] Anonymous. Letters: Responses on BYTE’s global perspective, the October editorial, and other topics. Byte Magazine, 18(1):20– ??, January 1993. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1993:LAB

[Ano93-60] Anonymous. Letters: The Amiga 3000T-040/200 and 4000-040/120, OS/2 2.0, Braincel defended, and other reader mail. Byte Magazine, 18(4):20–??, April 1993. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1993:LLF

[Ano93-61] Anonymous. Lotus 1-2-3 for OS/2 2.0 and Lotus Freelance Graphics for OS/2 2.0: Two major applications for IBM’s OS. Byte Magazine, 18(3):46– ??, March 1993. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1993:MP

[Ano93-62] Anonymous. Magazines without paper. Byte Magazine, 18(10): 108–??, September 1993. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1993:MPM

[Ano93-63] Anonymous. Many processors, many threads. Byte Magazine, 18(7):84–??, June 1993. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1993:MBU

[Ano93-64] Anonymous. Microbytes: Bringing UNIX into the NetWare fold presents Novell with a marketing dilemma: what to tell software developers who ask which platform to write applications for. Byte Magazine, 18(6):24– ??, May 1993. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1993:MHD

[Ano93-65] Anonymous. Microbytes: HP, DEC, and Sun unveil new highpowered workstations. Byte Magazine, 18(2):32–??, February 1993. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1993:MPI

[Ano93-66] Anonymous. Microbytes: Preliminary info on Motorola’s 68060 is impressive, but the PowerPC 601 will probably ship first. Byte Magazine, 18(1):32–??, January 1993. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1993:MTC

[Ano93-67] Anonymous. Microbytes: Tandy and Casio stand poised to compete with Apple’s Newton PDA. Byte Magazine, 18(4):28–??, April 1993. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1993:MFH

[Ano93-68] Anonymous. Microbytes: The first HDTV sets are not likely to appear in U.S. homes before 1995, yet companies from Microsoft to HBO are already involved in programming projects. Byte Magazine, 18(3):24–??, March 1993. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1993:MTWb

[Ano93-69] Anonymous. Microsoft tunes WFW: Microsoft fixes some deficiencies in its workgroup program. Byte Magazine, 18(13): 44–??, December 1993. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1993:MTI

[Ano93-70] Anonymous. microWriter: Texas Instruments’ low-cost laser printer. Byte Magazine, 18(4): ??, April 1993. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1993:MC

[Ano93-71] Anonymous. The Mips challenge. Byte Magazine, 18(7):80– ??, June 1993. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1993:MTWa

[Ano93-72] Anonymous. Moving toward Windows building blocks: Writing an OLE 2.0 application isn’t the easiest thing to do in the world. but the benefits are worth it. Byte Magazine, 18(13):32–??, December 1993. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1993:NNS

[Ano93-73] Anonymous. NetWare 4.0: The next step for a growing network. Byte Magazine, 18(6):42– ??, May 1993. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1993:NGS

[Ano93-74] Anonymous. A new graphics standard from Matrox: Matrox’s new 64-bit video card for the PC represents a new standard for accelerated graphics. Byte Magazine, 18(9):23–??, August 1993. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1993:NRC

[Ano93-75] Anonymous. New RISC chips for Windows NT: Windows NT is sparking hot competition among chip makers who want a piece of the growing market for highperformance desktop PCs and servers. Byte Magazine, 18(13): 42–??, December 1993. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1993:NCN

[Ano93-76] Anonymous. NoteJet 486:

Canon’s notebook/printer makes a dynamic package. Byte Magazine, 18(6):??, May 1993. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1993:NAS

[Ano93-77] Anonymous. NT’s architects speak. Byte Magazine, 18(7):92– ??, June 1993. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1993:UP

[Ano93-78] Anonymous. OS/2 2.1: A user’s perspective. Byte Magazine, 18 (12):97–??, November 1993. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1993:WN

[Ano93-79] Anonymous. OS/2 and Windows networks. Byte Magazine, 18(12):117–??, November 1993. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1993:GDS

[Ano93-80] Anonymous. OS/2 gets device support. Byte Magazine, 18(12): 146–??, November 1993. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1993:PUD

[Ano93-81] Anonymous. PagePlus 2.0: A US$59.95 desktop publishing package. Byte Magazine, 18(6): ??, May 1993. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1993:PNP

[Ano93-82] Anonymous. Painter 2.0: A “natural,” professional, and fun paint tool. Byte Magazine, 18(6): ??, May 1993. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1993:PAO

[Ano93-83] Anonymous. Pandora and the active office. Byte Magazine, 18 (10):76–??, September 1993. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1993:PSU

[Ano93-84] Anonymous. Partners seek to unite phone and PC. Byte Magazine, 18(12):107–??, November 1993. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1993:PAN

[Ano93-85] Anonymous. PCI: Apple’s new bus: The peripheral component interconnect local-bus architecture receives a powerful endorsement. Byte Magazine, 18(9):24– ??, August 1993. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1993:P

[Ano93-86] Anonymous. Pen for OS/2.

Byte Magazine, 18(12):112–??, November 1993. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1993:PC

[Ano93-87] Anonymous. Performance comparisons. Byte Magazine, 18 (7):138–??, June 1993. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1993:PCS

[Ano93-88] Anonymous. Personal communicator 440: Start-up EO does it right the first time. Byte Magazine, 18(3):??, March 1993. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1993:PPE

[Ano93-89] Anonymous. Political primer for enterprise networks. Byte Magazine, 18(13):78–??, December 1993. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1993:PSA

[Ano93-90] Anonymous. The PowerBooks of summer: Apple, TI, and Tadpole deliver a new wave of color notebook computers. Byte Magazine, 18(9):36–??, August 1993. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1993:PFS

[Ano93-91] Anonymous. PowerExec EL: Forgoing some notebook frills means big savings. Byte Magazine, 18(3):??, March 1993. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1993:QPS

[Ano93-92] Anonymous. QMS 1725 print system, the latest 17-ppm screamer. Byte Magazine, 18 (2):??, February 1993. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1993:QGC

[Ano93-93] Anonymous. QuickRing gets closer, expands to networking: Apple and National Semiconductor say QuickRing will be used to deliver 180-MBps data transfer over fiber-optic networks. Byte Magazine, 18(9):27–??, August 1993. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1993:RCA

[Ano93-94] Anonymous. Readers’ choice awards: BYTE readers name their favorite products. Byte Magazine, 18(7):65–??, June 1993. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1993:RRN

[Ano93-95] Anonymous. Reducing router network transmission costs. Byte Magazine, 18(13):70–??, December 1993. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1993:RB

[Ano93-96] Anonymous. Report from Birmingham. Byte Magazine, 18(12): 46–??, November 1993. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1993:RCI

[Ano93-97] Anonymous. Report from Cyprus: Information technology consultants in Cyprus get no respect. Byte Magazine, 18(13): 48–??, December 1993. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1993:RHE

[Ano93-98] Anonymous. Report from Hannover: The European Community takes a community approach to research. Byte Magazine, 18 (9):48–??, August 1993. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1993:RO

[Ano93-99] Anonymous. Report from Orlando. Byte Magazine, 18(12): 48–??, November 1993. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1993:RTT

[Ano93-100] Anonymous. Report from Taiwan: Taiwan is reengineering itself to be a designer of computer products. Byte Magazine, 18 (10):48–??, September 1993. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1993:RGM

[Ano93-101] Anonymous. Resource guide: Machine-Translation software. Byte Magazine, 18(1):185–??, January 1993. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1993:RGP

[Ano93-102] Anonymous. Resource guide: Plugging into wireless. Byte Magazine, 18(2):177–??, February 1993. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1993:RGVb

[Ano93-103] Anonymous. Resource guide: Virus protection for networks. Byte Magazine, 18(6):144–??, May 1993. CODEN BYTEDJ.

ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1993:RGVa

[Ano93-104] Anonymous. Resource guide: Visualization software. Byte Magazine, 18(4):148–??, April 1993. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1993:STB

[Ano93-105] Anonymous. Scanners turn business cards into database records: Is your desk cluttered with business cards that you still haven’t entered into your database? technology comes to the rescue. Byte Magazine, 18(10): 44–??, September 1993. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1993:SMC

[Ano93-106] Anonymous. Serving mobile clients. Byte Magazine, 18 (7):106–??, June 1993. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1993:SSM

[Ano93-107] Anonymous. Short subjects:

MicroPhone Pro for Windows, Dr. Floyd’s Desktop Toys, and more. Byte Magazine, 18(5):134– ??, Spring 1993. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1993:SMD

[Ano93-108] Anonymous. Sony’s MiniDisc for data: Future floppy: Sony’s new medium for storage. Byte Magazine, 18(12):32–??, November 1993. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1993:SLS

[Ano93-109] Anonymous. SPARCstation LX: Sun’s feature-packed little Unix box. Byte Magazine, 18(3):??, March 1993. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1993:SRT

[Ano93-110] Anonymous. Special report: Is there a better Windows 3.1 than Windows 3.1: Windows, OS/2, and NT offer distinct benefits and disadvantages. one excels at running standard Windows applications. Byte Magazine, 18 (12):85–??, November 1993. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1993:SDB

[Ano93-111] Anonymous. Spreadsheet or database? the best of both: A new breed turns the traditional spreadsheet metaphor on its head. Byte Magazine, 18(10): 29–??, September 1993. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1993:SL

[Ano93-112] Anonymous. Storage without limits. Byte Magazine, 18(9): 104–??, August 1993. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1993:SEB

[Ano93-113] Anonymous. Stylus 800: Epson is back in the ink-jet business. Byte Magazine, 18(4):55– ??, April 1993. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1993:TPP

[Ano93-114] Anonymous. Talking printer politics. Byte Magazine, 18(13): 110–??, December 1993. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1993:TIA

[Ano93-115] Anonymous. TCP/IP addressing. Byte Magazine, 18(13): 84–??, December 1993. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1993:TWD

[Ano93-116] Anonymous. Three ways to 3-D:

Creating 3-D graphics images in real time requires sophisticated and well-balanced computergraphics pipelines. Byte Magazine, 18(12):215–??, November 1993. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1993:TSW

[Ano93-117] Anonymous. Top software for Windows and OS/2. Byte Magazine, 18(12):151–??, November 1993. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1993:TPB

[Ano93-118] Anonymous. Turbo Pascal 7.0 and Borland Pascal with Objects 7.0: A new generation of Pascal. Byte Magazine, 18(1):54–??, January 1993. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1993:TPA

[Ano93-119] Anonymous. TyIN 2000, a packed adapter card. Byte Magazine, 18(2):??, February 1993. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1993:UDS

[Ano93-120] Anonymous. Unix database servers are not for everyone. Byte Magazine, 18(7):102–??, June 1993. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1993:VCS

[Ano93-121] Anonymous. Video compression standards vie for acceptance. Byte Magazine, 18(10): 72–??, September 1993. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1993:VWS

[Ano93-122] Anonymous. VideoSpigot for Windows, SuperMatch’s videocapture board. Byte Magazine, 18(2):??, February 1993. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1993:VNA

[Ano93-123] Anonymous. Visio: A new approach to graphics. Byte Magazine, 18(1):??, January 1993. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1993:WHE

[Ano93-124] Anonymous. A walk on the high end. Byte Magazine, 18(13): 156–??, December 1993. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1993:WNC

[Ano93-125] Anonymous. What’s new: A color and gray-scale scanner; external and internal frame grabbers; a 3-D manipulator; graphics software; and more. Byte Magazine, 18(13):243–??, December 1993. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1993:WNW

[Ano93-126] Anonymous. What’s new: A workstation line with zippy graphics, a notebook with an active-matrix display, portable multimedia, a RAID-5 array, CAD tools, and more. Byte Magazine, 18(10):236–??, September 1993. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1993:WNB

[Ano93-127] Anonymous. What’s new:

BriteLite LX puts SPARC power on the road, LANDesk simplifies network management, and more. Byte Magazine, 18(3):58– ??, March 1993. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1993:WNF

[Ano93-128] Anonymous. What’s new: Fargo’s Primera color printer, Octocom V.fast modems, and more. Byte Magazine, 18(9):220– ??, August 1993. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1993:WNT

[Ano93-129] Anonymous. What’s new: Tektronix offers printers for workgroups, SunConnect extends LANs to global networks, and more. Byte Magazine, 18(2): 72–??, February 1993. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1993:WND

[Ano93-130] Anonymous. What’s new: The DTR-1 is a notebook or a pen computer, the SmartLink V32bis FaxModem encrypts your data, LapCAD 5 for the Mac gives you finite modeling, and more. Byte Magazine, 18(6):57– ??, May 1993. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1993:WNL

[Ano93-131] Anonymous. What’s new: The latest Tektronix dye-sublimation printer, Smartcom for Windows from Hayes, Alps Electric’s wireless LAN adapter, and more. Byte Magazine, 18(8):226–??, July 1993. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1993:WNP

[Ano93-132] Anonymous. What’s new:

The Paperless 1 imaging system reduces paper consumption, the Raidion LT provides faulttolerant disk-array storage for OS/2, and more. Byte Magazine, 18(7):220–??, June 1993. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1993:WNSa

[Ano93-133] Anonymous. What’s new: The ScreenStar displays two fullsize documents; the Digibot II “reads” multidimensional objects; and more. Byte Magazine, 18(1):72–??, January 1993. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1993:WNX

[Ano93-134] Anonymous. What’s new:

X Window System and dual-Pentium systems; a digital, online video production system for the Mac; a cellular link for modems and faxes; project scheduling in Windows; and more. Byte Magazine, 18(12): 342–??, November 1993. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1993:WOD

[Ano93-135] Anonymous. Will OpenGL be 3-D standard for Windows NT? Byte Magazine, 18(12):144–??, November 1993. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1993:WG

[Ano93-136] Anonymous. Windows graphics. Byte Magazine, 18(7):90– ??, June 1993. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1993:WNV

[Ano93-137] Anonymous. Windows news and views: Claris Works, VL-Bus accelerators, WordPerfect Presentations, Windows CAD, and more. Byte Magazine, 18(5):10– ??, Spring 1993. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1993:WNM

[Ano93-138] Anonymous. Windows no Mac at DTP: In desktop publishing, Windows has come a long way. but DTP experts say it still has a long way to go. Byte Magazine, 18(10):35–??, September 1993. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1993:WNSb

[Ano93-139] Anonymous. Windows NT supports Posix, but does it matter? Byte Magazine, 18(12): 142–??, November 1993. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1993:WWI

[Ano93-140] Anonymous. WordPerfect 5.2 for Windows, an impressive upgrade. Byte Magazine, 18(2):??, February 1993. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1993:WWW

[Ano93-141] Anonymous. WordPerfect 6.0 for Windows: WordPerfect opts for bigger code and more features. Byte Magazine, 18(13): 46–??, December 1993. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1993:YNP

[Ano93-142] Anonymous. Your next printer might be for your TV. Byte Magazine, 18(13):98–??, December 1993. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1994:PID

[Ano94a] Anonymous. 1-2-3 proves DOS isn’t dead: Lotus’s latest DOS spreadsheet includes many slick features already proven in 1-2-3 for Windows, such as nameable index tabs and SmartIcons. Byte Magazine, 19(7):37– ??, July 1994. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1994:RCA

[Ano94b] Anonymous. 1994 readers’ choice awards: BYTE readers separate the best from the rest. Byte Magazine, 19(6):65–??, June 1994.

CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1994:SDV

[Ano94c] Anonymous. 35mm-Size display has VGA resolution: Thanks to Kopin’s Smart Slide technology, it may not be long before you can wear a head-mounted display that’s about the size of a 35mm slide yet sports 640-by 480-pixel VGA resolution. Byte Magazine, 19(4):32–??, April 1994. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1994:CP

[Ano94d] Anonymous. An 80x86-Compatible PowerPC? Byte Magazine, 19(6): 86–??, June 1994. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1994:AGS

[Ano94e] Anonymous. Adapting GUI software for the blind is no easy task: The widespread adoption of graphical applications adds a whole new set of challenges for applications developers and visually impaired users. Byte Magazine, 19(5):33–??, May 1994. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1994:ASZ

[Ano94f] Anonymous. Add seamless ZIP support to your Windows applications: DynaZIP lets you build Windows-based C/C++ or Visual Basic programs that can read and write standard ZIP files.

Byte Magazine, 19(8):44–??, August 1994. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1994:ALT

[Ano94g] Anonymous. Agents on the loose:

Telescript, General Magic’s communications-oriented programming language, lets developers write tools that permit casual users to create intelligent applications that seek out and retrieve important information. Byte Magazine, 19 (2):23–??, February 1994. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1994:ACM

[Ano94h] Anonymous. Alliances could mean better integration: Microsoft and DEC will integrate DEC’s Object Broker distributed-object system with Microsoft’s OLE technology, and the agreement will allow Windows- and OLE-supported Mac applications to seamlessly access data on a wide variety of platforms. Byte Magazine, 19(3):38–??, March 1994. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1994:ADW

[Ano94i] Anonymous. The Am386SC does DOS and Windows. Byte Magazine, 19(1):104–??, January 1994. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1994:AA

[Ano94j] Anonymous. The analog alternative. Byte Magazine, 19(9): 126–??, September 1994. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1994:AOM

[Ano94k] Anonymous. Apple opens the Mac OS: After years of carefully guarding its Macintosh system software, Apple (cupertino, CA) is taking steps to spread the Mac OS to several different platforms. Byte Magazine, 19(4):24– ??, April 1994. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1994:APP

[Ano94l] Anonymous. Apple provides PC on a Mac: Apple’s new card lets you run DOS and Windows applications and cut and paste among your Mac and PC programs. Byte Magazine, 19(1):19– ??, January 1994. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1994:ARM

[Ano94m] Anonymous. Apple redefines the Macintosh: Apple computer is embarking on a two-to three-year project to redefine the Mac’s proprietary hardware/software architecture to accommodate industry standards and eventually merge with IBM’s PReP (PowerPC reference platform). Byte Magazine, 19(11):56–??, November 1994. CODEN BYTEDJ.

ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1994:AMS

[Ano94n] Anonymous. Apple’s and Microsoft’s system software road map: Apple recently divulged new information on plans for its Mac operating system. Byte Magazine, 19(5):40–??, May 1994. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1994:AHE

[Ano94o] Anonymous. The AT&T Hobbit enters its second generation. Byte Magazine, 19(1):105– ??, January 1994. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1994:BCA

[Ano94p] Anonymous. Best of Comdex awards: BYTE editors worked morning, noon, and night picking the best products and technologies at Fall Comdex in Las Vegas. Byte Magazine, 19(2):24–??, February 1994. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1994:BIN

[Ano94q] Anonymous. The best of Interop + Networld: Editors from McGraw-Hill’s TechNet magazines — BYTE, Data Communications, LAN times, and Open Computing — recognized companies that have products that advance global networking. Byte Magazine, 19(7):36– ??, July 1994. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1994:BT

[Ano94r] Anonymous. A better OS/2 than OS/2? Byte Magazine, 19 (1):168–??, January 1994. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1994:BBR

[Ano94s] Anonymous. Bigger and better: Reality is setting in: Small too often translated into flimsy and underpowered products. subnotebook vendors’ renewed efforts will hit the market this summer and fall. Byte Magazine, 19 (9):40–??, September 1994. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1994:BB

[Ano94t] Anonymous. A bigger bus. Byte Magazine, 19(7):134–??, July 1994. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1994:BPHa

[Ano94u] Anonymous. Blasts from the past: Highlights from two decades of covering the computer revolution. Byte Magazine, 19 (10):41–??, October 1994. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1994:BPHb

[Ano94v] Anonymous. Blasts from the past: Highlights from two decades of covering the PC revolution. Byte Magazine, 19(11): 57–??, November 1994. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1994:BRD

[Ano94w] Anonymous. Borland readies dBase, takes aim at Microsoft: It looks like dBase for Windows, Borland’s long-delayed Windows version of the popular database development language, is close to release. based on a look at a prerelease version, the program will offer features that won’t be available in FoxPro for Windows for several months. also covered: the next release of Paradox for Windows. Byte Magazine, 19(7):28– ??, July 1994. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1994:BTH

[Ano94x] Anonymous. BPR tools help you work smarter: Applications for the desktop that let managers implement business process reengineering projects and, in some cases, create work-flow applications. Byte Magazine, 19 (10):24–??, October 1994. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1994:BBB

[Ano94y] Anonymous. Building a better BIOS. Byte Magazine, 19(9): 92–??, September 1994. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1994:BTB

[Ano94z] Anonymous. Businesses turn to BBSes: Amid all the information-highway headlines, business and government agencies are turning to an informationsharing technology that’s existed since the early days of personal computing: the electronic bulletin board system. Byte Magazine, 19(9):32–??, September 1994. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1994:BLB

[Ano94-27] Anonymous. BYTE Lab benchmarks the first P54C systems. Byte Magazine, 19(6):82–??, June 1994. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1994:CPS

[Ano94-28] Anonymous. The C-Phone solution: Out of band. Byte Magazine, 19(11):239–??, November 1994. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1994:CGO

[Ano94-29] Anonymous. CAD gets objective: CAD software vendors are beginning to exploit new technologies such as object-oriented programming and OLE 2.0. Byte Magazine, 19(5):38–??, May 1994. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1994:CPM

[Ano94-30] Anonymous. Can PC E-Mail be the wrong choice? Byte Magazine, 19(5):126–??, May 1994. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1994:CRD

[Ano94-31] Anonymous. CD-ROM drive prices drop: A combination of strong competition and new technology should force prices for CD-ROM readers to drop this fall. Byte Magazine, 19(11): 41–??, November 1994. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1994:CAC

[Ano94-32] Anonymous. Chicago: An ambitious compromise: Windows NT it’s not, but Microsoft’s next version of Windows will play well on low-end, Intel-based PCs. Byte Magazine, 19(3):22– ??, March 1994. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1994:CEB

[Ano94-33] Anonymous. Chicago enters beta testing: Microsoft is working on a new version of Windows that takes several steps toward rivaling IBM’s OS/2. Byte Magazine, 19(1):18–??, January 1994. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1994:CLM

[Ano94-34] Anonymous. Color laser moves toward mainstream: Color laser printers for under US$5000 may be appearing in offices soon. Byte Magazine, 19(9):25–??, September 1994. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1994:CAM

[Ano94-35] Anonymous. Competition for active matrix: The first commercial cold cathode field emission displays, which may compete with active-matrix screens, could show up this year. Byte Magazine, 19(1):24–??, January 1994. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1994:CBH

[Ano94-36] Anonymous. Compiler benchmarks: How useful? Byte Magazine, 19(2):80–??, February 1994. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1994:CNL

[Ano94-37] Anonymous. Connecting networks: Linking up multiple, farflung networks for an enterprise calls for high-bandwidth digital connections. Byte Magazine, 19 (9):98–??, September 1994. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1994:CCP

[Ano94-38] Anonymous. Controlling Cross-Platform objects. Byte Magazine, 19(7):155–??, July 1994. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1994:CAP

[Ano94-39] Anonymous. Converting an application for parallel processing. Byte Magazine, 19(2): 100–??, February 1994. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1994:CAB

[Ano94-40] Anonymous. CorelDraw 5.0 adds better Image-Editing tools: CorelDraw 5.0 offers an improved PhotoPaint image-editing application and numerous new imageediting tools. Byte Magazine, 19(8):48–??, August 1994. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1994:DCM

[Ano94-41] Anonymous. Data compression on the Macintosh. Byte Magazine, 19(2):62–??, February 1994. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1994:DHL

[Ano94-42] Anonymous. Data highway lags in Japan. Byte Magazine, 19 (3):56–??, March 1994. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1994:DLC

[Ano94-43] Anonymous. Data loss: A cautionary tale. Byte Magazine, 19 (2):64–??, February 1994. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1994:DWP

[Ano94-44] Anonymous. DEC and Wang put it all together. Byte Magazine, 19(7):106–??, July 1994. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1994:DAP

[Ano94-45] Anonymous. Developers announce PowerPC Mac applications: Apple has distributed the first software development kits for writing applications for the System 7 operating system on the PowerPC. numerous developers are now announcing support for the Mac PowerPC platform. Byte Magazine, 19(2): 38–??, February 1994. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1994:DC

[Ano94-46] Anonymous. Digital cash. Byte Magazine, 19(10):126–??, October 1994. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1994:DCT

[Ano94-47] Anonymous. Distributed computer telephony. Byte Magazine, 19(7):94–??, July 1994. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1994:DCR

[Ano94-48] Anonymous. Distributed-Computing resource guide. Byte Magazine, 19(6):206–??, June 1994. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1994:DDM

[Ano94-49] Anonymous. Distributed document management with OLE and OpenDoc. Byte Magazine, 19 (8):100–??, August 1994. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1994:DWI

[Ano94-50] Anonymous. Don’t write off the Internet. Byte Magazine, 19 (10):52–??, October 1994. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1994:DPM

[Ano94-51] Anonymous. DSPs and the PC mainstream. Byte Magazine, 19(5):84–??, May 1994. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1994:DS

[Ano94-52] Anonymous. Dual SuperSparcs.

Byte Magazine, 19(9):162–??, September 1994. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1994:MYP

[Ano94-53] Anonymous. E-Mail in your pocket. Byte Magazine, 19 (5):124–??, May 1994. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1994:EMM

[Ano94-54] Anonymous. The engines to make multimedia mainstream: As industry groups seek to standardize software programming interfaces, DSPs may hold the key to bringing voice video, and telephony to low-cost PCs. Byte Magazine, 19(5):22–??, May 1994. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1994:EL

[Ano94-55] Anonymous. The Eriang language. Byte Magazine, 19(8): 184–??, August 1994. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1994:ECM

[Ano94-56] Anonymous. Ethernet cabling methods. Byte Magazine, 19 (4):186–??, April 1994. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1994:ESA

[Ano94-57] Anonymous. Ethernet switching at a fraction of the cost: Even with the benefits of improved network performance, companies hesitate to adopt Ethernet switching, fearing expense and unfamiliarity. A new class of product, a switching hub on a PC card, eliminates both objections. Byte Magazine, 19(10):36– ??, October 1994. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1994:EMD

[Ano94-58] Anonymous. Europe’s many data highways. Byte Magazine, 19 (3):58–??, March 1994. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1994:ES

[Ano94-59] Anonymous. The evolution of a standard. Byte Magazine, 19(5):102–??, May 1994. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1994:FPB

[Ano94-60] Anonymous. Falling prices boost ISDN: The falling prices of ISDN in Europe may spark growth in telecomm and videoteleconferencing applications. Byte Magazine, 19(1):40–??, January 1994. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1994:FGF

[Ano94-61] Anonymous. Fax gets a Face-Lift: This spring and summer, a new wave of products will give fax a face-lift by adding binary-file-transfer capabilities. Byte Magazine, 19 (3):34–??, March 1994. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1994:FPV

[Ano94-62] Anonymous. First-Party vs.

Third-Party call control. Byte Magazine, 19(7):86–??, July 1994. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1994:FEE

[Ano94-63] Anonymous. Front ends ease Internet access: A number of companies are releasing software designed to make accessing the Internet easier. Byte Magazine, 19(5):30–??, May 1994. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1994:GDM

[Ano94-64] Anonymous. Get that data!: Modern business runs on information. new tools and standards help capture and process that information in nontraditional ways. Byte Magazine, 19 (10):98–??, October 1994. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1994:GCR

[Ano94-65] Anonymous. Getting CISC into RISC: The RISC-CISC debate is taking some interesting turns that could have a profound effect on the future of the personal computer industry. Byte Magazine, 19(9):38–??, September 1994. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1994:GLB

[Ano94-66] Anonymous. A giant leap for Borland C++: Borland now has the best C++ environment, but just barely. Byte Magazine, 19 (1):32–??, January 1994. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1994:GPD

[Ano94-67] Anonymous. Government policy on the data highway. Byte Magazine, 19(3):52–??, March 1994. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1994:GGB

[Ano94-68] Anonymous. Graphics gets down to basics: Though not as powerful as full-featured drawing programs, programs like Visio and SnapGrafx make it easy to generate professional graphics for business. Byte Magazine, 19(1):28– ??, January 1994. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1994:HPF

[Ano94-69] Anonymous. Help for patent fever: New products and services are out that can help a developer file a patent application or search for prior patents. Byte Magazine, 19(8):44–??, August 1994. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1994:HCV

[Ano94-70] Anonymous. The high cost of videoconferencing. Byte Magazine, 19(5):82–??, May 1994. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1994:HPD

[Ano94-71] Anonymous. High-Performance 3-D coming to PCs: Applications such as 3-D model animation, stock market visualization, and other high-end programs usually found on expensive workstations should start appearing in 1994 on less expensive 80x86based PCs. Byte Magazine, 19 (2):28–??, February 1994. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1994:HST

[Ano94-72] Anonymous. High-Speed 1394 train still at the station: First products based on the proposed high-speed serial-bus P1394 interface, also known as FireWire, likely won’t begin showing up until later this year. Byte Magazine, 19(7):37–??, July 1994. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1994:HSK

[Ano94-73] Anonymous. Highway safety: The key is encryption. Byte Magazine, 19(3):60–??, March 1994. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1994:HBS

[Ano94-74] Anonymous. How BYTE selected the best. Byte Magazine, 19(1):48–??, January 1994. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1994:HWT

[Ano94-75] Anonymous. How we tested SQL server on NetWare, OS/2, and NT. Byte Magazine, 19(4):150– ??, April 1994. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1994:HBW

[Ano94-76] Anonymous. Hubs branch out of the wiring closet: When combining hubs with devices that give users WAN access, vendors are making sure the components complement each other so that the combination is often better than what you’d get if you bought the pieces separately. Byte Magazine, 19(8):30– ??, August 1994. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1994:IPA

[Ano94-77] Anonymous. IBM plans ambitious network: This fall, IBM will introduce Intelligent Communications, a set of communications services that span the gulf between different access providers, mail systems, delivery media, and user devices. Byte Magazine, 19(8):34–??, August 1994. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1994:IRC

[Ano94-78] Anonymous. Image retrieval for compound documents. Byte Magazine, 19(8):104–??, August 1994. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1994:IFS

[Ano94-79] Anonymous. Inexpensive, fast, and slow to acceptance: When looked at from a technological viewpoint, Windows printers ought to be big sellers. but analysts don’t expect the category to take off for a couple of years. here’s why. Byte Magazine, 19 (9):40–??, September 1994. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1994:ID

[Ano94-80] Anonymous. Inheritance or delegation? Byte Magazine, 19(5): 60–??, May 1994. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1994:INC

[Ano94-81] Anonymous. Intel and Nestor to commercialize Neural-Net chip: An agreement between Intel and Nestor to commercialize a neural processor could affect applications ranging from handwriting recognition to air-traffic control. Byte Magazine, 19(3):32– ??, March 1994. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1994:IPE

[Ano94-82] Anonymous. Intel pushes the 80x86 envelope: In terms of sheer processing power, Intel is behind the RISC curve. its nextgeneration Pentium may only keep the gap from widening. but that may be enough. Byte Magazine, 19(4):22–??, April 1994. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1994:ITS

[Ano94-83] Anonymous. Intel, TI show off the 486DX4. Byte Magazine, 19(4):112–??, April 1994. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1994:ENS

[Ano94-84] Anonymous. Into the enterprise: New software releases by Lotus and announcements by Microsoft and WordPerfect reflect the companies’ different strengths and philosophies. Byte Magazine, 19 (9):24–??, September 1994. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1994:JLM

[Ano94-85] Anonymous. Just like magic:

After more than four years in gestation, a suite of technologies designed by General Magic stands poised to alter today’s mishmash of wired and wireless networks. Byte Magazine, 19 (2):22–??, February 1994. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1994:KBV

[Ano94-86] Anonymous. Kurzweil brings voice dictation to Windows: At least three Windows voicedictation programs will be available by year-end. Kurzweil Applied Intelligence’s Voice for Windows 1.0 is the first. Byte Magazine, 19(8):48–??, August 1994. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1994:LRCd

[Ano94-87] Anonymous. Lab report: 24 CD-ROM drives: Double-Speed rules: We test 24 double-speed and faster CD-ROM drives to identify the price/performance leaders. the best Double-Speed CD-ROM drives; do you need quad speed? how we tested; fast access to multiple CDs; CD-ROM for the road; honorable mentions; dubious achievements. Byte Magazine, 19(9): 176–??, September 1994. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1994:LRE

[Ano94-88] Anonymous. Lab report: 66 Ethernet adapters: We test Ethernet NICs, PCMCIA cards, and parallel-port adapters to find the best networking hardware for desktop and mobile applications. best 16-bit NICs; Hassle-Free NICs; how we tested; best PCMCIA cards; best portable Ethernet; adapters; honorable mention; dubious achievements. Byte Magazine, 19 (6):346–??, June 1994. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1994:LRCa

[Ano94-89] Anonymous. Lab report: 70 color monitors: We evaluate 70 15-to 21-inch color monitors and choose the best for important business applications. best monitors for general business; energy stars burn dimmer; how we tested; quality gauges; the keys to image quality; best monitors for spreadsheets and graphics; emissions overview; is bigger better?; best monitors for complex graphics presentations; Color-Matching monitors; Do-It-Yourself monitor testing; honorable mentions. Byte Magazine, 19(1):202–??, January 1994. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1994:LRCc

[Ano94-90] Anonymous. Lab report: 76 cards for fast graphics: Our applications-based tests identify the best PC and Mac accelerator boards for general-purpose and specialized applications. best for VL-Bus systems — how we tested — best for ISA-Bus systems — Windows NT drivers: Better late then never?–how to buy a graphics accelerator — best for Macintosh NuBus systems — best for EISA-Bus systems — how bus architecture affects graphics — honorable mentions — dubious achievements. Byte Magazine, 19(2): 146–??, February 1994. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1994:LOS

[Ano94-91] Anonymous. LAN Operating-System testing. Byte Magazine, 19(4):164–??, April 1994. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1994:LCPb

[Ano94-92] Anonymous. LANtastic 6.0 creates peer pressure: Peer LANs have evolved from limitedfunction packages to full-featured systems that integrate into enterprise LANs. Byte Magazine, 19(5):26–??, May 1994. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1994:LL

[Ano94-93] Anonymous. Learning to listen. Byte Magazine, 19(5):146– ??, May 1994. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1994:LCPa

[Ano94-94] Anonymous. Letters: Compression problems, the Richard III-MIS director comparison, and Mac SCSI hints. Byte Magazine, 19(4):18–??, April 1994. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1994:LCG

[Ano94-95] Anonymous. Letters: Computing in a global environment, computing and U.S. politics, and a few clarifications. Byte Magazine, 19(9):18–??, September 1994. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1994:LHP

[Ano94-96] Anonymous. Letters: How to preserve a sense of community in neighborhoods while still enjoying on-line communities. Byte Magazine, 19(8):18– ??, August 1994. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1994:LPP

[Ano94-97] Anonymous. Letters: Plug and Play new and old, split routing, Linux sources, OS/2 with Win32, and Ghost Busters. Byte Magazine, 19(11):18–??, November 1994. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1994:LPL

[Ano94-98] Anonymous. Letters: Power-PCs, the legality of E-mail contracts, a data-highway reality check, and Chicago’s effect on software development. Byte Magazine, 19(6):18–??, June 1994. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1994:LRCe

[Ano94-99] Anonymous. Letters: Reactions to Commodore’s demise. Byte Magazine, 19(10):18–??, October 1994. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1994:LRCb

[Ano94-100] Anonymous. Letters: Readers champion their favorite operating systems. Byte Magazine, 19 (2):18–??, February 1994. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1994:LRR

[Ano94-101] Anonymous. Letters: Readers resist hype and respond to coverage of object computing, object databases, and medical databases. Byte Magazine, 19(7): 18–??, July 1994. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1994:LRS

[Ano94-102] Anonymous. Letters: Readers share their views on PDAs, time synchronization, the software bulge battle, and more. Byte Magazine, 19(1):14–??, January 1994. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1994:LDS

[Ano94-103] Anonymous. Letters: The data superhighway, the “real” reasons for technical-support calls, and environmental concern. Byte Magazine, 19(5):18– ??, May 1994. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1994:LWP

[Ano94-104] Anonymous. Letters: WANs, printer technology, and statistics. Byte Magazine, 19(3):18– ??, March 1994. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1994:LA

[Ano94-105] Anonymous. Life at 28.8. Byte Magazine, 19(2):106–??, February 1994. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1994:LCV

[Ano94-106] Anonymous. Low-cost video acceleration arrives: Graphics accelerator cards that deliver improved video playback should arrive in force this fall at prices under US$500. Byte Magazine, 19 (10):34–??, October 1994. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1994:LEP

[Ano94-107] Anonymous. Low-End PowerPC leaves 68040 Macs in the dust: Preliminary test results have the least expensive PowerPC Mac completing computeintensive tasks at least three to four times faster than a Mac Quadra 900. Byte Magazine, 19(3):23–??, March 1994. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1994:LPC

[Ano94-108] Anonymous. Low-Power chip technology. Byte Magazine, 19 (8):68–??, August 1994. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1994:MT

[Ano94-109] Anonymous. Macintosh telephony. Byte Magazine, 19(7):89– ??, July 1994. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1994:MU

[Ano94-110] Anonymous. Made in the USA.

Byte Magazine, 19(10):104–??, October 1994. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1994:MCM

[Ano94-111] Anonymous. The memory-card manager grows up: The committee works on a new PCMCIA standard that will add bus mastering, a 32-bit data path, and 3.3-V operation. Byte Magazine, 19(6):32–??, June 1994. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1994:MOV

[Ano94-112] Anonymous. Motif offers variation on LCD theme: Motif says its active addressing offers the best of both worlds — the low cost of passive matrix and the fast response of active matrix. Byte Magazine, 19(4):32– ??, April 1994. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1994:MEF

[Ano94-113] Anonymous. Motorola’s envoy first to run magic cap: Despite initial disappointing results in the area of PDAs, companies continue to develop hand-held, communications-centric computing devices. Byte Magazine, 19(5):34–??, May 1994. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1994:NAT

[Ano94-114] Anonymous. New Access targets wider audience: Access 2.0 is what Microsoft wanted Access 1.0 to be: an easy-to-use, full-featured, relational database for Windows. Byte Magazine, 19(4):40–??, April 1994. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1994:NI

[Ano94-115] Anonymous. New and improved. Byte Magazine, 19(2): 108–??, February 1994. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1994:NCC

[Ano94-116] Anonymous. The new CPUs: As competition heats up both in the PC and the workstation marketplaces, many of the industry’s leading companies stake their futures on new ground-breaking CPUs. Byte Magazine, 19(11): 92–??, November 1994. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1994:NLR

[Ano94-117] Anonymous. New languages reflect different priorities: With their newest C++ products, Borland and Microsoft have taken fundamentally different focuses: Microsoft emphasizes cross-platform coverage and the ability to create OLE Custom Controls, and Borland introduces a highly modular set of OLE 2.0 tools. Byte Magazine, 19 (11):38–??, November 1994. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1994:NNS

[Ano94-118] Anonymous. NexGen Nx586 straddles the RISC/CISC divide. Byte Magazine, 19(6):76– ??, June 1994. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1994:NUN

[Ano94-119] Anonymous. NIC update: 22 new Ethernet cards: 16-bit NICs from Asia, evaluated by the Taiwan-based independent testing lab LANBit computer. Byte Magazine, 19(4):229–??, April 1994. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1994:ODS

[Ano94-120] Anonymous. The Object Database Standard. Byte Magazine, 19(4):82–??, April 1994. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1994:OW

[Ano94-121] Anonymous. Object wars. Byte Magazine, 19(5):52–??, May 1994. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1994:OC

[Ano94-122] Anonymous. Observing the conventions. Byte Magazine, 19(6): 96–??, June 1994. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1994:ODM

[Ano94-123] Anonymous. OLE 2.0: Death to monoliths? Byte Magazine, 19(3):122–??, March 1994. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1994:OPP

[Ano94-124] Anonymous. Optimizing with pre- and post-compilers. Byte Magazine, 19(2):82–??, February 1994. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1994:OND

[Ano94-125] Anonymous. Options for Notes developers to improve: Developers frustrated by a limited number of visual-programming tools should see a change this summer. Byte Magazine, 19(6):40– ??, June 1994. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1994:GLM

[Ano94-126] Anonymous. OS/2 gets lean and mean: IBM has released the first beta of a new version of OS/2 for Windows: a 32-bit operating system that will run well on 4-MB PCs. but support for APIs for future versions of Windows is uncertain. Byte Magazine, 19 (8):26–??, August 1994. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1994:PSY

[Ano94-127] Anonymous. OS/2 on the PowerPC slated for this year: By year-end, IBM plans to release its first microkernel-based workplace OS product for the PowerPC, which is, in essence, OS/2 running on the PowerPC. Byte Magazine, 19(6):33–??, June 1994. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1994:PC

[Ano94-128] Anonymous. PCSes are coming. Byte Magazine, 19(5):94– ??, May 1994. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1994:PGM

[Ano94-129] Anonymous. The Pentium goes mainstream: Intel has reduced prices on all but it 100-MHz Pentium chips. the result: a new line of affordable Pentium-based PCs, some with enhanced IDE drives, high-speed serial ports, and 64-bit graphics acceleration for US$2500 or less. Byte Magazine, 19(10):25–??, October 1994. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1994:PO

[Ano94-130] Anonymous. Pentium optimizations. Byte Magazine, 19(2): 86–??, February 1994. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1994:PPPb

[Ano94-131] Anonymous. Photoshop and Picture Publisher get a Face-Life: The top vendors of imageprocessing software are unleashing new versions of their highend programs. Byte Magazine, 19 (9):30–??, September 1994. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1994:PPA

[Ano94-132] Anonymous. PIMS are not so personal anymore: Proving that no person is an island, software developers are adding group-collaboration features to PIMs. Byte Magazine, 19(7):40– ??, July 1994. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1994:PH

[Ano94-133] Anonymous. Pipeline hazards.

Byte Magazine, 19(1):87–??, January 1994. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1994:PPD

[Ano94-134] Anonymous. Plug and Play with DMI. Byte Magazine, 19(9): 80–??, September 1994. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1994:PPB

[Ano94-135] Anonymous. Polysilicon: The path to better displays: New fabrication techniques for creating screens based on the polysilicon technology could result in bigger, higher-resolution LCD displays than currently found on laptop computers. Byte Magazine, 19 (11):34–??, November 1994. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1994:PAP

[Ano94-136] Anonymous. Porting Adobe photoshop: A case study. Byte Magazine, 19(2):92–??, February 1994. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1994:PRJ

[Ano94-137] Anonymous. Porting to RISC: Not just a recompile. Byte Magazine, 19(4):142–??, April 1994. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1994:PTL

[Ano94-138] Anonymous. Power2 takes the lead, for now. Byte Magazine, 19(1):77–??, January 1994. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1994:PAW

[Ano94-139] Anonymous. PRML at work. Byte Magazine, 19(3):93–??, March 1994. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1994:PPPa

[Ano94-140] Anonymous. Pumping up the parallel port. Byte Magazine, 19 (8):118–??, August 1994. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1994:PHT

[Ano94-141] Anonymous. Purchasing hints, troubleshooting tips. Byte Magazine, 19(8):112–??, August 1994. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1994:QHO

[Ano94-142] Anonymous. Quality: The hidden offshore advantage. Byte Magazine, 19(6):114–??, June 1994. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1994:RCG

[Ano94-143] Anonymous. Remote control gets redirected: Several programs offer an inexpensive, though less capable, alternative to dedicated hardware/software solutions for remote LAN access. Byte Magazine, 19(1):40–??, January 1994. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1994:SVI

[Ano94-144] Anonymous. SCSI vs. IDE. Byte Magazine, 19(3):99–??, March 1994. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1994:SSC

[Ano94-145] Anonymous. Simon says: Communicate: IBM’s new personal communication device combines a cellular phone, fax, E-mail, paging, and personal productivity applications. Byte Magazine, 19(2):28–??, February 1994. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1994:SR

[Ano94-146] Anonymous. Software in Russia. Byte Magazine, 19(6):118– ??, June 1994. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1994:SRN

[Ano94-147] Anonymous. Software roundup:

Network management systems:

BYTE evaluates software tools for managing complex LANs. the utilities range from desktop administration tools to highend products with enterprisewide device management. Byte Magazine, 19(7):121–122, 124, 126–127, 130, 132, July 1994. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1994:SRR

[Ano94-148] Anonymous. Software roundup: Remote-Control Windows: With remote-control software, you can access all the resources of your desktop computer system from just about anywhere. BYTE evaluates the six remote-control programs most widely used for running Windows applications. we test the programs for performance, features, usability and versatility. Byte Magazine, 19(9): 137–??, September 1994. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1994:SRV

[Ano94-149] Anonymous. Software roundup: Virus-prevention NLMs: As the computing world becomes increasingly interconnected through LANs, wide-area links, the Internet, and on-line services, corporations are more vulnerable to the threat of computer viruses. BYTE evaluates a convenient and effective solution: antivirus software that works as NetWare NLMs. we test seven products for performance, effectiveness, usability, and versatility. Byte Magazine, 19(8):129–130, 132– 134, 136, August 1994. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1994:SIT

[Ano94-150] Anonymous. Solutions to Internet traffic jams: The number of bytes transferred on the Internet has doubled. many Internet observers are busy working to prevent network overloads in the future. Byte Magazine, 19 (11):27–??, November 1994. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1994:SSL

[Ano94-151] Anonymous. Speaking the same language. Byte Magazine, 19 (5):118–??, May 1994. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1994:SCI

[Ano94-152] Anonymous. Speedy CDs improve video performance: Toshiba and others are developing new CD-ROM players to improve the performance of video playback on multimedia computers. Byte Magazine, 19(1):36–??, January 1994. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1994:SSN

[Ano94-153] Anonymous. SQL server for NT on CISC and RISC. Byte Magazine, 19(4):158–??, April 1994. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1994:SEA

[Ano94-154] Anonymous. Standards efforts aim to ease interoperability. Byte Magazine, 19(8):92–??, August 1994. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1994:SHS

[Ano94-155] Anonymous. Storage hot spots. Byte Magazine, 19(3):86–??, March 1994. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1994:SS

[Ano94-156] Anonymous. Subtyping or subclassing? Byte Magazine, 19(5): 58–??, May 1994. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1994:SWV

[Ano94-157] Anonymous. SunSelect’s Wabi vs. Insignia Solutions’ SoftWindows. Byte Magazine, 19(1):156– ??, January 1994. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1994:SHG

[Ano94-158] Anonymous. Switching hub get their due: Somewhat overshadowed by ATM products, Ethernet switching hubs are now getting the respect they deserve. Byte Magazine, 19(6):28– ??, June 1994. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1994:TVW

[Ano94-159] Anonymous. Taiwanese vendors wait for operating systems: They’re ready with the hardware, but major Taiwanese PC manufacturers continue to wait for a wider selection of operating systems than just AIX to load on the PowerPC clones. Byte Magazine, 19(10):30–??, October 1994. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1994:TPS

[Ano94-160] Anonymous. The taming power of the small. Byte Magazine, 19 (9):68–??, September 1994. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1994:TS

[Ano94-161] Anonymous. Telescript security. Byte Magazine, 19(10):64– ??, October 1994. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1994:TNW

[Ano94-162] Anonymous. Teutonizing the Newton: In working on the German version of the Newton PDA, Apple discovered that handwriting-recognition systems add a whole new set of factors in localizing a product. Byte Magazine, 19(3):26–??, March 1994. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1994:TPP

[Ano94-163] Anonymous. Tips for Plug and Play. Byte Magazine, 19(9): 82–??, September 1994. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1994:TNT

[Ano94-164] Anonymous. The tools for new TV. Byte Magazine, 19(3):62– ??, March 1994. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1994:TMT

[Ano94-165] Anonymous. Tut’s 10-Mbps telephone cord network could be Small-Office king: A new development brings 10-Mbps networking to companies that are put off by the complexities of network wiring. Byte Magazine, 19(3):28– ??, March 1994. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1994:TSA

[Ano94-166] Anonymous. TV services add value to desktop PCs: Cable TV may soon provide a lot more than clear reception of the Simpsons. companies are testing custom news services, on-line access, local discussion forums, and other services that will be delivered to your PC or a set-top box. Byte Magazine, 19(10):28– ??, October 1994. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1994:TPM

[Ano94-167] Anonymous. Types of parallel machines. Byte Magazine, 19 (2):98–??, February 1994. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1994:UGU

[Ano94-168] Anonymous. Users: Give us connected PDAs with PCM-CIA: BYTE interviewed and surveyed about 30 current or former PDA users and came up with a 10-point wish list for nextgeneration PDAs. Byte Magazine, 19(6):34–??, June 1994. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1994:VWC

[Ano94-169] Anonymous. Vendors work to cure incompatibility blues: Backup software and hardware vendors hope that a new standard, the System Independent Data Format, will make incompatibility problems obsolete. Byte Magazine, 19(6):32– ??, June 1994. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1994:VYL

[Ano94-170] Anonymous. Verifying your LAN cables. Byte Magazine, 19(4):188–??, April 1994. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1994:VAF

[Ano94-171] Anonymous. Video acceleration in the fast lane: Two coalitions, VESA and an Intel-ATI pairing, have offered new graphics standards to speed digital video across the desktop. Byte Magazine, 19(4):28–??, April 1994. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1994:WMB

[Ano94-172] Anonymous. The Web means business: The Internet’s popularity is booming, and it’s being partly driven by the World Wide Web’s red-hot popularity. Byte Magazine, 19(11):26–??, November 1994. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1994:WTS

[Ano94-173] Anonymous. What they’re saying about Plug and Play. Byte Magazine, 19(9):86–??, September 1994. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1994:WNC

[Ano94-174] Anonymous. What’s new: A cellular device transmits data; an applications development environment lets you create crossplatform workgroups; virtual reality becomes interactive; and more. Byte Magazine, 19(2): 204–??, February 1994. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1994:WNDa

[Ano94-175] Anonymous. What’s new: A desktop unit that faxes, prints, copies, and scans; a wireless device that coexists with your mouse; software that learns from experience; and more. Byte Magazine, 19(1):254–??, January 1994. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1994:WNPa

[Ano94-176] Anonymous. What’s new: A PCMCIA card with its own audio; Windows software for powerful imaging; Visual Basic tools; and more. Byte Magazine, 19 (3):220–??, March 1994. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1994:WNSb

[Ano94-177] Anonymous. What’s new: A stackable Ethernet switch; portable wireless LAN adapters; imaging in Windows NT; objectbased programming; and more. Byte Magazine, 19(7):212–??, July 1994. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1994:WNDb

[Ano94-178] Anonymous. What’s new:

DECtalk Express Speech Synthesizer goes where you go; Visual Thought communicates ideas graphically; and more. Byte Magazine, 19(10):224–??, October 1994. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1994:WNI

[Ano94-179] Anonymous. What’s new: InterActive Communicator talks back, HotDocs automates document creation, and more. Byte Magazine, 19(6):288–??, June 1994. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1994:WNPb

[Ano94-180] Anonymous. What’s new: PCM-CIA connects notebooks to Ethernet, Macs become multimediaready, open architecture brings visual development to C++, and more. Byte Magazine, 19(9): 220–??, September 1994. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1994:WNR

[Ano94-181] Anonymous. What’s new: RediDockit and DigiDial provide wireless control for your PC; Rapport Script brings objectoriented word processing to Unix; Internet-In-A-Box lets you access the Internet from your PC; and more. Byte Magazine, 19(4):252–??, April 1994. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1994:WNSc

[Ano94-182] Anonymous. What’s new: Scan color graphics in a single pass; converse and hold interactive conferences from a distance; turn workgroup objects into networked discussion forums; and more. Byte Magazine, 19(11):

308–??, November 1994. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1994:WNSa

[Ano94-183] Anonymous. What’s new: Sparcbook 3 ships with the Solaris OS; face to face provides real-time cross-platform document conferencing for PCs and macs; and more. Byte Magazine, 19(5):20– ??, May 1994. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1994:WNQ

[Ano94-184] Anonymous. What’s new: The Quartet provides four PCI ports; Enable for Windows integrates five applications; and more. Byte Magazine, 19(8):202–??, August 1994. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1994:WM

[Ano94-185] Anonymous. What’s on the market. Byte Magazine, 19(9): 114–??, September 1994. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1994:WS

[Ano94-186] Anonymous. Whiteboarding with software. Byte Magazine, 19(4):178–??, April 1994. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1994:WAS

[Ano94-187] Anonymous. Whose authentication systems? Byte Magazine, 19 (10):128–??, October 1994. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1994:WNW

[Ano94-188] Anonymous. Windows NT and Workplace OS: Plug it in. Byte Magazine, 19(1):166–??, January 1994. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1994:WEP

[Ano94-189] Anonymous. WordPerfect enters paperless-document arena: With WordPerfect Envoy, WordPerfect joins Adobe, Farallon Computing, and No Hands Software in offering portable electronicdocument solutions. Byte Magazine, 19(7):32–??, July 1994. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1994:WB

[Ano94-190] Anonymous. A world of benchmarks. Byte Magazine, 19(10): 68–??, October 1994. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Apiki:1990:CL

[Api90a] S. Apiki. A colourful luggable. Byte Magazine, 15(13):246, December 1990. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Apiki:1990:NCT

[Api90b] S. Apiki. New CAD test shuffles 34010 pack: New graphicsboard favorites emerge in a product focus update. Byte Magazine, 15(5):182–183, May 1990. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Apiki:1990:SLC

[Api90c]  S. Apiki.  Small, Low-Cost

UPSes: Small and inexpensive backup power systems make reliable power an individual choice. Byte Magazine, 15(12):258, 260, November 1990. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Apiki:1991:FCC

[Api91a] S. Apiki. Full color comes to LCDs: New portables from Toshiba and Dolch put full-color VGA on the road. Byte Magazine, 16(8):245–246, 248, August 1991. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Apiki:1991:JFL

[Api91b] S. Apiki. Journey to faraway LANs: Nine ways to connect remote sites to your local network. Byte Magazine, 16(7): 194–196, 198, 200, 202–203, 205– 206, 208, 210, July 1991. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Apiki:1991:VAR

[Api91c] S. Apiki. V-ATE revs up PC diagnostics: Vista Microsystems offers more than standard poweron self-test cards. Byte Magazine, 16(3):281–282, March 1991. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Apiki:1991:LDC

[Api91d] Steve Apiki. Lossless data compression: An explanation of two compression algorithms. Byte Magazine, 16(3):309–312, 314, 386–387, March 1991. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Apiki:1992:BOW

[Api92a] Steve Apiki. Bringing the outside into Windows: Outside in for Windows brings DOS documents into Windows. Byte Magazine, 17(9):281–??, September 1992. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Apiki:1992:CFP

[Api92b] Steve Apiki. Clocking the fastest PCs on the planet: Performance comparison of 66-MHz DX2 systems from Compaq, Dell, ALR, and NEC. Byte Magazine, 17(14):205–??, December 1992. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Apiki:1993:CF

[Api93a] Steve Apiki. Compressing with fractals. Byte Magazine, 18 (5):95–??, Spring 1993. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Apiki:1993:LNB

[Api93b] Steve Apiki. The littlest notebooks: BYTE tests three nextgeneration subnotebook machines, including CompUSA’s 4SL/25 subnote, HP’s Omni-Book 300, and the Zenith Data Systems Z-Lite 320L. Byte Magazine, 18(10):137–138, 140, 142, September 1993. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Apiki:1993:NGC

[Api93c] Steve Apiki. Next-generation code generators for Windows: Latest versions of case:W and WindowsMaker professional ease the task of generating Windows code. Byte Magazine, 18(1):225– ??, January 1993. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Apiki:1994:BSS

[Api94a] Steve Apiki. Big-Screen stars:

Capable 21-inch monitors from Nanao and Nokia push resolution to a flicker-free 1600 by 1200 pixels. both can support an 80-HZ refresh rate if you’ve got the right graphics card. Byte Magazine, 19(5):157–??, May 1994. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Apiki:1994:PPI

[Api94b] Steve Apiki. Paths to platform independence: With multiplatform toolkits, you can build applications for Windows, the Mac, X/Motif, OS/2 presentation manager, and a variety of other platforms — from a single set of sources. Apiki develops an application with Liant Software’s C++/Views, WNDX, XVT Software’s XVT, and Zinc’s Application Frameworks and evaluates each product for its programming environment and for its portability across multiple operating systems. Byte Magazine, 19 (1):172–??, January 1994. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Apiki:1994:QKO

[Api94c] Steve Apiki. Q+E: The key to ODBC: With ODBC drivers and Q+E’s 2.0 release of database library, you can build a database-independent applications. the drivers enable a transparent connection to multiple data sources, while the library of DLLs delivers features — such as transactions support and explicit record locking — not necessarily supported by the database engine. Byte Magazine, 19(3):139– ??, March 1994. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Apiki:1994:SWG

[Api94d] Steve Apiki. SPARC workstations to go: SPARC portables from RDI, Sun, and Tadpole put workstation computing on the road. these systems have at least a 50-MHz MicroSparc CPU, 32 MB of RAM, 340 MB of internal SCSI storage, and a color TFT display. all include software to handle such mobile problems as rapidly reconfiguring between different network situations. Apiki tests for performance, features, and portability. Byte Magazine, 19(8):153– ??, August 1994. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Apiki:1994:WRE

[Api94e]  Steve Apiki.  Windows on

RISC: Emulating Windows on even the fastest RISC workstation might mean disappointing performance. Byte Magazine, 19(4):109–??, April 1994. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Appleby:1991:CLP

[App91] D. Appleby. Classic language, part 1: FORTRAN: First in a series on language survivors. Byte Magazine, 16(9):147–148, 150, September 1991. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Andrews:1991:PPY

[AR91] D. Andrews and A. Reinhardt. A PC and 1-2-3 in the palm of your hand (HP 95LX palmtop computer). Byte Magazine, 16(5):44– 46, May 1991. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anderson:1992:OSC

[ASS92] Eric C. Anderson, Stephen Shepard, and Phil Sohn. Overview: Signal computing: Watch for hazards when moving information from the analog to the digital realms. Byte Magazine, 17(12):154–??, November 1992. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Apiki:1992:LEC

[AT92] Steve Apiki and Tom Thompson. Less expensive, or cheap? Byte Magazine, 17(12):??, November 1992. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Apiki:1992:NMD

[ATU92] Steve Apiki, Tom Thompson, and Jon Udell. Network modems dial in, dial out, and route packets: Microtest’s Lanmodem versus shiva’s NetModem/E. Byte Magazine, 17(12):269–??, November 1992. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Apiki:1990:PFM

[AWGY90] S. Apiki, S. Wszola, R. Grehan, and T. Yager. Product focus: Massive storage for multiple platforms: The BYTE Lab puts 15 high-capacity hard disk drives to the test across four operating systems: DOS, Unix, NetWare 386, and the Mac OS. Byte Magazine, 15(12):172, 174– 178, 180, 182, 184, 186, 188, November 1990. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Bacon:1994:CAC

[Bac94] David F. Bacon. Cache advantage: CPUs get the glory, but cache type and organization are just as critical in determining system performance. Byte Magazine, 19(8):78–??, August 1994. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Banks:1990:CU

[Ban90] M. Banks. Computers in the USSR. Byte Magazine, 15 (12):72IS–11–12, 14, 16, November 1990. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Banks:1991:LSD

[Ban91] Michael A. Banks. Are On-Line services delivering? Byte Magazine, 16(11):123–??, ?? 1991. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Banker:1992:RSP

[Ban92] Stephen Banker. Report from Sao Paulo: Over 450,000 people jam Brazil’s Fenasoft software show. Byte Magazine, 17 (12):45–??, November 1992. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Baran:1990:ASF

[Bar90a] N. Baran. Apple’s special fx (Mac IIfx). Byte Magazine, 15 (4):111–114, April 1990. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Baran:1990:PSG

[Bar90b] N. Baran. Putting the squeeze on graphics: Compression technologies for full-color graphics and full-motion video. Byte Magazine, 15(13):289–290, 292–294, December 1990. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Baran:1990:SLC

[Bar90c] N. Baran. Sun’s Low-Cost RISC: A powerful diskless version of Sun’s SPARCstation. Byte Magazine, 15(6):136–137, June 1990. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Baran:1990:IN

[Bar90d] Nick Baran. IBM in the nineties. Byte Magazine, 15(11): 62–??, Fall 1990. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Barker:1990:AG

[Bar90e] C. Barker. Amstrad’s generation 3. Byte Magazine, 15(12):72IS– 21–22, November 1990. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Barker:1990:FMD

[Bar90f] C. Barker. A fax machine for disks. Byte Magazine, 15(13): 72IS/11–12, December 1990. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Barker:1990:PCE

[Bar90g] C. Barker. Personal computing in Eastern Europe: Behind the crumbled Iron Curtain lie lands of high-technology disarray — and opportunity. Byte Magazine, 15(9):401–402, 404, 407– 408, 410, September 1990. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Barker:1990:CC

[Bar90h] Colin Barker. Continental computing. Byte Magazine, 15(11): 133–??, Fall 1990. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Barron:1990:CNB

[Bar90i] J. J. Barron. Chips for the nineties and beyond: New chips may make for higherperformance and unconventional ways of computing. Byte Magazine, 15(12):342–346, 348–350, November 1990. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Barron:1990:CHT

[Bar90j] J. J. Barron. Consortia: High-Tech co-ops: Computer and chip consortia are working hard to give the U.S. an edge over foreign technology. Byte Magazine, 15(6):269–270, 272, 274, 276, June 1990. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Barron:1990:WCS

[Bar90k] J. J. Barron. Want to catch some Z’s: Zero-slot LANs are a proven way to provide file transfer and peripheral-and file-sharing capabilities. Byte Magazine, 15(6): 217–218, 220–222, June 1990. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Baran:1991:DAP

[Bar91a] N. Baran. Data acquisition: PCs on the bench: Data acquisition on powerful personal computers and workstations. Byte Magazine, 16(5):145–149, May 1991. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Baran:1991:LBN

[Bar91b] N. Baran. LCDs and beyond: Nick Baran reviews the state of the art of display technologies and describes some fascinating alternatives. Byte Magazine, 16 (2):229–234, 236, February 1991. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Baran:1991:OSN

[Bar91c] Nicholas Baran. Operating systems now and beyond. Byte Magazine, 16(11):93–??, ?? 1991. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Baran:1991:W

[Bar91d] Nicholas Baran. This way to 1992. Byte Magazine, 16(11):16– ??, ?? 1991. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Barker:1991:ESH

[Bar91e] C. Barker. An easy solution for hard copy? (add-in boards for PC color printers). Byte Magazine, 16(2):72IS/21–22, 24, February 1991. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Barlow:1991:LCC

[Bar91f] John Perry Barlow. The law comes to cyberspace: John Perry Barlow, cofounder of the Electronic Frontier Foundation, tells his story. Byte Magazine, 16 (10):332–??, October 1991. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Barron:1991:BWL

[Bar91g] Janet J. Barron. A business wish list. Byte Magazine, 16(11):49– ??, ?? 1991. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Baran:1992:RGF

[Bar92a] N. Baran. Rough gems: First pen systems show promise, lack refinement: The first pen-input systems will need to overcome some technical hurdles. Byte Magazine, 17(4):212–214, 216, 220, 222, April 1992. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Baran:1992:OPC

[Bar92b] Nicholas Baran. The outlook for pen computing: Before it can deliver, pen-based computing has many rivers to cross. Byte Magazine, 17(9):159–162, 164, September 1992. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Barker:1992:ID

[Bar92c] D. Barker. Image doctors. Byte Magazine, 17(11):97–??, November 1992. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Barker:1992:SGU

[Bar92d] D. Barker. Sophisticated graphing under Windows: Delta-Graph professional brings good chart-making tools to Windows. Byte Magazine, 17(12):256–??, November 1992. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Barker:1992:TLP

[Bar92e] D. Barker. Tools for lassoing the paper cyclone. Byte Magazine, 17(9):??, September 1992. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Barker:1992:TTO

[Bar92f] D. Barker. TypeReader takes OCR toward better recognition: ExperVision’s TypeReader OCR software is a step toward better recognition. Byte Magazine, 17(14):213–214, 216, December 1992. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Barker:1992:BWD

[Bar92g] John Barker. Beyond DOS writing a device interface for Windows: The smart approach to writing Windows device drivers. Byte Magazine, 17(7):303–??, July 1992. CODEN BYTEDJ.

ISSN 0360-5280 (print), 1082-7838 (electronic).

Baran:1993:KHP

[Bar93a] Nicholas Baran. Keyboarding!:

How to put your mouse out to pasture. Byte Magazine, 18 (5):65–??, Spring 1993. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Baran:1993:PPT

[Bar93b] Nicholas Baran. Personable PIMs: Ten PIMs for Windows, DOS, and the Mac, ranging from simple organizers to full-blown project managers. Byte Magazine, 18(6):194–198, May 1993. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Barker:1993:RML

[Bar93c] D. Barker. Report from Mexico: Local programmers wait for a break. Byte Magazine, 18(7): 49–??, June 1993. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Barron:1993:PFL

[Bar93d] Janet J. Barron. Putting fuzzy logic into focus: Fuzzy-logic applications arrive on the desktop. Byte Magazine, 18(4):111– ??, April 1993. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Bogoch:1990:SGP

[BBWR90] S. Bogoch, I. Bason, J. Williams, and M. Russell. Supercomputers get personal: The i860based ComputeServer serves up power, not partitions. Byte Magazine, 15(5):231–234, 236–237, May 1990. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Benton:1991:MEL

[Ben91] P. M. Benton. The multilingual edge: A look at systems that translate human languages. Byte Magazine, 16(3):124–128, 130, 132, March 1991. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Berliner:1993:SBL

[Ber93] Hans Berliner. Stop bit: Losing the human edge: Computers are about to mount a new challenge in the chess world. Byte Magazine, 18(6):282–??, May 1993. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Betz:1994:SAR

[Bet94] David Betz. Some assembly required: Life with NewtonScript: Programming the game of life in the Apple Newton’s scripting language. Byte Magazine, 19 (3):191–??, March 1994. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Barker:1992:BLPb

[BEW92] D. Barker, D. L. Edwards, and S. Wszola. BYTE Lab product report: Writing in style: WYSI-WYG word processors give your text visual appeal. Byte Magazine, 17(6):306–315, June 1992.

CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Bellin:1990:TCW

[BF90] A. Bellin and P. Del Frate. True color for Windows: Windows 3.0 makes 24-bit color a realistic option. Byte Magazine, 15 (13):281–282, 284, 287, December 1990. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Barker:1992:BLPa

[BGE^{+}92] D. Barker, R. Ga Cote, D. L. Edwards, T. Thompson, and S. Wszola. BYTE Lab product report: Art for business’s sake: One of these packages is sure to make it easier for you to create good-looking pictures, charts, and graphs. Byte Magazine, 17 (4):226–254, April 1992. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Bissell:1991:WIC

[Bis91] D. Bissell. What-If CAD: Parametric Math migrates to Windows: Design View and Cedar offer CAD designers what-if analysis tools. Byte Magazine, 16(6): 305–306, 308, June 1991. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Baran:1990:FNS

[BL90] N. Baran and O. Linderholm.

Fast new systems from NeXT: Faster machines with lower prices and the long-sought floppy disk drive have arrived. Byte Magazine, 15(12):165–168, November 1990. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Boggan:1994:UHP

[BL94] Scott Boggan and Michael De Laurentis. Under the hood: The Panose Typeface-Matching system: A look at a font-matching system based on visual characteristics. Byte Magazine, 19(5):187– ??, May 1994. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Bond:1990:PNO

[Bon90a] G. Bond. The personal network: Organize companywide information into personal views with Lotus Notes. Byte Magazine, 15(7): 196, 198, 200, July 1990. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Bond:1990:SPI

[Bon90b] G. Bond. Strictly for personal information: A roundup of seven personal information managers shows that there is a way to get organized. Byte Magazine, 15 (9):196–198, 200, 202, 204–205, September 1990. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Bond:1990:STT

[Bon90c] G. Bond. The SX turns 20: Three 20-MHz SX systems push into 386DX territory. Byte Magazine, 15(10):197–200, October 1990. CODEN BYTEDJ.

ISSN 0360-5280 (print), 1082-7838 (electronic).

Bond:1990:AVY

[Bon90d] George Bond. Adding, value to your data. Byte Magazine, 15 (11):217–218, 220–221, Fall 1990. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Borsook:1993:OSS

[Bor93] Paulina Borsook. Overview:

Seeking security: Mainframetype security is coming to the client/server environment. Byte Magazine, 18(6):118–122, 124, 126, 128, May 1993. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Bosak:1993:NCR

[Bos93] Steve Bosak. The NetWare CD-ROM solution: Microtest’s Discport CD-ROM server makes the marriage of CD-ROM and Novell NetWare easy for all concerned. Byte Magazine, 18(13):151–152, December 1993. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Bryan:1993:PGA

[BR93] John Bryan and Bob Ryan. The power of graphics: Advanced graphics technologies are empowering new classes of applications. C-Cube marries video and RISC. Byte Magazine, 18(12):206–??, November 1993. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Bradley:1990:CIP

[Bra90] D. J. Bradley. The creation of the IBM PC: Design choices that culminated in the machine that conquered the microcomputer world. Byte Magazine, 15(9):414, 16– 18, 420, September 1990. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Brinkmann:1990:DGA

[Bri90] R. M. Brinkmann. 3-D graphics, from Alpha to Z-Buffer: The dedicated memory of graphics workstations speeds up their ability to render in three dimensions. Byte Magazine, 15(7):271– 272, 274, 276–278, July 1990. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Brown:1993:VA

[Bro93] Maxine D. Brown. Visualization applications. Byte Magazine, 18(4):??, April 1993. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Bryan:1992:SDY

[Bry92] John Bryan. Signals on the desktop: You don’t have to wait for DSP technology; it’s here today. Byte Magazine, 17(12): 167–??, November 1992. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Bryan:1993:DPM

[Bry93a] John Bryan. Desktop publishing made easy: Windows desktop publishing is more accessible than ever. Byte Magazine, 18(5):91–93, Spring 1993. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Bryan:1993:PEC

[Bry93b] John Bryan. Pumping up Ethernet: The competition is intensifying in the race to define a 100-Mbps Ethernet standard. Byte Magazine, 18(9):121– ??, August 1993. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Bryan:1994:ITL

[Bry94a] Joe Bryan. IDE takes off: A low-cost solution, enhanced IDE will allow you to connect more and different peripherals to your PC. Byte Magazine, 19(3):97– ??, March 1994. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Bryan:1994:DCS

[Bry94b] John Bryan. Data over cellular: Support builds for AT&T Paradyne’s enhanced throughput cellular. Byte Magazine, 19(9): 59–??, September 1994. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Bryan:1994:FCS

[Bry94c] John Bryan. Fibre channel speeds up: Here’s a new route to the fast lane on the serial superhighway, opening up the capabilities of optical-fiber connections and incorporating other buses and protocols. Byte Magazine, 19(8):123–??, August 1994. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Bryan:1994:LMS

[Bry94d] John Bryan. LANs make the switch: Three types of packetswitching hardware — sharedmemory systems, shared-bus designs, and multistage switch matrix units — make it easier to maintain high-speed connections among an organization’s LANs. Byte Magazine, 19(9): 113–??, September 1994. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Bryan:1994:PPP

[Bry94e] John Bryan. PCMCIA: Past, present, and promise: With Chicago on the horizon and adherence to standards improving, PCMCIA promises peripherals that truly let you plug and play. Byte Magazine, 19(11): 65–??, November 1994. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Badgett:1991:UVD

[BS91] T. Badgett and C. Sandler. “ultra” VGA debuts on the MicroPaq: Monolithic’s MicroPaq 452 Ultra uses the new Edsun chip to make VGA screens shine. Byte Magazine, 16(1):201–202, January 1991. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Buckley:1992:RPQ

[Buc92] William F. Buckley. Reflections on the “privacy” question: William F. Buckley, Jr. asks, who should have access to your electronic medical records? Byte Magazine, 17(6):426–??, June 1992. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Burke:1992:SBT

[Bur92] James Burke. STOP BIT: Technology and the new world order: Advancing technology is creating a new, more complex social order. Byte Magazine, 17(14): 324–??, December 1992. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Bushnell:1992:SBA

[Bus92] Nolan Bushnell. Stop bit: The artichoke theory: What applications can learn from PC games. Byte Magazine, 17(9): 378–??, September 1992. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Buxton:1990:SMI

[Bux90a] B. Buxton. Smoke and mirrors:

In an industry as new as ours, it’s too early to rest on our collective laurels. Byte Magazine, 15(7):205–210, July 1990. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Buxton:1990:SM

[Bux90b] Bill Buxton. Smoke and mirrors. Byte Magazine, 15(??):205– 210, July 1990. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic). (Keywords: user interface).

Cahill:1990:DEE

[Cah90] B. Cahill. Drawing on the 8514/A: An engineer exposes the inner workings of this graphics processor. Byte Magazine, 15 (3):279–280, 282–286, 288–289, March 1990. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Calabrese:1991:DDM

[Cal91a] B. Calabrese. 2-D and 3-D Mac CAD for less cost: DesignCAD’s new Mac program offers fast 2-D and 3-D design. Byte Magazine, 16(3):276–278, March 1991. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Calabrese:1991:PRT

[Cal91b] B. Calabrese. Photo-Realism for those with time (and RAM) to spare: Pixar’s MacRenderMan brings photo-realistic rendering to the Mac. Byte Magazine, 16 (1):223–225, January 1991. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Calem:1993:UPW

[Cal93] Robert E. Calem. Ultraportable PCs: Worth the trade-offs: Subnotebook or palmtop? BYTE looks at the alternatives. Byte Magazine, 18(4):164–166, 168, 170, 172, April 1993. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Carpenter:1991:NHP

[Car91a] S. Carpenter. The NeXTstation: A High-Performance graphical workstation with a PC price tag: Can NeXT take on both Sun and Apple? Byte Magazine, 16(6):297–299, June 1991. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Carr:1991:PPP

[Car91b] R. M. Carr. The point of the pen (PenPoint operating system). Byte Magazine, 16(2): 211–214, 216, 219–221, February 1991. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Carls:1992:ENK

[Car92] Jim Carls. Eight notebooks keep a tight grip on power: A thorough review of 386SL and 386SXL portables. Byte Magazine, 17(9):238–242, 244, September 1992. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Carls:1994:ABB

[Car94] Jim Carls. Access 2.0: The best of both worlds: With new easeof-use features and enhanced development tools, Microsoft Access 2.0 aims at database dabblers and developers. reviewer Jim Carls evaluates Access as both an end-user tool and a development environment. Byte Magazine, 19(9):157–??, September 1994. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Casasent:1992:WYS

[Cas92] David Casasent. Is what you see what you get: Access to optical information processing is now easier. Byte Magazine, 17(9): 189–??, September 1992. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Caudill:1991:ENI

[Cau91] M. Caudill. Expert networks: An innovative combination of technologies lets you build expert systems even when you don’t have an expert. Byte Magazine, 16 (10):108–112, 114, 116, October 1991. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Caudill:1993:NNA

[Cau93] Maureen Caudill. Neural net adds smarts to spreadsheets, slowly: If I only had a brain: Braincel aims to smarten spreadsheets with neural-network technology. Byte Magazine, 18(1): 221–222, 224, January 1993. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Chien:1992:SCY

[Chi92] Philip Chien. Smile for the computer: Your computer might be your camera’s best accessory. Byte Magazine, 17(12):139–??, November 1992. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Chiquoine:1993:DWW

[Chi93] S. Chiquoine. DOS dilemma: Word or WordPerfect? Byte Magazine, 18(11):145–146, October 1993. CODEN BYTEDJ.

ISSN 0360-5280 (print), 1082-7838 (electronic).

Christiansen:1994:SAR

[Chr94] Tom Christiansen. Some assembly required: Developing applications in Perl: This widely supported public domain language can easily solve many general programming problems. Byte Magazine, 19(4):231–??, April 1994. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Claiborne:1990:FGS

[Cla90a]  D. Claiborne.  Faster gets

smaller: Compaq’s speediest compact desktop machine yet. Byte Magazine, 15(8):170–172, August 1990. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Claiborne:1990:FGH

[Cla90b] D. Claiborne. Four 386SXes to go, hold the AC: Low power consumption makes Intel’s 386SX ideal for laptops. here are four models to consider. Byte Magazine, 15(5):161–166, May 1990. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Clarkson:1991:EII

[Cla91a] M. A. Clarkson. An easier interface: An innovative new user interface from Xerox’s Palo Alto Research Center makes use of color and 3-D graphics. Byte Magazine, 16(2):277–278, 280– 282, February 1991. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Clarkson:1991:EI

[Cla91b] Mark A. Clarkson. An easier interface. Byte Magazine, 16(??): 277–282, February 1991. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Clarkson:1992:ITX

[Cla92a] Mark Clarkson. The information theater: Xerox PARC presents a new way to view your data. Byte Magazine, 17(12):145–??, November 1992. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Clarkson:1992:OP

[Cla92b] Mark Clarkson. Objects and Penpoint. Byte Magazine, 17 (14):??, December 1992. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Clarkson:1992:WO

[Cla92c] Mark Clarkson. What’s in an object? Byte Magazine, 17(14):??, December 1992. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Clarkson:1992:SST

[Cla92d] Mark A. Clarkson. SONET: A standard for today. Byte Magazine, 17(9):??, September 1992. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Clarkson:1993:ATN

[Cla93a] Mark Clarkson. All-Terrain networking: ATM can span the network, from the desktop to the wide-area network. is it the answer to a network manager’s dreams? it could be — if all the pieces fall in place. Byte Magazine, 18(9):111–??, August 1993. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Clarkson:1993:MFS

[Cla93b] Mark Clarkson. The many flavors of SQL: Market forces complicate the search for a database access standard. while vendors pitch their versions of SQL, users suffer. Byte Magazine, 18(7):109–??, June 1993. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Clarkson:1993:PCP

[Cla93c] Mark Clarkson. Point-and-Click presentations: Two presentation programs for Windows square off. while Harvard Graphics solidifies its position as a market leader, newcomer WordPerfect Presentations exhibits some serious flaws. Byte Magazine, 18(12): 257–258, November 1993. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Clarkson:1993:HWS

[Cla93d] Mark A. Clarkson. Hitting warp speed for LANs: Collaboraative computing demands faster networks. Byte Magazine, 18 (3):123–??, March 1993. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Clarkson:1993:PGB

[Cla93e] Mark A. Clarkson. Presentations to go: the best presentation graphics package is the one you don’t have to think about. Byte Magazine, 18(5):99– 104, Spring 1993. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Clarkson:1994:PCN

[Cla94a] Mark Clarkson. Process control’s new face: Personal computers are a cost-effective tool for controlling industrial processes. Byte Magazine, 19(10):111–??, October 1994. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Clarkson:1994:SST

[Cla94b] Mark Clarkson. Seriously serial: Two new serial buses contend for desktop acceptance — the lowspeed access. bus and the highspeed P1394/Fire wire. each has special strengths. Byte Magazine, 19(8):117–??, August 1994. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Cook:1990:FCF

[Coo90a] R. Cook. Full circle: Finally, modern applications are as easy to customize as they are powerful. Byte Magazine, 15(8): 211–214, August 1990. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Cook:1990:CD

[Coo90b] Rick Cook. Color for the desktop. Byte Magazine, 15(11):175– 180, Fall 1990. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Cook:1990:DVS

[Coo90c] Rick Cook. Desktop video studio: Is desktop video going to be bigger than desktop publishing? Byte Magazine, 15(2):229–230, 232–234, February 1990. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Cook:1991:ESC

[Coo91] R. Cook. Embedded systems in control: The proliferation of, and new trends in, embedded systems. Byte Magazine, 16(6):153– 154, 156–158, 160, June 1991. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Cosentino:1994:CVL

[Cos94] Victor J. Cosentino. Commentary: Virtual legality: As the law struggles to catch up with technological changes, what happens to our legal, social, and ethical responsibilities? Byte Magazine, 19(3):278–??, March 1994. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Cote:1993:IWM

[Cot93a] Raymond G. A. Cote. Imagining the world: Macintosh software for simulating systems from the administrative office to the factory floor. Byte Magazine, 18 (3):181–??, March 1993. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Cote:1993:SAR

[Cot93b] Raymond G. A. Cote. Some assembly required: Processing magic on the Mac: How to exploit the System 7 Process Manager in your applications. Byte Magazine, 18(4):201–??, April 1993. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Counihan:1992:PSC

[Cou92] Tim Counihan. A platform for signal computing: The signalcomputing environment tries to set a signal-computing standard. Byte Magazine, 17(12):185–??, November 1992. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Covington:1990:SVA

[Cov90] M. A. Covington. Smooth views: Antialiasing lets you get better effective resolution out of VGA displays. Byte Magazine, 15 (5):279–280, 282–283, May 1990. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Cox:1990:TSB

[Cox90] B. J. Cox. There is a silver bullet: The birth of interchangeable, reusable software components will bring software into the information age. Byte Magazine, 15(10):209–210, 212, 214, 216, 218, October 1990. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Campbell:1994:OAC

[CP94] David K. Campbell and Kraig Proehl. Optical advances: A convergence of technologies is set to take the capacity of magnetooptical drives into the stratosphere. Byte Magazine, 19 (3):107–??, March 1994. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Crabb:1990:AMJ

[Cra90a] D. Crabb. AppleShare without a Mac: Jasmine’s DirectServe offers AppleShare file service without sacrificing a Mac. Byte Magazine, 15(3):151–152, 154–155, March 1990. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Crabb:1990:FEC

[Cra90b] D. Crabb. Fast and easy CAD on the mac: Deltasoft’s Origins provides fast competition to AutoCAD on the Mac. Byte Magazine, 15(3):185–186, 188, March 1990. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Crabb:1990:FFP

[Cra90c] D. Crabb. Fast fonts: Post-Script gets turbocharged: Han-Zon’s RISC-based controller turns LaserJet printers into Post-Script hot rods. Byte Magazine, 15(6):184–186, June 1990. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Crabb:1990:VRS

[Cra90d] D. Crabb. Voice recognition for a song: Covox’s Voice Master and Command Corp’s bug. Byte Magazine, 15(8):174, 176, 178, August 1990. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Crabb:1990:BFM

[Cra90e] Don Crabb. The big four for Mac databases: A survey of the four top relational database development systems. Byte Magazine, 15 (1):129–??, January 1990. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Crabb:1990:CMC

[Cra90f] Don Crabb. CAD: The Mac can do that: With processing power galore and large monitors, Don proves that the Mac can deliver. Byte Magazine, 15(7):91– ??, July 1990. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Crabb:1990:C

[Cra90g] Don Crabb. Crossing over. Byte Magazine, 15(11):41–??, Fall 1990. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Crabb:1990:FCF

[Cra90h] Don Crabb. The fruits of connectivity: A few innovative companies are showing what you can do with systems and software that can share information. Byte Magazine, 15(5):103–104, 106, May 1990. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Crabb:1990:HRM

[Cra90i] Don Crabb. Hit the road, Mac: The Mac Portable’s pluses outweigh its minuses. Byte Magazine, 15(2):167–171, February 1990. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Crabb:1990:IAY

[Cra90j] Don Crabb. Inspiration at the year’s end: A look at what Apple has accomplished over the year, and an inspirational new product. Byte Magazine, 15(13): 105–??, December 1990. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Crabb:1990:ENC

[Cra90k] Don Crabb. Is the end near?

not a chance: The U.S. computer industry is not failing. Byte Magazine, 15(2):117–118, February 1990. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Crabb:1990:MMA

[Cra90l] Don Crabb. A Mac melange:

Apple is suffering from the “not invented here” syndrome. Byte Magazine, 15(3):97–??, March 1990. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Crabb:1990:MED

[Cra90m] Don Crabb. Multimedia for everyone: Descriptions of realistic multimedia work done by students on Mac Pluses and SEs. Byte Magazine, 15(8):91–92, August 1990. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Crabb:1990:PDT

[Cra90n] Don Crabb. The place to be for DTP: Talking to professional desktop publishers reveals surprising facts about desktop publishing on the Macintosh. Byte Magazine, 15(9):87–88, 91, September 1990. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Crabb:1990:RAD

[Cra90o] Don Crabb. Rising from the ashes: Don hauls out his crystal ball and predicts Apple’s near future. Byte Magazine, 15(6):91– ??, June 1990. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Crabb:1990:SGV

[Cra90p] Don Crabb. Science goes visual on the Mac: Get a new view of scientific data with graphics tools from Spyglass. Byte Magazine, 15(10):99–??, October 1990. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Crabb:1990:TSS

[Cra90q] Don Crabb. Two sides of the same coin: A bright side with education, a darker side with software development. Byte Magazine, 15(4):81–??, April 1990. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Crabb:1990:WWM

[Cra90r] Don Crabb. Working with Windows 3.0 and a Mac: Don presents the Mac/Windows 3.0 user’s interoperability survival guide. Byte Magazine, 15(12): 107–108, 110, November 1990. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Crabb:1991:BMD

[Cra91a] Don Crabb. The business Macintosh: Don looks at Mac business simulation. Byte Magazine, 16(3):329–330, 332, March 1991. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Crabb:1991:DSE

[Cra91b] Don Crabb. Dealing with system errors: Don tries to fix his recurring System 7.0 crashes. Byte Magazine, 16(12):381–??, November 1991. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Crabb:1991:MPP

[Cra91c] Don Crabb. The Mac and personal programming: Don shows how the Mac’s oldest true personal programming system gets even better. Byte Magazine, 16 (1):107–108, 110, January 1991.

CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Crabb:1991:MRN

[Cra91d] Don Crabb. Mac realities: A new software concept, and tips for Mac happiness. Byte Magazine, 16(9):347–??, September 1991. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Crabb:1991:MNG

[Cra91e] Don Crabb. Macintosh: The next generation: Don converses with MacFolk to determine new directions for Mac evolution. Byte Magazine, 16(2):101–??, February 1991. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Crabb:1991:MYT

[Cra91f] Don Crabb. Making yourself truly Mac portable: Don’s problems with trying to find the “perfect” portable Mac computing environment. Byte Magazine, 16(13):285–??, December 1991. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Crabb:1991:NFM

[Cra91g] Don Crabb. New frontiers: Mac software finally gets serious about user scripting. Byte Magazine, 16(8):279–??, August 1991. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Crabb:1991:PPP

[Cra91h] Don Crabb. Portable pains and pleasures: When it comes to portable computers, Apple’s second attempt is a winner. Byte Magazine, 16(10):279–??, October 1991. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Crabb:1991:PDG

[Cra91i] Don Crabb. Professional 3-D graphics on the Mac: Don finds a graphics oasis. Byte Magazine, 16(5):334–??, May 1991. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Crabb:1991:RMQ

[Cra91j] Don Crabb. Roger and me: The quest for compatibility: Film critic Roger Ebert has Mac software compatibility problems. Byte Magazine, 16(4):334–??, April 1991. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Crabb:1991:SAD

[Cra91k] Don Crabb. System 7.0–Apple defines its future: The Mac moves into the 1990s with a new operating system. Byte Magazine, 16(6):357–??, June 1991. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Crabb:1991:WMH

[Cra91l] Don Crabb. The worldwide Macintosh: HyperCard makes CD-ROM development easier for small companies. Byte Magazine, 16(7):277–278, 280, 282, July 1991. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Crabb:1992:DMF

[Cra92a] Don Crabb. Designing macs: A firsthand look at the Mac’s new three-dimensional drawing prowess. Byte Magazine, 17(2): 271–??, February 1992. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Crabb:1992:MMU

[Cra92b] Don Crabb. Managing Mac upgrades: Don works up some Mac hardware upgrade strategies. Byte Magazine, 17(3):295– ??, March 1992. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Crichton:1993:CIH

[Cri93] Michael Crichton. Commentary: Installer hell: Software installation programs work in strange and mysterious ways, says the author of Jurassic Park. Byte Magazine, 18(10):291–??, September 1993. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Cronk:1993:UDC

[Cro93a] Randall Cronk. Unlocking data’s content: Tagging languages and compound document architectures code your documents for cross-platform access. Byte Magazine, 18(10):111–??, September 1993. CODEN BYTEDJ.

ISSN 0360-5280 (print), 1082-7838 (electronic).

Cronk:1993:EMY

[Cro93b] Randall D. Cronk. EISes mine your data: Client/server computing revolutionizes executive information systems. EISes can help you find and analyze your data, but it’s not always as easy as it might sound. Byte Magazine, 18(7):121, 123, 125, 127– 128, June 1993. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Crane:1993:PVU

[CRR93] H. D. Crane, D. Rtischev, and A. Rudnicky. Pen and voice unite. Byte Magazine, 18(11): 98–100, October 1993. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Crutchfield:1990:DDS

[Cru90] Robert J. Crutchfield. Data to the desktop: The SQL advantage. Byte Magazine, 15(11): 193–196, 198, Fall 1990. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Cryan:1992:NLN

[Cry92a] Shelley Cryan. A new Laser-Jet, A new standard: Hewlett– Packard’s LaserJet 4M should shake up the printer market. Byte Magazine, 17(14):209–210, December 1992. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Cryan:1992:PBP

[Cry92b] Shelley Cryan. A power boost for PowerPoint: Microsoft breaks new ground. Byte Magazine, 17 (9):267–??, September 1992. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Cryan:1994:MPT

[Cry94] Shelley Cryan. Multimedia presentations: Time-line-based presentation software can enliven a traditional computer-based presentation with motion, animation, video, sound, and interactive controls. cryan evaluates the best multimedia presenters for Windows and the Mac. Byte Magazine, 19(4):189–190, 192– 195, April 1994. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Cote:1992:SFP

[CS92] Raymond G. A. Cote and Ben Smith. Solutions focus: Profiles in document managing: The BYTE Lab looks at four PCbased document management systems. Byte Magazine, 17 (9):198–200, 202, 204, 206–208, 210–212, September 1992. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Cutrona:1991:CCW

[Cut91] Louis J. Cutrona. Class conflict: The Windows/C++ mismatch. Byte Magazine, 16(9): 345–??, September 1991. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Catchings:1990:GPB

[CV90a] B. Catchings and M. L. Van Name. Growing pains: Buying a LAN is kid’s stuff compared to the problems you can encounter trying to expand it and connect it to other networks. Byte Magazine, 15(6):203–204, 206, 208, 210, 214, June 1990. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Catchings:1990:PSF

[CV90b] B. Catchings and M. L. Van Name. Power servers: File servers from Zenith and Storage Dimensions face off. Byte Magazine, 15(7):167–170, July 1990. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Catchings:1990:MNH

[CV90c] Bill Catchings and Mark L.

Van Name. Microsoft’s network heavyweight: A look at the new version of OS/2 LAN manager: smaller, safer, and more secure. Byte Magazine, 15(5):97–98, 100, May 1990. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Catchings:1990:SPH

[CV90d] Bill Catchings and Mark L. Van Name. Serving the Power-Hungry: The age of the super server is upon us. Byte Magazine, 15(3):107–108, 110, March 1990. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Cote:1992:NWR

[CW92] Raymond G. A. Cote and Stanley Wszola. The new wave of removable storage: From MOs to docking bays, the BYTE Lab selects the best removable-storage solutions. Byte Magazine, 17 (10):198–??, October 1992. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Cook:1994:BCR

[CWVS94] Rick Cook, Lamont Wood, Dave Vislosky, and Ben Smith. Books and CD-ROMs: Entertaining math models: Mathematical modeling, CD-ROM databases, computer ethics, and 3-D graphics libraries. Byte Magazine, 19(5):41–??, May 1994. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Daney:1990:RCR

[Dan90] C. Daney. Rexx in charge: Rexx now can control and coordinate all aspects of the OS/2 environment. Byte Magazine, 15(8): 245–246, 248, 250, 252–253, August 1990. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Dauber:1991:FFD

[Dau91] S. M. Dauber. Finding fault: Don’t get caught with your network down. Byte Magazine, 16(3):207–208, 210, 212, 214, March 1991. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Davidovici:1990:RWL

[Dav90] S. Davidovici. On the radio:

A wireless LAN can provide a flexible alternative to its earthbound cousins. Byte Magazine, 15(6):224A–B, 224D, 226, 228, June 1990. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Davis:1994:DDC

[Dav94a] Andrew W. Davis. Desktop data conferencing: A new breed of multifunction DSP-based peripherals makes data conferencing inevitable and lays the groundwork for personal videoconferencing. Byte Magazine, 19(5):81– 84, 86, May 1994. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Davis:1994:WBD

[Dav94b] Andrew W. Davis. Windows becomes DSP-Aware. Byte Magazine, 19(11):151–??, November 1994. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Dawson:1994:DNT

[Daw94a] Bruce Dawson. Dial-up networking: Telebit’s NetBlazer remote routers enable telecommuting. the NetBlazer PN1, a small Ethernet router with a builtin V.32bis modem, supports IP, IPX, and Apple Ethernet protocols. Byte Magazine, 19(6):233– ??, June 1994. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Dawson:1994:PGF

[Daw94b] Bruce Dawson. Plug-and-Go file server: As simple to use as an appliance, the FAServer 400 is a fast, affordable NFS file server for TCP/IP networks. Byte Magazine, 19(6):227–??, June 1994. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Dawson:1994:PHR

[Daw94c] Bruce Dawson. PowerPC hits the road: IBM introduces the first PowerPC-based notebook. the RS/6000 N40 borrows many nice features from IBM’s ThinkPad line, but though strong on performance, it’s weak on battery life. Byte Magazine, 19(7):141– 142, July 1994. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Dawson:1994:PCL

[Daw94d] J. Bruce Dawson. Power of cooperation: Linux is a freely available version of Unix, developed by individuals from all over the world. but it isn’t a barebones Unix clone. it is full of the features you would expect to find only in a commercial implementation of Unix. Byte Magazine, 19(9):167–??, September 1994. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Dayton:1990:DMN

[Day90] D. Dayton. Document management on networks PCs: Imara and ProFound offer two approaches to keeping track of documents. Byte Magazine, 15 (12):251–252, 254–255, November 1990. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Davis:1993:MGM

[DB93] Andrew W. Davis and Joe Burke.

The Mac goes to the movies: A detailed look at Apple’s Quick-Time architecture. Byte Magazine, 18(2):225–??, February 1993. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Diehl:1990:WLP

[DE90] S. Diehl and H. Eglowstein.

When laser printers can’t cut it: A look at 27 alternatives to the popular laser printer. Byte Magazine, 15(13):156–162, 164–166, 168, 168B, D, December 1990. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Diehl:1991:PWP

[DE91a] S. Diehl and H. Eglowstein.

Penny-wise PostScript (printer market survey). Byte Magazine, 16(10):200–206, 208, 210, 212, October 1991. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Diehl:1991:TPT

[DE91b] S. Diehl and H. Eglowstein.

Tame the paper tiger: 14 OCR products designed to help you tame the paper tiger. Byte Magazine, 16(4):220–226, 228, 230, 232, 234, 236, 238, April 1991. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Diehl:1992:BLP

[DE92] S. Diehl and D. L. Edwards. BYTE Lab product report: Scanning the spectrum: The BYTE Lab evaluates 24-bit color scanners for the PC and the Mac. Byte Magazine, 17(7):230–236, 238–240, 242, July 1992. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

DeHaven:1993:SVA

[DeH93] John DeHaven. Stealth virus attacks: Anonymous attack software targets networked computers. Byte Magazine, 18(6):137– 138, 140, 142, May 1993. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Dejesus:1994:LDT

[Dej94]  Edmund Dejesus.  Linking

development teams: With a distributed-objects architecture, a full suite of programming tools, and seamless version control, DEC’s Cohesion WorX delivers a sophisticated environment for software development across heterogeneous networks. Byte Magazine, 19(7):153–??, July 1994. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Delonas:1992:SUS

[Del92] Nicholas Delonas. Spreadsheets under scrutiny. Byte Magazine, 17(11):70–72, 74–76, 78, 80, November 1992. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Delonas:1993:LRL

[Del93a] Nicholas John Delonas. Lotus 1-2-3 release 3.4 for DOS: Lotus’ latest high-end DOS spreadsheet is faster, but you still may prefer the low-end line. Byte Magazine, 18(6):209–??, May 1993. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Delonas:1993:LTA

[Del93b] Nicholas John Delonas. Lotus takes another run at Windows: Windows version of 1-2-3 is a serious challenger to Microsoft Excel. Byte Magazine, 18(7):22– ??, June 1993. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Dern:1992:PII

[Der92] Daniel Dern. Plugging into the Internet: If you’ve ever wondered how you can access the Internet, here’s how. Byte Magazine, 17 (10):149–??, October 1992. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic). See [Vos92].

Deutsch:1991:SYT

[DG91] L. P. Deutsch and A. Goldberg.

Smalltalk: Yesterday, today, and tomorrow: The trial balloon of a decade ago is now flying high. Byte Magazine, 16(8):108–110, 112–115, August 1991. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Duff:1990:MPS

[DH90] C. Duff and B. Howard. Migration patterns: Simulating objectoriented techniques with procedural languages can help you make the transition to tomorrow’s software technology today. Byte Magazine, 15(10):223, 226, 228, 230, 232, October 1990. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Diehl:1992:WFS

[DH92] Stanford Diehl and Dana Hudes. The Windows file shuffle: File conversion under Windows solves the complex problem of divergent graphics formats. Byte Magazine, 17(14):225–??, December 1992. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Diehl:1990:VDI

[Die90a] S. Diehl. Da Vinci does it penlessly: Da Vinci’s speedy Raster-Pro 720 delivers design-quality plots, with color printing on the side. Byte Magazine, 15(6): 178, 180, 182, June 1990. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Diehl:1990:GGL

[Die90b] S. Diehl. G is for graphics: Lotus finally gets graphical with 1-2-3/G. Byte Magazine, 15(9):

182–184, September 1990. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Diehl:1990:WTW

[Die90c]  S. Diehl.  Windows takes

on WingZ: Informix’s graphical spreadsheet puts Windows 3.0 through its paces. Byte Magazine, 15(12):221–222, 224, November 1990. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Diehl:1990:BAG

[Die90d] Stanford Diehl. Benchmarks at a glance: 1990. Byte Magazine, 15(11):75–81, Fall 1990. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Diehl:1990:MAE

[Die90e] Stanford Diehl. Mac adapters embrace Ethernet: Apple, Asante Technologies, and Compatible Systems adapters give Macs an easy entree into swift Ethernet networks. Byte Magazine, 15 (1):203–205, January 1990. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Diehl:1991:PPB

[Die91] S. Diehl. The perfect pitch:

The BYTE Lab picks the best presentation software for Windows, DOS, and the Mac. Byte Magazine, 16(13):206–212, 214, 216, 218, 220, 222, 224, December 1991. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Diehl:1993:AVC

[Die93a] S. Diehl. Acrobat vs. Common Ground (electronic documents distribution). Byte Magazine, 18 (11):133–134, October 1993. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Diehl:1993:CCS

[Die93b] Stanford Diehl. Complete communications for small businesses: A new fax and voice-mail card from Prometheus takes on the upgraded Complete Communicator. Byte Magazine, 18(2): 213–??, February 1993. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Diehl:1993:TMF

[Die93c] Stanford Diehl. Teaching Macs to fetch: Aldus introduces Fetch, a new multiuser, mixedmedia database for the Mac. Byte Magazine, 18(4):183–??, April 1993. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Diehl:1994:DDI

[Die94] Stanford Diehl. Desktop dictation: The IBM Personal Dictation System brings computerbased dictation services to a mainstream corporate audience. with its support of voice input of system commands and its sophisticated dictation application, the system can support text entry in a fully “hands free” environment. its accuracy and ease of use make the system viable for general business correspondence. Byte Magazine, 19(5):145–146, May 1994. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Dion:1994:LCS

[Dio94] Dany Dion. Low-Cost simulation: A circuit simulator lets you try before you fry. Byte Magazine, 19(9):171–??, September 1994. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Dlugosz:1991:QRW

[Dlu91] J. M. Dlugosz. Quick relief for Windows programming: Winpro/3 automates Windows 3.0 programming. Byte Magazine, 16 (3):265–266, March 1991. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Duntemann:1990:NOO

[DM90] J. Duntemann and C. Marinacci.

New objects for old structures: Using object-oriented techniques to convert existing applications has its advantages. Byte Magazine, 15(4):261–266, April 1990. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Donovan:1992:OOS

[Don92]  John Donovan.  Overview:

Operating-system trends: Desktop operating systems deliver mainframe/minicomputer features. Byte Magazine, 17(10): 158–??, October 1992. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Diehl:1990:LPG

[DW90] S. Diehl and S. Wszola. Laser printers get personal (buyer’s guide). Byte Magazine, 15(7): 138–144, 146, 148, 150–152, 154, 156, July 1990. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Dayton:1991:WAS

[DW91] D. Dayton and L. Wood. Windows 3.0 applications: The state of the market. Byte Magazine, 16 (9):153–154, 156, 158, 160, 162, 164, 168, 170, 172, 174, 176, September 1991. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Diehl:1991:RSD

[DWKS91] S. Diehl, S. Wszola, B. Kliewer, and L. Stevens. Rx for safer data. Byte Magazine, 16(8):218– 224, 226, 228, 230, 234–235, August 1991. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Diehl:1991:PPN

[DWN91] S. Diehl, S. Wszola, and M. Nadeau. Perfectly portable: These 15 notebook-size computers mix PC power with state-of-the-art portability. Byte Magazine, 16 (2):148–154, 156–157, 159–160, 162, February 1991. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Dyson:1992:SBA

[Dys92] Esther Dyson. Stop bit. artificial life and natural markets: Parallels between artificial life experiments and competitive markets demonstrate the advantages of a free-market system. Byte Magazine, 17(12): 372–??, November 1992. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Eisenstadt:1990:KETb

[EB90a] M. Eisenstadt and M. Brayshaw.

A knowledge engineering toolkit, part 2: The discussion continues, with a look at back ward and forward chaining. Byte Magazine, 15(12):364–368, 370, November 1990. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Eisenstadt:1990:KETa

[EB90b] M. Eisenstadt and M. Brayshaw.

A knowledge engineering toolkit: The first of a two-part series presenting a knowledge-engineering toolkit for building expert systems. Byte Magazine, 15(10): 268, 270, 272, 274, 278, 282, October 1990. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Ebert:1993:SBC

[Ebe93] Roger Ebert. Stop BIT: Cinema by computer: Movies no longer offer an escape from reality. Byte Magazine, 18(1):334– ??, January 1993. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Eglowstein:1990:BSU

[Egl90a]  H. Eglowstein.  Behind the

scenes: Understanding your programming interface can help you decide which user interface to support in a heterogeneous environment. Byte Magazine, 15 (4):215–216, 218, 220, 222, 224, April 1990. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Eglowstein:1990:HIS

[Egl90b] Howard Eglowstein. Hawk II soars: Club AT’s 25-MHz 80486 may actually be faster than most people need. Byte Magazine, 15 (2):96D–??, February 1990. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Eglowstein:1990:PPY

[Egl90c] Howard Eglowstein. PostScript in the palm of your hand: Pacific Data’s new cartridge gives HP LaserJet II printers easy PostScript compatibility. Byte Magazine, 15(1):197–200, January 1990. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Eglowstein:1990:RTY

[Egl90d] Howard Eglowstein. Reach out and touch your data: Three input devices, ranging from US$100 to US$15,000, let you “hand it to computers.”. Byte Magazine, 15(7):283–286, 288– 290, July 1990. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Eglowstein:1991:CGS

[Egl91a] H. Eglowstein. Can a grammar and style checker improve your writing: The BYTE Lab tests eight popular programs that promise to make you a better writer. Byte Magazine, 16(8): 238–242, August 1991. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Eglowstein:1991:FEN

[Egl91b] H. Eglowstein. Full Ethernet networking without a wire in sight: Motorola’s Altair LAN leaves Ethernet cabling behind. Byte Magazine, 16(7):229–230, July 1991. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Eglowstein:1991:NAW

[Egl91c] H. Eglowstein. A new angle on OS/2 and Windows: Wide Angle makes the virtual desktop a physical reality. Byte Magazine, 16(1):227–??, January 1991. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Eglowstein:1991:NBT

[Egl91d] H. Eglowstein. The next best thing to the paperless office: Eight forms packages help automate your paperwork. Byte Magazine, 16(12):290–292, 294, 298, 300, November 1991. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Eglowstein:1991:WDM

[Egl91e] H. Eglowstein. Windows display managers file rough edges from text: How’s your face? Windows font managers finally make true WYSIWYG a reality. Byte Magazine, 16(6):311–312, June 1991. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Eglowstein:1992:WRB

[Egl92a] H. Eglowstein. Windows on the road: The BYTE Lab tests portable systems and pointing devices with a flair for Windows. Byte Magazine, 17(3):208–213, 216, 218, 220, March 1992. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Eglowstein:1992:CIJ

[Egl92b] Howard Eglowstein. Color inkjet printers take whacks at wax: Hewlett–Packard and Canon release colorful new products. Byte Magazine, 17(10):233–234, October 1992. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Eglowstein:1992:CNN

[Egl92c] Howard Eglowstein. Compaq’s newest notebooks: The colorful LTE Lite/25c and inexpensive Contura make welcome additions to Compaq’s notebook line. Byte Magazine, 17(10):227–228, October 1992. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Eglowstein:1992:NT

[Egl92d] Howard Eglowstein. A new thumper. Byte Magazine, 17 (9):??, September 1992. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Eglowstein:1992:RNR

[Egl92e] Howard Eglowstein. Reviewer’s notebook: A Real-World notebook battery test: Notebook battery-life testing. Byte Magazine, 17(14):232–??, December 1992. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Eglowstein:1993:APP

[Egl93a] Howard Eglowstein. Applying the power of the pen: The promise of pen computers has been dulled by a lack of innovative, pen-centric, generalpurpose applications. here are nine software packages for Go’s PenPoint and Microsoft’s Windows for Pen Computing that challenge the notion that pen systems are only good for vertical markets. Byte Magazine, 18(8):132–??, July 1993. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Eglowstein:1993:BCS

[Egl93b]  Howard Eglowstein.  The

Business-Card shuffle: Microtek’s scan-in-Dex will appeal to any professional who makes numerous business contacts. the scanner reads business cards and drops the data into a searchable Windows database. Byte Magazine, 18(12):267–268, November 1993. CODEN BYTEDJ.

ISSN 0360-5280 (print), 1082-7838 (electronic).

Eglowstein:1993:HSL

[Egl93c] Howard Eglowstein. HP’s simple laser: HP’s new LaserJet 4L brings laser-quality output to the desktop, along with low cost and simple operation. the BYTE Lab examines the speed, quality, and expansion trade-offs of the 4L series. Byte Magazine, 18(7):141– 142, June 1993. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Eglowstein:1993:PNP

[Egl93d] Howard Eglowstein. Photography by the numbers: Professional electronic cameras from Kodak and Sony deliver instant results to your Mac or PC. Byte Magazine, 18(1):241– ??, January 1993. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Eglowstein:1993:PSC

[Egl93e] Howard Eglowstein. Pioneer’s super CD-ROM drive: Pioneer’s new minichanger can access six CDs at quadruple speed. Byte Magazine, 18(4):193–??, April 1993. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Eglowstein:1993:TBG

[Egl93f]  Howard Eglowstein.  Tape

backup on the go: A review of four parallel-port tape drives for backup. Byte Magazine, 18(6):205–??, May 1993. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Eglowstein:1994:AGB

[Egl94a] H. Eglowstein. Almost as good as being there (conferencing enhancements). Byte Magazine, 19(4):173–174, 176, April 1994. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Eglowstein:1994:ACD

[Egl94b] Howard Eglowstein. Advancing communications: Data communications programs with scripting languages can automate your on-line sessions. testing editor Howard Eglowstein looks at seven popular communications packages from four different environments, testing them with new 28.8-Kbps moderns. Byte Magazine, 19(2):104–??, February 1994. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Eglowstein:1994:DRO

[Egl94c] Howard Eglowstein. Due recognition of OCR: We compared the new high-end Windows products of major OCR vendors Caere and Calera with each other and with the international edition of less-know Recognita’s product. we also tested Xerox’s low-cost TextBridge 2.0 against all three. for serious OCR work, accuracy is critical, but if you don’t do high-volume OCR, TextBridge may be your best bet. Byte Magazine, 19(10):145–148, October 1994. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Eglowstein:1994:NBL

[Egl94d] Howard Eglowstein. Networking on a beam of light: Photonics’ Cooperative infrared LAN connects a roomful of Macs wirelessly. it’s simple to use: You plug the small transceiver unit into a system’s LocalTalk port. Byte Magazine, 19(10):143–144, October 1994. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Eglowstein:1994:OWO

[Egl94e] Howard Eglowstein. One world, one fax: Global Village’s One World fax server gives a Mac network an easy shared fax solution. the One World fax server installed on the BYTE editorial LAN is evaluated for its performance, flexibility, and ease of use. Byte Magazine, 19(10): 169–170, October 1994. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Eglowstein:1994:PFB

[Egl94f] Howard Eglowstein. PhotoStyler fights back: Since Aldus’s last release of PhotoStyler, the imageediting market has heated up with the introduction of Photoshop on the Windows platform. PhotoStyler 2.0 represents more than a simple face-lift. Aldus pumped up the performance and the features set and comes away with a serious competitor. Byte Magazine, 19(2):137–??, February 1994. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Eglowstein:1994:FAL

[Egl94g] Ira Eglowstein. Friendly acquisition: The latest Windows data acquisition software lets you build applications without writing a line of code. graphical programming saves time and money, allows more flexibility, reduces programming errors, and empowers nonprogrammers, Eglowstein evaluates seven current Windows tools. Byte Magazine, 19(7):147–??, July 1994. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Eglowstein:1993:SFM

[ES93] Howard Eglowstein and Ben Smith. Solutions focus: Mixed messaging: Multiplatform internetwork mail can link diverse clients on widespread networks. Byte Magazine, 18 (3):136–??, March 1993. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Eglowstein:1994:MAL

[ES94] Howard Eglowstein and Ben Smith. E-Mail from afar: Lotus cc:Mail and Microsoft Mail provide the necessary E-mail tools for communicating with offices in far-flung locations. we liked both user interfaces, but cc:Mail is easier on the administrator. Byte Magazine, 19(5):122–124, 126, 128, 130, May 1994. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Eglowstein:1994:BCR

[ESN94] Howard Eglowstein, Ben Smith, and Michael Nadeau. Books and CD-ROMs: Low-Cost data acquisition: All you want to know about data acquisition, an improved Encarta, open systems, and more. Byte Magazine, 19 (1):41–??, January 1994. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Essex:1994:BFI

[Ess94] David Essex. Big, fast IDE drives: New IDE drives provide performance suitable for localbus systems and capacities that exceed a gigabyte. we test eight of the latest. Byte Magazine, 19 (9):151–??, September 1994. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Eglowstein:1991:PMP

[ET91] H. Eglowstein and T. Thompson. Please, Mister Postman: The BYTE Lab tests nine Email packages for Mac and PC LANs. Byte Magazine, 16(3): 222–225, 227–235, March 1991. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Ewart:1993:HTD

[ET93] Norton Ewart and Larry Thayer. HP takes a dual approach. Byte Magazine, 18(12):216–??, November 1993. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Eugenides:1990:ESM

[Eug90] J. G. Eugenides. Easier strings for the Mac. Get a handle on Mac strings with this C++ class. Byte Magazine, 15(13):349–350, 352, December 1990. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Evans:1990:ESH

[Eva90] Ron Evans. Expert systems and HyperCard: HyperCard can be ideal for creating knowledgebased systems. Byte Magazine, 15(1):317–324, January 1990. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Evangelou:1994:CLS

[Eva94] Steve Evangelou. Commentary:

Layoff software: Let’s hope this doesn’t become a new growth category. Byte Magazine, 19(2): 262–??, February 1994. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Eglowstein:1992:BLP

[EW92] Howard Eglowstein and Stan Wszola. BYTE Lab product report. 486 systems for a graphical world: The best 33-MHz 486 systems for today’s demanding applications. Byte Magazine, 17(12):209–210, 212, 214, 218– 221, 224, 226, 230, 234, 236, November 1992. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Eglowstein:1990:WPB

[EWT90] H. Eglowstein, S. Wszola, and T. Thompson. Word processors that build character: The BYTE Lab evaluates 15 WYSI-WYG word processors for the Mac and the PC. Byte Magazine, 15(9):132, 134–136, 138, 140, 143–146, 148, 150, 152, September 1990. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Farris:1990:UNY

[Far90] R. Farris. Unix and 1-2-3: Now you can run Lotus 1-2-3 under Unix. Byte Magazine, 15(13): 233–??, December 1990. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Fried:1993:SCS

[FH93] Benjamin Fried and Othar Hansson. Sun’s C solution for Solaris: Sparcworks Professional C is a solid compiler with a few good tools. Byte Magazine, 18(4):189– ??, April 1993. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Fichtelman:1990:DWU

[Fic90] Mike Fichtelman. Don’t worry, use HLLAPI (programming tool). Byte Magazine, 15(11): 207–208, 210–212, 214, 216, Fall 1990. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Fiderio:1992:BGT

[Fid92] Jan Fiderio. BYTE’s guide to 128 top Windows products. Byte Magazine, 17(11):128–??, November 1992. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Fiedler:1990:ASG

[Fie90a] David Fiedler. Answers to some good questions: Our columnist answers the most commonly asked questions, including “which Unix for you?”. Byte Magazine, 15(1):115–??, January 1990. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Fiedler:1990:BWI

[Fie90b] David Fiedler. Back to the workstations II: Unix workstations and personal computers completely merge. Byte Magazine, 15(13):119–120, 122, December 1990. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Fiedler:1990:FSH

[Fie90c] David Fiedler. The free software hit parade: A quick review of the most popular free Unix software. Byte Magazine, 15(8): 85–86, 88, August 1990. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Fiedler:1990:FSW

[Fie90d] David Fiedler. Free software!:

When it comes to user-developed Unix programs, there is such a thing as a free lunch. Byte Magazine, 15(6):97, 100, June 1990. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Fiedler:1990:FHL

[Fie90e]  David Fiedler.  Future his-

tory: Looking at business software from the last 15 years and the next 15. Byte Magazine, 15 (9):81–??, September 1990. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Fiedler:1990:GUR

[Fie90f] David Fiedler. Getting UUCP running, and other stories: Our columnist details how to set up UUCP communications. Byte Magazine, 15(4):71–??, April 1990. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Fiedler:1990:GAM

[Fie90g] David Fiedler. Go ahead, make my day: Tips on installing freely available Unix software. Byte Magazine, 15(7):81–82, 84, July 1990. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Fiedler:1990:IUP

[Fie90h] David Fiedler. IBM and Unix: Perfect together? Byte Magazine, 15(11):51–??, Fall 1990. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Fiedler:1990:LYF

[Fie90i] David Fiedler. Let your fingers do the talking: Unix has the programs to communicate with the outside world. Byte Magazine, 15 (3):79–80, 83, March 1990. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Fiedler:1990:QUT

[Fie90j] David Fiedler. Not quite Unix:

The tribulations and treats of using a US$100 Unix clone. Byte Magazine, 15(12):119–??, November 1990. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Fiedler:1990:PNW

[Fie90k] David Fiedler. Prowling the networks: The who, why, and how of Unix network mail systems. Byte Magazine, 15(5):83– ??, May 1990. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Fiedler:1990:SIL

[Fie90l] David Fiedler. Sizing it down: A look at RISC versus CISC and the Coherent Unix-like operating system. Byte Magazine, 15 (10):105–??, October 1990. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Fiedler:1991:ANY

[Fie91a] David Fiedler. Anonymous on the net: You can have fast access to worldwide resources. Byte Magazine, 16(10):285–??, October 1991. CODEN BYTEDJ.

ISSN 0360-5280 (print), 1082-7838 (electronic).

Fiedler:1991:MPU

[Fie91b] David Fiedler. E-Mail for power users: Feature-rich mail management programs for Unix. Byte Magazine, 16(13):295–??, December 1991. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Fiedler:1991:GST

[Fie91c] David Fiedler. Gettying up to speed: Tips on talking to modems with the getty daemon. Byte Magazine, 16(9): 341–??, September 1991. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Fiedler:1991:HSL

[Fie91d] David Fiedler. Heed the standards: A look at some current standards battles and how they could affect Unix users. Byte Magazine, 16(2):107–??, February 1991. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Fiedler:1991:LMG

[Fie91e] David Fiedler. Let me down gently: Handling blackouts gracefully, and typesetting with troff. Byte Magazine, 16(8):281–??, August 1991. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Fiedler:1991:LWC

[Fie91f] David Fiedler. Lost in the woods: Changing the default scripts and parameters for new users. Byte Magazine, 16(4):338– ??, April 1991. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Fiedler:1991:NUN

[Fie91g] David Fiedler. Networking Unix:

Networking on LANs and the Internet. Byte Magazine, 16 (6):363–??, June 1991. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Fiedler:1991:PWH

[Fie91h] David Fiedler. PCs to workstations: How do you make a Unix workstation out of a 386 PC? Byte Magazine, 16(12): 375–??, November 1991. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Fiedler:1991:SHB

[Fie91i] David Fiedler. SCO hot: A brief look at the new SCO Unix and using PCs as X terminals. Byte Magazine, 16(1):101–102, 104, January 1991. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Fiedler:1991:SSV

[Fie91j] David Fiedler. Stars of the show:

A visit to UniForum. Byte Magazine, 16(5):318–??, May 1991. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Fiedler:1991:SGU

[Fie91k] David Fiedler. Survivalist’s guide to Unix: Methods of Unix system backup and protection. Byte Magazine, 16(7):285–286, 288, July 1991. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Fiedler:1991:TUG

[Fie91l] David Fiedler. Tricks of the Unix gurus: Some powerful hints for work and fun. Byte Magazine, 16(3):332, 334, 336, March 1991. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Fiezenman:1991:SPR

[Fie91m] M. Fiezenman. Smart printing: RISC is giving the latest generation of printers the punch it needs. Byte Magazine, 16(10): 139–142, 144–146, October 1991. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Fiedler:1992:MFN

[Fie92a] David Fiedler. My first network: Learning to set up a TCP/IP Ethernet for PCs running Unix. Byte Magazine, 17 (2):277–??, February 1992. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Fiedler:1992:XHS

[Fie92b] David Fiedler. X hits the spot: Setting up your PC Unix for the X Window System. Byte Magazine, 17(3):291–??, March 1992. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Finkelstein:1990:MDS

[Fin90a]  R. Finkelstein.  Multiuser

databases: The SQL: All SQL servers are not alike. here’s a look at the eight best-known products. Byte Magazine, 15(5):136– 140, 142, 144, 146, 148, 150, May 1990. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Finnie:1990:DPD

[Fin90b]  J. Scot Finnie.  Dual-page

duel: Two high-resolution monitors square off: New highresolution monitors from Cornerstone and Radius aren’t just for desktop publishing. Byte Magazine, 15(9):212–214, September 1990. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Fisher:1990:LGS

[Fis90a] S. Fisher. The latest GOSIP:

Sooner or later, the Federal Government’s new profile for procurements, called GOSIP, will affect us all. Byte Magazine, 15 (6):212–213, June 1990. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Fisher:1990:MMN

[Fis90b] S. Fisher. Mix-and-Match network adapters: Two specifications NDIS and ODLI simplify adapter driver chores. Byte Magazine, 15(8):277–279, August 1990. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Fisher:1990:NMF

[Fis90c] S. Fisher. Need more fiber: From FDDI to HDTV to broadband ISDN, fiber-optic technology is changing data communications. are you ready to make the move to fiber? Byte Magazine, 15(6): 233, 236–238, June 1990. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Fisher:1990:MMM

[Fis90d] Sharon Fisher. Making the Micro-to-Mainframe connection. Byte Magazine, 15(11):203–205, Fall 1990. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Fisher:1991:DPS

[Fis91a] S. Fisher. Dueling protocols: SNMP versus CMIP: Are they really more similar than different? Byte Magazine, 16(3):183– 184, 186, 188, 190, March 1991. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Fisher:1991:WNW

[Fis91b] S. Fisher. Whither NREN: What is the government’s proper role in providing a National Research and Education Network? Byte Magazine, 16(7):181–186, 188– 189, July 1991. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Fisher:1991:NPP

[Fis91c]  Sharon Fisher.  Networking:

Promises and problems. Byte Magazine, 16(11):117–??, ?? 1991. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Fox:1993:LRHa

[FJK^{+}93] Richard Fox, Alan Joch, Chandrika Krishnamurthy, Stephen Platt, and Leonard Presberg. Lab report: Hands-on testing 126 printers: Our testing shows which printer is best for your needs. Byte Magazine, 18(6): 146–149, 152, 154, 156–158, 160, 162, 164–166, 168, 170, 172, 174– 5, May 1993. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Fox:1993:BLR

[FJPR93] Richard Fox, Alan Joch, Leonard Presberg, and Leslie Reisz. BYTE Lab report: Desktop dynamite: 116 fast 486s: We tested 116 50-and 66-MHz 486s to tell you which is best for your DOS, Windows, and UNIX applications. how we tested; poor quality; bus choices. Byte Magazine, 18(7):156–??, June 1993. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Fleming:1993:SBB

[Fle93] Bill Fleming. Sun breaks the bottlenecks. Byte Magazine, 18(12): 218–??, November 1993. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Fountain:1991:MCP

[Fou91] ?. Fountain. Modula’s children, part II: Oberon. Byte Magazine, 16(??):??, ?? 1991. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Fowler:1991:PNI

[Fow91] D. Fowler. Perpetual networks: If the network is the computer, then you’d better be able to keep the power switch on. Byte Magazine, 16(8):205–208, 210, 212, August 1991. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Fox:1993:LRHb

[Fox93]  R. Fox.  Lab report: 90

High-Speed 486 systems: We evaluate 90 high-end 486 systems, with CPUs from Intel, Cyrix, and IBM, to find the best PCs for your applications. our rankings include ISA and EISA bus implementations, as well as local-bus designs. the best systems for Windows applications; the best systems for high-performance Windows; how we tested; Pentium performance; energy star systems; the best systems for Unix applications; honorable mentions; dubious achievements. Byte Magazine, 18(13):176–180, 183–184, 187–188, 190, 192, 193, 196, December 1993. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Fox:1993:HPN

[FP93] R. Fox and S. Platt. 62 highpower notebooks. Byte Magazine, 18(11):162–167, 170, 172, 174, 178, 181–182, 184, 189–190, October 1993. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Friedman:1990:SSP

[Fri90] D. Friedman. Sounds of success: Professional sound capabilities, once the exclusive domain of high-end recording studios, are now available to PC users. Byte Magazine, 15(9):429– 430, 434, 436, 438, 442, September 1990. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Fried:1991:PSI

[Fri91] S. S. Fried. Personal supercomputing with the Intel i860: Crunching numbers with the i860. Byte Magazine, 16(1):347– 348, 350–352, 356, 356, 358, January 1991. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Friedman:1992:RMP

[Fri92] Rich Friedman. Report from Moscow: Programming talent abounds here. Byte Magazine, 17 (9):41–??, September 1992. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Fritz:1994:CIR

[Fri94a] Jeffrey Fritz. Clearing away the ISDN roadblocks: ISDN still isn’t smoothly interoperable, but there’s hope. Byte Magazine, 19 (10):207–??, October 1994. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Fritz:1994:DRA

[Fri94b] Jeffrey Fritz. Digital remote access: No longer an expensive and hard-to-justify luxury, highspeed remote access to networks is now possible using ISDN and other digital services available from the phone companies. Byte Magazine, 19(9):125–??, September 1994. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Fritz:1994:SIS

[Fri94c] Jeffrey Fritz. Switching ISDN: Symplex communications’ DirectRoute is a flexible and powerful tool for remote LAN access. it provides aggregate bandwidth on multiple ISDN channels, data compression, and connectionoriented switching. Byte Magazine, 19(11):251–254, November 1994. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Fisher:1990:LVW

[FT90] Scott S. Fisher and Jane Morrill Tazelaar. Living in a virtual world: Head-mounted devices and stereoscopic viewers help add sound, sight, and touch to your computerized tool kit. Byte Magazine, 15(7):215–216, 218, 220– 221, July 1990. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Fersko-Weiss:1993:CMK

[FW93] Henry Fersko-Weiss. Contact managers: Keeping in touch.

Byte Magazine, 18(5):109–??, Spring 1993. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

GaCote:1992:LTD

[Ga 92a] Raymond Ga Cote. Lab tests:

Does brand matter? Byte Magazine, 17(14):??, December 1992. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

GaCote:1992:SXL

[Ga 92b] Raymond Ga Cote. Stepping up to XVT 3.0: The latest XVT marks interface library improvements and new development tools. Byte Magazine, 17(14):224–??, December 1992. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

GaCote:1993:MW

[Ga 93a] R. Ga Cote. Mac for workgroups.

Byte Magazine, 18(11):137–138, October 1993. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

GaCote:1993:FES

[Ga 93b] Raymond Ga Cote. A FirstClass experience: SoftArc’s FirstClass mail and conferencing system goes multiplatform. its elegant client interface, strong conferencing features, and multitasking communications engine make it a real contender in the E-mail market. Byte Magazine, 18(10):149– 150, 152, September 1993. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

GaCote:1993:MBT

[Ga 93c] Raymond Ga Cote. Mathcad: Better than paper. Byte Magazine, 18(5):115–??, Spring 1993. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

GaCote:1993:RIK

[Ga 93d] Raymond Ga Cote. The Renaissance of imaging: Kodak hopes to change the way graphics professionals process electronic images with Photo CD and supporting software. BYTE examines two of Kodak’s flagship Photo CD software products, PhotoEdge and Renaissance. Byte Magazine, 18 (7):146–147, June 1993. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

GaCote:1993:TWS

[Ga 93e] Raymond Ga Cote. Two ways to say VL-Bus: Testing two motherboards that mix VL-Bus and EISA. Byte Magazine, 18 (4):179–??, April 1993. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

GaCote:1994:CTM

[Ga 94a] Raymond Ga Cote. Commentary: Think multiplatform: A software engineer’s perspective of multiplatform applications development. Byte Magazine, 19 (6):344–??, June 1994. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

GaCote:1994:DTP

[Ga 94b] Raymond Ga Cote. Desktop telephony: PhonePro makes it easy to build phone automation applications for the Mac. Byte Magazine, 19(3):151–152, 154, March 1994. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

GaCote:1994:MPP

[Ga 94c] Raymond Ga Cote. Mac programming power tools: Cote finds mainstay’s new C and BASIC programming environments for the Macintosh hard to pigeonhole and harder to resist. Byte Magazine, 19(4):199– ??, April 1994. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

GaCote:1994:MHB

[Ga 94d] Raymond Ga Cote. Multiprocessor horsepower: BYTE compares three dual-processor PCs from ALR, Compaq, and VTech. running two processors — whether 486 or Pentium — gives a big boost to Microsoft’s SQL server under Windows NT. Byte Magazine, 19(3):129–??, March 1994. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Ga:1994:CPW

[Ga94]  Cote Raymond Ga.  Cross-

platform warrior: Metrowerk’s CodeWarrior is a powerful, exciting development environment for both 680x0 Macintosh and Power PC platforms. Byte Magazine, 19 (10):163–??, October 1994. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Grehan:1990:PHL

[GAM90] R. Grehan, S. Apiki, and R. Mitchell. 386SX PCs: Heirs to the low end. Byte Magazine, 15(8):152–164, 166, August 1990. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Garland:1994:SOS

[Gar94] Eric Garland. SparcStation overhaul: Sun has updated its graphics workstation lineup. BYTE looks at the inexpensive Sparc-Station 5, powered by Sun’s new MicroSparc II processor, and the SparcStation 20, which holds up to four SuperSparc processors. Byte Magazine, 19(9): 161–??, September 1994. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Gasteiger:1993:BTS

[Gas93] Daniel Gasteiger. The big three square off: All three of the market-leading Windows spreadsheets — Borland Quattro Pro for Windows, Lotus 1-2-3 for Windows, and Microsoft Excel — have undergone major product upgrades. Gasteiger sorts out the field and determines how each competitor stacks up. Byte Magazine, 18(13):138–140, 142, 144, 146, 148, December 1993. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

GaCote:1993:NFT

[GAW93] Raymond Ga Cote, Steve Apiki, and Stan Wszola. Network fax on tap: The BYTE Lab peers into 11 fax servers for PC, Mac, and Unix networks. Byte Magazine, 18(2):178–180, 182, 184–190, 192, 194, 196, February 1993. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Gantt:1993:PRG

[GB93] James D. Gantt and Catherine M. Beise. The public reacts to group decision support systems. Byte Magazine, 18(3): ??, March 1993. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

GaCote:1992:SCT

[GD92a] R. Ga Cote and S. Diehl.

Searching for common threads: Ten text management packages help you organize your files. Byte Magazine, 17(6):290–305, June 1992. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Grehan:1992:BLPa

[GD92b] R. Grehan and S. Diehl. BYTE Lab product report: DOS data at work: In our first BYTE Lab product report, we sample the DOS database market. Byte Magazine, 17(1):226–228, 230, 234, 236, 240, 242, 244–246, 248, January 1992. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Gershon:1993:DDV

[GD93] Nahum Gershon and Jeff Dozier. The difficulty with data: Visualization requires diverse data types and formats. Byte Magazine, 18(4):143–??, April 1993. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

GaCote:1992:BLP

[GE92] R. Ga Cote and D. L. Edwards. BYTE Lab product report: Captains of crunch: The top spreadsheet programs for DOS, Windows, and the Mac. Byte Magazine, 17(3):222–224, 228–230, 232, 234–238, March 1992. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Geary:1990:IWL

[Gea90] M. Geary. INSIDE WINDOWS 3.0 A long and winding road: Why your old Windows application may not work. Byte Magazine, 15(8):133–134, 136, 138, 140, August 1990. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Grehan:1991:GGS

[GETY91] R. Grehan, H. Eglowstein, T. Thompson, and T. Yager. Getting groups on schedule: The BYTE Lab tests 16 group schedulers. Byte Magazine, 16(9): 250–254, 256–258, 260, 262, 264, September 1991. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

GaCote:1994:BCR

[GFK94] Raymond Ga Cote, Rich Friedman, and Russell Kay. Books and CD-ROMS: Build Power Macintosh applications: Power Mac program development, a science CD-ROM, and object-oriented programming languages. Byte Magazine, 19(10):45–??, October 1994. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Gibson:1990:BMB

[Gib90a] C. R. Gibson. 9600-bps modem brings Apple networks closer together: Thanks to its AppleTalk connector, Shiva’s Net-Modem V.32 can serve as a shared network modem and an internetwork router. Byte Magazine, 15(9):185–186, 88, September 1990. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Gibson:1990:OBB

[Gib90b] E. Gibson. Objects — born and bred: Object behavior analysis provides a conceptual model for the first stage in creating an object-oriented application. Byte Magazine, 15(10):245–254, October 1990. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Gibson:1992:CES

[Gib92] C. R. Gibson. Claris enters the spreadsheet wars: A strong contender for the Macintosh spreadsheet crown elbows in on Excel’s and Wingz’s turf. Byte Magazine, 17(1):273–274, 276, January 1992. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Gillmor:1994:NSW

[Gil94] Steve Gillmor. A new synergy for Windows: With Prodea Synergy, you can easily automate interapplication Windows tasks. Gilmor builds cross-application projects with off-the-shelf Windows software and the Prodes tools, doing the same chores without ProdeaSynergy is difficult and sometimes impossible. Byte Magazine, 19(2): 115–??, February 1994. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Gephardt:1991:DLS

[GK91] D. Gephardt and M. C. Klonower. Destination laptop: Squeezing the components of an AT-class machine onto a single chip will have a great impact on the portable computers of the future. Byte Magazine, 16(2):239– 242, 246, February 1991. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Glass:1990:FWT

[Gla90a] L. B. Glass. Font wars (Type 1 and TrueType fonts). Byte Magazine, 15(8):289–290, 292–295, August 1990. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Glass:1990:GPM

[Gla90b] L. B. Glass. Gateways to protected mode: DOS extenders deliver 16-bit compatibility and 32bit performance. Byte Magazine, 15(4):287–288, 290, 292, 294– 295, April 1990. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Glass:1990:RDC

[Gla90c] L. B. Glass. Reeling in the data: Competing technologies are driving down the costs of tape backup units while increasing capacities. Byte Magazine, 15(5): 299–302, 304, 306, May 1990. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Glass:1990:SBPb

[Gla90d] L. B. Glass. The SCSI bus, part II: Brett looks at bus facilities, the common command set, the common access method, and SCSI devices. Byte Magazine, 15(3):291–294, 296, 298, March 1990. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Glass:1990:DPM

[Gla90e] L. Brett Glass. DPMI: The DOS protected mode interface. Byte Magazine, 15(11):113–114, 116, 118, Fall 1990. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Glass:1990:MCL

[Gla90f] L. Brett Glass. Math coprocessors: A look at what they do, and how they do it. Byte Magazine, 15(1):337–348, January 1990. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Glass:1990:SBPa

[Gla90g] L. Brett Glass. The SCSI bus, part I: The start of a two-part look at the SCSI I/O bus. Byte Magazine, 15(2):267–268, 270– 272, 274, February 1990. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Glassner:1990:RTR

[Gla90h] Andrew S. Glassner. Ray tracing for realism: Simulating light rays in a 3-D scene. Byte Magazine, 15 (13):263–264, 266, 268–271, December 1990. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic). nice introductory programming article.

Glass:1991:QSU

[Gla91a]  B. Glass.  QEMM-386 and

386Max square off under Windows: Contenders for the extended memory of Windows users. Byte Magazine, 16(5): 257–258, 260, May 1991. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Glass:1991:SRN

[Gla91b] B. Glass. SPARC revealed: A new breed of workstation evolves from a clonable RISC CPU.

Byte Magazine, 16(4):295–296, 298, 300–302, April 1991. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Glass:1991:RAA

[Gla91c] Brett Glass. The return of AR-Cnet: ARCnet plus is a fast alternative to Ethernet and Token Ring hardware, and it’s downward-compatible with AR-Cnet. Byte Magazine, 16(2): 119–??, February 1991. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Glass:1991:WNM

[Gla91d] Brett Glass. Windows 3.0 and networks: A marriage (almost) made in heaven. Byte Magazine, 16(4):343–344, 347, April 1991. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

GaCote:1993:PPP

[GN93] Raymond Ga Cote and Barry Nance. Pentium PCs: Power to burn: Fast and ready to roll, the first Pentium systems are now available. Byte Magazine, 18(8): 94–96, 98, 100, 102, July 1993. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Gelernter:1990:SYF

[GP90] David Gelernter and James Philbin. Spending your free time: Attached to a network, your existing computers can act as a powerful parallel computer. Byte Magazine, 15(5):213–214, 216– 219, May 1990. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Gehani:1990:CCA

[GR90] N. H. Gehani and W. D. Roome. Concurrent C: An AT&T language for programming multiprocessor systems. Byte Magazine, 15(13):327–330, 332, 334, December 1990. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Garland:1994:FFC

[GR94] Eric Garland and Dave Rowell. Face-to-Face collaboration: BYTE takes an in-depth look at a sampling of desktop videoconferencing systems. Byte Magazine, 19(11):233–237, 240–242, November 1994. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Grehan:1990:CDE

[Gre90a] R. Grehan. Cloak and data: An explanation of secret codes and a puzzle to test your skill. Byte Magazine, 15(6):311–312, 314, 316, 318, 320, 322, 324, June 1990. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Grehan:1990:FFA

[Gre90b] R. Grehan. FPU Face-Off: Not all FPUs are created equal. the BYTE Lab shows performance differences among FPUs from AMD, Cyrix, IIT, Intel, and Weitek. Byte Magazine, 15 (12):194–196, 198, 200, November 1990. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Grehan:1990:ECW

[Gre90c] R. Grehan. In any event: Comparing the ways that PC, Mac, and Unix software juggle input from mice, keyboards, and other sources. Byte Magazine, 15(5): 311–312, 314, 316, 318, 320, 322, May 1990. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Grehan:1990:JBF

[Gre90d] R. Grehan. Just between friends:

Talking tasks: The first part of a series on interprocess communications looks at Microsoft Windows and Desqview. Byte Magazine, 15(10):311–312, 314, 316, 318, 320, October 1990. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Grehan:1990:MTJ

[Gre90e] R. Grehan. More than just fast:

A look at programming SCSI devices on Macs and MS-DOS machines. Byte Magazine, 15 (13):361–366, 368–369, December 1990. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Grehan:1990:SARa

[Gre90f] R. Grehan. Some assembly required: Foreign file systems: Using special file systems from within standard file systems. Byte Magazine, 15(3):301–303, 306–308, 310, 312, March 1990. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Grehan:1990:SARb

[Gre90g] R. Grehan. Some assembly required. talking tasks, part 2: Introducing the most common methods of interprocess communications for Unix and OS/2. Byte Magazine, 15(12):403–404, 406, 408, 410–412, 414, 416, November 1990. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Grehan:1990:VVM

[Gre90h] R. Grehan. Virtually virtual memory: A memory management system for MS-DOS that lets you break the 640K-byte barrier. Byte Magazine, 15(9):455– 456, 458–460, 62, 64, September 1990. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Grehan:1990:HDM

[Gre90i] Rick Grehan. Hard drivin’ Mac: Utility software distinguishes 300-megabyte Mac hard disk drives from MicroNet, Racet, and Jasmine. Byte Magazine, 15 (1):183–188, January 1990. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Grehan:1990:MMA

[Gre90j] Rick Grehan. Multitasking for the masses: An analysis of different tools to put multitasking on your desk with just a PC or a Macintosh. Byte Magazine, 15(2):279–280, 282, 284, 286, 288, 334, February 1990. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Grehan:1990:SCG

[Gre90k] Rick Grehan. Stroke-Character graphics: Using stroke characters in PC graphics mode. Byte Magazine, 15(1):351–364, 414, January 1990. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Green:1991:RCR

[Gre91a] R. Green. Remote connections:

Remote sites can share common information and send messages to each other, courtesy of a widearea network. Byte Magazine, 16 (7):161–162, 164–168, July 1991. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Grehan:1991:EBP

[Gre91b] R. Grehan. Ethernet’s 32-bit players: EISA Ethernet adapters break LAN performance bottlenecks. Byte Magazine, 16 (3):240–243, March 1991. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Grehan:1991:SUA

[Gre91c] R. Grehan. “smart” UPSes alert LANs to power problems: New LAN-based UPSes prevent damage. Byte Magazine, 16(5):249– 250, 252, 254, May 1991. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Grehan:1992:SAR

[Gre92] Rick Grehan. Some assembly required: A shared resource access manager, part 2: Introducing a resource access manager server and C client functions that you can link into your applications. Byte Magazine, 17(9): 297–??, September 1992. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Grehan:1993:BSF

[Gre93a] Rick Grehan. Building SQL front ends: SQL databases enable sharing of critical information throughout the corporation, but creating client applications can be a tough row to hoe. BYTE examines client construction tools from Borland, Gupta, KnowledgeWare, and PowerSoft that provide technology that can ease the process. Byte Magazine, 18(12):240–??, November 1993. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Grehan:1993:CDW

[Gre93b] Rick Grehan. C++ does Windows: The BYTE Lab tests five C/C++ compilers, from Microsoft, Borland, Symantec, MetaWare, and Watcom, to determine how well each addresses the complexities of Windows. Byte Magazine, 18(10):130–??, September 1993. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Grehan:1993:MWR

[Gre93c] Rick Grehan. Making Windows rock and roll: The BYTE Lab looks at 16 accelerators that speed up Windows operations (for as little as US$139). Byte Magazine, 18(1):202–??, January 1993. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Grehan:1993:MIR

[Gre93d] Rick Grehan. Mips inside: The RISC PC: DeskStation technology is not the first to build a PC for NT based on a Mips processor, but it is the first to price machines directly against 486s. test results based on BYTE’s preliminary Portable Benchmarks illustrate the speed you can expect from the Evolution RISC PC. Byte Magazine, 18(9):145– ??, August 1993. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Grehan:1993:NTS

[Gre93e] Rick Grehan. New tricks for slow Macs: The BYTE Lab tests 17 Mac accelerator boards. Byte Magazine, 18(2):198–??, February 1993. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Grehan:1993:RNH

[Gre93f] Rick Grehan. Reviewer’s notebook: How to give Windows a workout: The BYTE Lab introduces new Windows benchmarks. Byte Magazine, 18(1):

246–??, January 1993. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Grehan:1993:RNI

[Gre93g] Rick Grehan. Reviewer’s notebook: It worked fine a minute ago: Reflections on Macintosh compatibility problems. Byte Magazine, 18(3):183–??, March 1993. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Grehan:1993:TTA

[Gre93h] Rick Grehan. A tale of two Alphas: An under-the-hood exploration of two Alpha-based systems from DEC: The DEC 3000 Model 300, a Unix box, and the DECpc AXP 150, which runs Windows NT. Byte Magazine, 18(13):169–??, December 1993. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Grehan:1994:OOC

[Gre94a] Rick Grehan. Object-Oriented COBOL: Object-oriented — hold on to your hat — COBOL is in the works. how could such a thing remain true to the spirit of the language — and what would it look like? Byte Magazine, 19 (9):197–??, September 1994. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Grehan:1994:PMP

[Gre94b] Rick Grehan. Poet in motion: Poet 2.1 is a true object-oriented database that also includes all the features of a full-blown multiuser database: compound indexes, locks, even multilevel transactions. Grehan explains how programmers can use the same structures of C++ code to support the database. Byte Magazine, 19(5):137–??, May 1994. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Grehan:1994:PMA

[Gre94c] Rick Grehan. Port Mac applications to the PowerPC: Apple finds a way to ease the task of moving programs over to the Power Mac. Byte Magazine, 19(6):273–??, June 1994. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Grehan:1994:PPW

[Gre94d] Rick Grehan. Processors proliferate: While most CPU designers seek bigger processors with more power, others are looking for less. small is a strategy that’s working remarkably well in niche markets. Byte Magazine, 19(9): 67–??, September 1994. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Grehan:1994:WCC

[Gre94e] Rick Grehan. Watcom C/C++ get a new face: With version 10.0 of its C/C++ compiler, Watcom has added a complete graphical development environment and enhanced cross-platform capabilities. from a DOS, OS/2, Windows, or Windows NT host, you can now generate executables for DOS, Windows, Windows NT, OS/2, Novell NetWare, and AutoCAD. and from a single host you can produce both 16-and 32bit code. Byte Magazine, 19 (10):153–??, October 1994. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Griswold:1994:SAR

[Gri94] Ralph E. Griswold. Some assembly required: The Icon Programming Language: Icon does string and structure processing and numerical computation. Byte Magazine, 19(5):193–??, May 1994. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Grudin:1990:FCS

[Gru90] J. Grudin. A foolish consistency: A software engineer argues that consistency isn’t always the best policy when it comes to user interfaces. Byte Magazine, 15(3): 364, March 1990. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

GaCote:1993:BCR

[GT93] Raymond Ga Cote and Tom Thompson. Books and CD-ROMs: Entrepreneurial enterprise: Starting a business, a classic dictionary on CD-ROM, and more. Byte Magazine, 18(13): 49–??, December 1993. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Grehan:1994:BCRa

[GTGW94] Rick Grehan, Tom Thompson, Raymond Ga Cote, and Scott Wallace. Books and CD-ROMs: Levels of secrecy: A look at cryptography, Marvin Minsky on CD-ROM, “plugs,” and visualization of scientific data. Byte Magazine, 19(6):41–??, June 1994. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Grehan:1994:BCRb

[GTN94] Rick Grehan, Tom Thompson, and Michael Nadeau. Books and CD-ROMs: Embedded systems programming: Developing embedded systems, a Holocaust story on CD-ROM, and a look at the future of publishing. Byte Magazine, 19(8):49– ??, August 1994. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Gutmann:1993:CPP

[Gut93] Michael J. Gutmann. Cluster PCs for power: A look at network high-end PCs able to run applications that were once too big for your server. Byte Magazine, 18(8):57–??, July 1993. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Guttag:1994:MPD

[Gut94] Karl M. Guttag. Multimedia powerhouse: The designer of the TI MVP discusses the multimedia functions that drove the design of the chip. Byte Magazine, 19(6):57–??, June 1994. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Grehan:1992:BLPb

[GW92] Rick Grehan and Stan Wszola. BYTE Lab product report workhorse computers: The Lab evaluates the new low end of computing: 386SXes. Byte Magazine, 17(9):215–217, 220, 222, 226–229, 232, 234, September 1992. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Grehan:1993:SFS

[GW93] Rick Grehan and Stan Wszola. Solutions focus: Shrink to fit: The BYTE Lab tests on-the-fly data compressors for Macs and PCs. Byte Magazine, 18(4):150– ??, April 1993. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Halstead:1990:DAE

[Hal90] Rodd Halstead. Develop advanced expert systems: Gold Hill’s new expert-system shell works with Microsoft Windows. Byte Magazine, 15(1):219–224, January 1990. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Halfhill:1993:BCR

[Hal93a] Tom Halfhill. Buying a CD-ROM drive. Byte Magazine, 18(2):??, February 1993. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Halfhill:1993:CMC

[Hal93b] Tom R. Halfhill. ClarisWorks 2.0 for Macintosh: ClarisWorks is already established as the leading integral package for the Macintosh, but it’s not resting on its existing modules. version 2.0 adds new features and applications to this seamlessly integrated software. Byte Magazine, 18(8):151–??, July 1993. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Halfhill:1993:ILR

[Hal93c] Tom R. Halfhill. Intel launches rocket in a socket: Intel’s new Pentium CPU doubles the speed of the fastest 486. Byte Magazine, 18(6):92–94, 96, 98, 100, 102–104, 106, 108, May 1993. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Halfhill:1993:PAP

[Hal93d] Tom R. Halfhill. A peek at PowerOpen: The first PowerPC chip may not equal the Pentium’s performance, but at onefifth the price, you may not care. Byte Magazine, 18(7):46– ??, June 1993. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Halfhill:1993:RWM

[Hal93e] Tom R. Halfhill. Ruling won’t mean lower prices for 486 chips: Despite AMD’s winning the latest round in its legal battle with Intel, don’t expect a big price drop in 486 systems. Byte Magazine, 18(8):28–??, July 1993. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Halfhill:1993:VBS

[Hal93f] Tom R. Halfhill. Visual Basic 3.0 strengthens connectivity: The newest version of Visual Basic inherits Microsoft’s Access database engine. Byte Magazine, 18(7):48–??, June 1993. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Halfhill:1994:HSD

[Hal94a] Tom Halfhill. How safe is data compression: On the fly data compression can double your storage space — but not without some hassles and headaches. Byte Magazine, 19(2):56–??, February 1994. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Halfhill:1994:WIB

[Hal94b] Tom R. Halfhill. 80x86 wars: The Intel-based 80x86 continues to dominate the desktop, and it will likely continue with the anticipated introduction of new designs from Intel and its competitors. Byte Magazine, 19(6):74– ??, June 1994. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Halfhill:1994:AVS

[Hal94c] Tom R. Halfhill. AMD vs. Superman: The AMD K5 is an innovative marriage of RISC and CISC technology that will challenge Intel’s strangehold on the high end of the PC market. Byte Magazine, 19(11):95–??, November 1994. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Halfhill:1994:CRP

[Hal94d] Tom R. Halfhill. Commentary: R.I.P. Commodore 1954-1994: A look at the company that introduced millions to personal computing. Byte Magazine, 19(8):252–??, August 1994. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Halfhill:1994:CIA

[Hal94e] Tom R. Halfhill. Commentary:

The introversion of America: Are virtual communities on the Internet taking the place of neighborhood communities? Byte Magazine, 19(5):278–??, May 1994. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Halfhill:1994:ERS

[Hal94f] Tom R. Halfhill. Emulation:

RISC’s secret weapon: Emulation is beginning to appear as an integral feature in some new operating systems, and its ambitious goal is to shield users from the incompatibilities of different hardware and operating-system platforms. Byte Magazine, 19 (4):119–??, April 1994. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Halfhill:1994:MCC

[Hal94g] Tom R. Halfhill. Managing color in Chicago. Byte Magazine, 19(11):215–??, November 1994. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Halfhill:1994:TBF

[Hal94h] Tom R. Halfhill. T5: Brute force: Using aggressive superscalar techniques, the T5 gives a big boost to the Mips architecture as it battles Windows NT competitors Alpha and Pentium. Byte Magazine, 19(11): 123–??, November 1994. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Halfhill:1994:TPP

[Hal94i] Tom R. Halfhill. Transforming the PC: Plug and Play: For over a decade, users have struggled with the arcana of expanding a PC system. Plug and Play promises to make system configuration — and reconfiguration — a simple and painless affair. getting from here to there, however, won’t be nearly as simple or painless. Byte Magazine, 19(9):78–??, September 1994. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Hamilton:1990:MTM

[Ham90] Douglas A. Hamilton. Mastering OS/2 threads: Mastery of OS/2 threads taxes developers but rewards users. Byte Magazine, 15(9):101–102, 110, September 1990. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Hamilton:1991:SFT

[Ham91] Douglas Hamilton. A smaller, faster OS/2: OS/2 1.3: technical success, marketing flop. Byte Magazine, 16(4):347–??, April 1991. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Hamilton:1992:GWI

[Ham92] Douglas A. Hamilton. OS/2 2.0 goes down to the wire: IBM’s OS/2 2.0 limited availability release is less than palatable. Byte Magazine, 17(5):301– ??, May 1992. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Hansson:1992:BTW

[Han92a] Othar Hansson. Borland targets Windows developers with latest C++ release: Borland C++ 3.1 contains some features we’ve been waiting for. Byte Magazine, 17(12):253–??, November 1992. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Hansson:1992:MLN

[Han92b] Othar Hansson. Microsoft’s lucky number: Microsoft joins the C++ minions with its newest compiler. Byte Magazine, 17(9): 273–??, September 1992. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Harbison:1990:MIO

[Har90a] S. Harbison. Modula-3: An introduction to the OOP language that grew from Pascal and Modula-2. Byte Magazine, 15 (12):385–388, 390, 392, November 1990. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Harvey:1990:OSP

[Har90b] David A. Harvey. Optical storage primer. Byte Magazine, 15 (11):121–122, 124, 126, 128, 130, Fall 1990. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Harvey:1990:SM

[Har90c] David A. Harvey. State of the media. Byte Magazine, 15(??): 275–282, November 1990. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Harvey:1991:CWD

[Har91a] D. A. Harvey. Catch the wave of DIP: Document image processing brings the goal of a paperless office closer to reality — today. Byte Magazine, 16 (4):173–174, 176, 178–180, 182, April 1991. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Harvey:1991:CRD

[Har91b] D. A. Harvey. CD-ROM drives: How good is the third generation: Will CD-ROM’s promise be fulfilled? Byte Magazine, 16(9):

268–270, 272, 274, 276, September 1991. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Harvey:1991:HSF

[Har91c] D. A. Harvey. Health and safety first: Ergonomic design can make using personal computers safer and more productive. Byte Magazine, 16(10):119–120, 122, 124– 126, 128, October 1991. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Hardman:1993:NAT

[Har93] Harriett Hardman. New authoring tools for Windows: Two new Windows applications deliver different approaches for authoring multimedia titles. Authorware Professional 2.0 is more suitable for interactive training and information kiosks, while Microsoft Viewer 2.0 has stronger support for indexing, searching, and linking large textural databases. Byte Magazine, 18(9): 153–154, 156, August 1993. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Hayes:1990:TVF

[Hay90a] F. Hayes. From TTY to VUI:

Frank Hayes discusses the past, present, and future of userinterface design. Byte Magazine, 15(4):205–206, 208, 210– 211, April 1990. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Hayes:1990:SCB

[Hay90b] F. Hayes. The spirit of ’86s: The competition between PCcompatible CPUs heats up, as faster 80286s, 80386 clones, and the high-powered 1486 emerge. Byte Magazine, 15(3):266–270, March 1990. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Hayes:1993:PTB

[Hay93] Frank Hayes. The printers talk back: The Network Printing Alliance Protocol will provide communications capabilities to your network printer. now you’ll know when the printer is low on toner, out of paper, and more. Byte Magazine, 18(13):103–??, December 1993. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Hayes:1994:PPM

[Hay94a] Frank Hayes. Personality plus: Multiple operating-system personalities are here to stay. Byte Magazine, 19(1):155–??, January 1994. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Hayes:1994:TCC

[Hay94b] Frank Hayes. Today’s compilers: Challenged by multiple new processors, parallel machines, and operating systems, compiler builders must rework their mainline products to produce faster, leaner code. Byte Magazine, 19(2):76–??, February 1994. CODEN BYTEDJ.

ISSN 0360-5280 (print), 1082-7838 (electronic).

Hurwicz:1993:MAT

[HC93] Mike Hurwicz and Dan Carroll. Merlin and AlertView: Two packages for managing PC networks have similar capabilities but drastically different orientations. AlertView takes action on workstation events, while NLMerlin automates procedures with a focus on NetWare servers. Hurwicz and Carroll question the long-term viability of each. Byte Magazine, 18(12):271–??, November 1993. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Hedberg:1993:NKT

[Hed93a] Sara Hedberg. New knowledge tools: Combining knowledge systems with other technologies can improve your cost/ performance figures. Byte Magazine, 18(8):106–??, July 1993. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Hedberg:1993:SHL

[Hed93b] Sara Hedberg. See, hear, learn:

With smart multimedia and virtual reality, you can create virtual Cheshire cats to answer your questions. projects at Northwestern University and Andersen Consulting are putting this technology to use. Byte Magazine, 18(8):119, 121, 123, 125, 127– 128, July 1993. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Hedberg:1994:DLP

[Hed94] Sara Reese Hedberg. Design of a lifetime: Product designers today face new requirements: They must account for the entire life cycle up front. Byte Magazine, 19 (10):103–??, October 1994. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Heller:1990:JAW

[Hel90a] M. Heller. Just add water: Rational Systems’ Instant-C 4.0 promises instant software gratification. Byte Magazine, 15 (6):188, 190, June 1990. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Heller:1990:ZSB

[Hel90b] M. Heller. OS/2 1.2: A zaftig system: Beauty goes more than skin deep in IBM’s newest OS/2 1.2. Byte Magazine, 15(3):167– 168, 170–171, March 1990. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Heller:1990:BWI

[Hel90c] Martin Heller. Beyond DOS: Windows and OS/2. I’ve got DIBs device-independent bit maps and palette management. Byte Magazine, 15(13):101–102, 104, December 1990. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Heller:1990:PB

[Hel90d] Martin Heller. Programming 32bit OS/2. Byte Magazine, 15(11): 97–100, 102, 104, Fall 1990. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Heller:1990:TPR

[Hel90e] Martin Heller. Through the OS/2 porthole: OS/2 rolls out the red carpet for Windows applications. Byte Magazine, 15 (10):127–??, October 1990. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Heller:1991:APR

[Hel91a] M. Heller. AI in practice: A real company’s real-world use of AI techniques and methods. Byte Magazine, 16(1):267–270, 272, 274, 276, 278, January 1991. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Heller:1991:FDC

[Hel91b] M. Heller. Future documents:

Combine Windows text, graphics, voice, and more. Byte Magazine, 16(5):126–129, 132, 134– 135, May 1991. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Heller:1991:BWW

[Hel91c] Martin Heller. Beyond DOS:

Windows and OS/2. Windows programming made easy. Byte Magazine, 16(3):326–328, March 1991. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Heller:1991:HYB

[Hel91d] Martin Heller. Hedging your bets: Survival strategies for cross-platform developers. Byte Magazine, 16(8):291–??, August 1991. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Heller:1991:SSS

[Hel91e] Martin Heller. Setting up shop: Setting up to program for Windows. Byte Magazine, 16(12): 367–??, November 1991. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Heller:1991:WMA

[Hel91f] Martin Heller. Windows meets AI: AI applications come to Windows. Byte Magazine, 16(6):351– ??, June 1991. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Heller:1992:BWT

[Hel92a] Martin Heller. 32-bit Windows today: Watcom and MetaWare deliver 32-bit Windows programming toolkits. Byte Magazine, 17(3):281–??, March 1992. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Heller:1992:BIE

[Hel92b] Martin Heller. Beyond DOS inside 386 enhanced mode: Under the covers of Windows 386 enhanced mode. Byte Magazine, 17(4):297–??, April 1992. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Hettler:1994:SRS

[HH94a] Mark Hettler and Scott Higgs.

Software roundup: SQL front ends for Windows: NSTL evaluates three high-end SQL packages for developing sophisticated applications in a client/server environment. PowerBuilder, SQL Windows, and Object View are tested for performance, versatility, power, and usability. Byte Magazine, 19(10):129–??, October 1994. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Hsu:1994:WFL

[HH94b] Meichun Hsu and Mike Howard.

Work-flow and legacy systems: Legacy systems are good at what they do; it’s what they don’t do that’s the problem. Byte Magazine, 19(7):109–??, July 1994. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Holzbaur:1993:LRP

[HHP93] H. Holzbaur, J. Hurd, and S. Platt. Lab report: 176 printers face off: We stress-tested 176 printers to choose the best ones for eight important business applications. the best printers for general business; enhanced drivers; how we tested; dye sublimation; color for tomorrow?; the best printers for: Highquality color; General-Purpose color; workgroups; draft quality; large format; CAD and DTP; portable; 600 vs. 300: Trading speed for quality; honorable mentions; dubious achievements. Byte Magazine, 18(12): 276–280, 282, 284–285, 289–291, 293, 295, 299, 301–302, 304–7, November 1993. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Hinton:1992:SPF

[Hin92] H. Scott Hinton. Smart pixels and free-space interconnection. Byte Magazine, 17(9):??, September 1992. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Higgs:1994:LRP

[HK94] S. Higgs and J. Kane. Lab report: 21 pyrotechnic Pentiums: Our application tests identify the best Pentium systems for general business and high-performance computing. best Pentiums for general purpose Windows; 486DX4: A 100-MHz alternative to Pentiums; best Pentiums for highperformance Windows; best Pentiums for Unix applications; how we tested; beating the heat; honorable mention. Byte Magazine, 19(8):164–167, 169, 171, 174, 176, 178–181, August 1994. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Hsu:1993:OCC

[HL93] Jeffrey Hsu and Tony Lockwood. Overview: Collaborative computing: Computer technology brings workgroups closer. Byte Magazine, 18(3):112–114, 116, 118, 120, March 1993. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Higgs:1994:LRH

[HM94] Scott Higgs and Chandrika Mysore. Lab report: Head to head: 71 printers: We use our PC and Mac printer tests to choose the best of today’s laser, inkjet, dot-matrix, and color printers for six important business applications. best for general business; as we went to press; how we tested; best for workgroups; best for CAD and DTP; best for color; best for draft quality; best for listings and forms; honorable mentions; dubious achievements. Byte Magazine, 19(5):164, 166– 169, 171, 173, 175, 177–178, 180– 184, May 1994. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Hogg:1990:PPB

[Hog90] T. Hogg. Primed for performance: Balancing the need for resources against a changing environment is a shared problem. can the solution be a shared one as well? Byte Magazine, 15(6):241–244, 246, 248, 250, June 1990. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Holtzman:1990:ELU

[Hol90] J. Holtzman. Expanding the limits: Unix and OS/2 are not the only solutions to memory problems. Byte Magazine, 15 (3):205–206, 208–210, 212, 214, March 1990. CODEN BYTEDJ.

ISSN 0360-5280 (print), 1082-7838 (electronic).

Holloway:1991:OOA

[Hol91]  T. Holloway.  The Object-

Oriented Amiga Exec: The design of the Amiga operatingsystem kernel follows the rules of object-oriented programming. Byte Magazine, 16(1):329–332, 234, January 1991. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Holub:1993:SARa

[Hol93a] Allen Holub. Some assembly required: The power of inheritance: How to take advantage of multiple inheritance in your C++ class designs. Byte Magazine, 18(6):221–??, May 1993. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Holub:1993:SARb

[Hol93b] Allen I. Holub. Some assembly required: Using C++ for directory management: Build C++ tools to manage your directories and perform other useful tasks. Byte Magazine, 18(13): 213–??, December 1993. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Holzbaur:1994:LRM

[Hol94] H. E. Holzbaur. Lab report: 26 modems: Faster than 14.4 kbps: High-speed modems flout standards but promise faster transmission and lower costs. we evaluate 26 modems to see which are ready for data and fax applications. best V.34 modems; V.34 versus V.32terbo; how we tested; impaired lines: Crash tests for modems; best V.32terbo modems; portable modems; honorable mentions. Byte Magazine, 19(7):168–173, 175–176, 178– 181, July 1994. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Hovy:1993:HMW

[Hov93a] Eduard Hovy. How MT works:

There is more than one way to perform machine translation. Byte Magazine, 18(1):167–??, January 1993. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Hovy:1993:MAY

[Hov93b] Eduard Hovy. MT at your service. Byte Magazine, 18(1):??, January 1993. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Hawkins:1993:BVD

[HP93] Bill Hawkins and Ed Puckett. Beyond DOS: Virtual device drivers for DOS: Virtual device drivers aren’t just for Windows applications. Byte Magazine, 18(10):217–??, September 1993. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Holzbaur:1994:LRP

[HP94] H. E. Holzbaur and S. M. Platt.

Lab report: 57 PCs that set the pace: Should you buy a highend 486 or a Pentium? we evaluate 37 66-MHz 486DX2 systems and 20 Pentium systems to find the best performers for Windows, Unix, and DOS applications. best Pentiums for Windows; how we tested; best Pentiums for Unix; best ISA 486s for Windows; best EISA 486s for Windows; best 486s for Unix; best 486s for DOS; honorable mentions; dubious achievements. Byte Magazine, 19(4):206–211, 215, 217, 219, 221–222, 224–228, April 1994. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Halfhill:1993:PAA

[HPPR93] T. R. Halfhill, E. Perratore, D. Pountain, and B. Ryan. PDAs arrive but aren’t quite here yet. Byte Magazine, 18(11):66–69, 72, 74, 76, 78, 80–82, 84, 86, October 1993. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Harvey:1990:SML

[HR90] D. A. Harvey and A. Reinhardt. State of the media: A look at the conflict between traditional magnetic mass storage devices and optical technologies. Byte Magazine, 15(12):275–281, November 1990. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Harvey:1991:PPW

[HR91] D. A. Harvey and B. Ryan. Practically paperless: Where need and technology fit, document image processing helps to streamline your business. Byte Magazine, 16(4):185–186, 188–190, April 1991. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Harvey:1994:WGR

[HS94] David A. Harvey and Richard Santalesa. Wireless gets real: Buoyed by new products, services, and access standards, wide-area wireless communications is ready for prime time. Byte Magazine, 19(5):90–??, May 1994. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Hubley:1992:GAU

[Hub92] Mary Hubley. GUIs, applications, and Unix: Vendor unity and Unix. Byte Magazine, 17 (10):186–??, October 1992. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Hurd:1993:BLR

[Hur93a] Jim Hurd. BYTE Lab report V.32 or better: 69 modems: We run line-impairment nd datathroughput tests to measure the efficiency of 9600-bps and faster modems. results reveal the best for high-speed communications, portability, data-only applications, and all-around communications. Byte Magazine, 18(8): 172–176, 178, 180–182, 184, 186– 187, 192–197, 189–190, 192–195, July 1993. CODEN BYTEDJ.

ISSN 0360-5280 (print), 1082-7838 (electronic).

Hurd:1993:MVC

[Hur93b] Jim Hurd. Microsoft’s visual C++: Is visual C++ a more powerful Visual Basic or a more graphical C++? Jim Hurd checks out its new tools and capabilities. Byte Magazine, 18 (7):133–??, June 1993. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Heller:1991:TWW

[HWS91] M. Heller, P. Wayner, and B. Smith. Tools for window workers: Development systems make it easier to create GUI programs. Byte Magazine, 16 (6):139–140, 142, 144, 146–148, June 1991. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Hyon:1994:UHS

[Hyo94] Jason Hyon. Under the hood: A standard for writing recordable CDs: A look at a CDrecordable standard, ISO 13490, that supports adding sessions and support for new operating systems. Byte Magazine, 19 (1):231–??, January 1994. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Ismail:1994:FTL

[IC94] Amin R. Ismail and Rhonda Copley. Fine-tune LANtastic: The LANtastic API gives you an unprecedented degree of control over your network configuration. Byte Magazine, 19(8):55– ??, August 1994. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Islam:1992:LSB

[Isl92] Mohammed N. Islam. Light switches: The best way to switch an optical signal is with another optical signal. Byte Magazine, 17 (9):183–??, September 1992. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Jacobson:1991:DIE

[Jac91] Robert Jacobson. Designing the information environment. Byte Magazine, 16(??):??, ?? 1991. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Jeffries:1993:SAR

[Jef93a] Thomas Jeffries. Some assembly required: Divide and conquer: Here’s how to debug interrupt service routines. Byte Magazine, 18(3):187–??, March 1993. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Jeffries:1993:UHM

[Jef93b] Thomas Jeffries. Under the hood:

Multimedia infrastructures: A developer’s view of Microsoft’s Windows Multimedia Extensions and IBM’s MMPM/2 architecture. Byte Magazine, 18(9):193– ??, August 1993. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Joch:1990:NBJ

[Joc90a] A. Joch. New Bubble-Jet outpaces portable printers: Canon’s new portable printer bubbles over with sharp resolution and flexibility. Byte Magazine, 15 (12):235–236, November 1990. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Joch:1990:NMN

[Joc90b] A. Joch. No-Muss, No-Fuss, Low-Cost PostScript printer: QMS’s new laser printer offers speed and automatic emulation for under US$3000. plus, two new Apple LaserWriters. Byte Magazine, 15(8):144–148, August 1990. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Joch:1990:LII

[Joc90c] Alan Joch. The LaserJet IIP: Inexpensive, not cheap: Hewlett– Packard brings affordable laser printers to the desktop. Byte Magazine, 15(2):179–183, February 1990. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Joch:1991:CPD

[Joc91] A. Joch. Color printing, Diconix style: Vibrant but slow: Kodak’s new ink-jet printer graces business documents. Byte Magazine, 16(5):287–288, 290, May 1991. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Joch:1992:RTV

[Joc92a] Alan Joch. Report from Taiwan:

Vendors want to move from being lowcost providers to technology leaders. Byte Magazine, 17 (9):44–??, September 1992. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Joch:1992:RNB

[Joc92b] Alan Joch. Reviewer’s notebook: The BYTE Lab, behind the scenes: An invitation to look over the shoulders of our testing editors. Byte Magazine, 17(12): 276–??, November 1992. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Joffe:1993:GYK

[Jof93] Daniel Joffe. Get your kicks with Switched 56: Faster than a speeding modem, Switched 56 service may be the answer to your communications needs. Byte Magazine, 18(12): 57–??, November 1993. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Kenner:1993:BCRd

[KAY93] Hugh Kenner, Steve Apiki, and Tom Yager. Book and CD-ROM reviews: From the mind of MIT: Windows NT, multimedia, MIT, Compton’s on CD-ROM, and more. Byte Magazine, 18(6):210–??, May 1993. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Kay:1994:BBC

[Kay94a] Russell Kay. Back of the bus: Connecting add-on devices to your computer can be an exercise in frustration. new buses promise to simplify the process. Byte Magazine, 19(8):108– ??, August 1994. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Kay:1994:DSS

[Kay94b] Russell Kay. Distributed and secure: Securing distributed systems starts with controlling access. Byte Magazine, 19(6):165– ??, June 1994. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Kay:1994:OUO

[Kay94c] Russell Kay. Objects in use: The object-oriented approach looks great in theory — but what’s it like out in the field? here’s a look at how object database technology is being used in a variety of real-world applications — the reasons, the benefits, the drawbacks, product selection, and wish lists. Byte Magazine, 19 (4):99–??, April 1994. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Kay:1994:SGG

[Kay94d] Russell Kay. Software goes global: U.S. software developers scramble for a piece of the international pie. Byte Magazine, 19(6):90–??, June 1994. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Kenner:1993:BCRc

[KCU^{+}93] Hugh Kenner, Raymond G. A.

Cote, Jon Udell, Tom Thompson, and Rob Mitchell. Book and CD-ROM reviews: Market as virtual reality: The death of money, Windows 3.1 insider, and other titles. Byte Magazine, 18(4):194– ??, April 1993. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Kenner:1993:RBCb

[KDG^{+}93] Hugh Kenner, Stanford Diehl, Raymond Ga Cote, Michael Nadeau, and Rick Grehan. Reviews: Books and CD-ROMs quest for the silicon grail: Tales of AI hackers, the green PC, art on CD, and other subjects. Byte Magazine, 18(8):49– ??, July 1993. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Keating:1990:FSS

[Kea90] C. Keating. A fearful symmetry: Sophisticated multiprocessing machines deserve an operating system to match. Byte Magazine, 15(5):221–222, 224, 226, 228, May 1990. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Kearns:1991:BCM

[Kea91a] S. Kearns. Borland C++ 2.0 moves into Windows territory: Borland’s latest compiler produces DOS and Windows applications. Byte Magazine, 16 (6):301–302, June 1991. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Kearns:1991:MDM

[Kea91b] S. Kearns. The MultiScope debuggers make debugging easier: The MultiScope debuggers for DOS provide GUI power. Byte Magazine, 16(5):271–272, 274, May 1991. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Kearns:1991:UIC

[Kea91c] S. Kearns. User interfaces, C++ style: Zinc’s class library brings text and graphical interfaces to your C++ applications. Byte Magazine, 16(1):218–220, January 1991. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Kenner:1990:FAA

[Ken90] H. Kenner. Flirting with assembly: Armed with a few general concepts, you can make assembly language improvements without knowing assembly. Byte Magazine, 15(4):297–298, 300, 302, April 1990. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Kenner:1992:PWW

[Ken92] Hugh Kenner. The pick of Windows: Windows books for every skill level. Byte Magazine, 17 (11):63–??, November 1992. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Kenner:1993:BCRf

[Ken93a] Hugh Kenner. Books and CD-ROM’s: Cyber worlds: William Gibson’s virtual light; the hip guide to NetWare; learning about computers on CD-ROM; nanotechnology; and other titles. Byte Magazine, 18(10): 49–??, September 1993. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Kenner:1993:BCRe

[Ken93b] Hugh Kenner. Books and CD-ROMs: The Pocket Godzilla: The impact of Nintendo, Mac networking, nanotechnology, artificial life, OS/2 programming, and other subjects. Byte Magazine, 18(9):49–??, August 1993. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Kenner:1993:CEB

[Ken93c] Hugh Kenner. Commentary:

Electronic books: Our eminent critic doesn’t share the excitement of books on CD-ROM. Byte Magazine, 18(12):404–??, November 1993. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Kendall:1994:UHI

[Ken94] Guy W. Kendall. Under the hood: Inside the PCI local bus: This new PC bus provides high throughput and self-configuring add-in cards. Byte Magazine, 19 (2):177–??, February 1994. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Keyooka:1992:BOO

[Key92] Gen Keyooka. Beyond DOS: Object-Oriented DLLs: Build reusable objects with Windows DLLs. Byte Magazine, 17(14): 257–??, December 1992. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Kenner:1993:RBCa

[KGEM93] Hugh Kenner, Raymond Ga Cote, Howard Eglowstein, and Rob Mitchell. Reviews: Books and CD-ROMs computer ruminations: Unix power tools, things that make us smart, and other titles. Byte Magazine, 18(7):51– ??, June 1993. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Kenner:1992:BCRb

[KGTD92] Hugh Kenner, Raymond Ga Cote, Tom Thompson, and Stanford Diehl. Book and CD-ROM reviews. amok in cyberspace: A look at the hacker crackdown, Macintosh programming secrets, support on site, and other selections. Byte Magazine, 17(12): 278–??, November 1992. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Kiyooka:1993:BGH

[Kiy93a] Gen Kiyooka. Beyond DOS: Getting a handle on NT: Windows NT offers outstanding exceptionhandling tools. here’s why you need them. Byte Magazine, 18 (6):229–??, May 1993. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Kiyooka:1993:SAR

[Kiy93b] Gen Kiyooka. Some assembly required: Elegant Windows dialog boxes: Reusable classes let you construct special Windows dialog boxes. Byte Magazine, 18 (9):203–??, August 1993. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Kiyooka:1994:SAR

[Kiy94] Gen Kiyooka. Some assembly required: Subclassing in OLE 2.0: On the road to object-oriented system services. Byte Magazine, 19(1):237–??, January 1994. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Kliewer:1990:MTU

[Kli90a] B. D. Kliewer. More than upsidedown mice: New trackballs from CH Products, Kensington, Logitech, MicroSpeed, and Mouse Systems. Byte Magazine, 15 (5):175–180, May 1990. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Kliewer:1990:PLC

[Kli90b] B. D. Kliewer. A paradox for LANs and C: Borland’s Paradox Engine turns the key to better performance. Byte Magazine, 15 (8):187–188, 190, August 1990. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Kliewer:1990:VMN

[Kli90c] B. D. Kliewer. VGA to the max:

A new set of extension breathes life into Super VGA hardware. Byte Magazine, 15(13):355–356, 358–360, December 1990. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Kane:1994:PGB

[KM94] J. Kane and J. McDonough. 92 printers go to battle: Ninetytwo laser, color, ink-jet, and fast dot-matrix printers go head-to-head in our latest round of realworld tests. our rankings identify the best models for six important business uses, ranging from general business to high-end color applications. the best printers for general business; how we tested; the best printers for workgroups; CAD and DTP; highquality color; draft quality; listings and forms. Byte Magazine, 19(11):262–265, 268–269, 271, 273, 277, 279–283, November 1994. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Kofer:1993:QQN

[Kof93] Chris Kofer. A quicker Quicken: A new Mac version of intuit’s personal-finance software. Byte Magazine, 18(8):40–??, July 1993. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Kogan:1994:RSW

[Kog94] Michael Kogan. Retrofitting OS/2 for SMP: OS/2 well exploit symmetric multiprocessing hardware. Byte Magazine, 19(6):267– ??, June 1994. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Korzeniowski:1993:MWD

[Kor93] Paul Korzeniowski. Make way for data: Middleware — such as remote procedure calls and message-passing systems — invisibly aids data exchange. with it, you can save development time implementing your client/ server applications. programmers don’t have to modify applications to accommodate network protocols. Byte Magazine, 18(7):113– ??, June 1993. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Kenner:1993:BCRb

[KPC93] Hugh Kenner, Dick Pountain, and Raymond G. A. Cote. Book and CD-ROM reviews: The AI debate revisited: What computers still can’t do, guides to the Internet, and updated Grolier’s multimedia encyclopedia, and more. Byte Magazine, 18(3):184– ??, March 1993. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Kenner:1993:ASG

[KPUG93] Hugh Kenner, Dick Pountain, Jon Udell, and Raymond Ga Cote. Again the swinging gates: Another look at Microsoft’s leader, a controversial Windows book, how to program in Oberon, and more. Byte Magazine, 18(2):222–??, February 1993. CODEN BYTEDJ.

ISSN 0360-5280 (print), 1082-7838 (electronic).

Kenner:1993:BCRa

[KRCT93] Hugh Kenner, Bob Ryan, Raymond G. A. Cote, and Tom Thompson. Book and CD-ROM reviews: Going-on at the edge: The science of complexity. Byte Magazine, 18(1):248– ??, January 1993. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Krechmer:1991:HSS

[Kre91] K. Krechmer. High-Speed safety: How to ensure integrity and maximum serial communications. Byte Magazine, 16(4):143–144, 146, 148, 150, 152, April 1991. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Kron:1990:SDL

[Kro90] P. Kron. A software developer looks at OS/2: OS/2 represents a rich arena for developers, but with pitfalls. Byte Magazine, 15 (8):269–272, August 1990. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Kenner:1992:BCRa

[KUD92] Hugh Kenner, John Unger, and Stanford Diehl. Book and CD-ROM reviews: The origins of hypertext: Vannevar Bush and hypertext, street maps on CD-ROM, and Unix primers. Byte Magazine, 17(9):286–??, September 1992. CODEN BYTEDJ.

ISSN 0360-5280 (print), 1082-7838 (electronic).

Kenner:1992:ALP

[KYCD92] Hugh Kenner, Tom Yager, Raymond G. A. Cote, and Stanford Diehl. Is artificial life possible: Steven Levy’s artificial life, Visual Basic for beginners and pros, and corporate snapshots on CD-ROM. Byte Magazine, 17 (10):256–??, October 1992. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Kenner:1992:BCRc

[KYT^{+}92] Hugh Kenner, Tom Yager, Tom Thompson, Jon Udell, and Stanford Diehl. Book and CD-ROM reviews: Photographic lies: The search for visual truth, design tips, TCP/IP explainer, the times of physicist Richard Feynman, and more. Byte Magazine, 17(14):234–??, December 1992. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Lab:1992:RNP

[Lab92a] The Byte Lab. Reviewer’s notebook: A personal librarian for Windows and a low-cost bundle of applications for the Macintosh. Byte Magazine, 17(9): 285–??, September 1992. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Lab:1992:RNB

[Lab92b] The Byte Lab. Reviewer’s notebook: The BYTE Lab finds a solution for testing removable storage devices: CorelSCSI and Adaptec controller. Byte Magazine, 17(10):254–??, October 1992. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Lindhorst:1991:PTT

[LAD91] G. Lindhorst, A. Anderson, and D. Dahms. Programming the 68040: Tricks and traps of software design for this high-powered processor. Byte Magazine, 16 (8):121–122, 124, 126, 128, August 1991. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Laffey:1991:RTE

[Laf91] T. J. Laffey. The Real-Time expert: Expert systems designed to work in real-time environments can make complex systems easier to handle. Byte Magazine, 16(1):259–260, 262, 264, January 1991. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Lancaster:1990:HPP

[Lan90a] D. Lancaster. High-performance PostScript (word processing package). Byte Magazine, 15 (8):297–300, August 1990. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Lancaster:1990:PIS

[Lan90b] D. Lancaster. PostScript insider secrets: Don Lancaster, a pioneer of the microcomputer revolution, reveals PostScript tricks for better output. Byte Magazine, 15 (7):293–294, 296, 298, 300, 302, July 1990. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Lane:1990:GMY

[Lan90c] Alex Lane. Get the max from your 80386: 386Max breaks through MS-DOS’s 640K-byte barrier. Byte Magazine, 15(2): 191–194, February 1990. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Linderholm:1992:PGM

[LAN92] O. Linderholm, S. Apiki, and M. Nadeau. The PC gets more personal: Apple and traditional notebook PC vendors take divergent paths toward the ultimate personal computer. Byte Magazine, 17(7):128–134, 136, 138, July 1992. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Lane:1994:DRR

[Lan94a] Alex Lane. Developing for RISC:

RISC development tools must offer more variety and greater ease of use. Byte Magazine, 19(4):139–??, April 1994. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Lane:1994:OTC

[Lan94b] Alex Lane. Optimizing for today’s CPUs: Making code run faster and smaller is an important goal of modern compilers. Byte Magazine, 19(2):81–??, February 1994. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Lavin:1991:OBP

[Lav91] P. Lavin. Olivetti broadens its product line. Byte Magazine, 16(7):72IS/55–56, 58, 60, July 1991. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Lawrence:1994:TSD

[Law94] Bill Lawrence. Three suits deals: Office application suites are taking the market by storm. but do you really get extra functionality and smoother integration by turning to a single vendor for all your major applications. Byte Magazine, 19(3):120– 126, March 1994. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Lazzaro:1990:ODD

[Laz90] J. J. Lazzaro. Opening doors for the disabled: How personal computers offer disabled users professional opportunities. Byte Magazine, 15(8):258–260, 262–264, 266, 268, August 1990. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Lazzaro:1993:CDS

[Laz93] Joseph J. Lazzaro. Computers for the disabled: Off-the-shelf products help you meet the needs of disabled workers. Byte Magazine, 18(7):59–60, 62, 64, June 1993. CODEN BYTEDJ.

ISSN 0360-5280 (print), 1082-7838 (electronic).

Linderholm:1990:NMB

[LB90] O. Linderholm and J. Bertolucci.

The new Macs on the block at last, lower prices — and a new design. Byte Magazine, 15(12): 146–150, 152, November 1990. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Lacey:1991:FES

[LB91] S. Lacey and R. Box. A fast, easy sort: A novel enhancement makes a bubble sort into one of the fastest sorting routines. Byte Magazine, 16(4):315– 316, 318, 320, April 1991. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Links:1994:IBO

[LDH94] Cees Links, Wim Diepstraten, and Vic Hayes. The IEEE brings order to the chaotic world of wireless LANs by agreeing upon a foundation protocol for media access. Byte Magazine, 19(5):99– ??, May 1994. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Ledler:1990:DDA

[Led90] David Ledler. Dealing with devices: Answers to readers’ questions about floppy disk drives, printers, and upgrades. Byte Magazine, 15(2):127–??, February 1990. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Lee:1992:IP

[Lee92] Sing H. Lee. Interconnecting the pieces. Byte Magazine, 17(9):??, September 1992. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Lent:1993:WNT

[Len93a] Anne Fischer Lent. Workgroups by the numbers: Ten essential tips and techniques for getting the most out of Windows for Workgroups. Byte Magazine, 18 (5):32–35, Spring 1993. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Lent:1993:WPG

[Len93b] Anne Fischer Lent. Workgroups partners: A guide to products that integrate seamlessly with Windows for Workgroups. Byte Magazine, 18 (5):37–??, Spring 1993. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Lennon:1994:LRP

[Len94a] A. Lennon. Lab report: 46 powerful portables: 486 processing power now comes in two portable-computer form factors. see our exclusive ranking to find the best notebook and subnotebook for your application. best Windows notebooks; new pointers; how we tested; essential features; best color notebooks; dual scan: An economical compromise?; best desktop replacement notebooks; mobile multimedia; best subnotebooks; SL-Enhanced CPU’s: Power misers, but rare; lithium promises longer battery life; honorable mentions; dubious achievements. Byte Magazine, 19(3):156–160, 164, 166, 169, 171–172, 174, 177–178, 180– 181, March 1994. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Lennon:1994:LRN

[Len94b] A. J. Lennon. Lab report:

31 No-Compromise portables: We choose the best highperformance notebooks and subnotebooks. we also rate the leading portable printers for speed and print quality. best High-Performance notebooks; Power-Book update; best 486 based subnotebooks; best protable printers; how we tested; honorable mentions.; dubious achievements; roll call. Byte Magazine, 19(10):174–177, 181, 183, 186, 188, 192–195, October 1994. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Lee:1990:SWR

[LHR90] Kai-Fu Lee, A. G. Hauptmann, and A. I. Rudnicky. The spoken word: Researchers at Carnegie Mellon report on voice interfaces for computers. Byte Magazine, 15(7):225–226, 228–232, July 1990. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Liebowitz:1993:RYO

[Lie93] Jay Liebowitz. Roll your own hybrids: Emerging technologies — such as neural networks and genetic algorithms — can add robustness to knowledge-based systems. stand-alone expert systems could go the same route as the dinosaurs. Byte Magazine, 18(8):113–??, July 1993. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Linderholm:1990:PCW

[Lin90a] O. Linderholm. Portable chips: When it comes to chips, small can mean powerful. Byte Magazine, 15(13):312–316, 318–319, December 1990. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Linderholm:1990:SSO

[Lin90b] O. Linderholm. Solbourne S4000 outguns SPARCstation 1+: The S4000 uses Solbourne’s own integrated, 64-bit CPU. Byte Magazine, 15(13):144, 146, December 1990. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Linderholm:1991:AIL

[Lin91a] O. Linderholm. Apple introduces low-cost laser quality with style. Byte Magazine, 16(5):48, 50, May 1991. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Linderholm:1991:GER

[Lin91b] O. Linderholm. A graphical environment that runs where Windows can’t: GeoWorks ensemble runs on the smaller, slower PCs in your office. Byte Magazine, 16(4):276–277, April 1991. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Linderholm:1991:ZIT

[Lin91c]  O. Linderholm.  Z-486/25E:

the i486 and TIGA video make a powerful personal workstation. Byte Magazine, 16(5):48– ??, May 1991. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Linderholm:1991:MM

[Lin91d] Owen Linderholm. Mind melding. Byte Magazine, 16(11):41– ??, ?? 1991. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Lindstrom:1993:VMT

[Lin93] Bob Lindstrom. Video machine: True desktop video: Fast electronic of Munich, Germany, presents the Video Machine, a software/hardware combination that links with VCRs and video monitors to turn a PC into a Windows-based desktop video editing system. Byte Magazine, 18(10):153–??, September 1993. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Lindstrom:1994:ONP

[Lin94] Bob Lindstrom. Opening night for Premiere 3.0: For multimedia audio and video, Adobe Premiere 3.0 offers impressive editing capabilities — if you’ve got the hardware to handle it. Lindstrom evaluates the latest version of Premiere as a professional tool for video development. Byte Magazine, 19(1):179– 180, 182, January 1994. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Lion:1990:DSD

[Lio90a]  K. Lion.  DAT’s a solu-

tion: Digital-audiotape technology comes of age. Byte Magazine, 15(12):323–324, 326, 328, November 1990. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Lion:1990:DS

[Lio90b] Karina Lion. DAT’s a solution. Byte Magazine, 15(??):323–328, November 1990. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Lippincott:1990:BHL

[Lip90a] Rob Lippincott. Beyond hype: Lotus’s multimedia point man tackles the question: How do we get there from here? Byte Magazine, 15(2):215–218, February 1990. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Lippman:1990:HSD

[Lip90b] A. Lippman. HDTV sparks a digital revolution: In the 1990s, the shift will be to high-definition and digital pictures. Byte Magazine, 15(13):297–301, 303, 305, December 1990. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Loveria:1990:MDA

[LK90] Greg Loveria and Don Kinstler. Multimedia: DVI arrives. Byte Magazine, 15(11):105–108, Fall 1990. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Lahti:1990:SDF

[LM90] Walter Lahti and Dean McCarron. Store data in a flash: The flash-memory disk offers a fast and rugged replacement for both hard and floppy disk drives. Byte Magazine, 15(12):311–313, 315, 317, 318, November 1990. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

LaRiviere:1993:HDT

[LM93] Kathleen LaRiviere and Stan Miastkowski. How to deal with taxing questions: Tax-preparation software for DOS, the Mac, and Windows. Byte Magazine, 18 (3):165–??, March 1993. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Locke:1991:DSD

[Loc91] C. Locke. The dark side of DIP:

Christopher Locke discusses the complexities of indexing, an essential element in retrieving imaged documents. Byte Magazine, 16(4):193–194, 196, 200, 202, 204, April 1991. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Locke:1992:OMK

[Loc92] C. Locke. Overview: Making knowledge pay: How to mine the knowledge scattered throughout an organization. Byte Magazine, 17(6):244–252, June 1992. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Loeb:1990:TDA

[Loe90] L. H. Loeb. Two different approaches to Mac portability: The Outbound and Dynamac take opposing approaches. Byte Magazine, 15(9):169, 172, 174, September 1990. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Loveria:1990:PPH

[Lov90a] G. Loveria. Pumping pixels: Hercules flexes its muscles in 24bit color: Hercules’s economical graphics card offers 24-bit color display on standard VGA monitors. Byte Magazine, 15(10): 192, 194, October 1990. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Loveria:1990:WWV

[Lov90b] G. Loveria. Window wonderland: VideoLogic’s multimedia board turns VGA screens into video playgrounds. Byte Magazine, 15 (6):170, 172, 174, June 1990. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Loveria:1991:HQI

[Lov91a] G. Loveria. High-Quality image editing develops on the PC: Image-editing software turns PCs into electronic darkrooms. Byte Magazine, 16(5):266–268, May 1991. CODEN BYTEDJ.

ISSN 0360-5280 (print), 1082-7838 (electronic).

Loveria:1991:LCD

[Lov91b] G. Loveria. Low-Cost 3-D animation materializes for PC users: 3D studio brings affordable animation to the PC. Byte Magazine, 16(4):259–262, April 1991. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Loveria:1991:PRR

[Lov91c] G. Loveria. Photo-Realism reigns when AutoShade meets Render-Man: The RenderMan extension in new AutoShade brings photorealism to AutoCAD users. Byte Magazine, 16(7):219–220, 222, 0224, July 1991. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Loveria:1991:TLC

[Lov91d] G. Loveria. TARGA+ lowers cost of High-End graphics: Truevision’s new 32-bit TARGA+ board makes raster graphics more affordable. Byte Magazine, 16(1):204–206, January 1991. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Loveria:1992:BDA

[Lov92a] Greg Loveria. 24-bit display adapter roundup. Byte Magazine, 17(11):84–??, November 1992. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Loveria:1992:CUN

[Lov92b] Greg Loveria. Compaq unveils a new network printer: Compaq enters the printer market with the powerful Pagemarq line. Byte Magazine, 17(12):249–??, November 1992. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Loveria:1992:TTZ

[Lov92c] Greg Loveria. True type A to Z: Why TrueType is a key part of Windows 3.1. Byte Magazine, 17(14):129–??, December 1992. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Loveria:1993:FAW

[Lov93a] Greg Loveria. The fine art of Windows printing: How to speed up printing from Windows. Byte Magazine, 18(5):52– ??, Spring 1993. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Loveria:1993:SFM

[Lov93b] Greg Loveria. Solutions focus: Making the MPC upgrade: Some of the best options available for bringing multimedia to the PC. Byte Magazine, 18(6):176– ??, May 1993. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Loveria:1994:EWG

[Lov94] Greg Loveria. Easing Windows’ graphics bottleneck: Greg Loveria tests new 64-bit graphics adapters from five companies. all are blazingly fast with 24-bit graphics; one is surprisingly affordable. Byte Magazine, 19(7):133–??, July 1994. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Lubeck:1991:HEP

[LS91a] J. H. Lubeck and B. D. Schatzman. High-End printing on a Low-End budget: Champagne output on a beer budget. Byte Magazine, 16(9):315–316, 318– 320, September 1991. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Lubeck:1991:NWM

[LS91b] J. H. Lubeck and B. D. Schatzman. Networking Windows: Making Windows 3.0 work on a network. Byte Magazine, 16 (3):299–300, 302, 304, 306–307, March 1991. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Lu:1992:OOE

[Lu92] Cary Lu. Overview: Objects for end users: Document-oriented computing is the logical culmination of the object-oriented revolution. Byte Magazine, 17(14): 142–??, December 1992. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Lu:1993:IHS

[Lu93a] Cary Lu. Is ITV here to stay: Interactive TV’s survival and prosperity are in question. Byte Magazine, 18(2):139–??, February 1993. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Lu:1993:PIE

[Lu93b] Cary Lu. Publish it electronically: Electronic publishing lets you build enterprise-wide knowledge bases. Byte Magazine, 18 (10):94–97, 100, 102, 106, 108– 109, September 1993. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Lussmyer:1990:WRN

[Lus90] J. Lussmyer. Windows rides a new wave: With NewWave, Hewlett–Packard expands Windows, but it’s not easy. Byte Magazine, 15(4):171–172, 174, 176, April 1990. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

MacDonald:1991:SUC

[Mac91] L. MacDonald. Smart use of color in displays. Byte Magazine, 16(13):84IS–35–36, 84IS– 40, 84IS–42, 84IS–44, 84IS–46, December 1991. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Mallett:1991:LAR

[Mal91a] M. Mallett. A look at remote procedure calls. Byte Magazine, 16(5):309–312, 314, 316, 317, 384–385, May 1991. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic). See [Mal91b].

Mallett:1991:NIC

[Mal91b] M. Mallett. Networkwide interprocess communications. Byte Magazine, 16(5):309–??, May 1991. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Malloy:1993:CBT

[Mal93] Rich Malloy. Comdex: Bigger than ever: Picking the best products at the show was even harder than usual. Byte Magazine, 18 (2):41–??, February 1993. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Mandel:1990:WAM

[Man90] R. Mandel. The world according to micros: Need to know some exotic geographic facts? try one of these 24 world atlas packages. Byte Magazine, 15(7):256–260, 262, 264, 267, July 1990. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Marshall:1990:CRC

[Mar90a] T. Marshall. A calculating RISC:

Coprocessors based on RISC engines will soon deliver supercomputer performance to your desktop. Byte Magazine, 15(5): 251–254, 256, May 1990. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Marshall:1990:CCC

[Mar90b] Trevor Marshall. Creating custom chips: EPLDs are fast becoming the device of choice for fast turnaround or rapidly changing design tasks. Byte Magazine, 15(1):271–280, January 1990. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Marshall:1992:FTN

[Mar92a] Trevor Marshall. Fast transit: New buses dramatically increase speed and will be showing up in systems soon. Byte Magazine, 17 (10):122–??, October 1992. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Marshall:1992:PFS

[Mar92b] Trevor Marshall. Portable fax software rated for DOS and Windows. Byte Magazine, 17(13): ??, ?? 1992. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Marshall:1992:WPF

[Mar92c] Trevor Marshall. WorldPort Palmtop Fax/data modem: US-Robotics designs modems for hand-helds. Byte Magazine, 17 (13):??, ?? 1992. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Mashyna:1990:CCD

[Mas90a] M. Mashyna. C compilers have different strengths: Apple and Symantec bring object-oriented C compilers to the Mac. Byte Magazine, 15(4):179–180, 182, 184, April 1990. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Mastrianni:1990:TTD

[Mas90b] Steve Mastrianni. Tales from the trenches: An OS/2 device-driver specialist talks shop. Byte Magazine, 15(12):127–??, November 1990. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Mashey:1991:BCB

[Mas91a] J. R. Mashey. 64-bit computing:

64-bit micros may be coming to desktops near you. Byte Magazine, 16(9):135–138, 140, 142, September 1991. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Mashey:1991:BC

[Mas91b] John R. Mashey. 64-bit computing. Byte Magazine, 16(??): 135–142, September 1991. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Mastrianni:1991:DDP

[Mas91c] S. J. Mastrianni. OS/2 device drivers: A practitioner’s guide to OS/2 driver development. Byte Magazine, 16(7):241–242, 244, 246, 248, 348–350, July 1991. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Mastrianni:1992:PTA

[Mas92] Steve Mastrianni. OS/2 2.0 programming tools arrive — finally: OS/2 programmers now have a wide choice of development tools. Byte Magazine, 17(10):277–??, October 1992. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Mastrianni:1993:BCD

[Mas93] Steve Mastrianni. Beyond DOS: Confessions of a DDK developer: IBM’s OS/2 DDK is a good start. Byte Magazine, 18(8):203– ??, July 1993. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Mathisen:1994:NNN

[Mat94a] Terje Mathisen. Novell’s newest DOS: Is Novell DOS 7 a better DOS than Microsoft’s? Byte Magazine, 19(6):241–??, June 1994. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Mathisen:1994:PSU

[Mat94b] Terje Mathisen. Pentium secrets: Undocumented features of the Intel Pentium can give you all the information you need to optimize Pentium code. Byte Magazine, 19(7):191–??, July 1994. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

May:1994:KYW

[May94a] Thornton May. Know your workflow tools: The four basic classes of work-flow tools vary widely from one another. Byte Magazine, 19(7):103–??, July 1994. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

May:1994:CSW

[May94b] Thornton A. May. Commentary:

Shakespearean wisdom: Richard III and information technologists have much in common, but there are differences. Byte Magazine, 19(1):312–??, January 1994. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Mezick:1993:PCC

[MB93] D. Mezick and D. Bricklin. Pen computing catches on. Byte Magazine, 18(11):105–106, 108, 110–112, October 1993. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

McAvinney:1990:TGL

[McA90] P. McAvinney. Telltale gestures:

A look at devices that will change how you manipulate 3-D design applications. Byte Magazine, 15 (7):237–240, July 1990. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

McGinnis:1990:MEW

[McG90a] B. McGinnis. Of monitors and emissions: What’s being done about magnetic fields from monitors? Byte Magazine, 15(9): 445–446, 448, 451–452, September 1990. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

McGrath:1990:WOC

[McG90b] W. T. McGrath. Who owns the copyrights: Who owns the copyrights on independently developed programs? an attorney discusses recent developments. Byte Magazine, 15(4):269–271, April 1990. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

McNutt:1994:SBN

[McN94] Dinah McNutt. SCSI and beyond: New standards clarify the future direction and higherspeed capabilities of this longestablished workhorse interface for PC, Macintosh, and Unix platforms. Byte Magazine, 19 (8):111–??, August 1994. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Mello:1993:FCB

[Mel93a] John P. Mello, Jr. Future communications: Beyond file and print sharing, networks are evolving into the preferred medium for all sorts of communications — voice, text, graphics, and video. Byte Magazine, 18(9):94–??, August 1993. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Mello:1993:PTP

[Mel93b] John P. Mello, Jr. Printers in transition: People want faster, easier-to-use printers that produce higher-quality images. those demands are starting to be met. Byte Magazine, 18(13): 94–??, December 1993. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Mitchell:1990:CGP

[MG90] R. Mitchell and R. Grehan.

Cheetah’s golden performer: Cheetah’s new 33-MHz 486 machine is one fast cat. Byte Magazine, 15(6):132–133, June 1990. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Miastkowski:1991:GCA

[MG91] S. Miastkowski and R. Grehan. The 386 gets a competitor: AMD’s innovative 386 “clone” chip sparks controversy. Byte Magazine, 16(3):44–45, March 1991. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Manson:1991:DDM

[MH91] C. Manson and J. S. Haugdahl.

Dynamic and distributed: Managing distributed systems in general, and doing it from a centralized location in particular, are major problem areas. Byte Magazine, 16(3):167–168, 170– 172, March 1991. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Miastkowski:1990:AMM

[Mia90a] S. Miastkowski. The ALR MPS:

Modular Micro Channel: ALR gambles that it can take a bite out of the True Blue market. Byte Magazine, 15(12):162–164, November 1990. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Miastkowski:1990:EMT

[Mia90b] S. Miastkowski. Eccentric mouse tames complicated GUIs: Cal-Comp’s mouse-digitizer hybrid simplifies PC and Mac use. Byte Magazine, 15(8):204–??, August 1990. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Miastkowski:1990:MFG

[Mia90c] S. Miastkowski. A ‘more filling’ generation of tape backup. Byte Magazine, 15(13):235–237, December 1990. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Miastkowski:1990:NFD

[Mia90d] S. Miastkowski. New floppy drive puts 20-MB disk in your pocket: Q/Cor’s new floppy disk drive leads the 20-megabyte vanguard. Byte Magazine, 15(9):188–190, September 1990. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Miastkowski:1990:PHD

[Mia90e] S. Miastkowski. Pricey hard disk drive portability: The Disctec 20 provides floppy disk convenience with hard disk storage in a very small package. Byte Magazine, 15(12):266–??, November 1990. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Miastkowski:1990:TTT

[Mia90f] S. Miastkowski. Thoroughly totable Tandy: The Tandy 1100 FD offers all the necessities — and then some. Byte Magazine, 15(6):190, 192, June 1990. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Miastkowski:1990:WSA

[Mia90g] S. Miastkowski. Windows shopping (authorware). Byte Magazine, 15(8):114–115, August 1990. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Miastkowski:1990:PGG

[Mia90h] Stan Miastkowski. PC GUIs go head to head. Byte Magazine, 15(11):82–87, Fall 1990. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Miastkowski:1990:TSL

[Mia90i] Stan Miastkowski. Time to switch: A look at five application switchers for DOS that let you keep multiple programs in RAM. Byte Magazine, 15(2): 185–188, February 1990. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Miastkowski:1991:WSM

[Mia91] S. Miastkowski. A whale of a system: Moby Brick is a loaded 486 system. Byte Magazine, 16 (8):39–40, August 1991. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Miastkowski:1992:KST

[Mia92a] S. Miastkowski. Keyboardless sense: A touchscreen computer that makes sense. Byte Magazine, 17(2):36–37, February 1992. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Miastkowski:1992:TAH

[Mia92b] Stan Miastkowski. Tag-Along hard copy: Seven portable printers offer near-laser-quality output in surprisingly small packages. Byte Magazine, 17(13):41– 45, ?? 1992. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Miastkowski:1993:BWP

[Mia93a] Stan Miastkowski. Beyond word processing: word processors aren’t just for juggling text anymore. Byte Magazine, 18(5): 85–86, 88, 90, Spring 1993. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Miastkowski:1993:FBT

[Mia93b] Stan Miastkowski. Forging a business tool: Three fax software packages for Windows: Three next-generation fax software packages bring new abilities to communications. Byte Magazine, 18(2):209–210, 212, February 1993. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Miastkowski:1994:CFF

[Mia94a] S. Miastkowski. Collecting facts from fax. Byte Magazine, 19 (12):153–154, 156, 158, December 1994. CODEN BYTEDJ.

ISSN 0360-5280 (print), 1082-7838 (electronic).

Miastkowski:1994:WPH

[Mia94b] Stan Miastkowski. WinFax Pro hits the network: Delrina’s Win-Fax Pro for Networks lets workgroups share fax modems. Byte Magazine, 19(2):141–145, February 1994. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Michalski:1991:WDY

[Mic91] G. P. Michalski. The world of documents: You can’t choose a document imaging system without understanding the base technology. Byte Magazine, 16(4): 159–160, 162–164, 166–168, 170, April 1991. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Miller:1993:BDM

[Mil93] L. Chris Miller. Babelware for the desktop: Many machinetranslation systems are available on workstations and personal computers. Byte Magazine, 18 (1):177–178, 80, 182–183, January 1993. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Miller:1994:TTT

[Mil94a] Chris Miller. Transborder tips and traps: A look at the many technical issues and problems involved in moving software. Byte Magazine, 19(6):93– ??, June 1994. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Millikin:1994:DBD

[Mil94b] Michael D. Millikin. DCE: Building the distributed future: The distributed computing environment standard provides the services to make distributed computing manageable, multiplatform, and transparent to users. Byte Magazine, 19(6):125–??, June 1994. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Minasi:1990:DHM

[Min90a] Mark J. Minasi. Digging into HPFS: The more you examine the High Performance File System, the better it looks. Byte Magazine, 15(5):109–110, 112, May 1990. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Minasi:1990:FLA

[Min90b] Mark J. Minasi. A first look at HPFS: Performance file system allows bigger, faster, and safer hard disk drives. Byte Magazine, 15(1):137–??, January 1990. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Minasi:1990:FAL

[Min90c] Mark J. Minasi. Free at last!: Thanks to OS/2’s new memory architecture, you no longer have to squeeze code into 64K-byte segments. Byte Magazine, 15(7): 97–??, July 1990. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Minasi:1990:LDM

[Min90d] Mark J. Minasi. A letter from a dissenter: Mark defends OS/2 and Presentation Manager against a reader’s criticisms. Byte Magazine, 15(2):121–124, February 1990. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Minasi:1990:LLL

[Min90e] Mark J. Minasi. Living with OS/2 1.2: Life with OS/2 1.2 is a lot like life with version 1.1, with some welcome changes. Byte Magazine, 15 (4):85–??, April 1990. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Minasi:1990:MLM

[Min90f] Mark J. Minasi. Managing LAN Manager 2.0: The latest version of OS/2 LAN Manager offers integrity, security, and somewhat easier administration. Byte Magazine, 15(8):95–96, August 1990. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Minasi:1990:MSS

[Min90g] Mark J. Minasi. Money, SQL, and spreadsheets. Byte Magazine, 15(11):37–??, Fall 1990. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Minasi:1990:MDL

[Min90h] Mark J. Minasi. OS/2 marries Desqview: The latest version of OS/2 lets you run multiple DOS sessions, à la Desqview. Byte Magazine, 15(6):103– ??, June 1990. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Minasi:1990:HHC

[Min90i] Mark J. Minasi. To HPFS or not to HPFS: Can OS/2’s HPFS and the DOS file allocation table live together on the same disk? Byte Magazine, 15(3):101– ??, March 1990. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Mindell:1990:IDU

[Min90j] D. A. Mindell. Images from the deep: Using a fiber-optic network and oceanographic vehicles, Woods Hole scientists are probing the wonders of the ocean. Byte Magazine, 15(6):256–260, June 1990. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Minasi:1991:CMM

[Min91a] Mark J. Minasi. The case of the missing memory: DOS 5.0 meets Windows 3.0. Byte Magazine, 16 (10):273–??, October 1991. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Minasi:1991:ERR

[Min91b] Mark J. Minasi. Embarrassment of riches: A report from the future: living with OS/2 2.0 and Windows 3.0. Byte Magazine, 16 (1):95–96, 98, January 1991. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Minasi:1991:LFW

[Min91c] Mark J. Minasi. LaserJets, fonts, and Windows: Getting fancy fonts from Windows is tricky but doable. Byte Magazine, 16 (7):271–273, July 1991. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Minasi:1991:WTT

[Min91d] Mark J. Minasi. Windows tips and tricks: Answers to some frequently asked Windows 3.0 questions. Byte Magazine, 16(5): 328, 330–332, May 1991. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Minasi:1992:ATS

[Min92a] Mark Minasi. The OS/2 alternative: There is some kind of 32-bit OS in your future. Byte Magazine, 17(11):55–??, November 1992. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Minasi:1992:ACO

[Min92b] Mark Minasi. OS/2 at the crossroads: Overcoming market resistance. Byte Magazine, 17 (10):179–??, October 1992. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Minasi:1992:CCS

[Min92c] Mark J. Minasi. Cures for C sickness: The “doctor” prescribes three Windows programming toolkits. Byte Magazine, 17 (2):259–??, February 1992. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Minasi:1992:EAP

[Min92d] Mark J. Minasi. Exorcising the A20 poltergeist: Here’s what to do when keystrokes seem to appear on-screen at random in your DOS and Windows applications. Byte Magazine, 17(8):293– ??, August 1992. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Minasi:1993:BEA

[Min93a] Mark J. Minasi. Beyond DOS: OS/2 extended attributes: How to fix problems with OS/2 extended attributes. Byte Magazine, 18(12):313–??, November 1993. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Minasi:1993:NIS

[Min93b] Mark J. Minasi. A new OS/2: IBM’s ServicePak and the Professional Developer’s Kit CD-ROM are dissected. Byte Magazine, 18(2):237–??, February 1993. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Mitchell:1990:SFB

[Mit90a] R. Mitchell. Small footprint, big impression: Emerald Computers’ LANstation — small size, big price. Byte Magazine, 15(6):186– ??, June 1990. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Mitchell:1990:GSZ

[Mit90b] Robert Mitchell. A good sport:

The Zenith MinisPort is appealing, but not perfect. Byte Magazine, 15(2):173–177, February 1990. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Mitchell:1991:CSL

[Mit91a] R. Mitchell. The Compaq SLT:

A laptop fit for the desktop: The BYTE Lab tests how well the Compaq SLT laptop performs with its new 386SX engine and other enhancements. Byte Magazine, 16(1):210–211, January 1991. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Mitchell:1991:IAE

[Mit91b] R. Mitchell. IBM and AT&T enter the fray of 386SX notebook computers: The BYTE Lab tests and compares IBM’s PS/2 model L40 SX and AT&T’s Safari NSX/20. Byte Magazine, 16 (8):252–254, August 1991. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Mitchell:1991:LRC

[Mit91c] R. Mitchell. LAN remote-control software: Better than being there: The BYTE Lab examines eight communications programs that let you use workstations on a LAN via remote control. Byte Magazine, 16(1):186– 189, 192, 194, 196, January 1991. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Mitchell:1993:BHT

[Mit93a] David H. Mitchell. Being here and there. Byte Magazine, 18(3): ??, March 1993. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Mitchell:1993:APG

[Mit93b] Rob Mitchell. AST’s PowerExec goes modular: Upgradability meets portability in AST’s new PowerExec. Byte Magazine, 18 (1):209–??, January 1993. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Mysore:1994:PAM

[MM94] C. Mysore and J. McDonough. 19 Pentiums at 90 MHz. Byte Magazine, 19(12):192–195, 198, 201, 203–204, 206–209, December 1994. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Moore:1990:MXW

[Moo90]  David Moore.  The migra-

tion of the X Window System. Byte Magazine, 15(11):183–185, Fall 1990. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Mossberg:1993:CTJ

[Mos93] Walter S. Mossberg. Commentary: They just don’t get it: Needless complexity has alienated many businesspeople from the personal computer revolution. Byte Magazine, 18(9):268– ??, August 1993. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Martin:1990:SHP

[MPW^{+}90] G. Martin, J. Pittman, K. Wittenburg, R. Cohen, and T. Parish. Sign here, please: Interactive tablets enable free-form communication between you and your computer. Byte Magazine, 15 (7):243–244, 246–248, 250–251, July 1990. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Meisel:1993:TYC

[MR93] W. S. Meisel and D. B. Roe. Talk to your computer. Byte Magazine, 18(11):113–116, 118, 118, October 1993. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Miastkowski:1993:TRW

[MS93] Stan Miastkowski and Marc Schnapp. Two roads to Windows databases: Paradox and FoxPro migrate to Windows: Borland’s Paradox for Windows wraps its powerful features in an object-oriented package. Microsoft’s FoxPro 2.5 for Windows maintains its solid Xbase underpinning and is built for speed. Byte Magazine, 18(7):136–138, June 1993. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Mullins:1994:GDS

[Mul94] Craig S. Mullins. The great debate: To store complex data objects, relational database systems must perform expensive and CPU-intensive transformations because of their simplified storage capabilities. object database systems and object-oriented programming languages make it possible to store and access such data in an easy, efficient manner. Byte Magazine, 19(4):85– ??, April 1994. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Murray:1993:DGD

[Mur93] Philip Murray. Documentation goes digital: A confusing array of tools await those who want to publish reference material electronically, but the benefits might be worth the trouble. Byte Magazine, 18(10): 121–??, September 1993. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Manber:1992:APM

[MW92a] Udi Manber and Sun Wu. Approximate pattern matching. Byte Magazine, 17(??):??, ?? 1992. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Manber:1992:SAR

[MW92b] Udi Manber and Sun Wu. Some assembly required. approximate pattern matching: Agrep’s algorithms let you perform text searches using an approximate pattern. Byte Magazine, 17(12): 281–??, November 1992. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Mello:1993:WMC

[MW93a] J. P. Mello and P. Wayner.

Wireless mobile communications. Byte Magazine, 18(??):147–154, ?? 1993. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Mello:1993:OWM

[MW93b] John P. Mello, Jr. and Peter Wayner. Overview: Wireless mobile communications: The ability to communicate anytime, from anywhere, is almost here. Byte Magazine, 18(2):146–??, February 1993. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

McParland:1994:RFR

[MW94] Michael McParland and Ethan Wilansky. Running the Frame-Relay race: Frame relay is an affordable way to get your LANs talking to each other. Byte Magazine, 19(11):297–??, November 1994. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Miastkowski:1991:WA

[MYT91] Stan Miastkowski, Tom Yager, and Tom Thompson. Windowing: Not by DOS alone. Byte Magazine, 16(11):172–??, ?? 1991. CODEN BYTEDJ.

ISSN 0360-5280 (print), 1082-7838 (electronic).

Nadeau:1990:APP

[Nad90a] M. Nadeau. ALR pumps up PowerFlex (PowerFlex 20CSX Model 110). Byte Magazine, 15(12): 164–??, November 1990. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Nadeau:1990:CNU

[Nad90b] M. Nadeau. Compaq Notebook ups the ante. Byte Magazine, 15(12):140–142, November 1990. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Nadeau:1990:FKG

[Nad90c] M. Nadeau. The fast keep getting faster: New 33-MHz 486 machines from AST and ALR are fast and upgradable. Byte Magazine, 15(5):131–133, May 1990. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Nadeau:1990:NPS

[Nad90d] M. Nadeau. The NEC ProSpeed SX/20: Take it and leave it: This 13-pound laptop can double as a powerful desktop system. Byte Magazine, 15(9):128–129, September 1990. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Nadeau:1990:NNV

[Nad90e] M. Nadeau. Notebook newcomer: The VH-286 from Airis offers features, power, and low cost. Byte Magazine, 15(7):133– 135, July 1990. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Nadeau:1990:TLN

[Nad90f] M. Nadeau. TravelMate 2000 lives up to its name: Texas Instruments puts AT-class power in a 4-pound package. Byte Magazine, 15(12):262, 264, November 1990. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Nadeau:1991:PCN

[Nad91] Michael Nadeau. Portable computing: Notebooks coming of age. Byte Magazine, 16(11):194– 200, 202, 204, 206, ?? 1991. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Nadeau:1992:ZPP

[Nad92] Michael Nadeau. Zeos Pocket PC one-ups Poqet PC. Byte Magazine, 17(13):??, ?? 1992. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Nadeau:1993:BAB

[Nad93] Michael Nadeau. The 1992 BYTE awards: BYTE editors pick the best products of the past year. Byte Magazine, 18(1):116– ??, January 1993. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Nadeau:1994:BAB

[Nad94a] Michael Nadeau. 1993 BYTE awards: The best products of 1993 provide a window to the trends of 1994. Byte Magazine, 19(1):46–??, January 1994. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Nadeau:1994:RCA

[Nad94b] Michael Nadeau. Remote connections: Access performance, and data integrity are some of the issues of connecting remote clients. Byte Magazine, 19 (6):197–??, June 1994. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Nagy:1992:SAR

[Nag92] Randall A. Nagy. Some assembly required: Writing a Windows DLL: Here’s a simple DLLbased debugging aid for Windows. Byte Magazine, 17(14): 247–??, December 1992. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Nagy:1993:HIE

[Nag93] Randall A. Nagy. Handling input events using C++: Use the event class to handle keyboard and mouse input across platforms. Byte Magazine, 18(2): 231–??, February 1993. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Nakakoji:1994:CCB

[Nak94] Kumiyo Nakakoji. Crossing the cultural boundary: Cultural factors impinge on software design. Byte Magazine, 19(6):107– ??, June 1994. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Nance:1990:OSF

[Nan90a] B. Nance. One-Size-Fits-All code with Lattice C: A royaltyfree DOS extender is standard with Lattice’s new C compiler. Byte Magazine, 15(12):245–246, November 1990. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Nance:1990:SNL

[Nan90b] B. Nance. Speaking OS/2’s native language: Object-1 speaks to OS/2’s presentation manager in object-oriented terms. Byte Magazine, 15(9):206, 208, 210, September 1990. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Nance:1991:DBP

[Nan91a] B. Nance. A disk-based print spooler: Free timesaving utilities for DOS, Mac, and Unix systems. Byte Magazine, 16(7):267– ??, July 1991. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Nance:1991:MBB

[Nan91b] B. Nance. Managing big blue:

A look at IBM’s network management tools. Byte Magazine, 16(3):197–198, 200, 202, 204, March 1991. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Nance:1991:NNV

[Nan91c] B. Nance. New NetWare versions, options broaden appeal:

Dramatic improvements in connectivity and ease of installation. Byte Magazine, 16(9):293– 294, 296, September 1991. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Nance:1991:AAC

[Nan91d] Barry Nance. The AIX alternative: Choosing a Unix-based architecture. Byte Magazine, 16(5): 324, 326, 328, May 1991. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Nance:1991:AAP

[Nan91e] Barry Nance. The AIX alternative, part 2: The search for the right network-support software. Byte Magazine, 16(6):343– 344, 346, 348, June 1991. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Nance:1991:BAN

[Nan91f] Barry Nance. The black art of networking: There’s nothing simple about managing today’s LANs. Byte Magazine, 16(13): 303–??, December 1991. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Nance:1991:FST

[Nan91g] Barry Nance. The future of software technology. Byte Magazine, 16(11):69–??, ?? 1991. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Nance:1991:GMN

[Nan91h] Barry Nance. Getting the most from NetBIOS: Barry offers sage advice — and a free utility — for NetBIOS programmers. Byte Magazine, 16(10):269–??, October 1991. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Nance:1991:PRP

[Nan91i] Barry Nance. In praise of remote procedure calls: RPCs harness the power of the client/server architecture. Byte Magazine, 16 (3):338–340, March 1991. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Nance:1991:ITB

[Nan91j] Barry Nance. Interoperability today: Barry Nance looks at how far we’ve come and how far we have to go. Byte Magazine, 16(12):187–??, November 1991. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Nance:1991:LTR

[Nan91k] Barry Nance. LAN Tune-Up: Revitalizing the office network. Byte Magazine, 16(8):287– ??, August 1991. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Nance:1991:NTW

[Nan91l] Barry Nance. NetWare troubles:

Whom do you call when Net-Ware acts up? with the right tools, you can do the job yourself. Byte Magazine, 16(1):119– 120, 122, January 1991. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Nance:1992:NDD

[Nan92a] Barry Nance. 486 notebooks double as desktop: Power to go. Byte Magazine, 17(9):249–250, 256, 258, September 1992. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Nance:1992:HOW

[Nan92b] Barry Nance. How OLE works: Some sobering experiences with OLE. Byte Magazine, 17(11): 45–??, November 1992. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Nance:1992:LAM

[Nan92c] Barry Nance. LAN analyzers move to AI: AI is redefining the role of LAN analyzers. Byte Magazine, 17(3):287–??, March 1992. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Nance:1992:LEL

[Nan92d] Barry Nance. LANlord evicts LAN problems: Microcom’s high-level LAN manager uses OS/2 to manage DOS and Windows workstations. Byte Magazine, 17(12):265–??, November 1992. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Nance:1992:WNC

[Nan92e] Barry Nance. Windows NT and OS/2 compared: Windows NT and OS/2 2.0 have a lot in common. Byte Magazine, 17(6):359– ??, June 1992. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Nance:1993:BIU

[Nan93a] Barry Nance. Beyond DOS, IBM unleashes a new OS/2: IBM delivers a full-featured successor to OS/2 2.0. Byte Magazine, 18(7):193–??, June 1993. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Nance:1993:CSB

[Nan93b] Barry Nance. Compaq stakes out both ends of the server spectrum: Compaq’s new high-end Systempro/XL and low-end ProSignia servers. Byte Magazine, 18 (3):169–??, March 1993. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Nance:1993:GLB

[Nan93c] Barry Nance. A giant leap to OS/2 2.1: OS/2 2.1 brings IBM’s 32-bit operating system forward, beyond the reach of current desktop operating systems. bulletproof multimasking and support for Windows 3.1 applications are among the improvements. Byte Magazine, 18(10): 145–??, September 1993. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Nance:1993:LVI

[Nan93d] Barry Nance. LANtastic 5.0 vs. invisible LAN 3.4: Upgraded versions of peer-LAN operatingsystem staples from Invisible Software and Artisoft offer improved performance, better Windows integration, and options for connectivity. testing speed and ease-of-use features set the two apart. Byte Magazine, 18(10): 163–??, September 1993. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Nance:1993:SW

[Nan93e] Barry Nance. Serving up the works. Byte Magazine, 18(5): 105–??, Spring 1993. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Nance:1993:SAR

[Nan93f] Barry Nance. Some assembly required: OS/2’s system object model: The OS/2 2.0 system object model offers a languageneutral approach to objectoriented programming. Byte Magazine, 18(1):261–??, January 1993. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Nance:1993:STI

[Nan93g] Barry Nance. Stacking up TCP/ IP for Windows: Linking Windows clients with Unix hosts via TCP/IP stacks from Beame and Whiteside, Frontier Technologies, and NetManage. Byte Magazine, 18(2):215–??, February 1993. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Nance:1993:WOL

[Nan93h] Barry Nance. WordPerfect Office 4.0: The latest version of WordPerfect’s groupware product supports multiple operating systems. Byte Magazine, 18(8): 22–??, July 1993. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Nance:1994:BIN

[Nan94a] Barry Nance. Beyond DOS: IPX and NetBIOS for OS/2: Overthe-wire message passing with NetBIOS and IPX in an OS/2 environment. Byte Magazine, 19(5):201–??, May 1994. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Nance:1994:EEP

[Nan94b] Barry Nance. Easier Ethernet:

Plug Tut Systems’ US$99 Silver Streak adapter onto your Ethernet card, and you can string inexpensive, easy-to-install silversatin phone cord instead of thin Ethernet. the Silver Streak allows full 10-Mbps Ethernet speed over the UTP phone cord. Byte Magazine, 19(4):185– ??, April 1994. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Nance:1994:WBL

[Nan94c] Barry Nance. Windows becomes LAN friendly. Byte Magazine, 19(11):207–??, November 1994. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Nance:1994:PZD

[Nan94d] Barry Nance. Without peer:

Zenith Data Systems’ new US$999 Z-Stor Personal Server is a dedicated file server that makes sense for small workgroups. bundled with Novell’s Personal Net ware, it almost installs itself on Ethernet LANs. Byte Magazine, 19(5):133–??, May 1994. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Nelson:1991:HCT

[Nel91a]  M. D. Nelson.  Hot col-

ors: Thermal-transfer and dyediffusion printing expand your desktop color printing options. Byte Magazine, 16(10):177–178, 180, 182, October 1991. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Nelson:1991:SNF

[Nel91b] V. P. Nelson. Safety in numbers: Fault-tolerant computing can help you increase the useful lifetime of your system. Byte Magazine, 16(8):175–178, 180, 182–184, August 1991. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

New:1993:DCP

[New93] What’s New. The Dauphin 5500 Color Pentop flips its display, the pocket faxxer sends paperless faxes, ParaSet helps you develop and maintain software, and more. Byte Magazine, 18(4):62– ??, April 1993. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Niezgoda:1993:SAR

[NHW93] Steve Niezgoda, Lloyd Holt, and Derek Wojciech. Some assembly required: NT’s structured exception handling: The reality of structured exception handling in Windows NT may not live up to its promise. Byte Magazine, 18(12):317–??, November 1993. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Nicholls:1990:LAG

[Nic90] Bill Nicholls. Looking at the graphical user interface. Byte Magazine, 15(11):161–162, 164, 166, Fall 1990. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Nicholson:1993:SGT

[Nic93] Andy Nicholson. Software gluttony: It’s time for programmers to rein in today’s bloated, resource-hogging applications. Byte Magazine, 18(2):286–??, February 1993. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Niezgoda:1994:CUM

[Nie94] Steve Niezgoda. Charting the uncharted: Market dynamics may force you to port your Unix applications to Windows NT. here’s a guide to translating calls from one operating system to the other. Byte Magazine, 19 (10):203–??, October 1994. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Nova:1994:BCR

[Nov94a] Craig Nova. Books and CD-ROMs: The joy of the Internet: A look at 10 guides to the Internet and a CD-ROM on morphing. Byte Magazine, 19 (11):61–??, November 1994. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Nova:1994:CST

[Nov94b] Craig Nova. Commentary: Slouching toward the Internet: A world of information can seem just out of reach. Byte Magazine, 19(10):282–??, October 1994. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Nance:1992:SCA

[NTS92a] Barry Nance, Tom Thompson, and Ben Smith. Software corner: Audit your LAN: Easy LAN inventory, a crash disk for System 7.0, and face-saving Unix utilities. Byte Magazine, 17(14): 256–??, December 1992. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Nance:1992:SCS

[NTS92b] Barry Nance, Tom Thompson, and Ben Smith. Software corner: Software with curves: Bézier curves under Windows, synchronizing PowerBooks, and the language of bit maps. Byte Magazine, 17(9):305–??, September 1992. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Nance:1993:SCY

[NTS93a] Barry Nance, Tom Thompson, and Bem Smith. Software corner: Your own devices: DOS device drivers, hexadecimal editing in Unix, and monitoring System 7.0 memory partitions. Byte Magazine, 18(1):269–??, January 1993. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Nance:1993:SBE

[NTS93b] Barry Nance, Tom Thompson, and Ben Smith. A small browser with everything: A powerful DOS browser, a faster finder, and Perl-based recursive grep. Byte Magazine, 18(2):235–??, February 1993. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Nance:1993:SCA

[NTS93c] Barry Nance, Tom Thompson, and Ben Smith. Software corner: Automatic NetWare Log-Ins: Let your applications log in to NetWare; a Mac text editor; and a graphics file viewer. Byte Magazine, 18(3):208–??, March 1993. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Nance:1993:SCL

[NTS93d] Barry Nance, Tom Thompson, and Ben Smith. Software corner: LAN remote control: Remotecontrol programs for NetWare and Apple Talk. Byte Magazine, 18(4):210–??, April 1993. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Nudelman:1991:STN

[Nud91] M. Nudelman. Symmetry, thy name is Unix: Unix SVR4/MP: A new standard for multiprocessing with Unix? Byte Magazine, 16(6):245–246, 248, 250– 253, June 1991. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Nadeau:1992:BEG

[NW92] Michael Nadeau and John Wolfskill. BYTE’s essential guide to notebook PCs: The BYTE Lab compares 65 notebooks. Byte Magazine, 17(13):15–18, 20, 22– 23, 26, 28, 30, 32–40, ?? 1992. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Obermeier:1990:NSN

[Obe90] K. K. Obermeier. Natural selection: Natural-language front ends access databases without a formal query language. Byte Magazine, 15(8):217–218, 220, 222, August 1990. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Olson:1994:DMP

[Ols94] Douglas K. Olson. Developing for multiple platforms: Writing applications to run on multiple platforms is an art that software engineers are just beginning to master. Byte Magazine, 19 (2):91–??, February 1994. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

OMalley:1994:SPP

[O’M94] C. O’Malley. Simonizing the PDA (personal digital assistant). Byte Magazine, 19(12):145–146, 148, December 1994. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Oney:1991:WWN

[One91] Walter Oney. Why wait for NT and Win32: Writing 32bit Windows applications today. Byte Magazine, 16(13):307–??, December 1991. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Orr:1994:HBD

[Orr94] Joel Orr. Hoops 4.0: Beyond 3-D: The common API to most 2-D and 3-D libraries and hardware now offers a real-time mode and a powerful font engine. Byte Magazine, 19(2):125–??, February 1994. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Osher:1990:OAN

[Osh90] H. Osher. An open approach:

With its new Distributed Applications Architecture, Data General challenges IBM and DEC. Byte Magazine, 15(4):246–247, April 1990. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Osmundsen:1990:BIS

[Osm90] S. Osmundsen. Blueprints for the 1990s: IBM’s SAA versus DEC’s NAS — how do they compare? Byte Magazine, 15 (4):237–238, 240–242, 244–245, April 1990. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Othmer:1992:UHI

[Oth92] Konstantin Othmer. Under the hood: Inside QuickDraw: A detailed look inside Apple’s imaging engine. Byte Magazine, 17 (9):291–??, September 1992. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Pack:1994:CTT

[Pac94] Hughes Pack. Commentary: Teacher training is key: Too many schools are buying computers and stinting on the training. Byte Magazine, 19(11): 366–??, November 1994. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Parish:1990:CCS

[Par90] Tom Parish. Crystal clear storage: The holostore, a new mass storage device with supercomputer performance, could eliminate the I/O bottleneck. Byte Magazine, 15(12):283–288, November 1990. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Parshall:1992:MOE

[Par92] James H. Parshall. Mac OOP explained: Build a tool palette as you learn object-oriented programming techniques. Byte Magazine, 17(10):267–??, October 1992. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Pepper:1991:QPV

[Pep91] J. Pepper. Quattro Pro 3.0 vs.

Lotus 1-2-3 release 2.3: Two worthy GUI spreadsheets. Byte Magazine, 16(9):279–280, 282, September 1991. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Perratore:1992:FIN

[Per92] Ed Perratore. First impressions.

new systems … new IBM: Big Blue comes out swinging. Byte Magazine, 17(12):50–??, November 1992. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Perratore:1993:CMF

[Per93a] Ed Perratore. The case of the missing file. Byte Magazine, 18 (5):81–??, Spring 1993. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Perratore:1993:HEC

[Per93b] Ed Perratore. A higher end for Compaq notebooks. Byte Magazine, 18(1):??, January 1993. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Perratore:1993:NBC

[Per93c] Ed Perratore. The new breed of CD players. Byte Magazine, 18(2):??, February 1993. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Perratore:1994:LPP

[Per94a] Ed Perratore. Low-Priced Pentium PCs: General Pentium price wars have started. this review examines three Pentium systems, priced between US$4000 and US$4500, from ALR, Gateway, and Ambra. Byte Magazine, 19(2):129–130, 133–134, February 1994. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Perratore:1994:PAW

[Per94b] Ed Perratore. Printer at work: The first at work laser printer. Lexmark’s WinWriter 600 provides good performance and an excellent user interface. Byte Magazine, 19(5):149–??, May 1994. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Pettersen:1990:TMM

[Pet90] Richard Hans Pettersen. The tongues of men and machines: Do computer languages reflect the language and culture of the people who created them? Byte Magazine, 15(8):356–??, August 1990. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Peters:1991:GSV

[Pet91] R. A. Peters. Giga-Storage: A variety of storage technologies exist, and conflicting considerations are involved in choosing among them. Byte Magazine, 16 (5):201–202, 204–206, May 1991. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Platt:1992:FFS

[PGRA92] Stephen Platt, Tadesse Giorgis, Leslie Reisz, and Steve Apiki. Fast 486 file servers: Tests of 12 486 file servers. Byte Magazine, 17(10):212–??, October 1992. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Pickover:1993:SBF

[Pic93] Clifford A. Pickover. Stop bit:

Fractal fantasies: Fractals add a new dynamic to game design. Byte Magazine, 18(3):256– ??, March 1993. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Pietrek:1993:SAR

[Pie93] Matt Pietrek. Some assembly required: Debugger support in Windows 3.1: How to exploit debugging support in Windows 3.1. Byte Magazine, 18(10): 207–??, September 1993. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Piersol:1994:UHC

[Pie94] Kurt Piersol. Under the hood:

A close-up of OpenDoc: A new vendor-neutral document standard provides a natural way to work with information while operating on different platforms. Byte Magazine, 19(3):183–??, March 1994. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Pleas:1994:ERO

[Ple94] Keith Pleas. Extending the reach of OLE. Byte Magazine, 19(11): 199–??, November 1994. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Pountain:1990:HSN

[Pou90a] D. Pountain. HM Systems’ new Minstrel 486 workstation. Byte Magazine, 15(6):64IS–17, 18, June 1990. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Pountain:1990:IMO

[Pou90b] D. Pountain. Indexing with Micro-OCP. Byte Magazine, 15 (7):64IS–15–16, 62IS–18, 64IS– 20, 64IS–22, July 1990. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Pountain:1990:ODI

[Pou90c] D. Pountain. Opus Datasafe 120 (IBM-compatible machine). Byte Magazine, 15(5):64IS/11–14, 16, May 1990. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Pountain:1990:PL

[Pou90d] D. Pountain. The PCL language. Byte Magazine, 15(6):64IS–3, 4, 6, 8, 10, 12, June 1990. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Pountain:1990:RWA

[Pou90e] D. Pountain. A RISC workstation from Acorn. Byte Magazine, 15(13):72IS/14–18, 20, December 1990. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Pountain:1990:STP

[Pou90f]  D. Pountain.  Short takes

(pocket PC and word processor). Byte Magazine, 15(5):64IS/19– 20, May 1990. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Pountain:1990:SPP

[Pou90g] D. Pountain. Simulating a primitive parallel computer. Byte Magazine, 15(13):72IS/25–26, 28, 30, 32, 34, 36–37, December 1990. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Pountain:1990:CPP

[Pou90h] Dick Pountain. Configuring parallel programs, part 2: The Netherlands has a C compiler for parallel processing with the IN-MOS transputer. Byte Magazine, 15(1):327–334, January 1990. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Pountain:1990:OOP

[Pou90i] Dick Pountain. Object-Oriented programming: You can use Turbo Pascal 5.5 to learn the principles of OOP. Byte Magazine, 15(2):257–260, 262, 264, February 1990. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Pountain:1990:VCN

[Pou90j] Dick Pountain. Virtual channels: the next generation of transputers. Byte Magazine, 15(4):3–4 (or 3–12??), April 1990. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Pournelle:1990:BFM

[Pou90k] Jerry Pournelle. Backups, fax, and Mac disasters: A Mac mishap prompts Jerry to look at archiving software and backup devices. Byte Magazine, 15(5): 65–??, May 1990. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Pournelle:1990:CCN

[Pou90l]  Jerry Pournelle.  Casting a

chaotic network. Byte Magazine, 15(11):15–??, Fall 1990. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Pournelle:1990:CMA

[Pou90m] Jerry Pournelle. Chaos Manor awards: Find out if your favorite product has been honored. Byte Magazine, 15(4):53– ??, April 1990. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Pournelle:1990:DYP

[Pou90n] Jerry Pournelle. Double your pleasure: A hard disk drive saga and a Comdex report. Byte Magazine, 15(3):65–??, March 1990. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Pournelle:1990:FYC

[Pou90o] Jerry Pournelle. Fifteen years and counting: Jerry looks back at 15 years of BYTE. Byte Magazine, 15(9):65–??, September 1990. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Pournelle:1990:LMJ

[Pou90p] Jerry Pournelle. A lesson in maintenance: Jerry remembers the solution for an old problem and examines the prospects for a new DOS. Byte Magazine, 15 (10):81–??, October 1990. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Pournelle:1990:LLP

[Pou90q] Jerry Pournelle. Life, liberty and the pursuit of CD-ROMs: Jerry examines new CD-ROMs of classics, including the Bible and Shakespeare’s works. Byte Magazine, 15(7):65–??, July 1990. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Pournelle:1990:MSG

[Pou90r] Jerry Pournelle. A matter of style and grammar: Seeking a new word processor, and it’s upgrade time at Chaos Manor. Byte Magazine, 15(1):99–??, January 1990. CODEN BYTEDJ.

ISSN 0360-5280 (print), 1082-7838 (electronic).

Pournelle:1990:MVJ

[Pou90s] Jerry Pournelle. Multimedia video: Jerry looks at multimedia video boards, a new Modula-2, and assorted gadgets. Byte Magazine, 15(12):73–??, November 1990. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Pournelle:1990:ODD

[Pou90t] Jerry Pournelle. Optical disk daze: Jerry looks at some new CD-ROMs and educational software. Byte Magazine, 15(2): 99–114, February 1990. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Pournelle:1990:OMM

[Pou90u] Jerry Pournelle. Our man in moscow: Glasnost gives Jerry an inside look at Soviet computing. Byte Magazine, 15(8):65– ??, August 1990. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Pournelle:1990:TTC

[Pou90v] Jerry Pournelle. A tale of three conventions (and two cities): A look at CD-ROMs and a report on the American Association for the Advancement of Science. Byte Magazine, 15(6):65– ??, June 1990. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Pournelle:1990:WSJ

[Pou90w] Jerry Pournelle. Working smart:

Jerry looks at portable computers and an outlining program. Byte Magazine, 15(13): 73–??, December 1990. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Pountain:1991:BRP

[Pou91a] D. Pountain. Beyond RISC: the PgC7600 microprocessor. Byte Magazine, 16(3):90IS/109–114, March 1991. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Pountain:1991:IHP

[Pou91b] D. Pountain. Inside a heterogeneous parallel computer. Byte Magazine, 16(2):72IS/27–28, 30– 32, 34, February 1991. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Pountain:1991:LOS

[Pou91c] D. Pountain. Learning OOP style by playing poker. Byte Magazine, 16(7):72IS/43–44, 46, 48, 50, 52, July 1991. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Pountain:1991:OLA

[Pou91d] D. Pountain. Oberon: A look at the latest language to come out of Niklaus Wirth’s workshop. Byte Magazine, 16(3):135–136, 138, 140, 142, March 1991. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Pountain:1991:EV

[Pou91e] D. Pountain. Out of Eden: the VPi386. Byte Magazine, 16(8): 72IS/7–8, 10, August 1991. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Pountain:1991:PDE

[Pou91f]  D. Pountain.  Programming

databases the easy way. Byte Magazine, 16(1):721S–15–16, 721S–18, 721S–20, 721S–22, January 1991. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Pountain:1991:RMF

[Pou91g] D. Pountain. Research Machines’ first notebook PC. Byte Magazine, 16(9):88IS–15, September 1991. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Pountain:1991:TIO

[Pou91h] D. Pountain. Taos: an innovation in operating systems. Byte Magazine, 16(3):90IS/117– 126, March 1991. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Pountain:1991:CA

[Pou91i] Dick Pountain. Chorus of approval. Byte Magazine, 16(5):7, 8, 10, 12, 13, 16, May 1991. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Pournelle:1991:ARJ

[Pou91j] Jerry Pournelle. Atari revisited: Jerry looks at the Atari TT030. Byte Magazine, 16(5): 93–??, May 1991. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Pournelle:1991:DPF

[Pou91k] Jerry Pournelle. Desktop publishing fever: Jerry looks at some new games, including Chris Crawford’s Balance of the Planet, Mac software and hardware, and simulation software. Byte Magazine, 16(3):91– ??, March 1991. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Pournelle:1991:DJW

[Pou91l] Jerry Pournelle. DOS decisions: Jerry works with the Arche Legacy 486/33 and Digital Research’s DOS 5.0. Byte Magazine, 16(6):89–??, June 1991. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Pournelle:1991:DTC

[Pou91m] Jerry Pournelle. Dredging through the chaos: Jerry looks at new CD-ROMs, file utilities, and a multilink database. Byte Magazine, 16(7):73–??, July 1991. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Pournelle:1991:ELJ

[Pou91n] Jerry Pournelle. An exercise in logistics: Jerry looks at the latest version of Q and A. Byte Magazine, 16(10):81–??, October 1991. CODEN BYTEDJ.

ISSN 0360-5280 (print), 1082-7838 (electronic).

Pournelle:1991:ADJ

[Pou91o] Jerry Pournelle. It’s all digital: Jerry looks at multimedia products. Byte Magazine, 16 (9):89–??, September 1991. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Pournelle:1991:JCJ

[Pou91p] Jerry Pournelle. Jukebox computing: Jerry looks at new CD-ROMs and a CD-ROM drive, a brick of a computer, and a new trackball. Byte Magazine, 16 (1):73–74, 77–78, 80, 84, 86–88, January 1991. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Pournelle:1991:RAJ

[Pou91q] Jerry Pournelle. On the road again: Jerry looks at a new portable computer. Byte Magazine, 16(8):73–??, August 1991. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Pournelle:1991:PLJ

[Pou91r] Jerry Pournelle. A pack of laptops: Jerry picks the ideal portable computer. Byte Magazine, 16(2):73–74, 77–78, 80, 84, 86–88, February 1991. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Pournelle:1991:RCD

[Pou91s] Jerry Pournelle. The revolution continues: Did personal computers bring down the evil empire? Byte Magazine, 16(13): 85–??, December 1991. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Pournelle:1991:SAJ

[Pou91t] Jerry Pournelle. Stellar attractions: Jerry builds a new computer. Byte Magazine, 16(12): 117–??, November 1991. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Pournelle:1991:UCA

[Pou91u] Jerry Pournelle. User’s choice awards: The best products of 1990. Byte Magazine, 16(4):91– ??, April 1991. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Pountain:1992:FIP

[Pou92a] D. Pountain. First impressions Psion’s powerful pocketable: A full-fledged hand-held computer. Byte Magazine, 17 (1):40–41, January 1992. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Pountain:1992:PSW

[Pou92b] D. Pountain. Psion Series 3:

the whole story. Byte Magazine, 17(3):92IS7–8, 92IS10, 92IS12, 92IS14, 92IS16, March 1992. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Pountain:1992:FWD

[Pou92c] Dick Pountain. Four Windows desktops (desktop enhancers). Byte Magazine, 17(11):113–116, November 1992. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Pountain:1992:UHC

[Pou92d] Dick Pountain. Under the hood. A call to ARM: The 32-bit ARM610 is a high-performance, power-saving RISC CPU in a tiny package. Byte Magazine, 17(12):293–??, November 1992. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Pournelle:1992:CFJ

[Pou92e] Jerry Pournelle. Computer ferment: Jerry looks at Windows software. Byte Magazine, 17 (7):101–??, July 1992. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Pournelle:1992:HEJ

[Pou92f]  Jerry Pournelle.  The high

end: Jerry discusses technology trickle-down. Byte Magazine, 17 (2):67–??, February 1992. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Pournelle:1992:IBC

[Pou92g] Jerry Pournelle. Interrupts and big cats: Jerry configures a new 486 computer. Byte Magazine, 17(3):93–??, March 1992. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Pournelle:1992:MGS

[Pou92h] Jerry Pournelle. The Mac goes to school: Jerry sends a computer to college. Byte Magazine, 17(6):101–??, June 1992. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Pournelle:1992:MMJ

[Pou92i] Jerry Pournelle. Multimedia medley: Jerry looks at the latest in multimedia products and upgrade kits. Byte Magazine, 17(8):95–??, August 1992. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Pournelle:1992:RAL

[Pou92j] Jerry Pournelle. On the road again: Laptops, flashdrives, and power supplies at Chaos Manor. Byte Magazine, 17(10):101–??, October 1992. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Pournelle:1992:UMW

[Pou92k] Jerry Pournelle. Unsolved mysteries: A Windows mystery revealed; and wrapping up the annual user’s choice awards at Chaos Manor. Byte Magazine, 17(5):85–??, May 1992. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Pournelle:1992:UCAa

[Pou92l] Jerry Pournelle. User’s choice awards: Jerry issues his annual user’s choice awards. Byte Magazine, 17(4):83–??, April 1992. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Pournelle:1992:UCC

[Pou92m] Jerry Pournelle. User’s column: A computer in the hand: Palmtops, PCMCIA, virus newsletters, and the BBS scene. Byte Magazine, 17(14):97–??, December 1992. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Pournelle:1992:UCP

[Pou92n] Jerry Pournelle. User’s column. pondering OS/2: OS/2 invades Chaos Manor. Byte Magazine, 17(12):109–??, November 1992. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Pournelle:1992:UCAb

[Pou92o] Jerry Pournelle. User’s column: The Amiga revisited: Jerry looks at the Amiga 3000T. Byte Magazine, 17(9):113–??, September 1992. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Pountain:1993:AMP

[Pou93a] Dick Pountain. Amstrad’s modest PDA: The first PDA is somewhat less ambitious than systems (like newton) still on the drawing board. dick works with Amstrad’s PDA 600 Pen Pad, which offers pen input with character recognition. Byte Magazine, 18(9):161–162, 164, August 1993. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Pountain:1993:CCA

[Pou93b] Dick Pountain. Computing without clocks: Asynchronous processors turn computing on its head. Byte Magazine, 18(1):145– ??, January 1993. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Pountain:1993:EVB

[Pou93c] Dick Pountain. Extending visual basic: Visual Basic extensions let you produce quality programs faster than ever before. Byte Magazine, 18(5):43– ??, Spring 1993. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Pountain:1993:OGA

[Pou93d] Dick Pountain. Oberon: A glimpse at the future: A radically object-oriented design previews future operating systems. Byte Magazine, 18(6):111–??, May 1993. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Pountain:1993:TPA

[Pou93e] Dick Pountain. Track people with active badges: Developed by Olivetti and DEC, the Active Badge network improves at-work communications by tracking your colleagues’ whereabouts. Byte Magazine, 18(13):57–??, December 1993. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Pountain:1993:UHC

[Pou93f] Dick Pountain. Under the hood: Computing on wheels: This month’s column literally goes under the hood to examine a new generation of automotive-control computers. Byte Magazine, 18 (6):213–??, May 1993. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Pountain:1993:UHP

[Pou93g] Dick Pountain. Under the hood: Pentium: More RISC than CISC: Why the Pentium’s architecture doesn’t measure up to its RISC competitors. Byte Magazine, 18(10):195–??, September 1993. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Pountain:1993:UHM

[Pou93h] Dick Pountain. Under the hood the multiprocessor solution: Multiprocessor architectures lead the charge to improve I/O performance. Byte Magazine, 18 (7):185–??, June 1993. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Pournelle:1993:QOC

[Pou93i] Jerry Pournelle. The DOS 6 question: Our columnist finds DOS 6 is the least expensive route to disk compression and memory optimization. Byte Magazine, 18(8):209–??, July 1993. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Pournelle:1993:LWW

[Pou93j] Jerry Pournelle. LAN wars:

Windows for Workgroups battles LANtastic for domination at Chaos Manor. Byte Magazine, 18 (2):97–??, February 1993. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Pournelle:1993:PBI

[Pou93k] Jerry Pournelle. Pournelle: Basic instinct: Programming a Quick-Basic application, tax software, and the search for the ideal word processor. Byte Magazine, 18 (9):209–??, August 1993. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Pournelle:1993:PCW

[Pou93l]  Jerry Pournelle.  Pournelle:

Clean water and dirty keyboards: The water is in Stockholm; the keyboards are in Chaos Manor. Byte Magazine, 18(13):225–??, December 1993. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Pournelle:1993:PFF

[Pou93m] Jerry Pournelle. Pournelle: Fenasoft and furniture: Jerry travels to Brazil to attend Fenasoft, a world-class computer exposition. Byte Magazine, 18(12): 325–??, November 1993. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Pournelle:1993:PIP

[Pou93n] Jerry Pournelle. Pournelle:

IBM’s preemptive strike: OS/2 2.1 is technically excellent. what’s needed is widespread device-driver support. Byte Magazine, 18(10):221–??, September 1993. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Pournelle:1993:PSP

[Pou93o] Jerry Pournelle. Pournelle: Statistics programs help businesses work: Mysteries revealed. Byte Magazine, 18(7):201–??, June 1993. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Pournelle:1993:UH

[Pou93p] Jerry Pournelle. Unsung heroes. Byte Magazine, 18(5):136–??, Spring 1993. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Pournelle:1993:UCC

[Pou93q] Jerry Pournelle. User’s column: CD-ROM secrets: The trick to integrating CD-ROM and Windows for Workgroups. Byte Magazine, 18(3):73–??, March 1993. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Pournelle:1993:UCO

[Pou93r] Jerry Pournelle. User’s column: Once more into the breach: User’s choice awards wrap-up. Byte Magazine, 18(6): 75–??, May 1993. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Pournelle:1993:UCP

[Pou93s] Jerry Pournelle. User’s column:

The principle of pursuit: Microsoft moves to dominate with Access and Windows for Workgroups. Byte Magazine, 18 (1):97–??, January 1993. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Pournelle:1993:UCW

[Pou93t] Jerry Pournelle. User’s column:

What’s hot, what’s not: Pournelle’s annual orchid and onion parade arrives. Byte Magazine, 18(4):81–??, April 1993. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Pountain:1994:CMP

[Pou94a] Dick Pountain. The Chorus Microkernel: Pountain takes a look at Chorus/MiX, a microkernelbased distributed Unix operating system from france. Byte Magazine, 19(1):131–??, January 1994. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Pountain:1994:DKR

[Pou94b] Dick Pountain. A different kind of RISC: Hewlett–Packard’s PA-RISC 7200 superscalar processor is not typical, and neither is its performance: It’s likely to hold the title of “fastest RISC in town” for the immediate future. Byte Magazine, 19(8):185– ??, August 1994. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Pountain:1994:FAC

[Pou94c] Dick Pountain. The fine art of CD-ROM publishing: The electronic version of the National Gallery collection demonstrates how traditional material can be adapted to new publishing media. Byte Magazine, 19(6):47– 48, 50, 52, 54, June 1994. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Pountain:1994:FPC

[Pou94d] Dick Pountain. Functional programming comes of age: Following a decade of crucial research breakthroughs, functional programming languages are catching on, even in the realms of parallel programming and real-time systems. Byte Magazine, 19(8):183– ??, August 1994. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Pountain:1994:LBC

[Pou94e] Dick Pountain. The last bastion: “Computer on a chip” is old hat; are you ready for a “supercomputer on a chip”? Byte Magazine, 19(9):47–??, September 1994. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Pountain:1994:MTS

[Pou94f] Dick Pountain. Microprocessor trends: Several trends converge to threaten the near monopoly the Intel 80x86 architecture enjoys on the desktop. Byte Magazine, 19(1):74–??, January 1994.

CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Pountain:1994:PCT

[Pou94g] Dick Pountain. Parallel course:

With Taos — a radically different, object-oriented, parallel operating system — it’s possible to harness together the power of different types of processors. Byte Magazine, 19(7): 53–??, July 1994. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Pountain:1994:SCS

[Pou94h] Dick Pountain. Starting with a clean sheet: Objects are the future of programming, and the CleanSheet spreadsheet construction set gives you a picture of what that future holds. Byte Magazine, 19(11):75–??, November 1994. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Pournelle:1994:PLF

[Pou94i] Jerry Pournelle. Pournelle: A look to the future: Jerry highlights some major trends he sees in the computer industry. Byte Magazine, 19(11):299–??, November 1994. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Pournelle:1994:PPS

[Pou94j] Jerry Pournelle. Pournelle: A Pentium is sounded out: Sound boards clash with hard drive controllers on new Pentium system. Byte Magazine, 19(6):277– ??, June 1994. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Pournelle:1994:PET

[Pou94k] Jerry Pournelle. Pournelle: An educational trip: A journey to England finds Jerry attending a conference on education and a computer show. Byte Magazine, 19(7):197–??, July 1994. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Pournelle:1994:PBB

[Pou94l]  Jerry Pournelle.  Pournelle:

Booting, benchmarking, and Bob’s your uncle: A potpourri of activity at Chaos Manor as Jerry continues his exploration of the Windows machinery. Byte Magazine, 19(3):201–??, March 1994. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Pournelle:1994:PCB

[Pou94m] Jerry Pournelle. Pournelle: Crash, bang — quake: Jerry survives an earthquake, installs a LAN server, and gives out more user’s choice awards. Byte Magazine, 19(5):205–??, May 1994. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Pournelle:1994:PDB

[Pou94n] Jerry Pournelle. Pournelle: Don’t blink: Jerry tackles the trinity of SCSI, video cards, and sound cards; as usual, he wins. Byte Magazine, 19(9): 205–??, September 1994. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Pournelle:1994:POE

[Pou94o] Jerry Pournelle. Pournelle: Odds and ends: Jerry looks at a dramatic morphing program, a neural network, and more. Byte Magazine, 19(10):211–??, October 1994. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Pournelle:1994:PTL

[Pou94p] Jerry Pournelle. Pournelle:

Traveling light: Jerry experiments with minimalist computing. Byte Magazine, 19(8):193– ??, August 1994. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Pournelle:1994:PTT

[Pou94q] Jerry Pournelle. Pournelle: Travels and travails: An especially busy month finds Jerry roaming the country but also active at Chaos Manor. Byte Magazine, 19 (1):243–??, January 1994. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Pournelle:1994:PUH

[Pou94r]  Jerry Pournelle.  Pournelle:

Upgrades from hell: One assumes Jerry didn’t have a great month. Byte Magazine, 19(2): 185–??, February 1994. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Pournelle:1994:PWH

[Pou94s]  Jerry Pournelle.  Pournelle:

What’s hot, what’s not: It’s time once again for the annual user’s choice awards from Chaos Manor. Byte Magazine, 19(4):241–??, April 1994. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Page:1990:LFD

[PP90] Matt Page and Mary Page. Laying out the future (DTP). Byte Magazine, 15(11):169–170, 172, Fall 1990. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Pountain:1993:PPP

[PP93] Dick Pountain and Amstrad’s Affordable Pen Package. The Pen Pad PDA600 may not offer the dazzle of Newton, but it’s affordable and it works. Byte Magazine, 18(7):24–??, June 1993. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Prasnikar:1993:RCR

[Pra93] Neven Prasnikar. Report from Croatia: Recovery through technology: Technology helps a troubled country rebound. Byte Magazine, 18(4):41–??, April 1993. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Pountain:1994:ESS

[PS94] Dick Pountain and Clemens Szyperski. Extensible software systems: New programming tools are needed to develop software systems tha can be easily extended with new modules. Byte Magazine, 19(5):57– ??, May 1994. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Psaltis:1992:POM

[Psa92] Demetri Psaltis. Parallel optical memories: A performance breakthrough. Byte Magazine, 17(9): 179–??, September 1992. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Perratore:1993:FFB

[PTUM93] Ed Perratore, Tom Thompson, Jon Udell, and Rich Malloy. Fighting fatware: Bloated software slows you down, but help is on way. Byte Magazine, 18(4):98–102, 104–106, 108, April 1993. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Raich:1990:UNH

[Rai90] K. E. Raich. Using a nonstandard hard disk drive. Byte Magazine, 15(1):8015–8013–8010, January 1990. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Ransen:1990:ART

[Ran90] Owen F. Ransen. The art of ray tracing: The perfect match: complex 3-D image generation and parallel microprocessors. Byte Magazine, 15(2):

238–242, February 1990. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic). Occam timings for ring of transputers.

Ranade:1993:RIH

[Ran93] Jay Ranade. Report from India: High-Tech opportunities abound: The government is high on growth but down on imports. Byte Magazine, 18(3):37– ??, March 1993. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Redfern:1992:OPF

[RARH92] Andy Redfern, Dave Andrews, Andy Reinhardt, and Tom Halfhill. Other players find niches. Byte Magazine, 17 (14):??, December 1992. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Rash:1990:BBI

[Ras90a] Wayne Rash, Jr. Backing up the biggies: It takes more than a box of floppy disks to meet today’s backup needs. Byte Magazine, 15 (2):133–134, February 1990. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Rash:1990:BTX

[Ras90b] Wayne Rash, Jr. Born to travel: XT-class laptops from GRiD and Sharp offer the right mix of features for computing en route. Byte Magazine, 15(1):177–180, January 1990. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Rash:1990:CRR

[Ras90c] Wayne Rash, Jr. CD-ROM to the rescue: CD-ROM databases can provide your business with valuable information in a hurry. Byte Magazine, 15(4):77–78, April 1990. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Rash:1990:CCS

[Ras90d] Wayne Rash, Jr. Charting the course: Sophisticated presentation software can produce highquality slides. Byte Magazine, 15 (8):79–80, 82, August 1990. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Rash:1990:CEP

[Ras90e] Wayne Rash, Jr. Cheap and easy publishing: You may not need all the bells and whistles to look like a pro. Byte Magazine, 15 (1):123–??, January 1990. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Rash:1990:DIY

[Ras90f] Wayne Rash, Jr. Do-It-Yourself CD-ROMs: If you need access to vast quantities of your own data, you could make your own CD-ROM. Byte Magazine, 15(5):89– 90, 92, May 1990. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Rash:1990:FJM

[Ras90g] Wayne Rash, Jr. The family jewels: To make sure your data is secure, choose a strategy and see that it’s carried out. Byte Magazine, 15(3):85–86, 88, March 1990. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Rash:1990:GBG

[Ras90h] Wayne Rash, Jr. Getting bigger groupware: With groupware, you can communicate with almost anyone, almost anywhere. Byte Magazine, 15(13):93–94, 96, December 1990. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Rash:1990:GGW

[Ras90i] Wayne Rash, Jr. The growth of groupware: Wayne addresses groupware and how to determine what capabilities your company needs. Byte Magazine, 15(12): 89–90, 92, November 1990. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Rash:1990:MMP

[Ras90j] Wayne Rash, Jr. Moving down to micros: Powerful decisionsupport systems, once used only on mainframes, are now migrating to micros. Byte Magazine, 15(9):93–94, 96, 98, September 1990. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Rash:1990:PNY

[Ras90k] Wayne Rash, Jr. Picture this: Now you can make professional presentations quickly and easily. Byte Magazine, 15(10):111– ??, October 1990. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Rash:1990:PFP

[Ras90l] Wayne Rash, Jr. A Poqet full of power: It’s small. it’s innovative. but is it practical? Wayne Rash, Jr. takes the Poqet PC on the road. Byte Magazine, 15(12): 239, 240, November 1990. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Rash:1990:TWD

[Ras90m] Wayne Rash, Jr. Through a window, darkly. Byte Magazine, 15(11):47–??, Fall 1990. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Rash:1990:CPP

[Ras90n] Wayne Rash, Jr. Up close and personal: Personal organizers let you calculate, compute, and transfer data on the fly. Byte Magazine, 15(7):87– 88, July 1990. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Rash:1990:WWW

[Ras90o] Wayne Rash, Jr. Who, what, when, and why not: A business scheduling package can restore your sanity. Byte Magazine, 15 (6):85–86, 88, June 1990. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Rasmus:1990:MSM

[Ras90p] Daniel W. Rasmus. The Mac state of mind: A look at some expert-system shells and AI languages for your Macintosh. Byte Magazine, 15(1):305–314, January 1990. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Rash:1991:BFS

[Ras91a] Wayne Rash, Jr. Beyond file sharing: What fax servers, CD-ROMs, and other add-ons do for LANs. Byte Magazine, 16(13): 103–??, December 1991. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Rash:1991:CCW

[Ras91b] Wayne Rash, Jr. Corporate connections: Wayne Rash, Jr. describes how networks that span entire companies will change the way you work in the 1990s. Byte Magazine, 16(9):215–220, 222–223, September 1991. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Rash:1991:CSH

[Ras91c] Wayne Rash, Jr. Corporate style: How consistent is the language of your business documents? Byte Magazine, 16(2):89–??, February 1991. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Rash:1991:DYL

[Ras91d] Wayne Rash, Jr. Defending your LAN: Be prepared for secondguessers when proposing a LAN spec. Byte Magazine, 16(10):97– ??, October 1991. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Rash:1991:LHY

[Ras91e] Wayne Rash, Jr. A LAN away from home: Your LAN: You can take it with you. Byte Magazine, 16(6):107–??, June 1991. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Rash:1991:MLL

[Ras91f] Wayne Rash, Jr. Managing the LAN: LAN management doesn’t get enough respect. Byte Magazine, 16(7):91–??, July 1991. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Rash:1991:MLH

[Ras91g] Wayne Rash, Jr. The missing link: Heterogeneous LANs can communicate despite a shortage of solutions from major platform vendors. Byte Magazine, 16(5):111–??, May 1991. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Rash:1991:PMC

[Ras91h] Wayne Rash, Jr. The power man cometh: The big orange power truck pulls up again, but this time Wayne’s ready. Byte Magazine, 16(1):89–90, 92, January 1991. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Rash:1991:PKB

[Ras91i] Wayne Rash, Jr. That pesky 640-KB barrier: DOS 5.0 and a memory manager make a good team. Byte Magazine, 16(12): 133–??, November 1991. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Rash:1991:WNN

[Ras91j] Wayne Rash, Jr. WaveLAN: A network with no strings attached: NCR’s product does away with LAN cabling — for a price. Byte Magazine, 16(6):294– 296, June 1991. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Rash:1991:WEI

[Ras91k] Wayne Rash, Jr. Whose enterprise is it: Make sure that your enterprise network needs are indeed your own. Byte Magazine, 16(3):107–108, 110, 112, March 1991. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Rash:1991:WBW

[Ras91l] Wayne Rash, Jr. Windows and the business workstation: Windows 3.0 will change the way you buy PCs. Byte Magazine, 16(4):103–??, April 1991. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Rash:1991:YFLb

[Ras91m] Wayne Rash, Jr. Your first LAN: do it yourself? Byte Magazine, 16(9):105, 108, 110, September 1991. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Rash:1991:YFLa

[Ras91n] Wayne Rash, Jr. Your first LAN:

Tips for first-time network buyers. Byte Magazine, 16(8):93–94, 96, August 1991. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Rasmus:1991:PEW

[Ras91o] D. W. Rasmus. Putting the experts to work: The 1990s will see the walls between intelligent applications and conventional applications crumble. Byte Magazine, 16(1):281–282, 285, 287, January 1991. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Rash:1992:MCS

[Ras92a] Wayne Rash, Jr. Making connections: Seven portable Ethernet adapters reviewed. Byte Magazine, 17(13):61–??, ?? 1992. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Rash:1992:MMB

[Ras92b] Wayne Rash, Jr. Multimedia moves beyond the hype: Can multimedia succeed as a legitimate business application? Byte Magazine, 17(2):85–??, February 1992. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Rash:1992:NNC

[Ras92c] Wayne Rash, Jr. NEC’s notebook compromises for color: NEC’s active-matrix color notebook is colorful but cumbersome.

Byte Magazine, 17(8):253–254, August 1992. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Rash:1992:WMB

[Ras92d] Wayne Rash, Jr. Windows moves out: Better notebook computers make traveling with Windows a workable proposition. Byte Magazine, 17(3):109–??, March 1992. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Rasmus:1992:OOC

[Ras92e] Daniel W. Rasmus. Object-Oriented case. Byte Magazine, 17(14):??, December 1992. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Rasmus:1992:ROO

[Ras92f] Daniel W. Rasmus. Relating to objects: Object technology finds a place in database management. Byte Magazine, 17(14): 161–??, December 1992. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Redfern:1991:KRR

[Red91a] A. Redfern. Kyocera rustles up the Refalo (pocket computer). Byte Magazine, 16(9):88IS–7– 8, 88IS–10, 88IS–12, September 1991. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Redfern:1991:OE

[Red91b] Andy Redfern. The outlook for Europe. Byte Magazine, 16(11):

58–??, ?? 1991. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Redfern:1992:MRC

[Red92a] Andy Redfern. Make the right CPU move: New CPUs confuse PC buying decisions. Byte Magazine, 17(14):114–??, December 1992. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Redfern:1992:RIC

[Red92b] Andy Redfern. Report from Israel: Coming in from the cold: As Mideast tensions wane, Israel can sell its wares more openly. Byte Magazine, 17(14): 41–??, December 1992. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Redfern:1993:CIS

[Red93] Andy Redfern. Computers Italian style: The Italians love stylish computers but crave innovation from abroad. Byte Magazine, 18(2):47–??, February 1993. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Reff:1994:SRD

[Ref94]  Bobby Joe Reff.  Software

roundup: Dynamic-Viewing spreadsheets: Lotus Improv gained exposure and praise as a tool for viewing strategic data in a variety of useful ways. although Improv has been discontinued, other programs deliver similar functionality, slicing and dicing data to create analytical models. NSTL evaluates Microsoft Excel, Quattro pro, and TM/1 perspectives. Byte Magazine, 19(11): 255–258, November 1994. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Reinhardt:1990:NSQ

[Rei90a] A. Reinhardt. A new status quo for Quattro Borland adds 3-D graphics to its spreadsheet. Byte Magazine, 15(12):156–157, November 1990. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Reinhardt:1990:PPN

[Rei90b] A. Reinhardt. Power to the portables: With new options coming on strong, the battery battle for portables is heating up. Byte Magazine, 15(5):273– 276, May 1990. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Reinhardt:1990:SES

[Rei90c] A. Reinhardt. Suddenly, everything’s smaller in Texas: TI’s 5.7-pound 368SX notebook. Byte Magazine, 15(13):151–153, December 1990. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Reinhardt:1990:PC

[Rei90d]  Andrew Reinhardt.  Playing

catch-up. Byte Magazine, 15 (??):279, November 1990. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Reimer:1991:MMP

[Rei91a] J. Reimer. Memories in my pocket: Solid-state memory cards provide the size, weight, and capacity necessary to be practical in a notebook computer. Byte Magazine, 16(2): 251–252, 254–256, 258, February 1991. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Reinhardt:1991:NKN

[Rei91b]  A. Reinhardt.  NCR knows

notepads: The System 3125A tablet computer runs PenPoint or PenWindows. Byte Magazine, 16(8):37–38, August 1991. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Reinhardt:1991:NEE

[Rei91c] A. Reinhardt. New extras for Excel: Microsoft’s new version can do spreadsheet outlining. Byte Magazine, 16(2):136–138, February 1991. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Reinhardt:1991:DOT

[Rei91d] Andy Reinhardt. Disk, DAT, and the optical thing. Byte Magazine, 16(11):102–??, ?? 1991. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Reid:1992:RTL

[Rei92a] T. R. Reid. Report from Tokyo:

In the land of the double byte: Japan’s software developers work wonders. Byte Magazine, 17 (10):47–??, October 1992. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Reinhardt:1992:PPP

[Rei92b]  Andrew Reinhardt.  Penny-

Pinching PCs: How they did it: Low-priced PCs bring benefits — and risks — to buyers. Byte Magazine, 17(12):128–136, November 1992. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Reinhardt:1992:ADT

[Rei92c]  Andy Reinhardt.  Acquiring

data through Windows: Visual programming for data acquisition. Byte Magazine, 17(9): 63–??, September 1992. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Reinhardt:1993:RHK

[Rei93a] Andrew Reinhardt. Report from Hong Kong: Hong Kong hopes to profit by its ties to mainland China. Byte Magazine, 18(1):41– ??, January 1993. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Reinhardt:1993:HTC

[Rei93b] Andy Reinhardt. HP takes color mainstream: Hewlett– Packard’s DeskJet 1200C meets the demand of users who covet color. Byte Magazine, 18(7):38– ??, June 1993. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Reinhardt:1993:OF

[Rei93c] Andy Reinhardt. Optical flavors. Byte Magazine, 18(2):??, February 1993. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Reinhardt:1993:SMC

[Rei93d] Andy Reinhardt. Smarter E-Mail is coming: Intelligent E-mail delivers more than messages. it will change how your business works by improving communications and automating workflow. Byte Magazine, 18(3):90, 92, 94, 96–98, 100, 102, 105–106, 108, March 1993. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Reinhardt:1993:VCD

[Rei93e] Andy Reinhardt. Video conquers the desktop: Integrated video capability has arrived for desktop systems — and with it, intriguing new possibilities. Byte Magazine, 18(10):64–67, 70, 72, 76– 77, 80, September 1993. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Reinhardt:1994:DCT

[Rei94a] A. Reinhardt. Desktop conferencing takes off: Although not as flashy as full-length feature films on CD-ROM or 3-D virtual reality, desktop conferencing applications can save you time and money. Byte Magazine, 19(6):24– 25, June 1994. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Reinhardt:1994:BDH

[Rei94b] Andy Reinhardt. Building the data highway: Building the data highway will require a blending of diverse technologies. what’s the best way to send large amounts of data over today’s mix of copper, coaxial cable, and fiber-optic lines? should the transport protocol be ATM? TCP/IP? TCP/ IP over ATM? the major players — the phone companies, the cable companies, and proponents of the Internet — each have their own vision of what the data highway should be. whose vision should prevail? Byte Magazine, 19(3):46–??, March 1994. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Reinhardt:1994:CCF

[Rei94c] Andy Reinhardt. Commentary: Call in the Feds: The government must be involved with the data highway. Byte Magazine, 19(7):270–??, July 1994. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Reinhardt:1994:MSH

[Rei94d] Andy Reinhardt. Managing storage: How do you get the most recent, important data where it is needed in a distributed environment? Byte Magazine, 19(6):153–??, June 1994. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Reinhardt:1994:MND

[Rei94e]  Andy Reinhardt.  Managing

the new document: As compound document architectures make their way into the desktop computing mainstream, the way we manage documents is fundamentally changing. Byte Magazine, 19(8):90–??, August 1994. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Reinhardt:1994:NSI

[Rei94f] Andy Reinhardt. The network with smarts: Intelligent networks from AT&T and IBM could dramatically change the way you work and may set the model for a future of mobile software agents. Byte Magazine, 19(10):50–52, 56, 58, 60, 62, 64, October 1994. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Reinhardt:1994:SCP

[Rei94g]  Andy Reinhardt.  Smarter

copiers, printers, and fax devices are coming: Microsoft and Novell believe that peripherals and PDAs need their own operating system. Byte Magazine, 19 (11):81–??, November 1994. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Romkey:1991:AAP

[RF91] J. Romkey and S. Fisher. All about packet drivers. Byte Magazine, 16(5):297–298, 300, 302– 304, 306, May 1991. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Reinhardt:1993:PCP

[RH93] Andy Reinhardt and T. R. Halfhill. Pentium changes the PC: The Intel Pentium CPU demands subsystems and I/O that can keep pace and that call for a fundamental rethinking of how to build everything from the expansion bus to memory architecture. Byte Magazine, 18(8):80–82, 84, 86–87, 90, 92–93, July 1993. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Ribarsky:1993:NDF

[Rib93] William Ribarsky. Navigating the data flood: Find your way through large data sets visually. Byte Magazine, 18(4):129– ??, April 1993. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Richter:1991:XNG

[Ric91]  J. Richter.  XGA: A new

graphics standard: The Extended Graphics Array offers full graphics functionality as well as bus mastering. Byte Magazine, 16(2):285–286, 288–290, February 1991. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Richter:1992:DTD

[Ric92] Jake Richter. DGX takes the direct approach to graphics performance: Dell’s direct graphics accelerator speeds Windows and CAD. Byte Magazine, 17 (10):235–??, October 1992. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Richter:1994:DDA

[Ric94] Jane Richter. Distributing data:

Although complex, businesses are relying more and more on distributed databases. Byte Magazine, 19(6):139–??, June 1994. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Riezenman:1991:CWF

[Rie91] M. Riezenman. Chips that work:

Without fault tolerance, you’d be paying a lot more for your next computer. Byte Magazine, 16 (8):187–190, August 1991. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Reinhardt:1991:CF

[RLU^{+}91] Andy Reinhardt, Owen Linderholm, Ellen Ullman, D. Barker, David Andrews, and Rich Malloy. Cornerstones of the future. Byte Magazine, 16(11):27– ??, ?? 1991. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Rash:1992:FMG

[RM92] Wayne Rash, Jr. and T. Marshall. Fax modems to go: These 10 portable packages deliver varying degrees of performance when you’re traveling. Byte Magazine, 17(13):49–51, 54, 56, ?? 1992. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Ryan:1994:BCR

[RNG94] Bob Ryan, Michael Nadeau, and Raymond Ga Cote. Books and CD-ROMs: A Cache-Memory primer: The theory and practice of cache-memory design, a fantastic world on CD-ROM, multimedia production, and strange patterns. Byte Magazine, 19 (3):41–??, March 1994. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Robinson:1990:FMG

[Rob90a] Phillip Robinson. The four multimedia gospels: Multimedia is taking the computer world by storm, and it’s more available than you think. Byte Magazine, 15(2):203–212, February 1990. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Robinson:1990:HOS

[Rob90b] Phillip Robinson. The High-Octane semiconductor: Chip makers move gallium arsenide from curiosity to practicality. Byte Magazine, 15(1):251–258, January 1990. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Robinson:1991:PPV

[Rob91a] M. Robinson. Popular and parallel: Various approaches to creating truly scalable shared-memory architectures. Byte Magazine, 16 (6):219–224, 226, 228, June 1991. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Robinson:1991:TLS

[Rob91b] M. Robinson. Through a lens smartly: Information lens can simplify electronic communications. Byte Magazine, 16(5):177– 178, 180, 186–187, May 1991. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Rogers:1991:IJT

[Rog91] A. J. Rogers. Ink jet takes off:

Phase-change ink-jet color printing combines brilliant color with the ability to print on plain paper. Byte Magazine, 16(10):163– 164, 166, 168, October 1991. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Rosenberg:1990:AMA

[Ros90a] S. Rosenberg. Art in motion:

Autodesk’s animator lets anyone create animated graphics. Byte Magazine, 15(3):173–174, 177, March 1990. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Rosenberg:1990:NAG

[Ros90b] S. Rosenberg. New adventures in graphic design: Micrografx’s designer 3.0 rivals competitor Corel Draw. Byte Magazine, 15(8): 198–199, August 1990. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Rowell:1993:WUP

[Row93] Dave Rowell. Windows under 4 pounds: Two 486-class subnotebooks for Windows: Epson’s ActionNote 4000 and IBM’s ThinkPad 500. Byte Magazine, 18(12):249–250, 252, November 1993. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Ryan:1990:OFK

[RP90] Bob Ryan and B. Passavanti. The once and future king: Hard disk technology will be your primary computer storage medium for years to come. Byte Magazine, 15(12):301–306, November 1990. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Reinhardt:1992:GCC

[RPRM92] Andy Reinhardt, Ed Perratore, Andy Redfern, and Rich Malloy. The greening of computers: The computer industry wakes up to ecological concerns. Byte Magazine, 17(9):147–??, September 1992. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Reinhardt:1990:SRS

[RS90] A. Reinhardt and B. Smith. Sizzling RISC systems from IBM: IBM’s RISC System/6000 family sets a new standard of performance. Byte Magazine, 15 (4):124–128, April 1990. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Reekes:1991:MSR

[RT91] J. Reekes and T. Thompson. Macintosh sound revealed: A look at the Mac’s built-in sound capabilities. Byte Magazine, 16 (7):249–250, 252, 254, 256, 258, 260, 262, July 1991. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Ryan:1994:PWI

[RT94a] Bob Ryan and Tom Thompson. PowerPC 604 weighs in: IBM and Motorola’s latest CPU outperforms the Pentium. Byte Magazine, 19(6):265– ??, June 1994. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Ryan:1994:RGR

[RT94b] Bob Ryan and Tom Thompson.

RISC grows up: RISC vendors expand their offering to respond to the needs of a wider variety of applications. Byte Magazine, 19(1):91–??, January 1994. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Reinhardt:1993:BCR

[RUHG93] Andy Reinhardt, Jon Udell, Frank Hayes, and Raymond Ga Cote. Books and CD-ROMs: Big Blue tales: Inside the walls of IBM, technological change, constructing your own robots, and more. Byte Magazine, 18(12): 49–??, November 1993. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Ryan:1990:CDI

[Rya90a] B. Ryan. Coping with diversity:

Incompatibility between computers with different architectures doesn’t have to be an obstacle. Byte Magazine, 15(3):257– 260, March 1990. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Ryan:1990:SUN

[Rya90b] B. Ryan. Scripts unbounded: New, improved graphical scripting languages may make stand alone applications obsolete. Byte Magazine, 15(8):235–236, 238– 240, August 1990. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Ryan:1990:SAB

[Rya90c] B. Ryan. Separated at birth: Although they’re the same age, PCs and supercomputers are now sharing more than just birthdays. Byte Magazine, 15(5):207–208, 210, May 1990. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Ryan:1990:SCW

[Rya90d] B. Ryan. The succession crisis: Will DOS yield its crown to OS/2 or Unix? Byte Magazine, 15(3): 199–200, 202, March 1990. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Ryan:1990:ENP

[Rya90e] Bob Ryan. Entering a new phase: Optical and magnetic are at opposite ends of the spectrum. can phase-change technology bridge the gap? Byte Magazine, 15(12): 289–290, 292, 294, 296, November 1990. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Ryan:1990:FCS

[Rya90f] Bob Ryan. Farewell to chips:

Semiconductor technology is approaching its theoretical and practical limits. where do we go from here? Byte Magazine, 15 (1):237–249, January 1990. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Ryan:1990:MMS

[Rya90g] Bob Ryan. A marriage made in silicon: BiCMOS proves that good things come in pairs. Byte Magazine, 15(1):261–266, January 1990. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Ryan:1991:AIC

[Rya91a] B. Ryan. AI’s identity crisis: Can AI provide the kind of intelligent systems that will make all the work, and all the introspection, worthwhile? Byte Magazine, 16 (1):239–246, 242, 244, 246, January 1991. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Ryan:1991:DST

[Rya91b] B. Ryan. The data swamp: Together, hardware and software developments are making it easier to handle massive amounts of data. Byte Magazine, 16(5):153– 156, May 1991. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Ryan:1991:DBB

[Rya91c] B. Ryan. Downsizing: Bane or boon: Downsizing saves money, increases efficiency, and changes the role of MIS. Byte Magazine, 16(9):227–230, 232, September 1991. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Ryan:1991:MSP

[Rya91d] B. Ryan. Multiprocessor surf’s up: A parallel technology that’s gone from real soon now to here now! Byte Magazine, 16(6):199– 202, 204, 206, June 1991. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Ryan:1991:FTF

[Rya91e] Bob Ryan. On the fast track: Frame-relay services are creating a stir among network managers. here’s why. Byte Magazine, 16(12):361–??, November 1991. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Ryan:1991:PW

[Rya91f] Bob Ryan. Processor wars. Byte Magazine, 16(11):85–??, ?? 1991. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Ryan:1993:CGP

[Rya93a] Bob Ryan. Communications get personal: AT&T’s Hobbit powers a new generation of personal communications devices. Byte Magazine, 18(2):169–176, February 1993. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Ryan:1993:RDP

[Rya93b] Bob Ryan. RISC drives PowerPC: The PowerPC puts all the best features of RISC — pipelining, branch prediction, and plenty of registers — into a scalable, low-price package. Byte Magazine, 18(9):79–??, August 1993. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Ryan:1994:ARH

[Rya94a] Bob Ryan. Alpha rides high:

The 21164 is head and shoulders above the rest. Byte Magazine, 19(10):197–??, October 1994. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Ryan:1994:MCP

[Rya94b] Bob Ryan. M1 challenges Pentium: Cyrix will compete with Intel’s Pentium using an innovative 80x86 superscalar processor. Byte Magazine, 19(1):83–??, January 1994. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Rydberg:1993:BSN

[Ryd93]  John Rydberg.  Banyan’s

“StreetTalk for NetWare”: Banyan’s distributed directory meets server-based NetWare. Byte Magazine, 18(6):199–??, May 1993. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Sachs:1990:CSC

[Sac90] E. Sachs. Coming soon to a CAD lab near you: 3-Draw lets you sketch out your ideas in 3-D. Byte Magazine, 15(7):238– 239, July 1990. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Saffo:1993:CSD

[Saf93]  Paul Saffo.  A conspiracy

of silence: The dangers of electromagnetic-field radiation are evident. so why isn’t the industry doing anything? Byte Magazine, 18(8):278–??, July 1993. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Salus:1994:UAQ

[Sal94] Peter H. Salus. Unix at 25: For a quarter of a centry, Unix has defined the design of commercial operating systems. here’s the story of what gives Unix its staying power. Byte Magazine, 19 (10):75–??, October 1994. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Sandler:1990:NFF

[San90a] C. Sandler. The new 486s: Are faster FPUs enough: Spear and Dyna Micro’s 25-MHz i486-based systems capitalize on the chip’s integrated FPU. Byte Magazine, 15(7):160–163, July 1990. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Sandler:1990:NCM

[San90b] C. Sandler. New controller makes SCSI palatable to PCs: Distributed processing technology’s SmartConnex/ISA hides SCSI’s incompatibility from PCs. Byte Magazine, 15(12):205–206, 208, November 1990. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Satchell:1990:MFN

[Sat90] S. Satchell. Megafloppies: Four new technologies are in the bidding to be the next standard A drive. Byte Magazine, 15(10): 301–302, 304, 306, 308, 310, October 1990. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Sharp:1994:MMM

[SB94] Oliver Sharp and David F. Bacon. Measure for measure: The more you know about benchmarks, the more you can use them intelligently when evaluating systems. Byte Magazine, 19(10):65–66, 68, 70, 72, October 1994. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Sheldon:1991:WWL

[SBS91] K. M. Sheldon, J. J. Barron, and B. Smith. Window wars: The leading graphical user interfaces go head to head. Byte Magazine, 16(6):124–132, 134, June 1991. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Sejnowski:1992:SBI

[SC92] T. J. Sejnowski and P. S. Churchland. Silicon brains: Innovative computer devices are coming from studies of the human brain. Byte Magazine, 17(10):137–??, October 1992. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Scapicchio:1992:LGL

[Sca92] Mark Scapicchio. Looking good for less (windows word processors). Byte Magazine, 17(11): 104–106, 109–110, 112, November 1992. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Schmidt:1990:NPS

[Sch90a] Paul Schmidt. Notebook PCs set the portable standard. Byte Magazine, 15(11):153–156, 158, Fall 1990. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Schulman:1990:UUP

[Sch90b] Andrew Schulman. DOS unbounded: uses of protected mode. Byte Magazine, 15(11): 250–256, 1990. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Schulman:1990:GPC

[Sch90c] Andrew Schulman. Glockenspiel puts C++ to work: Common-View applies C++ to graphical user interface programming. Byte Magazine, 15(1):213–217, January 1990. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Schulman:1991:UVD

[Sch91a] Andrew Schulman. Undocumented DOS: Valuable details about MS-DOS. Byte Magazine, 16(3):287–288, 290, 292, 294, 296–298, March 1991. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Schwerdtfeger:1991:MGT

[Sch91b] R. S. Schwerdtfeger. Making the GUI talk: New technology addresses the GUI access barrier for visually impaired and learningdisabled people. Byte Magazine, 16(13):118–120, 122, 124, 126–128, December 1991. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Schatzman:1992:BOW

[Sch92a] Bruce D. Schatzman. Beyond DOS. an objective way to compute: The object-based model in Windows NT forms the foundation for Microsoft’s future operating systems. Byte Magazine, 17(12):301–??, November 1992. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Schifreen:1992:PPD

[Sch92b] R. Schifreen. Practical PC data security. Byte Magazine, 17 (8):94IS–23–24, 94IS–26, 94IS– 28, 94IS–30, August 1992. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Schnapp:1992:ARX

[Sch92c]  M. Schnapp.  Arago raises

the Xbase ante: With Arago Professional providing a dBasecompatible interpreter and a compiler, Wordtech becomes a dark horse in the Xbase race. Byte Magazine, 17(7):273–274, 278, 280, July 1992. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Schatzman:1993:BNG

[Sch93a] Bruce D. Schatzman. Beyond DOS: Next-Generation OLE: A faster, simpler OLE looms on the horizon. Byte Magazine, 18 (3):209–??, March 1993. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Schmidt:1993:DDF

[Sch93b] Robert Schmidt. Dynamic documents: Folio views 3.0 comes to Windows and brings with it some exciting new features, including an open client/server architecture, concurrent multiuser editing, embedded graphics and multimedia support. Byte Magazine, 18(8):145–146, 148, July 1993. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Schneier:1993:DS

[Sch93c] Bruce Schneier. Digital signatures. Byte Magazine, 18(??): ??, ?? 1993. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Schneier:1993:UHD

[Sch93d] Bruce Schneier. Under the hood:

Digital signatures: Digital signatures will enable electronic documents to serve as legal instruments. Byte Magazine, 18(12): 3309–??, November 1993. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Scott:1993:FLA

[Sco93] Bernard E. Scott. The five layers of ambiguity. Byte Magazine, 18(1):??, January 1993. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Sproule:1991:AIP

[SE91] William Sproule and Jon Edwards. AppleTalk over the Internet: Princeton and Penn State interconnect their AppleTalk LANs. Byte Magazine, 16(7): 293–294, 296, 298, July 1991. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Smith:1993:GEF

[SE93a] Ben Smith and Howard Eglowstein. In good electronic form: Two companies prove how valuable electronic forms can be. Byte Magazine, 18(12):67–68, 70, 72, 74, November 1993. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Smith:1993:SFP

[SE93b] Ben Smith and Howard Eglowstein. Solutions focus: Putting your data on the map: Connecting maps with data helps visualize information. Byte Magazine, 18(1):188–??, January 1993. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Smith:1994:SAE

[SE94] Ben Smith and Howard Eglowstein. Scheduling across the enterprise: Group-scheduling software relieves some of the tedium and frustration of corporatewide scheduling by accessing the combined calendars of all the people across an organization. the BYTE Lab evaluates six cross-platform schedulers. Byte Magazine, 19(6): 216–220, 222, June 1994. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Seifert:1991:ETY

[Sei91] R. Seifert. Ethernet: Ten years after: Rich Scifert, one of Ethernet’s designers, talks about its first 10 years. Byte Magazine, 16 (1):315–316, 318, 320–321, January 1991. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Seligman:1992:TPC

[Sel92] Daniel Seligman. Trivial pursuits: Computing diversions offer their own rewards. Byte Magazine, 17(10):348–??, October 1992. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Seybold:1993:SBL

[Sey93] Patricia Seybold. Stop bit:

The learning organization: Distributed computing won’t work unless companies are willing to change. Byte Magazine, 18 (4):264–??, April 1993. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Simmel:1992:OSO

[SG92a] Sergiu S. Simmel and Ivan Godard. Objects of substance: Object storage is a natural for persistent-data servers. Byte Magazine, 17(14):167–??, December 1992. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Smith:1992:SFS

[SG92b] Ben Smith and Raymond Ga Cote. Solutions focus. stalking the ultimate workstation: Eight powerhouses reviewed. Byte Magazine, 17(12):192–??, November 1992. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Stone:1990:DWR

[SH90] C. M. Stone and D. Hentchel.

Database wars revisited: The relational and object-oriented camps do battle for database honors. Byte Magazine, 15(10): 233, 236, 238, 240, 242, October 1990. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Simonyi:1991:HRD

[SH91] C. Simonyi and M. Heller. The Hungarian revolution: A developing standard for naming program variables. Byte Magazine, 16(8):131–132, 134–138, August 1991. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Shapiro:1994:CFB

[Sha94a] Fred Shapiro. Commentary: The first bug: Exposing the myth of the first bug. Byte Magazine, 19(4):308–??, April 1994. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Sharp:1994:NPE

[Sha94b] Oliver Sharp. NT programming’s early leader: Microsoft’s Visual C++ 32-bit edition shows flaws, but overall, it delivers effective tools for Windows NT programming and for porting 16-bit Windows applications. Byte Magazine, 19(1):189–??, January 1994. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Sharp:1994:CPC

[Sha94c] Olvier Sharp. Compilers for parallel CPUs: In converting applications for parallel processing, how much of the job can the compiler do without programmer intervention? Byte Magazine, 19 (2):97–??, February 1994. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Sheldon:1990:MEB

[She90a] Kenneth M. Sheldon. Micro Edsels: Besides winners, we’ve seen our share of duds in 15 years. Byte Magazine, 15(2):245–248, February 1990. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Shetler:1990:BBM

[She90b] Tim Shetler. Birth of the BLOB:

Multimedia databases and “binary large objects” will revise the way you store, access, and manipulate information. Byte Magazine, 15(2):221–222, 224, 226, February 1990. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Sheldon:1991:AGG

[She91] K. M. Sheldon. ASCII goes global: Computer companies and international committees wrestle with making ASCII large enough to cover the globe. Byte Magazine, 16(7):108–112, 114, 116, July 1991. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Shepard:1992:LVR

[She92] Jeffrey D. Shepard. Lower the voltage, raise the power. Byte Magazine, 17(14):137–??, December 1992. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Sheldon:1993:LIW

[She93a] Kenneth Sheldon. Lotus Improv for Windows: A radical application comes to Windows. Byte Magazine, 18(2):52–??, February 1993. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Sheldon:1993:DYK

[She93b] Kenneth M. Sheldon. Do you know where your money is? Microsoft Money and Quicken for Windows let you get a grip on your personal-and not so personal-finances. Byte Magazine, 18(5):75–76, 78, 80, Spring 1993. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Shepard:1993:JLF

[She93c] Jeffrey D. Shepard. Japanese leaders in fuzzy logic. Byte Magazine, 18(4):??, April 1993. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Sheldon:1994:MBC

[She94] Tom Sheldon. MAPI blooms in Chicago. Byte Magazine, 19(11): 163–??, November 1994. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Singleton:1994:SAR

[Sin94] Andy Singleton. Some assembly required: Genetic programming with C++: With an appropriate algorithmic structure, you can develop programs that are self-modifying and evolving. Byte Magazine, 19(2):171–??, February 1994. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Schatzman:1992:LMG

[SL92] Bruce D. Schatzman and Jeffrey H. Lubeck. LAN manager gets a foot in the door: What the new capabilities mean, and how to integrate LAN manager with NetWare. Byte Magazine, 17 (2):265–??, February 1992. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Sloman:1991:CCP

[Slo91] J. Sloman. Control central:

Providing centralized LAN-based services often involves major changes at each client’s workstation. Byte Magazine, 16(3):175– 176, 178, 180, March 1991. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Smarte:1990:G

[Sma90] Gene Smarte. Guideposts. Byte Magazine, 15(11):8–??, Fall 1990. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Smarte:1991:RPI

[Sma91a] Gene Smarte. R.I.P. IBM. Byte Magazine, 16(11):8–??, ?? 1991. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Smarte:1991:SS

[Sma91b] Gene Smarte. Surveys say.

Byte Magazine, 16(11):143–??, ?? 1991. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Smarte:1992:PAR

[Sma92] Gene Smarte. PowerExec: AST research’s new notebook features two PCMCIA 2.0 slots. Byte Magazine, 17(13):8–??, ?? 1992. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Smarte:1993:FIT

[Sma93a] Gene Smarte. First impressions: Two Toshiba systems to go: The T4500 notebook and the dynapad T100X pen-based computer. Byte Magazine, 18(1):46–??, January 1993. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Smarte:1993:TGA

[Sma93b] Gene Smarte. Toshiba gets aggressive with passive color: Toshiba’s T1900C could change the way you look at passivematrix color displays. Byte Magazine, 18(8):32–??, July 1993. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Smith:1990:AOSb

[Smi90a] B. Smith. Alternative operating systems, part 2: From a tiny kernel: When you’re building a real-time operating system, it helps to start small. Byte Magazine, 15(9):423–424, 26, September 1990. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Smith:1990:AOSd

[Smi90b] B. Smith. Alternative operating systems, part 4: Pick: OS or DBMS: What do you get when you build an operating system around a database? Byte Magazine, 15(12):381–382, 384, November 1990. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Smith:1990:AWT

[Smi90c] B. Smith. Around the world in text displays: Computing with nonroman characters presents some formidable obstacles. Byte Magazine, 15(5):262–266, 268, May 1990. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Smith:1990:BUB

[Smi90d] B. Smith. The BYTE Unix benchmarks: Before you jump into the Unix pool, see how your favorite system stacks up against the rest of the pack. Byte Magazine, 15(3):273–277, March 1990. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Smith:1990:MWB

[Smi90e]  B. Smith.  Microsoft Word

brings PC-Style word processing to Unix: Unix word processing takes a turn for the better. Byte Magazine, 15(13):209–210, December 1990. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Smith:1990:PID

[Smi90f] B. Smith. Personal Iris: The dream maker: A living-color look at the Silicon Graphics Personal Iris 4D/25. Byte Magazine, 15 (7):174–176, 178, July 1990. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Smith:1990:ASW

[Smi90g] Ben Smith. Apollo shrinks the workstation price tag: Apollo introduces the world’s most affordable workstation. Byte Magazine, 15(1):94–??, January 1990. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Smith:1991:AOS

[Smi91a] B. Smith. Alternative operating systems, part 6: FlexOS’s muscle: Digital Research’s FlexOS closes out our series. Byte Magazine, 16(1):323–326, January 1991. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Smith:1991:PEW

[Smi91b] B. Smith. Printing with electrons: Will electron-beam printing capture the high-volume market? Byte Magazine, 16(10): 185–186, 188, 190, 192, October 1991. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Smith:1993:BMT

[Smi93a] Ben Smith. A beefier MKS toolkit: MKS Toolkit 4.1 is a bigger and better collection of Unix tools for DOS and OS/2. Byte Magazine, 18(4):191– ??, April 1993. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Smith:1993:DDE

[Smi93b] Ben Smith. Data from the depths: Engineers at the Woods Hole Oceanographic Institution deploy inexpensive, autonomous data loggers on small underwater vessels. BYTE Lab editors study the Woods Hole solution and the latest trend in data acquisition. Byte Magazine, 18(8): 69–??, July 1993. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Smith:1993:GEP

[Smi93c] Ben Smith. Going to extremes:

With a powerful new graphics processor, Silicon Graphics’ Indigo2 was born for data visualization. Byte Magazine, 18(7): 34–??, June 1993. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Smith:1993:HHU

[Smi93d] Ben Smith. A Heavy-Hitting Unix for PCs: Solaris for Intel is the most complete PC Unix. Byte Magazine, 18(6):37– ??, May 1993. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Smith:1993:NIN

[Smi93e] Ben Smith. NextStep for Intel: Next is back, but not in black. NextStep for Intel processors brings Next’s strong objectoriented environment to PCs. ben tests NextStep 3.1, concluding that Next’s move to “white” hardware was a wise move. Byte Magazine, 18(9):141– ??, August 1993. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Smith:1994:BPD

[Smi94a] Ben Smith. Blazing the path: DEC’s LinkWorks delivers a multiplatform — Unix, Open-VMS, PC, and Macintosh — work-flow system. if groups in your organization collaborate on the creation and development of documents, images, or data, LinkWorks can provide an effective set of tools for automating your most complex work-flow tasks. Byte Magazine, 19(8):147– 148, 150, August 1994. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Smith:1994:CSM

[Smi94b] Ben Smith. Client/server made easy: NobleNet’s EZ-RPC takes the sting out of writing client/ server applications. whether converting existing stand-alone programs of starting from scratch, applications programmers can write distributed applications without any network programming skills. Byte Magazine, 19 (3):143–??, March 1994. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Smith:1994:DMP

[Smi94c]  Ben Smith.  Digital-media

power: Imaginative packaging and start-up software add some fun to SGI’s new low-price workstation. the fun doesn’t detract from the Indy’s computing price/performance, 2-D graphics strengths, and ability to work with both Macs and PCs. Ben Smith’s hands-on testing finds the new Indy serious about 2-D graphics and SGI’s concept of digital media. Byte Magazine, 19 (1):183–??, January 1994. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Smith:1994:HTL

[Smi94d] Ben Smith. From here to there:

A look at the growing popularity of the Point-to-Point Protocol. Byte Magazine, 19(6):271– ??, June 1994. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Smith:1994:MPS

[Smi94e] Ben Smith. Multiprotocol print server: A US$695 multiprotocol print server from Axis Communications connects as many as three printers directly to Ethernet LANs, providing equal access to PCs, Macs, and Unix machines. the NPS 550’s virtual printer technology simplifies configuration and administration. Byte Magazine, 19 (4):181–??, April 1994. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Smith:1994:NPS

[Smi94f] Ben Smith. Network problem solver: Fluke’s LANMeter 675 tests and analyzes protocols, networking hardware, and cables. this tool can quickly isolate problems on both Ethernet and Token Ring LANs. Byte Magazine, 19(11):247–248, 250, November 1994. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Smith:1994:PWA

[Smi94g] Ben Smith. Power workstation at a Pentium price: The HP 9000 Series 700 Model 712/60 uses innovative hardware packaging and the superscalar PA-7100LC CPU to achieve breakthroughs in low cost and high performance for a Unix workstation. Byte Magazine, 19(7):161– ??, July 1994. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Spicer:1990:CPZ

[Spi90a] S. Spicer. C++, plus: Zortech’s C++ 2.0 may make an OOP believer of you yet. Byte Magazine, 15(7):186, 188, 190–191, July 1990. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Spicer:1990:OOC

[Spi90b] S. Spicer. Object-Oriented C that goes VROOMM: Borland’s Turbo C++ promises to bring object-oriented programming to the masses. Byte Magazine, 15 (10):186, 188–189, October 1990. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Stallings:1994:SVS

[SS94] William Stallings and Ben Smith.

SNMP version 2: SNMP is maturing as evidenced by the added functionality of SNMP 2. Byte Magazine, 19(8):191– ??, August 1994. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Slivka:1993:UHI

[SSF93] Benjamin W. Slivka, Eric Straub, and Richard Freedman. Under the hood: Inside MS-DOS 6: MS-DOS 6’s designers examine the inner workings of MemMaker and DoubleSpace. Byte Magazine, 18(8):197–??, July 1993. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Shapiro:1993:SAR

[ST93] Eric Shapiro and Tom Thompson. Some assembly required: The Mac extended: Savvy programmers can write their own Mac extensions. Byte Magazine, 18(8):205–??, July 1993. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Stallings:1990:MWT

[Sta90] W. Stallings. Managing the Well-Tempered LAN: ISO standards signal that network management help is on the way. Byte Magazine, 15(4):275–278, 280–283, April 1990. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Stallings:1991:PGQ

[Sta91a] W. Stallings. A practical guide to queuing analysis: William Stallings, an authority on data communications, presents some simple tools for analyzing many kinds of problems. Byte Magazine, 16(2):309–316, February 1991. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Stallings:1991:FPN

[Sta91b] William Stallings. Faster packet networks: William Stallings explains the workings of a new and exciting data communications technology. Byte Magazine, 16(12):173–??, November 1991. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Smith:1992:SCB

[STA92] Ben Smith, Tom Thompson, and Steve Apiki. Software corner bash, the Bourne Again Shell: A free Korn-shell replacement, a JPEG viewer for Macs, and a Windows file utility. Byte Magazine, 17(12):299–??, November 1992. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Stallings:1993:UHF

[Sta93] William Stallings. Under the hood FDDI speaks: The FDDI II standard mixes voice and data on a single medium. Byte Magazine, 18(4):197–??, April 1993. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Stallings:1994:PGP

[Sta94a] William Stallings. Pretty Good Privacy: Privacy and security are important issues to commercial users of public E-mail systems. PGP, an E-mail security package, is finding acceptance as the way to achieve protection. Byte Magazine, 19(7):193– ??, July 1994. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Statt:1994:IVJ

[Sta94b] Paul Statt. Intel/VLSI join the PDA fray: The Intel/VLSI polar chip set brings the 80x86 architecture to the PDA realm. Byte Magazine, 19(1):101–??, January 1994. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Stewart:1990:DDT

[Ste90] G. A. Stewart. A database developer that’s different from the rest: The Clarion Professional Developer provides a complete development environment. Byte Magazine, 15(10):201–202, October 1990. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Stein:1991:BTT

[Ste91a] R. M. Stein. Browsing through terabytes: Wide-area information servers can distill vast archives of data. Byte Magazine, 16(5):157–160, 162–164, May 1991. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Stein:1991:RAL

[Ste91b] R. M. Stein. Real artificial life: Nature’s skill and craftsmanship, when harnessed toward the creation of artificial life, presents a virtually unlimited reservoir of possibilities for engineering solutions. Byte Magazine, 16(1): 289–290, 292, 294, 296–298, January 1991. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Stein:1991:SGM

[Ste91c] R. M. Stein. Scaling up: Get the message: Message-passing multicomputers bypass shared memory. Byte Magazine, 16(6):231– 234, 236, 238–240, June 1991. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Stephenson:1991:CWW

[Ste91d] P. Stephenson. Create a WAN: What you need to know and do to implement a wide-area network in your business. Byte Magazine, 16(7):169–176, 178, July 1991. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Stephenson:1991:MML

[Ste91e] P. Stephenson. Mixing and matching LANs: The thorniest network management problem is how to tie different networks together. Byte Magazine, 16(3):157–158, 160, 162, 164, March 1991. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Stein:1992:SFD

[Ste92] R. M. Stein. Safety by formal design. Byte Magazine, 17(??):157, August 1992. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Stein:1994:ODO

[Ste94a] Richard Marlon Stein. Object databases: Object-oriented DBMSes solve the problem of storing persistent data from an object-oriented application. Byte Magazine, 19(4):74–??, April 1994. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Stewart:1994:FFB

[Ste94b] George Stewart. Forecasting the future: Business forecasting software predicts future trends — sales projections, inventory levels, energy usage — by incorporating traditional forecasting methods. the four programs reviewed here guide businesspeople through a forecasting session and generate accessible reports and graphs. Byte Magazine, 19 (6):237–240, June 1994. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Stoll:1993:CCE

[Sto93] Clifford Stoll. Commentary: The Cuckoo’s Egg revisited: Five years after a hacker broke into a computer at Lawrence Berkeley Laboratory, the repercussions still echo across the Internet. Byte Magazine, 18(7):274– ??, June 1993. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Stading:1992:CBM

[STS92] James Stading, Tom Thompson, and Ben Smith. Creating Bit-Mapped buttons: A Windows file manager, Telnet for Macs, and dired sans emacs. Byte Magazine, 17(10):275–??, October 1992. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Stys:1991:PWP

[Sty91] P. K. Stys. Putting waveforms to paper: Here’s how to get data from a Mac screen into a file or printout. Byte Magazine, 16(1): 339–342, 344, January 1991. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Smith:1993:LAF

[SU93] Ben Smith and Jon Udell. Linking aids: The four key issues of connecting networks to one another are bandwidth, protocols, management, and cultural conflicts. Byte Magazine, 18(13): 66–??, December 1993. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Summers:1990:PDE

[Sum90] E. Summers. ES: A public domain expert system: Develop your own expert systems — or experiment with some sample knowledge bases — with this free program. Byte Magazine, 15(10): 289–290, 292, October 1990. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Swart:1993:CHY

[Swa93] Edward R. Swart. Commentary: “How are you at interfacing?”: Computers are rapidly changing the English language. Byte Magazine, 18(13):302–??, December 1993. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Smith:1990:BBU

[SY90] B. Smith and T. Yager. Battle for the best Unix V/386: New releases from Interactive Systems and The Santa Cruz Operation. Byte Magazine, 15(13):206–207, December 1990. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Tabaza:1993:RJL

[Tab93] Khaldoon Tabaza. Report from Jordan: Localizing software in Arabia. Byte Magazine, 18(8): 48–??, July 1993. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Tazelaar:1991:VTW

[Taz91] J. M. Tazelaar. Visions of tomorrow: What will the office of the future be like? Byte Magazine, 16(9):207–208, 210, September 1991. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Thompson:1991:PPB

[TG91] T. Thompson and R. Grehan. Paint for the pros: The BYTE Lab tests eight professional-level paint programs for Macs and PCs. Byte Magazine, 16(6):258– 261, 264, 266–268, 272, 274, 276, June 1991. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Thompson:1994:BCRa

[TGN94] Tom Thompson, Raymond Ga Cote, and Michael Nadeau. Books and CD-ROMs: The genesis of the Mac: The birth of a computer, an astronomy CD-ROM, legal requirements in business, and laptop repair. Byte Magazine, 19(2):41–??, February 1994. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Thompson:1992:FIA

[TH92] Tom Thompson and Tom R. Halfhill. First impressions: Apple’s performas: Macs for the home: Apple tackles the home market. Byte Magazine, 17 (10):52–??, October 1992. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Thompson:1993:FIS

[TH93] Tom Thompson and Tom R. Halfhill. First impressions: A spring harvest of Apple Macintoshes: With street prices of about US$1000 to US$4500, these new Macs are priceconscious and powerful. Byte Magazine, 18(3):40–??, March 1993. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Thelen:1994:UHP

[The94]  Randy Thelen.  Under the

hood: The Power Mac’s Run-Time architecture: The RISCbased Power Mac uses a dramatically different application architecture that provides compatibility with past applications and future applications. Byte Magazine, 19(4):131–??, April 1994. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Thompson:1993:EUR

[THN93] T. Thompson, T. R. Halfhill, and M. Nadeau. Ease of use is relative. Byte Magazine, 18(11):89– 90, 92, 94, October 1993. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Thompson:1990:DQM

[Tho90a] T. Thompson. A digital “Quill”

for Mac video displays: Data Translation’s VideoQuill combines text, graphics, and video. Byte Magazine, 15(13):229–231, December 1990. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Thompson:1990:DTC

[Tho90b] T. Thompson. Director takes charge of Mac multimedia: MacroMind’s new Director 2.0 provides interactive control and input from more audio and video sources. Byte Magazine, 15(10): 178–180, October 1990. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Thompson:1990:MAMb

[Tho90c] T. Thompson. The Mac at 40 MHz: The Mac IIfx is a powerful number cruncher in the Mac or Unix environment. Byte Magazine, 15(9):162–164, September 1990. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Thompson:1990:MAMa

[Tho90d] T. Thompson. Mac at the minimum: Some suggestions and hints for running all you can on your 1-megabyte Mac. Byte Magazine, 15(3):219–222, 224, March 1990. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Thompson:1990:MMCb

[Tho90e] T. Thompson. More 16-Million-Color fireworks: SuperMac Technology’s ColorCard/24 and Spectrum/24 series III video boards. Byte Magazine, 15(6):174, 176, June 1990. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Thompson:1990:MMCa

[Tho90f] T. Thompson. Motorola’s 68040 microprocessor: This CISC processor for the 1990s offers new features and boosted performance. Byte Magazine, 15(2): 96A–96C, February 1990. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Thompson:1990:NUM

[Tho90g] T. Thompson. The Norton Utilities for the Mac: More than just warmed-over versions of the Norton Utilities for DOS. Byte Magazine, 15(13):179–??, December 1990. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Thompson:1990:SSN

[Tho90h] T. Thompson. Svelte scanner is no fistful of dollars: Sharp’s lowcost scanner delivers high-quality color images to those who can afford to wait. Byte Magazine, 15(4):151–152, 154, April 1990. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Thompson:1990:SSL

[Tho90i] Tom Thompson. Stranger in a strange land (Macintosh file sharing). Byte Magazine, 15(11):89– 92, 95, Fall 1990. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Thompson:1991:OCB

[Tho91a] B. Thompson. Overturning the category bucket: Categorizing knowledge is one of the primary ways that an AI system can acquire “understanding.”. Byte Magazine, 16(1):249–250, 252, 254–256, January 1991. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Thompson:1991:DPB

[Tho91b] T. Thompson. Desktop PCs: The buyer’s market continues. Byte Magazine, 16(11):157–164, 166–168, ?? 1991. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Thompson:1991:MVR

[Tho91c] T. Thompson. Macintosh video revealed: Tom Thompson, BYTE’s resident Mac guru, takes us on a trek to the depths of the Mac display. Byte Magazine, 16 (4):305–308, 310, 312–314, 390– 391, April 1991. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Thompson:1991:WFM

[Tho91d] T. Thompson. A Wallet-Friendly Mac that delivers performance: The high-performance Mac IIsi. Byte Magazine, 16(3):257–258, 260, 262, March 1991. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Thompson:1991:NSS

[Tho91e] Tom Thompson. Networking with System 7.0: System 7.0 provides peer-to-peer capabilities. Byte Magazine, 16(9): 337–??, September 1991. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Thompson:1992:CAR

[Tho92a] T. Thompson. Color at a reasonable price: The BYTE Lab tests seven color PostScript printers for PCs and Macs. Byte Magazine, 17(6):316–324, June 1992. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Thompson:1992:PIF

[Tho92b] T. Thompson. The Phaser III fires dazzling colors: Tektronix’s dazzling new color printer. Byte Magazine, 17(1):285–286, January 1992. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Thompson:1992:UMO

[Tho92c] Tom Thompson. A/UX 3.0:

The Mac’s odd couple: Apple’s latest version of Unix. Byte Magazine, 17(9):283–??, September 1992. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Thompson:1992:FIN

[Tho92d] Tom Thompson. First impressions: New Macs for the desktop and road: Apple is updating computers and introducing new ones as rapidly as possible. Byte Magazine, 17(14):44–??, December 1992. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Thompson:1992:FSA

[Tho92e] Tom Thompson. The future of System 7.0: Apple’s plans for a smooth transition to RISC. Byte Magazine, 17(10):182–??, October 1992. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Thompson:1992:MR

[Tho92f] Tom Thompson. The Mac on the road. Byte Magazine, 17(13): ??, ?? 1992. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Thompson:1992:PIS

[Tho92g] Tom Thompson. The Phaser II SD prints dazzling dyes: Tektronix’s new dye sublimation printer costs under US$10,000. Byte Magazine, 17(14):217–??, December 1992. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Thompson:1992:RAM

[Tho92h] Tom Thompson. Racing at 33 MHz: Quadra 950 and Radius Rocket 33: Apple’s most powerful Macintosh and Radius’s latest accelerator board provide ample computing power. Byte Magazine, 17(10):241–??, October 1992. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Thompson:1993:AMM

[Tho93a] Tom Thompson. Apple’s midrange Mac a heavy hitter: Benchmark results and hands-on experience with two of Apple’s newest Macs, the Centris 650 and Power-Book 165c. Byte Magazine, 18 (7):129–130, June 1993. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Thompson:1993:BCP

[Tho93b] Tom Thompson. Bargain color printers: Inexpensive color printing from Star Micronics, Hewlett–Packard, and Fargo. Star Micronics’ SJ-144 and Fargo’s Primers break the US$1000 price barrier for thermal-transfer color printing by leaving the image processing to the host system. Byte Magazine, 18(13):153–154, 156, 158, 160, December 1993. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Thompson:1993:BRU

[Tho93c] Tom Thompson. BASIC for the rest of us: Zedcor’s FutureBasic provides a powerful development environment for the Macintosh that’s actually easy to use. Thompson builds a project with FutureBasic and a helpful addon, PG:PRO. Byte Magazine, 18(7):143–??, June 1993. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Thompson:1993:PLA

[Tho93d] Tom Thompson. PostScript Level 2: Adobe takes the driver’ seat: You may already have a PostScript Level 2 printer, but now you can take advantage of it with Level 2 drivers. Byte Magazine, 18(7):28–??, June 1993.

CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Thompson:1993:PPN

[Tho93e] Tom Thompson. PowerBook peripherals: New hardware makes your Apple notebook more useful. Byte Magazine, 18(4):173– ??, April 1993. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Thompson:1993:PPL

[Tho93f] Tom Thompson. PowerPC performs for less: Will your next desktop PC be RISC-based? the PowerPC 601 has the performance, low cost, and support for multiple operating systems needed to make that a possibility come true; PowerOpen gives users freedom of choice; the PowerPC does Windows; Pentium Out-Powered; translation tool ports programs in a flash. Byte Magazine, 18(9):56–??, August 1993. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Thompson:1993:QSC

[Tho93g] Tom Thompson. QMS strikes with color laser printer: ColorScript Laser 1000 brings color laser printing into a more affordable price range. Byte Magazine, 18(8):24–??, July 1993. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Thompson:1993:RNR

[Tho93h] Tom Thompson. Reviewer’s notebook: Retooling a classic: The BYTE Lab updates its Macintosh benchmarks. Byte Magazine, 18(2):221–??, February 1993. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Thompson:1993:TPG

[Tho93i] Tom Thompson. Two Power-Books great and small: The PowerBook 180 and PowerBook duo 230 show different design directions. Byte Magazine, 18 (3):173–??, March 1993. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Thompson:1994:LCC

[Tho94a] T. Thompson. Low-cost color lasers. Byte Magazine, 19 (12):139–140, 142, 144, December 1994. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Thompson:1994:ARN

[Tho94b] Tom Thompson. Apple redefines the notebook: The latest PowerBooks set a new standard for notebook computers: built-in Ethernet, an innovative trackpad, optional PCMCIA expansion, 16-bit color, stereo sound, and a fast 68040 processor upgradable to PowerPC. Tom Thompson tests the new PowerBooks. Byte Magazine, 19 (8):143–145, August 1994. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Thompson:1994:BCRb

[Tho94c] Tom Thompson. Budget CD recording: Cutting your own CD-ROMs has become affordable now that CD-R drives with recording software go for around US$4000. Tom Thompson tests three, from JVC, Microboards, and Pinnacle Micro, on both Mac and PC platforms. Byte Magazine, 19(3):145–??, March 1994. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Thompson:1994:MSU

[Tho94d] Tom Thompson. Mac SCSI utility sampler: Two SCSI utilities that let you attach third-party SCSI drives to your Mac. Byte Magazine, 19(10):159–??, October 1994. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Thompson:1994:MAM

[Tho94e] Tom Thompson. The Macintosh at 10: On the Mac’s tenth birthday, BYTE looks at its profound influence on computing. Byte Magazine, 19(2):47–??, February 1994. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Thompson:1994:NMB

[Tho94f] Tom Thompson. New Mac blazes technology trails: Apple’s new Mac Quadra 840AV makes the move into video and voice communications with a faster CPU, a built-in DSP, video connections, and software for voice recognition and text-to-speech conversion. Byte Magazine, 19(1): 197–200, January 1994. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Thompson:1994:PMC

[Tho94g] Tom Thompson. Power Mac code optimizations: Avoid performance pitfalls in Mac OS. Byte Magazine, 19(11):291–??, November 1994. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Thompson:1994:PNM

[Tho94h] Tom Thompson. Power to the new Macs: Affordable Power Macs run old 680x0 software surprisingly well. the new Macs save their performance punch for a small but growing number of programs compiled for the PowerPC 601 processor. Byte Magazine, 19 (6):209–210, 212, June 1994. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Thompson:1994:SSA

[Tho94i] Tom Thompson. SITcomm is serious: Apple events support is one reason Thompson finds Aladdin Systems’ SITcomm a standout among Mac communications packages. Byte Magazine, 19(5):161–??, May 1994. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Thompson:1994:SST

[Tho94j] Tom Thompson. System 7.5:

A step toward the future: System 7.5 is a significant evolutionary step toward a new Mac OS. Byte Magazine, 19(8):187– ??, August 1994. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Thompson:1994:DAN

[Tho94k] Tom Thompson. “the” debugger is aptly named: This program is an essential tool for developing native PowerPC programs. Byte Magazine, 19(8): 159–??, August 1994. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Toperczer:1991:PPD

[Top91] T. Toperczer. From pyramids to peers: Data management applications strengthen the integrity of network data. Byte Magazine, 16(5):191–192, 194–196, 198, May 1991. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Thompson:1994:AIB

[TR94a] Tom Thompson and Bob Ryan. Apple, IBM bring PowerPC to the desktop: The first PowerPC systems for the desktop provide many pleasant surprises, especially low price. Byte Magazine, 19(4):44–47, 50, 52, 54, 56, 58, 60–63, 66–68, 70, April 1994. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Thompson:1994:PSN

[TR94b] Tom Thompson and Bob Ryan. PowerPC 620 soars: The newest member of the PowerPC family targets the workstation market with fast throughput and speedy floating-point performance. Byte Magazine, 19(11): 113–??, November 1994. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Trask:1990:CVP

[Tra90] Matt Trask. Creating virtual PCs on the 386. Byte Magazine, 15(11):137–140, 142–144, 146, Fall 1990. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Trask:1992:UTU

[Tra92] Matt Trask. Univel’s trim Unix:

A thinner, more attractive Unix. Byte Magazine, 17(10):192–??, October 1992. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Treinish:1993:IMD

[Tre93] Lloyd A. Treinish. Inside multidimensional data. Byte Magazine, 18(4):??, April 1993. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Thompson:1990:SCN

[TS90] Tom Thompson and Ben Smith.

Sizing up the Cube: The NeXT computer – advanced features, fair performance. Byte Magazine, 15(1):169–176, January 1990. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Thompson:1993:ASB

[TS93a] Tom Thompson and Ben Smith. Apple, SGI blaze video trail: With integrated video and sound features, new systems from Apple and Silicon Graphics herald a new era of desktop computers. Byte Magazine, 18(10): 81–??, September 1993. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Trask:1993:PD

[TS93b] M. Trask and D. Sorensen. Personal databases. Byte Magazine, 18(11):124–128, October 1993. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Turner:1990:MBC

[Tur90] S. E. Turner. Modem business: Confused by modern standards like 212A, V.22, and V.32bits? here’s help. Byte Magazine, 15(12):353–354, 356, 358, 360, November 1990. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Udell:1990:BTW

[Ude90a] J. Udell. Bridging troubled waters: Thriving in a diverse computing environment is a lot easier if you have the right tools. Byte Magazine, 15(4):225–226, 228– 230, April 1990. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Udell:1990:CNJ

[Ude90b] J. Udell. CAD and NetWare 386 join forces: IsiCAD’s CADVance 4.0 makes good on the promise of multiuser CAD applications. Byte Magazine, 15(13):182–184, December 1990. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Udell:1990:CDL

[Ude90c] J. Udell. CompuAdd delivers a Low-Cost SPARCstation: The SS1 is a faithful clone. Byte Magazine, 15(13):142–143, December 1990. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Udell:1990:LMF

[Ude90d] J. Udell. LAN manager 2.0: A force to be reckoned with: Microsoft’s network flagship proves it is a viable alternative to NetWare. Byte Magazine, 15 (13):221–222, 224, 226, December 1990. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Udell:1990:MCC

[Ude90e] J. Udell. Macintosh CAD comes of age: Radius’s display-list processor makes Macintosh CAD more competitive. Byte Magazine, 15(5):187–188,190, May 1990. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Udell:1990:NLP

[Ude90f] J. Udell. NetWare 386: less pain, great gain. Byte Magazine, 15(3):160E–F, 160H, 162, March 1990. CODEN BYTEDJ.

ISSN 0360-5280 (print), 1082-7838 (electronic).

Udell:1990:FAM

[Ude90g] J. Udell. OS/2 2.0: It’s a family affair: Microsoft’s long-awaited 32-bit OS/2 forges ahead, with DOS and Windows in tow. Byte Magazine, 15(4):119–120, 122– 123, April 1990. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Udell:1990:TOW

[Ude90h] J. Udell. Three’s the one (windows 3.0). Byte Magazine, 15 (6):122–124, 126, 128, June 1990. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Udell:1991:AEE

[Ude91a] J. Udell. Access Extended Edition and dBASE Data with InfoAlliance: A solution for network users with disparate data sources. Byte Magazine, 16 (4):271–274, April 1991. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Udell:1991:CNM

[Ude91b] J. Udell. Citrix’s new multiuser OS/2: OS/2-based workgroup computing without a LAN. Byte Magazine, 16(1):134–136, 138, January 1991. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Udell:1991:DWS

[Ude91c] J. Udell. Debuggers for Windows shows promise, lacks polish: MultiScope’s latest debugger does Windows programs. Byte Magazine, 16(9):297–298, September 1991. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Udell:1991:PPG

[Ude91d] J. Udell. Powerfusion provides the glue for networking DOS and Unix: One answer to the dilemma of interconnecting DOS and Unix LANs. Byte Magazine, 16(3):267–268, 270, 272, 274, March 1991. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Udell:1991:NTN

[Ude91e] Jon Udell. Networks: Trends in network management. Byte Magazine, 16(11):208–210, 212, 216, 218–220, ?? 1991. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Udell:1991:WWP

[Ude91f] Jon Udell. Whither Windows:

Putting Windows 3.0 and OS/2 2.0 on the scales. Byte Magazine, 16(2):95–??, February 1991. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Udell:1992:MWD

[Ude92a] Jon Udell. Microsoft’s Windows database: Microsoft access, the happy union of SQL and Visual Basic. Byte Magazine, 17 (14):51–??, December 1992. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Udell:1992:UWN

[Ude92b] Jon Udell. Using Windows NT: How will NT affect you and your environment? Byte Magazine, 17 (11):27–??, November 1992. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Udell:1992:WWP

[Ude92c] Jon Udell. Windows for Workgroups: Peer-to-peer networking and more with Windows for Workgroups. Byte Magazine, 17 (12):52–??, November 1992. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Udell:1992:WNC

[Ude92d] Jon Udell. Windows NT up close: Microsoft’s next-generation operating system could live up to its hype. Byte Magazine, 17 (10):167–??, October 1992. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Udell:1993:BCW

[Ude93a] Jon Udell. Beyond DOS: Connecting Windows to data with ODBC: Open Database Connectivity is an API for databaseenabled Windows applications. Byte Magazine, 18(1):271–??, January 1993. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Udell:1993:BEN

[Ude93b] Jon Udell. Beyond DOS: Exploring NetDDE: The power of this peer-to-peer protocol in Windows for Workgroups remains largely unappreciated. Byte Magazine, 18(9):199–??, August 1993. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Udell:1993:BSM

[Ude93c] Jon Udell. Beyond DOS: Simple MAPI delivers: Microsoft’s firstrelease messaging API is easily supported. Byte Magazine, 18(4):211–??, April 1993. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Udell:1993:DCR

[Ude93d] Jon Udell. Desktop CD-ROM publishing: Philips’ CDD521 ushers in the second CD-ROM revolution. Byte Magazine, 18 (1):217–??, January 1993. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Udell:1993:MWS

[Ude93e] Jon Udell. E-Mail from The Workplace Shell. Byte Magazine, 18(3):??, March 1993. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Udell:1993:FIE

[Ude93f] Jon Udell. First impressions:

Easy does it with MS-DOS 6.0: Microsoft adds compression and memory management to the venerable operating system. Byte Magazine, 18(4):44– ??, April 1993. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Udell:1993:ISE

[Ude93g] Jon Udell. India’s software edge: India has contributed programming talent to the microcomputer world. now it wants to contribute products. Byte Magazine, 18(10):55–??, September 1993. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Udell:1993:NGG

[Ude93h] Jon Udell. NetWare goes global: NetWare 4.0 has arrived, claiming support for serious enterprise networking. NetWare, directory service brinks NetWare beyond the LAN, and 4.0 adds other features like file compression, CD-ROM sharing, and data migration. Byte Magazine, 18(8):141– ??, July 1993. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Udell:1993:OTO

[Ude93i] Jon Udell. One thumb up, one thumb down: Release 3 of Lotus Notes delivers long-awaited features, including Macintosh support and full-text indexing, but it lacks development tools needed to build effective groupware. our reviewer finds some significant improvements and some significant disappointments. Byte Magazine, 18(8):161–??, July 1993. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Udell:1993:SPC

[Ude93j] Jon Udell. Start the presses:

CD-ROM publishing comes to the desktop. Byte Magazine, 18 (2):116–??, February 1993. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Udell:1993:VA

[Ude93k] Jon Udell. The Vines advantage. Byte Magazine, 18(3):??, March 1993. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Udell:1993:WWE

[Ude93l] Jon Udell. Windows, Windows everywhere: Microsoft wants you to someday use a version of Windows for every computing platform. the unique demands of each platform, however, might make this goal unrealistic. Byte Magazine, 18(7):72–??, June 1993. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Udell:1993:WNW

[Ude93m] Jon Udell. WorkMan needs work: WorkMan, Reach Software’s work-flow applications development package, is the first system to present a usable model for the creation of work-flow applications. unfortunately, the implementation fails to deliver on the architecture’s promise. Byte Magazine, 18(9):167–??, August 1993. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Udell:1994:AOS

[Ude94a] Jon Udell. Advanced operating systems: Introduction: The great OS debate. Byte Magazine, 19(1):117–??, January 1994. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Udell:1994:BER

[Ude94b] Jon Udell. Beyond DOS: Essential reading: A look at four recent books on Windows NT, Win32, and OLE 2. Byte Magazine, 19(4):237–??, April 1994. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Udell:1994:BVB

[Ude94c] Jon Udell. Beyond DOS: Visual Basic custom controls: Exploring visual C++ 1.5’s powerful OLE support. Byte Magazine, 19(3):197–??, March 1994. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Udell:1994:BWA

[Ude94d] Jon Udell. Beyond DOS: Wide-Area Windows networking: Experimenting with routable protocols for Windows networking. Byte Magazine, 19(1):227– ??, January 1994. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Udell:1994:BWW

[Ude94e] Jon Udell. Beyond DOS: Windows for Workgroups: This is the most advanced version of DOS-based Windows now available. Byte Magazine, 19(2): 181–??, February 1994. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Udell:1994:CWI

[Ude94f] Jon Udell. Commentary: Why IBM should license Win32: To survive as an operating-system vendor, IBM will need a credible Win32 strategy. Byte Magazine, 19(9):278–??, September 1994. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Udell:1994:CCS

[Ude94g] Jon Udell. Componentware:

Component software, as exemplified by Visual Basic’s custom controls, is succeeding where object-oriented computing has failed. Byte Magazine, 19(5):46– ??, May 1994. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Udell:1994:CTM

[Ude94h] Jon Udell. Computer telephony:

The means to link your phone and computer is now available to nearly everyone. Byte Magazine, 19(7):80–82, 86, 88–90, 94, 96, July 1994. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Udell:1994:ECD

[Ude94i] Jon Udell. Exploring Chicago and Daytona. Byte Magazine, 19(11):132–??, November 1994.

CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Udell:1994:FCC

[Ude94j]  Jon Udell.  The fix is in

for Chicago: Chicago will be the most radical Windows upgrade ever. here’s what’s important and improved in Microsoft’s newest operating system. Byte Magazine, 19(9):193–??, September 1994. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Udell:1994:JNA

[Ude94k] Jon Udell. Justifying NT: Available on RISC, CISC, and SMP platforms, advanced server and SQL server make NT a viable choice to provide file and database services to your LAN. Byte Magazine, 19(4):149– ??, April 1994. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Udell:1994:TUT

[Ude94l] Jon Udell. A Taligent update: Taligent’s long-awaited developer’s kit debuts this summer. Byte Magazine, 19(7):183– ??, July 1994. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Udell:1993:ACR

[UE93] Jon Udell and Howard Eglowstein. Affordable CD-R drives. Byte Magazine, 18(2):??, February 1993. CODEN BYTEDJ.

ISSN 0360-5280 (print), 1082-7838 (electronic).

Udell:1994:BCRb

[UFC94] Jon Udell, Rich Friedman, and Rick Cook. Books and CD-ROMs: PowerPC: Cultural and technological perspective: A chronicle of the PowerPC revolution, college selection via CD-ROM, and an encyclopedia of computer cracking via a network. Byte Magazine, 19(9): 41–??, September 1994. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Ullman:1993:CSF

[Ull93] Ellen Ullman. Client/server frees data: Client/server brings data to your desktop. Byte Magazine, 18(7):96–??, June 1993. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Udell:1990:NPL

[UM90] J. Udell and R. Mitchell. Networks of peers: Low-cost alternatives to dedicated server LANs. Byte Magazine, 15(6):142, 144– 146, 148, 150, 152–154, 156, 158, 160, 162, June 1990. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Udell:1990:WSA

[UMR^{+}90] J. Udell, R. Malloy, A. Reinhardt, H. Eglowstein, and G. Bond. Windows shopping: 3.0 applications take shape. Byte Magazine, 15(7):116–118, 120, 122, 124, 126, 128, July 1990.

CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Unger:1990:PPU

[Ung90a] J. Unger. Plug-and-Play Unix machine: Dell’s Intel-based Unix workstation. Byte Magazine, 15(13):213–214, 216, December 1990. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Unger:1990:VAP

[Ung90b] J. Unger. Video analysis on a PC. Byte Magazine, 15(5):64IS/3–4, 6, May 1990. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Unger:1991:AMM

[Ung91] J. Unger. ALR’s multiprocessing monster uses six i486 processors: The MultiAccess system brings multiprocessing to the masses. Byte Magazine, 16 (4):279–282, April 1991. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Upstill:1990:GGD

[Ups90] S. Upstill. Graphics go 3-D: Creating photo-realistic 3-D images is a real challenge. Byte Magazine, 15(13):253–256, 258, December 1990. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Udell:1994:BCRa

[UPSW94] Jon Udell, Hughes Pack, Ben Smith, and Scott Wallace. Books and CD-ROMs: Put fuzzy logic to work: A calculus of everyday concepts, a CD-ROM about the space shuttle, cross-platform connection, and electronic imaging. Byte Magazine, 19(7):51– ??, July 1994. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Urlocker:1990:OOP

[Url90] Z. Urlocker. Object-Oriented programming for Windows: Using the Actor OOP environment to develop applications for Microsoft Windows. Byte Magazine, 15(5):287–292, 294, May 1990. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Udell:1991:TBD

[UT91] J. Udell and T. Thompson. Two bumbling detectives: Dariana Technology Group’s WinSleuth and MacSleuth miss the mark. Byte Magazine, 16(1):228–229, January 1991. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Udell:1991:AAF

[UY91] J. Udell and T. Yager. Atlantix, Altos fill DOS-to-Unix connectivity gaps: Approaching the LAN connectivity problem from different angles. Byte Magazine, 16(4):253–256, April 1991. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

VanDyke:1990:SSE

[Van90] Bruce Van Dyke. SCSI: The I/O standard evolves. Byte Magazine, 15(11):187–191, Fall 1990. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

vanHorn:1991:RGT

[van91] G. A. van Horn. The right graphics tool for the job: Are you using the right graphics programs for your projects? Byte Magazine, 16(7):123–124, 126, 128, 130, July 1991. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

VanHorn:1992:CSA

[Van92a] G. Armour Van Horn. Corel-Draw 3.0: A Swiss army knife for illustrators: More than just an illustration package. Byte Magazine, 17(10):223–224, 226, October 1992. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Vannatta:1992:BBB

[Van92b] R. Vannatta. Borland builds a better Quattro Pro: With Quattro Pro 4.0, Borland builds an even better DOS spreadsheet. Byte Magazine, 17(7):267–268, 270, July 1992. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

VanHorn:1993:CWM

[Van93a] G. Armour Van Horn. CorelDraw 4.0: The word is more: Corel releases an ambitious upgrade. CorelDraw 4.0 features tighter integration across modules, multiple page illustrations, an underlying object database, an impressive bundle of fonts and clip art, path-based animation, and even OCR and forms generation. Van Horn finds out how it all comes together. Byte Magazine, 18(10): 169–??, September 1993. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

VanHorn:1993:NRD

[Van93b] G. Armour Van Horn. A new resolution for desktop lasers: A comparison of the latest 600-dpi laser printers. Byte Magazine, 18(3):156–??, March 1993. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

VanHorn:1993:PVQ

[Van93c] G. Armour Van Horn. Page-Maker 5.0 vs. Quark 3.1: Recent releases of these two popular page-layout packages duke it out both on the Mac and under Windows. Van Horn determines which of these aggressive competitors currently has the upper hand. Byte Magazine, 18(8): 157–158, 160, July 1993. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

VanHorn:1994:FCP

[Van94a] G. Armour Van Horn. Flatbed color professionals: New color flatbed scanners from Agfa, Microtek, and Umax provide a price/performance balance that should appeal to graphics professionals. you may not need your local color service’s drum scanner for image processing. Byte Magazine, 19(8):137–141, August 1994. CODEN BYTEDJ.

ISSN 0360-5280 (print), 1082-7838 (electronic).

VanHorn:1994:WIS

[Van94b] G. Armour Van Horn. Windows illustration: Similarly priced but miles apart in the features they offer, CorelDraw 5 and FreeHand 4 for Windows both hold their own in the competitive Windows design and graphics illustration arena. Byte Magazine, 19(11): 243–246, November 1994. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Varhol:1994:SKH

[Var94a] Peter Varhol. Small kernels hit it big: Microsoft, IBM, USL, and others differ in their opinions on how best to implement microkernel architecture into new operating systems. Byte Magazine, 19 (1):119–??, January 1994. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Varhol:1994:QFA

[Var94b] Peter D. Varhol. QNX forges ahead: New uses for QNX, a venerable, micro-kernel-based operating system. Byte Magazine, 19 (10):199–??, October 1994. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Varhol:1994:VPM

[Var94c] Peter D. Varhol. Visual programming’s many faces: A quick climb through a growing family tree of visual programming products and paradigms. Byte Magazine, 19(7):187–??, July 1994. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Vasconcellos:1993:MRY

[Vas93a] Muriel Vasconcellos. Is MT right for you? Byte Magazine, 18(1): ??, January 1993. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Vasconcellos:1993:OMT

[Vas93b] Muriel Vasconcellos. Overview:

Machine translation: Machine translation is coming to your computer. Byte Magazine, 18 (1):152–??, January 1993. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Vaughn:1990:CWC

[Vau90] F. Vaughn. Color WYSIWYG comes of age: Matching the colors you see on-screen with your printed output. Byte Magazine, 15(13):275–277, 279, December 1990. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

VanName:1990:EMS

[VC90a] M. L. Van Name and B. Catchings. 486 EISA machines: A slow start in the fast lane: Three 486 vendors deliver EISA bus machines, but where are the EISA add-ins? Byte Magazine, 15 (10):172–176, October 1990. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

VanName:1990:CHS

[VC90b] M. L. Van Name and B. Catchings. Color hits the streets: NEC brings color to a laptop, but is it worth it? Byte Magazine, 15(4):145–149, April 1990. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

VanName:1990:ERC

[VC90c] M. L. Van Name and B. Catchings. Easing the RAM-Cram blues: Take an active role in managing your applications and TSRs and their use of your memory. Byte Magazine, 15(3):227– 228, 230, 232, 234, March 1990. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

VanName:1990:HPA

[VC90d] M. L. Van Name and B. Catchings. High-Performance 486 ATs: The great performance of three 33-MHz 486s shows there’s still life in the old AT bus. Byte Magazine, 15(12):190–192B, November 1990. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

VanName:1990:ISM

[VC90e] M. L. Van Name and B. Catchings. Inexpensive SXes by mail: Two 80386SX systems that provide 80386 power at low prices. Byte Magazine, 15(3):143–146, 148, March 1990. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

VanName:1990:MBY

[VC90f] M. L. Van Name and B. Catchings. More bang for your buck: Four integrated software packages that won’t strain your budget. Byte Magazine, 15(3):245– 248, 250, 252, March 1990. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

VanName:1990:TPP

[VC90g] M. L. Van Name and B. Catchings. Transparent and portable: By providing a consistent framework, applications architectures let software run on different machines and operating systems. Byte Magazine, 15 (4):199–202, April 1990. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

VanName:1990:TGC

[VC90h] M. L. Van Name and B. Catchings. Two to grow on: CPUupgradable systems from AST and ALR. Byte Magazine, 15 (6):164–167, June 1990. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

VanName:1990:WFL

[VC90i] M. L. Van Name and B. Catchings. World’s fastest lunchbox: Dolch is first to sell an i486based portable. Byte Magazine, 15(5):155–158, May 1990. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

VanName:1990:ATP

[VC90j] Mark L. Van Name and Bill Catchings. Apple Talk Phase 2 and you: How will AppleTalk Phase 2 affect your LAN? the answer depends on what you’re using and what your needs are. Byte Magazine, 15(1):145– ??, January 1990. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

VanName:1990:FLY

[VC90k] Mark L. Van Name and Bill Catchings. Faraway LANs: You don’t have to be in the office to take advantage of the office LAN. Byte Magazine, 15(4):97–98, 100, April 1990. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

VanName:1990:HCN

[VC90l] Mark L. Van Name and Bill Catchings. Hard choices for network managers: You can’t always wait for the perfect network management solution. Byte Magazine, 15(12):97–98, 100, November 1990. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

VanName:1990:KSP

[VC90m] Mark L. Van Name and Bill Catchings. Kicking and screaming into the present: DEC slowly embraces PC networking standards. Byte Magazine, 15(13): 125–126, 128, December 1990. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

VanName:1990:NMS

[VC90n] Mark L. Van Name and Bill Catchings. A natural match: Sharing CD-ROMs over a LAN seems like a natural idea. so why isn’t it easier? Byte Magazine, 15 (6):109–110, 112, June 1990. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

VanName:1990:NP

[VC90o] Mark L. Van Name and Bill Catchings. Network perestroika. Byte Magazine, 15(11):57–??, Fall 1990. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

VanName:1990:NSH

[VC90p] Mark L. Van Name and Bill Catchings. Networks shouldn’t be this hard: Getting PCs up and running on a LAN is too much of a hassle. Byte Magazine, 15 (7):105–106, 108, July 1990. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

VanName:1990:NWM

[VC90q] Mark L. Van Name and Bill Catchings. New ware’s missing links: NetWare products connect disparate systems, but pieces don’t always fit perfectly. Byte Magazine, 15(2):137–140, February 1990. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

VanName:1990:SPM

[VC90r] Mark L. Van Name and Bill Catchings. Space patrol: Managing thousands of files on today’s big server hard disks doesn’t have to be a nightmare. Byte Magazine, 15(8):101, 104, August 1990. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

VanName:1990:UTD

[VC90s] Mark L. Van Name and Bill Catchings. Unite or die: Three developing application areas must unite before LANs can become a part of everyday life in the 1990s. Byte Magazine, 15(9): 113–114, 116, September 1990. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

VanName:1990:WLD

[VC90t] Mark L. Van Name and Bill Catchings. Welcome to Lilliput, DEC: DEC once again visits the land of microcomputers with LanWorks for Macintosh. Byte Magazine, 15(10):115–??, October 1990. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Veljkov:1990:MMA

[Vel90] M. D. Veljkov. Managing multimedia: Authoring systems let nonprogrammers create powerful multimedia applications. Byte Magazine, 15(8):227–232, August 1990. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Vislosky:1994:BCR

[VFGZ94] Dave Vislosky, Rich Friedman, Roger Goode, and Lee Zaslow.

Books and CD-ROMs: Internet resource guide: Thousands of Internet listings, government information on CD-ROM, computer design, and repetitive strain injuries. Byte Magazine, 19(4):41– ??, April 1994. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Vaughan-Nichols:1990:BDF

[VN90a] S. J. Vaughan-Nichols. A better dBASE: FoxPro may have outdone all other dBASE systems, including dBASE IV. Byte Magazine, 15(4):163–164, 166, 168, April 1990. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Vaughan-Nichols:1990:GYBb

[VN90b] S. J. Vaughan-Nichols. Getting your byte’s worth: Hardwarebased data compression gives you more bang for your QIC, DAT, and hard disk buck. Byte Magazine, 15(12):331–332, 334–336, November 1990. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Vaughan-Nichols:1990:RDR

[VN90c] S. J. Vaughan-Nichols. Relational databases: The real story: Is that a relational database manager or not? Byte Magazine, 15(13):321–322, 324–325, December 1990. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Vaughan-Nichols:1990:SSW

[VN90d] S. J. Vaughan-Nichols. Saving space: Whatever size hard disk you have, it’s probably nearly full. data compression can help. Byte Magazine, 15(3):237–238, 240, 242–243, March 1990. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Vaughan-Nichols:1990:XSM

[VN90e] S. J. Vaughan-Nichols. X.400: Standardizing E-Mail: E-mail is ready to live up to its promise. Byte Magazine, 15(13):341–342, 344, December 1990. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Vaughan-Nichols:1990:GYBa

[VN90f] Steven J. Vaughan-Nichols. Getting your Byte’s worth. Byte Magazine, 15(??):331–336, November 1990. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Vaughan-Nichols:1991:CCC

[VN91a] S. J. Vaughan-Nichols. Catch as cache can: Cache coherency is a critical element of sharedmemory multiprocessor systems. Byte Magazine, 16(6):209–210, 212–214, 216, June 1991. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Vaughan-Nichols:1991:DIF

[VN91b] S. J. Vaughan-Nichols. Disk insurance: Fault tolerance isn’t an option; it’s a necessity. Byte Magazine, 16(8):195–202, August 1991. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Vaughan-Nichols:1991:NCN

[VN91c] S. J. Vaughan-Nichols. No-Compromise notebooks with 386SX power: From a torrent of SX notebook introductions, one clear winner emerges. Byte Magazine, 16(6):282–284, 286, 288, 290, 292, June 1991. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Vaughan-Nichols:1991:SUB

[VN91d] S. J. Vaughan-Nichols. SX upgrade boards: Not for the fainthearted: Boards that upgrade your 286 system can be difficult to install. Byte Magazine, 16(4): 283–284, 286, April 1991. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Vaughan-Nichols:1991:WLM

[VN91e] S. J. Vaughan-Nichols. When less is more: Making Mac images more manageable: Storm Technology debuts an effective imagecompression duo. Byte Magazine, 16(5):263–264, May 1991. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Vaughan-Nichols:1991:TDE

[VN91f] Steve Vaughan-Nichols. Transparent data exchange: Steve Vaughan-Nichols surveys the current state of networkwide data transparency. Byte Magazine, 16(12):211–??, November 1991. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Vogelgesang:1990:DDG

[Vog90] Peter Vogelgesang. Drowning in data: The gathering deluge of information calls for new approaches to data storage. Byte Magazine, 15(2):251–256, February 1990. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Voorhies:1993:DT

[Voo93] Douglas Voorhies. Damn the torpedoes! Byte Magazine, 18(12): 224–??, November 1993. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Vornberger:1992:NIW

[Vor92] Cal Vornberger. A new Illustrator for Windows: A few problems temper the enthusiasm for Adobe’s latest release of Illustrator for Windows. Byte Magazine, 17(9):263–??, September 1992. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Vornberger:1993:BBM

[Vor93] Cal Vornberger. Beyond bit maps: Multiple floating objects deliver new power and flexibility to bit-map image processing. Micrografx Picture Publisher 4.0 and Fractal Design Painter X2 feature object layers and other useful innovations. Byte Magazine, 18(13):165–166, 168, December 1993. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Vose:1990:NTY

[Vos90a] G. Michael Vose. New tricks for your laser printer: Dan Bricklin’s PageGarden takes laser printing beyond most application programs. Byte Magazine, 15(1): 225–227, January 1990. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Vose:1990:HLG

[Vos90b] M. Vose. Hot links to go: A look at Windows’ and OS/2’s dynamic data exchange facility. Byte Magazine, 15(12):373–377, November 1990. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Voswinkel:1992:LIA

[Vos92] Merik Voswinkel. Locating Internet access in Europe. Byte Magazine, 17(10):149–??, October 1992. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Walsh:1993:CLG

[Wal93] Birrell Walsh. Correspondence that looks good globally: Multi-Lingual scholar, a word processor for the global market. Byte Magazine, 18(2):219–??, February 1993. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Wallace:1994:AED

[Wal94a] Scott Wallace. Accelerating engineering design: A sophisticated engineering design system — combined with concurrent engineering methodology — is compressing time on a Lockheed Missile and Space Corp. missile development program. Byte Magazine, 19(7):62–??, July 1994. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Wallace:1994:CPR

[Wal94b] Scott Wallace. The computerized patient record: Computerized patient records will improve health care and reduce costs. Byte Magazine, 19(5):67–68, 70, 72, 74, 76, May 1994. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Wallace:1994:MMS

[Wal94c] Scott Wallace. Managing mass storage: Advances in storage technology struggle to keep up with the burgeoning pace of storage requirements. Byte Magazine, 19(3):78–??, March 1994. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Wallace:1994:SFE

[Wal94d] Scott Wallace. Solutions focus: Experts in the field: Expertsystem technology can pay off by capturing and codifying knowledge from throughout an organization. Byte Magazine, 19(10): 86–88, 90, 94–96, October 1994. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Wallace:1994:WSP

[Wal94e] Scott Wallace. Working smarter:

Process-analysis and modeling tools provide core support for workgroup and work-flow applications. experience in the trenches shows that technology is only part of the solution. Byte Magazine, 19(7):100–101, July 1994. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Warnock:1992:NAD

[War92] J. E. Warnock. The new age of documents: Establish a viable document-interchange strategy. Byte Magazine, 17(6):257– 260, June 1992. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Watterson:1994:CWE

[Wat94] Karen Watterson. The changing world of EIS: The role of EIS in a distributed environment is uncertain. Byte Magazine, 19(6):183– ??, June 1994. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Waurzyniak:1992:FIA

[Wau92] Patrick Waurzyniak. First impressions AutoCAD gets a facelift: And AST’s Power Premium 4/50d. Byte Magazine, 17(9): 46–??, September 1992. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Waurzyniak:1993:FIA

[Wau93a] Patrick Waurzyniak. First impressions: AutoCAD draws on Windows: Autodesk’s bestselling CAD software goes Windows. Byte Magazine, 18(6):33– ??, May 1993. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Waurzyniak:1993:HSS

[Wau93b] Patrick Waurzyniak. HP’s superior subnotebook: Hewlett– Packard packs a lot, including Windows and applications in ROM, into its 3-pound omnibook. Byte Magazine, 18(8):32– ??, July 1993. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Waurzyniak:1993:WGG

[Wau93c] Patrick Waurzyniak. WordPerfect goes GUI with DOS update: It’s taken WordPerfect more than three years to update its DOS word processor, but version 6.0 may be worth the wait. Byte Magazine, 18(7):42– ??, June 1993. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Wayner:1990:TMP

[Way90a] P. Wayner. Time and money: A program called Spawn uses auctions to fairly allocate precious computer time. Byte Magazine, 15(4):252–254, 256, 258, April 1990. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Wayner:1990:ZSM

[Way90b] P. Wayner. The zen of symbolic math: Theorist takes the low road to equation solving by leading you to the right answer on your Mac. Byte Magazine, 15 (6):193–194, 196, June 1990. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Wayner:1990:MMP

[Way90c] Peter Wayner. Mainframe math on a PC: Macsyma, the grande dame of computer algebra, is finally available for PCs. Byte Magazine, 15(1):207–210, January 1990. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Wayner:1991:BIF

[Way91a] P. Wayner. The big index (file indexing). Byte Magazine, 16 (6):317–318, 320, 322, 412–413, June 1991. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Wayner:1991:GAN

[Way91b] Peter Wayner. Genetic algorithms: A novel technique crossbreeds algorithms to find the best programming solution. Byte Magazine, 16(1):361–364, 366, 368, January 1991. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Wayner:1991:SMH

[Way91c] Peter Wayner. Smart memories: How content-addressable memory chips and other special memories can speed searches and solve surprising problems. Byte Magazine, 16(3):147–150, 152, March 1991. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Wayner:1991:TDL

[Way91d] Peter Wayner. True data: A look at techniques for ensuring the authenticity of the data you send, receive, or store. Byte Magazine, 16(9):122–124, 126, 128, September 1991. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Wayner:1992:BND

[Way92a] Peter Wayner. Brave new desktop: Object technologies will let you shape your computing environment. Byte Magazine, 17(14): 153–??, December 1992. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Wayner:1992:ISC

[Way92b] Peter Wayner. Inside signal computing: The architecture of DSP chips mirrors the functions they perform. Byte Magazine, 17(12): 177–??, November 1992. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Wayner:1993:OCR

[Way93a] P. Wayner. Optimal character recognition. Byte Magazine, 18 (??):??, ?? 1993. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Wayner:1993:SE

[Way93b] P. Wayner. Stretching the ether. Byte Magazine, 18(??):159–165, ?? 1993. CODEN BYTEDJ.

ISSN 0360-5280 (print), 1082-7838 (electronic).

Wayner:1993:ECD

[Way93c] Peter Wayner. Encryption chip draws fire: A new encryption chip promises to protect your electronic messages, but there’s a catch: A trapdoor lets the government eavesdrop. Byte Magazine, 18(8):36–??, July 1993. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Wayner:1993:IBL

[Way93d] Peter Wayner. Image building: A look at the core of modern visualization software. Byte Magazine, 18(4):137–??, April 1993. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Wayner:1993:IWA

[Way93e] Peter Wayner. Inside Windows accelerators: Speeding up the operations of your graphical interface is the job of some very specialized technology. Byte Magazine, 18(12):229–??, November 1993. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Wayner:1993:PPF

[Way93f] Peter Wayner. Print pages faster:

In the near future, you will be asking your printer to produce pages ever more quickly. help is on the way in the form of better processors and more efficient software. your computer’s CPU might throw its weight behind processing print images, too. Byte Magazine, 18(13): 115–??, December 1993. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Wayner:1993:SER

[Way93g] Peter Wayner. Should encryption be regulated: U.S. law enforcers want to limit your use of data encryption. Byte Magazine, 18(6):129–??, May 1993. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Wayner:1993:SET

[Way93h] Peter Wayner. Stretching the ether: Technology expands the wireless spectrum. Byte Magazine, 18(2):159–??, February 1993. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Wayner:1993:UHO

[Way93i] Peter Wayner. Under the hood: Optimal character recognition: Reliable optical character recognition faces many hurdles, including odd typefaces and marginal output. expert systems, machine learning, and other techniques can help you tackle the problem. Byte Magazine, 18(13):203–??, December 1993. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Wayner:1994:ATP

[Way94a] Peter Wayner. Agents away: Telescript provides the glue that lets personal communicators access the world. it could change the way you talk to the world. Byte Magazine, 19(5):113–??, May 1994. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Wayner:1994:ATI

[Way94b] Peter Wayner. Automating TCP/IP in NT. Byte Magazine, 19(11):189–??, November 1994. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Wayner:1994:DHD

[Way94c] Peter Wayner. Digital hard drives: Digital-read-channel technology will double the capacity of a standard hard drive, letting you pack more data on your hard disk. Byte Magazine, 19(3):91–??, March 1994. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Wayner:1994:DVG

[Way94d] Peter Wayner. Digital video goes Real-Time: Real-time video might soon be on your desktop thanks to C-Cube’s VideoRISC Compression Architecture, which can encode video on the fly using either MPEG 1 or MPEG 2. Byte Magazine, 19(1):107–??, January 1994. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Wayner:1994:EMD

[Way94e] Peter Wayner. EDI moves the data: Eliminating the paper chase can speed up business and cut costs. Byte Magazine, 19 (10):121–122, 124, 126, 128, October 1994. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Wayner:1994:OMT

[Way94f] Peter Wayner. Objects on the march: The trend is toward an object-oriented approach to the design of operating systems. Byte Magazine, 19(1):139– ??, January 1994. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Wayner:1994:RAA

[Way94g] Peter Wayner. On the road to ATM: ATM is a strong contender for connecting enterprise networks cost-effectively. among the first to adopt ATM will be small workgroups that need to move large blocks of data. Byte Magazine, 19(9):103–??, September 1994. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Wayner:1994:SDI

[Way94h] Peter Wayner. Silicon for 3-D: Inexpensive 3-D hardware is coming to a PC near you, and games, slide presentations, virtual reality — even spreadsheets — will never be the same. Byte Magazine, 19(9):191–??, September 1994. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Wayner:1994:SRR

[Way94i] Peter Wayner. Silicon in reverse: Reversible logic circuits promise to radically decrease the power requirements of future VLSI chips. Byte Magazine, 19 (8):67–??, August 1994. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Wayner:1994:SSB

[Way94j] Peter Wayner. SPARC strikes back: The latest incarnation of SPARC can perform with the big boys and is a multimedia screamer to boot. Byte Magazine, 19(11):105–??, November 1994. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Wayner:1994:VQD

[Way94k] Peter Wayner. VLIW questions:

Does the Intel/HP alliance spell the beginning of the end for RISC? Byte Magazine, 19(11): 287–??, November 1994. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Wood:1990:SBN

[WB90] Lamont Wood and Dana Blankenhorn. State of the BBS nation: Whatever your electronic appetites, you can feed them on a BBS. Byte Magazine, 15(1): 298–304, January 1990. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Wszola:1992:PMC

[WE92] S. Wszola and D. L. Edwards.

PostScript’s middle class (Post-Script printers). Byte Magazine, 17(8):224–230, 232, 234, 236–237, August 1992. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Weber:1992:OPR

[Web92] Jack Weber. Overview: Photonics: Revolution or evolution: Optical technologies are changing the way we process information. Byte Magazine, 17(9): 168–??, September 1992. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Weber:1993:OVS

[Web93] Jack Weber. Overview: Visualization: Seeing is believing: Visualization lets you see the meaning of numeric data. Byte Magazine, 18(4):120–??, April 1993. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Webster:1994:WNC

[Web94] Bruce F. Webster. Whither NextStep: Concepts of the NextStep operating system are finding new life in OpenStep. Byte Magazine, 19(11):289–??, November 1994. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Wszola:1990:JWH

[WET90] Stan Wszola, Howard Eglowstein, and Tom Thompson. Just what the hard disk doctor ordered (hard disk utilities). Byte Magazine, 15(1):152–164, January 1990. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Wiggins:1990:HCS

[Wig90] M. Wiggins. Help for the C sick:

Microsoft sets its sights on professional developers with BASIC PDS 7.0. Byte Magazine, 15(5): 193–194, 196, May 1990. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Wiggins:1992:WBT

[Wig92] Mike Wiggins. Windows BA-SICs: Three Windows BASIC programs square off. Byte Magazine, 17(14):196–??, December 1992. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Williamson:1994:HTT

[Wil94] M. Williamson. High-tech training (Anderson Consulting). Byte Magazine, 19(12):74–75, 78, 80, 84, 88, December 1994. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Whang:1990:JEE

[WK90] Min-Hur Whang and Joe Kua.

Join the EISA evolution: Bus mastering lets desktop CPUs spend more of their time processing data rather than dealing with I/O. Byte Magazine, 15(5):241– 244, 246–247, May 1990. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Wood:1990:DOH

[Woo90a] L. Wood. DR DOS offers hope for the RAM-Crammed: Digital Research’s new MS-DOS competitor promises to make more memory available for applications. Byte Magazine, 15(13): 197–198, 200, December 1990. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Wood:1990:JAT

[Woo90b] L. Wood. Jack of all trades: IBM’s current is a personal information manager with deskaccessory-style functionality. Byte Magazine, 15(3):179–180, 182, March 1990. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Wood:1990:WST

[Woo90c] L. Wood. Windows 3.0 software tool for end users: Asymetrix’s TookBook lets you create Windows 3.0 applications without learning C. Byte Magazine, 15(9):159–160, September 1990. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Wood:1990:WPW

[Woo90d] L. Wood. Word processing in Windows: Ami Professional, Legend, and Word for Windows are the first WYSIWYG word processors for Microsoft Windows, but are they fast enough? Byte Magazine, 15(4):157–160, April 1990. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Wood:1991:DPN

[Woo91a] L. Wood. Desktop prototyping: New technologies make CAD drawings into solid models — from PCs! Byte Magazine, 16(5):137–138, 140, 142, May 1991. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Wood:1991:SLB

[Woo91b] L. Wood. Script languages: The BASIC of the 1990s: An examination of quick-and-dirty programming products on Macs and PCs. Byte Magazine, 16(4):244– 246, 248, 250, April 1991. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Wright:1991:DCN

[WS91] D. P. Wright and C. L. Scofield.

Divide and conquer: Neural networks offer the promise of quick and effective automatic character recognition. Byte Magazine, 16(4):207–210, April 1991. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Ward:1993:UHD

[WS93] Jean Renard Ward and Debra Schultz. Under the hood: Digitizer renaissance: Pen computers are reinvigorating digitizer tablet technologies. Byte Magazine, 18 (1):251–??, January 1993. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Wszola:1990:FYF

[Wsz90]  S. J. Wszola.  Flashdisk:

Not your father’s RAM disk: Digipro’s Flashdisk plugs up to 8 megabytes of nonvolatile memory into any available 16-bit slot.

Byte Magazine, 15(9):214–215, September 1990. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Wszola:1991:SBU

[Wsz91] S. Wszola. Server-based UP-Ses promise order during power problems. Byte Magazine, 16 (7):235–236, July 1991. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Wszola:1993:FPO

[Wsz93a] Stan Wszola. Fax plus OCR: More than meets the eye: The BYTE Lab reviews eight fax packages with OCR built in. OCR on faxed documents is not quite automatic, but these systems can speed the transition from graphical image to usable text data. Byte Magazine, 18 (9):130–132, 134, 136, 138, 140, August 1993. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Wszola:1993:ISC

[Wsz93b] Stan Wszola. Info Select clears clutter: this PIM for Windows works the way you do. Byte Magazine, 18(5):97–98, Spring 1993. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Wszola:1993:WDW

[Wsz93c] Stan Wszola. Windows dressing: The Windows user interface isn’t for everyone. the BYTE Lab examines 12 accessory packages that plug the gaps in the Windows user interface, including New Wave, Power Launcher, and Norton Desktop for Windows. Byte Magazine, 18(7):148– ??, June 1993. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Yager:1990:AOSa

[Yag90a] T. Yager. Alternative operating systems, part 1: The QNX operating system: The first installment of a six-part series. Byte Magazine, 15(8):281–283, August 1990. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Yager:1990:AOSc

[Yag90b] T. Yager. Alternative operating systems, part 3: Theos: Serious business: A unique multiuser operating system for business applications. Byte Magazine, 15(10): 295–296, 298, October 1990. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Yager:1990:AOSe

[Yag90c] T. Yager. Alternative operating systems, part 5: Unix with a microscope: Minix, a lowcost Unix, runs on ordinary personal computers. Byte Magazine, 15(13):345–346, December 1990. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Yager:1990:CRB

[Yag90d] T. Yager. Compaq’s reason to believe in EISA: Compaq’s newest high end system, the Systempro, may be its best yet. Byte Magazine, 15(3):122–124, March 1990. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Yager:1990:DLR

[Yag90e] T. Yager. DEC’s latest RISC: Digital Equipment makes a play for the serious work station user with its revved-up DECstation. Byte Magazine, 15(9):154–156, September 1990. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Yager:1990:UST

[Yag90f] T. Yager. DOS and Unix: On speaking terms: A productive link between DOS and Unix systems begins with an understanding of Unix network basics. Byte Magazine, 15(6):281–282, 286, 288, 290, 292, June 1990. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Yager:1990:PGM

[Yag90g] T. Yager. DOS on a pedestal: DOS goes multiuser with Theo + DOS. Byte Magazine, 15(8): 194, 196, August 1990. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Yager:1990:MII

[Yag90h] T. Yager. Mac-ish interfaces for Unix: Looking Glass and X desktop provide point-and-click ease of use to Unix. Byte Magazine, 15(12):227–228, 230, November 1990. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Yager:1990:BCW

[Yag90i] T. Yager. On becoming a clock wise scheduler: Phase II Software’s ClockWise helps manage your time. Byte Magazine, 15(13):201–202, December 1990. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Yager:1990:ODR

[Yag90j] T. Yager. Open Desktop: Relief for the Unix-Wary: SCO’s Open Desktop may be the shrinkwrapped Unix that DOS users have been waiting for. Byte Magazine, 15(9):176–178, 180, September 1990. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Yager:1990:US

[Yag90k] T. Yager. OS/2, Unix style.

Byte Magazine, 15(2):8015– 8013–8016, February 1990. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Yager:1990:PSU

[Yag90l] T. Yager. Power and the single user: Opus weds big-iron performance to PC prices. Byte Magazine, 15(5):169–172, May 1990. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Yager:1990:SNM

[Yag90m] T. Yager. Sony NeWS and MIPS magnum: A double shot of RISC: Two RISC workstations join the low-end Unix market. Byte Magazine, 15(13):172–175, December 1990. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Yager:1991:BMP

[Yag91a] T. Yager. Build multimedia presentations with MacroMind’s MediaMaker: MediaMaker sets the pace with captivating presentations. Byte Magazine, 16(9): 302–304, September 1991. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Yager:1991:CGV

[Yag91b] T. Yager. Computers go video with NEC’s PC-VCR: There’s more to this high-tech VCR than meets the eye. Byte Magazine, 16(9):307–308, September 1991. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Yager:1991:FNS

[Yag91c] T. Yager. Five new SPARCbased workstations compete with Sun: Five non-Sun systems put SPARC’s promise to the test. Byte Magazine, 16(7):210–214, July 1991. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Yager:1991:IHD

[Yag91d] T. Yager. Information’s human dimension: Multimedia technologies provide the tools that can help you make your presentations come alive. Byte Magazine, 16(13):153–156, 158, 160, December 1991. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Yager:1991:LSS

[Yag91e] T. Yager. The littlest SPARC:

Sun’s SPARCstation IPC squeezes RISC power onto even the smallest desk. Byte Magazine, 16 (2):169–170, 172, 174, February 1991. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Yager:1991:NVT

[Yag91f]  T. Yager.  Newtek’s Video

Toaster makes professional video affordable: “desktop video” now a reality. Byte Magazine, 16 (3):245–246, 248–250, 252, 254, March 1991. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Yager:1991:WMC

[Yag91g] T. Yager. A workstation in a Mac’s clothing: A/UX and the X Window System turn a Macintosh into a workstation in a near-seamless way. Byte Magazine, 16(1):213–214, 26, January 1991. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Yager:1991:XTW

[Yag91h] T. Yager. X terminals for workstation power at PC prices: Seamless Unix and X window networking. Byte Magazine, 16 (5):238–242, 244, May 1991. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Yager:1992:MPH

[Yag92a] T. Yager. The multimedia PC: High-Powered sight and sound on your desk: A look at the first crop of MPC hardware and software. Byte Magazine, 17(2): 217–220, 222, 224, 226, February 1992. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Yager:1992:PDV

[Yag92b] T. Yager. Practical desktop video, part 1: The BYTE Multimedia Lab’s series on desktop video production begins with a look at video formats. Byte Magazine, 17(4):106–110, 112, 114, April 1992. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Yager:1992:PGD

[Yag92c] T. Yager. Presentation graphics that deserve an extra bow: Curtain call makes high-impact presentations easy and affordable. Byte Magazine, 17(5):275– 276, May 1992. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Yager:1992:AER

[Yag92d] Tom Yager. Animation energizes Rio’s world-class graphics: Rio Animator brings a slick look to presentation graphics. Byte Magazine, 17(9):279–??, September 1992. CODEN BYTEDJ.

ISSN 0360-5280 (print), 1082-7838 (electronic).

Yager:1992:CVF

[Yag92e] Tom Yager. Cut to video: Four programs for moving presentations: The BYTE Lab looks at four video-presentation programs. Byte Magazine, 17(12): 238–240, 242, 244, 246, November 1992. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Yager:1992:FPP

[Yag92f]  Tom Yager.  FrameMaker:

Power publishing for Windows: FrameMaker brings its workstation-bred document publishing power to Windows. Byte Magazine, 17(10):247, 249–250, October 1992. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Yager:1992:INS

[Yag92g] Tom Yager. IBM’s new system speaks for itself: The PS/2 Ultimedia. Byte Magazine, 17(9): 269–270, September 1992. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Yager:1992:NSN

[Yag92h] Tom Yager. NextStep: The sleeper: NextStep quietly fills the need. Byte Magazine, 17 (10):190–??, October 1992. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Yager:1992:SMS

[Yag92i] Tom Yager. Style meets substance in Matrox Studio: Studio turns your PC into a video powerhouse. Byte Magazine, 17(12): 261–??, November 1992. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Yager:1992:WSQ

[Yag92j] Tom Yager. Windows speaks: A quick guide to using Windows Multimedia Extensions. Byte Magazine, 17(11):33–??, November 1992. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Yager:1993:BBV

[Yag93a] Tom Yager. A BASIC breakthrough: Visual Basic for DOS makes powerful programs easy to write. Byte Magazine, 18 (2):207–??, February 1993. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Yager:1993:CYC

[Yag93b] Tom Yager. Color you can carry anywhere: combine a 486 CPU with a color active-matrix LCD, and you get a dynamite Windows notebook. Byte Magazine, 18(5): 69–70, 72–73, Spring 1993. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Yager:1993:CGT

[Yag93c] Tom Yager. Commodore gets tough: Two new 69040-based Amigas. Byte Magazine, 18 (1):239–??, January 1993. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Yager:1993:IVW

[Yag93d] Tom Yager. Inside video for Windows: Microsoft’s new extensions for desktop video open new horizons to Windows computing. Byte Magazine, 18(5):57– ??, Spring 1993. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Yager:1993:MWP

[Yag93e] Tom Yager. Macs and Windows PCs share control: Timbuktu for Windows makes cross-platform remote control possible, but it can be slow. Byte Magazine, 18(4):185–??, April 1993. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Yager:1993:MEI

[Yag93f] Tom Yager. OS/2’s multimedia extensions: IBM builds a strong multimedia foundation for OS/2. Byte Magazine, 18(4):177– ??, April 1993. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Yager:1993:PND

[Yag93g] Tom Yager. Photoshop now does Windows: Version 2.5 is new for Windows and improved for the Mac. Byte Magazine, 18(4):49– ??, April 1993. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Yager:1993:SPP

[Yag93h] Tom Yager. The second premiere: Premiere 2.0 for the Mac offers QuickTime movie editing at its best. Byte Magazine, 18 (2):205–??, February 1993. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Yager:1993:TSD

[Yag93i] Tom Yager. Topas 4.0 simplifies 3-D: Topas 4.0 makes quick work of realistic 3-D graphics. Byte Magazine, 18(1):231– ??, January 1993. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Yager:1993:UNH

[Yag93j] Tom Yager. UnixWare: New hope for Unix: The friendliest Unix around. Byte Magazine, 18 (1):51–??, January 1993. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Yager:1993:VBW

[Yag93k] Tom Yager. Visual Basic for Windows gets a face-lift: Microsoft improves its programming package with version 2.0. Byte Magazine, 18(3):177–??, March 1993. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Yager:1993:WVR

[Yag93l] Tom Yager. Windows video recorders: The hardware side of video for Windows. Byte Magazine, 18(6):201–??, May 1993.

CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Yahdav:1992:TEP

[Yah92] Daniel Yahdav. Tracking the elusive project (project management packages). Byte Magazine, 17(11):119–122, 124, 126, November 1992. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Yavin:1994:ONR

[Yav94] David Yavin. Optimizing Notes replication: The issue of replication scheduling is moving to the forefront. buyers and vendors of distributed data systems should raise their level of awareness about this potentially critical issue. Byte Magazine, 19(9): 201–??, September 1994. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Yager:1992:SFG

[YG92] Tom Yager and Rick Grehan. Solutions focus: Grab your audience with audio: With current sound boards, sequencers, and editors, you don’t have to be a recording engineer to add audio to your presentations. Byte Magazine, 17(14):174–194, December 1992. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Yager:1993:BTB

[YM93] Tom Yager and D. H. Mitchell.

Better than being there: Desktop video conferencing is on its way. Byte Magazine, 18(3):129– 130, 132–134, March 1993. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Yourdon:1990:ALS

[You90] E. Yourdon. Auld Lang Syne: Ed Yourdon tackles the question of whether you should jump on the object-oriented bandwagon. Byte Magazine, 15(10):257, 260, 262, 264, October 1990. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Young:1993:SAR

[You93] Michael J. Young. Some assembly required animation for Windows applications: CSprite class forms the foundation for adding animation to Windows programs. Byte Magazine, 18 (7):197–??, June 1993. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Yourdon:1994:DSO

[You94] Edward Yourdon. Developing software overseas: The developing countries are also developing competitive software industries. Byte Magazine, 19(6):113– ??, June 1994. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Yager:1990:SSS

[YS90] T. Yager and B. Smith. Son of SPARCstation: Sun Microsystems ups the midrange workstation ante. Byte Magazine, 15 (13):140–141, 143, 146, December 1990. CODEN BYTEDJ.

ISSN 0360-5280 (print), 1082-7838 (electronic).

Yager:1992:UDU

[YS92] Tom Yager and Ben Smith. Is Unix dead: Unix: A child of a thousand parents: The history of Unix says much about its future. Byte Magazine, 17 (9):134–??, September 1992. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Yager:1990:NUS

[YT90] T. Yager and T. Thompson. The Norton Utilities for System V. Byte Magazine, 15(13):178–??, December 1990. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Zeis:1993:CBA

[Zei93] Michael Zeis. Color becomes affordable: At both the high and low ends, color printing continues to improve dramatically. whether you want photographicquality output or to spice up a document, you’ll find a color technology that fits the bill. Byte Magazine, 18(13):125–??, December 1993. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Zanellato:1992:DS

[ZV92] Georges Zanellato and Bart Verhaeghe. Digitally speaking. Byte Magazine, 17(12):??, November 1992. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).